A Complete Bibliography of Publications in Communications in Statistics: Theory and

Methods: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.10

Title word cross-reference

(−∞, ∞) [CC16a]. (1, 2] [Uya17]. (2n − 1) [Len17]. (GARMA(1, 1; δ_{1}, δ_{2})) [PSP12]. (k, l) [RS19, TSG19]. (μ_{x}) [SS15a]. (n, f, k) [Kam14, KS14a]. (n − k + 1) [RSA13, SHB17]. (n − r + 1) [BBHN18]. (n > p) [SC18]. (NML_{m}(m)) [Gar10]. (r − k) [AK13]. (s, S) [Ali17]. 1 [BB12, CO10]. 1/α [Sel17b, Sel17a]. 100 [LZH15]. 1: m [He18]. 2 [PBY^{+}14]. 2^{k} [Nor16]. 2^{m} [HYK15b]. 2^{m−k} [Jac11]. 2^{m−p} [YWZ18]. 2^{n−(n−k)} [Hil15]. 2^{n}4^{1} [ZYLZ17]. 2l + 1 [HYK15b]. 2 × 2 [Bar17b, BSY12]. 3 [Are14, KS12a, TIA11]. 3^{m} [KHY17, YH11]. > [IBA18]. ^{k} [TN13]. ^{Q} [Gue10]. ^{X} [KC19a, Ye19, KC19b]. k [ZXM16]. A [Gra12, HS19, ROM16, ZLZ19]. α [AD17a, AD17b, KAS17, KAPA19, Lia10, Lia11, LLY14, LO17, QL19, RASR17, She18a, SW19a, SK10b, WLF11, Yeh17, YSC17b, ZC17b, dSE11]. α = 2 [TIA11]. AR(1) [MZ15, PZ15, WG15, ZLF14, Pop12]. ARMA(p, q) [RB11]. B [FZ10b]. X̄ [PKM17, PMKF17, PMFK18, SMSF19, SKK18, Sag15, WWZC18, YKL^{+}17]. β [RM10, SK10b, YM18]. c [FN09, KNHM17, WCK16]. C_{pk} [AB13]. C_{P L} [CC15a]. C_{pm} [AK19]. C_{P U} [CC15a]. χ^{2} [BK09, Huh16]. D [AW10, CG19, DDM10, DDM14, Jaf13, KS12a, LRC17, Sco09, Tin10, YA10, ZY14, DK17a, Sco11, Šir15]. D_{A} [Tan10a]. δ [BMLY19, PMW17, TE18a]. E [CG14a, Han17, JJ16, JKD17, Ros19, You17]. ϵ [GVGP11]. F [BDM15, Hot16]. G [AE23b, DSC^{+}14, GDSC15a, Hu17, JCY10, JC12, MAYH17, NCO15, NAYC17, RS13, ZZ15, CPW19, HZW17]. \frac{G_{1}}{G_{2}} [KC19a, KC19b]. Gamma(λ, 2) [CY10]. GAR(1) [SP13]. GARCH(1, 1) [BL13, KSBN17, PC16a]. GARCH(2, 2) [KTBN17]. GN_{2} [Oll17]. GQARCH(1, 1) [AL15]. h [SSS18]. H > 1/2 [MS11b]. H ∈ (2/3, 1) [FNM19]. I [FK19, Min11, RAB15]. I(d) [Nar13]. INAR(1) [BIP^{+}10, KP13, Nas12, SW14b]. INAR(p) [PS13, TW14]. ∞ [Bur19]. K [BSS19, LCA17, RZ14, VBATK11, AC16, CKLC17, DK17a, DK17b, DSC^{+}14, Gau17, GXB18, Jas16, Jas19, JCY10, JC12, KQ17, Kom17, KN16a, LB16, MT14, MP15, MGP14, Min10, MA14, SBA17, SSB17b, ŠS12c, Šir14, SNN16, Wan16e, Wu16c, YL18d, Zha11, ZT17, ZL15]. L [Dut17, Hui10, KR17, Ryc14b, SV11, MP15]. L^{1} [PTV18]. L^{2} [Qum12, ZLL11]. L^{p} [BD16a, KL17b, XDWC18]. L_{1} [BK17b, Has19]. L_{2} [GQC17, OQC14, YYS19]. L_{p} [Che19b, FZ17a, Li14a, MYM11]. L_{r} [SS17b, XTL^{+}17]. λ [MKD11]. ⟨n, f, k⟩ [KS14a]. ⟨n, k, f⟩ [Kam14]. M [BS15, Cah12, Dan18, ES14, Han12a, HS11, Kon15a, Kos14, Lin19, Shi10, SL15, Sun17b, Sun17a, SV11, WL10, WT15, WH17b, WJ11, YC12, ZJC10, ZWL16, ZDS16, CDS19, CKLC17, GK13, HP10, MK13, MC11, MX19, NHV17, SYW16, SY16, SY17c, SNN16, WYY16]. MA(1) [DK13, Dos11, KNH17, MS10c]. E[∏_{i=0}^{k} Tr{W^{m_{i}}}] [PVS17]. L_{2}(R^{d}) [BR17]. N [KC19b, Rez17, SY17a, BBHN18, DSC^{+}14, GXB18, JCY10, JC12, Kom17, Len17, MK13, RSA13, SHB17, Wan16e, Wys19, YL18d, Zha11, ZL15]. N ≡ 0 (mod 4) [KS12a]. n ≡ 1 (mod 4) [JJ16]. np [WCK16]. O [LG10]. P [CM16c, CFA17, BC13, BL15, Bos18, CS12, Fel10, FBS10, GF17, HXL13, Lin14b, LN13a, MCX10, Pao10, RP15, ZXM16, ZXW16]. P (X > Y) [PG18]. P (X_{(t)} > Y_{(t)}) [AHEA18]. P (Y < X) [MADSAS10, Noo16, PKJ14]. p = 3, 4, 5 [LN17]. PAR(1) [Mer16]. p ≥ 2 [BN18a]. ϕ [KAW^{+}18, LSZ16, LQ17b, LSL18]. PINAR(1)_{T} [JWZ14]. PSETAR(2, 1, 1) [Mer16]. ψ [HS14, WdX17, XY19]. q [Afe13, HG19, KV17]. QQ [HT16b]. R [CR13, LSSS14, Ros19, SA16, DK17a, ŠS12c, Šir14, Šir15, Tor12, Wu16c]. R/S [Fu11a]. R = P [X < Y] [BK17a]. R^{[X]} [PBC17]. RCINAR(1) [ZWZ12]. ρ [hCXxShY18]. ρ^{∗} [HWW15]. S [Tsi15, LCT12, LK17d]. S^{2} [COK17]. S_{B} [Fly10]. ∼ W_{√}(I, \) [PVS17]. SINAR(1, 1) [GSB12]. T

[HHC16, CY14, Cha18, CW19a, CHP14, Cos18, DYXW19, HUK19, Hod14, HNV13, HLP17, HST19, JZTN17, Kim11, PK11, RLL12, SF12b, SD14, TP14, WHZ17, Wil09, Xu10, Xu12, Xu17a, XO19, ZX16]. T^{2} [BG19a, CSM16, DW18, FCSH16, KSF17, Lee13, LK19a, MCY14, PM18, SFSH16, SI17, TN19]. τ [SSS14a]. θ^{r} [Mah12b]. T L [Dut17]. U [BQC17, EE14, FZ10b, Fu11b, Fu12a, Fu12b, LG10, Wan17e, YYS19, YMF15]. U(θ, mθ) [PR11]. υ [RP10]. V [EE14, IBA18]. φ [Guo14, SWL14, SWLW14, WWW17, XYL14, YWH14, ZDW17]. W [EW16, Fly10, Vin13, Tee10, TS19]. X [LJ10, NGT13, TCA^{+}16, CY10, SS15a, YWLY11]. X^{−} [ML13]. X_{(t)} [AHEA18]. Y_{(t)} [AHEA18]. z [YSD^{+}16]. Z_{+} [SSA10]. |S| [LK17c].

-Almost [BR17]. -analog [KV17]. -asymptotically [NHV17]. -Asymptotics [YA10]. -Bar [CY10, YWLY11, LJ10, NGT13]. -Bayesian [SK10b, Han12a, Han17, You17]. -bifurcating [Dan18]. -branch [Dan18]. -Brownian [Hu17, ZZ15]. -characteristic [Sel17b]. -chart [KNHM17, Sag15]. -Circulant [MGP14]. -Contaminated [GVGP11]. -control [PKM17, PMKF17, PM18, PMFK18]. -Copula [PK11]. -copulas [Wys19]. -D [PBY^{+}14]. -Dependence [dSE11]. -Dependent [HP10, MC11, MX19, WYY16]. -Dimensional [MK13]. -discrepancy [GQC17, OQC14, YYS19]. -distance [BD16a]. -distorted [SSB17b]. -distribution [Cos18, Huh16]. -Distributions [Wil09, Tsi15]. -Divergence [BK09, SK10b, She18a]. -Efficient [RS13]. -error [PTV18]. -Esscher [HG19]. -estimates [WH17b]. -Estimation [ZDS16, Kon15a, SL15, Sun17b, Sun17a, WT15]. -Estimator [HS11, WL10]. -Estimators [ES14, SV11, YC12, Lin19]. -expectation [CPW19, HZW17]. -Functions [Tee10, TS19]. -Generalized [GK13]. -group [LCA17]. -hurdle [RS19]. -inflated [TSG19]. -Laplace [Yeh17]. -Level [Tor12]. -Max [RP15]. -Maximin [Tan10a]. -median [ZLL11]. -medoids [BSS19]. -Mixing [Guo14, HWW15, Lia10, Lia11, LLY14, RM10, SWL14, SWLW14, WLF11, AD17a, AD17b, hCXxShY18, KAW^{+}18, LSZ16, LQ17b, LSL18, LO17, WWW17, WdX17, XYL14, XY19, YWH14, ZDW17]. -moments [Dut17]. -Nearest [VBATK11]. -Norms [Li14a]. -occasion [SSS18]. -of- [CKLC17, Len17]. -optimal [FN09, AW10, CG14a, DDM10, DDM14, Gra12, Jaf13, KS12a, Tin10, ZY14, CG19, LRC17, RZ14, ROM16, ZLZ19]. -optimality [HS19, JJ16, JKD17, Ros19]. -ordered [Rez17]. -Out-of- [JCY10, RSA13, BBHN18, DSC^{+}14, GXB18, JC12, Kom17, SHB17, SNN16, Wan16e, YL18d, Zha11, ZL15]. -overlapping [MP15]. -plots [HT16b]. -policy [KC19b]. -Pólya [Min11]. -Principal [MK13]. -probabilities [Sel17a]. -Quantile [MCX10, BS15]. -record [AC16, SBA17]. -records [KQ17, MT14, MA14]. -rooted [SY17c]. -sample [Gau17]. -series [KAS17, KAPA19]. -shock [PMW17, TE18a]. -Skeleton [YM18]. -Spherical [RP10]. -Stable [dSE11, SW19a, YSC17b]. -State [SY17a]. -Statistic [Xu10, Wan17e]. -Statistics [FZ10b, Fu11b, Fu12a, Fu12b, Hui10, Ryc14b, Shi10, EE14, KR17, YMF15]. -Test [Sco09, BDM15, Hot16]. -th [DK17b, Jas16, Jas19, RAB15]. -th-Quantile [Pao10]. -tuple [LB16]. -Type [JZTN17, BQC17, HHC16, YYS19]. -Value [BC13, Fel10, GF17, ZXM16, CM16c]. -Valued [FZ10b, SSA10]. -Values [FBS10, HXL13, LN13a, BL15, ZXW16, CFA17]. -Variance-Component [Are14]. -Weak [HS14]. -Wright [Cah12].

/1 [KC19b, ZXM16]. /1/ [KC19a]. /M/1 [Ye19]. /R/1 [PBC17].

1 [BL19, CM13a, PBC17, SLOT17, SA19b, SRK11, SU14, SK16e, XZD10, YL18c, Ye19, YFV19, ZG19a, ZXM16].

2002/03 [AZN15]. 2022 [AE23c]. 2072 [AE23a]. 2404 [GK19a]. 2828 [Mül12]. 2859 [NN15].

35 [MBS17].

4095 [Ano23]. 41 [GK19a]. 4240 [AE23b]. 43 [SG17d]. 44 [Ano23, AE23b]. 45 [AE23c]. 46 [AE23d]. 47 [AE23a]. 4755 [AE23d]. 4th [MSD19].

= [IBA18].

Aalen [sSW18, She16]. AANA [XDWC18]. Aberration [Bis12, BDM13, HLL15, Wan10a, EIM18, Fon17, LWWW16, ZL16e]. abilities [VV17]. abnormality [SI17]. Abrupt [WK13, BKR19]. absence [YAA^{+}17]. Absolute [Chr10, Tsi14, VSS12, WLW15, WSlT15, ÇE17a, ZZL16]. absolutely [Gen17]. absorbing [HL18b]. Abstract [FP12]. Accelerate [sK11, Ko14c]. Accelerated [BLN13, hFmH14, JMEM14, Ko14a, Ko15b, RM14, TGXX12, UM09, WS11, WXPS12, XFC11, XF12, ZS15, CSY19, DE17a, DW19, Ko16a, LLD17, LSCW17, XK18]. Acceptability [KLS12]. Acceptance [AYC^{+}12, Gov11, GB15, NN13, WT18, ANJ17, KAJH18, NS18, WAJ17]. Accident [MKS15]. Accordance [Aca15]. Accumulated [SMS14]. Accumulation [Pul13]. Accuracy [LH14b, BPV19]. accurate [Cos18]. Acoustic [GGL14]. Across [LCG12]. activation [MOC16, SBLC17]. Active [CS11, YL18a, YL18d]. actuarial [Hon17, JZ16, OL19b]. Acute [HS15c]. ad [Hu16]. Adapted [YC12]. Adaptive [AC11, AAP11, BD11a, BM15, CSC17, CDS19, EM13, GLZW12, GFS16, GMP11, HH15, JS13, KNHM17, Lee13, LW13b, MZ10a, NAD15, PS18, RZ13, WS11, YS19, ZK15a, ZY19, ZX16, ZLWZ10, AN17d, BSL^{+}19, Bak17, BAKZ18, BS19, Haq18, HJC17, KM19a, LSCW17, Mah17, MMT17, OOT19, OTA17, ÖK17, QKCH19, SB17, SFSH16, WWZC18, WH16b, ZX18, ZXW16, ZGK17, ZWZwZ17, CCY17]. Adaptive-Cox [CSC17]. Adding [AW10]. addition [IBA18]. additional [CHM16, MT16]. Additive [BJVV15, BD17a, CZ15c, FDL17, GP12, GL10b, LR14, LCO14, Li13a, Lim11, LSZ13, LLQ13, Mat14, Oli13, PBY^{+}14, PYB^{+}15, She15a, SSM12, WCL10, WK12, WTZS15, WJH15, Yu10, ZZS13, Ai17, AS18, ÁF16, BJVV17, BM16, BCL16, CY17b, Fan18, FC18b, HP17b, HZW17, HAZ16, KN18a, LXL17, Nak19, Pop16, She16, ST16e, WHC16, YZP17, ZF17, ZZ18c]. Additive-Multiplicative [Yu10]. Additivity [BF10]. adept [TS17a]. adept-stratified [TS17a]. adjacent [DM18]. Adjusted [DLP15, GK11, LLC10, ML12, MF14, PDS09, Ree12, RPJ15, SSS17c, WHC16, YZ17a, ZLF14, KOFA18, LLF19, SA17c, Wan17a, WF17, dAdLA18]. Adjusting [FBS10]. Adjustment [Oga15]. Administrative [CGL13]. Admissibilities [TOT15, KT17a]. Admissibility [HX11, Sen16a, Tak10b, TT10, Tan10b]. Admissible [Bay16, LR10b, Mah12b, Sen16b, Sen18b, ZFL13, CS16a, CX17, CH19a]. Admixture [MS10a]. Advances [CS11, Edi10b, HH15]. AEPD [GT15]. Aeppli [Min10, MB14]. Affiliation [LD18b, ZQQY17, LQHZ18, LQW19]. Affine [YDC15, WSW16]. AFT [GSA14]. AFT-Generalized [GSA14]. After [JHP10, QNP12, RPH17, Wu16d]. Against [BC13, Bis12, BDM13, DJ15, DLL13, PS13, Sur12, BDM15, CK17b, Gau17, Hot16, Man17a, Mer16, PMA14]. Age [BR13, CSC13, Che14, CC10d, SCC10, SCC11, CC19, ZW17c, ZMN17]. Age-dependent [CSC13, CC19]. age-replacement [ZW17c]. ageing [HN16, NHAMG17, SR18, YYL17]. ages [JZ16]. Aggregate [HQW15, BNB18, FL17, FL18b, GWZ17, GWP19, LG16b, LYG17, LSL19, YS16]. Aggregated [XYH12, CS18, LW17b]. Aggregation [Hou12, CY17a]. Aging [IRAB13, KRB13, LL17a, MP10b, RAP13, TK14, AAE^{+}19, FPPRG16, SBLC17, SI19]. agnostic [LDL18]. Agreement [Aca15, FGTS12]. Ahead [GB12]. AIC [KYS13]. Aids [ASMZ14, SG10c]. AIR [MSS14a, LVBS10]. Ait [JZ17b]. Akahira [AK10, Edi10b]. Akaike [KS12b]. al [GK19a, LF17]. al. [ST17b]. Alert [DJ15]. Algebraic [GFR12, ZB12b]. Algorithm [Afs13, AW10, BZAZ15, Cha14, CL10b, DL11, GDS^{+}14, HK11, NF11b, NF11a, SBL13, SLL14, SG11, TTZ14b, Zho10, AD17b, AD16, AD17c, BTC16, CSM16, DTT^{+}16, zF16, LZ18, OOT19, OTA17, PC16b, RD16, SW19a, Van17, WYG17, XZD19]. Algorithms [SMS14, GFS16, SA17b]. Aljarrah [Ano23]. all-or-none [GZ16]. ALLHAT [LZY^{+}17]. Allocate [sK12a, Ko14a]. Allocating [ZL15, sK16c]. Allocation [BD11a, CZZZ18, Hou12, HW14, LYH11, LWZ13, YLL19, BS19, CBSN16, HN17a, KAR16, LL17b, SB17, SCYR17, SCYR19, SH17b, TS17a, TS17d, WW17b, YL18a]. allocations [YLS17]. allometric [Sch17]. allometry [NAAC16]. allowing [Bar19]. Almost [BR17, BR11, LYX11, Li13b, Wu17c, XY11a, XM14, Yan18, YZG17, Zan17, ZZ15, AR17b, Cha18, CW19a, HWY17, Li18, LYX18, LPH16, LY16c, LO17, MGN16, MX19, NHV17, SS17b, zWDD19, Wu16b, Wu17d, WAA18, XTL^{+}17, Zha16b]. ALNQD [Ko17b]. Alpha [Gui15, Jon18, KG14b, MFBG13, NAMAK17, LARS19, LAF17, MFPG16, SSB17a, Uya17]. Alpha-power [MFBG13]. alpha-skew-normal [LAF17, SSB17a]. Alpha-stable [KG14b, Uya17]. Alternative [Arn11, PS13, Par13b, RL11a, SA15b, Sur12, ZWLI13, AG17, AE23a, Bic18, BH17c, ÇE17a, DGMC17, FK16, GMMA19, GN16, GH18, GHBM18, GG16a, Has19, Hut19, Kvå17, Kvå18, NRI17, OL19a, Vis18, WCZP16]. Alternatives [BC13, DLL13, PMA14, Sto11, Gau17, GMH16]. Altman [Aca15]. Always [LJ10]. ambiguity [Mal17]. amended [AAL17]. American [LS16b, SL16]. Among [Con12, OQL12, BAR16, Ima18, PF19, ZXSZ18]. Amoroso [IK18]. Amount [MP12, PM12, PL12, LYG17]. amounts [BNB18]. Ampère [Sel17a]. ANA [YJS19]. Analog [NAB12, KV17]. analogs [DW18, DO19]. Analogue [MXD13]. analogues [Rao18]. analyses [Gro18, PG16]. Analysis [AT15, AK14a, ASAK10, AGVGP^{+}11, BDS11, BL15, BIJ17, Ber12, BYDM15, Bru10, CSL17, Chi09, Chi12a, CDG12, Cra14, DBC14, DC14, DS09, DSO12, Den13, DM10, Eid12, FMS15, Fu12a, GT15, GVGP11, GW18, Haj19, HS15a, HD13b, HYK15a, Hu11b, HS13b, HLX14, HS12b, JFS14, JZQ15, KHM12, KMC10, KC19a, KS10a, Kaz11, KY16, Kem11, Ko10b, sK11, Ko14b, KVV11, Koz12, LMCH12, LP10, LSD10, LMC18, Liu11a, MSSU14, MN15, MS10e, NKS15, NRS^{+}14, Oga11, OFS13, PBY^{+}14, PYB^{+}15, QZY12, RSV10, RGL12, Roy16, RVP15, SCSMN11, She10b, SWW^{+}12, SK12d, SS12d, SSM12, SZD12, TSG19, Tsi10, Tsi14, TB13, UM09, VBAR15, Vas14a, WP11, WZY14, WTZS15, WST10, WILL13, WYL12, XLPL12, XLL15a, XT11, XT12, YS12, YC14, YY14, YL18c, YW13, YD14, ZG19a, ZHLZ14, ZWZZ15, ZSY13, ZS15, ZZSP14, AGS18]. analysis [Alv16, dCA19, BSL^{+}19, BP17, BH17b, BCB19, BK17b, CYF^{+}18, CR16, CK17b, CS16c, Che18, CLC^{+}14, CH19b, CTAL17, CM17, D’A17a, DAC17, DMM16, DGMC17, Ery16a, FZ17b, FZ18, FK16, FN18, FA19, FKR16, FTF17, HL18a, HWZ19, IKH19, JX19, JKL19, KAT18, KP19, KYW16, KA17, KDG^{+}18, KS18a, KN17b, LW17a, LSCW17, LDL18, LWH19, Man18, ML19, Maz19, MSD19, MLD17, NS16b, OU16, Pel17, PLDS17, QC19, QCK19, RAM16, RTT17, SS16a, SK18a, SF16, SR17, Sha18c, She17, She18b, SLSC18, SUS19, Tat16, VN17, Vas16, WXS18, WH18, WSS17, YHS17, Yam18b, YL17, Ye19, YM17, YYLM18, ZJLO17, ZX18, ZCC16]. Analysis-of-variance-type [JZQ15]. Analytic [OQ17, PAvS^{+}19]. Analytical [CMN^{+}18, OL19b, ZZ19c]. analytically [Cos18]. analytics [RBK^{+}19]. Analyzing [Afs13, BG11a, Coo12, EM18, Gui15, HX14, Ju15, KH19, SG16b]. Anbar [SR11]. ANCOVA [SAJ13]. Andersen [FWZ^{+}17b, HJ14]. Anderson [Lie16]. Angular [SO14b]. Annuities [QWZ14, YSK15, QWWW19]. Anomalies [MN15]. ANOVA [HAP17, Hot16, Jeo12, MGG^{+}17, SAJMF12, WZ16, WK19]. Answer [BBP14]. Any [Ell11, KK14, DZH^{+}19, Pro18, TH17]. APND [SAB18]. appearances [Kon16]. Application [AJGG14, BCG11, BEKL13, Bru10, CLC^{+}14, CDZ14, DS09, Far12, Fu12a, GJ15, GL19, GL15b, Gue10, Hou12, Hu11a, HS13b, KO10a, Kim12, KT17b, Lef10, LVBS10, LSD10, LS11, MKS15, Mae10, MA11, Mea15, NAMD14, Oga11, OJR15, ÖD10, PP12, PKM17, RPY^{+}13, Raz14, Sei13, SU11, SDJ15, SG10c, SS10b, Smi19, SR11, TWH10, TV11, Tsi10, VT15, VLD15, VBATK11, WS12a, Wee12, Xu10, AXLJ17, AJT16, ANS17, BKM18, BAF19, BKGD17, BBS16, CR16, DTT^{+}16, DA17b, Ery16b, FFN^{+}19, HF19, JN18, JPAB17, JCN17, Jon18, KD16a, KCY^{+}19, KN16a, LGLX18, LS16b, MK17, ML19, Mar16, MST19, NC18a, PT16, RLSL19, RMN16, Sel17a, TSG19, TME19, TE18b, WHYW18, XY19, ZRL19b, ABTR19]. Applications [AFL13, AO12, ABLR15, AGVGP^{+}11, BS12b, CNC14, CL10c, Coo13, CVC^{+}12, DO13, Gia14, GDCOS13, HW11, HWM12, KSDR15, KDK14, KR15, Kur10, LX13, Lop11, MZ11, Min11, Muk11, NCO15, NSV12, OBPL15, PT10, PSS^{+}15, RLL12, SGD11, TT11, Vij11, XY11b, YVCA17, Yan19, YW11b, YMF15, Yar14, Zha15a, ACO^{+}18, AFL19, AABL17, AG16, ARB19, AE23d, BP16, BB19b, dSBCO18, CKS17, CXC19, CLC^{+}16, CARO17, CAPO17, CBOR18, CY16, DYW18, GDGG17, GDDC19, Hon17, INOC17, Jon18, KD16b, KD17b, LPC17, LK19b, Lou19, MAYH17, MML17, MT19, NAMAK17, OMFB17, OL19b, PWLO17, Rah18, RMNA17, RSAN17, SP16a, Sel17b, Sha18b, SC19, SW19b, TCA^{+}16, TCM^{+}16, WWW17, WCW17, WH17a, WER^{+}19, WV16, Yan18, ZMMX17, ZWZ17]. Applied [GAK14, MJ15, ST13, SS14a, Bic17, Van16]. Applying [Chi09, Cui12, HHC^{+}17, Lu16c]. Appreciation [NS15b]. Approach [Alm11, AH14, BG11a, BG11c, CKLS11, Chi10, DZ14, DC10a, DMST12, FZ10a, GC14, GHY11, HT10, HC12, IRAB13, JAA11, JHP10, KR14, Ko14a, LW13a, Mån12, MS12a, ÖD10, Pao10, RSV11, SAJ13, SO14a, SW14c, Sim14a, SE13, SR11, SS12d, WP11, WL14, YOD10, YCL14, YC11b, ZZSP14, AE17a, AS19, AAK16, Alv16, BT18a, Bib16, BRAS17, BH17c, BCB19, CO16a, CLZ18, CTAL17, ESAKK18, FSSW17, FPM19, FKNR17, FK16, Fon17, FGL18, GN16, GR18b, GPP19, Hen17, Hot16, JJ19, Kam16, KM19a, KCL^{+}16, Kon16, Mah17, MN17, MN18, Mna17, MSEN18, MGG^{+}17, NSSS17, OU16, OH16, PMF17, PAvS^{+}19, PNPB19, PM17, RSC16, SKK18, SNAN17, SNWL17, SFSH16, WLZ17, WW17b, WF17, YC18, ZXM16, ZY16]. Approaches [BM10, BPSV14, DTE13, sK12a, HMT^{+}17]. Approximate [BL15, CL10b, FLK12, FA19, KXX11, KZ15, LVVN11, Miy10, SP13, BT18a, ESAKK18, XJW19]. Approximated [Hua17a]. Approximation [CG11b, Dai15, ES11b, HKR15, KRH12, Lie16, Lim11, RN13, SSC18, SF12b, Sch12, SY14a, SW14b, SI13, Tee10, Tee14, Tsa10, YHS13, YS15a, YLW17, ZSL17, AD16, CN17c, JN19, KNC18, LL17e, aSaCN18, SL16, SN18, TS19, Udo18, YHW17, YZ17b, ZK16, ZZ19c]. Approximations [AE13, AAE13, BEZ11, CN15, DK13, EdSMR14, FR10, Gar14, Lin14a, MZ14, RS10, ABES18, CSS16, Miy17, NM19, ZYH18]. ARA [CDG12]. Arbitrary [FP12, Bar17c]. ARCH [CH17, KSDB18, XHL18b, YC18]. ARCH-M [XHL18b]. Archimedean [LL17a, MKD11, MP10b, RAP13, Wys19]. arcsin [Hom17]. Area [BAK13, Con10, PC12, Tun15, CS18, CM16d, KWM17, NNvRS17, NVS19]. Area-Level [PC12]. Areas [Jor13, PMC13]. Arguments [ZB12b, Dru17]. arietinum [AIIK17]. ARIMA [DK17b, Ho16, YC18]. ARIMA-ARCH [YC18]. arises [Rah18]. Arising [LBL^{+}10]. ARL [Rav19]. ARMA [GS13, AH18b, BB12, El 15, GS15a, NV15, Pan19, TT11, Tun15]. Array [HK11, RN13, BM17, Mat18, ZQZY18]. Arrays [GFR12, HYK15b, LZH15, SWLW14, Sun12, WHVH13, CLLW16, Fon17, HL16, MWW18, PYL17, SAB18, SS17b, Wan17f, WW19, WV16, YJS19, fZWD19]. Arrival [Kem11, El 19, KY16, NC18a]. Arrivals [MGF14, FL18b, FL19, TSOT19, ZG19a]. Artificial [ZA14]. ascent [Kia17]. Aspects [ASB13, FM11, GEA^{+}10, NSP12, KN18a]. Assays [Dji10, SB16b]. Assembly [MS10b, Mat11]. Assess [MS12b, RM15]. Assessing [Con10, GB12, KCK10, RS10, SYP13, UPA13a, VRC13, YAA^{+}17, YY16, NKG16, NSR^{+}19, PF19]. Assessment [AB13, CD12, FU10, GdC12, Lin14b, LC17, TEF12, BSGS16, CCCT16, FG17, LW18a]. Assessments [ZZL12, AP19, LW17a]. Asset [LPS11, ST13, BBS16]. assets [WQW16]. assignable [PMKF17, SMSF19]. Assignment [SW14a]. Assignments [OQL12]. Associate [SG18b]. Associated [AT19, Bos15, GZSS14, LWY10, LQ13, LW14, LW15, QL11a, SD14, VEKH15, XY09, XY10, dXcY12, XY14, XM14, AOÖ17, AD16, AD17c, Car19, FST16, GT17, HA18, Ko16c, Ko18, LQ17a, Men19, NHV17, PFD19, PL17, SG18a, SG16d, SS17b, SYW16, WP18, Wu17c, XTL^{+}17]. Association [APM10, CVC^{+}12, KG14b, Li13b, YL15, Coo19, HSW17, Kvå17, Kvå18, LYR^{+}18, PCL16, WH16c]. assumed [XMP17]. Assumption [BDM15, HYK15a, KK14]. assumptions [LLM16]. assuring [BJU18]. Asymmetric [ADP10, CGW15, Han15, Ko13, MSS14a, OPSGB12, Tak10b, Web12, WH16c, Zha13, AO18c, BQC17, BMFS18, HWH18, KM16a, KOA16, LGLX18, MS17a, PV16, RWY16, SB16b, XQ17b, You17]. Asymmetrical [LZH15]. asymmetry [FPZ16]. Asymptotic [AS19, Afs13, AEHQ17, AKK09, AAK16, BS12a, BNE15, BCH15, BS15, BL13, BD17b, BG16b, CC10a, CD17, CM12, DDD17, ES11a, EdSMR14, FJ11, Fat13, FF19, Fu11a, FLY16, FL19, GLZW12, GFA10, GMS18, GW15b, HN14, HS15d, HGZG15, HYS17, IS11, JYX16, JXW17, KAW^{+}18, KKS18, KS18b, LH11a, Lem14, Li17c, Lia11, Lin19, Liu11b, LN13b, LS12, Lu14, LS15b, Lu15, LQHZ18, LQW19, MFS12, MCX10, Miy10, NY11, Oga11, Oga15, Oga19, PZ15, Par12, Rah18, SY11, SY10, SNWL17, She18a, SS16b, SP13, Su17b, Tak10b, TK13, TW14, TN15, WCL11, WLF11, XWJ10, XJW19, XX18, XHL18a, XZA18, Xu10, YHS13, Yan12, YWS14, YMF15, YWL17, YDL14, ZLN11, ZL12, ZY13, ZC17a, ZDW17, AHKO16, AO17a, ABK17, AK19, AABL17, CMN^{+}18, CZ19b, GWP19, HZS17, KKKÖ17, KSBN17, KTBN17]. asymptotic [KSDB18, KVR17, KH18, Li16, LG16b, Lu16b, LQYZ17, Sch16, Sel17b, SNW19, SYW16, SY17c, SA19a, TY17, Tch19, WdX17, YHW17, ZZ17d]. asymptotical [WCW17]. Asymptotically [ABC12, Kor11, Mar15, VEKH15, YD14, Ko18, LG16b, NHV17, SS17b, Sze17, XTL^{+}17]. Asymptotics [DA16, EZ14, GF17, HJ14, HL15, Li14a, Lii17, LM14, Tak14, Tan18, Wan13a, WLY14b, Yan15, YA10, BSW15, CWW16, GCW19, LGG16, LSL19, TN19]. Attainability [Raq10]. attenuation [Bur17]. Attraction [RP15, Gra17, Sch16]. Attribute [AGR15, SHANA18, SK15b, JBA18, KM19a, LHS19, ST17b, SSSK18, SSS19b, TS17b]. Attributes [BQDH12, LC15, BMQ16, UB17]. AUC [ZZ19a]. Auge [GEA^{+}10]. augmentation [Kia17]. Augmented [AAG12a, AAG12b, dBCR15, DT12, LOHW19, RAAJ17]. Augmenting [RSV11]. Australia [BHKHM13]. Australian [HB17a, HB17b]. auto [KM19b]. auto-covariance [KM19b]. Autocorrelated [Alk10, AMH14, NKS15, ÖA15, HJT19, MATC18, MB17a]. Autocorrelation [LIPMF13, SN14, FJS^{+}18, Lu16c, Sun17b, WT18]. Autocorrelations [Sim12, BBC18]. Autocovariance [GS15a]. Automated [GDS^{+}14]. Automatic [YSC17a]. automorphic [LYX18]. Autonomous [Fan15]. Autopersistence [Hud14]. Autoregression [MT12, SG15, Are16, HCW18, ZL18]. Autoregressions [BCH12, Cav17]. Autoregressive [AR13, AT09, BH12, BM11, CCL14, CTQ11, CG11b, CSF12, DP13b, Fat13, GSB12, GB12, HL15, JRM10a, JP13, JT11, Kac14, Li14a, LH11c, LR10c, Mia11, MXD13, MSP12, PSP12, Raï15, RON12, RZ13, RNB12, SC10, SA15a, SSS14a, SP11a, WW13a, WW15a, XLPL12, XLL15a, hYhW14, ZY10, ZWZ11, ZW11, ZWPZ15, ZZC13, ZB12b, BP16, BPS16a, Bib16, BBS16, BCK17, Che16d, DA16, GDPR17, HZS17, HL18a, HS18b, JDSS16, KM18, KM19b, KLKK17, LPC17, MSL17, MR16b, NRP16, Pro18, SN17, VN17, YW17a, YWLZ17, ZDZ17, ZWP17, ZWPB17, ZL16g]. Auxiliary [ARD14, AC15, DM11, Gro10, GSSCR12, GS15c, LY11b, MS15, QL11a, SK13a, SS15a, SV10, SS13b, SS13d, SS13e, SSS14b, SST14, SS14b, SS15b, TSS15, WCGZ12, XLL15b, YUSC14a, AHH16, ASeMJ17, FDL17, GK19a, HKH17, Haq18, KJ17, MSK17, Ozg19b, RK19, SG17a, SG17b, Sha18a, SLF18, SHM16, SP16d, SSS17b, SP17b, SKM17, SSS18, SS16g, SS16i, YAA^{+}17, YSG17]. Availability [QC19, QCK19]. Available [SS15a, SMK17]. Average [CTQ11, Cox15, ER11, Hua14a, HWM12, Lim11, MZ10a, Men10b, Men13a, MS10c, MS14c, OP11, PSP12, SA12, SJS15, SL13, SHH13, SP11a, Sun11, TKW12, WW15a, XLPL12, pZ13, AP17, AAJ16, Das17a, DTWW17, GFM17, GR18a, Ko17b, KPZ16, Lu18, NO18, NP17b, OPK18, RSA19, STD18, Sun17c, WT18, WN16b, XN19, YYS19, YKL^{+}17]. average- [YKL^{+}17]. averaged [KWM17, LCZA17a]. Averages [HWM12, ZLZ15, BP18b, Roo17, ZSL17, ZZ17d]. Averaging [DZX13, Fat13, LN13a, Liu19, WZ13, AJ18b, AIIK17, Bic18, CSC17, FG17, HN17b, ZCHR19]. Axis [MF10, Gey18]. axis-parallel [Gey18].

B [AE23d, SG17d, JHM19, KN17b, LWD^{+}18]. B-spline [JHM19, LWD^{+}18]. B-splines [KN17b]. Back [RBK^{+}19]. Background [CXS11]. backtesting [ZN18]. Backtracking [YW13]. Backup [Lin14b]. Backup-routing [Lin14b]. Backward [NY11, CJ17, LW18b]. Bagging [KLS10]. Bahadur [CLX14, XY19]. Baker [BB14b]. Balakrishnan [NCO15, AEN16, CB16, SG18a, SD14, TN16]. Balance [CG12, CG14a, Gra12, KS12a]. Balanced [AJV11, AA11a, AIA14, BA10, CL14, CS14a, FM12, GHY11, HYK15b, ITG09, MGP17, MGM13, NS15b, Par13b, QLZ14, VJV14, VB13, YH11, ZAV14, ASR^{+}19, BP17, BAT^{+}18, CS16a, CX17, CH19a, HH16, HLY16, Kha19a, Kha19b, KHY17, MPG16, MD17a, NNZA16, ÖK17, PY16a, RAST18, RRRA18, Su17b, VV17]. Balancing [KDGV16, OH19]. balking [KS18a]. Ballot [NIN17]. balls [NA18]. Banach [STD18]. Band [HKLW12, SLSY14]. Bands [AKL15, Zha14, Ker17, Lin16, Son16]. Bandwidth [CT13, HKV12, ALR19, BAKZ18, KH16, Pak17, Ten17]. Bandwidths [MFS12, DS16b, DS17e]. Banerjee [HME17, PKM17, PMKF17, PM18, PMFK18, RMM17]. Bar [CY10, LJ10, NGT13, YWLY11]. Barnard [AA11b, AASRZE17]. barrier [FNM19, hXZ17, XDW19, ZZ19c]. Barriers [CCH^{+}10]. Barroso [MBS17]. Bartlett [Wan17a, ZJC10]. Bartlett-corrected [Wan17a]. Based [Afe13, ADP10, AC11, AH13, AK13, AF10, AALS12, BTK10, BS12c, Bar10, BR12, BT13, BB14b, BAT15, Ben14, Bic12, Bur09, CYE15, Cey11, CT13, CB14, CW14, CS15a, CS15b, CF13, CS14b, CB10, CB12, CM13a, CS11, CL15c, CDZ14, CD12, CVC^{+}12, DC10a, sDsWcCjK14, EDAAS12, EWRZ14, FNDM12, Fon14, GMN13, GC12, GB12, GK11, GG10b, HME12, HC12, HS12a, HCS10, HWM12, JCY10, JC12, JHP10, JO10, KRH12, KC12, KS15a, Kha12, KS14b, KRB12, KJD^{+}15, Kvå15, LMCH12, LWL13, LWY10, LPLW13, Li14b, Mån12, MP14, MGAf14, Men10a, Men10b, Men13a, MKD11, MN15, Nas12, NN13, NKS15, OM13, PS12, PMC13, RR09, RMC11, Sag15, Sal15, SBR14, SG10c, Sha10b, Shi10, SK15a, SW14c, Ste11, SYP13, SZ13, SL15, TTZ14b, Tor12]. Based [cTtL14, VML^{+}10, VLD15, VB12, WWL13, WLY14a, WQ14, WWT15, WH15a, WLW15, XLL13, YXG11, YSR11, YDC15, ZLZ15, ZXD10, ZSY13, ZHL15, AO18a, AM17a, Aga17a, AMP17, AR17b, AHAH17, AEHQ17, ABK17, AAO17, AHP19, AH14, AB19, AAJ16, BAL18b, Bal18a, BBC18, BK18, Bar17c, BH17a, BT18a, BAB^{+}18, BSGS16, BMR17, BT16, Bic17, BRY17, Bow17, BAW19, CZCM19, Çel19, CB17, CCCT16, CM19, CCN17, CCM19, DWW18, DZH^{+}19, DW19, EIM18, EDKAEN17, EP17, FGP16, FY17a, FPM19, FPPRG16, FGGL18, Fon17, GN17, GL19, GFM17, GF17, GFL19, GX18, GK19c, HP16, HB17a, HP17a, HB17b, HP17b, HBMAO16a, HKH17, HW18a, HA17, Hom17, HZ18a, HJC17, HZ18b, Huh16, HBAJ17, HN17b, IES^{+}17, JJ19, JBA18, KKS19, Kay16, Kay18, KM16b, KLJP16, KQ17, KL17a, KDG^{+}18, KL17b]. based [Kum17, KDN17, LK17c, LK17b, LHS19, LLHZ16, LB16, LX17, LLLL18, LPDD18, Li19, LPDD19, LTZ19, LC17, LLWK18, LWD^{+}18, LAW^{+}19, Lu18, LQ19, MFJ17, MC18, MM17, ML19, MBV17, MS17b, MST19, Men13b, MAA17, MRA16, MHF16, MR17, NLRD17, NC18a, NB16, NP17a, NS17, NS18, NRS^{+}14, Nou17, OPK18, OH19, Ozg19b, ÖGGB16, PSF16, PB17, PK18, Qom17, QK17, RBK^{+}19, RAA17, RAM16, RTT17, Sab19, SYC19, SCYR17, SCYR19, SDS17, SS17a, Sha17, SBA17, SV17, Sha19, SA16, Sha16, SC19, She16, SI17, SSS17a, Sre18, Sun17b, Sun17a, SR18, TS17f, Tsi15, UB17, UU17, Wan16b, Wan16c, WZX17, WT18, WXS18, WGW19, WGG^{+}19, WYMZ17, Wie18, WBGS19, WCK16, WZZ17, Wu18a, Wu18b, XQS19, XX18, YC18, YLG16a, YSG17, YCAJ18, YYLM18, You17, ZBA17, ZLT17, ZT17, ZBTL17, ZX18, ZWPZ15, ZL16c]. based [ZCCW18, ZH16, ZDS16]. Baseline [HS15c, HD15, TWH10, FK19, HP17a, PBL19]. Basis [FG11, KK11a, YS15a]. Basket [GDW18]. Bassett [IS11]. Batch [Kem11, Par13a, KY16, TSOT19]. Batch-Oriented [Par13a]. Bathtub [CZ15a, SK15a, AE17c, JC18, MKSS17, NSR17, NB17, SSSM16]. Bathtub-shaped [SK15a, AE17c, JC18, MKSS17, NSR17]. batting [Das17a]. Bayes [AB09, AB17, Ahm16, ADP10, AHAH17, BA10, BG19b, Bic17, Bic19, CL10a, Che11, CW13, CLC^{+}14, Che17a, CBM14, EP11, GMR16, GKP16, HK15, HASA10, JMEM14, JSC17, LQ17a, Mar15, MT16, PJ10, QLZ14, SA19b, SUS19, UM09, WS14, WW17a, YW17b, ZK13a, Zha17d, ZXSZ18, ZRL19a, Zin18]. Bayesian [ANJ17, AM17b, AXLJ17, Ami18, AD19, ADO13, Aoy10, ASNT15, AdBPdBP18, AIIK17, AGVGP^{+}11, Ayk19, BS12c, BQC17, BD16b, BN18b, BAKZ18, Bic18, BH17c, CYF^{+}18, CO16a, CG14b, CS14a, CS16c, Che10a, Che18, CM16c, CTAL17, Con10, CDG12, DS09, Das16a, DP10a, DE17b, DMST12, DPS11, DPS12, DR17, DGMC17, EDAAS12, EDKAEN17, Far11, FM10a, FA19, FG17, GT15, GC14, GN17, GVGP11, GPP19, Gun16, Gun17, Han11, Han12a, Han17, Han19, HS15b, HD17, He17, IP12, JN18, JHP10, JSC17, JC18, KHM12, KP19, KS15a, KR14, KQ17, KLKK17, KvdVvZ13, Kos14, LKK17, LGY17, LS15a, LS17a, LSC13, LWHB16, MSSU14, MN17, MV13, MA18, MS12a, MLD17, MAA17, Miy17, ML13, MS10e, Mur13, Mur16a, NN11b, NS18, Ng12, ÖD10, PG18, PNPB19, Qom17]. Bayesian [RSV10, RvdM14, RGL12, RM15, RM14, SB17, SG16b, Sco11, SB12, SK17, SK18a, SA12, SA15a, Sha17, SBA17, Sha19, SR17, Sha18c, ST13, SCC11, SUKK19, Sil18, SJ10, Slj17, ST18, SK12d, SS12d, SA14, SSH10, SK10b, SBLC17, TLT17, Tsa10, Tsi10, Tsi14, VN17, VBAR15, WLC11, WW15a, WNL18, WLS18, Wei14, WILL13, XLPL12, XLL15a, Xin15, XQ17b, XZ16, XT11, XT12, YOD10, YCL14, YW17a, YJSL19, Yao12, YA17, YM17, YC11b, YCL17, You17, YD14, ZY14, ZJLO17, ZHW17, ZLT17, vZvdM17, vZvdM19]. BDR [ÇE17a]. be [PT16, Tan10a]. BeadArray [CXS11]. Beanplot [JTR15]. before [Hot16]. Behavior [BS12a, FZ10b, FZ10c, LPN14, LS12, MSW13, Par12, Sal15, TW14, BG16b, CY17a, FL17, HZS17, HC16b, KVR17, Li16, LG16b, Roo17, SA19a, SA17c, VFP17b, ZC17a]. Behavioral [DS09, PAvS^{+}19]. Behaviors [GFA10, MLGZ14]. behaviour [KAW^{+}18]. Behnken [AAG12b, RAAJ17]. Behrens [SP11b]. Beliefs [GGFL11]. Ben [PM18]. Benefits [Pel17]. Benford [Bal15]. Bennett [Zhe18]. Berkson [WF17, WCF18]. Berkson-type [WF17, WCF18]. Bernoulli [AD10, DDGM13, HF19, HP14, Ili16, JX19, NBS13, NRI17, PHH18, SO17, Tee14, ZZ17a, Zör14]. Bernoulli-schedule-controlled [JX19]. Bernstein [Kre18]. Berry [Hui10, LPLW13, LLM16, MXD13, Sze17, TN13, TCN15, WWH15a, YWH14, ZL16b]. Bessel [GW18]. Best [BDKM12, CHP14, GGS13, HBMAO16a, jK15, LWR11, SU14, YCL14, HBM16, Len17, WHZ17]. best- [Len17]. Beta [BSCS11, BREA11, DC13, Duf10, GN12, HSMK13b, JBH12, LH14a, MM13, NABA15, PBS12, RCC12, SGD11, TWH10, TK14, VMVD14, Wil09, Zin12, CM18, KKKÖ17, MBS^{+}19, NAN17a, OHKK17, SS19, SG16e, TSG19, ZWZ17]. Beta- [TK14]. Beta-Binomial [TWH10]. Beta-Dagum [DC13]. Beta-Exponential [NABA15]. Beta-Poisson [LH14a]. Better [LJ10, LMM15, BAW19, Ram17, YSD^{+}16]. Between [Aca15, APM10, Atk12, DK09, DI10, HN14, Kit15, KLM10, KZ15, Mah12a, MCY14, MCX10, MS11c, Qum12, Tia12, Tun15, Vij11, VB13, Yu12, ZC11, AdBPdBP18, BN18a, BBHN18, BBG18, FLK12, HNB19, Ima17, LSY17, MMN16, MHA19, Nag18, OQ17, SH17b, Sil18, Sin16, SA19b, SN14, SR11, Tri17, WT18, WH16c, XHH17, YF17, ZMV^{+}17]. BFGS [YW13]. Bhattacharyya [HK15]. Bhattacharyya-Type [HK15]. bi [LQYZ17]. bi-degree [LQYZ17]. Biadditive [BF10]. Bias [BGG11, BG19b, Chi11, Gil12, GFG13, GFG16, God16, GMP11, KYS13, KNH17, LG14, Liu11a, Oga15, Roo18, SG16c, SHHH09, XY11a, Bie16, EAE18, HJW19, IK18, KNC18, MF19]. Bias-Corrected [KYS13, GFG16, EAE18]. Bias-reduced [SG16c, IK18]. Biased [CS15a, CS15b, CS14b, ET13, Ahm16, ARJF19, CCN17, CLZ19, Fak15, FZ17a, GK19c, JRF17, KL17b, KA18, MZ17, PB18, ZZ18a]. Biases [MF19, Oga19]. biasing [ANA17, KA18]. BIC [ZX18, ZP17]. bidimensional [Bib16, FL19, Li17b]. bifurcating [Dan18]. big [RBK^{+}19]. Bilateral [Kim11, LD14]. BILINEAR [BG18, BB17a, BG19c, LGY17]. bilinear- [BG19c]. Bimatrix [BREA11, GN12, NAN17a]. Bimodal [AT17, BMFS18, Cah15, HEB16, OMFB17, dSBCO18]. BINAR [JKS18]. Binary [BBB10, BCCR18, Dem14, Fed14, KL10, KZW09, Lu10, Lu11, MP15, SRT14, VML^{+}10, VMVD14, Wit14, YS15a, YC14, AO18b, AEA17, CS16b, Den17, Gen19, HH16, LC17, LCZA17a, LWZ17, MA18, OD19, PP17, SB17, TR16a, TH17, ZY17b]. Bingham [Tsi18]. BINMA [KJS19]. Binned [sDsWcCjK14]. Binning [MS10b]. Binomial [AMV15, AO12, CHM16, DZ15, FWH^{+}15, GRV15, Han12a, IO13, Imo15, KXX11, KZ15, MO14a, OCIS12, PD10, RNB12, SCMR15, SG14, SSON11, TWH10, Tee12, WCGZ12, XYH12, YHA10b, YCL17, Zör14, AJ18a, CRTdC17, CO16b, CG16, CG17, CCOB16, CW19c, DDD17, Gen19, INOC17, JM16, NC18b, Oga16, SG16b, TMZ^{+}17, Xio16, ZRL19b]. Binomial-Poisson [CHM16]. binormality [WNL18]. bioassays [Che16a]. Biological [Pap13]. Biomarker [sK11, Ko14c, LK14]. Biomarkers [Ko10b, Ko12, Ko14b, Ko16a]. Biparametric [CTAL17]. Bipartite [AJV11]. Bird [LS16b, El 17]. Birnbaum [CLC^{+}16, AALS12, CCOB16, Kam14, KLJP16, LF10, MFFMA17, OMFB17, UM09, ZHZA18]. Birnbaum-Saunders [CLC^{+}16, ZHZA18]. Birth [Ano13b, MS10d, BER16a, DP18, KD16c, OAOQ16]. birth-immigration [DP18]. births [DP18]. Biting [Far12]. Bivariate [AT11, AAO17, AH17a, BG11a, BCS10, BB14b, BCK15, CSC13, CJ19, DMS14, Dem14, DC10b, DAL12, Dom11, DJ10, DS14, FI17, FNDM12, FV11, Fer14, GN15, HD13a, HS15a, HS15b, IAA15, JBH12, JRM10a, KP13, KK17b, KK18, LD13, LSC13, MFSAS10, Mey13, MKD11, MB14, Mod12, MP10b, NH15a, NSL14, PKJ14, PJL13, PJ14, RR09, RR13, RR14, SG14, SO14a, SDJ15, Sha10a, She14a, She10a, sS12b, She13, TV11, TEF12, VRB13, VEKH15, XFC11, YSHL11, dCOSW13, Ala17, AGA17b, AG17, Bar19, BAF19, BCK17, CSL17, Ery17, FF19, GAS17, Gen17, GDGG17, Gup16, HV17, HB16, HL17, IAB16, JCN17, Jon16, Key16, KLJP16, KG18b, KN19, LAF17, MFJ17, MDA16, NRP16, NA18, PG18, Quo17, RPH17, SS19, SX18, She18b, SLSE19, SM16, Sre18]. bivariate [SI19, TSG19, TS17e, VT17, YYL17, ZYH18]. Blackwood [HOK17]. Bland [Aca15]. blank [BSH17]. Blest [RPJ15]. Blind [Ben19]. blinded [ST18]. BLINEX [CBM14]. Block [AJV11, BJVV15, Bis12, CL14, CM13b, DP14, DDM10, FM14, GCR12, GN10, Hu11a, HS14, Igl13, Ju15, KLS12, LX10b, Man17a, PPMD19, Pes17, QL18, Shi10, TIA11, BJVV17, FM17, IBA18, Kha19a, Kha19b, MGP17, Man17b, RMN16, Ros19, SMN18, ST16b, SB16b, SM18, ZS17a, ZZ18e, ZZ19e]. Block-Block [Igl13]. Block-Scalar [CM13b]. block-sum [Kha19a]. Block-Thresholded [LX10b]. Blocked [CJ15, Jac11, Pre11, WW13b, BO17, JJ16, JKD17, ZL16e]. Blocks [AA11a, Hua10, DS17a, HH16, JKD17, MEZ16, Oll17, Wan17c]. BLUE [LX13]. BLUEs [GPÖ14, HMP12]. BLUP [LX13]. BLUPs [WL13]. BMS [MSB19b]. Board [Ano11e, Ano12d, Ano18, Ano17a]. bond [KYW16]. bonds [XDW19]. Bonferroni [WX15]. Bonferroni-type [WX15]. boosting [ZZ17c]. Bootstrap [Bon14, DL11, GMP11, HNV13, Igl13, JP13, JZZZ18, Ju15, MS17a, MG15, Qum12, SAJ13, SP11b, Tat16, TEF12, VV12, WH14, YH19, BRAS17, FLY16, GP16a, Li17a, Pes17, QL18, Xu17b, XT17]. Bootstrapping [DW18, DO19, HS14]. Borel [TN13]. Both [Abd15, DZ15, GN15, RL12, AHEA18, FWLL19, FP17b, Gil14, LL19c, QK17, SKM16a]. Both-Sided [GN15]. Botswana [AZN15]. Bottom [BCCH14]. Bottom-Up [BCCH14]. Bound [BC13, CEK11, CP14, Hay10, Hui10, LPLW13, LDY15, Lu13a, NR12, OQC14, RN13, Tee12, TN13, Yam12, YWH14, EG18, GR18a, Kus18, LLM16, SN16, SN18, SNS19, TS19, YYS19]. Boundary [Lu14, ZB12a, CW19c, Ten18]. Boundary-Respecting [ZB12a]. Bounded [BCG11, IF18, Par13a, TN10, Tun15, Gen17, TKP17]. Bounds [AJ11, Bia17, BCDLA12, CN15, EE11, KO10a, KM15, Kvå15, MFSAS10, Raq10, Ryc11, SASMF13, SB14, Tee10, TCN15, VYZ14, WWH15a, ZZP10, AA19, BAB^{+}18, CN17c, FF16, Gor17, KNC18, Kar17, KR17, OQC17, Raj18, Sun19, ZL16b]. Bowley [SSS14b]. box [Çel17, AAG12b, GPÖ14, GE11, Hu11b, LH11b, RAAJ17]. Bradley [HOK17, Yan16]. Brain [WS12b]. branch [Dan18]. Branching [Li11, MO14b, Rah12, Yar11, Yar13, Yar14, FN18, Fro17b, KD16a, Rah16]. break [Sen18a, XP18]. Breakdown [MX15, ZLL11, SSY18]. breakdowns [JX19, SLOT17, YL18c]. breaking [SAD16]. breaks [BKR19, HS18b, JZZZ18]. Breeding [LS16b]. Breslow [LN13b]. BRICS [PLC17]. Bridge [HSHW14, WSW11, ZC17b]. Bridging [Alv16, KM16b]. brief [AH17a]. Broad [BCH12]. Brownian [Ai17, AS18, CLV14, Dai15, Dai16, HAP17, HW11, Hu17, Kaw13, LL17e, LTC17, LS12, Lu14, LS15b, Lu15, Lu16b, Man11b, MS11b, RS16, SSC18, SY14a, SY14b, SXZ17, Sti17, XML16, Yu19, Zha13, ZZ15, ZLD15, ZLD16, ZC17b]. BSDEs [SJ19]. Building [PP10, SA14]. Built [Bru10]. Built-in [Bru10]. burden [WER^{+}19]. Burn [CF12, CF14]. Burn-In [CF12, CF14]. Burr [ACO^{+}18, BAT15, GdSC15b, HME17, RAK15, SMSF19, STRD17, WL10, Wat11]. Buys [NIN17]. Buys-Ballot [NIN17]. by-claim [Li17c].

C [Ano23, AE23b, AE23c, Vas16, KS18a]. C. [GPÖ14]. Cadence [KS10a]. Calculated [Aca15]. Calculating [KRMG19]. calculation [Bar17a, CC16c]. calculations [LCZA17a, ZRL19b]. calculus [LTC17]. calendar [MC19]. Calibrated [DSSL15, Par15b, SSS19a]. Calibration [DB13, DFH16, Koy18, NSSS17, RSC16, SKHL10, SHLK10, SK15b, CO16a, Cle18, HA16a, Ozg19b, SS16d, ZXW16]. Caliński [Koz12]. call [JN19]. Calving [GL15b]. CAN [SU14]. Cancer [BCD12, BHKHM13, Con12, KC12, MSSU14, RPY^{+}13, Sha18c, She17]. Canonical [FN09, Hu11b, OFS13, AO18c, Car19, Oga19]. canopies [DBC16]. Cantelli [CYY14]. Capabilities [UC15]. Capability [AO18a, CC15a, CC15b, DZ12, Gra10, Gra11, NNA12, Sim14a, WH10, Yen12, AK19, BAL18b, Bal18a, CC16b, Has19, HBAJ17, MBG19, NS17, TS17e, UB17, YCAJ18]. capacities [Vas16]. Capacity [Vas14a]. capital [TMZ^{+}17, YWX17]. Capture [WBH10, ZSS15]. Capture-Recapture [WBH10]. Capturing [Otr15]. Carcinogenesis [KC12]. card [RZHC16]. Cards [ASN12, BSH17]. care [BKGD17]. Career [AK10]. Carlo [Chi12a, Hol15, OOT19, ÖD10, XM17a, Zha17c]. Carter [DHP^{+}16]. Cascading [QZZN10]. Case [AZN15, BOT17, Ber12, BBP14, Che11, CHP14, CDZ14, DC10a, DK13, EWRZ14, GCR12, LKN15, Lin14b, Liu11a, LSC13, MTM13, MM15, MADSAS10, QL11b, RVD14, SA14, TTC17, Tan10b, WDZ11, BG19a, CY17b, CSY19, Che17a, CW15b, DKY16a, DCZL17, Gol16, Hen17, MZ17, OTA17, SC18, SBLC17, Van17, YC18, WTZS15]. case-cohort [CY17b, CSY19, MZ17]. Case-Control [WDZ11, DCZL17]. Case-Crossover [LKN15]. Cases [KM15, LL11, DKY16b, Sha16, Zar19]. casex [WN16b]. Cash [PK11]. Casting [BCH12]. catastrophes [Bur19]. Categorical [AK14a, AH14, Bon14, Cag12, CDZ14, GCR12, GD11, LB11, SHHH09, BH17a, GLY17, HSW17]. categorizing [Gol16]. Category [FGTS12, WNL18]. Cauchy [AGVB10, Coo13, CARO17, – Dor11, KP12a, NY16, Pun13]. Causal [Chi09, Chi10, GZ16, MMP14, RMP12, GZ17, Wan19a]. Causality [Mån12, MSS14b, KA17, Zha17c]. cause [SDS17, WP18]. Causes [Con12, RM14, PMKF17, SMSF19, YCL17]. causing [CLZ16]. Cayley [Dan18, HYS17, SY16, SY17c]. CBPS [GX18]. CBPS-based [GX18]. CCC [BK19, GFM17]. CDF’s [SKG18]. CDO [CR14]. CDS [GDW18]. Ceded [LLSM14]. Cell [Far12, KG14a, MGM13, SGKB17]. cell-mean [SGKB17]. Cells [BCD12, MP10a]. cellular [LDS18]. Censored [ADP10, ABI12, BS12c, BS12b, Bas14, CMCCO10, CB10, CB12, Coo12, Cra14, EDAAS12, FR10, FMS15, GT15, GFA10, HN14, Hua11, JFA12, KPS10, Kha12, Kha13, Lem14, Lop11, MA14, NH15a, Par15a, RAB15, SG10c, She12a, sS12b, She14b, She15a, She15b, SK15a, Sug13, TTZ14a, WY12, WWL13, WWT15, YXC11, YCL14, YSHL11, ZSY13, ZW09, ZS10, AN17d, AAK17, BN17, BP17, BMLY19, BG16a, CSC17, CLZ18, CT16, Che19b, FGP16, FZZ16, FZS19, FC18b, FA19, FST16, FGL18, GN17, GFL19, Har17, He17, KCY^{+}19, LL17d, LS17c, LY17, LM18b, MAA17, NB16, NC18c, PHY19, PMW17, Par16, Qom17, RSV17, Rez17, SS16a, Sha17, She16, sSW18, SSR19, TZW16, TR16b, WTZS15, Wan16b, Wan16c, Wan16d, WH17a, WP18, ZZ18a, ZJZ17, ZL16f, ZFD17]. Censoring [Abd15, BHLK15, Bas14, BGT10, CW15a, CW15b, CB11, CT14, DP10a, DP13a, Den13, GK11, HS15b, JMEM14, JY14, LK14, LJ13, Pao10, PGF09, PGF11, RD10, RB10, SS12a, SBK13, She10a, Wan12, WK13, ZS15, AA17a, AHAH17, BSL^{+}19, BP18a, BIJ17, CSL17, CBSN16, CK19, GP16b, GC19, JJ19, KTR18, MK19, SK18a, Sha19, SR17, STRD17, WL17a, WSS17, XZ16, XMP17]. Censorship [DPP15b, HS15d, WZY14, YLS10, GMS18, SB16a]. Census [GG13, GBC12, Jor13, CM16b]. Centered [DO13, LLV10, Xie16, YYS19]. Central [CZCM19, GLS12, IO13, JFA12, Li13b, MZ14, NR12, OCIS12, PP10, UC15, ZC14, Bar17a, hCXxShY18, DeR19, Ko16c, LPH16, LMT17, MGN16, MX19, OPK17, SN16, SNS19, Wu17c, Wu17d, Zan17, Zha16b, ZLZ19]. Centrality [Iez12]. Centrally [CMCCO10]. Centroid [SF12a]. Certain [Aka10, HV11, WLL13, DZL18]. certainty [LO17]. Chain [JC12, LD14, LDY15, ÖD10, Pap13, SS16i, XYH12, Bal18a, DPP17, HHC^{+}17, HL18b, KCL^{+}16, PAvS^{+}19, PC16b, SKK18, SFSH16, SG17c, SF16, SPY18, UB17, ZCC16]. Chain-type [SS16i]. Chained [Gov11]. Chains [Hun14, Sen14, WYL12, YWS14, BKKP19, Dan18, Geo17, HZ17a, HYS17, HYS18, LW18b, Pen17, SF16, SYW16, SY16, SJY17, SY17b, SY17c, Sze17, TY17, XHY16, YYS^{+}18, ZZ17d]. challenge [RBK^{+}19]. Chance [AKK09, AAK16, KKKÖ17]. Change [AVCL13, CTQ11, CJ17, DMT^{+}15, GBC12, ML13, Nos14, PJ10, PZ15, PS13, QTD10, SGPA12, Sur12, WZY14, WK13, AHZ17, CN17a, CKT18, Ciu17, DPZ17, Dvo17, FPM19, Far17b, GLY17, GMH16, He17, Huh16, JSC17, KL17a, LW17b, LCA17, MATC18, QL19, TZW16, VFP17b, XHY16, YAA^{+}17, ZBTL17, ZT18, ZG17b, ZFZ17]. Change-Over [SGPA12]. Change-Point [Nos14, Sur12, CKT18, Ciu17, Far17b, He17, JSC17, TZW16, ZFZ17]. change-points [ZBTL17]. changed [BVJV17, GWK19, SJ19]. Changepoint [Men13b, RK16]. Changes [AR14, DP13b, HTN15, QTD15, WILL13, XGZ11, BKR19, CS16c, CN17a, LTCQ17, PAvS^{+}19, QL18, Wu19]. changing [JZZZ18]. Character [Arn11, KK11b, KZ19, SSKP18, SG19]. characteristic [BT18a, MA18, PMFK18, Sel17b, SSS17c, Van16, YL16]. Characteristics [AS13, MJ15, SD14, SS14a, Arn19, CM19, CH18, Gup16, HME17, KLJP16, LPHC18, Shu18, SSS17c, TS17c, TS17e]. Characterization [AGAM17, AB11b, DJ10, FZ10b, Jaw14, Jeo12, Kay16, KRB13, Kum17, LWL13, MR16a, RR13, Tak10b, YMF15, YH11, ZFL13, AG18, NA19]. Characterizations [AH13, Bar10, CZ19b, Hom17, HS12a, IS16, IS17, KRB12, NSL14, NP17a, RR09, Yeh12, ZL16a, AH17a, BJ18, Jas19, UU17]. characterize [KG17]. Characterizing [MB15, Peg12, TV11, Tia12]. Charlier [OCIS12, SSON11]. Chart [ARD14, AMR14, BP15a, BTK10, CC11, CS12, FNDM12, Hua14a, Hua15, KS10b, KS14b, KTW10, LJ10, MC14, MZ10a, ML13, Par13a, SL13, SHH13, TKW12, AO18a, AH18a, AAJ18, COK17, CSM16, CKLC17, DHKM17, FCSH16, Haq18, HUK19, HK19b, HCZT19, JBA18, KNHM17, LK17a, LK17c, LK17b, LK17d, LK19a, LC19, PS17b, RSA19, RFM18, SKK18, Sag15, SFSH16, SHANA18, SA16, SG17d, WWZC18, WH16b, WZX17, YKL^{+}17, ZLT17, ZWZwZ17, ZG19b, vZvdM17]. Charts [AGR15, Bi15, CC15a, CY10, CR13, HHW14, HC10, HCS10, KHC14, LCT12, Lee13, LSSS14, Mah12a, MCY11, MN13a, Men10a, Men10b, Men13b, Men13a, NGT13, NAD15, WLY14a, YWLY11, ZARH14, AR17a, AP17, CC16b, Cox19, FPM19, GFM17, KSF17, KM19a, KCL^{+}16, KOFA18, Lu16c, Lu18, Mah17, Nen18, PKM17, PMF17, PMKF17, PM18, PMFK18, Pou17, RMM17, RC16, Rav19, SMSF19, WCK16, ZSJ16]. Chebyshev [Aga17a, Bud16, Cho10, Nav16, Rao17a]. Checking [Gar14, Iqb13, Bic18, DZL18, Ima18]. Chemical [CG12, KS12a]. Chemostat [JLV13]. Chen [BSL^{+}19, CZ15a]. Chernoff [Kay18, Kun17a, WZL10]. Chervonenkis [Gey18]. Chi [BLN13, GSA14, HJ14, SHH13, Voi15, RFM18, VPMV16]. Chi-process [HJ14]. Chi-Square [SHH13, RFM18]. Chi-Squared [BLN13, GSA14, Voi15, VPMV16]. chickpea [AIIK17]. China [Wei14, YC18]. Chisquared [FFN^{+}19]. Choice [BSB12, CT13, HKV12, Jaf13, KZW09, NC14, PS12, Rab13, DM18, LR19, PGCT19, SLZ18]. Choosing [Atk12]. Choquet [Aga17a, Wan15]. Cicer [AIIK17]. cigarette [TE18b]. Circle [Lin14a]. Circulant [LR10c, MGP14]. Circular [AT09, CC15b, FM12, FM14, ITG09, KS15b, QZY11, TJ13, Yam12, YWS14, Abu18, ASR^{+}19, BAT^{+}18, HH16, HYS17, KS16b, MPG16, RAST18, RRRA18, SYW16, YX18]. Circular-Linear [KS15b]. Circular-Linear/Linear-Circular [KS15b]. Cited [Koz12]. Claim [BS12a, EIJK13, NB15, PHW11, SW14b, BNB18, Ery18, Li17c, LYG17, LWG18b]. Claims [GW15b, XGZ15, BSW15, CY17a, DLH^{+}18, FS17, FL17, FL18b, FL19, GWL16, GCW19, GWZ17, GWP19, Li17b, LG16b, LSL19, WQW16, hXZ17, YS16]. Class [AK13, ASNT15, ABC12, BBG14, BCH12, BPSS12, CC10a, DM11, GDCOS13, Gue10, Hui10, Kak12, Kim11, MGAf14, MSN14, Neg09, Pet10, RP15, SASMF13, Sco11, SP11a, SS13b, ŠS12c, Šir14, Šir15, SS14b, SS15b, WWW13, WZT15, YA13, YUSC14a, Ala17, Are16, CK17c, DK17a, DK17b, DTWW17, Ery17, EEE16, Gau17, GA18, Gue13, HHS18, Has19, HHC16, KG18a, KN18a, LTV17, MKSS17, PP18, PT19, RGBG19, RMNA17, RCdCN16, RN17, ST16a, SK19a, SJT17, SGA19, SYW16, SJY17, SP17a, SPS17, SY17d, Ten18, WHV16, Wu16c, WZ19, Yad16]. Classes [EE14, IRAB13, KJ17, LLV10, SS13e, SS15c, TE11, FL17, JCN17, Li18, SKS17]. Classical [Aca15, BL15, GV11, KHM12, Koz12, VS19, LTZ19, Par17b]. Classification [CL10a, CD12, FU10, jK15, LR10a, LR14, LR10c, RL12, SF12a, SU11, SA14, VRC12, AXLJ17, BMQ16, FEP18, XM17a, YJSL19]. Classifier [Ell11]. Classifiers [Gue10, MC18, PR17]. Clatworthy [SMN18]. Clear [WYZ15, YWZ18, ZZ19e]. clearing [BL19]. Clinical [Ani11, CLC^{+}14, LMCH12, WTB11, Ko17a, MLD17, MR17, Wu16a, Yad16]. Clipped [Dos11]. Clock [PJL13]. Closed [CCH^{+}10, LdM17, SM18, WSW16, Moh19b, SS18]. Closed-Form [CCH^{+}10, LdM17, WSW16]. Closeness [AB14, BDKM12, EWRZ14, MA14, RL09, AMP17, Sab19]. Closure [LMH17, WGW19]. CLT [Gra17]. clumps [SU17a]. Cluster [CGL13, DC10a, Koz12, SA15b, VBATK11, WST10, Bak17, CS16b, Haq17, He18, MSB19a, MV16b, QKCH19, VN18]. cluster-head [MV16b]. cluster-specific [He18]. Clustered [Man11a, CS16b, DL19, Fan18, LH17a, QKCH19]. Clustering [AH14, Cey11, DBC16, DG12a, GD11, Iez12, Kaz11, MSB19b, Yao12, ZLDT14, Con16, KC16, MBV17, MV16a]. Clusters [DI10, Jor13]. CLWD [GCW19]. Co [JLL13]. Co-Volatility [JLL13]. Coefficient [Aca15, Aoy10, CC10c, CYL13, DSO12, FH13a, FHZ12, GBD10, Gro11, Hu11a, Hu14a, Hua12, KK11b, KG14b, Men10a, RB11, SSS14b, Tak10b, Uya17, WW15a, WH15a, Wei12, Xu13, YA13, YX15, ZY10, ZWZ11, ZX10, ZWPZ15, ZZLL15, AV18, ASG17, Bib16, CH19a, FTT16, HH19, Hu14b, HLY16, HJT19, JQZ18, JHM19, KM18, KS18b, Kvå18, LLHZ16, LZS17, LLWK18, LL18a, Lu13b, LLF19, MJH19, Mat18, MKY16, PY16a, RTT17, SG17b, TT17, VN18, WX19, WH16c, WXZ18, XQS19, XZ19, Xu16, XN19, XFWC19, YLG16a, YLG16b, YSC17a, ZWL16, ZT17, ZL16d, ZWPB17, ZXF17, ZSS18, ZP17, ZL16g, cZzXgL18, ZY19, vZvdM17]. Coefficients [Afs13, AJA14, BS15, CS14a, Gho15, GR11, Jaf15, LS09, LWR11, Sen14, SGM11, LARS19, LWHB16, Nag17, mNK17, Sun17b, Yad16, ZJLO17, ZSWA19]. Coherent [AA14, Bic12, ZM13, Che16b, KS19, MSR17, PBA17, PBA18, Tav16]. Cohort [WR10, CY17b, CSY19, MZ17]. cohorts [WER^{+}19]. Cointegration [DPS12, KA17, OL19a]. Cold [BC12a, WZT15, Ery16a, FK16, LCWG16, MSR17, VM17, Wan16e, YYLM18, ZW17b]. Collaborative [CGC17]. collapsibility [Bur17]. collect [YSS19]. Collective [SW14b]. Collinearity [IT12, Spa19, WF17]. column [DJVV17, GPM18, VJV14, VV17, Wan17d]. Columns [KLS12]. COM [CO16b, JKS18, WP18]. COM-Poisson [WP18]. combat [WCF18]. Combination [CC10d, CDZ14, LR10b, Mar16, Sal15, SLL14, Che16a, IF18, ZW19b]. Combination-Based [CDZ14, Sal15]. Combinations [AE13, GK13, KRT13, WL13, KJ18, SMK17]. Combinatorial [Ery16a, Ind11, NR12, Dut17, Fro17a, SN16, SNS19, VZ19]. Combined [EMI14, LCT12, LK17a, MCY11, Men13a, MKY16, NAD15, OQC14, OQC17, Wu19, YGWQ17]. Combining [CL14, Gue13, KMC10, New11b, VV17]. Comment [KD17a]. Comments [Kaç19, Li19, LF17, MC14, SG17d]. Common [Bi15, CHP14, EIJK13, KL11, KT11, PT15, ST10, AB19, CX17, Gun17, JT19, LW17c, MH16b, Meh17, NRN^{+}17, PK19, TKJ17, Wu19, vZvdM17]. Commonly [Hol15]. Communication [Gup14, KJS19]. communications [MSD19, Ano23, AE23a, AE23b, AE23c, AE23d, MBS17, SG17d, Mül12]. community [Bar17c]. commutative [MT19]. comorbidity [WER^{+}19]. Compactly [KOP14, PTV18]. comparative [VPMV16]. Comparing [AJA14, BB12, EE15b, GGS13, HMP12, JFS14, KS15a, LWWH14, LS17b, Tun15, WW15c, WH10, Hay16, JZ17a, KG18a, LCA17, SS18, SE17, VV17]. Comparison [AMP17, AAR11, AH14, BDKM12, BNE15, BAR16, Bie16, CB12, Dji10, FJS^{+}18, Gru12, HN14, HS15c, HD15, JA10, JAA11, Kak10, Kat11b, KG14b, LX11, LH11b, Li12b, Lin16, Qum12, Raj18, RSA13, SN17, SAD16, Sha10b, SBK13, ST13, SS10a, SA10, SSD19, TE11, UC15, ÜY13, UM09, Xu10, XF12, YLX10, YF17, ZVS14, ZTSL15, ZLWZ10, AHKO16, AAL17, AEA17, BMR17, CSM16, CK19, FM16, FRC16, Ima18, KH19, LPDD18, LCZA17b, SB17, Sen17a, Sen17b, SW19b, TE18b, WXZ16]. Comparisons [BC15, CKD14, CMCSC12, EE11, FWH^{+}15, GW15a, Gro11, HNB19, HW14, sK16c, KT15, Kur10, LS15a, MWW14, OAOQ16, Shi10, SD12, XY11a, ZS12, ZCCW18, Bic17, CVL19, FLB16, Hot16, Ima17, KA18, LL19a, LLZ17, Meh17, MN18, RSA19, Ros19, Say17, SM18, Sin16, Sou17]. Compatibility [DS14, GN16]. compatible [Sou17]. Competing [Ko14b, RON12, RM14, SS12d, WY12, XT11, AN17d, BSL^{+}19, Bel17, CSL17, CK17b, FZZ16, LXL17, LS16a, LS17c, SS16a, SDS17, SX18, SLSE19, SSD19, Wan16b, WXS18, WP18, WSS17, ZZ18a]. competition [GDDC19, Len17]. Competitive [Atk12]. Complementary [LH13]. Complete [AA15b, CKS17, DP17, DWM^{+}17, DTWW17, FWW19, GCR12, Guo14, HL15, Ko14e, Ko18, Liu11a, LSYL17, MWW18, SAB18, SWL14, SWLW14, SXW17, Sun12, TWZ16, WHVH13, WLHW14, WW19, WV16, WP19, XTL^{+}17, XLL15b, Yan19, YQC17, Yu12, YW18, pZ13, ZMMX17, BZH19, CZM17, HZZ19, LH17b, WHV16, XDWC18]. Complete-Case [Liu11a]. Completely [MH15]. Complex [ASN12, AZN15, Cha14, Duf10, Eid12, HTN15, Kon10, Mon12, SG11, FN18, LL16, NAAC16]. Complexity [Gue10, KDG^{+}18, Sha16]. complexity-type [KDG^{+}18]. compliance [GZ16, GZ17]. Complicated [Lu15, Lu16b, SA17a]. Component [Are14, BYDM15, hFmH14, JB15, OFS13, SE13, ARA16, Ali17, BPV19, CNL19, Li18, MAA17, MSR17, PG16, Wan16e, Wu17b, YH16, YLG16b]. Components [ASAK10, BHB15, Bur15, CL10a, Che11, CW13, CZ15c, DZ12, Eid12, EZ14, Ery12, GE11, Ko14d, KSBM14b, KSBM14a, NKS15, Par13b, Raq10, RSA13, SGR^{+}12, SS10a, Tsa10, WY13, YLL19, Yil12, ZL15, BNB17, BNB19, BCL16, Che17a, CFK17, FLB16, HNB19, Ko16b, KJ18, LL19a, Mao18, NI16, PSF16, SHB17, ST16c, YL18d, Zha17e, ZZ19e]. Composite [CTT15, FTT16, Fed13, LLP10, LL11, PP10, UC15, WXPL13, AD19, JQZ18, JHM19, LD18a, LY17, OPK17, PT19, TLT17, YWLW18, ZQZY18]. Compositional [FH12, Gal15, ZSD^{+}15, DFH16]. Compound [ZYN11, AAO17, BP17, CLW17, GH19, GWZ17, GWP19, HS19, IS16, Kou16, Mat18, RCdCN16, TMZ^{+}17, Yan17a, ZL16a]. Comprehensive [Cui12]. compressed [YGWQ17]. Compressive [MN15]. comprising [BNB17, BNB19]. compromise [SH17b, YWZ18]. Computation [Gho15, DZH^{+}19]. Computational [JX19, SMS14, MGG^{+}17]. Computations [CR14, SJJ19]. computed [Nag17]. Computer [CS11, SGPA12, SWW^{+}12]. Computer-Generated [SGPA12]. Computing [Ham12, KC16, RS16, YGWQ17]. Concave [LLSM14]. Concentration [Rat13, AMP18]. Concept [PK12, Zha15a, AAE^{+}19, Bow17, ZBA17]. Concepts [HT13, Sre18, SZL14, ZC11, IAB16]. concerning [DAS16b, Jon18]. concerns [HW18b]. Concomitant [PT17, CLZ19]. Concomitants [AAR11, KO10a, RJN16, TB12, TV11, VT15, Ery16b, JPAB17, VT17]. Concordance [TAY11]. Concordant [Lu11]. Condition [KvdVvZ13, Lia10, RZ14, Yar14, Gra17, RD16, Yan16]. Conditional [AT19, AH13, BCS10, BBR17, CCL14, CW19c, Fau12, GGFL11, HB13, He18, Hod14, HCZT19, HL13b, KHM12, KSBM14b, LZY^{+}17, LMR13, LWL13, Lia10, LLY14, MN10, NSL14, PGF11, PBA17, Pen17, Rah12, SY11, SH11, She12a, ST13, TL16, WLF11, Wan15, WT15, Wan16e, Wu11, XYL14, XXMZ18, YL15, Zha11, ZWP17, ZZSP14, BN17, BH17b, BLT16, CZ19c, DLS17, FC18a, GN16, GDPR17, HA18, Hen17, HL17, KOA16, KNH17, LLM16, Liu19, PBA18, SK16a, SB16a, TS17f, WHYW18, WBGS19, XZA18, YWL17, YL16, ZBA17, ZY17b]. Conditionally [(Ba14, Pie11, RG19, SCMR15, YL15, KLMD16, KS17, Rez17, Wan17f, ZC17a]. Conditioning [WN16a, AGA17b]. Conditions [Gra12, Guo14, Hu13, HYK15b, KP12b, LS09, Lia11, Nge11, Tan10a, Tan10b, WLF11, WTZ15, BBHN18, BMR17, BH17b, DP17, HZZ19, KHY17, Kyr17, Nie17]. Condom [TWH10]. Conference [Sty10, MSD19]. Confidence [AKL15, BHLK15, BKL19, BL15, Bic12, Bic17, BSY12, Bur15, DC14, DZ12, EE11, FLK12, HKLW12, HNV13, KG14a, KL10, Ker17, KV10, KZ15, KX18, LQ13, Liu11c, LVVN11, Moh19b, New11a, New11b, Par13b, PY16a, PGF09, PR11, PJL13, QQL14, Rub10, SP11b, SLSY14, WCL11, WXZ13, WQ14, WH14, WWH15b, WH10, XZD10, XM17b, YXG11, ZB12a, Zha14, ABK17, AK19, BAB^{+}18, BG19b, Che16b, Fro17b, GP16a, GW19, Hua17a, KWM17, KPZ16, Kus18, Lin16, LH17c, ML19, PA18, PG16, PY16b, Raj18, SH17a, Son16, Sou17, Tar17, YX17, YCH17, ZLD16]. Configuration [HK11]. confirmatory [YM17]. conflicting [CW19a]. Conforming [KS14b, NN13, FY17a, UK16, YF17]. Confounded [Chi12a, FM10a, DJVV17]. Confounder [WZ15]. Confounding [Chi09, Chi11, ZYLZ17, ZS17a, ZZ18e]. Conjecture [GGL14, NM17]. conjugacy [KM16b]. conjugate [GDCO16, SUS19, ZRL19a]. Connection [VB13, Had17, Yan16]. Connections [Dem14, OQL12, LD17, OQ17]. Connectome [PVB13]. Consecutive [JCY10, Ste11, MR17]. Considering [sK11, Lu11, MBN14, ANJ17, BU17, FY17a, Xio16]. Consistencies [Sug13]. Consistency [AD17a, AD16, CM16d, DS16b, EMH16, HK13, LN13a, LMR15, Tak10a, WLC11, WY12, Wan13b, WS14, WJ11, XYL14, Xin15, YC12, YC11b, ALR19, Bin17, BCL16, Che17b, Che19b, Das16a, HJT19, Rej16, SB16a, WH17b, XW16, YYY17, Yan17b]. Consistent [BNS11, FP12, KKS19, NBK14, WW13c, WX17, WF17, ZP17, HW18b, NB16, Nas18, VPMV16]. Consistently [LSX14, DW17, WGW19]. consisting [FY17b, NO18, XN19]. constancy [AV18]. constansistent [ZL16d]. Constant [Ell11, hFmH14, HN14, Hun14, JMEM14, Raï15, Sur12, ZS15, BSW15, DW19, Kha19a, Kha19b, Li17b, XX18]. Constant-stress [HN14, ZS15, DW19]. constant-sum [Kha19b]. Constants [TN13]. Constrained [AKL15, Fly10, UC14, EE16, Fon17, LSL16, LL19c, MP17, Pat16]. Constraint [LH14b, MS10b, SCC11, SKG18]. Constraints [AJN15, CY14, FZ10a, FH12, HX11, Nos14, PC10a, SN12, ZFL13, Che16c, LL16, WZ16, YY16]. Constructing [AW10, DDÚF11, KXX11, Sha10a, ZK15a, LGLX18, Ram17, ZS17a, ZZ18e]. Construction [AA11a, AGS18, AGA17b, BAT15, DDGM13, FF13, GFR12, HS19, HKLW12, IBA18, KT18, KS19, LIPMF13, LLV10, LL16, MGP11, MGP14, MKD11, Mud16, QYLL17, SO14a, Voi10, WLY14a, ZYLZ17, ZLYZ17, fZWD19, ZL16e, FPPRG16, KLJP16]. Constructions [DFS10, JSWL17, Oll17, RAST18]. Consumer [GB15, LIPMF13]. consumption [LG16a, TE18b, WWWX19, ZY17a]. contagion [PF19]. Containing [WYZ15, GAS17, JD16, ZZ19e]. Contaminated [GVGP11, GR15, WZK17]. Context [BGG11, Sim14b, SS12e, Bos18, NS16b]. contextualizing [Bur17]. Contingency [BSY12, CC14, DBC14, EPGTH15, FZ10a, GN10, THV16, YSR11]. contingent [Hon17, WQW16]. Continuity [ASAK10, Hu17]. Continuous [AMR14, AH13, BCS10, Bia15, CCH^{+}10, Cho10, DDGM13, DT12, EF11, JRM10b, PMC10, RJ15, SB11, TW13, Vas14a, WQ14, XY11b, YSHL11, ZK15b, Bak17, BH17a, Bos18, BD17a, CC19, CZ19b, Cuc16, EM18, GWL16, Gen17, LQ17a, LH17a, NRN^{+}17, PAvS^{+}19, SG16b, TS17f, VFP17b, WWWX19, WG16a, XHH17, ZF18]. Continuous-Scale [WQ14]. continuous-time [CZ19b, PAvS^{+}19, WWWX19]. Continuum [WTZ15, Bic17]. Contoured [ASNT15, Ric13, ANA17, Rez17]. Contracted [HW15]. Contractive [BCG11, PYL17]. Contracts [DW13]. Contrast [RMC11, FKR16]. Contrasting [YC14]. Contribution [PM16a, SLZ18, ZMCM16]. Contributions [AK10, AA17b]. Control [ARD14, AGR15, BP15a, BTK10, BQDH12, Bi15, Bis12, CC15a, CY10, CC11, CS12, FNDM12, GGS13, GW15a, HHW14, HC12, Hua14a, Hua15, HCS10, JFS14, KS10b, KS14b, KMM12, LJ10, LCT12, Lee13, LSSS14, Mah12a, MZ10a, MN13a, Men10a, Men13b, ML13, NGT13, NAD15, SHH13, SZ13, WDZ11, WLY14a, YXG11, YWLY11, ZARH14, ZS12, AO18a, AR17a, AH18a, AAJ18, BMQ16, CC16b, CKLC17, Cox19, DHKM17, DCZL17, FPM19, Haq18, HUK19, HK19b, He18, Hot16, JBA18, KSF17, KM19a, KOFA18, Ko16b, Kyr16, KD16c, LK17c, LK17b, LC19, Lu16c, Mah17, Nen18, PKM17, PMF17, PMKF17, PM18, PMFK18, PS17b, Pou17, RMM17, RC16, RFM18, Rav19, Ros19, SMSF19, SKK18, SFSH16, SHANA18, SA16, TMZ^{+}17, UK16, VV17, WH16b, WZX17, YWX17, ZLT17, ZSJ16, ZWZwZ17, ZG19b, vZvdM17]. control-limit [Kyr16]. Controlled [Chi12a, Muk10, GICRR17, JX19, Pen17]. Controller [Par13a]. Controlling [GW15a, MWW14, WX15, QK17]. Convariances [PV14]. Convergence [AA15b, ABC12, BR11, Din15, DMT^{+}15, ES11b, EdSMR14, FM13, Fu11b, Guo14, HH13, HF19, HL15, Ko14e, KOP13, KOP14, LZPJ11, Mel13, Nad10, QC16, RB11, SWL14, SWLW14, SWZ16, Ste15, Sun11, Sun12, TL15b, WHVH13, WLHW14, WG16b, Xin10, dXcY12, XY14, XM14, Yan19, pZ13, ZL17, AD17b, BR17, BZH19, CSS16, CLLW16, CZM17, CKS17, CL17b, CZW16, DP17, DWM^{+}17, DTWW17, DS17e, FWW19, HV17, HZZ19, Ike17, Ko17b, Ko18, LH17b, LP18, LSYL17, LO17, MWW18, Men19, PC16b, SAB18, SZV16, SS17b, SXW17, SXZ17, STD18, SY17d, Sym19, TWZ16, WHV16, zWDD19, WW19, WV16, WP19, XTL^{+}17, XDWC18, Xie16, Yan18, YZG17, YJS19, YQC17, YW18, ZMMX17]. convergency [NE17]. Convergent [Sha10a, UC14]. convertible [SK16d]. Convex [AA11b, GK13, Mat11, SLL14, AAE^{+}19, Had17, LWH19, Rej16, TS17f]. Convolution [CXS11, CF13, Imo15, PW16, WY11, CG17, KC16, NM19, XZD19]. convolution-defined [NM19]. Convolution-Equivalent [WY11]. Convolution-invariant [PW16]. Convolutions [BCH15, ZL14a, LMH17]. Conway [INOC17]. Cooperation [GGFL11]. coordinate [SSR19]. Coordinates [GKW14]. coordinatewise [Ko18]. Coordination [TS12]. cope [SKM16b]. Copula [(Ba14, BK09, BEZ11, BZ13, BG11d, FNDM12, Fau12, GGL14, JCY10, JC12, LD14, LSC13, Mey13, PK11, YYLM18, Ber16b, Coo19, Den17, FGGL18, HTW17, JSWL17, LGLX18, Rez17, SSD18, Sre18, XZ19, XMP17, ZYH18]. Copula-Based [FNDM12, YYLM18, FGGL18, Sre18]. Copulas [DDÚF11, DFS10, DJ10, DS14, IAA15, Jaw14, KP13, MKD11, PL17, SASMF13, SLL14, Ber18, Key16, MBV17, PV16, SS19, ST16a, Wys19]. Cordeiro [AE23b, AE23c, MBS17]. corneal [Abu18]. Cornish [Are16]. Correct [FM11, YZFM17]. Corrected [CHH15, DDP18, KYS13, BG19b, EAE18, FPPRG16, GFG16, JZTN17, Wan17a]. corrected-loss [JZTN17]. Correcting [Çel17, ÇE17a]. Correction [AR17c, Ano19a, Ano23, AE23a, AE23b, AE23c, AE23d, CXS11, CT14, Gro10, SHHH09, Kvå17]. Corrections [LF10, Roo18]. Correctness [SNN12]. Correlated [BB14a, CG14a, DP14, ER14, KL10, LX10b, LY11a, LY11b, MS14c, NGT13, OJR15, RBB13, RM15, SG14, SGPA12, Wit14, Zad15, AP17, dCA19, BQC17, BO17, DeR19, FL17, HZS17, Hay17, JKS18, LMC18, MHA19, MKY16, NS19, NC19a, PBC17, RSG17, SA19a, TSG19, TK17a, TH17, Xia17b, YZFM17, Zha17e, ZY17b]. Correlation [CC10c, GR11, GHY11, Hu11b, Kim12, KCS12, LWHW11, Mat18, MBN14, OFS13, RL12, SC10, Sal15, SMM10, SP15, SO14b, TEF12, VML^{+}10, VMVD14, YA13, YSR11, AO18c, Bar19, CS16b, KS18b, LLHZ16, LTGK17a, LTGK17b, Pro18, Sun17a, WH16c, XZ19, Xu17b, Yad16, ZJLO17]. Correlations [AD10, Dem14, Bar17c, JD16]. Correspondence [BDS11, BL15, DBC14, DC14, LSD10, D’A17a, DAC17, Sil18]. corresponding [SUS19]. Corrigenda [Ano11a]. Corrigendum [Ano10a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano13a, Ano13b, Ano14, Ano15a, Ano19b, Ano19c, Ano19d, Ano19e, GK19a, MBS17]. cosine [ZG17a]. Cost [CC10b, EF11, Hil15, IP12, JHP10, SCSMN11, CSC13, Ko17a, PM18, SH17b]. Costa [SG17d]. Costing [MGF14]. Costs [GB15, CLZ18]. Count [BG11c, Den13, KM14a, KS14b, OBPL15, ÖD10, YHA10a, YW15, ZVS14, BW17, BKGD17, BB19b, CM16b, Gen19, LYR^{+}18, LCZA17b, LCA17, NC19a, Quo17, TSG19, TE18b, YF17]. Countable [YWS14, SJY17, YYW17]. Counter [SAB15]. Counter-Type [SAB15]. Counterparty [DMS14, LD14]. Counties [Con12]. Counting [Nas12, Ili16, NRI17]. countries [PLC17]. Counts [Cox15, WW15b, Con16, FGL18]. Coupling [ES11b]. couplings [KNC18]. Course [HL11, Kim15]. covariables [ZL17]. Covariance [AV15a, AMZ15, BBC13, CK10, CBB06, DMP11, GLZW12, HP10, KK14, Kon10, LR14, Liu17b, LR10c, MBS17, MZ11, MS11c, RL11b, RVP15, Sch12, SP13, SAvR12, SAC13, SSH10, Tsu14, WDM14, Yam12, YDC15, Ahm17, Bar17c, Bud16, DLS17, DC17, FZ17b, FZ18, FI17, FPCP16, HZS17, HJC17, HSSP16, KM19b, KT18, Liu19, Mal17, NKG16, OU16, PR17, Rao17a, SÁHLvR17, TME19, WYY16, YX18, ZKS18, ZBHW18]. covariances [GS17, WDLZ16]. Covariate [Ala14, BC12d, DE17a, DLP15, DDM10, DDM14, LLC10, Liu11c, ML12, OH19, SSM12, SK10b, XLL15b, ZSS15, DS17c, DS17d, FZ18, LAW^{+}19, LLF19, Nak19, WY16]. Covariate-Adjusted [DLP15, LLC10, ML12, LLF19]. Covariate-balancing [OH19]. covariate-dependent [WY16]. Covariates [CG11a, CHH15, DPB13, Den13, LY11b, LWH12, Pic15, TWH10, WJH15, YWO^{+}15, ZZ11, BIJ17, CLZ16, FDL17, GGO17, GHL17, KRMG19, LY16a, LWZD17, LXT17, OW19, OD19, PLC17, Qiu17, Roy16, SLF18, Sun17c, YGL16, YCZ17, YD19b, ZY19]. Covariation [KG14b]. Coverage [WWH15b, LL17b, YLS17]. Cox [Abd15, BG16a, CCZ15, CSC17, CS19c, FNM19, FZS19, FWZ^{+}17b, GHL17, GE11, HA14, HSHW14, IT12, LH17c, sSW18, XLZ^{+}12]. Cox-Aalen [sSW18]. Coxian [MGF14, ZMCM16]. Cramér [Gri17, YDL14, ES11b, HK13, Kal19, Kvå18, Mar11, SS11b]. Cramér-type [Gri17]. Cramér-von [Mar11]. Credibility [EIJK13, NAN17b, SLL14, WW11, WWW13, WYMZ17]. Credible [Han12a].

Credit [BCCH14, CG14b, DMS14, FG17, HF14, LD14, BKKP19, BH17c, CCCT16, LMdO18, Zen17a, ZZ19a]. credit-rating [BKKP19]. Credits [GRV15]. cricket [Das17a]. Criminal [NRS^{+}14]. Crisis [Cui12, KYW16]. Criteria [APM10, Lem14, NRS^{+}14, OQL12, ST13, AL16, DGMC17, EMH16, EQ17, HS19, KDG^{+}18, LPDD19, Miy17, Oga19, OQ17, Sha16, WQ18, ZZJ19]. Criterion [DZX13, FH13b, GHY11, HS12b, KS12b, OFS13, RL09, Sha10b, Tra11, ÜY13, Wei14, YLX10, EIM18, KT18, KH18, WLS18, ZLYZ17]. Criterion-Robust [FH13b]. Critical [MS12b, RCdC14]. Criticism [Ras17]. Cronbach [LARS19]. Cross [BDM13, MD17b, NRS^{+}14, Slj17, TWH10, JKS18, Mao18, Oga19, PBC17, ST16b, Ten17, ZKS18, ZL16c]. cross-correlated [JKS18, PBC17]. Cross-entropy [Slj17]. cross-equation [ZKS18]. Cross-Over [TWH10]. cross-section [ZL16c]. cross-sectional [Mao18]. Cross-validation [MD17b, NRS^{+}14, Oga19, Ten17]. crosses [VV17]. Crossing [SM09]. Crossings [LM13]. Crossover [BM15, CC10c, LKN15, LS14, Liu11a, Zho10, LS17a, MPG16, Wu16a]. CSTM [GK19a]. CTE [LLSM14]. Cuadras [GEA^{+}10]. CUB} [Ian12]. Cube [FH13b, NC14, Pic15]. cubic [AOÖ17]. Cuboidal [UC15]. Cum [TC14, LTS16, RM16, SSS16]. Cumulant [KCK10]. cumulants [Shu19]. Cumulative [Bar10, BR12, BDS11, CB14, CSC13, CC11, Che14, CW15b, DBC14, D’A17a, DM18, EE15b, HKNL15, JRM10b, KS14b, BK18, CC19, Che17b, Kum17, KN16b, KK17b, KK18, Min17, MB17b, MBF17, NP17a, PNK18, PE17, SK16a, SS17a, SSD19, YF17]. Cure [ABI12, BP15b, XLZ^{+}12, ZZ10, BKM18, CRTdC17, GZ17, POS^{+}19, She18b, WP18, YCL17]. currency [HZ17b]. Current [RPY^{+}13, She12b, ZTZ^{+}13, FDL17, SKM16b, SHM16, SSD19]. Curtailed [ZLD15]. curtailment [ZLD16]. curvature [Abu18]. Curve [DG11, DTE13, GFA10, Li15, CM16d, LX19, LWD^{+}18, LQ19, NS16b, SS16f, WNL18, XQ17a, Xu17b, SS17c]. Curved [Aka10, GCSB10]. Curves [AB11b, EZ14, Kaz11, LX11, RGM15, VV12, ZLWZ10, MBAJ19, Nie17, SJT17]. Cusp [Web12]. Customer [Fon14, BCB19]. customers [KS18a, SLOT17]. CUSUM [AR17a, AH18a, AAJ18, AR14, HHW14, JK19, RSA19, WH16b, WZX17, Wu19, XGZ11]. Cut [ZB12b, LPDD19]. cut-point [LPDD19]. cuts [Gey18]. CV [Gro11]. CVA [CR14]. CVSSIDS [NAD15]. cycle [ZMN17]. Cyclic [MPG16, TIA11]. cyclostationary [MHA19].

D [AE23c, HM15, MBS17, SK16e, CO10, PBY^{+}14, XM17b]. D-minimax [HM15]. D-trace [XM17b]. da-Silva [AE23b]. Dagum [DC13, TCM^{+}16]. daily [BBS16, El 19, TE18b]. Damage [Che14, CC19]. Danish [ANS17]. Dantzig [ZHLZ14]. Darling [Dvo17, Lie16]. Data [AT19, AK14a, AAG12a, AV15a, ADP10, AC11, ABI12, AH14, AT09, BCH12, BG11a, BS12c, BC12d, BC14, Bon14, BG11c, CMCCO10, CT13, CL10a, CS15a, CS15b, CY10, CXS11, CZ15c, CS14b, CB10, CB12, CMCSC12, CCSCME13, Coo12, DNL12, DS09, DZ14, DC10a, Dem14, Den13, sDsWcCjK14, DM10, DG12a, DP10b, DL11, Eid12, Eid14, EDAAS12, ER11, FR10, FJ11, FMS15, FV11, FH12, FDHH14, Fu12a, Gal15, GFA10, GD11, GHY11, GL15b, GS15b, Gui15, GBC12, Ham10, HN14, HD13a, HS15a, HS15c, HD15, HYK15a, HL11, HS15d, Hu11a, Hu13, Hua11, HQW15, Ian14, IP13, Jia12, JLL13, Ju15, KG14a, KM14a, KL10, KR14, KSDR15, KNK13, jK15, Kim15, Ko12, Ko14a, Ko14b, Ko14c, Ko15b, Ko15a, KN17a, LK14, LR14, LJ13, LS11, LSSS14, Li14b, LX15, LLY14, Liu11b]. Data [Liu11a, LY11b, LB11, Lu10, Lu11, MKS15, MSSU14, Man11a, MZ11, NB15, NA13, Par15a, QZY12, QQL14, QNZ14, RPY^{+}13, Rao10, RM15, RL11b, RL12, SS13a, SU11, SO14a, SG10c, She10b, She12a, sS12b, She12b, She13, She14b, She15a, She15b, She15c, SCdSP10, SSP^{+}12, SS10b, SS14a, SP15, SS12d, Sug13, SSH10, TK13, TWH10, TTZ14a, TX14, TT11, Tsa10, TJ13, VML^{+}10, VMVD14, VSS12, WR10, Wan10b, WH11, WP11, WY12, WCGZ12, WQ14, WG14, WTZS15, WL15, Wit14, Wu11, WS12b, Wyl13, zXyL13, XT11, XYH12, YHA10a, YXC11, YC14, YA10, YWLY11, YSHL11, Yu12, hYhW14, ZPZ10, ZZ11, ZTZ^{+}13, ZHLZ14, ZX10, ZZS13, ZVS14, ZSY13, ZW09, ZS10, ZZSP14, Ahm17, AD17a, AD17b, AXLJ17, AT17, Alv16, AHP19, ABTR19, ANS17, AN17d, BN17, BW17, BJ19, BP18a, BP17, BANC17, BB19a, BH17b, BIJ17]. data [BLT16, BKGD17, BB19b, BT16, BP18c, BCCR18, Bos18, BG17, BG16a, BJC16, BK17b, CSL17, CS16b, CSC17, CCN17, CHM16, CG16, CT16, Cuc16, DMM16, DE17a, DDP18, Den17, DSA16, DR17, DDS16, DL19, EMH16, EM18, FZ17a, FZZ16, FZS19, FEP18, FDL17, FC18b, FJS^{+}18, FST16, FP17b, GAS17, GZ16, GZ17, Gen19, GFL19, Gro18, GGtH^{+}19, GK19c, GW18, HB16, HB17a, HB17b, HL18a, Har17, HH19, HDS16, He17, HW18a, HJW19, Hen17, HL17, HSW17, HNB17, HSSP16, IK18, JSSO18, JCN17, JYZ19, KD16a, KHY18, KOFA18, Kia17, Ko16a, KCY^{+}19, KL17b, Kus18, KT17b, LW18a, LW17b, LM18a, LXL17, Li17d, LG18, LC17, LZS17, LL17d, LXT17, LS17c, LSCW17, Liu17a, LTGK17a, LTGK17b, LDL18, LL18a, LAW^{+}19, LM18b, LMdO18, Lu16c, LWZ17, ML19, Mao18, MBV17, MSD19, MT16, MMP17, NC18a]. data [Nag17, NB16, Nak17, NAAC16, NNvRS17, NVS19, NS19, NC19a, NC18c, OU16, PHY19, Par16, PR17, Phi18, PP17, Qom17, Quo17, RSV17, RBK^{+}19, RLSL19, RKMG19, RTT17, Roy16, SAD16, SS16a, SV17, Sha19, Sha18b, SSSM16, Sha18c, She16, sSW18, She18b, SLSE19, SS16e, Sla17, Spe17, SSD19, SSY18, Tak18, TZW16, TR16b, TK17b, TH17, TE18b, VN17, Wan16b, Wan16d, WZK17, WH17a, WXS18, WH18, Wu19, XQS19, XTS16, XXMZ18, XZA18, XT17, XFWC19, YSS18, YAA^{+}17, YJSL19, YWL17, YWLW18, YSS19, YWLZ17, YXZ17, YH19, ZCHR19, Zha17e, ZZ18a, ZWNS18, ZL16c, ZJZ17, ZG17b, ZMV^{+}17, ZY17b, ZFD17, ZL17, JPAB17]. Data-Based [AC11, CT13]. Data-Driven [Wyl13]. Datasets [AA13, BBB14, ABES18, TSG19]. Davidson [AAL17]. Davis [WH17a]. Daya [PM18]. DD [Gor17]. DDA [Gor17]. Deal [SKK10]. Death [Ano13b, GS11, MS10d, PDS09, CY16, KD16c]. Decades [HH15]. Decision [AAL17, Bic12, CHLP09, GGS13, LWWH14, Bic19, CCCT16, FEP18, KD16c]. Decision-making [AAL17]. Decisions [Fal14]. Decks [ASN12, JSSO18].

Decomposition [BDS11, DC14, DAC17, MCY10, Voi10, HOK17, PS17a, SK16b, YZFM17]. Deconvolution [TPTT14, Ben19, PTV18]. Decreasing [LM14, AE17c, GR18a, LMM18, MKSS17, PKM17]. Deductibles [HW14, LL17b]. Default [LD14, CCCT16, GDW18, LMdO18, WRZ19, XDW19]. defaultable [MWY18, XDW19]. Defectives [TD14]. Defence [NN11a]. deficits [TMZ^{+}17]. defined [LPDD19, NM19, SLZ18]. definite [DTT^{+}16]. Definiteness [DMP11]. Definition [DC13, MK13, FWZ^{+}17a, SY17b]. Deformed [Mur13]. Degenerate [DN14, FZ10b]. Degradation [RE10, She14a, cTtL14, XT12, BPS16b, DW19, GPP19, aYwSlDmS17]. degrading [PSF16]. Degree [HTN15, KMC10, YMF15, LQYZ17, LD18b, LQW19, ZQQY17]. Delay [DJ15, GS11, UZ19, ZW17b]. Delayed [SB11, XGZ15, ZW13, DLH^{+}18, GWL16, KC19a, KC19b, Wan17f, hXZ17, ZW17a, ZW17c, ZW18]. delaying [JX19]. Delays [PJL13, LDS18, SLOT17]. deletion [EP17, Zar19]. Delivered [SA10]. Delta [FM13, GG18]. Demand [SYP13]. Demographic [GBC12]. Demonstration [Sco11]. Dendrite [Koz12]. Denoising [NAMD14]. Dense [SLSY14]. Densities [AE17a, BC14, Ham10, JTR15, RP10, SS10b, BS17, MHA19]. Density [AE13, ABLR15, BYDM15, CS15b, CF13, Che13, FJ11, Fau12, HWM12, HS14, Jeo12, JFA12, LMR13, LX10a, Li14a, Li14b, Lia10, LCP^{+}14, LLY14, MSW13, Nad10, Pak14, Par12, Rao10, RR14, Ryc14a, VLD15, WH15a, Woj11, WHY18, AE17b, ARJF19, AM16, BN18a, BR17, BR19, BT18b, BD17b, CD17, Che19b, DS16b, DS17e, FZ17a, FPZ16, FST16, GTW17, GR18a, Iga16, IK18, Kam19, LA17, Li17a, LWZ18, MMP17, NE17, NS16a, Pak17, Pen17, Rao17b, SAD19, SCYR17, Sla17, Ten17, VN18, WZK17, XW16, XZA18, Yan17b, ZL18, ZL17, ZHZA18]. Departure [Kem11]. Dependence [AT11, ASB13, BS12a, Dok12, Dom11, DFS10, EIJK13, Fer14, FF13, FMPF16, GEA^{+}10, HS14, IS11, KG14b, LPS11, LBL^{+}10, MP10b, Nge11, PP12, RAP13, RGL12, SO14b, TH15, TAY11, VV12, Web12, WW11, ZY13, ZC11, dSE11, dCOSW13, BR19, Ber16b, BS17, BB17b, DZL18, DRAS17, FPZ16, FF19, FL18b, GWP19, IAB16, LL17a, LSY17, Mer16, NY16, PV16, SO17, WHHW17, WHYW18, Wie18, XZ17, XHH17, ZBA17, ZL16c]. Dependency [CCL^{+}13, AHP19]. Dependent [AGR15, BP15a, Ber12, DNL12, DS09, DP10a, DPP15b, Den13, FJ11, GS11, HS15d, HP10, JFA12, KL11, KY10b, Li13a, MC11, RSA13, RAB15, She12a, Sun11, Sun12, Sun15, TW13, Tee14, WY11, WHVH13, WW13c, WLY14b, WLHW14, WJ11, YMF15, ZL14b, Zör14, AHP19, AMP18, BR17, BN17, BAL18b, BD19, BKKP19, CSL17, CNL19, CS18, CSC13, CC19, CF16, CL17a, CXC19, CZW16, CDS19, DWM^{+}17, Ery18, FWW19, FS17, FL17, GHSZ19, GWZ17, Had17, Haj19, HDS16, HJW19, Hen17, HF19, Ili16, JSWL17, KM18, KTBN17, KSDB18, LM18a, Li17b, LL19a, LPH16, LWWH16, LG16b, LGG16, LSYL17, LYG17, LWG18a, LWG18b, LSL19, MWW18, MX19, MZ14, NRI17, OL19a, PSF16, Pes17, QC16, RSV17, RK18, ST16c, SZV16, SXW17, SX18, SS16e, SSR19, Tan18, TR16a, WP16, WH17b, WYY16, WV16, WY16, XX18]. dependent [XMP17, YVCA17, Yan17b, YCAJ18, YQC17, YWO^{+}15, YC17, YW18, ZC17a, ZZ17a, Zha17a, ZWZwZ17]. depending [HL18a]. Depth [IP13, YM18]. Derivation [ZB12b]. Derivative [Hua14b, Nge11, Rao17b, Son16]. Derivatives [CS15b, Sug13, Bos18, CCN17, GK19c, MDG19, WZK17, WZJH19]. Derived [Han12a, HYK15b, AAE^{+}19, SS19]. Deriving [WH14, LH17c]. descent [HT16a]. Descents [CN15, CN17c]. describing [Ali17, AMP18, BG16b]. Description [GV11]. descriptive [BOT17, OOT19, OTA17]. Design [AZN15, ASeMJ17, AMH14, BD11a, BM15, Ben14, BQDH12, BDM13, CL10b, CY10, CC10c, CC11, Cho10, CS11, CR13, DP14, DDM10, DD13, DSSL15, FG11, FN09, GCR12, GL11, GS15c, Hol15, HC10, Jaf13, JBA18, KSG13, KAJH18, sK13b, KLS12, LKN15, LCT12, LSSS14, LC15, NGT13, NFZ11, PP10, Par13a, PK12, PLDS17, QL11b, Sag15, Shi10, TGXX12, cTtL14, TN15, Wan17c, XFC11, YWLY11, BJU18, BS16c, CY17b, CSY19, CS17, DW19, FY17a, FPM19, FCSH16, HCZT19, JKJ16, Kha19b, Ko17a, Koy18, LK17b, LK19a, LS17a, LX17, LZ18, Liu17b, MB19, MV16b, OPK17, OQ17, PKM17, PMF17, PMKF17, PM18, PMFK18, PS17b, RFM18, SMSF19, SKK18, SM18, SS16d, WWZC18, Wu16a, YKL^{+}17, ZLT17, ZQZY18]. Designing [BU17, NS17, NN18, UB17, Bal18a, BAA^{+}18, NS18]. Designs [AJV11, AAG12a, AA11a, AIA14, AW10, ASN12, AZN15, AAG12b, Atk12, BJVV15, Bis12, BDM13, BDM15, BCDLA12, BSB12, CL14, CG12, CG14a, CT11, CL15a, CJ15, CYL13, DPA10, DPP15a, DDM10, DDM14, FHS15, FM12, FM14, FH13b, Fon14, FN09, GLZW12, Gar10, Gra12, Hil15, HLL15, Hua10, HM15, HYK15b, ITG09, Jac11, JR18a, KS12a, KMM12, LW13b, LHHC14, MGP11, MP12, MGP14, MS12a, MGM13, MH15, NC14, OQC14, PMA11, PM12, RS13, RZ14, RN13, SGPA12, SAC13, Sty10, TIA11, TE11, Tan10a, Tin10, TS12, UC15, VJV14, Wan10a, WW13b, WLL13, WYZ15, XY11b, YA13, YH11, ZZP10, Zha13, ZY14, ASR^{+}19, AGS18, ADKZ16, BQC17, BAT^{+}18, BO17, BJVV17, CG19, CSM16, CK17a, CZ19a, DS17a, DJVV17, DT17, DS17c, DS17d, FM17, GPM18, GQC18, HH16, He18, HME17]. designs [IBA18, JYZ19, KSF17, KM19a, Kha19a, KHY17, LK17c, LRC17, LH17a, LWWW16, LL16, LOHW19, LL19c, MPG16, Man17a, MD17a, MGP17, Man17b, MP17, MEZ16, MN19, Nor16, OPK18, Oll17, OQC17, PGCT19, PPMD19, PM17, QYLL17, RMM17, RAST18, RRRA18, RAAJ17, RAZ19, RK19, ROM16, Ros19, SMN18, ST16b, SG18b, SB16b, VV17, WZ16, Wan17d, WQ18, XK18, YZP17, YYS19, YZ18, ZF18, ZYLZ17, ZLYZ17, ZL16e, ZS17a, ZZ18e, ZZ19e, ZLZ19]. desirability [KDGV16]. destructive [POS^{+}19]. Detect [SR11]. Detecting [AHZ17, AR14, BANC17, Bos18, Dok12, GBC12, Ko11, LF17, NNN10, DY16, LW17b, Par17a, SI17, Wu19, ZT18, ZG17b]. Detection [AVCL13, CGL13, DPZ17, DC10a, DI10, ET13, Eid14, HTN15, JDSS16, KJD^{+}15, LX19, LSC12, MCY11, Nos14, PS13, QTD10, SA15b, WG15, BRZR16, Bar17c, CJ17, Dvo17, Gha19, GLY17, LWD^{+}18, PK18, QL19, SA19a, SL18, SW19a, WL17c, XHY16, ZDZ17, ZFZ17]. deteriorating [PM18]. Deterioration [KLK14, SK16d]. Determinacy [Sto12, BH17a]. Determination [BAL18b, DSO12, IP12, Ko14d, Ko15a, ZK15a, Ko16b, WAJ17]. Determine [Ko14a]. Determining [KAR16, VT17]. Deterministic [KY10b, AG18, Sti17]. Detrended [Ban13, Ai17, Gho17]. Developing [AEA17, WAJ17]. Development [AO18c, SS10b]. Developmental [Raz14]. Developments [ZL14a, CG19, DSA16]. Deviation [Aka10, CL15b, GM14, Hay10, Kor11, MCX10, Mia11, SI13, WLW15, WSlT15, CF16, LSW16, MGLX14, SYW16, SJY17, Tar17, WGW19, ZC17b]. Deviations [CGW15, Kon15a, KL11, Li14b, LSX14, VSS12, WW13c, BN18a, CS19c, DW17, FS17, FL18b, Gri17, GWZ17, HSLQ17, JY16, LJW18, LS17d, LYG17, Lou19, LSZ16, LQ17b, LSL18, MG19, ST16c, SHLQ17, XP18, YS16]. device [YSS19]. devices [Bow17, LL19a, TP19]. Diagnosis [ASMZ14, HTN15, RVD14]. Diagnostic [Iqb13, MAS16, NIN11, WQ14, BPV19, EP17, Sha16, TP19]. Diagnostics [CTW10, ER14, LSC13, SL15, UPA13b, XJW17, ZZC13, BD16b, CYF^{+}18, EE16, LLWK18, Sun17b, SBLC17, YCL17, ZVG19]. Diagonal [DDÚF11, LL14, KS16b, ST16a]. Diallel [BDM13, VV17]. dichotomization [NRN^{+}17]. Dichotomous [BD11a]. DIF [TK19]. Difference [FLK12, HWM12, KLM10, KZ15, Lim11, SR11, TJ13, WJY15, ZL16c, AR17b, AGS18, BAW19, CZM17, FJW17, FP17b, HSLQ17, KR17, Li19, Liu17b, LZF17, LCZA17a, LSW16, MMN16, MHA19, MKK17, Par17a, PK18, SHLQ17, Wu18a, Wu18b]. Difference-based [ZL16c, Li19, PK18, Wu18a, Wu18b]. difference-type [MKK17]. Differences [NN11a, Ryc11, Ima18, STD18, Spe17, Wan17g, WBGS19]. Different [FM16, FHZ12, HEHM11, Him10, HT13, OQL12, WBH10, AIIK17, JKD17, LW17c, OQ17, RRRA18, SCYR17, SCYR19, SX18, SBLC17, TPN18, XO19, ZW18]. differentiable [Has19]. Differential [Con10, Jaw14, MS11b, ZZ15, ZHL15, dSE11, FP17a, Gol16, LYX18, LL17e, LW18b, Rah18, RG16, XKW19]. Difficulty [Aca15]. Diffuse [Piz13]. Diffusion [BKP11, DW13, Han12b, Han15, Kor11, Kos14, KG14c, Lim11, Neg09, QWZ14, WL11, YSK15, CZCM19, CZ15b, DLZ17, HZ17b, LLLL18, LSW17, LMC18, Mud16, QJWC18, WRZ19, XDW19, Zho17]. Diffusions [NY11, SW19a]. Digraph [Cey11]. Dimension [Him10, HS12b, YS12, YHS13, YA10, YL10, ZYX12, ZZ17c, Zhu10, ZYZ11, DLS17, DBG17, FP17b, Gey18, HN17b, SA17b, XTS16, YHS17, Yam18b]. Dimensional [Afs13, AV15a, BNS11, KK14, KVV11, MTM13, MK13, SCSMN11, SLZ11, WSlT15, Yam12, ZHLZ14, Ahm17, AXLJ17, AJ18a, BCCR18, Das16a, Ery18, FJW17, FLL17, HKR15, JSWL17, LWZD17, LX19, LTZ19, LGG16, NAML16, PR17, PP17, ST16c, SLSC18, TN19, WL17c, WX19, XX17, YJSL19, YM17, YJ17, YX18, ZBHW18, ZG17b, ZZZ19].

Dimensionality [Lu13a, WSW11, YA10, CS17]. Dimensions [KS12c, RAP13, SL14, DY16, Mao17, SS17c, YCH17]. Direct [Chi12a, YC11b, Sen17a]. Directed [Cha14, LQYZ17]. direction [Wie18]. Directional [DPA10, Li14b, Sei13, SU11, SO14a, Sim14a, SO14b]. Directions [Vel14]. Dirichlet [Chi12b, Gho09, KJ18, LCG18, SUS19, TZ17, YM17, YJ17, ZHW17]. Disability [DGM13]. Disaggregation [ZA14]. disasters [YFV19]. discontinuous [LWD^{+}18]. discount [DLH^{+}18, WL17b]. Discounted [Mia11, MXD13, FL17, GWP19, LG16b, LWG18a, LSL19, XZ17]. Discounting [Don11]. Discoveries [MWW14]. Discovery [De 11, Mel13, Bic17, DZL18, WX17, YSC17a]. discrepancies [OQC17]. discrepancy [CFA17, GQC17, LOHW19, OQC14, YYS19]. Discrete [AA15a, CC12, Cha15, CCH^{+}10, DP13a, Ery16a, GL15a, Gue14, Gup15, HQW15, Jaf13, JRM10a, JB18, Kaw13, KRB13, KLK14, Lim11, LDY15, MB15, NC18a, NSP12, NS15a, NAB12, NABA15, PHW11, TN10, TW13, TEF12, WN17a, YSK15, Yan14, ZLZ15, AER16, AG18, BAKZ18, BB19b, BMLY19, Bos18, BD17b, CF19, CC16a, CC19, CXC19, DLH^{+}18, DS16a, DSA16, Djo17, Ery18, GG18, GN16, JS18, JCN17, KN18b, KN19, KH19, LWG18a, LWG18b, LD18b, LQW19, Mal17, MR16a, NSR17, NA18, OAOQ16, PFD19, RZHC16, SRN17, SSB17b, SSD19, SI19, TN17, Vas16, VFP17a, WN16b, Xia17b, ZL16a]. Discrete-Time [GL15a, Gue14, Lim11, PHW11, Yan14, CXC19, LWG18a, Mal17, OAOQ16, TN17, Vas16]. discretely [CY16]. Discriminant [GKW14, HS13b, HS12b, Kaz11, YHS13, ZHLZ14, SLSC18, Xia17a, YHS17, Yam18b]. discriminate [NRN^{+}17]. Discriminating [DK09, AdBPdBP18]. Discrimination [CB14, KK11a, Kit15, BCCR18, Pat16]. Discussion [LW17c, PMF17]. Disease [Ken11, Con16, NRN^{+}17, WER^{+}19]. Disequilibrium [WYG17]. disjoint [FY17b]. disk [Far17a]. disk-type [Far17a]. Disorders [Muk10]. Dispersion [CT11, CTW10, CBB06, GP12, HLL15, NA12, SS13a, Tin10, XWJ10, ZARH14, GOLM19, HK19b, LPC17, MBS17, MBS^{+}19, SA16]. disruption [SF16]. Dissimilarity [KB13]. Distance [AH14, CVC^{+}12, Gar14, GG10b, TIA11, BD16a, BK17b, FZ17a, GTW17, Kap18, Kay18, Kun17a, Tri17, WH17a]. Distance-Based [CVC^{+}12, AH14]. distances [DHK17, SSM16]. Distinct [ZZP10]. Distinguishability [FGTS12]. Distinguishing [Fle13]. Distorted [DFS10, SBRGD16, SSB17b]. Distortion [Yan15, Ric17, WZ19, ZF17, ZWNS18]. Distributed [AKK09, CY10, LSSS14, RF12, SYP13, VT15, YWLY11, Ali17, BZH19, BJU18, Çel17, GHSZ19, GC19, GW19, KKKÖ17, RK19, SHLQ17, TWZ16, TS17e, UZ19]. Distribution [AT15, AE13, AA15a, AT11, Afs13, AJGG14, AK14c, AFL15, AFL13, AO12, Ano13b, ASAN12, AVCG10, AGVB10, BTK10, BDKM12, BS12b, BHLK15, BZAZ15, BR12, BSCS11, BNS11, Bas14, BREA11, BBT13, BCK15, Can15, CLGFB15, Cha10, CG15, Cha15, CRG13, CG11b, CW15b, CJ19, Coo13, CT14, DPP15b, DN14, DC13, – Dor11, Duf10, EdSMR14, Fly10, Fu12c, FZ10c, GMB^{+}13, GL10a, GL15a, Gho09, GN15, Gil12, GFG13, GSA14, GL15b, GC12, GY13, Gui15, GG16b, Han12a, HN14, HTN15, HS15c, HD15, HSMK13b, HL13a, HKNL15, HHW14, HASA10, HME12, HGZG15, HP10, HLR11, IO13, IANT13, JBH12, JMEM14, Jia12, JRM10b, KSK15, Koi10, KG15, KDK14, KCK10, KP12a, KJAR13, KLM10, KA15, KR15, LMG13, LKK15, LH14a, LVBS10, LLP10, LCO14, LWY10]. Distribution [LL12, LQ13, LH13, LPN14, LJ12, Lop11, LF17, LVVN11, MEG15, MZ10b, MM13, MFSAS10, MV14, MH11, Mea15, Men10a, Men10b, Men13a, Min10, MB14, MO14a, Mit13, Mon14, MS10d, Mur13, MADSAS10, NH15a, NBK14, NSV12, Neg09, NAB12, NAA13, NABA15, NP15, NNN10, NMJY15, OJR15, OBPL15, PKJ14, Pao10, PGF09, PR11, Peg12, PBS12, Pet10, PCN15, Pop12, Por15, PDS09, QKP12, RAK15, Raz14, RCC12, RAB15, RLL12, RP12, Sag15, Saj11, SH11, SF12b, Sco09, SD14, SBK13, SJS15, sS12b, She12b, She13, SK15a, SA10, SK12c, SK15b, SB11, Sug13, SM14, TT10, TD14, TTZ14a, VRB13, VV15, VEKH15, Vin13, WXPS12, WL14, Wat11, Wit14, Xin15, Xu10, Xu12, YOD10, YSHL11, ZY13, ZS15, Zin12, AE17a, AE17b, AE17c, Abu18, AM17a]. distribution [ACO^{+}18, AGAM17, AA17b, AOÖ17, AT17, ABK17, AFL19, Alm18, ATO19, AG16, AAO17, AR17c, AD19, AJ18a, Ano23, ANA17, AEN16, ANS17, AEB17, AN17d, AAJ17, AAJ18, AGB17, AE23a, AE23c, AE23d, BK17a, BN18a, BSL^{+}19, BKN16, BKM18, Bar19, BP17, Bar17a, Bay16, BN18b, BAB^{+}18, BJ18, BB19b, BT16, BMLY19, BN16, dSBCO18, CSL17, ÇG18, CO16b, CC16a, CK19, CHM16, CG16, CD17, CG17, CCZR17, CXC19, CCM19, Che17b, CM16c, CFA17, Coi17, CLC^{+}16, CAPO17, Cos18, DPP17, DYXW19, DY18, DS16a, DMA17, DS17b, DBA18, DRAS17, DZH^{+}19, EW16, EDKAEN17, EM16, EMHK17, EN17, Fab16, FPR17, FYLC16, FTF17, zF16, GSM16, Gau17, Gen17, Gen19, GT19, GMMA19, Gho17, GH18, GHBM18, GA18, GFG16, God16, GRSN18, GDCO16, GDGG17, GT17, GW18, HG19]. distribution [Han19, HP17a, HA16b, HEB16, HT16a, He17, HME17, HLP17, HC16b, HR17, HWLP18, Huh16, Hut19, INOC17, IABG17, IVVG17, ICBG18, IR19, IS16, Jas19, JB18, JS18, JYX16, JY16, Jon18, JM16, JJ18, JSC17, JC18, Kal17, KAS17, KAPA19, KTR18, KJT17, KM19b, KY17a, KKH17, KI19, KKS18, Kon19, KD16b, Kou16, KX18, KN16a, KDN17, KD17b, KN18a, KVR17, KN19, KJ18, KS18b, LK19b, LHS19, LM16, LN17, LS17c, LMT17, LAF17, LM18b, LQW19, MFJ17, MK17, MTC^{+}19, Mar16, MS17b, MPD^{+}19, MKSS17, MMD18, MAS16, MOC16, MA17, MAA17, MML17, MDA16, MBAJ19, Mod17, MHF16, Mur16a, NCA16, NAC19, NRCG16, NB16, NORV19, NHN16, NAMAK17, NAB16, NB17, NM17, NM18, NY16, NA19, NN16b, Noo16, Nou17, OSX^{+}16, OMFB17, OHKK17, ÖGGB16, PBL19]. distribution [PWLO17, PN16, QJKC17, QK17, RS16, RLSL19, Rao18, RAM16, Ras17, RGV16, RMN16, RKMG19, RPH17, RK18, SGKB17, SP16a, SG18a, SK16a, SG16a, SS19, SBRGD16, SG16c, SK17, SBA17, Sha19, Sha18b, SSB17a, SSSM16, SC19, SUKK19, SM16, STRD17, ST17b, SSSK18, SSS19b, SK19b, SW19b, SU17a, TC17, TCM^{+}16, Tak18, Tar17, TS17b, TS17c, Ten18, The19, TWG18, TS17f, TR16b, TPJ18, Tsi18, Van16, VS19, Wan16b, Wan16c, WH16a, WC16, WL17a, WW17a, WWY17, WYY16, WN16b, WSS17, XQ17a, XQ17b, XZ16, Xu17a, YVCA17, YSS19, Yil16, YW17b, ZMCM16, ZRL19a, ZRL19b, ZFZ17, ZRN17, vSK16, vZ17, vZvdM19, RAA17]. Distribution-Free [BTK10, SH11, SBK13, WL14, BAB^{+}18, Gau17, Mar16, XQ17a]. Distributional [Cha10, KM10, LTCQ17]. Distributions [Afe13, AH13, AJGG14, AMGK13, AMGK15, AAE13, ADO13, AMP18, AF10, AALS12, BCS10, BNE15, BB14b, Bur15, Bur09, Cah15, CMCCO10, CC12, CC15a, CZ15a, Che11, CM13b, CA14, CdSBB15, DK09, DC10b, DGV11, DAL12, EB12, hFmH14, FLK12, FWH^{+}15, Fer14, FFA09, FF13, GL10a, GDCOS13, GCSB10, GC12, GN12, Gup15, HV11, HM12, HS12a, HLX14, IRB13, JAA11, JT11, KRB12, KRB13, Kim11, Kit15, Kon15b, KL11, KQ11, Kur10, LWL13, Lef10, LL11, Lem14, LCG12, LBL^{+}10, LD13, LHZR14, LG14, MS13, MS10a, MN10, MB15, Mat11, MMV12, MRS12, MSN14, MF10, Miy10, Mod12, NCO15, NSP12, NS15a, Ng12, OPSGB12, OCIS12, PT10, PS11, PJ14, PT15, Pun13, RP15, Ric13, RR09, Ryc11, SCMR15, ST10, SGD11, SSA10]. Distributions [SAB15, SDJ15, Sha10a, SBK13, SB14, SA10, Sto12, SSON11, TN10, TGXX12, TV11, TP14, TK14, TEF12, TH15, VLD15, WY11, Wan13b, WH13, Wil09, WW15c, XFC11, XY13, YS15a, Yeh12, Yeh15, Yu12, ZK15b, Zör14, AA19, AA17a, Ahm16, AHKO16, AO17a, Ala17, AJT16, AJ19, ARB19, ABTR19, AN17b, AGA17b, AdBPdBP18, BNZ17, Bar17c, BBG18, BH17b, Ber18, Bic17, Bob19, CL16, CK17c, CM19, CZ19b, CTAL17, CB16, CARO17, Das17a, DeR19, Djo17, DA17b, DL16, EW17, Ery17, FA19, GG18, GGDVP16, GMMA19, GCA18, GN16, GP16a, GN17, GR18a, Gup16, HOAH17, Hom17, HS17, HST19, Imo16, IF18, IS17, IAB16, JT16, JPAB17, JCN17, Jon16, Jon18, Kam19, KM16b, KK17a, KPZ16, Kun17a, KG17, KN18b, LKK17, Lie16, LW17c, LCWG16, Lou19, LQ17b, LQYZ17]. distributions [LQHZ18, MK17, MM17, MAYH17, MIG17, MAAI17, NC18a, NAN17a, NSR17, mNK17, NA17, NAYC17, NA18, NY17, OL19b, OGD17, PBC17, PP18, Par17b, PMFK18, PK19, PW16, PT16, PE17, RGBG19, RMNA17, RSAN17, RCdCN16, RJN16, RN17, SGKB17, SK19a, Sch16, SG16d, SHANA18, SSB17b, Shu18, Shu19, Smi19, Su17b, SI19, TCA^{+}16, TZ17, TKP17, Tsi15, TH17, VPCG16, VT17, Wan17b, WG16a, WCF18, XXMZ18, YYL17, Yeh17, ZVG19, ZL16a, ZWZ17, ZCCW18, ZKW16, ZY16, dPD16]. Disturbance [CK10, ÜY13]. disturbances [CS16c]. Divergence [BK09, KM10, KS15a, MP14, NSL14, Pak14, SK10b, Tan13, KHY18, PHY19, She18a]. Diverging [EM13, WSK14, CCY17, GYL17, LY17, ZL16d]. Dividend [hXZ17, TMZ^{+}17, XZ17, YWX17, ZWY16, ZDDH17]. Dividends [YSK15]. dividing [WYG17]. Divisia [Bia15, Bia17]. divisible [BB19b, SMN18]. DMFA [LP10]. DNA [GL19, SU13, SU17a]. Document [MSB19b]. Dodge [YF17]. Does [MC17]. Domain [LS12, LS15b, SNN12, Gen17]. Domains [BCG11, Lu14, Lu15, RP15, Sch16, Zad15, Lu16b, Mal17]. Dominance [SW14c, BCB19]. dominated [LSL18]. dominatedly [SHLQ17]. Domination [Cey11, Lu13a]. Don’t [SS13a]. dose [ZF18]. dose-response [ZF18]. Double [GS15c, KTW10, Kit15, LCT12, LL11, NAD15, OH19, WXPL13, ZM13, Zho17, Ami18, BAA^{+}18, BAW19, COK17, Çel17, Cle18, FNM19, HBMAO16a, HBM16, HBMAO16b, HZ18a, KOA16, LK17a, LK17d, LK19a, Lu18, MF16, NS17, NSSS17, OQ17, PBA17, PBA18, SKK18, Sha18a, ST16b, SK19b, SS16g, Wan16c, ZZ19c]. double-objective [SKK18]. double-ranked [HBMAO16a, HBMAO16b]. double-sampling [NS17]. Double-smoothed [Zho17]. double-truncated [SK19b]. Doubly [AB11b, DBA18, KPS10, Kha12, KRB12, RF12, SSD18, SG10c, She10b, sS12b, She13, She14b, Sug13, ZW09, ZS10, Car17, FA19, God16, GOLM19, Kun17a, Kun17b, PR17, SS16a, She16, ZL16f]. doubly-truncated [God16]. down [AE17c, QCK19, SSSM16]. Downtime [JHP10]. drawing [NA18]. DRHR [WZ10]. Drift [LS12, SY14b, OL19a, Sti17, TK17b, Zho17]. Driven [PW11, Wyl13, ZZ15, dSE11, LL17e, SJ19, SW19a, TMZZ19, YSC17b, ZLZ15]. dropout [BJ19]. Dropouts [LK14]. DRPFR} [OL19b]. drug [Liu17b]. drug-placebo [Liu17b]. Dual [EB12, El 15, Ken11, LP10, MOC16, AEHQ17, HKH17, KI19, LSY17, YSG17, ZLT17, ZT17, ZWY16]. due [Bur17]. Dunn [MSB19b]. Dunnett [Hot16]. duplex [VM17]. Duration [MH15, OBPL15, PV14, SS12a, ZZSP14, CK17b, DPP17, GDPR17, LS16b, Liu17c]. during [El 17, LL17f]. Dynamic [AB15, BCCH14, CM16a, CB14, CD15, Fal14, Gar14, KC12, KR14, MB17b, NSL14, NN11b, RON12, SBR14, TN15, Wu11, Wu13, BW17, BK18, BCB19, GSLZ16, HH19, HJT19, KAR16, Kum17, KN16b, KK17b, KG18b, LL18b, Min17, NSR^{+}19, Phi18, QJWC18, RASR17]. Dynamical [Rab13]. dynamically [AN17a]. Dynamics [PMC10, Tun15, TMZZ19]. Dzhaparidze [Voi10].

E-Bayesian [Han19, Han11, Sha19]. Each [YSHL11, Xu13, Xu16, XN19]. Early [KJD^{+}15, LJC18, PKM17, PM18]. Earthquake [VLTP14]. Economic [BJU18, CY10, CR13, HC10, KSF17, LK19a, MB19, NGT13, Par13a, PS17b, SMSF19, WWZC18, YWLY11, ZA14, ZLT17, CSM16, FY17a, FCSH16, GFM17, HME17, KM19a, MBAJ19, NN18, PKM17, PMF17, PMKF17, PM18, PMFK18, RMM17, RFM18, SKK18]. Economic-statistical [LK19a, KSF17]. Economical [BQDH12]. EDF [ZXD10]. Edge [CDE10, SRT14, AGS18]. Edgeworth [Aga11, DK13, EE14, Mae10, Qum12, Yam12]. editor [AR17c, SG17d, Aki11, HSMK13a, Jon18, Özg19a]. Editorial [Ano11e, Ano12d, Ano17a, Ano18]. educational [SSM16]. Effect [CEK11, CK10, HEHM11, HYK15a, sK12a, LX15, Liu11c, Otr15, PL12, SK10a, TWH10, XFC11, YW15, Yu12, ZSS15, ZB12a, Bel17, Ber16b, DT17, DRAS17, FEP18, sK16c, LL17c, LWZ18, MN19, OMB17, RC16, SKM17, Su17b, Wan17d, WN16a]. Effective [SMPH15, SMMS15, KYW16, PM16b, SP16c, SMK17, SK18b]. Effectiveness [CC10b, Ko12, SAS11, CSM16, LJC18]. Effects [AIA14, BF12, BJVV15, BC12d, BSB12, CT11, CS12, CW13, Chi10, Chi12a, Con10, DSO12, EIJK13, Eid12, GE11, Hua10, KR14, Li15, LR10b, Mån12, MMP14, Mon14, MH15, PMA11, PMFK18, Pul13, RMP12, RM15, SL15, TK17b, WLL13, WS14, WYZ15, BTC16, BJVV17, CS18, DAC17, DHKM17, EMI14, FZS19, FPCP16, FM17, GZ17, GPM18, JJ16, JKD17, LCG18, Liu17a, Man17a, MAS16, MC19, Nak17, RS19, SG16b, SE17, Sun17b, Sun17a, SJJ19, TLT17, XJW19, XXMZ18, YWZ18, Zha17e]. Effectual [SSB17c]. Efficiency [Aka10, AR17b, FF16, KV10, Ko14c, LKN15, Li18, Neg09, YDL14, CE17c, FF18, GPM18, Ko17a, ST16f, ZH16]. Efficient [AESO19, DLS17, Hu14a, Hu14b, JKJ16, KAD14, LX15, LL18b, LTV17, MGP14, Mao15, MSL17, MS12a, NC19a, RS13, Sha18a, SS13b, SS13c, SG17e, SSS17b, SS14b, Tu17, VEKH15, WR10, ZTZ^{+}13, KM18, MDG19, SP16c, ST16d, SG17f, SG17g, ST17a, SG17h, SKS17]. EGARCH [MS17a]. EGB1 [KM15]. EGB2 [KM15]. Eggenberger [SAB15]. eigenproblem [WW17b]. Eigenvalues [Sch12]. Eigenvectors [Din15]. either [CS16c]. EIV [AS17]. Elastic [KS10a, Kaz11, MTM13, YS19]. Elasticity [VFP17a, VFP17b]. Electricity [CCL14]. Elementary [LIPMF13, NC18b, Pes15]. Elements [HP10, Rao17a]. elicitation [MF19, SUS19]. Eliminated [GH13]. Eliminating [CF14]. Ellipses [DC14]. Ellipsoidal [AKL15, Pre11]. Elliptic [AKTS13]. Elliptical [APA14, ABLR15, DGV11, GG10a, MF14, OPSGB12, PJ14, SDJ15, VBAR15, AN17b, AN17c, BBR17, HL16, RGBG19, RJN16, Shu19, YY16]. Elliptically [ASNT15, Ric13, ANA17, Rez17]. Embeddable [Gue14]. Embeddings [BCG11]. Emergency [Cui12]. Emerging [KJD^{+}15]. emphasis [LS16a]. Empirical [Bak11, BT13, BEZ11, BZ13, CL10a, Che11, CSF12, CW13, CLC^{+}14, CW16, Che16c, Che17a, CLW17, CLL17, EE15b, FLX11, FLW12, FH13a, FJW17, FMS15, HME12, HSHW14, Hu11a, HGZG15, HXF18, Hua11, Hua14b, LQ17a, LQ13, LLHZ16, Lia10, LW14, LW15, LLWK18, Mar15, MT16, QL11b, QL15, Qum12, SZ13, TX14, WDZ11, WCL10, WWL13, WQ14, WWT15, WW15c, XTG15, XTS16, zXyL13, YXC11, YCH17, YCZ17, Yu10, hYhW14, YHC14, ZK13a, ZWZ11, ZZ11, ZX10, ZW11, ZWPZ15, ZSY13, ZL18, AABL17, BG19b, Bic17, CNG17, CL17b, CW19b, FLL17, FM16, HZ18b, JSZL17, LCZ16, LD18a, LXT17, LQ19, MS17b, SNWL17, SLF18, SQCL19, TS17f, Van16, WHC16, Wan16d, Wan17a, WZZ17, YZ17a, YW17b, YD19b, ZT17, ZBTL17, ZRL19a]. Empirical-Likelihood-Based [SZ13]. Empirically [KYW16]. Encouragement [RMP12]. End [Ano18, Koi10, WH16a, DWX17, DW17, HSLQ17, Yan19]. Endogeneity [DMRO14, LG18]. Endogenous [LCLJ14, CLZ16, YCZ17]. endpoints [Tar17, ZF18]. Energies [GGL14]. Engel [LS17d]. Engineering [Ben14]. English [GDDC19]. enhanced [HUK19, HL18b]. enhancing [JR18a]. Ensemble [Gue13]. Entrance [XLL13, XX18]. entries [WYY16]. entropic [MBG19]. Entropies [SBR14, KN16b, SS17a]. Entropy [AA10, AB15, AM17a, Bar10, BR12, BEKL13, CB14, GL19, GIT14, JY14, LLV10, PS12, AC16, BK18, Das17b, EIM18, GGtH^{+}19, KS16a, Kay16, KR16, Kum17, KK17b, KK18, Min17, MB17b, MBF17, Mur16a, NP17a, NN16b, NY17, PM16a, Pen17, PE17, Qiu18, RASR17, SY17a, SK19b, Slj17, Too17, YYS^{+}18]. Entropy-Based [PS12, AM17a, GL19, EIM18]. entropy-type [KS16a]. Entry [WK13]. Environment [CDZ14, SY17b]. Environments [MO14b, HYS18, WCZP16, ZZ17d]. Enzyme [Atk12]. EOV [Ano11e, Ano12d, Ano17a, Ano18]. EPC [HC12]. epidemic [EKZ16, FP17a, GMH16]. Epidemiology [KJD^{+}15]. EPMC [YHS13]. Epsilon [AVCG10, MH11, MN17]. Epsilon-Skew-Normal [AVCG10, MH11, MN17]. Equal [KQ17, Cle18, WW17b]. equal-precision [WW17b]. Equal-tailed [KQ17]. Equalities [WL13, Ren16]. Equality [Bea15, Him10, Mod12, PS11, Sch12, VV12, YW11a, ESAKK18, MMP18, PK17, YSS18, YY16]. Equation [GHY11, Han15, KvdVvZ13, LY11b, MS11b, DTT^{+}16, FP17a, LL17e, RG16, Sel17a, Wan17g, XKW19, ZKS18]. Equations [BCG11, Fan15, Fed14, Jaw14, KJL11, Lim11, LX13, MS10a, NY11, XGZ11, YTX14, ZZ15, ZHL15, dSE11, CCM19, LYX18, LW18b, PCL16, Rah18, SÁHLvR17, Tu17]. Equicorrelated [LR10a, RL12]. Equilibrium [Kyr16, LRZ17, NSP12, SY17a]. equipartition [HYS17]. Equipower [Sal15]. equipped [LL19a]. Equity [QWZ14, YSK15, QWWW19, ZWY16]. Equity-indexed [QWZ14, YSK15, QWWW19]. Equivalence [GST17, HB13, SA19b, JSZL17, Kam16, MH16b, ZZ19a]. Equivalent [Guo14, HZZ19, TE11, WY11, LO17]. Equivariant [CR16, SEA15, YDC15]. Erdos [Dvo17]. Erdos-type [Dvo17]. Ergodic [zXyL13, BLT16, BD17a, BD17b, LL17d, Tra17]. Ergodicity [JZ17b, JB15, LS09, Sen14, TMZZ19, ZWPZ15]. Erlang [EMHK17, YOD10]. Errata [Ano13c]. Erratum [Ano11f, Ano12e, Ano13d, Ano13e, Ano15b, Ano15c, Ano16a, Ano16b, Ano17b, Ano19f]. Error [AB09, Aoy10, BCDLA12, CY14, CKD14, DPB13, GG10a, GT15, GB12, GR15, Hu14a, Hua11, KY10a, KSBM14b, KSBM14a, KE10, KS12b, Li12b, LPN14, Liu11c, Mah12b, MV13, MS10b, Muk10, Nad10, Oga15, Otr15, Par13b, SGM11, SE13, SL15, Tan13, Tee12, Tsi14, WS13, WX15, WJH15, Xu12, XY13, YA13, Yam12, hYhW14, ZZ11, ZL14b, AH17b, BRZR16, CS19a, CD17, Che19b, CN17b, Coi17, DHKM17, DWW18, DA16, GAS17, GB18, Gha19, HL18a, HR17, JZTN17, KA18, LXT17, LL19c, Mao18, MB17a, MAG18, MGLX14, MGN16, Nak19, PY16a, PTV18, RK16, Roy16, WK19, WF17, XTS16, Xu17a, XO19, Yad16, YX17, YAA^{+}17, Zar19, ZZL16, Zha17e, ZWNS18, ZL16g, ZL18]. error-based [DWW18]. Error-Contaminated [GR15]. error-in-responses [XTS16]. Error-in-variable [hYhW14, JZTN17]. Error-in-Variables [AB09]. Error-Prone [ZZ11, WJH15, LXT17]. Errors [Aga11, Alk10, ADWR11, AMH14, BQDH12, CG14a, CHH15, CYL13, DB13, DP14, DG12b, DMT^{+}15, ER14, FLW12, FH13a, FHS15, GdC12, GH13, Gil14, HS11, LY11a, LS11, LPLW13, LW14, LW15, LH11c, LWHW11, MF14, MW10, ÖA15, QL15, RF12, SP11a, SSM12, SZD12, WW13a, Wan13c, XLL15a, YC12, Yan12, YX15, YS15b, ZK13a, ZL12, ZZC13, AD16, AD17c, BQC17, BMQ16, BO17, Cha18, CW19a, FJW17, FPCP16, FP17b, Hay17, HJW19, HJT19, HZ19, JKL19, KGH19, LLY16, LCZ16, LX17, LLWK18, MB17a, MKY16, MR16b, NS18, NVS19, Pes17, Pop16, RWY16, Sun17b, TK17a, WP16, WH17b, WZK17, WL19, Wie18, XLS19, XQ17a, XFWC19, YZP17, YWH17, YZFM17, ZL16b, ZDZ17, ZF17, ZDW17, cZzXgL18, ZZL19]. Errors-in-Covariates [CHH15]. Errors-in-Variables [FLW12, FH13a, GdC12, LWHW11, MF14, SZD12, YX15, LLWK18, Pes17, XLS19, XFWC19, ZL16b, ZZL19]. Esscher [HG19, LLLL18]. Esséen [EdSMR14, Fro17a, Hui10, LPLW13, LLM16, MXD13, Sze17, TN13, TCN15, WWH15a, YWH14, ZL16b]. Establishing [CCCT16]. Estimable [Hua10]. Estimate [BAK13, DMT^{+}15, EdSMR14, Fly10, Gue13, Lia11, XYL14, Yan14, ZYX12, ALR19, BR17, BLT16, CS16b, CLZ19, HK19a, Hen17, HHC16, Rah18, WP16, XZD19, ZZ18a]. Estimated [DI10, KAD14, LY11b, Mar11, RON12, Xu13, COK17, ÇE17a, HCZT19, KNHM17, KN17b, Rao17a, WYMZ17, Xu16, XN19, YKL^{+}17]. Estimates [CKD14, FP12, GLZW12, GG13, GW15b, Li12b, LMR15, TEF12, UPA13a, ZLN11, BH17a, Bic17, FST16, Lu16a, ML19, SS16f, WH17b, Zen17b]. Estimating [AC15, BR13, BM16, Bur17, Che10a, Chi10, Fed14, FCJ^{+}17, GZ17, GHY11, GE11, Hug09, JT19, JLL13, JY14, KS15a, KT11, LY11b, MS10a, MAG18, Mod17, MADSAS10, PK19, RPY^{+}13, RL11b, SG10a, SK13a, Sha19, TSS15, TD14, TTZ14a, TKJ17, WL10, WLL13, Yad16, YOD10, YW15, ASG17, Bic17, Bos18, CC16c, CCM19, El 19, Gho17, GPM18, JKL19, LHS19, MKY16, MKK17, PGCT19, RSC16, Sen16b, Sen17a, Sen17b, SSS16, SP16d, ST17b, SSSK18, SSS19b, TS17b, Van16, WW17a, Yan17a]. Estimation [Abd15, AT19, AHH16, ADP10, AC11, AMV15, ABI12, Alk10, AN17a, AKTS13, ADWR11, Arf13, AS13, AZN15, AB19, AF10, BN17, BG11a, BH12, BZAZ15, BS16a, BAT11, BGG11, Ber12, BL13, Bob19, Cah12, CYF^{+}18, CG12, CW14, CSB19, CS14a, CK17c, CK19, CS15b, CL10c, CW13, CZ15b, CF13, Che13, CW15b, CA12, CM13a, Ciu17, CBM14, CLV14, DNL12, DMA17, DG12b, DN14, Dji10, DCL12, Dos11, DD13, ES14, ER11, EP11, Fed13, FHZ12, GG10a, GSY10, GZSS14, GL15a, GMRN13, GR11, GP12, Gil14, GD12, GMP11, GT17, GJM14, Gro10, GY13, GHL17, GSSCR12, GS15c, HA18, Han11, Han12b, Han15, HT16a, HH19, HASA10, HF14, HS15d, HA14, HL17, Hu13, Hu14a, Huh16, HLR11, HSSP16, Isl17, JZ16, KS16a, KD16a, KTR18]. Estimation [Ken11, KGH19, KK11b, KY10a, KG14b, Koi10, KVV11, Kon10, KZW09, LMCH12, LMR13, LH11a, Li11, LL12, LW13a, LG18, LCG18, LLY14, LR10b, Liu11a, LWH12, Lop11, LAW^{+}19, LSC13, Lu13b, LWWR18, MSB19a, MR15, MP12, Mao15, Mat14, MN13b, MRS12, ML13, MBN14, MB10, MW10, MSK17, Mur13, Nad10, NBS13, NBK14, NF11b, NF11a, NRP16, NP17b, Nge11, NRW15, NK15, NH15b, NP15, OU16, Pak14, PMA11, PM12, Pan19, PJ10, PGF11, PBY^{+}14, PYB^{+}15, PMC13, PD10, PJL13, PW11, PDS09, PT15, QL11a, QNP12, QKCH19, Rah16, RASMN14, RSV17, Rao10, RAK15, RAA17, RvdM14, RNB12, RMP12, RM15, RPH17, RL09, RMC11, SC10, SBL13, ST10, Saj11, SG15, SG17a, SG17b, SG10c, SGM11, SSS14a, sS12b, She13, She15b, SRT14, SV10, SKK10, SK12b, SS13d, SSS14b, SK14]. Estimation [SMMS15, SKM16a, STRD17, SSS18, SG19, SS15c, SRK11, SU14, SB11, SS14c, SÁHLvR17, TN10, TR16b, TPJ18, TL15b, VT15, VEKH15, VLD15, VLTP14, WR10, WL11, Wan12, WK12, WCGZ12, WZ13, Wan13c, WG15, WTZ15, WH15a, WLW15, Wan16b, WL17a, Wee12, Woj11, Wu13, WY16, WL17d, XLS19, YA13, YLS10, YW11a, YSG17, YA10, Yu12, YWLZ17, ZZA12, ZK13a, ZWLI13, ZTZ^{+}13, Zha17e, ZSWA19, ZZ10, ZJZ17, ZW09, ZHL15, ZDS16, ZYZ11, ZZL19, Zil11, ASS18, AM16, Ale16, ATO19, ÁF16, AH17b, BD16a, BTC16, BS16b, BBA16, BS16c, BSH17, BT18a, Ber16b, BT18b, BP18c, BG18, BK19, BRAS17, BCL16, BG17, BJC16, Car19, CS18, CR16, CCN17, CLZ16, CKS17, CN17a, CCZR17, CCM19, CKT18, CDS19, CLC^{+}16, CW19c, DLS17, DPZ17, Das17a, DAS16b, DWW18, DA17b, DDS16, DLZ17, EW17]. estimation [FPM19, GTW17, Geo17, GMR16, GFG16, GP19, GK19b, GICRR17, GX18, GK19c, Han19, HBMAO16a, He17, HZ18a, HZ19, HLR16, HAZ16, Iga16, IKH19, IF18, JZTN17, JHM19, Jun17, Kam19, KAPA19, KM18, KKH17, KJS19, KOA16, Kom17, Kon15a, KSBN17, KTBN17, KSDB18, KT17b, LA17, LPdDTH17, LQ17a, LM18a, LY16b, LLY16, LWZD17, LS17a, Li17d, LZZL17, LX17, LLM16, LZS17, LWZ18, LL17d, LARS19, LY17, LP17, LL18b, LWD^{+}18, LL18a, Liu19, LCZA17b, LCA17, LLF19, MM19, Mal17, MATC18, MA18, MT16, MMP17, MRA16, Mur16a, NE17, NLRD17, NB16, NNvRS17, NSSS17, NC19b, Noo16, Oga16, OGD17, OTA17, Pak17, PC16a, PMW17, PNPB19, Phi18, Pop16, Qin17, Qom17, RC16, RLSL19, Rao17b, Rej16, RMN16, RG19, SNSA17, SAD19, SKG18, SGGdML17, SCYR17, SCYR19, SG16c, Sen18b, Sha17, SV17, SA17a]. estimation [SX18, SQCL19, SS16c, SM16, SP16b, SKM16b, SHM16, SG17e, SSB17c, SP17b, SP17c, SK18b, SSKP18, Slj17, ST18, SS16h, Son16, SSR19, SK16e, SC18, SL15, Sun17b, Sun17a, Sun17c, SBLC17, Tak18, TZW16, Ten17, Ten18, TT17, Tu17, TME19, VS19, WT15, WSW16, WZ16, Wan16c, WC16, Wan17e, WNL18, Wan19a, Wan19b, WX19, WGG^{+}19, WF17, Wu17a, WHY18, WZ19, XTT17, XZA18, XZ16, XMP17, XM17b, XKW19, YHS17, Yam18b, YLG16a, YW17a, Yan17b, YL18b, YGM19, YWLW18, YCL17, YSC17b, ZW19a, ZZ19b, ZZL16, ZB17, ZWNS18, ZL16c, ZL16d, ZWPB17, ZXF17, ZZ19d, Zho17, cZzXgL18, ZY19, ZX16, ZRN17, dAdLA18, vZ17, GK19a]. Estimations [CC12, Tia12, WJY15, Han17, KL17b, Li17a, Ren16, WL19, You17]. Estimator [AAP11, AK13, AK14b, AS15, BNS11, BAT15, BEKL13, CS15a, CG11a, CG11b, CLX14, EY10, ES11a, EM13, Fak15, Fau12, GFG13, GÖ12, GK11, HSHW14, HS11, HL15, HWM12, IT12, IS11, JO10, KMSS14, KCC14, KP12a, LG10, LWY10, LYX11, LY11a, LPLW13, Li14b, LL14, LCP^{+}14, LG14, LG11, LGX12, LN13b, LM14, MFS12, MYM11, MC11, NO12, ÖK15, Par12, PS12, QLZ14, QNZ14, RS14, SY11, She14b, SS15a, ŠS12c, Šir15, SHLK10, SK15b, TC14, Tsu14, TB13, WL10, WLF11, WH15a, WW11, WJ11, WY13, Xu12, XY12, Xu13, XY13, YC10, YC11a, YDC15, ZLZ15, Zha15a, ZTSL15, ZL14b, dCOSW13, AR17b, ABK17, Alg19, dCA19, AJ18b, ANA17, AW16, AEA17, Bak17, BAKZ18, Bin17, BB17b, CW18, Cha18, CM16b, Che16a, CE17b, CLZ18, Che19b, CN17b, CE17c, DDD17, EW16, EAE18, FZ17a]. estimator [FAT17, Fat13, FJS^{+}18, FRC16, FP17b, FTF17, GR17, GB18, Gha19, Ghr19, God16, GGO17, GMS18, HKH17, HWH18, HWY17, HLR16, IES^{+}17, JSSO18, JXW17, KD17a, Kaç19, KAT18, Koy18, KNH17, LP19, Li16, Li19, LY16c, LL19b, LTS16, LO17, MO18, RM16, MGLX14, MGN16, MHF16, Nas18, OH19, OOT19, ÖK17, Ozg19b, SB16a, SGGdML17, Sen16a, SA16, She18a, SNW19, SSS16, SMP18, SPY18, SA19b, SS16i, TK17a, VW18, Vis18, WH16a, Wu16b, Wu16c, WF17, Wu17b, WA17, Wu18a, Wu18b, WAA18, WCF18, Xia17a, XJW19, Xu16, Xu17a, YH16, YYY17, YWH17, YWL17, Yil18, ZC17b, ZDW17, ZL17, FM16]. Estimators [Arn11, ABC12, BGG11, BS15, Bur09, CGW15, CY14, CK10, Chr10, DM11, DK13, DP10b, DSSL15, ES14, FJ11, Far11, Gil12, GBD10, GMN13, Gru12, HS13a, HB13, HYK15a, Hon17, HK13, HL13b, HS14, JRM10b, JFA12, KMS13, Kir12, LMG13, Lec11, LX10b, LY10, LvR12, Li13a, Li14a, MKS15, Mah12b, MS15, MSW13, Neg09, Oga11, ÖÇO14, Pet10, Rah12, RB11, SN12, SG10b, SEA15, She10a, She12b, SY14b, SP13, SS13b, SS13c, SS13e, SST14, ŠS12c, Šir14, SS14b, SS15b, SS15c, SKHL10, Sug13, SV11, TTC17, TT10, TN15, ÜY13, WSW11, WZ15, WWH15b, WWW13, XY11a, YUSC13, YUSC14a, YUSC14b, YC12, YLX10, Yan12, YMF15, ZL12, ZFL13, AS19, Ahm16, AMP17, ARJF19, AESO19, AEA17, BN18a, Bay16, BBR17, BD17b, CZCM19, CS16a, CX17, CH19a, CS19a, CW19a, CM19, Che17b]. estimators [CM16d, DK17a, DBA18, DA16, DS16b, DS17e, EEE16, FM16, HBM16, HW18a, HJC17, IK18, Ima17, JKJ16, JYX16, JY16, KT17a, KJ17, KR16, KKS18, KZ19, KA18, LY16a, LN17, Li18, Lin19, Liu17b, LTC17, LTV17, MSK18, MBS^{+}19, MMT17, MMD18, MKK17, NO18, NIN17, NVS19, PK18, PM16b, PT19, Ram17, SG16a, SSS19a, SG17c, SGA19, SKT16, SP17a, SSS17a, SPS17, SKS17, Sla17, SS16g, SA17c, SSY18, SJJ19, TKP17, Wan16a, WZK17, Wu16c, XHL18a, XN19, XO19, YGL16, ZB19, ZXSZ18, ZRL19a, ZL18, Zin18, ZHZA18]. Euclidean [Lii17, dCOSW13]. Euler [LL17e]. European [CMN^{+}18]. EV [YX15, CKS17, Liu11b, MLGZ14, WCL10, Yan12, ZSS18]. Evaluate [Cui12, DZ14, Ko14c, Ko17a]. Evaluating [Chi11, TWH10, AL16]. Evaluation [BAK13, CKD14, CDZ14, Cui12, DDGM13, DK17a, EI11, FHS15, Fon14, LH14b, ÖK15, SW14c, WZT15, ZSD^{+}15, AHEA18, DTT^{+}16, GAS17, Kha16, LdM17, NKG16, NRN^{+}17, OPK18, RWY16, WCZP16, aYwSlDmS17]. Evaluations [AA13, Liu17b]. even [SY17c, TY17]. even-odd [TY17]. evenness [HK19c]. Event [DS09, DC10a, Ko14a, Ko15a, MH15, ZZS13, ZWZZ15, Ber12, BG16a, DDS16, EMI14, Fan18, HDS16, HNB17, LL17c, She18b]. Event-Dependent [DS09]. Events [Ber12, DS09, DPB13, FNDM12, KJD^{+}15, KNK13, KCS12, LSZ13, ZWZZ15, FMPF16, GZ16, Kon16, OAOQ16, PF19, WN16a]. every [Van17]. everywhere [YZG17]. Evidence [MC19]. Evolution [Fan15, Kor11, MA11, SY17a]. Evolutionary [GG10b]. Evolutions [Lim11]. EWMA [ARD14, AR17a, Haq18, HUK19, HK19b, HCZT19, KAJH18, KTW10, KCL^{+}16, MN13a, SZ13, ZSJ16]. EWMA- [HUK19]. EWMA-Type [ARD14]. EWMAMS [DHKM17]. Exact [AA17a, BF10, BF12, CC16c, CG17, CM13b, CA14, Cox15, Fu11b, GC12, Ham12, IUT12, KJD^{+}15, Kir12, Mah12a, Pao10, Pie11, RB10, RZ14, Tak10a, Wan12, Xu12, pZ13, vZ17, Bar17a, Bar17b, KC16, RK18, ZS17b]. Example [Afs13]. Examples [Car17]. Exceedance [Sto11]. Exceedance-Type [Sto11]. Excel [Pel17]. exceptional [KY16]. Excess [XZNL15, Zoi09, FPZ16, LRZ17]. excess-of-loss [LRZ17]. Exchange [LLLL18]. Exchangeability [Mod12]. Exchangeable [Ery12, MS13, RF12, Wit14, BG17, PR17, SHB17, YYL17]. Exciting [MSP12, WWWX19]. Exhaustive [Kem11]. exist [Nie17]. Existence [HYK15b, KHY17, MS11b, SY17b, Zen17b]. existing [Kia17]. Exit [SS10b, SS14a]. exogeneity [Tch19]. Exogenous [XLPL12, XLL15a]. Exotic [Vas13]. Expansion [Aga11, Mae10, OCIS12, Qum12, SY10, Xu10, Yam12, Ai17, AS18, Are16, CMN^{+}18, Cos18, HWLP18, KKKÖ17, She18a, ZC17b]. Expansions [AKK09, BCH15, EE14, EdSMR14, KK11a, KOP13, KOP14, NY11, Oga11, WH13, CD17, Hac18, HLP17, WN17a]. EXPAR [MDC16]. expectancy [SS16f]. Expectation [AH13, KN18c, LWL13, LMM15, Wan15, ZRL19b, CF16, CHZ17, CPW19, GHSZ19, HZW17, Slj17, WHYW18]. Expectations [Ryc14a, Ryc14b, ZC14, CL17a, FWW19, Gor17, GR18a, HC16a, KR17, KG17, WW19, YW18]. Expected [DHK17, EF11, MWW14, VYZ14, Bia17, Kar17, NCA17, XZ17, YZ17a]. expectile [WZ19, XQ17b, ZZ18d]. Expenditure [AZN15]. Experiment [Jaf13]. Experimental [Atk12, FH13b, FG11, PK12, ADKZ16, Bow17, LZ18, MP17]. Experimentation [RSV11]. Experiments [AGVGP^{+}11, BO14, BSB12, CEK11, CS11, NFZ11, Pre11, Sty10, WW13b, Web11, ZB12a, BO17, BVJV17, CE17b, EIM18, KB17, LL16, MEZ16, Mat18, Nor16, PLDS17, ST16b, Wan17c, Wan17d]. Expert [SA14]. Explanatory [BSSJ11, ZW11, YW17a]. Explicit [ES11b, Shu19, Tia12, YJ19]. exploration [LDL18]. Exploratory [Bea13, WH18]. Exploring [FTF17, GBC12]. Exponential [Aka10, AMZ15, AM15, BDKM12, BHLK15, BHB15, BZAZ15, BR12, Bas14, BBT13, Bos15, CXS11, CN17c, CT14, DB13, DB15, DJ15, DLL13, DC10b, DAL12, GT19, GY13, GW15b, HA16a, HN14, HS15b, HD15, JMEM14, KG15, Kum15, LL11, Lem14, LJ13, LDS18, MN10, MADSAS10, NH15a, NA12, NAB12, NABA15, PJ10, Pao10, Pet10, PJL13, PCN15, QKP12, RD16, RB10, SKG12, TC14, TTC17, TTZ14a, VZ19, WTB11, WWW17, XY09, YUSC14b, YCL14, ZLD11, ZS15, AA17a, AHKO16, AO17a, AN17a, AB19, AAK17, AAJ17, AGB17, BNB18, BKGD17, Bob19, BN16, CBSN16, CR16, DP17, FYLC16, GCA18, GOLM19, HD17, HT16a, He17, Hut19, IF18, IS16, JN18, JT19, Jon18, KNC18, KQ17, Kou16, KX18, KDN17, KZ19, LQ17a, LW17c, LWWH16, LTS16]. exponential [LQYZ17, MDG19, MFJ17, MK17, MP16, MTC^{+}19, MDA16, NP17b, Noo16, NC18c, PC16a, PK19, PM16b, PY16b, QJKC17, QK17, RMM17, SNSA17, SK18a, SG17c, Shu18, SP17a, SSS17a, SP17c, SS18, SPY18, Sun17a, TR16b, Wan16b, WN16a, YVCA17, YWX17, vSK16, vZvdM19]. Exponential-Gamma [CXS11]. Exponential-Geometric [DAL12, BZAZ15, BN16]. exponential-type [SP17a, SSS17a]. exponential-weighted [GCA18]. Exponentiality [BT13, PMA14]. Exponentially [AP17, MZ10a, MRS12, Men10b, Men13a, AAJ16, GW19, Kre18, TS17e, WT18, YKL^{+}17]. Exponentiated [AF10, AE23d, CG15, Coo13, GDSC15a, KSDR15, LK19b, PWLO17, AE23b, BNB17, BNB19, HNB19, Han19, KTR18, KD17b, MOC16, MAYH17, NAB16, RAA17, RMNA17, Smi19]. exponentiated-log-logistic [MOC16]. Expository [Ham13]. expressed [KR17]. Expression [HL11, ML19]. expressions [AIIK17, NC18b, vZ17]. Extended [AJGG14, AO12, Bos15, Can15, CS14a, CYY14, DAL12, GMB^{+}13, GP12, GdSC15b, KCS12, KA15, KDN17, LVH19, OPK17, PCN15, SCC12, SSON11, WL10, WHVH13, WLHW14, AMV18, AAL17, BSL^{+}19, BKN16, BNB18, CL16, CLC^{+}16, GQC18, MML17, NA17, OPK18, ÖK17, Sab19, SXW17, WH17b, YQC17, YW18, ZW17a, Zha17a, ZW18]. Extending [AASRZE17, Zhu10]. Extension [AVCG10, CZ15a, Li13b, OPSGB12, PP10, YWLY11, ATO19, Bak17, Bud16, Coo19, Djo17, GQC17, HME17, JSS16, KJT17, KK17b, KK18, NA19, PM18, PN16, SS19, Sch17, Sha18b, Wu17d, Zha16b]. Extensions [BM10, Dji10, Hay10, LRC17, OCIS12, BAT^{+}18]. Extent [GH13]. Extinction [VYZ14, Fro17b]. Extra [DZ15]. Extract [LH11b]. extragalactic [DBC16]. Extrapolation [BCDLA12, HM15, Wan17e, Sab19]. Extremal [BNE15, Kon11, AEHQ17, CY17a]. Extreme [AT11, BGG11, CP14, Cra14, Fer14, GD12, GMN13, HLR16, RL11a, Ryc14a, Yan15, AEB17, Ber16b, BBS16, DYL17, FMPF16, GGO17, HWLP18, LK19b, LYR^{+}18, LL19b, OGD17, PF19, RMN16, Roo18, Sch16, SK17, WDLZ16, ZWZ17]. Extreme-Value [GMN13, OGD17]. Extremes [FZ10c, LZPJ11, TW13, CD17, FL18a, HLP17, LP18]. Extremum [Kom17].

F [Ano23, SG17d]. facilitate [Cox19]. Factor [ADWR11, Bea13, Bea15, Bi15, DD13, FHS15, Hua10, LP10, LH11b, OQL12, SBL13, SNN12, WS14, BH17a, CMN^{+}18, FZ18, JSC17, MH16b, SAKK15, YWZ18, YM17, ZZ19e, ZH16, ZS17b]. factor-covariate [FZ18]. Factorial [CT11, CEK11, CJ15, Fon14, HYK15b, Jac11, KG14a, LHHC14, MGM13, MH15, Tin10, TT11, WW13b, WYZ15, YH11, BVJV17, CZ19a, DT17, GPM18, KHY17, MH16b, Nor16, PPMD19, Wan17c, Wan17d, ZL16e]. Factorials [OQC14, GQC17]. factorization [TZ17]. Factors [AA13, CT11, Con12, HLL15, MGM13, SCC12, Tin10, WYZ15, Web11, ZZP10, Bic19, CRTdC17, FP17b, LWG18a, PPMD19, XK18, ZS17b, ZLZ19]. Fail [BHB15, FTF17]. Fail-Safe [BHB15, FTF17]. fails [Yan16]. Failure [AA14, BLN13, Bas14, CY10, CB12, COL15, DN14, FMS15, Han11, Jia12, KRB12, sK11, Ko14a, Ko14c, Ko15b, LY11b, PS14, QZZN10, QNZ14, RM14, Ryc14a, SJS15, TWH10, TK14, WS11, ZW13, AE17c, CSY19, CB17, CF19, DE17a, HA16b, KJT17, Ko16a, MKSS17, NB17, PKM17, RLSL19, SDS17, TPN18, YCL17, ZW17a]. Failures [BC12a, CGSM10, Sco11]. Fair [DW13]. Faithful [AT17]. falls [Kon19]. False [De 11, DZL18, Mel13, MWW14, Bic17]. Familial [SC10, Den17]. Families [AMZ15, FFA09, MS15, MA14, Ryc11, SA10, BMFS18, CM18, DL16, HBMAO16a, LQ17a, SSS17a, WN16a, fZWD19]. Family [Aka10, AJGG14, Ano13b, AF10, Cey11, Cha10, CK10, CZ15a, Coo12, GL10a, GEA^{+}10, GCSB10, GC12, HS13a, IAA15, LCG12, LHZR14, LG14, MS10d, NCO15, OPSGB12, Pun13, SSON11, TB12, TV11, Wan12, YUSC13, AA19, AMP17, AHKO16, AO17a, AN17a, AL16, BKGD17, CM19, CB16, CARO17, CBOR18, GGDVP16, Ghr19, Gup16, HOAH17, Han19, Har17, JKJ16, KCY^{+}19, KN18b, KH18, MN17, MM17, MAYH17, MAAI17, NORV19, NAYC17, RSAN17, SSB17b, Shu18, SKT16, SS16g, TCA^{+}16, TWG18, TKP17, Wan17b, ZY17b]. Familywise [WX15]. Famoye [Ano23]. Far [Koz12, CMS17]. Farlie [Gup16, PV16, TB12, YYL17]. Farlie-Gumbel-Morgenstern [TB12]. Fast [Afs13, SSS14a, Lu16c]. Fast- [SSS14a]. Faster [TL15b]. Fault [Ber12]. Fay [HEHM11]. FBSDEs [CL10c]. FDR [GW15a]. feature [HSW17]. features [Lu16c, Ram17]. feedback [GMR16, ZG19a]. Feller [MEG15, NMAA19]. FEXP [Nar13]. Feynman [BCG11]. FFT [FSSW17]. FGM [TB12, Coo19, LSC13]. Fiducial [KZ15, NORV19, TL15a, Bic17, Bic18, ESAKK18, NO17]. Field [GV11, Jia12, Ko16c, PM16a, ZSL17]. Fields [Ko13, Ko14e, LG10, Li13a, RMC11, SMM10, WDM14, Ham19, Ko17b, STD18, Tan18, Wu17d]. Filter [SÖ12, Nas18, TL17]. Filtered [SSS14a, AESO19]. Filtering [Piz13, TL17, Yam17b, Yam18a]. filters [GS19]. Final [EKZ16]. Finance [CL10c, KDK14, MC19]. Financial [Bin14, CGW15, Hua14b, PT10, SDJ15, TT11, Yan14, KYW16, MDG19, Quo17, ZT18]. financing [ZDDH17]. Find [MF10, ZK13b]. Finding [FF18, Kus18, SKK16]. Finite [BR19, CG11a, CP14, DG12b, Dra13, EY10, Fal14, GS15c, HJ14, MS14c, NF11b, NF11a, Par17b, QZY12, SG10a, SG10b, SK13a, SKK10, SS13d, Sub14, TC14, TSS15, TTZ14a, Vas14a, XLL13, Xu16, YUSC14b, Yan14, YMF15, AHH16, ASG17, BRAS17, CS19b, DYXW19, Dut17, zF16, HKH17, HZ17a, HLY16, JN18, KGH19, LWG18a, LQW19, RM16, MKK17, MSK17, OE16, Sen16b, Sen16a, Sen17a, Sen17b, Sen18b, SG17a, SSS16, SPS17, SPY18, Vas16, WYY19, XM17a, YM17, YJ17, ZCC16]. finite-dimensional [YM17, YJ17]. Finite-Population [SKK10]. Finite-Sample [CG11a, Par17b]. finite-state [HZ17a]. Finite-time [HJ14, Yan14, LWG18a]. Finitization [LCG12]. Fire [NB15, ANS17, DTT^{+}16]. First [Aka10, Bar10, Geo17, GSB12, GL15b, GB12, IO13, ITG09, LLV10, Nge11, SP11a, SS10b, SS14a, Yan15, hYhW14, ZY10, ZWZ11, ZZC13, Are16, BP16, BCK17, Cha16, Ery18, FP17b, HZS17, KM19b, KLKK17, SN17, SW19a, WN16b, YW17a, ZL18]. First- [Aka10, ITG09, Yan15]. first-difference [FP17b]. First-Exit-Time [SS10b]. First-Order [Bar10, GSB12, GB12, SP11a, ZY10, ZWZ11, ZZC13, hYhW14, BP16, BCK17, HZS17, KM19b, KLKK17, WN16b, YW17a, ZL18]. Fiscal [HH13]. Fisher [AAR11, Are16, BER16a, Fre14, GCA18, Ham12, JR18a, Kaw13, LPC17, LYR^{+}18, MEG15, NH15a, PL12, SP11b, Voi10, WH11, YSD^{+}16]. FISSARMA [GS13, GS15a]. Fit [BLN13, BR12, Cag12, DTO11, DTO13, HXL13, HME12, HLR11, KM10, LLP10, LL11, Li12a, PS12, SA10, VP10, Voi15, WDZ11, YXC11, ZXD10, AM17a, AP19, BB19a, BRY17, CL15c, Coi17, GL19, GF17, JRF17, Kou16, LM16, MGAf14, Nou17, ÖGGB16, RAB15, She16, TS17f, VPMV16]. fitted [AT17, YH19]. Fitting [ABES18, AAG12b, DG11, Eid12, NIN17, She15c, SLSE19, SA10, TH15, XM17a]. Five [SA10, SMN18]. Fixed [BF12, CLV14, Eid12, GL11, HM12, LR10b, MF10, PR11, QL11b, Ber16b, sK16c, LX17, Liu17c, Liu17a, SBA17, SE17]. Flat [Ham10]. Flatland [Dru17]. flattened [SC19]. Flexibility [MH11]. Flexible [BP15b, BG11d, GP12, KP13, NAA13, BSGS16, FI17, KN18a, PP18, Yil16]. flooding [ZH16]. Flow [CL10b, PK11]. flowgraph [SF16]. Flows [AB11a]. flu [KD16a]. Focused [DZX13]. Fold [Par13b, PY16a]. Folded [NB15]. folding [CJ15]. Foldover [Sty10, EQ17]. Foldovers [Jac11]. Follow [Xu12, YSK15, Liu17c]. follow-up [Liu17c]. following [LCWG16, YL16]. Food [KLS12]. football [JCN17, NC18a]. force [Li17b]. Forecast [MS12b, MS17a]. Forecasting [AP12, BHKHM13, Ho16, Kha11, ZK15a, ARA16, HHC^{+}17, LGY17]. foreign [HZ17b]. forest [BH17c, WHY18]. forests [DLB17]. Form [AAP11, CCH^{+}10, RR14, Wil09, ZI12, ZI14, Ala17, LdM17, MIG17, WSW16]. form-invariant [Ala17]. Forming [Jor13]. Forms [Are14, RLL12, SK12c, BBC18, KK17a, RK18, Zha17b, ZKW16]. Formula [BBC13, BCG11, BM10, Bia12, CCH^{+}10, CBB06, Kal17, MBS17, NHN16, WLZ17]. Formulae [Chi11]. Formulas [Kir12, Lu14, LS15b, Lu15, Xu12, KN18c, Lu16b, Shu19, YJ19]. formulation [Bow17, Hut19]. Fortunato [Mon12]. Fortune [MS14b]. Forward [MS10c, MS14c, ZG17a]. Four [WYZ15, ACO^{+}18, Wan17c, ZLYZ17]. Four-level [WYZ15, ZLYZ17]. four-parameter [ACO^{+}18]. Fourier [FM10b, KL17a, ZG17a]. Fourier-cosine [ZG17a]. Fourth [KV10, Cav17, Wie18]. Fourth-Order [KV10]. Fractal [RMC11]. Fraction [ABI12, ZZ10, CRTdC17, GZ17, She18b, UK16]. Fractional [BKP11, CT11, CLV14, Fon14, GWK19, GIT14, HYK15b, Jac11, Kaw13, LHHC14, MS11b, Tin10, WYZ15, YH11, Zha13, ZLD15, Ben19, CZ19a, FNM19, JZ16, KHY17, LL17e, LZF17, LTC17, RAAJ17, SSC18, SXZ17, Sti17, XML16, Yu19, ZLD16]. Fractionally [GS15a, GS13]. Fractions [RSV11]. Frailty [CRTdC17, DS09, EMI14, HD13a, HS15a, HS15b, HS15c, HD15, Ko10b, sK11, Man11a, Wan10b, WK12, HP16, HB16, HB17a, HP17a, HB17b, HP17b, PBL19, ZSWA19]. frame [Haq17]. Framework [BBB10, CCL^{+}13, KS10a, RL11a, CMS17, DHP^{+}16, GHSZ19, GL19, HA18, HZW17, Kam16, ZLDT14]. Frank [GGL14]. Franklin [JSS16]. Franses [dBCR15]. Fréchet [BSCS11, KJAR13, NA19, AT15, TC17]. Free [BTK10, CC10d, DM10, Hil15, SH11, SBK13, WW13b, BAB^{+}18, Gau17, Mar16, WL14, XQ17a]. Freedom [KMC10]. Frequencies [BDS11, GEM^{+}18]. Frequency [sDsWcCjK14, JLL13, Kaw13, SNN12, GR18b, KAW^{+}18, LG18, LR19, Quo17, WdX17]. Frequentism [Bic12]. Frequentist [BAK13, Far11, WZ13, ZK13b, BN18b, BH17b, Bic18, Sil18]. fresh [PM17]. Friedman [LVH19]. Frisch [Yam17a]. Frontier [Ko11]. Fubini [HZW17]. Full [Dem14, RGL12, Vij11, CYF^{+}18, GPM18, HK19a, KM19a]. Fully [CC10d, KM18]. Function [Abd15, AT11, Arf13, AB11b, BA10, BG11d, Cah12, CS14a, CZ15a, CS15a, CW15b, DL11, Eid14, FG11, GS15a, GP12, HKR15, HKNL15, HX11, Hod14, Hu11b, Hu13, HGZG15, JRM10b, KRH12, LH11a, LWY10, Lia11, Lop11, LLSM14, Mah12b, MCY14, MC11, MKD11, MSN14, NS15a, Neg09, QTD10, QLZ14, QNP12, RASMN14, SG10c, She10a, sS12b, She12b, She13, SS14a, SB11, Sug13, TN10, TT10, Tan10b, TB13, Vin13, WG15, WWH15b, YSHL11, ZZA12, ZFL13, ARJF19, Alv16, BT18a, CRTdC17, CS16a, CX17, CH19a, CCN17, CDS19, Fab16, GH19, GKP16, GT17, HA18, Han19, HW18a, HLY16, Huh16, JN18, JSSO18, JN19, KM19b, KT18, KN17b, LdM17, LA17, LM18a, LX17, LWD^{+}18, MC17, MS17b, MAG18, Mod17, MKY16, NP17b, OTA17, ÖK17, Pak17, PMKF17, PMFK18, Roz18, RK18, SAD19]. function [SYC19, SK16a, SK19a, SQCL19, SR18, Ten18, TS17f, Van16, VFP17a, WL17b, WZK17, WYMZ17, XZ17, Yan17a, YSD^{+}16, You17, ZJZ17]. function-based [BT18a]. Functional [DNL12, ER11, Fu12a, GG10a, GKW14, HL11, HS15d, Hu13, IP13, Kaz11, LLY14, MS12a, Rao10, RGM15, THHL15, WW15a, zXyL13, ZI12, ZI14, ZPZ10, AS17, ALR19, Ayk19, BLT16, HA18, Hen17, HXF18, LM18a, Li17a, LL17d, Nas18, OW19, Qin17, SK16a, SB16a, TT17, XZA18, XML16, XHH17, ZWL16, ZZL16, ZDZ17, ZDS16, ZZL19]. Functional-Coefficient [WW15a, TT17]. Functionals [Lim11, SY10, KS16a, Lii17, NS16a]. functioning [Gol16]. Functions [AE13, ADP10, AYC^{+}12, BCS10, BYDM15, DMP11, EB12, EE15b, Fel10, FP12, Gro11, Gup15, HM12, Hud14, LX10a, LQ13, MFSAS10, Mat11, MSW13, NS15b, PT15, QQL14, RvdM14, Tak10b, Tee10, TS12, TAY11, WDM14, WK13, zXyL13, AE17a, AE17b, BU17, BP17, Bic19, CH19a, Car17, CK17c, CNG17, Che17b, DRAS17, FPZ16, IS17, JZ17a, KM16a, KG17, LWD^{+}18, Mna17, Mur16a, NSS18, NP17a, PA18, PL17, PY16b, SG16a, Sel17b, SV17, SW19c, SSD19, TS19, VS19, VN18, YM18, YL16, ZW18, ZZJ19]. fund [QJWC18]. Fundamental [LX13, MV14]. Further [AER16, DCL12, HWW15, IAA15, Qiu18, YYX14, YH16]. fused [SUKK19]. Future [BSSJ11, PK11, RA11, CCMA18, Tri17]. Futures [HL18b]. Fuzzy [Cui12, MN13b, Tor12, WLY14a, Yen12, HA17, SV17, Sha19, Sym19, THV16].

G [AE23a, AE23c, AE23d, BL19, MBS17, SLOT17, SG17d, YFV19, ZG19a]. G.M [AE23b]. gain [LSY17]. Gait [KS10a]. Gait-Cadence [KS10a]. galaxies [DBC16]. Galton [Fro17b, VYZ14]. Gamma [AE17a, AA15a, AKK09, AG16, AD19, AE23a, BP15b, CC12, Cha15, CXS11, CJ19, CA14, EMI14, GH18, GDCOS13, HS15b, HP16, HME12, IANT13, KCK10, KL12, LF17, Mea15, NBK14, NNN10, PS14, Pop12, She14a, TK14, cTtL14, VRB13, WP18, AE17b, AB17, AMV18, BAF19, Ber18, BPS16b, CBOR18, DeR19, DW19, GWK19, GPP19, HNB19, HB17b, KT17a, KAS17, Kvå17, LdM17, MC17, MMD18, MHF16, NAN17a, PSF16, PN16, SP16a, SS16b, YW17b, Zha17d, ZZ17b, ZRL19a, BIJ17, IABG17]. gamma-beta [NAN17a]. Gamma-Generalized [GDCOS13]. Gamma-generated [TK14]. Gamma-Kumaraswamy [AE23a, GH18]. Gamma-Pareto [AG16, AD19]. gamma-Weibull [SP16a]. Gap [PB17, Alv16, CT16, HHS18, She18b]. Gap-based [PB17]. GARCH} [BG18, BG19c, BK19, ARA16, Che13, GK19b, Iqb13, LS09, LS11, LCZ16, LTCQ17, LWG18a, Mån12, MC19, Rao17b, XLL15a, ZWLI13, ZWL16, ZX16]. GARCH- [ZWL16]. GARCH-in-Mean [ZWLI13]. GARCH-Type [Iqb13, Che13, Rao17b]. gateway [MV16b]. Gauge [GV11, RWY16]. Gauss [SN12]. Gaussian [Ale16, Alg19, AK14c, AAM14, BNZ17, BP18b, Ben19, BR11, CJ17, DMST12, Den17, FR10, Fat13, GMMA19, GDCOS13, GDPR17, Gra17, HD13a, HS15a, HB16, HP17b, HT10, HW15, HH13, HL15, JS13, JD16, Kal17, KK11a, KOP13, KOP14, KT18, KRMG19, KNH17, Lef10, LN17, Lu16a, MR15, MBV17, Mud16, Nak17, NSR^{+}19, NH15b, NPC17, SSD18, TW13, Tan18, THHL15, VEKH15, WK12, Wu17d, XZD19, XLS19, XZ19, ZX18, Zoi09]. Gaussians [DB13]. GEE [CE19, LCZA17b, WP11, XXMZ18, ZX18]. GEE1 [Nak17]. Geiger [SAB15]. gene [AIIK17, ML19]. General [Ben14, Bia12, CC10c, CBOR18, GL11, HN17a, HC16a, IRAB13, IAA15, LWZ13, LPN14, LJ12, LZ14, LR10b, LWR11, LR10c, Mat11, Men19, NP15, OM13, PNK18, RL09, Sho15, SS15b, Tak14, Tia12, VV15, Vas14b, Wan12, WL13, WZ16, WHV16, WW11, Xu12, Yeh12, Yeh13, Yu10, ZB12a, ZK15b, AAK16, BIJ17, Bin17, CK17c, CD17, EE16, FWLL19, Fon17, JN18, Kam16, KP19, KM18, KRMG19, KG18a, LB16, LMT17, Mar16, MIG17, OH19, PGCT19, PP18, QC19, Ren16, SLOT17, SHB17, SP17a, SKK16, Sun17c, SBLC17, Wan17b, WG16a, XZ16, XHH17, YFV19, YZ17b, ZYLZ17, ZB17, ZS17a, ZZ18e]. generalised [KJS19]. Generalization [Aoy10, Bia15, CG15, HW11, JRM10a, MM13, She14b, VRB13, CO16b, DC17, FF16, JS18, KH18, NAB16, NB17, OB16, SSB17a, Smi19]. Generalizations [– Dor11, OCIS12, TY17]. Generalized [AB15, AH13, AMV18, Ala14, AK14b, AN17b, AW16, ANS17, BNE15, BEY18, BBC13, BBT13, Bur09, CLGFB15, Cha10, Cha15, CRG13, CW14, CC10c, CM12, CKD14, CA14, CBB06, CdSBB15, COL15, DDGM13, DJVV17, Deb12, DLL13, EB12, EDAAS12, Fam15, Fed14, FWZ^{+}17a, FV11, FBS10, Fon17, Fu12c, GLZW12, GGL14, GP12, GBD10, GMN13, GDCOS13, GK13, GHY11, GSA14, GC12, GY13, GW19, GL10b, GMSD13, HA16b, HS11, HLR11, JMEM14, JZQ15, JT11, KM10, Kát11a, KS17, KG14b, KG15, KA15, Li12a, LX15, LMZ17, LL14, LG14, LY11b, LXT17, LTS16, MEG15, MZ10b, MS10a, Mea15, MN15, NAB12, NAA13, Oli13, OBPL15, PSP12, PC10b, RP10, Ryc14b, SBL13, SBR14, She12a, SCC10, SP11a, SP11b, SCdSP10, SW19b, Sou17, SSON11, Tan10a, Tan13, TX14, Tor12]. Generalized [TK14, UDC12, VML^{+}10, WZ10, WY12, WXZ13, WSC13, WX15, Wee12, WWW13, WH10, XZNL15, XYH12, YUSC13, YHC14, ZI12, ZI14, ZK13a, ZLN11, ZWZ12, ZWPZ15, Zha15a, ZZLL15, ZS15, Zör14, dPD16, AE17b, Aga17a, Ahm16, AMP17, AR17b, Ala17, AEHQ17, AFL19, ATO19, ARB19, AEB17, AGB17, BK18, BD16b, BT16, Bib16, CY17b, CSY19, CCZR17, CCM19, CM18, Coi17, CARO17, CAPO17, CFK17, CCY17, Das17a, Das17b, DY18, DA16, EP17, FI17, GSM16, GA18, Ghr19, GFG16, GXB18, GDGG17, Gor17, GSLZ16, GYL17, HOAH17, HNB19, HN16, Hom17, HC16b, HR17, HS17, HWLP18, IVVG17, JZ17b, KD17a, Kay16, KJT17, KKKÖ17, KI19, KD16b, KN16a, KK18, KG18b, LP19, LK19b, LM18a, LN17, Li19, LSW17, LZF17, LP17, LWWR18, MBS17, MFJ17, MP16, MC17, Min17, MML17, MBAJ19, Mur16a]. generalized [NHAMG17, NP17a, NA17, NN16a, NAYC17, Nor16, OHKK17, PBL19, PMF17, PV16, Pen17, POS^{+}19, PW16, PE17, RMM17, RMN16, RTT17, SG18a, SBRGD16, Sch16, SK18a, SG17c, SGA19, Sha17, Sha16, SSSM16, Sha18c, SUKK19, SSB17b, Shu18, SSS16, SS16g, The19, TP19, TR16b, VW18, WC16, Wu18a, Yan16, YYS^{+}18, ZXM16, ZZZ19, ZHZA18]. generalized-difference-based [AR17b]. Generalized-order [BEY18]. Generalizing [MM11b]. Generally [CL14, Hua14a, SL13, SHH13, TKW12, Lu18]. Generated [El 15, Pun13, SGPA12, Web12, AHAH17, BKGD17, CBOR18, DTWW17, NA18, RSAN17, SS19, TK14, WN17a, WN17b, WZ19]. Generating [Dem14, EB12, Xia17b, Jon18, MK17, SAD16, SW19c, WYMZ17, ZY17b]. Generation [Fer14]. generators [NO17, SJ19]. Genetic [HL11, LS14, SG11, CSM16, PC16b]. Geo [YFV19]. Geo/G/1 [YFV19]. Geographic [BAK13]. Geometric [BBT13, CG15, CCL14, DAL12, FWT13, GIT14, HW11, JB15, LS09, LD13, Nas12, Par12, RASMN14, Ric13, SI13, Tee10, WXPS12, BZAZ15, Bar19, BNB17, BJ18, BCCR18, BN16, Das17a, FWZ^{+}17a, GDGG17, Ili16, Jas19, JB18, MOC16, NRI17, NAB16, NB17, SBRGD16, TS19, YSS19, ZW17a, ZW17b, ZW18, ZW19b]. Geometry [EPGTH15, JLV13]. Gerber [GH19, WL17b]. get [Tar17]. getting [GDDC19]. GEV [EN17]. geyser [AT17]. Ghosh [AE23a]. Gibbs [JB15, Van17]. Gini [Gho15, KR17, Nag17, Nou17, TJ13]. Given [Can15, DDÚF11, KPS10, Kha12, Kha13, MFSAS10, ST16a, Tar17]. Glivenko [CYY14]. GLM [HTW17]. GLMs [WY16]. Global [UZ19, El 19, LM18b]. Gloves [HC12]. GLR [HD17]. GMC [ZLYZ17]. GMM [CLZ16, GP19]. Godambe [CH17]. GOF [LJ13]. Gomes [AE23b]. Gompertz [AJGG14, HS15c, JT16, LM16, Rao18, Sha19, Smi19]. Gompertz-power [JT16]. Good [Ell11, SSS17b, SP17c]. Goodman [Kvå17]. Goodness [BLN13, BR12, BRY17, Cag12, DTO11, DTO13, HXL13, HME12, HLR11, JRF17, KM10, LLP10, LL11, LM16, MGAf14, Mur16b, ÖGGB16, PS12, RAB15, SA10, VP10, WDZ11, YXC11, ZXD10, AM17a, BB19a, Coi17, GL19, GF17, Kou16, Nou17, She16, VPMV16]. Goodness-of-Fit [BLN13, BR12, Cag12, DTO13, HXL13, HME12, KM10, LLP10, LL11, PS12, VP10, WDZ11, YXC11, BRY17, JRF17, LM16, MGAf14, RAB15, AM17a, BB19a, Coi17, GL19, GF17, Kou16, Nou17, She16, VPMV16]. Goodness-of-link [Mur16b]. governing [BKKP19]. Govindarajulu [NSV12]. GPS [JAA11]. GQL [KJS19]. Gradient [Sei13, EP17, HT16a]. gradient-based [EP17]. graduation [YJ19, YD19a]. Graeco [LHHC14]. Granger [Mån12, Zha17c]. Graph [RP12, LQYZ17]. Graphical [HT10, Kou16, OPK18, JD16, LC17, NPC17]. Graphs [Hud14, Car17, Kre18]. Greater [RAB15]. Greece [SSM16, ZK19]. Greek [ABTR19]. Green [BG11d, Wan16a]. Greenwood [Wan16a]. Grid [WW13a]. Gril [EM13]. Gross [GH13]. Group [CKLC17, HK11, JFS14, LPHC18, NRS^{+}14, Oga11, SMN18, TD14, Zha13, CM17, JBA18, Kam16, Liu17c, LL19c, LCZA17b, LCA17, TK19, Xio16, Yam18b, ZZZ19]. Group-based [NRS^{+}14]. group-specific [TK19]. Grouped [Coo12, SCdSP10, FGL18, YH19, ZMV^{+}17]. Groups [Him10, HS12b, JFS14, ZZP10]. growing [Kre18, YCH17]. Growth [BR13, BC12d, BCD12, Li15, BTC16, She17, XQ17a, Xu17b, Xue17, SS17c]. GSH [PP12]. Guarantee [AM15, Che17a, JT19]. Gumbel [Gup16, PV16, TB12, YYL17, dBSCD16]. Gupta [GK19a]. Guttman [BBB14]. GWMA [AH18a, Hua15, Lu16c]. GWMA-CUSUM [AH18a].

Hajek [DWX17]. Hajnal [Sen14]. Half [CG11b, Gil12, Gui15, Kha13, NP15, aSaCN18, Smi19, Abu18, AHAH17, AE23c, CARO17, CAPO17, OSX^{+}16, POS^{+}19, SN18, TP19, WL17a]. half-Cauchy [CARO17]. Half-Life [CG11b]. Half-Logistic [Gil12, AHAH17, AE23c, OSX^{+}16, TP19, WL17a]. Half-Normal [Kha13, NP15, Smi19, aSaCN18, CAPO17, POS^{+}19, SN18]. Hall [Mar12]. Hamedani [AE23a]. Hand [SCSMN11]. Handling [Par15b]. Hankel [BT13]. Harabasz [Koz12]. harm [LJC18]. Harmonic [PYB^{+}15]. Harmonics [PBY^{+}14]. Harmonizable [OP11]. Harrell [WH17a]. Harris [AA19, PCN15]. Hartley [RS14, SST14]. Hastings [GFS16, Van17]. Having [AM15, IUT12, LCWG16, MKSS17, Mud16]. Hayter [MF16]. Hazard [AJGG14, Arf13, BCS10, CZ15a, Dom11, GK11, HSHW14, IT12, MSN14, RA11, RGI15, RR13, SBK13, Smi19, Sur12, WZY14, WK13, Yil12, Yu10, AMP17, BBHN18, BBA16, CY17b, GP16b, HA18, HK19a, HP16, HB17a, HP17a, HB17b, LXL17, LLD17, LCW^{+}18, MC17, PKM17, PK19, Qiu17, SAD16, SRN17, SV17, VS19, XMP17]. Hazard-Product [Smi19]. Hazards [CL15b, HA14, SS12a, SSM12, WTZS15, Zha15a, ÁF16, BB19a, CW16, DSA16, FDL17, FC18b, GHL17, HP17b, LP17, NSS18, PA18, SS16a, WHC16]. Head [MSSU14, MV16b, Sha18c]. Health [Con12, SS14a, BPV19, BKGD17, SS16e, SS16f]. health-related [BPV19]. healthy [SS16f]. hearing [GEM^{+}18]. Heat [KvdVvZ13, Wan17g]. Heavy [ADO13, BS12a, GW15b, WLY14b, WW15c, Yar13, Ali17, DPZ17, FL19, GGDVP16, GS17, GCW19, HSLQ17, LCZ16, LSZ16, QL18, Roo17, ST16c, WHHW17, YC17]. Heavy-Tailed [ADO13, BS12a, WLY14b, WW15c, GW15b, Ali17, DPZ17, FL19, GGDVP16, GCW19, LCZ16, QL18, Roo17, WHHW17]. Hedging [WQW16, HZ17b, QWWW19]. Heine [KV17]. Hellinger [Gar14, XW16]. Henderson [YJ19, YD19a]. Hermite [Zoi09]. Herriot [HEHM11]. Heston [ZZ19c]. Heterogeneity [Bon14, sK13b, Ko15b, Ko15a, Tsi10, GDPR17]. Heterogeneous [CF12, CF14, sK12a, NO12, SAC13, ZLD11, ZL14a, ZL15, BNB17, BNB18, BNB19, DYL17, FLB16, HD17, HS18b, sK16c, ZZ17b, ZWZ17]. Heteroscedastic [CYL13, CF13, FLX11, FLW12, FH13a, FV11, GdC12, GC14, Hu14b, LY11a, Wan13b, WT15, ZL12, CA18, Nak17, NAML16, PTV18, ROM16, WZK17, WL19, YZP17, ZL16b, ZSS18]. Heteroscedasticity [AAP11, KSBM14b, KSBM14a, SÖ12, Wu11, Çel19, ESAKK18, LLWK18, MGG^{+}17, Sun17a, WHZ17, XK18, Yam18b, ZWP17]. Heteroskedastic [BBC13, Hod14, XTT17]. Heteroskedasticity [Ric17, AHZ17, Arv18, ÇE17a, HJC17]. Heteroskedasticity-robust [Ric17, HJC17]. Heuristic [BM10, DGMC17]. HEWMA [AAJ18]. HEWMA-CUSUM [AAJ18]. Hidden [AR13, GN15, Pie11, VLTP14, Ale16, BKKP19, HJY18, YYW17]. Hierarchical [BAK13, CTT15, EIJK13, GDS^{+}14, Han11, HS13b, Ian12, KM14a, LS15a, MS13, Raz14, RM15, VMVD14, Han17, JSC17, Sha19, You17, Zha17d]. hierarchy [PGCT19]. High [AV15a, CC11, Das16a, HJW19, JLL13, KK14, Kaw13, KVV11, MN13a, MX15, RVD14, RAP13, SLZ11, SL14, SSY18, TN19, WSW11, WSlT15, Yam12, YS12, YHS13, YA10, ZHLZ14, Ahm17, AXLJ17, BCCR18, DY16, FJW17, FLL17, LWZD17, LG18, LTZ19, Mao17, NAML16, PR17, PP17, Quo17, SS17c, WL17c, WX19, XX17, YJSL19, YX18, ZBHW18, ZRL19b, ZZZ19]. High-Breakdown [MX15]. High-Dimension [YA10]. High-Dimensional [KK14, KVV11, SLZ11, WSlT15, Yam12, Das16a, Ahm17, AXLJ17, BCCR18, FJW17, FLL17, LWZD17, LTZ19, NAML16, PR17, PP17, WL17c, XX17, YJSL19, YX18, ZBHW18, ZZZ19]. High-Frequency [JLL13, Kaw13, LG18, Quo17]. High-order [HJW19, ZRL19b]. Higher [HLP17, HWLP18, JP13, LR14, RS10, JSWL17, PR17, SG17g]. higher-dimensional [JSWL17]. Higher-Order [JP13, HLP17, LR14, PR17]. Hilbert [Ko18, Par15a, Rao17a, Yeh15]. Hilbert-space [Rao17a]. Hilbertian [HZS17]. Hill [BB17b]. Histogram [Eid14, KB13, EAE18, Yan17b]. Histogram-valued [KB13]. history [CB17]. Hitting [RCdC14, Geo17]. HIV [SG10c]. HIV/AIDS [SG10c]. HMS [MTM13, Vas16]. HMS/c [Vas16]. hoc [Hu16]. Hodges [KP12a]. Hodrick [Nas18, Yam17b, Yam18a]. Hodrick-Prescott [Nas18]. Hölder [RS16, XML16]. Holm [LMZ17]. holonomic [HT16a]. Homogeneity [GVGP11, MO10, Mar12, SKG12, ZJC10, Ahm17, CGS16, DCZL17, HD17, RPH17, SKG18, SG16e, UU17]. Homogeneous [Ano13b, DDGM13, LS14, MTM13, MS10d, PS14, Vas14a, Vas13, AMP18, Dan18, GPP19, HYS18, PC16b, SJY17, Sym19, TY17, Vas16, YYW17, YYS^{+}18, ZSL17, ZCC16]. Honor [Edi10b]. Horizon [Fal14, LW18b]. Horvitz [Bak17, GÖ12, Sen17a]. Hosmer [YXZ17]. hospital [ZMCM16]. Hotelling [BG19a, DW18, FCSH16, Lee13, SFSH16, TN19]. Household [AZN15]. HPT [Xu13, Xu16, Xu17a]. Huang [BB14b]. Human [KS10a, SU11, PHH18]. hurdle [BW17, zF16, RS19]. Hurst [CLV14, FNM19, MS11b]. Hybrid [AAE17, AA15b, BS12c, MJ15, Tsa10, ZS15, AN17d, BP18a, CSL17, GP16b, GC19, HK19a, Haq17, KTR18, MAA17, NC18c, Par16, SK18a, Sha17, TR16b, WC16, WL17a, WSS17]. hydrocarbon [AA17c]. Hyper [LHHC14, CYF^{+}18, KN16a]. Hyper-Graeco-Latin [LHHC14]. hyper-Poisson [CYF^{+}18, KN16a]. hyperbolic [ARB19, DY18, PW16]. Hypercube [CL15a]. Hypercubes [HM15]. hyperexponential [QJWC18]. Hypergeometric [HP14, TD14, Bar17a, LMT17, TS17c]. hyperparameters [SUS19]. Hyperspectral [CVC^{+}12]. hypoexponential [SKK16]. Hypotheses [BL10, KHM12, LLP10, LL11, MN13b, PS13, PSS^{+}15, SLZ11, Tak14, Tor12, WXPL13, HA17, Kha16, LMZ17, OL19a, SE17]. Hypothesis [AGVGP^{+}11, ER14, KG14c, KE10, Rab13, SY11, SR11, XZD10, Zha15b, CQ17, CZ19c, CM16c, CFA17, DKY16a, DKY16b, Far17a, Far17b, Isl17, JZ17a, KLKK17, KT18, LLY16, LWHB16, MR17, Roz18, Tat16, WBGS19].

I.G. [DN14]. i.i.d [BCH15, CKS17, DP17]. ICA [NAMD14]. Ideas [Voi10]. Identical [Che11, Che17a, GHSZ19, PHH18]. Identically [VT15, BZH19, SHLQ17, TWZ16]. Identifiability [LS16a, CMS17, OGD17]. Identification [BKKP19, CMCCO10, KS10a, Ko12, Ko14b, Ko16a, MMP14, Raï15, SA15a, HZ18a, LL18a, Tch19, TK19, YLG16b, ZL16c, ZP17]. identified [CLZ16]. Identify [Ko10b, sK11, Ko14c, BJ19]. Identifying [BBB14, GGFL11, AIIK17]. identities [Dut17, VZ19]. Identity [WZL10, ZRL19b]. idiosyncratic [ZS17b]. If [Pel17, Pel17]. IFRA [IKL16].

IGFR} [OL19b]. Ignorability [BH17b]. ignorable [RSG17, YD19b]. II [GN15, AA17a, AHAH17, AH17a, BS12c, BS12b, BHLK15, CK19, CS12, CHP14, CB10, CB12, Cra14, DS17d, EDAAS12, FH12, IKL16, Kha13, LJ13, LH13, LJC18, MA14, MK19, NB16, NNI15, Pao10, PGF11, Qom17, RB10, RAB15, Rez17, SNAN17, SK18a, Sha19, SBK13, TTZ14a, Wan12, Wan16c, WCK16]. iid [Xin10, QKP12]. III [Ko15b, RSV11, STRD17, The19, Tin10, WL10, YH11]. ill [AD17a, AD17b, LSL16, MDR17]. ill-posed [AD17a, AD17b, MDR17]. Illness [PDS09]. Illness-Death [PDS09]. Illumina [CXS11]. Illustration [KJD^{+}15]. image [WLS18]. Images [CVC^{+}12]. imbalance [SA17b]. Imbalanced [LLV10]. immediate [LL17c]. Immigration [GS11, Li11, LS11, MO14b, Rah12, CY16, DP18, FP17a, KD16c, Rah16]. immigration-birth-death [KD16c]. Impact [Chi11, Nak19, TK19, ZW13, MCGMS17, YAA^{+}17]. Imperfect [Che14, CDG12, Fre14, PL12, SCC12, Cha16, ZW17b, ZW18]. Implementations [XY11b]. implementing [PMF17]. Implications [NSR^{+}19, MH16b]. Implicit [GJM14, Tia12]. implied [KYW16]. Importance [DSC^{+}14, Kam14, LW18a, SDC^{+}12, ZSD^{+}15, BCB19, MDG19, WGG^{+}19]. Important [WLL13]. Imprecision [New11b]. improper [TL16]. Improve [HT10]. Improved [BHB15, Bur15, CA12, DM11, DP10b, HS13a, HBM16, Kha12, KHC14, KZ19, LPdDTH17, LM14, Lu13a, MMD18, Pet10, PY16b, RJ15, SKK10, SS13d, SM16, SMP18, SS15b, TSS15, TTC17, Wan16a, YUSC14b, YSS19, ZS12, AR17a, CM16b, CLZ18, Das17a, DA16, KJS19, MSB19b, SS16c, SS16h, WER^{+}19, Zhe18]. Improvement [AC15, Gro10, LS11, Li19, NR12, RN13, SCC12, SP17b, Tee14, Wu18a, SNS19]. improves [BPV19]. Improving [LKN15, Kia17, ST16f]. impulse [Roz18, YWX17]. Imputation [SK10a, SSP^{+}12, XLZ^{+}12, KT17b, Qiu17, SGKB17]. Imputations [GR15]. Imputed [WWL13]. Imputing [SHHH09]. in-hospital [ZMCM16]. inaccuracy [KS17, Kun17b, RSV17, TD18]. Inactivity [KN10, AHEA18, KP18b, NK19, Tav16, ZBA17]. INAR} [Djo17, LJW18, ZT17]. Incidence [BHKHM13, EE15b, Ken11, RPY^{+}13, SSD19]. Include [Cah15]. included [PT16]. includes [GGDVP16]. including [DS16a]. Inclusion [GÖ12]. Income [AZN15, RS17, MBAJ19, SLZ18]. incomes [DLH^{+}18]. Incomplete [BBB14, CL15b, GPM18, HF14, HYK15a, KMM12, SB16b, TK13, Tsu14, XLL15b, BH17b, CSL17, CHM16, Ham19, Kha19a, Kha19b, KB17, LG16a, MD17a, MGP17, MMP17, PPMD19]. incongruent [ZG19a]. Inconsistent [YSHL11]. Incorporate [BSB12]. incorporating [VV17, dPD16]. Incorrectly [Koz12]. Increase [SNN12, LL17f]. Increased [LYR^{+}18]. increases [Bic17]. Increasing [CZ15a, Ryc14a, SJS15, YK11, AE17c, AAE^{+}19, DWW18, Had17, LQHZ18, LD18b, LQW19, MKSS17, PKM17, ZQQY17]. Increments [GJ15, Kor11, Wan13a, WLY14b, WHHW17, XML16]. Independence [BD11b, BK09, EPGTH15, Ham13, HKLW12, SLZ11, YSR11, YL15, CHZ17, Mao17, SDS17, WCW17, Zha17b, ZZJ19]. Independent [AASR12, Are14, (Ba14, GJ15, HP14, Kor11, KLM10, LSX14, MP15, MSW13, New11b, NRW15, VT15, ZL15, BN18a, BOT17, BKL19, CS19b, FLB16, GSM16, GG16b, KLMD16, LL17b, LPDD18, LG16b, MMN16, MMP18, MNM19, MHA19, PHH18, Wan17f, YLS17, ZVG19]. Index [BB14a, DTO13, GD12, GMP11, Hua11, Hua12, HL18b, JO10, LIPMF13, LW15, LLY14, LWHW11, LWH12, MS11b, Sim14b, TCN15, VPMF14, WCL11, ZT13, AO18a, AK19, BAL18b, Bal18a, BPV19, Bia17, CA18, FNM19, FTT16, GW19, HA18, HZL16, HZ19, HLR16, HBAJ17, HK19c, JZQ15, LLHZ16, LY16a, LZZL17, LPDD18, LL18b, LL19b, MJH19, MB17a, MSB19b, NS17, Nou17, SK16a, TS17e, UB17, WT18, XHL18b, YLG16a, YCH17, YCZ17, YHC14, ZZ19b, ZF17, ZXF17]. index-based [AO18a]. indexed [Dan18, HYS17, HYS18, Pen17, QWZ14, QWWW19, SY16, SY17b, SY17c, TY17, YSK15]. Indexing [ZLDT14]. India [CM16b, KA17]. Indicators [CW15a]. Indices [Bia12, Bia15, CC15b, Gra10, Gra11, NNA12, Sim14a, WW15c, Yen12, BNZ17, BEY18, CC16b, LWWR18, MBG19, SNWL17, YCAJ18]. indifference [EM18]. indifference-zone [EM18]. indirect [Gro18, Xue17]. Individual [GGFL11, HQW15, sK12a, LH14b, Xu13, DHK17, sK16c, Liu17a, Xu16, XN19]. Indoor [CDZ14]. Induced [CY14, EIJK13]. inducing [FFN^{+}19]. Industrial [OJR15]. ineffectiveness [LJC18]. Inequalities [DB13, DB15, JD16, KG17, Ryc14b, SP15, Spe17, Wan13d, YW11b, FC18b, Fro17a, HA16a, LWWH16, NHV17, RD16, WWW17, YWH17, ZXSZ18]. Inequality [Bea15, FZ10a, HK15, HX11, MM11b, Muk11, Nos14, Por15, Rat13, Sei13, WZL10, XY09, ZFL13, Aga17a, BBG18, Bud16, CLMW19, Che16c, DWX17, DY18, Kre18, MBAJ19, MT19, Nav16, Rao17a, RS17, TK17a, Udo18, Zhe18]. Inequivalent [GFR12]. Inexact [YW13]. Infection [HD15]. Inference [Afe13, AAP11, AARN10, ASNT15, AN17d, BK17a, Bak11, BS12b, BP15b, BPSS12, CHM16, Che10b, CCL^{+}13, CHLP09, CJ19, CCSCME13, DP10a, DP13a, Edi10b, FGP16, FZ10a, FH12, FM13, GN15, GP16b, GCSB10, GL15b, HZS17, HS15b, HL11, Hu14b, Hua12, JMEM14, KPS10, Kha12, Kha13, KSDR15, Kos14, LX15, LS17c, LCW^{+}18, MFPG16, MF14, mNK17, NN11b, Ng12, OM13, PKJ14, Pao10, PT10, PVB13, RB10, Sco11, Sta17, Sto12, SSH10, TK13, TX14, Tsi15, VRC13, WCL10, WXPS12, WW15a, WWT15, WWY17, Wei12, WJH15, Wu16d, zXyL13, ZZ11, ZK13b, ZT17, ZWPZ15, ZS10, AA17a, AHAH17, Ami18, AAK17, Bar17b, BD16b, Bar17c, BS17, Bic17, BPS16b, CSY19, CW16, CCMA18, CY16, EDKAEN17, GZ16, GN17, GFL19, Gun16, Gun17, HD17, HXF18, HZL16, JC18, KAS17, LKK17, LGY17, LXT17]. inference [Liu17a, MSL17, NS18, NC19b, PG18, Qiu17, SG18a, SGGdML17, SNWL17, Sel17b, SK17, Sti17, Tsi18, TH17, WHC16, Wan17b, WLC17, Wan19a, WG16a, WP18, WN16a, XTS16, XHL18b, XFWC19, YCZ17, ZXM16, ZZ17a, ZSS18, ZL18]. Inferences [AMGK13, AMGK15, CMCCO10, HKNL15, Jaf15, PZ15, WSK14, Wan16d, ZX10, LQ19, Xu17b, XT17]. Inferential [Pic15, RAM16, Zha14]. Inferiority [AASR12, Far12, AASRZE17]. Infinite [CTQ11, HB13, LW18b, PZ15, ZLF14, CLW17, CLL17, FLY16, Li17b, Van16, YYWH19, ZL16g]. infinite-time [Li17b]. infinitely [BB19b]. Inflated [FWH^{+}15, Fu12c, HJ14, Mon14, ÖD10, PS11, PBS12, RBN19, YHA10b, ZI14, ZZLL15, BW17, BD16b, BG16a, CYF^{+}18, DDD17, FI17, LMdO18, Mur16a, RC16, SG16c, SSD18, SG16e, TSG19]. Inflation [PS11, YHA10a, MCGMS17]. Influence [APA14, BO17, BSSJ11, CTW10, ER14, EE16, GdC12, Gro11, LSC13, LSC12, MP14, MCY14, PC10a, SE15, TB13, UPA13a, UPA13b, YY14, ZVG19, AH17b, CYF^{+}18, LM18b, Pak17, SBLC17, YY16, YCL17]. Influencing [AA13]. Influential [NAAC16, NIN11, ÖA15, BANC17, Gha19, Sha16]. Information [ARD14, AC15, AAR11, BER16a, CL14, CB14, CC15a, CL15b, CB11, DZX13, Fre14, Gro10, GSSCR12, Ham12, HK15, HF14, KG14a, Kaw13, KS12b, MT14, MEG15, NH15a, NSL14, PL12, QL11a, Sha10b, SV10, SS13b, SS13d, SS13e, SSS14b, SS14b, SS15b, TB12, TSS15, WBH10, WH11, XLL15b, YUSC14a, ASeMJ17, BRY17, CH17, DGMC17, EMH16, EDT18, GCA18, GK19a, HKH17, Haq18, JR18a, KDG^{+}18, KG18b, LYR^{+}18, Miy17, Oga19, Par16, PNK18, PK17, RK19, Sha16, She17, SLF18, SSS17b, SP17b, SMK17, SS16g, SS16i, THV16, Tri17, TH17, YSG17]. Informative [DMST12, Den13, SS12a, YD14, ZZSP14, DDS16, JJ19]. INGARCH [CW19c]. Ingersoll [CCZ15, FNM19]. inherited [MH16a]. Inhibition [Atk12]. Inhomogeneous [Faz11, Sen14, BD16a]. inhomogeneously [DLZ17]. Initial [KvdVvZ13, MT12, Kyr17, Lu16c, Nie17]. injection [YWX17]. injections [TMZ^{+}17]. innovation [Sen18a]. Innovational [BIP^{+}10]. Innovations [BD11b, JKS18, KTBN17, KSDB18, QL18, SNSA17, Sab19]. Innovative [RPM18]. Input [DDGM13, GLP14]. Inputs [XLPL12]. insights [PS17a]. inspected [GFM17]. inspection [BMQ16, NS18, QCK19]. instants [Bos18]. Instrument [MMP14]. Instrumental [RON12, SGM11, VS10, HZ18b]. instruments [LR19]. Insurance [DGM13, DW13, KY10b, Yan14, ANS17, BR19, BG16b, LG16a, WWWX19, YLS17]. insurer [CPW19, FWLL19, LRZ17, WRZ19, ZYH17]. intake [JKL19]. Integer [GSB12, Kac14, MGP11, MSP12, RNB12, ZWZ11, BB17a, BCK17, CC16a, Fon17, JM16, KN18c, LPC17, MPG16, MGP17, NRP16, Quo17, SW19c, YZ18]. Integer-Valued [GSB12, Kac14, MSP12, RNB12, ZWZ11, BB17a, BCK17, LPC17, NRP16, Quo17, YZ18]. integrable [Cos18]. Integral [ABLR15, Par13a, Pri15, Wat11, Aga17a, DZH^{+}19, Kal17, KS16a, LZF17, NS16a, SSC18]. integrals [LdM17, SY17d]. Integrated [GS13, GS15a, JLL13, LCLJ14, SDC^{+}12, AABL17, BCB19, LG18, PKM17, PMFK18, RFM18, WWZC18, ZS17b]. Integrates [SSON11]. Integration [HC12, PMKF17]. Intensities [BCCH14, AESO19, Far17a, GDW18, SI19]. Intensity [CM13a, DJ15, DL11, Bos18, CS19c, SR18]. inter [LSY17, NC18a]. inter-gain [LSY17]. Interacting [Man11b]. Interaction [MSS14a, SAKK15, XFC11, DAC17, DT17, DLH^{+}18, FZ18, ZZ19e]. Interactions [CC14, Hua10, MH15, WLL13, YWZ18]. interactive [LDL18]. Interacts [Ko15a]. interarrival [KV17]. intercept [CE17b, SQCL19, YZ18]. intercepts [PK17]. intercession [PM17]. Intercorrelation [Wu13]. Interdependence [Kim12]. interdependent [FL18a, YL18d]. Interest [CCZ15, KL10, LDY15, QWZ14, XGZ15, BSW15, FNM19, FSSW17, JZ17b, KYW16, Li17b, MCGMS17, SLZ18, XX18]. interests [ZYH17]. Interference [AKK09, BJVV15, FM12, FM14, LS14, Zho10, AAK16, BJVV17, FM17, KKKÖ17]. Intergenerational [SSM16]. Interim [KMC10, LMCH12, LL17f]. intermediate [Gri17, LL17c]. Internal [KMC10]. international [LW18a, MSD19]. Interpolation [BCDLA12, HM15, AA17c, SN17]. Interpretation [MCY11]. Interpreting [BYDM15]. Interrupted [KCS12]. Intersection [RP15, LS17e, Yu19]. Interval [ABI12, CW15b, Far11, FMS15, HD13b, jK15, Ko14d, LCT12, Liu11c, Mah12a, NRW15, NGT13, NAD15, PSS^{+}15, PGF11, PR11, She14b, She15a, TB15, WY12, Wan12, WXZ13, WZ13, WJY15, WTZS15, Wan16c, WWH15b, YHS17, Yam18b, YXG11, YCL14, Yu12, ZB12a, Zho10, ABK17, AK19, AJ18b, BSL^{+}19, BMQ16, FZS19, FC18b, GMMA19, GGtH^{+}19, Ko16b, LK17a, Lin16, Oga16, PMW17, Pou17, PL17, SKK18, Tar17, Uya17, WP18, YX17, ZFD17]. Interval-Censored [FMS15, She14b, She15a, WY12, WTZS15, FZS19, FC18b, WP18]. Intervals [BS12c, BHLK15, BSY12, Bur15, Chv13, DZ12, EDAAS12, HNV13, KG14a, KL10, KV10, KLM10, KXX11, KZ15, LQ13, NN11a, New11a, New11b, Par13b, PGF09, PJL13, QQL14, RA11, Rub10, SP11b, VB12, WCL11, WQ14, WH14, WH10, BAR16, BAB^{+}18, BG19b, Fro17b, GP16a, KWM17, Ker17, KQ17, KPZ16, KX18, MS17a, ML19, Moh19b, PA18, PY16a, PY16b, QK17, Raj18, Sou17, XM17b, ZLD16]. Intervention [AMH14]. Intra [YA13, CS16b, Yad16]. Intra-Class [YA13, Yad16]. intra-cluster [CS16b]. intraclass [Xu17b]. Intrinsic [YA10]. Introduction [VEKH15, KD16c, Sre18]. intuitionistic [HA17]. Invariance [BZ13, Fu12b, MGN16, YL10, Kam16, Kyr17, MH16b]. Invariant [BDKM12, JRM10b, Kon10, LGV13, MR15, Neg09, SEA15, Zen17a, Ala17, HBMAO16a, LL19b, MR16a, PW16, RPM18, VPMV16]. Inventory [BS14, SK16d, Ali17, Kyr16]. Inverse [AD19, CKLS11, DB15, El 15, GDCOS13, GSA14, Gro11, HD13a, HS15a, HSMK13b, HWYW14, JS13, KS15b, KLS10, Lec11, LY16a, Mea15, Pop12, SW14a, WK12, YHS13, Zhu10, Alg19, ANS17, BSS19, CGC17, GMMA19, GDPR17, HB16, HP17a, HP17b, Moh19b, MHF16, PB17, RLSL19, SSSM16, Sha18c, TP19, XLS19, Xio16, XZ16, YGL16, YHW17, Zha17d, ZRL19a]. Inverse-Gamma [Pop12]. Inverse-Probability-Weighting [Lec11]. Inverse-Weibull [HSMK13b]. Inversion [HWM12, WH15a]. Inversions [CN15, CN17c]. Inverted [Gho09, IANT13, KTR18, TR16b]. Invertible [Hu14a]. Investigation [Aca15, Mat14, WWH15b]. Investment [AC11, GW15b, ZYN11, CPW19, GWP19, HW18b, LRZ17, LG16a, MWY18, WWWX19, ZY17a, ZYH17]. investment-consumption-insurance [WWWX19]. Involving [FFA09, Gri15, MS11b, Sto12, Wan13d, AO18c, CS16c, KG17, LdM17, Spe17, VV17, Xia17b]. IPPS [TS12]. Irreconcilability [CM16c]. irregular [QYLL17]. Irregularities [DG11]. IRT [BBG14, MMV12, Mat14]. ISE [JFA12]. ISO [NN15]. Isotonic [SZD12, Hua17b, Ima17]. Isotropic [SMM10, WDM14, Tan18]. issuance [ZWY16]. Issue [Edi10b, Ko15b, Ko17a, MSD19]. Issues [DG12a, Pic15]. Itô [LCLJ14]. Item [Aca15, BBG14, BC12a, MMV12, Gol16, HS18a, SK16d, TK19]. Items [AA13, Bas14, CF14, NN13, GFM17]. Iterated [CL15b, XY10]. Iteration [MF10]. Iterative [DB15, HT10, Sha10a, ÜY13, Udo18, XM17a]. Iteratively [SSS14a]. IV [AG16, GHBM18, AM16, BREA11, Hil15, JBH12, RPY^{+}13]. Izawa [CJ19].

J [AE23c]. Jabbari [LF17]. Jackknife [CNG17, DL19, GMN13, LQ19, WZZ17]. Jackknife-Based [GMN13]. Jackknifed [LY11a, Yil18]. Jackknifing [KCC14]. Jacobian [Sei13]. James [FRC16, ZTSL15]. Jeffreys [OB16, VW17]. Jensen [CLGFB15, Muk11]. Jensen-Logistic [CLGFB15]. Johansen [OL19a]. John [Sen14]. Johnson [AALS12, Fly10]. join [Con16]. Joint [AH17b, DDS16, EB12, FM10b, GEM^{+}18, HDS16, KS10a, Ko14c, Kon19, KSBM14a, MZ11, PGF11, QKP12, RB10, RS19, SDC^{+}12, TIA11, Wu11, ZDZ17, AA17a, AAK17, BAB^{+}18, HNB17, JJ19, Maz19, Moh19a, NHN16, TH17, ZWZwZ17, ZYH17]. joint-adaptive [ZWZwZ17]. Jointly [JTR15]. judgment [DAS16b]. Jump [CCL14, CZ15b, DG11, DW13, Han12b, Kos14, LWD^{+}18, NY11, ÖD10, QWZ14, WG15, YSK15, AG18, BJ19, DYW18, HZ17b, LLLL18, LX19, LSW17, LMC18, QJWC18, SW19a, WRZ19, XDW19, Zho17]. Jump-detection [LWD^{+}18]. Jump-Diffusion [Han12b, Kos14, CZ15b, HZ17b, LLLL18, LMC18, WRZ19, XDW19, Zho17]. jump-diffusions [SW19a]. Jumps [JLL13, Kos14, Bos18, LG18, LZF17, Tra17, ZG17a, ZZ19d].

Kad [SKK16]. Kalman [Piz13, TL17]. Kantorovich [Wan13d, YW11b]. Kaplan [Zha15a]. Karhunen [Ai17, AS18, Hac18]. Karhunen-Loeve [AS18, Hac18]. Kármán [SMM10]. Katz [GL10a]. Kemeny [Hun14]. Kendall [HL17, Sim12]. Kernel [AP12, CLX14, ES11a, FJ11, GBD10, HWM12, HS14, JFA12, KLK14, LMR13, LWY10, Li13a, Li14b, Li17a, LCP^{+}14, Par12, Par15a, Pun13, QZY11, SY11, WH15a, ZHL15, AM16, BN18a, BAKZ18, BLT16, CLZ18, Che17b, DS17e, FZ17a, FST16, GMS18, Iga16, IK18, KOA16, KH16, LL17d, LQ19, Pak17, SCYR17, ST16a, SS16b, Sla17, Ten17, XLS19, YYY17, YJSL19, ZHZA18].

Kernel-Based [ZHL15]. kernel-type [BN18a]. Kernels [Han15, YMF15, Ten18]. Kerridge [KS17]. Khintchine [Car19]. Kidney [HD15]. Kiefer [AD17c]. Kies [KD17b]. Kim [ST16f]. Kin [WR10]. Kin-Cohort [WR10]. Kind [IO13, pZ13, WDLZ16, YSC17b, ZZ17d]. Kinds [LY10]. Knot [DM10]. Knowledge [SA14]. Known [CHP14, GK11, KK11b, MFSAS10, SST14, ASG17, CW19c, SASMF13, YX17]. Koenker [IS11]. Kolmogorov [BP18a, HK13, NY11, ST13]. Kolmogrov [NMAA19]. Koziol [BG11d, Wan16a]. Krätzel [Pri15]. Kriging [JZ16, RGM15]. Kronecker [SAvR12, SÁHLvR17]. Kruskal [Kvå17]. Krylov [YGWQ17]. kth [EWRZ14]. Kuk [SSS17c, TS17a]. Kullback [BRY17, KS15a, Par16, PNK18]. Kumaraswamy [AE23a, AG17, CdSBB15, DMA17, EN17, GT19, GN17, GH18, MDA16, Mit13, WWY17]. Kummer [BREA11, JBH12]. Kurtosis [KM15, RPJ15, Zoi09, vSK16, FPZ16, LTZ19, NHN16, Sta17]. Kuwait [AA17c].

L [MBS17]. Labeled [Ind11]. Labeling [Yao12]. labels [NA18]. Lack [CL15c, Li12a, MS14a, RR14, Rao18]. Lack-of-Fit [Li12a, CL15c]. LAD [TZW16, ZW09, ZL16f]. LAD-Lasso [ZL16f]. lag [Zha17c]. lagged [Bel17]. Lagrange [OCIS12, Par17b, Sen18a]. Lagrangian [IO13, Imo16, LBL^{+}10]. lambda [Har17, GC12]. Lambert [Vin13]. Lamé [SJT17]. LAN [Tra17]. Land [ST17b]. Landscape [SG11]. Language [GV11]. Laplace [AR17c, ASAN12, Ber18, CGW15, Djo17, DA17b, FM10b, HG19, HME12, MH16a, Miy17, NY16, PW11, Sco09, Yeh17, Yil16]. Laplacian [ASNT15, JT11]. Large [Aka10, BOT17, BAK13, Bur15, CY14, CF16, GM14, Hay10, Him10, HP10, Ko11, Ko13, KL11, KB14, Kor11, KS12c, LG10, LJW18, LS17d, LSX14, LSZ16, Mia11, WW13c, dXcY12, XY14, YWS14, ZC17b, ABES18, BZH19, CHZ17, CL17a, CS19b, Dan18, DBC16, DWX17, DZL18, FL18b, GHSZ19, GWZ17, HC16a, HSLQ17, HN17b, KPZ16, LW17a, LW18a, LYG17, Lou19, LSW16, LQ17b, LSL18, NMAA19, NHV17, Pen17, ST16c, SZV16, SY16, SY17c, SHLQ17, Su17b, TY17, WGW19, WYY16, WP19, YHS17, Yam18b, YS16, YYW17, YYS^{+}18, YYWH19, YXZ17, Zha17a]. Large-Deviation [Aka10]. Large-Sample [Bur15, BOT17]. large-scale [DZL18, LW17a, LW18a]. Larger [AM15, HS12b]. Largest [KT15, Kur10, CY17a]. Laspeyres [MB17a]. Lasso [FC18b, HL11, Kim15, Rej16, RZ13, WS11, ZGK17, ZL16f, Das16a, SUKK19, Yam17a, ZZZ19, NAML16, PS18, YL18b, ZX18, CCY17, XM17b]. LASSO-penalized [ZX18]. Lasso-type [Kim15]. last [Cha16, MC17]. Latent [AK14a, BBG14, BC12d, BC14, Cag12, CD15, Fed13, GN10, NAMD14, RSG17, RMP12, She14a, ZLDT14, Ayk19, CRTdC17, DM18, MM19, YCL17]. Latin [CL15a, Gar10, LHHC14, MD17a]. Law [Bal15, CGW15, DCG13, Ko13, KB14, LL11, MRS12, Ree12, WXZ13, XY10, dXcY12, XY14, YWS14, YL10, Zoi09, BZH19, Dan18, DWX17, GHSZ19, Ike17, NMAA19, NHV17, Pen17, SZV16, SY16, TY17, YYS^{+}18, KA17]. Laws [CL15b, HW15, HWW15, NA12, RP15, CHZ17, CL17a, CS19b, HC16a, SY17c, WP19, YYW17, YYWH19, Zha17a]. Layer [Vij11, XZ17]. Layouts [SAKK15]. lead [SB17]. lead-in [SB17]. League [GDDC19]. Learning [Aoy10, CS11, LLV10, SB12, SZL14, Hen17, WYG17, YA17]. Least [AG11, CHP14, Chr10, CKD14, FP12, MA11, Nie17, PYB^{+}15, QZY11, Rah12, SJS15, Tee12, Tsi14, TB13, VSS12, WH15a, WLW15, WSlT15, ZLZ15, ZZL16, AJ18b, CCZR17, LSL16, OU16, OOT19, PCL16, Ten17, Yam18a, YWH17, ZWPB17, vZ17]. Least-squares [Nie17, AJ18b, OU16, PCL16, Ten17, Yam18a, YWH17, vZ17]. Lee [Ano23, LOHW19, DHP^{+}16]. Leffler [GG18, JRM10a, KUWG19, SG16d]. Left [FJ11, LK14, Lia10, PJ10, Par15a, SY11, She10a, She12a, WLF11, WH11, WTZS15, BG16a, HL17, She16, sSW18, VPCG16, YWL17, YWLW18, ZJZ17]. left-censored [BG16a]. Left-Truncated [PJ10, Par15a, She12a, WLF11, WH11, WTZS15, She16, sSW18, YWL17, YWLW18, ZJZ17]. left-truncating [VPCG16]. Left-Truncation [FJ11, SY11, HL17]. leftmost [ZCC16]. Lehmann [KP12a, Sto11, YWO^{+}15]. Leibler [BRY17, KS15a, Par16, PNK18]. Lemeshow [YXZ17]. Lemma [Tor12, LH17b]. Length [CM13a, Fak15, MP15, SH17a, ARJF19, CLZ19, FZ17a, FY17a, JRF17, LK17c, LK17b, Len17, MZ17, ZZ18a]. Length-biased [Fak15, ARJF19, CLZ19, FZ17a, JRF17, MZ17, ZZ18a]. Lengths [Cox15, KV10, Mah12a, RSA19]. Leone [Bay16, Gen17, KI19, RSAN17]. Letter [Aki11, AR17c, HSMK13a, Jon18, Özg19a, SG17d]. Leukemia [HS15c]. Level [CT11, HLL15, Hua10, LM13, MGP14, PC12, RL11b, SM09, Tin10, Tor12, Vij11, WILL13, CBSN16, CS18, CJ15, CZ19a, Fon17, GQC17, Kam19, LOHW19, MA18, OQC14, Pou17, WLL13, WYZ15, Wan17c, Wan17d, WQ18, Yad16, ZLYZ17, ZT17, ZL16e, ZS17a, ZZ18e, ZZ19e]. Level-Crossing [SM09]. Levels [ADWR11, FHS15, AA17c, EIM18, FJS^{+}18, PG16, Su17b]. Leverage [MP10a, RVD14]. Leverages [ÖA15]. Lévy [Car19, DO13, GW15b, Hac18, KUWG19, LYX18, SJ19, ZLZ15, CWW16, GWK19, Gra17, YSC17b]. Lévy-driven [ZLZ15]. Lévy-Khintchine [Car19]. Liénard [YTX14]. Life [ADP10, CG11b, CW15a, DCL12, DW13, Fak15, hFmH14, HS12a, IRB13, JMEM14, Jia12, KPS10, KRB12, KRB13, LM14, MO18, MB15, RG15, RR09, RR13, RM14, RE10, SS10b, SS14a, TGXX12, WXPS12, XFC11, XF12, ZS15, BJU18, BP18a, BT16, CNG17, FZZ16, GXB18, JBA18, KP18b, LG16a, MSR17, SV17, SS16f, VS19, XK18, ZJZ17, ZMN17]. Lifetime [BMLY19, CMCSC12, hFmH14, KRB12, NS15a, NRB14, PDS09, Raq10, SK15a, Wan10b, AE17c, AER16, BIJ17, CM19, Che17a, CAPO17, Ery17, GRSN18, GW19, JC18, KHY18, LSCW17, LM18b, NI16, NY17, PBA18, RCdCN16, SHB17, TPN18, Tav16]. Lifetimes [BP15b, LMM15, RSA13, RCdC14, Zha11, CNL19, GC19, KG18b, LL17a, LMM18, MAA17, PBA17, Rez16, WP18]. Light [Wan13a]. Light-Tailed [Wan13a]. like [LYX18, YSD^{+}16]. Likelihood [Ala14, AS15, Bak11, BP15b, BGG11, BG17, CG11a, CSF12, CM13b, CY16, EY10, EE15b, FLX11, FLW12, FH13a, Fed13, FZ10a, GLZW12, GAK14, Gil12, GFG13, GG13, GFL19, GL10b, HB13, Hod14, HSHW14, Hu11a, Hu11b, Hua11, HLR11, KM14a, KP18a, LvR12, LQ13, Lia10, LW14, LW15, LG14, Lu11, MF14, PSS^{+}15, PGF11, PD10, QL11b, QL15, RB10, SAC13, Sug13, SZ13, Tak14, TX14, Tsi15, TH17, WDZ11, WCL10, WWL13, WSC13, WQ14, WSK14, WWT15, Wan18, WW15c, XTG15, zXyL13, YXC11, YC14, Yu10, hYhW14, YHC14, ZLN11, ZWZ11, ZZ11, ZX10, ZW11, ZWPZ15, ZSY13, dCOSW13, AA17a, AS19, Ale16, Alv16, BT18a, BJC16, CW16, Che16c, CNG17, CW19b, CLW17, CLL17, CW19c, DDP18, DDD17, FJW17, FLL17, GFG16, God16, HXF18, HZ18b, HLR16, IKH19, JJ19, JXW17, JSZL17, KSBN17, KTBN17, KNH17]. likelihood [KT17b, LLHZ16, LCZ16, Li17d, LN17, LD18a, LXT17, LLWK18, LAW^{+}19, LQ19, MZ17, MBS^{+}19, MMD18, NKG16, PC16a, PS18, Par17b, PAvS^{+}19, Phi18, SG16c, She18a, SLF18, SQCL19, SA19b, Tak18, TME19, WHC16, WSW16, Wan16d, Wan17a, WNL18, WZZ17, XJW17, XJW19, XTS16, XX17, XMP17, YZ17a, YCH17, YCZ17, YX18, YH19, YD19b, Zen17b, ZT17, ZBTL17, ZX18, ZC17b, ZZ19d, ZL18, ZX16, dAdLA18]. Likelihood-Based [WQ14, GFL19, Tsi15, ZWPZ15, HZ18b, LLHZ16, LQ19]. likelihood-ratio [DDP18, NKG16]. Likelihoods [Cha14, RSV10]. LIL [FZ10b, ZLF14]. Lilliefors [SS11b]. Limit [AL15, BNZ17, CSC13, FR10, GLS12, HW15, JFA12, Kát11a, KY10a, Li13b, LPN14, MC11, MWM13, MLGZ14, MZ14, NA12, NR12, SSA10, Sun13, Sun15, XP18, ZY10, ZC14, ZZ17a, ZL16g, Arv18, hCXxShY18, CZCM19, DPP17, DeR19, Ike17, Ko16c, Kyr16, LPH16, MGN16, MX19, SN16, SNS19, Wan17f, Wu17c, Wu17d, Xie16, XML16, Zan17, Zha16b, ZLT17]. Limitation [KLS12]. limited [DeR19]. Limiting [HP10, IK18, Roo17, RP12, SA17c, SU17a, Dru17, KM19b, WYY16, ZWPB17]. Limits [FLK12, Han12a, HW14, LVVN11, YXG11, BN18b, CC16b, GW19, LL17b, Pou17, YLS17]. Lin [BB14b, KHY18, PHY19]. Lin-Wong [KHY18]. Lindley [AE23d, AGAM17, AMGK13, AMGK15, ANS17, BP16, EM16, EMHK17, GDCO16, IR19, JJ18, KY17a, MZ10b, MA17, PWLO17, RLSL19, RAM16, SSSM16, Sha18c, SW19b, VW17]. Line [BQDH12, Dji10, ES11a, Eid14, HT13, UC14, EAE18, HD13b, SB16b, VV17]. Linear [AE13, Aga11, AKL15, AA11a, Ala14, Alk10, ASMZ14, AG11, ABC12, BDKM12, BBC13, BC12b, BD11b, BSSJ11, Bru10, Bur09, CCN17, CZ15b, CLX14, Chv13, CBB06, COL15, DG11, DLP15, ES14, Ell11, ER14, FLX11, FLW12, FH13a, FQZ14, Fel10, FZ10a, FHZ12, FMS15, FBS10, FH12, Fu11a, GLZW12, GH13, Gil14, GL11, Gue13, GPÖ14, GW15b, GR15, GE11, Han12b, HMP12, HKLW12, HX11, Hil15, Hu14a, HLY16, HS11, Hua11, HWW15, IP12, JP13, KRT13, jK15, Ko13, KVV11, KE10, KQ11, LLC10, LY10, LY11a, Li12a, LPLW13, LW13a, LX15, LL14, LW14, LR10b, Liu11b, LWR11, LG11, LWHW11, LWH12, LGX12, LX13, LN13b, MGP11, MZ11, Mao15, Mas15, MYM11, MLGZ14, MSN14, MM15, MN15, NKS15, NIN11, OM13, PMA11, PHW11, PYB^{+}15, PC12, Pie11, Por11, PC10b, QL11b, QZY11, QL15, QLZ14]. Linear [QNZ14, RvdM14, RL09, RF12, SBL13, SGR^{+}12, Say14, SN12, SO14a, She12a, She15b, She17, SS10a, ŠS12c, SN14, SS14c, SL15, Tan10a, TT10, Tia12, TX14, TTZ14b, TCN15, UPA13a, ÜY13, UDC12, VML^{+}10, WCL10, WSW11, WL11, WLC11, WS12a, WL13, WWL13, WW13a, WS13, Wan13c, WSC13, WSK14, WWT15, WLW15, Wei12, WJH15, WJ11, WY13, XTG15, XT12, XY12, XY13, YHS13, YC12, YC10, YW11a, YC11a, Yan12, YYX14, YY14, YX15, YS15b, YHC14, ZLN11, ZY13, ZFL13, ZTZ^{+}13, ZX10, Zha15b, Zho09, ZZC13, AE17a, AE17b, AM17b, AD17b, AR17b, AS17, AW16, ALR19, BRZR16, BTC16, BT18b, BCL16, CH19a, CW18, Cha18, CS16c, CS15a, Che16c, CLZ16, CCZR17, CHZ17, Che19b, Ciu17, CCY17, DP18, DTT^{+}16, El 17, EE16, EP17, FZZ16, FJW17, FPCP16, FWW19, GST17, GAS17, GS17, GB18]. linear [Gha19, GSLZ16, GYL17, GX18, GK19c, HA16a, HHS18, HBMAO16a, HBM16, HA16b, HXF18, HZL16, HHC16, HR17, HWY17, HZ19, Isl17, IF18, JN18, JSSO18, JXW17, JQZ18, JHM19, KJT17, KY17b, KOA16, KRMG19, KJ18, KN17b, LZY^{+}17, LM18a, LSL16, LLY16, Li17d, LZZL17, LCG18, Li18, LD18a, Li19, LX19, LY16c, LXT17, LP17, Liu17a, LTGK17a, LTGK17b, LLWK18, LWG18b, LWWR18, LLF19, MBS17, MDR17, MGLX14, MGN16, NNZA16, NNZA18, Nak17, Nak19, NN16a, NNvRS17, NVS19, Nor16, OL19a, Oll17, OOT19, PS18, PDdHNMR17, Par17b, Qin17, Ren16, RTT17, RA16, SNAN17, Sha16, SNW19, SP16b, SE17, Spa19, Sun17b, Sun17a, Sun17c, SJJ19, SÁHLvR17, TZW16, TLT17, TT17, Wan16d, WW17a, WH17b, WW19, WH16c, WB18, Wu16c, WF17, WL17d, Wu17a, Wu17b, WA17, Wu18a, Wu18b, WCF18, Xia17a, XQS19, XTS16, XLS19, XX17, XZA18, Xu16, XFWC19]. linear [Yan17a, YY16, YH16, YSC17a, YCZ17, YWH17, YL18b, YA17, Yil18, YW18, YZ18, Zar19, ZCHR19, Zha16a, Zha17b, ZF17, ZDS16, ZY17b, ZZZ19, ZZL19, dPD16]. Linear-Circular [KS15b]. linear-exponential [JN18]. Linear/Linear [KS15b]. linearities [DY16]. linearity [AO18c]. Linearized [LG11, LGX12, Oga11]. Linearly [CS16a, CX17, CH19a, Shu19, YC17]. Lines [DDGM13, LS17b, Meh17, VV17]. linex [Mur16a, Han19, IP12, LH11a, NP17b, TN10, TT10, Tan10b]. Linguistic [GGL14]. link [Mur16b]. linked [DYW18, HA18, WZJH19]. Links [BBB10]. Liouville [Dai15]. Lipschitz [SJ19]. Liu [Kaç19, Li19, AK13, ANA17, AEA17, CE17b, GB18, Gha19, Gru12, IES^{+}17, LYX11, MKS15, ÖK15, UPA13a, Wu16b, Wu18a, Wu18b, WAA18, XY12, Yil18]. Liu-type [Kaç19, AK13, ANA17, AEA17, CE17b, IES^{+}17, Wu18b, Yil18]. Lives [KM14b]. Liyanage [AE23d]. LM [ZS17b]. LM-type [ZS17b]. LMS [Kir12]. load [Gun16, SNN16, ZB17]. load-sharing [SNN16, ZB17]. Local [CZ15b, DG11, ER11, GD12, Han12b, KL17a, KOA16, LM18a, LZS17, LLQ13, LSC12, LLF19, MFS12, MMP14, Nad10, NE17, Nge11, PC10a, QZY11, RJ15, SBL13, UPA13b, WL11, WW13a, WT15, WLW15, YY14, ZT13, ALR19, BMRSE17, Bic17, Con16, CWW16, GTW17, GGO17, OL19a, QWWW19, WH18, XW16, XZA18, YY16, Yu19, Zan17, ZSL17, ZG17b]. Localization [HT10]. Localized [EZ14]. Locally [CB12, Jaf13, MR15, RD10, ZF18, Mal17, PR17, Wan19b]. Locating [SG11]. Location [ADO13, CT11, CBM14, GR11, Kha12, KV10, MA14, New11a, Pet10, QTD15, SKG12, XFC11, YLS10, AB19, CR16, DL16, EM16, Gau17, GCA18, HBMAO16a, HN17b, IF18, LX19, LL19b, LM18b, MM17, MS17b, NP17b, PDdHNMR17, SS18, TKP17, WH16b, YW17b, ZG19b]. location-invariant [LL19b]. Location-Scale [YLS10, MA14, CR16, DL16, GCA18, HBMAO16a, MM17, PDdHNMR17, TKP17]. Loci [Zho10]. Locking [SM09]. Loève [Hac18, Ai17, AS18]. Log [BS12b, BSY12, DK09, DGV11, Ery12, GL15b, Jeo12, KRB13, LB11, Mah12b, NB15, PD10, SCdSP10, SW19b, XFC11, CAPO17, CBOR18, GFL19, HOAH17, Iga16, LM18b, MFPG16, MFFMA17, MAG18, MOC16, PS18, POS^{+}19, Sha18b, WZ15, vZ17]. log-alpha-power [MFPG16]. Log-Binomial [PD10]. log-Birnbaum [MFFMA17]. Log-Density [Jeo12]. Log-Elliptical [DGV11]. log-gamma-generated [CBOR18]. log-likelihood [PS18]. Log-Lindley [SW19b]. Log-Location-Scale [XFC11, LM18b]. Log-Logistic [BS12b, DK09, Ery12, GL15b, CAPO17, GFL19, HOAH17, POS^{+}19, Sha18b, vZ17]. Log-Normal [DK09, SCdSP10, Iga16]. Log-odds [KRB13]. Log-Odds-Ratio [BSY12]. Log-Rate [LB11]. log-skew-normal [MFPG16]. Log-Transformed [NB15]. Logarithm [CL15b, XY10]. Logarithmic [JA10, KR15, LPN14, CD17, FYLC16]. Loggerhead [Far12]. Logistic [BS12b, CLGFB15, CHH15, DK09, Ery12, FN09, Gil12, GL15b, Gro18, KK11a, KCS12, KDK14, MS11a, MP14, NNI15, TOT15, WDZ11, WJY15, AOÖ17, AHAH17, AEA17, AEN16, AE23c, CCM19, CAPO17, DDP18, DM18, GA18, GFL19, HOAH17, He18, JYZ19, Ker17, KD16b, MTC^{+}19, MN19, MOC16, Nie17, Nou17, OSX^{+}16, POS^{+}19, Sha18b, SC19, TCA^{+}16, The19, TP19, VW18, WXZ16, WL17a, WAA18, Xue17, Zen17b, Zen17a, vSK16, vZ17]. logistic- [TCA^{+}16]. logistic-exponential [MTC^{+}19, vSK16]. Logit [CT10, Jaf13, MKS15, VMVD14, Jun17, MSK18, PGCT19]. Logit- [VMVD14]. logits [DM18]. Loglinear [MP10a]. Lognormal [MSSU14, Pak14, Tsi10, Yu12, GGDVP16, GT17, KAPA19, KVR17, NB16, mNK17]. Lomax [GFG13]. Long [Aga11, Ban13, IS11, KL11, LSC13, WH15b, Ber16b, BS17, BT18b, CCOB16, Gho17, Gra17, HW18a, HS18b, LQ17b, MR17, Quo17, SBLC17, Wan19b]. Long-Memory [Aga11, Ban13, BT18b, HW18a, HS18b, Quo17]. Long-Range [IS11, Ber16b, BS17]. Long-Tailed [KL11, LQ17b]. Long-Term [LSC13, CCOB16, MR17, SBLC17]. longevity [WZJH19]. longevity-linked [WZJH19]. Longitudinal [CGL13, DPB13, FDHH14, GHY11, Hu11a, Ju15, KNK13, Ko10b, sK11, Ko12, Ko14b, Ko14c, LK14, LX15, Liu11b, MO10, MZ11, RTT17, SG14, SLSY14, SD12, TX14, Tsa10, VML^{+}10, WL15, YC14, Zad15, ZZ11, ZX10, ZZSP14, BJ19, DDS16, FJS^{+}18, HL18a, HDS16, HNB17, JJ19, JYZ19, Ko16a, Li17d, LC17, LZS17, LXT17, LL18b, LL18a, LCZA17a, MV16a, Nak17, PM17, RSG17, RS19, SG16b, XXMZ18, ZX18]. longitudinally [GEM^{+}18]. look [EQ17, HT16b, VFP17b]. Lorenz [AB11b, GFA10, LQ19, SJT17]. Loss [ADP10, AYC^{+}12, BA10, CS14a, CBM14, Faz11, Fel10, HX11, HQW15, IP12, Kon10, LH11a, LLSM14, Mah12b, MS10b, Mat11, NS15b, QLZ14, QNP12, TN10, Tak10b, TT10, Tan10b, ZFL13, AO17a, BU17, BP17, Bic19, CS16a, CX17, CH19a, CS19a, GKP16, Han19, HWH18, HLY16, JZTN17, KM16a, LRZ17, MAG18, Mur16a, NP17b, ÖK17, PMKF17, PMFK18, Rej16, TSOT19, XM17b, You17, Zha17d]. losses [YLS17]. lot [BAL18b, Bal18a]. lot-size [BAL18b, Bal18a]. lots [MB19]. Lovell [Yam17a]. Low [YA10, SK16b, WH16a]. low-end [WH16a]. low-order [SK16b]. Lower [AJ11, BCH15, CEK11, Kar17, OQC14, YYS19, FLB16, Lu13a, MRA16, NK19, OQC17, TKP17, Wan16c, Wan17b, ZYLZ17, ZS17a, ZZ18e]. lower-bounded [TKP17]. Lower-order [BCH15, ZYLZ17]. lower-truncated [FLB16, Wan16c, Wan17b]. LQS [Kir12]. LR [Yam12]. LS [Fat13, SG15, Yan12]. LS-estimator [Fat13]. Lundberg [ES11b]. Lung [Con12].

M [AE23c, AE23d, Bur19, CM13a, HZ18a, KC19b, KS18a, LF17, MBS17, Qin17, SG17d, SA19b, SRK11, SU14, Vas14a, XHL18b, XZD10, YGM19, YL18c, Ye19, CDS19, SG18b, BL19, Bur19, CM13a, KC19a, KS18a, SLOT17, SA19b, SRK11, SU14, SK16e, Vas14a, XZD10, Ye19, ZG19a, ZXM16]. m-Associate [SG18b]. m-dependent [CDS19]. M-estimation [HZ18a, Qin17, YGM19]. M.A [Ano23]. M/D/1 [SK16e]. M/E [ZXM16]. M/G/1 [BL19, SLOT17, ZG19a]. M/M/ [Bur19]. M/M/1 [CM13a, SA19b, SRK11, SU14, XZD10]. M/M/c [KS18a]. M/M/k/N/N [Vas14a]. MA [DK17b, MS14c]. MA-ARIMA [DK17b]. Macdonald [NRCG16]. Machado [SG17d]. machine [HWZ19, SA17b]. Machines [NN11b]. Mack [GG16a]. Mahalanobis [DHK17]. Mail [DSSL15]. Main [CT11, CEK11, Hua10, MH15, WLL13, XFC11, DAC17, GPM18, JJ16, JKD17, Su17b, Wan17d, YWZ18]. main-effect [Wan17d]. maintained [Kyr16]. Maintenance [AEH11, Che14, CGSM10, JHP10, QZZN10, SCC10, SCC12, Cha16, CC19, PM18, RFM18, WWZC18]. Majorization [Kvå15, PK12]. makers [Bic19]. making [AAL17, NAN17b]. Malaysian [MC19]. Malliavin [LTC17]. Malta [CE19]. Management [Cui12, PP12, NSR^{+}19, WWZC18]. MANET [MV16b]. Manifold [Kaz11, KN17a]. Mann [HH15, Cuc16, DT17, WZX17]. MANOVA [DO19, HMT^{+}17, YSS18]. Manpower [Gue14, KDGV16]. Manufactured [RVP15]. manufacturing [LS17b]. Many [KHM12, KM14b, Web11]. Map [Web12, YL18c]. MAP/M/1 [YL18c]. Mapping [Che10a, Zho10]. Maps [Sei13]. Marcinkiewicz [HWW15, Ko13, Yan19]. Marginal [CC14, Chi09, Chi12a, FPR17, HM12, Miy10, Pop12, Tan10a, VML^{+}10, VMVD14, WJY15, ZZS13, Djo17, Fan18, Gho17, JM16, Kam16].

Marginalized [EMI14]. Marginals [RNB12, Ber16b, Mud16]. Margins [MFSAS10]. Marine [Far12]. Marked [SY10]. Market [SI13, LG16a, MWY18, MC19, NSR^{+}19, RBK^{+}19]. marketing [YC18]. Markets [JPAB17, SDJ15, NN18]. Markov [ARA16, Ale16, AH18b, AR13, AMP18, AG18, BP15a, BC12a, BG18, BKKP19, BR11, Che16d, DDGM13, DGM13, DMS14, DPP17, Dan18, DT12, FGGL18, FPR17, GL10a, GL15a, Geo17, Gue14, HZ17a, HYS17, HHC^{+}17, HL18b, HJY18, HYS18, Hun14, KCL^{+}16, Kor11, KD16c, LD14, Lim11, LDY15, LW18b, MTM13, MO14a, MR16a, MZ14, ÖD10, Pap13, PV14, Pen17, PAvS^{+}19, PC16b, Pie11, RBB13, SKK18, ST16a, SN12, SFSH16, Sen14, SYW16, SY16, SJY17, SY17b, SY17c, SMS14, SI13, Sym19, Sze17, TN17, TMZ^{+}17, TY17, UU11, Vas14a, Vas16, Vas13, Vas14b, VLTP14, WW10, WQW16, WCZP16, WYL12, XHY16, YSK15, YWS14, YYW17, YYS^{+}18, ZZ17d, ZCC16]. Markov-Correlated [RBB13]. Markov-Dependent [BP15a, AMP18, MZ14]. Markov-Modulated [WW10, YSK15, WQW16]. Markov-switching [AH18b, BG18, FPR17]. Markovian [AGR15, AKK09, AK14c, El 19, HYS18, KKKÖ17, TMZZ19, ZZ17d, ZDDH17]. Markovian-Dependent [AGR15]. Marshall [KCY^{+}19, AJGG14, BSL^{+}19, BNB18, Ber18, CSL17, CL16, GGDVP16, GT19, HS15b, KJAR13, MS13, MTC^{+}19, MML17, MM11b, NH15a, NA17, SP16a, SX18]. Marshall-Olkin [BNB18, Ber18, GT19, MTC^{+}19, SX18]. Martingale [Hua14b, AS19, CZM17, FJW17, SDS17, STD18, Wan17g]. martingale-based [SDS17]. Martingales [RJ15]. Masafumi [AK10, Edi10b]. Masked [CL10a, WY12, XT11, LSCW17]. masking [LS16a]. matched [He18, ZY17b]. matched-pair [ZY17b]. Matches [GG10b]. Matching [Lec11, Oga16, XZ19, ZK13b, CMS17, GGtH^{+}19]. Mathematical [NCO15, CBOR18, Rah18]. mathematics [VW17]. Matrices [APM10, AV15a, GFR12, HP10, Kon10, Kur10, SAC13, SSH10, Ahm17, Bar17c, DLS17, FMPF16, KH19, NKG16, SÁHLvR17, TME19, WYY16, Xie16, ZBHW18]. Matrix [Aoy10, AMZ15, ASNT15, BBC13, CLGFB15, CK10, CA12, CBB06, GLZW12, Ham12, HX11, HKV12, IANT13, KK14, KA18, LX10a, MEG15, NRCG16, Ng12, RL11b, Sch12, SP13, SK12c, SAvR12, Tia12, Tsu14, WS12a, WDM14, YHS13, YW11b, YDC15, ZFL13, Zin12, AH17a, Bud16, DLZ17, FM16, HSW17, HJC17, HSSP16, MBS17, Nag18, SKK16, SS17c, Tak18, TZ17, XM17b, ZKS18, ZKW16]. Matrix-Type [Aoy10]. Matrix-Variate [Zin12, TZ17]. Max [HC10, RP15, DHKM17]. Maxima [GM14, HW15, HL16, HM12, MZ15, FF19, KVR17, RMN16, Tan18, WHHW17, ZZ17b]. Maximal [GG10b, NHV17, Ryc14a, SEA15, YSR11, AEB17, Li17d, NMAA19]. Maximin [RD10, Tan10a]. Maximization [Fly10]. Maximize [CC10b]. Maximizing [Cha14]. Maximum [AS15, BGG11, BR13, CG11a, CB11, DPA10, EY10, GLZW12, Gil12, GFG13, GG13, GGtH^{+}19, GIT14, HB13, HLR11, KO10a, LvR12, LPN14, LG14, Lu14, Lu15, PGF11, PD10, QKP12, SHH13, ZZP10, ZLN11, ZZ19d, AS19, Ale16, AEB17, BT18a, CW19c, DDD17, El 17, EDT18, FC18a, FF18, GFG16, God16, HLR16, Jas16, JXW17, KSBN17, KTBN17, KSDB18, KNH17, LN17, Lu16b, MBS^{+}19, MMD18, NAC19, PC16a, PM16a, Phi18, SG16c, She18a, SA19b, Tak18, TME19, WN16b, Wu17d, XJW19, Zen17b, ZC17b, ZX16, dAdLA18]. maximum-likelihood [DDD17, KSBN17, KNH17, LN17, MMD18, ZC17b]. maximums [Ko17b]. Maxwell [INOC17, HC16b, HWLP18, IABG17]. McDonald [AE23c, OSX^{+}16, dBSCD16]. MCEM [GDS^{+}14]. MCLE [AHKO16]. MCMC [BJ19, Mat14]. McMillan [HYS18, SY16]. McNemar [Lu10, LWZ17]. Mean [ARD14, AS13, AB11b, BC13, BC12c, CW15a, CLZ19, DP10b, DSSL15, Fak15, Far11, FWT13, GP12, GÖ12, GSSCR12, GS15c, HX14, Hua15, HAZ16, IUT12, IRB13, KK14, KK11b, Kha11, KTW10, KRB12, KT11, LH11a, LR14, LWZD17, LM14, MZ11, MS15, Men10b, Men13b, Men13a, MS11c, Nad10, NK19, NVS19, Oga15, Rah12, Raq10, RG15, RR13, SGKB17, SG10a, SG10b, SK13a, SHH13, SS15a, SV10, SKK10, SK12b, SS13b, SK14, SST14, SMMS15, SJ10, SS14b, SS14c, TC14, TSS15, TTC17, TKW12, TJ13, WLY14a, WRZ19, XZD10, YUSC14b, ZWLI13, AO18a, AB17, ASG17, AH18a, AER16, AH17b, BJU18, BS16b, BSH17, BG19b, BP18c, CS16a, CX17, CSB19, CNG17, CN17b, CFA17, CJ17, ESAKK18, EEE16, FWZ^{+}17a, GXB18, GK19a, GGtH^{+}19, HKH17, Haq18, HUK19, HW18a, HR17, JPAB17]. mean [KGH19, KJ17, Koy18, KT18, KR17, KZ19, KA18, LW17b, LRZ17, LTV17, LMM18, MSB19a, RM16, MF16, MSR17, MT16, MSK17, NLRD17, NSSS17, PTV18, PT19, PY16b, QL19, Rah16, Raj18, Rao17a, Rez16, SHB17, SK19a, Sen16b, SG17a, SSS16, SS16c, SKM16b, SHM16, SP16d, SG17e, SPS17, SKS17, SK18b, SSS18, SS16g, SS17c, Tav16, TS17f, TKJ17, UK16, VS19, WXS18, WZZ17, YSS18, YSG17, YKL^{+}17, ZBA17, ZRL19a, ZMV^{+}17, ZRN17]. Mean-Covariance [MZ11]. Mean-Squared [NVS19]. Mean-variance [WRZ19, JPAB17, LRZ17]. Meaning [GIT14]. Means [CHP14, DDGM13, Him10, KG14a, MGM13, Ryc11, ASS18, BS16a, Bie16, CS19a, CS19b, GS17, GP16a, Gri17, Hon17, Hu16, Ima17, KPZ16, Par17a, WXZ16, WBGS19, YYWH19]. Measure [CEK11, DSC^{+}14, Kit15, MP10a, NIN11, RPJ15, SDC^{+}12, Yan15, ZB12a, Aga17a, AMP17, BK18, BP18b, Car19, CM16c, DT17, EP17, KS17, Kun17b, Kvå17, LO17, Min17, RSV17, Sab19, XHH17]. Measured [DPB13, ZHL15, GEM^{+}18]. Measurement [CKD14, CD15, DZ12, DG12b, DMT^{+}15, GG10a, HS11, Hua11, LH11c, MV13, SGM11, WS13, Wan13c, YS15b, ASR^{+}19, AH17b, BRZR16, CK17a, DHKM17, GB18, Gha19, Haj19, KGH19, LLY16, RWY16, Roy16, WP16, WZK17, WF17, XQ17a, YX17, YAA^{+}17, YZFM17, Zar19, ZF17, ZWNS18, cZzXgL18]. Measurements [ITG09, Ko12, RL12, WL14, BAT^{+}18, DFH16, HCW18, JKL19, NLRD17, OMFB17, RAST18, RRRA18]. Measures [ADKZ16, CVC^{+}12, Dom11, EDT18, IP13, Iez12, KM10, KB13, Kvå15, LLSM14, Man17b, MP14, NSL14, New11a, RL12, SE15, SRK11, SAC13, TD18, Tan13, TT11, TAY11, ZSD^{+}15, BBG18, BSGS16, BMR17, BO17, Car17, CM16a, FZ17b, FZ18, HZW17, KM16a, KG18b, Nen18, NNvRS17, NVS19, OMB17, PV16, SLOT17, SLL17, Sta17, Wie18, WZ19, YZFM17, Zar19, ZXM16]. Measuring [CLMW19, YW15, BCB19, HK19c, WER^{+}19]. Mechanics [GV11]. Mechanism [ET13, BJ19, PT19]. mechanisms [SBLC17]. Median [AC15, AARN10, AG11, GB12, Kha12, SS13e, SS15c, VSS12, ZS10, HCZT19, JKJ16, LK17c, LK17b, PM16b, ZLL11, ZMV^{+}17, ZL16f]. Median-Ranked [AARN10]. medians [AJ19, YM18]. Mediators [Chi12a]. Medical [Ko11, CLZ18]. Medicine [SA14]. Medium [MTM13]. medoids [BSS19]. Meier [Zha15a]. members [ST16a]. Memory [Aga11, Ban13, NH15b, RR14, WH15b, BT18b, Gho17, HW18a, HS18b, Quo17, Rao18, Wan19b]. Merging [Vij11]. Meta [CLC^{+}14, DSO12, DMST12, FTF17, ZJLO17]. Meta-Analysis [DSO12, CLC^{+}14, FTF17, ZJLO17]. Meta-Modeling [DMST12]. methane [AA17c]. Method [Aca15, BR13, Bon14, CYE15, Che10a, CLC^{+}14, Cui12, DB15, Dji10, Fat13, FF13, FM13, Gil14, GD11, GL10b, HKLW12, Hu11a, Ko14c, Koz12, LSC12, MSS14a, MKD11, NBK14, ÖD10, SG15, SU11, SSS14a, SCC11, SNN12, Tee10, TN13, TCN15, Wan10a, WLC11, WSC13, Woj11, WYL12, Yu10, YW13, YK11, ZB12a, ZSY13, Zho09, ZT13, AMV18, Bib16, CMN^{+}18, ÇE17a, CS16b, DZH^{+}19, DL19, Fab16, HK19a, HJC17, IBA18, Jon18, JN19, KNC18, KJS19, KT17b, LSL16, LGLX18, LD18a, LWWW16, LH17c, MK17, NB16, NIN17, OW19, PYL17, SI17, SP17c, Slj17, Smi19, SA17c, WC16, WW17a, WG16a, WYY19, WHY18, Xue17, YD19a, YGWQ17, YL18b, ZZ19b, ZG17a, ZX18, ZT18, ZZ19c, ZRL19a]. Methodology [Chi12b, FV11, GMP11, KM14a, Men13b, SWW^{+}12, TEF12, ZS12, Con16, FEP18, YGM19]. Methods [BZAZ15, DTO13, Gue13, KS15a, KLK14, Kim12, LM13, MA11, SC10, SK10a, XGZ11, YS15a, YC14, ZS10, Zhu10, BP18c, BCCR18, CT16, GGtH^{+}19, KJS19, LPdDTH17, LH17a, LC17, LWD^{+}18, MSD19, Nag17, NRN^{+}17, PF19, RPM18, SN17, SB17, SX18, SL16, SS16e, ZB19, ZFZ17, Ano23, AE23a, AE23b, AE23d, MBS17, SG17d, AE23c, Mül12]. Metric [LMR13, SZL14]. Metropolis [GFS16, Van17]. MI [HYK15a]. Microarray [AGVGP^{+}11, DG12a, Kim15]. Microstructure [TL15b, TL17]. Mid [Fel10]. middle [Wan16b]. Midrange [BNE15]. migration [El 17]. migrations [BKKP19]. migratory [El 17]. Mileage [Pul13]. MIMO [BAF19]. min [ZYH18, Sou17]. Minima [GM14, HM12]. Minimal [ASR^{+}19, CSC13, Lin14b, RRRA18, LRC17, MAA17, Oll17]. Minimally [BVJV17]. minimand [LH17c]. Minimax [Mah12b, MM11a, Nor16, TN10, TKP17, Wan13c, Xu13, ZRN17, Bay16, HLY16, HM15, Wu17a, Xu17a, Zin18]. Minimaxity [Fel10]. minimization [CPW19, QWWW19, SH17a]. Minimize [SK10a]. Minimizing [LDY15, Oga15, RZ14, HZ17b]. Minimum [CB11, EIM18, GTW17, Hil15, HLL15, JFS14, JRM10b, Lu14, Lu15, MYM11, MB10, Pak14, RMC11, Wan10a, WY11, BD16a, BM17, CC16c, Fon17, GF17, JT19, LWWW16, Lu16b, Ric17, ZYLZ17, ZL16e, ZS17a, ZZ18e]. Mis [Yu12]. Mis-Specification [Yu12]. Misclassification [BQDH12, WH18, Xio16]. Misclassified [LW13b, TR16a]. MISE [CS15a]. Mises [HK13, Kal19, MV14, Mar11, SS11b]. mismeasured [PCS18]. Missing [AAG12a, BG11a, CW15a, DP10b, Fu12c, GR15, JWZ14, Liu11a, LWH12, MMP14, QL11b, QQL14, SG14, SW14a, SSP^{+}12, SHHH09, SP15, SSH10, SZD12, WP11, WL14, WQ14, Wei12, WJH15, XTG15, YX15, ZZ11, BP18c, BM16, BJC16, DLB17, EMH16, FEP18, zF16, GAS17, GHL17, GX18, HMT^{+}17, HXF18, HSSP16, KT17b, LY16a, LWZD17, LXT17, LAW^{+}19, OU16, Qiu17, RAZ19, RSG17, RKMG19, SNSA17, SLF18, Sla17, Spe17, SSY18, Sun17c, Tak18, VN17, WXZ18, XQS19, XFWC19, YSS18, YGL16, YD19b, ZCHR19, ZL17]. Missingness [SW14a]. Mission [WYL12, YGWQ17]. missions [DLC16]. Misspecification [DTE13, Šir15, BMRSE17]. misspecified [Far17b, GP19, Miy17, Wu16c, Xu16]. Misspecifying [CK10]. Mittag [GG18, JRM10a, KUWG19, SG16d]. Mittag-Leffler [GG18, JRM10a, KUWG19, SG16d]. mix [XXMZ18]. mix-GEE [XXMZ18]. Mixed [AR17a, Ala14, AAJ16, BF12, BC12d, Bru10, DSO12, DZ15, DN14, FQZ14, FM14, GB18, GE11, Li12b, LW13a, LS15a, LR10b, MGF14, MGM13, OQC14, PC12, RL09, SCMR15, SG14, SGR^{+}12, Sha10b, SL15, Tan10a, TTZ14a, VML^{+}10, WL13, YYX14, ZZ19e, Zho09, AM17b, AR17b, AH18a, BRZR16, BTC16, BAL18b, Bal18a, BKGD17, BJVV17, DLB17, EG18, EP17, FPCP16, FFN^{+}19, Fon17, GR18b, HLP17, JYZ19, JDSS16, Kaç19, LC19, Li17d, LLD17, LCG18, LDS18, Li19, LR19, LOHW19, MBV17, MAS16, Nak17, PS18, RSG17, RS19, SG16b, Sha16, ST16f, SE17, Sun17b, Sun17a, TSG19, TLT17, Too17, UB17, WQ18, Wu18a, ZLYZ17]. mixed-chain [UB17]. mixed-effects [Nak17, Sun17b, Sun17a, TLT17]. mixed-frequency [GR18b]. Mixed-level [ZZ19e, Fon17, WQ18]. Mixed-Models [BF12]. Mixing [AR14, CS15a, FR10, GFA10, Guo14, Hu13, HWW15, KB14, LPLW13, Lia10, Lia11, LLY14, RM10, SY11, SWL14, SWLW14, WLF11, ZL12, AD17a, AD17b, hCXxShY18, CM16a, KLMD16, KAW^{+}18, KM16b, LX17, LLM16, Lii17, LSZ16, LQ17b, LSL18, LO17, Noo16, QL19, RD16, VPCG16, WWW17, WdX17, WY16, XYL14, XY19, YWH14, YYY17, YWH17, YM17, Zha16b, ZDW17]. mixingale [Ike17, ZMMX17]. Mixture [AA13, AR13, BC12b, BM11, BPSS12, DS17a, DC10b, DN14, FDHH14, GMB^{+}13, GC14, GDPR17, GRV15, HL13a, KQ11, LZ10, LH13, LZ14, MP12, NF11b, NF11a, NN11b, OGD17, PMA11, PM12, PJ10, Pet10, Pre11, Pri15, QZY12, RS13, RD10, RG15, YOD10, Yao12, ZM13, AIIK17, CF19, CE17b, DYXW19, DCZL17, DA17b, DBG17, FA19, zF16, FGL18, GDCO16, JPAB17, LRC17, LZ18, LL16, MP17, MN18, NM19, OE16, PLDS17, SGKB17, WYY19, Wu16a, WCF18, XZD19, XM17a, YZP17, YH19, ZZ18b, ZZ18c, ZLZ19]. Mixture-Amount [MP12, PM12]. Mixtures [Bal15, CMCCO10, Che10b, EY10, JA10, JAA11, LWL13, Ng12, NFZ11, VLD15, YS15b, CM16a, DA17b, GHBM18, KK17a, MM19, MM17, WY16, WL17d, ZHW17]. ML [GMR16, HYK15a, LG10, SK16e]. MLE [AHKO16, AO17a, CL15b, CT14, WY12, XWJ10, ZRL19a]. MM [DK13]. Mobile [YC18]. mobility [SSM16]. modal [XQS19, YLG16a]. Mode [GSY10, BN17, FST16, GMS18, KY17b, LLM16, SB16a, ZMV^{+}17]. Model [Abd15, AB09, AC11, AH14, AO18b, AVCL13, AT09, AL15, BA10, Bas14, BAT15, Ben14, BSSJ11, BJVV15, Bia15, BCCH14, BAK13, BG11c, BG11d, CCL14, CY14, Che10a, CT10, CXS11, CG11a, CG11b, CW13, Che14, CL15b, CZ15b, CW15a, CZ15c, CF13, Che13, CA12, Chv13, COL15, CD15, DJ15, DTE13, DCG13, Din15, DW13, DZX13, DP13b, EMI14, EY10, Eid12, Eid14, Fal14, Fam15, FLX11, FLW12, Fed14, FHZ12, FU10, FM12, FM14, FH12, Fon14, FN09, FDHH14, Fu12c, GdC12, Gal15, GT15, Gar12, GGLV14, GSB12, GS13, GDSC15a, GN10, GJM14, GE11, HD13a, HS15a, HS15b, HJ14, HX11, HEHM11, Hod14, HA14, HSHW14, Hu11a, Hu11b, Hu14a, HS11, Hua11, Hua12, HS12b, IT12, IS11, Jaf13, Jeo12, JZQ15, JFA12, KC12, KJL11, KPS10, Kha12]. Model [Kha13, KSDR15, Ko11, sK11, Ko14a, Ko14c, Ko15b, KT15, KCS12, Kos14, KSBM14b, KSBM14a, KG14c, KE10, KS12b, KY10b, LS09, LPS11, LS14, LY11a, LH11b, LvR12, LS15a, LX15, LD14, LDY15, LLY14, LWR11, LX13, LLQ13, Lop11, LS17e, LB11, LSC13, Lu13b, MKS15, MSSU14, Man11a, MP12, MZ11, Mao15, MBV17, MFBG13, Mat14, MJ15, MGM13, MF14, MYM11, MLGZ14, MS12b, NN15, ÖA15, PM12, PZ15, PP10, Par13b, PHW11, PMC13, PD10, Pet10, PSP12, Pop12, PDS09, QTD15, QZY12, RA11, Raz14, Rej16, RZ13, RG15, RGL12, RM15, RB11, SS12a, Say14, SN12, SA12, SGM11, Sha10b, She14a, She12a, She12b, ST13, She15b, She15c, SCC10, SCC12, SW14b, Sho15, SCdSP10, SK12b, SSS17a, ŠS12c, SE13, SSH10, SSM12, Tak14, TOT15, TWH10, Tia12, Tra11, TH15, ÜY13, VBAR15, VBATK11]. Model [WW10, WSW11, WLC11, WK12, WWL13, WZ13, Wan13c, WS14, WSK14, WJY15, WTZS15, Woj11, WJ11, XLPL12, XLL13, XGZ15, XYH12, XLZ^{+}12, XY13, XLL15b, YSK15, YC18, YXC11, YC11a, Yan12, Yan14, YX15, YDL14, YWO^{+}15, ZI12, ZI14, ZSK12, ZCHR19, ZL12, ZWLI13, ZFL13, ZW11, ZZS13, ZWPZ15, Zha15a, ZZLL15, Zho09, ZL14b, ZZC13, ACO^{+}18, ASG17, AXLJ17, AR17b, AHAH17, AS17, ARA16, Alg19, Ali17, AAL17, ÁF16, AHP19, AJ18b, AW16, AEA17, AIIK17, AO18c, AE23b, BRZR16, BW17, BTC16, BSL^{+}19, BP16, BBHN18, BPS16a, BD16b, BR19, Bel17, BB19a, Ben19, BB17a, BKGD17, Bic18, BCL16, BD17a, BG16a, CYF^{+}18, CY17b, CSY19, CS18, CSC17, CW18, Cha18, CS17, CSB19, CS16c, CW16, Che16c, CKS17, CXC19, Che16d, Ciu17, CCOB16, DK17b, DLH^{+}18, DZL18, DSA16, DTT^{+}16, DBG17]. model [DM18, DR17, El 17, EKZ16, EMH16, Ery18, FNM19, FY17a, FJW17, FSSW17, FZS19, Fan18, FP17a, FI17, FWZ^{+}17b, FC18b, FG17, FP17b, zF16, FS17, FL17, FL18b, FL19, GWL16, GH19, Gen17, GCW19, GP16b, GDPR17, GDDC19, GOLM19, GFL19, GK19b, GMS18, GHL17, GWZ17, GDW18, Haj19, HK19a, HP16, HB16, HBM16, HME17, Ho16, HWH18, HXF18, HZ18a, HZL16, HHC^{+}17, HWY17, HL18b, HJY18, HJT19, HZ19, Ili16, JSSO18, JZ17b, JZTN17, JM16, JKS18, KWM17, KYW16, KM19b, KJS19, KLJP16, sK16c, Ko16a, KCY^{+}19, KSDB18, KRMG19, KH18, KH19, KNH17, KS18b, KN17b, KD16c, LHS19, LGY17, LXL17, LRZ17, Li17b, Li17c, LSY17, LZZL17, Li18, LJW18, Li19, LX19, LH17c, LMC18, LWZ18, LY16c, LG16b, LGG16, LXT17, LS17c, LY17, LYG17, LTGK17a, LL18b, LWG18a, LL18a, LWG18b, Lou19, LMdO18]. model [LCW^{+}18, LSL19, MJH19, MSK18, MS17a, MBS^{+}19, Mao18, MST19, MA18, MC19, NSS18, Nak19, NRI17, NNvRS17, NVS19, NS19, NC18c, PA18, PG18, PC16a, PMW17, PS18, PGCT19, PY16a, Par17b, PKM17, PMKF17, PMFK18, POS^{+}19, PCS18, PK17, Qiu17, Quo17, RMM17, Rao17b, RFM18, RSG17, Ren16, Rez16, Roy16, SNSA17, SMSF19, SK16a, SG16b, SS16a, Sch17, SSD18, Sen18b, SSSM16, Sha18c, She16, sSW18, SX18, SQCL19, SK16d, ST16d, ST16e, ST16f, SP16b, ST16g, SG17f, SG17g, SG17h, ST17b, SSSK18, SSS19b, SS17c, SSS17c, SJJ19, Sun19, SNN16, SBLC17, TPN18, TMZ^{+}17, TZW16, TS16, TS17a, TS17b, TS17c, TS17d, TT17, TE18a, Wan16a, WL17c, WRZ19, WWWX19, WZJH19, WP18, Wu16a, Wu16c, Wu17a, WSS17, Wu17b, WA17, Wu18a, Wu18b, WAA18, XZD19, XQS19, XX18, XLS19, XHL18a, hXZ17, XZ17, XX17, XQ17a]. model [XTT17, XHL18b, Xu16, XDW19, YX17, YZP17, YH16, YW17a, YCH17, YWH17, YJSL19, YM17, Yil18, YCL17, ZVG19, Zen17a, ZMCM16, ZL16b, ZWL16, ZZL16, Zha17d, ZW17a, ZW17b, ZW18, ZX18, ZT18, ZZ19c, ZW19b, ZL16c, ZWY16, ZL16d, ZWP17, ZWPB17, ZXF17, ZDW17, ZP17, ZDS16, Zho17, ZS17b, ZLZ19, dBSCD16]. Model-Based [Fon14, PMC13, AH14, MBV17, SSS17a]. modeled [CG16]. Modeling [AA13, Ani11, AEB17, BGT10, BG11d, BCD12, CE19, CCL14, DMST12, Den17, DTE13, FPCP16, GGL14, GMN13, GL15b, GRV15, Ham10, HD13a, HB16, HB17a, Har17, Jia12, KLK14, KG14b, LZY^{+}17, LD13, LS16b, LSC13, MV13, Mon14, OD19, PP12, Pap13, PK11, Pul13, RL11a, RCdC14, SBLC17, TB15, VLTP14, WL15, WS12b, ZSS15, Abu18, AD19, ABTR19, Ayk19, BB19b, BBS16, DHP^{+}16, DA17b, DDS16, GEM^{+}18, Gen19, HB17b, HDS16, JJ19, LW18a, LC17, LM18b, MSD19, NC19a, PSF16, PC16b, RPM18, Roo18, SUKK19, SO17, YAA^{+}17]. Modelling [ADO13, AP19, BF10, Bin14, DPB13, FGTS12, Ian14, LLQ13, WL14, Zie13]. Models [AA13, AB11a, AK14a, ADP10, Aka10, Ala14, AR13, Alk10, AKTS13, ASNT15, ADWR11, Are14, AG11, Atk12, AMH14, BBG14, BLN13, BH12, BP15b, BIP^{+}10, BBC13, BC12b, BF12, Bea13, BM11, BR13, BC12d, BC14, BL13, BPSS12, Bos15, BK09, Bru10, Cag12, CNC14, CO10, CY10, CTW10, CL10c, CSF12, CHLP09, CYL13, Chi09, Chi12a, Cho10, CBB06, CDG12, CR13, DGM13, DS09, DTO13, DSO12, DK13, DMST12, DZ15, DLP15, DI10, DDM14, EIJK13, El 15, ER14, FH13a, FQZ14, Fed13, FMS15, FU10, FBS10, GG10a, GLZW12, GC14, GDS^{+}14, GS15a, GP12, GLP14, GdSC15b, Gue14, GPÖ14, GL10b, GB12, GBC12, GMSD13, Ham12, HS15c, HD15, Han12b, HT10, HMP12, HKLW12, HF14, Hu14b, HQW15, Ian12, Iqb13, Jia12, KYS13, KR14, KV10, Ko10b, KVV11, KZW09]. Models [KS12c, KP12b, Lan14, LR14, LLP10, LL11, LLC10, LBL^{+}10, LZ10, Li12b, Li12a, Li13a, LW13a, Li14a, LS15a, Li15, LW14, LW15, LZ14, LR10b, Liu11b, LWHW11, LWH12, LSZ13, LSC12, MS14a, MZ11, MP14, MP10a, MMV12, MS12a, MT12, MG15, MM15, MN15, NB15, NA12, Nar13, Nas12, NSL14, NN11b, NV15, Oli13, Otr15, ÖD10, OM13, PMA11, PC12, Pic15, Pie11, PW11, Por11, Pri15, PC10b, QL11b, QL15, QNZ14, RSV10, Raï15, RS13, RON12, Ree12, RM15, RL09, RAG12, RF12, RE10, SBL13, SGR^{+}12, SE15, SO14a, SSS14a, ST13, She15a, SS10a, SD12, SZD12, SL15, SS12e, Tan10a, TW14, THHL15, TX14, TT11, UU11, UPA13a, UM09, UDC12, VRC13, VML^{+}10, VMVD14, Vij11, WDZ11, WCL10, Wan10b, WS11, WL11, WLC11, WS12a, Wan13b, WL13, WW13a, WW13c, WSC13, WZY14]. Models [WW15a, WWT15, WT15, WLW15, WSlT15, Wei12, WJH15, Wu11, Wu13, XWJ10, XLL15a, XTG15, XF12, XT12, XY12, YC12, YLS10, YHA10b, YW11a, YYX14, YWLY11, Yu10, hYhW14, YHC14, ZK13a, ZLN11, ZTZ^{+}13, ZY14, ZK15a, ZX10, ZZ10, ZW09, ZS10, ZLF14, ZB12b, AM17b, AV18, Ale16, AH18b, AAO17, Ami18, ANA17, AN17c, AG17, BRZR16, BKM18, BBA16, BMLY19, Bib16, BG18, BG19c, Bic18, BM16, BG16b, CZCM19, CRTdC17, CE19, CR16, CCCT16, Cha16, CK17b, CK19, CLZ16, CN17a, CZ19c, Con16, CFK17, CCY17, DMM16, DYXW19, DK17b, DCZL17, Djo17, DA16, Dvo17, EE16, EP17, FTT16, FZ18, FPCP16, FFN^{+}19, FM17, GST17, GB18, Gha19, GP19, GR18b, GSLZ16, GYL17, GX18, HTW17, HHS18, HB17a, HP17a, HB17b, HP17b, HH19, HAP17, HZ17b, HHC16, HR17, HZ18b, HJT19, HNB17, HS18b]. models [Isl17, IKH19, JXW17, JQZ18, JYZ19, JDSS16, JHM19, JD16, KM18, KS17, Ker17, KLKK17, LPC17, LLHZ16, LY16a, LLY16, LCZ16, LRC17, LTCQ17, Li17d, LLD17, LLLL18, LCG18, LD18a, LS16a, LZS17, LP17, Liu17a, LTGK17b, LLWK18, Liu19, LS17e, LAW^{+}19, LWWR18, LLF19, LQYZ17, MBS17, MMN16, MMP18, MHA19, MN19, MB17a, MFPG16, MFFMA17, Maz19, MAS16, Mer16, MDC16, MGLX14, MGN16, MN18, Miy17, MKY16, Mur16b, Nak17, NSR^{+}19, NN16a, NM19, NAML16, Nor16, NPC17, OH16, OE16, OAOQ16, OB16, Pan19, PBL19, PDdHNMR17, PM18, PNPB19, Phi18, QJKC17, RS19, ROM16, RTT17, RA16, SN17, Sab19, SAD16, SDS17, Say17, SG16d, SM17, SF16, Sha16, She17, SNW19, SLSE19, Slj17, SE17, Su17b, Sun17b, Sun17a, Sun17c, SÁHLvR17, Tak18, Tat16, TP19, TLT17, TK17b, Tu17, TE18b, VN17]. models [WZ16, WQW16, Wan16d, WH17b, Wan18, WX19, WYY19, WL17d, WXZ18, WCF18, XJW17, XJW19, XTS16, XXMZ18, XM17a, Xu17b, XT17, XFWC19, XKW19, Xue17, Yan16, Yan18, YY16, YLG16a, YLG16b, YSC17a, YCZ17, YYW17, YL18b, YA17, YWLZ17, YH19, YZ18, Zar19, ZCHR19, ZZ19b, ZF18, Zha17e, ZF17, ZZ18b, ZQZY18, ZZ18c, ZSWA19, ZL16c, ZSS18, ZL16f, ZL18, cZzXgL18, ZY19, ZZZ19, ZX16, ZZL19, dPD16]. Moderate [CGW15, CL15b, DW17, FS17, JY16, Kon15a, Li14b, MG19, MGLX14, Gri17, LJW18, LS17d, XP18]. Modes [Miy10]. Modification [AT11, CW14, Nak17, Yam17b, YD19a]. Modified [AW10, AK13, APA14, BP15a, DPP15a, FN18, FP17b, Gar10, HK13, Hua14b, ICBG18, IKH19, JR18b, KMS13, KPZ16, KR15, LY11a, LL17e, Mar12, Mur13, RGV16, SS11b, She12b, Šir14, UPA13b, ZB19, ZSK12, Alm18, BN16, CE17b, CTAL17, EEE16, GG16a, HP17a, Har17, INOC17, JK19, KAJH18, LS17d, RM16, Mur16a, NNZA16, NNZA18, OPK17, ST16e, TC17, YXZ17]. modifying [PMFK18]. modular [KAR16]. Modulated [WW10, YSK15, Man18, WQW16]. modules [FY17b]. Modulus [BB12, Hu17, NE17]. Moment [Afe13, CL17b, DK13, EB12, FM13, Fu11b, Guo14, HWYW14, HL15, LS09, LCG12, LL12, LP18, Oga11, Sto12, WTZ15, pZ13, Bib16, CZ19c, CLW17, CLL17, DP17, DWM^{+}17, DTWW17, DA16, Fab16, GGO17, HLR16, JYX16, JY16, Kap18, LH17b, LSYL17, LL19b, MWW18, ST17a, SW19c, WHV16, WYMZ17, Wie18, WV16, XTL^{+}17, Xie16, YHW17, YQC17, ZRL19a]. Moment-Based [Afe13, Wie18]. moment-generating [WYMZ17]. moment-type [LL19b]. Moments [AKK09, AK14c, AR13, BS12b, BZAZ15, BM11, ÇG18, CdSBB15, Gil14, GIT14, HS12a, KO10a, KRH12, KK17a, SMS14, SM14, Wu13, Xu12, ZKW16, AGAM17, AMV18, Cav17, Dut17, HWLP18, KI19, Mna17, NC18b, NA17, RMN16, Shu19, SG17g, SA17c, ZRL19b]. Moments-Based [KRH12]. MOMW [DPP15b]. Monge [Sel17a]. monitor [SNAN17]. Monitoring [ARD14, ASMZ14, CTQ11, CC11, Cox15, FNDM12, GS17, HD13b, Hua15, KTW10, LTCQ17, MCY10, MN13a, Muk10, NGT13, NKS15, NNI15, QTD15, SHH13, SN14, TKW12, XGZ11, ZARH14, ZM13, AO18a, AR17a, AH18a, CG16, FKNR17, FK19, Haq18, HUK19, HK19b, JBA18, Lu16c, MATC18, PBA17, PBA18, ZSJ16, vZvdM19]. monomodality [PFD19]. monomode [PFD19]. Monotone [SSH17, Tsu14, YW13, BG11a, HSSP16, KNC18, KN17b, PL17, YSS18, ZZ18b]. monotonic [SY17a, Zen17a]. Monro [AD17b, GFS16, RD16]. Monte [Chi12a, Hol15, OOT19, ÖD10, XM17a, Zha17c]. Monte-Carlo [Chi12a]. Moore [YHS13]. Morgenstern [Gup16, PV16, SM16, TB12, YYL17]. Mortality [ABTR19, Con10, AHP19, Ayk19, CE19, DHP^{+}16, DYW18, WZJH19, WER^{+}19, ZK19, ZMCM16]. mortality-linked [DYW18]. Mortem [WS12b]. Most [CB12, cTtL14, BMRSE17, SSS18]. Motion [GV11, HW11, Kaw13, LS12, Lu14, LS15b, Lu15, MS11b, PW11, SY14a, SY14b, Zha13, ZZ15, ZLD15, Ai17, AS18, Dai16, GWK19, HAP17, Hu17, LL17e, LTC17, Lu16b, RS16, SXZ17, Sti17, Yu19, ZLD16]. Motions [CLV14, YSC17b]. motivated [GRSN18]. Moving [CTQ11, Hua14a, Ju15, MZ10a, Men10b, Men13a, MS10c, MS14c, PSP12, SA12, SL13, SHH13, SP11a, Sun11, TKW12, WW15a, XLPL12, pZ13, ZLZ15, AP17, AAJ16, BP18b, DTWW17, Ko17b, LZY^{+}17, Lu18, STD18, WT18, WN16b, YKL^{+}17, ZZ17d]. Moving-Average [pZ13, DTWW17, Ko17b]. MSE [NE17, NO12, NO18, Xu13, Xu17a, XN19]. Multi [Arn11, BCH12, BD11a, BO14, CM13b, DSC^{+}14, DD13, EI11, Gue10, Gue13, sK12a, sK13b, Ko15b, Ko15a, LH14b, MGP14, MBN14, MZ14, NN11b, Oga11, SDC^{+}12, SG11, UU11, WW13c, WZT15, ZBHW18, ZH16, AHH16, Ahm17, BS19, CBSN16, DLC16, JSS16, KS19, LAW^{+}19, Lou19, LSL19, MSEN18, PNPB19, SH17b, SSS17b, SKM17, SS16i, Su17b, Tat16, WYG17, XZ17, ZZ19b, ZZ19e]. multi-auxiliary [AHH16, SSS17b, SKM17, SS16i]. Multi-Casting [BCH12].

Multi-Character [Arn11]. Multi-Class [Gue10, Gue13]. multi-dividing [WYG17]. Multi-Factor [DD13, ZH16]. Multi-Group [Oga11]. multi-index [ZZ19b]. multi-layer [XZ17]. Multi-Level [MGP14, CBSN16]. multi-objective [SH17b]. Multi-Optima [SG11]. multi-phased [DLC16]. Multi-Population [SG11]. multi-proportions [JSS16]. Multi-Regional [Ko15a, sK12a, sK13b, Ko15b]. Multi-Response [BO14, MBN14, MSEN18]. Multi-Risk [WW13c, LSL19]. Multi-Sample [CM13b, ZBHW18, Ahm17]. Multi-Stage [NN11b, UU11, PNPB19]. Multi-State [DSC^{+}14, EI11, SDC^{+}12, LH14b, KS19, LAW^{+}19, Tat16]. Multi-Treatment [BD11a, BS19]. Multi-type [MZ14]. Multi-Unit [WZT15]. multi-way [Su17b]. multiauxiliary [SKM16b]. Multicollinearity [MSS14b, HJC17]. Multicomponent [RAK15, DMA17, RAA17]. Multidimensional [BBG14, MM11a, CN17b, Yu19]. multifactor [AHP19, ROM16]. Multifractional [Dai15]. Multilevel [AAG12a, MG15, RM15, TME19, LW18a, OH16, ZMCM16]. Multinomial [Far12, Lan14, MP10a, AXLJ17, BG17, CGS16, FGL18, Ghr19, Kar17, MSK18, PGCT19, YJSL19]. Multinomial-Poisson [Lan14]. Multiple [AKL15, ASMZ14, AG11, AGVGP^{+}11, BL10, BB14a, CHLP09, CB10, CB12, Con12, CMCSC12, DHP^{+}16, DC14, DC10a, El 17, El 19, hFmH14, FN18, GH13, GGS13, GW15a, HKLW12, Ima18, JFS14, Kem11, Ko14d, LP10, LWWH14, LS15a, Lin14a, LH14b, MWW14, NNZA16, RVD14, Sha10b, SL14, She15b, Shi10, SYP13, SD12, SM14, TS12, WCL11, Wan19a, WBGS19, XQ17a, XLZ^{+}12, XY12, YCAJ18, Zar19, ZS12, ZW09, Zho10, BAL18b, BL19, CW18, CQ17, Che19a, CKT18, CFK17, DAC17, GEM^{+}18, He17, Ko16b, LMZ17, LS17b, LPHC18, LCWG16, NC19b, PK18, PMKF17, SMSF19, Say17, Sen16b, Sen16a, SY17d, TN17, WX17, Wan18, WH18, XK18, ZBTL17, ZT18, ZWZwZ17]. Multiple-Component [hFmH14]. multiple-outlier [Wan18]. Multiple-response [XQ17a]. Multiple-robot [FN18]. Multiple-Sample [SL14]. Multiple-start [NNZA16]. Multiple-Stream [WCL11]. Multiplicative [DL11, EE11, LR14, Otr15, PBY^{+}14, PYB^{+}15, Yu10, Pop16, SN17, SQCL19, ZZL16]. multiplier [Par17b, Sen18a]. Multiply [EDAAS12, LJ13, HD17]. Multipower [LZF17]. Multiregional [Ko14a, sK16c]. multiresponse [KB17]. Multiscale [XZD19]. Multistage [BL10, PC16b, SAS11, FF18, MSEN18]. Multistate [AEH11, OAOQ16]. Multitype [GDS^{+}14]. multivariable [HW18a]. Multivariate [Aga17a, AV15a, ASMZ14, AKTS13, BQC17, BF12, BC13, BC12d, BBS16, BYDM15, BEZ11, BD17b, CY14, CLC^{+}14, CA12, CM12, CDZ14, DGV11, DLB17, Fam15, Fed13, Fer14, FFA09, FF13, FAS11, GDS^{+}14, GCR12, Gia14, HL13a, HHW14, IP13, Jaf15, KSK15, Kim11, jK15, Ko12, KVV11, Kon10, KS12c, LG10, LZ14, LX13, LWH19, Lop11, MZ10a, MV14, MFFMA17, MCY10, MCY11, MF14, Men10b, MS10c, MS14c, NF11b, NAMD14, NNA12, NK15, PT10, QNZ14, RL11a, RAP13, RLL12, RR09, RL11b, RL12, SRN17, SGD11, SA15a, She15a, Sim14a, Sim14b, SN14, SC18, SSH10, SZ13, TT10, VPMF14, WSC13, XZNL15, Xu12, YS12, Yeh12, Yeh13, Yeh15, Yeh17, ZFL13, ZVS14, ZHZA18, AR17a, AP17, AA17b, AJT16, AH17a, AGA17b, BR17, BS16c, BAKZ18, BT18b, Ber18, BK19]. multivariate [Bud16, CS16a, CR16, Che16a, CN17a, CM18, DY18, DC17, DBA18, DLZ17, DS16b, DS17e, ESAKK18, FPPRG16, GS17, GEM^{+}18, Gra17, GICRR17, Had17, HK19b, HMT^{+}17, HME17, Hu16, Isl17, JPAB17, JYZ19, KP18a, Ko16a, KKS18, KRMG19, KN18c, LK17c, LL17a, Li17a, Lu16a, MC18, MFPG16, MV16a, Maz19, MAS16, MA17, MF16, MN18, MAAI17, Mur16b, NAC19, NKG16, Nav16, NNvRS17, NVS19, OW19, PM18, PM17, RMM17, RJN16, Shu19, SS16e, SÁHLvR17, TWG18, VN17, VPMV16, WH16b, WB18, XQ17a, Xu17a, XO19, ZLT17, ZSWA19, ZRN17]. Murthy [Sen17a]. mutually [LL17b].

N [Vas14a, Vas14a]. nabla [GG18]. Nadaraya [Li16]. naive [TN19]. Nakagami [ÖGGB16]. Narrowest [YXG11]. national [WER^{+}19]. Natural [AMZ15, TN15]. Navigation [Pap13]. NBAFR [PMA14]. NBINAR [RNB12]. ND [WW19]. Near [CM13b, CA14, GLS12, GC12, Spa19, Tch19, KVR17]. Near-collinearity [Spa19]. Near-Exact [CM13b, CA14, GC12]. near-maxima [KVR17]. Near-Records [GLS12]. Nearest [SF12a, TS12, VBATK11]. Nearly [DS17c, DS17d, Lef10, LJW18, QYLL17]. Nearness [AB14, ÜY13, YLX10]. Necessary [BBHN18]. Neck [MSSU14, Sha18c]. needed [Tar17]. NEG [SUKK19]. Negative [ASB13, FWH^{+}15, Gui15, HWYW14, IO13, Imo15, OCIS12, RNB12, Ryc11, SG14, SSON11, Tee12, YCL17, YL15, Zör14, AJ18a, Bar19, CRTdC17, CO16b, CG16, CG17, Coo19, CCOB16, CW19c, DR17, FWZ^{+}17a, IK18, IF18, JM16, NC18b, OMFB17, SG16b, TS17c, WH16c, YHW17]. Negatively [GZSS14, KL11, LWY10, LQ13, LW14, LW15, QL11a, Sun15, WHVH13, WLHW14, WJ11, XY09, XM14, CF16, CL17a, FWW19, GHSZ19, Ko18, MWW18, Men19, NHV17, SZV16, SS17b, SXW17, WH17b, WV16, Wu17c, XTL^{+}17, YQC17, YW18, Zha17a]. Neglected [BC12a, ZW13, QCK19]. negligible [WN16a, ZZSP14]. Neighbor [AJV11, AA11a, AIA14, FM12, VJV14, HH16, IBA18, Oll17]. Neighbor-Balanced [AJV11, AA11a, VJV14]. Neighboring [BJVV15]. Neighbors [VBATK11]. Nested [CL14, CC14, CL15a, KLS12, KE10, KS12b, MGM13, Par13b, YA13, LS17e, MST19, PGCT19, PY16a, Say17, SE17, Su17b, Yad16, fZWD19]. net [YS19]. net-penalized [YS19]. Network [Bar17c, CL10b, GLP14, Gup14, LH14b, SB12, FKNR17, FK19, Gun16, RBK^{+}19, TSG19]. Networks [CCL^{+}13, HTN15, HL11, KS15a, MV13, MS14b, PJL13, SWW^{+}12, SA14, ZA14, HCW18, LD17, LDS18, LQHZ18, LD18b, LQW19, ZQQY17]. Neural [GLP14, ZA14, LDS18, RBK^{+}19]. neuroscience [Rah18]. Neurotoxicology [Raz14]. neutralize [SKM17]. Newsboy [SYP13]. Newsboy-Type [SYP13]. Neyman [TS17a, TS17d, Tor12]. NGINAR [RBN19]. Nikulin [Voi10]. NMAR [KT17b]. NMF [ZLDT14]. NMF-framework [ZLDT14]. No [Mül12]. NOD [CLLW16]. Node [SRT14]. Node-Sampling [SRT14]. Noise [LX10b, PBY^{+}14, PYB^{+}15, QTD10, Rab13, TL15b, TN15, WG15, Zil11, ABK17, Ben19, CS19c, GDW18, LYX18, Mat18, SW19a, TL17, TK17b, ZT17]. noises [Haj19, Pan19, PTV18, SJ19]. Noisy [JLL13, NAMD14, SG11, GGtH^{+}19, LG18]. Nominal [NNI15, HSW17]. Non [AV15a, AASR12, AD10, Atk12, BG11a, BL15, BD11b, BJVV15, Car19, CS15a, CY10, Che11, Che17a, CCMA18, DC14, DDGM13, DY16, DMST12, DO13, Fan15, Fro17b, FZ10b, Gra10, Gui15, GK19c, HH16, HL18a, Han15, Hay10, HMT^{+}17, Hen17, HS15d, HWYW14, IO13, JZ17a, JSZL17, JHP10, KNC18, KY17b, KK11b, KJD^{+}15, Koi10, KL17b, LA17, LR10a, Lem14, LS14, Lu18, Mas15, MMP14, MST19, MM15, MT19, MW10, MSW13, NGT13, NK15, OCIS12, PJ10, Par15b, PS14, PYB^{+}15, Raï15, Rat13, RN13, Say17, SK13a, SR17, She18b, SN16, SKK10, SK10a, SK14, SMPH15, Son16, SSR19, TSS15, TWTG14, THHL15, Tee10, TCN15, UC14, Vas13, VT15, WW13a, WG15, WZ15, XTG15, Xin15, Xu10, XMP17, YWLY11, YW13, ZZ18b, ZZ18a, Zha15b, ZZSP14, AB17, Ahm17, ASG17, AA17c, AASRZE17]. non [AMP18, Arv18, AO18c, AH17b, BTC16, BZH19, BQC17, Bar17a, BAR16, BP18b, BH17b, BS17, BJVV17, Bur17, CSB19, CCZR17, CM16d, Ciu17, CMS17, DAC17, Dan18, DAS16b, DR17, El 17, FWZ^{+}17a, FJS^{+}18, FRC16, FY17b, FL18b, FL19, GZ17, GHSZ19, GS17, GCW19, Gol16, Gor17, GPP19, HK19b, HZW17, HHC16, Hua17a, Hua17b, HJY18, IK18, Isl17, JXW17, KJ17, Kon16, KG18a, KZ19, KN17b, LP19, Li17d, LD18a, LL19b, LMT17, LS17e, LMdO18, MSB19a, Man17b, MHF16, Mur16a, NSR^{+}19, OMFB17, OB16, OTA17, PC16a, PDdHNMR17, PS17b, PC16b, PHH18, PK17, Rah16, RSC16, RSG17, SB16a, SNAN17, SGA19, SJ19, SKM16a, SKM16b, SSB17c, SSS17a, SMK17, SKS17, SKM17, SK18b, SSS18, SN18, SE17, SHLQ17, SSS17c, SNS19, Sun17b, Sym19, Sze17, TWZ16, TS19, Tu17, UK16, WQW16]. non [WP16, WZX17, Wie18, XTS16, XTT17, Yan18, Yan17a, YHW17, YYY17, YYW17, YWH17, YYS^{+}18, YWL17, YSG17, YD19b, Zha16a, ZDW17, ZFZ17]. non-central [LMT17]. Non-commutative [MT19]. Non-Degenerate [FZ10b]. Non-Homogeneous [DDGM13, LS14, GPP19]. non-identically [BZH19, TWZ16]. non-inferiority [AASRZE17]. non-linear [XTS16]. non-Lipschitz [SJ19]. non-negative [IK18]. Non-negligible [ZZSP14]. Non-nested [MST19]. Non-Normal [BD11b, HMT^{+}17]. non-normality [Ahm17]. non-overlapping [Kon16]. Non-parametric [CCMA18, HMT^{+}17, JSZL17, SSR19, ZZ18a, HK19b, HJY18]. Non-Positive [AD10, LL19b]. Non-Regular [Hay10, Koi10, MW10]. Non-Renewing [JHP10]. non-response [SK18b]. non-stationary [FL19]. Non-Uniform [Rat13, SNS19]. Nonadjacent [EB12]. Noncentral [SSON11]. Noncompliance [Chi10]. none [GZ16]. nonexistence [Kha19a]. nonhomogeneous [SY16, ZW19a]. Nonignorability [ZT13]. Nonignorable [SG14, WCGZ12]. Nonlinear [AAM14, CTW10, Li14a, MS12a, VRC13, XWJ10, El 19, Haj19, KLJP16, XJW17, XJW19].

Nonlinearity [DMRO14]. Nonnegative [FWT13, CS19b]. Nonnested [RON12]. Nonnormal [Bur15, HX14]. Nonparametric [ARJF19, AJA14, ABI12, Arf13, BG11c, CC10a, CL10c, CYY14, CS14b, CMCSC12, CCSCME13, DNL12, Fak15, FZS19, Fu12a, GL15a, GFR12, Hu13, Hua12, HCS10, JO10, KS10b, KLK14, KNK13, KHC14, LLP10, LL11, LX11, Li12b, Lop11, Man11a, Mar15, MW10, NBS13, Nge11, OW19, PS12, QTD10, QQL14, RASMN14, Rao10, RA11, RM10, Ryc11, She10a, sS12b, She12b, She13, SS10a, VLD15, VV12, VB12, VB13, WR10, Wan13b, WZY14, zXyL13, XT11, YMF15, ZZA12, ZVS14, ZL14b, ZLWZ10, BT18b, Che16b, SNW19, ZG19b]. Nonresponse [WCGZ12, PM17]. Nonstationary [AP12, CTQ11, SWL14, ZY19]. Nooghabi [LF17]. Norm [MYM11, YW11b, Has19, XML16]. Normal [AT17, AR17c, ASAN12, AVCG10, AVCL13, AGVB10, BG11a, BM15, BC13, BD11b, CHP14, CN15, DK09, EMI14, Far11, GC14, Gra10, Gui15, HKNL15, HYK15a, IUT12, Jaf15, JN19, KSK15, Kha13, Kon10, KLM10, KT11, KA15, LMG13, LVBS10, LX10a, LSSS14, MM13, MH11, Mey13, NAA13, Ng12, NRW15, NGT13, NP15, Peg12, QNP12, RCC12, RN13, SCdSP10, Smi19, SSH10, SI13, Tak14, TT10, TP14, VMVD14, WLC11, WS13, WG14, AM17b, AFL19, ATO19, AJT16, AJ19, AH17a, BK17a, BMFS18, BCP17, Can15, CFA17, CN17c, CAPO17, DeR19, DS17b, DA17b, DZH^{+}19, ESAKK18, HV17, HEB16, HMT^{+}17, Hu16, HS17, Iga16, Ima17, Ima18, JPAB17, KNC18, KK17a, KKS18, LP18, LAF17, MM19, MN17, MM17, MFPG16, MF16, MT16, NAC19, NKG16, NAAC16]. normal [OHKK17, POS^{+}19, RS19, Raz14, RG19, RPH17, RN17, SG18a, aSaCN18, SSB17a, SN18, SC18, Tak18, Tar17, TN16, TKJ17, VLD15, WYY19, Wie18, XQ17b, ZVG19, Zha17d, Zha17b, ZRL19a, ZZJ19, ZG17b, ZKW16, ZRN17, AT17]. normal-inverse-gamma [ZRL19a]. Normal-Laplace [AR17c, ASAN12]. normal/independent [ZVG19]. Normality [AV15a, BDM15, BEKL13, DLL13, HS15d, Lia11, Liu11b, MFS12, NA13, RB11, SY11, SS11b, Xu10, YC12, Yan12, ZL12, Ahm17, Isl17, Kal19, KP18a, KR16, KKS18, KTBN17, LTZ19, NHN16, Par17a, SNW19, VPMV16, WdX17, XZA18, YWL17, ZDW17]. Normalized [GS13, Hui10, hCXxShY18, CS19a, CD17, Kal17, LPH16, Wu17c, Zha16b]. Normalizing [Mae10, GR18b]. Normally [CY10, SYP13, YWLY11]. Norms [Li14a, RS16]. North [LS16b]. Norwegian [NB15]. Note [Bea13, Bea15, Bi15, Cag12, Can15, CS17, CP14, sDsWcCjK14, DS14, Fle13, Fre14, Gro10, HSMK13b, Hol15, HWYW14, Hug09, Lec11, LX10b, Li11, LH13, MB10, PMC10, PT15, RS14, RM10, Rub10, SKG12, SP15, Tsi18, VP10, Voi15, WZ10, Web12, dXcY12, YWH14, YL10, ZJC10, ZTSL15, ZC11, AL16, Arv18, Bar17a, BM17, Bic18, Cle18, DP18, HZ17a, Ike17, JM16, Kam19, KM19b, Kha19a, Li17c, Lu16a, LQYZ17, MH16b, MX19, MSR17, NAB16, PBA18, Sin16, SO17, WK19, Wie18, WB18, WZZ17, Yam17b, YJ17, ZWPB17, ZY17b]. Notes [GY13]. Notions [Dom11, FPPRG16]. Novel [Che10a, DZ14, GCR12, HHC^{+}17, SL18, WCW17, ZZ19b]. np [JBA18, LK17a, LK17b, ZWZwZ17, KM19a]. NPMLE [YSHL11]. NQD [Sun13, TWZ16, WG16b, WP19, YZG17]. NSD [HZZ19, SWZ16]. Nuisance [Gri15, Kak12, LN13a, CM16c]. Null [Lem14, Xu10, Zha15b, MR17]. Number [Cey11, CT13, Dra13, EM13, GLS12, Gov11, GB15, KJD^{+}15, Kon16, MWW14, MZ14, NN13, RZ14, RAB15, TD14, WLY14b, WSK14, ZZP10, Bic17, CN15, CN17c, CCY17, CW19c, DBC16, DWW18, EW16, El 17, FTF17, GHSZ19, GFM17, GYL17, Kon19, LY17, MB17a, SN18, SU17a, Su17b, Van16, ZL16d, ZCC16]. Numbers [Ko13, KB14, dXcY12, XY14, YWS14, BZH19, CHZ17, CL17a, CS19b, Dan18, DWX17, HC16a, NMAA19, NHV17, Pen17, SZV16, SY16, SY17c, TY17, WP19, YYW17, YYS^{+}18, YYWH19, Zha17a, ZMN17]. Numerical [Afs13, BR13, RG16, WYL12, CMN^{+}18, Spa19].

Object [MV13]. Object-Oriented [MV13]. Objective [KLKK17, LKK17, SJ10, TS12, WS13, XT12, HT16b, LS17a, SKK18, SH17b, ZW18]. Objects [KS12a]. observable [BER16a]. Observation [DS09, Den13, Sco11, ZZSP14, DDS16, NAAC16, RAZ19, SSD19]. Observational [YW15]. Observations [BKP11, HB13, KB13, Kum15, Li13a, Lu11, MLGZ14, MSW13, NIN11, ÖA15, RAB15, SGPA12, SB11, Xin10, ZLZ15, BR17, BANC17, Gha19, HA18, KVR17, MV16a, Pop16, PT16, QL19, Van16, WXS18]. Observed [JTR15, MS10a, CY16, LAW^{+}19, PG16]. Obtain [FN09, LWWW16]. Obtained [LN13a, WBH10, PYL17]. Obtaining [LKK15, Wan10a, HT16b]. Occasion [SK10a, SSP^{+}12, SK12b, SK14, SMPH15, SMMS15, BS16a, PM16b, SKM16a, SKT16, SS16c, SKM16b, SP16d, SSB17c, SMK17, SP17c, SKM17, SSS18]. occasional [ZT17]. occasions [SP16c, SKM16a]. Occurrence [VLTP14]. occurrences [Ery18]. occurs [SKM16a]. Odd [Coo12, CARO17, CAPO17, HOAH17, POS^{+}19, SY17c, Shu19, TY17]. odd-even [SY17c]. Odds [BSY12, LZ10, NS15a, TWH10, Bur17, KRB13, KN18b]. Off [LLV10, PVB13, SH17b]. Off-Centered [LLV10]. offered [Gun16]. Offset [PJL13]. Offspring [Rah12, Rah16]. Often [Koz12]. Old [AT17]. Old-Faithful [AT17]. Oliveira [AE23c]. Olkin [AJGG14, BSL^{+}19, BNB18, Ber18, CSL17, CL16, GGDVP16, GT19, HS15b, KCY^{+}19, KJAR13, MS13, MTC^{+}19, MML17, NH15a, NA17, SP16a, SX18]. OLS [CO10]. OLSE [Arv18]. Oluyede [AE23d]. Omega [GH19]. omission [CM16b]. omnibus [HOK17]. On-Line [BQDH12, HD13b]. Oncology [LSD10]. One [Afs13, BP15a, BS12c, BDKM12, CHP14, Eid12, GAK14, GB12, KL10, KTW10, KX18, Muk10, PHW11, SAJMF12, SAJ13, She12b, SE13, Tak14, WXPL13, WS14, XFC11, Yen12, Zha14, ZSJ16, DO19, GGDVP16, HN17b, KRMG19, LQ17a, Lin19, LWG18b, LL19c, MGG^{+}17, RC16, RAZ19, Sou17, TSOT19, WSW16, Wu16d, YSS18, YCAJ18, YL18d]. One- [BS12c, BDKM12, KX18]. One-Dimensional [Afs13]. one-parameter [LQ17a]. One-Sample [KL10, HN17b]. One-Shot [Muk10]. One-Sided [GAK14, PHW11, SE13, WXPL13, Yen12, Zha14, Tak14, ZSJ16, LWG18b, RC16, Wu16d, YCAJ18]. One-Step [BP15a, Lin19]. One-Step-Ahead [GB12]. One-Way [Eid12, SAJ13, SAJMF12, WS14, MGG^{+}17, YSS18]. Ones [Cah15, DZ15, MP15]. Online [BMQ16, XHY16]. ontology [WYG17]. Open [Hug09, SF12a]. Open-Set [SF12a]. Operating [MJ15]. operational [LCWG16]. Operations [Mas15]. Operator [ER11, JS13, ZWZ12, HZS17, KU17, Nas18]. Operators [Yar14, Dut17, Nag18]. Opposite [DDÚF11]. optical [Gun16]. Optima [SG11]. Optimal [ANJ17, AKL15, AIA14, AW10, AEH11, Bal18a, BAA^{+}18, BJVV15, BJVV17, BSB12, CY17b, CG12, CG14a, CK17a, CBSN16, CT11, CSC13, Cha16, CC10b, Che14, Che19a, CYL13, CC10d, CGSM10, DW19, DDM10, DDM14, FG11, Gra12, HW14, HK11, Jaf13, JYZ19, KM19a, KS12a, LH11a, LK17c, LK17b, LG16a, LL19c, LLSM14, MWY18, Mal17, Mar15, MS10b, Mat11, MS14b, Mer16, NC14, OQL12, PGCT19, RK19, SB17, ST16b, TMZ^{+}17, Tin10, TS12, Tra11, cTtL14, VW18, WWWX19, XK18, YZP17, YWX17, YLS17, YKL^{+}17, ZY14, ZY17a, ZWY16, ZDDH17, CG19, DK17b, DS17c, DS17d, EQ17, FWLL19, FPM19, FM17, FN09, GFS16, KAR16, KS19, KD16c, LRC17, LZ18, LPDD19, MN19, NS17, ÖK17, Pak17, QCK19, RZ14, ROM16, SCYR17, SCYR19, Sen16b, ST16e, TK17b, WX17, WW17b, Xio16, YL18a, ZF18]. optimal [ZW17c, ZYH17, ZLZ19]. Optimality [BB14a, BP18c, CEK11, FM12, FM14, Tak10b, HS19, JJ16, JKD17, OPK17, OQ17, Ros19]. Optimization [CYE15, CC19, DGMC17, FY17a, IES^{+}17, KDG^{+}18, MSEN18]. optimization-based [KDG^{+}18]. optimized [LO17]. Optimizing [KV10, PVB13]. Optimum [Atk12, DD13, MP12, MP17, PMA11, PM12, TGXX12]. Option [CCH^{+}10, CMN^{+}18, FNM19, FSSW17, LLLL18, LMC18]. Optional [AO17b, Arn19, GMRN13, GMSD13, GMR16, KC19b, MA18, TS17b]. Options [WW10, HZ17b, LSW17, SL16, ZG17a, ZZ19c]. Oracle [FC18b, GP19, NS19]. Order [AA10, AAG12a, AB14, AH13, Aka10, ADWR11, AAG12b, BTK10, BS12b, BZAZ15, Bar10, BBC13, BB14b, (Ba14, BC12c, Bur09, CTQ11, Cra14, DPA10, DPP15a, Deb12, EB12, EDAAS12, GSB12, GD12, GK13, GB12, Han15, HLX14, ITG09, IUT12, IRB13, JPAB17, JP13, Kak10, Kak12, KO10a, KV10, KT15, KP12b, LWL13, Len17, MCX10, Min10, MA14, MS14c, NC14, OM13, PP10, PJ14, Raï15, RA11, RS10, RGI15, RAB15, Ryc14a, SA12, SDJ15, SP11a, Ste11, SM14, TB12, Tak10b, TOT15, TV11, VT15, WZ10, WL11, XZNL15, ZY10, ZWZ11, ZK15a, ZLD11, ZZC13, AMP17, AHKO16, AO17a, AEHQ17, ARB19, Are16, BP16, BNZ17, BEY18, BD19, BAR16, BCH15, BCK17, CMN^{+}18, ÇG18, CS17, CZ15b, CCM19, CCMA18, DPP17, DYL17, Ery16b, Gor17]. order [GR18a, Had17, HZS17, HD17, HJW19, HLP17, Hu16, HYS17, HWLP18, HM15, KT17a, KM18, KD16a, Kay16, KM19b, KI19, KLKK17, KN16a, KDN17, LR14, LMH17, MH16a, Moh19a, NE17, NA17, PR17, Pro18, RASR17, RAZ19, RJN16, SN17, Sch16, SK16b, SBA17, SYW16, SY16, SY17c, SG17g, TN16, VT17, VZ19, WDLZ16, Wan18, WN16b, YJ19, Yan15, YW17a, hYhW14, ZBA17, ZYLZ17, ZQZY18, ZRL19b, ZS17a, ZWZ17, ZZ18e, ZL18]. Ordered [AT11, BC13, CDZ14, GCR12, EDKAEN17, EW17, Gau17, Ima17, JT19, NCA16, Rez17, Roy16, Sin16]. Ordering [BHB15, BNB19, BPSV14, CC10b, DC10a, FL18a, GW15a, JA10, JAA11, MT12, Por15, XZNL15, YL18d, Zha11, ZL14a, AAE^{+}19, AJ19, CNL19, IKL16, LL17a]. Ordering-Based [DC10a]. Orderings [BNB17, HN16, NHAMG17]. Orders [BC12c, BBHN18, SYC19]. Ordinal [BBG14, BC14, CCSCME13, DBC14, DZ14, FGTS12, GD11, Ian12, Ian14, Kat11b, SG14, AP19, CO16a, CM17, DM18, KOFA18, Kvå17, LLZ17, RSG17, WNL18]. Ordinary [WH15a, ZHL15, OOT19, XKW19]. Oriented [MV13, Par13a]. Origin [AG11, Kum15, SN18]. original [OQ17, dPD16]. Origins [MA11]. Ornstein [Tra17, WTZ15, YSC17b, ZW19a, ZZ19d]. orthant [DWM^{+}17, FPPRG16, LWWH16, LSYL17, QC16, SZV16]. orthant-dependent [DWM^{+}17, LWWH16, SZV16]. Orthogonal [AV15b, FPZ16, FH12, GFR12, HZ18b, LZH15, RN13, SLL17, WLL13, Wan17d, WZL10, WN17b, ZQZY18, BM17, D’A17a, DAC17, DS17a, FFN^{+}19, Fon17, LA17, PYL17, SAD19, ST16e, WZ16, YZFM17, fZWD19]. Orthogonal-array [ZQZY18]. Orthogonality [HR17, Por11, Man17b]. Orthogonally [Pre11]. Osahan [ST16g]. Other [SS14a, Sou17]. Out-of-Sample [HL13b]. Outcome [MMP14, MH15, YW15, EKZ16, KM16b, dPD16]. Outcomes [CGL13, EMI14, dCA19, BS19, CM17, DLB17, HMT^{+}17, KOFA18, LH17a, LCZA17a, LCZA17b, LCA17, PCS18, RS19, SB17]. outgoing [FY17a]. Outlier [LSC12, PK18, WL17c, SL18, Wan18]. Outliers [BIP^{+}10, BO14, GB12, Ko11, KL12, KB17, LF17, NNN10, SM14, WL10, BRZR16, JDSS16, NN16b, WCF18]. outward [Çel17]. over-identified [CLZ16]. over-parameterized [GST17]. Over-Rejection [LS11]. Overall [HM12, LS17b, LL19c]. Overdispersed [EMI14, YHA10a, NC19a]. Overdispersion [Ian14, YHA10b]. Overlapping [Kyr17, UC14, ZZZ19, Kon16, MP15, ZCC16]. Overshoot [Wan13a, CWW16]. overview [ADKZ16, CMS17].

P [MBS17]. PA [Yan12]. Padmanabhan [Mar12]. pair [GG16b, LGLX18, MZ18, ZY17b]. pair-copula [LGLX18]. Paired [HBMAO16b, Kur10, Lu10, Lu11, Wan10b, ZB12a, AAL17, BG19a, HOK17, LWZ17]. paired-samples [HOK17]. Pairs [AAG12a, Mod17]. Pairwise [GJ15, LLZ17, Sun13, KH19, SM18, TWZ16, WCW17, WG16b, WP19, YZG17]. Palindromes [SU13, SU17a]. Pandora [GPÖ14]. Panel [BG11c, Den13, KR14, Wu11, Wu13, ZVS14, FP17b, HH19, Liu17a, Mao18, Phi18, Wu19, XT17, Zha17e, ZL16c]. panels [Hay17]. Parabolic [BCG11, LS15b]. paradox [Dru17, VW17]. Paradoxes [BG11b]. Parallel [Dji10, GSLZ16, LYH11, LWZ13, NRB14, QZZN10, Raq10, YLL19, BNB17, BNB19, CNL19, CZZZ18, FLB16, FY17b, Gey18, GRSN18, HNB19, HN17a, LL19a, Liu17b, NI16, PK17, SB16b, Wan18, YYLM18]. Parallel-Line [Dji10]. Parameter [AMV15, BNS11, CC12, Che10a, CCZR17, CCM19, CA12, CBM14, CT14, DGV11, DA17b, DLZ17, FF13, GFG13, GD12, GY13, HF14, HL15, Kak12, KAPA19, KPS10, Kha12, KG14b, Mah12b, Mat14, MRS12, NBK14, NH15b, PMA11, PMW17, Pao10, PR11, PC10a, PBY^{+}14, Pet10, Pop12, PT15, QZY12, RMC11, ST10, Saj11, SSS14a, SX18, SK15a, SV11, TN10, TOT15, TN15, WCL10, WXPS12, WCGZ12, WXZ13, WSC13, WWT15, XKW19, YC10, YC11a, YSC17b, AE17c, AA17a, ACO^{+}18, AO17a, AR17b, AN17a, ANA17, AB19, AEA17, BHLK15, BBA16, Bay16, BH17a, Bob19, CX17, CW18, Cha18, CW19a, CC16c, CH18, CM16c, CW19c, EM16, GMMA19, GP16b, Gun17, Han17, Han19, HWY17, JC18, KY17a, KNHM17, KX18, KVR17, LQ17a, LY16b, LTC17, MAG18, NB16, NP17b]. parameter [OPK18, PNPB19, RC16, RAA17, RAM16, Sha19, Sha18b, SP17a, SS18, Sti17, SJJ19, TGXX12, TKP17, TPJ18, WZ16, Wan19b, WG16a, YH16, You17, YW17b, Zha17d, ZB17, ZP17, vZvdM19]. parameterized [GST17]. Parameters [AARN10, AVCL13, Che10a, CLV14, DN14, Dos11, Duf10, DDM10, DD13, DP13b, ES14, EM13, FLK12, Fly10, GMRN13, Gil12, Gri15, HASA10, LMG13, LG10, Li11, LL12, LN13a, LG14, MS11a, Mar11, MMV12, Nos14, PM12, PGF09, PS11, SP13, SKG12, SST14, SS15b, WL10, WS13, WSK14, WTZ15, YOD10, YLX10, ZSK12, Zil11, Ahm16, BKL19, CR16, CS16c, CHM16, CZ19c, CCY17, DWW18, DBA18, Dvo17, GFG16, GKP16, GYL17, HCZT19, HS18b, IF18, JYX16, JY16, KKS18, KG18a, KA18, LKK17, LYR^{+}18, LY17, MMD18, MSL17, Oga16, OAOQ16, SK18a, SV17, SM16, Van16, WCK16, YWLW18, YKL^{+}17, ZY17a, ZRL19a]. Parametric [ABI12, BLN13, BPS16b, Eid14, El 15, Han15, HS15d, KR14, KJD^{+}15, KLK14, NK15, PDS09, RAK15, SAJ13, SD12, THHL15, Xu17b, XT17, ZAV14, ZL12, BQC17, BAR16, BH17b, BS17, BG16a, Car19, CO16a, CQ17, CM16d, CMS17, CCMA18, DAS16b, DY16, FJS^{+}18, Fro17b, GT17, HK19a, HL18a, HK19b, HMT^{+}17, Hen17, Hu14b, Hua17a, Hua17b, HJY18, JZ17a, JSZL17, KY17b, KL17b, KG18a, KN17b, LA17, Lu18, MHF16, OE16, OTA17, PS17b, SB16a, SNAN17, Say17, SR17, She18b, Son16, SSR19, Tu17, WG15, WP16, WZX17, XHL18a, Xin15, XTT17, XMP17, Yan18, YLG16b, YYY17, YWL17, ZZ18b, ZZ18a, ZDW17, ZFZ17]. Parametrization [KSK15, Oga19]. Pararai [AE23d]. parent [VT17]. Pareto [Ano23, AM17a, AFL15, AFL13, AG16, AD19, BKL19, Bur09, CCZR17, CM16c, EB12, EDKAEN17, FGP16, FWLL19, GN15, GHBM18, GFG16, Gun17, HLR11, JO10, LWL13, LK19b, LH13, MEG15, MWM13, NA17, NN16b, PBL19, PT10, PGF09, SBA17, SLSE19, TPJ18, VS19, WC16, YYWH19, Yeh15, ZCCW18]. Pareto-optimal [FWLL19]. Pareto-type [YYWH19]. parity [PLC17]. Parshvanath [Kha19b]. Part [FH12, AS17, DS17c, DS17d]. Partial [AE13, BKP11, CMCCO10, CL15c, FLX11, FLW12, LW14, MA11, MZ15, ST16d, TB13, WWL13, WWT15, ZF17, Zha17c, ZLF14, CLZ16, HXF18, JD16, KN17b, LPH16, MX19, RG16, RA16, SA17c, Wan16d, Wu17c, Wu17a, Zha16a].

Partially [BCH12, DLP15, FH13a, FHZ12, FDHH14, Hua11, JMEM14, JS13, LLC10, LX15, Liu11b, LWHW11, LWH12, MZ11, MMT17, Par15b, QL11b, SS10a, UC15, WCL10, WSW11, WSK14, Wei12, WJH15, XTG15, YX15, YHC14, ZX10, ZHL15, AR17b, BER16a, BCL16, FJW17, HH16, HHS18, HZL16, HHC16, HZ19, JQZ18, JHM19, Kha19b, KDGV16, LLHZ16, LZZL17, LH17a, Li19, LXT17, LSCW17, Liu17a, LTGK17b, LLWK18, RTT17, TT17, WL17d, WA17, Wu18a, Wu18b, XQS19, XTS16, XLS19, XFWC19, YSC17a, YCZ17, YL18b, ZCHR19, Zha17e, ZDS16]. Participant [CKD14]. participation [LS16b]. Particle [CYE15, GV11, IES^{+}17, KDG^{+}18]. Particles [Man11b, Zie13]. particular [CM16b]. Partition [Ko15a, FGGL18]. Partitioned [Por11, Nag17, Ren16]. Partitioning [Mat11]. Partitions [APM10]. Partly [YC12]. Parts [Hua12, RVP15]. Pascal [CLGFB15, IS16]. passage [HZ17a]. Past [KRB12, Raq10, RSA13, AER16, Ham19, KHY18, KN16b, KK17b, KK18, Min17, NI16, Rez16, Tri17]. Patched [CSS16]. Patent [Tsi10]. Patent-R&D [Tsi10]. Path [Edi10b, Kia17]. Paths [Lin14b, Maz19]. Pathway [GMB^{+}13, Tho14]. Patient [GMB^{+}13]. pattern [GQC18, SP16c, SP16d, TN19, WQ18]. Patterns [SK10a, SSP^{+}12, SK12b, SK14, SMPH15, SMMS15, PM16b, SKT16, SSS17b, SP17c]. Pavement [KLK14]. payments [TMZ^{+}17]. PBIB [Gar10, SG18b]. PC [Wei14]. PDF [SKK16, MHF16]. Peak [DG11]. peakedness [EW17, Sta17]. peakedness-ordered [EW17]. Peaks [BGT10]. Peaks-Over-Threshold [BGT10]. Pearson [Afe13, Kim12, Lan14, Tor12, Voi10, WH16c]. peculiar [Mud16]. Peña [Hu11b, LH11b]. penalization [HZ18a]. Penalized [Ale16, GL10b, HA14, HSHW14, HNB17, LY17, Mao15, SLZ11, SLF18, Smi13, WSC13, WSlT15, YGL16, ZYX12, LH17c, XMP17, XM17b, YS19, ZX18]. penalties [Has19, TMZ^{+}17]. Penalty [HL11, Has19, Jun17, XZ17, ZWY16]. Penrose [YHS13]. pension [SLZ18]. Percentage [MF10]. Percentile [CW14, HNV13, NRB14, ZJZ17]. Percentile- [HNV13]. percentiles [Hua17a]. Perfect [VB13, ZAV14]. perform [PG16]. Performance [Alg19, CW18, CS12, HWZ19, HL13b, KAT18, KM16a, Kha16, KS14b, LY11a, MKS15, MSK18, NKG16, NO12, NSR^{+}19, ÖK15, RWY16, RM15, Sab19, SLOT17, SWW^{+}12, TEF12, Wu18b, Xu13, YL17, Zha16a, ZSD^{+}15, DK17a, GW19, HW18b, KWM17, LC19, NO18, Nen18, ÖK17, RC16, Rav19, Wu16b, Xu17a, XN19, XO19, Yil18, ZXM16]. Performance-Related [SWW^{+}12]. Performances [XY13]. Period [HNV13, ASR^{+}19, LS17a]. Periodic [BM11, BB17a, BL13, DL11, Ham12, LS09, SCC10, ZZ15, AH18b, Ben19, BG19c, GK19b, MSL17, Pan19]. Periodically [MS14c, HZS17, MHA19, SA19a]. periodicities [Yan17a]. Periodicity [Alm11, SA19a, MDC16]. Periodogram [GS13, HNV13, RMC11]. periodograms [SA19a]. periods [CK17a, FK19, RRRA18]. Permutation [BF12, BPSS12, BPSV14, Bon14, CDZ14, CD12, GFR12, GCR12, Gia14, Pes15, Rat13, Sal15, SGR^{+}12, SB12, UU17]. Persistence [SE13, QL18]. persons [TE18b]. Perspective [FG11, Ham13, AS17, EIM18, ST16a, Spa19].

Perturbation [Sun17b, HJC17, LH17c]. perturbations [KU17]. Perturbed [HJ14, ZYN11, BPS16b, LSY17, ZWY16]. Pesarin [Mon12]. petition [EW16]. Phase [DM11, FK19, GBC12, KOFA18, Ko15b, MGF14, NNI15, Nos14, RPY^{+}13, SK13a, SKK10, SHLK10, AO17b, BS16b, GMRN13, GMR16, JKJ16, KZ19, LJC18, LCWG16, RSC16, SG17b, SG17c, SGA19, SKS17, SK18b, WCK16, aYwSlDmS17, ZMCM16, MATC18, Sag15]. Phase-I [Sag15]. Phase-type [MGF14, LCWG16, ZMCM16]. phased [DLC16, YGWQ17, WYL12]. phased-mission [YGWQ17, WYL12]. phases [SNAN17, CS12]. Phi [MP14]. Phi-Divergence [MP14]. Pickands [AT11]. Piece [ASAN12, MM17, DS17b]. Piecewise [WWT15, AG18, LWD^{+}18, PL17]. piecewise-deterministic [AG18]. Pilot [CT13, KMC10]. pipeline [JKL19]. Pitman [AB14, AMP17, BDKM12, EWRZ14, RL09, Sab19, ÜY13, YLX10]. pivot [FFN^{+}19]. pivotals [Raj18]. placebo [Liu17b]. plan [AAJ16, ASeMJ17, AAJ17, BAL18b, Bal18a, BAA^{+}18, BAW19, HBAJ17, KAJH18, NS17, NN18, NS18, Sen16a, SLZ18, UB17, WT18, WAJ17, YF17]. plane [Ham19]. Planning [Gue14, Liu17c, KDGV16, RFM18]. Plans [AYC^{+}12, CEK11, CB11, EF11, JFS14, LC15, EQ17, FY17a, JJ16, JKD17, KS19, Sen17b, Wan17d, YF17, YWZ18, YCAJ18]. plant [AIIK17, BTC16]. Plants [GG13]. Plausible [Eid14, LW17a]. Plot [KMM12, FPM19, FWZ^{+}17b, LWWW16, WYZ15]. Plots [KCK10, Oli13, HT16b, NN16a]. plotting [HT16b]. PLS [MA11, LvR12]. Plug [CT13, WX15]. Plug-in [CT13, WX15]. PMSE [XO19]. Point [AVCL13, Deb12, DMT^{+}15, HKLW12, MF10, ML13, MRA16, Nos14, PJ10, PZ15, QTD10, RL11a, RCdC14, SY10, Sur12, WJY15, Wee12, CKT18, Ciu17, FPM19, Far17b, He17, Ho16, Huh16, JSC17, LPDD19, MATC18, SSY18, TZW16, WSW16, ZFZ17]. Points [AW10, Koi10, KQ11, MFSAS10, MK13, MF10, MS12b, SASMF13, WZY14, WG15, Xu13, YS15a, ZLL11, CN17a, GDDC19, Xu17a, ZBTL17, ZT18]. Pointwise [DS17e, CN17b]. Poisson [AE23b, AE23d, AB17, AP17, AA17b, AESO19, AAO17, BD16a, BKN16, BKM18, BN18b, BKGD17, CYF^{+}18, Cha10, CO16b, CRG13, CS19a, CHM16, CG11a, CC16c, CLW17, CKT18, Chr10, COL15, Cox15, DHP^{+}16, DKY16a, DKY16b, DS16a, Fam15, FK19, Far17a, Far17b, FI17, FC18a, Faz11, FWH^{+}15, FYLC16, zF16, FGL18, GH19, God16, GDSC15a, GDCO16, Gri15, HL13a, Hay16, HHW14, INOC17, JM16, JKS18, KYS13, Kap18, KDG^{+}18, Kou16, KXX11, KPZ16, KN16a, KS18b, KV17, Lan14, LH14a, LH11a, LPC17, LHS19, LS14, MZ10b, MATC18, MGF14, Mon14, NA12, Oga16, PWLO17, POS^{+}19, RvdM14, RBB13, SGD11, SH17a, SG16c, SSD18, ST17b, SSSK18, SSS19b, SJ10, SK15b, SI13, TS17b, Tee12, Tee14, Tra17, Tsi10, UZ19, VPCG16, WP18, WZT15, YHA10b, ZI12, ZI14]. Poisson [ZY14, ZL16a, ZZLL15, ZWP17, ZSJ16, ZYN11]. Poisson-conjugate [GDCO16]. Poisson-multinomial [FGL18]. Poisson-Negative [Tee12, JM16]. Poisson-type [CO16b]. policies [Cha16, Che19a, ZMN17].

Policy [AEH11, CSC13, Che14, HW14, JHP10, KY10a, NN13, QZZN10, SCC11, ANJ17, JX19, KC19a, KC19b, KY16, Kyr16, NS18, QCK19, SFSH16, YL17, Ye19, YLS17, ZW17c, ZW19b, ZDDH17]. Polled [GL15b]. Pollution [LVBS10]. Poly [WWH15a]. Poly-Variograms [WWH15a]. Pólya [FCJ^{+}17, CW14, CZ19b, FLK12, GL10a, LVVN11, Min10, Min11, MB14, SAB15]. Polygon [sDsWcCjK14]. Polygonal [MGP11, TIA11, NM18]. polygons [KAW^{+}18, WdX17]. Polyhazard [TH15]. Polynomial [AOÖ17, BB12, KDK14, KP12b, LLQ13, MFS12, WWT15, Cos18, HT16a, KD16b, Lin16, NIN17, WT18, ZDZ17, ZLZ19]. polynomial-logistic [KD16b]. Polynomials [WZL10, Zoi09, D’A17a, DAC17, FPZ16, SLL17, WN17b]. Polytomous [BBG14, MP14, Gol16]. Pool [BAK13, GAK14, Ken11]. Pool-Testing [Ken11]. Pooled [TD14]. popular [SHANA18]. Population [AB14, AMV15, AC15, AM15, Bak11, CKLS11, CP14, DP10b, DG12b, Din15, DSSL15, GÖ12, GBC12, GS15c, Hug09, Jaf15, KG14a, KK14, KK11b, LWWH14, ÖÇO14, SAS11, SG10a, SG10b, SK13a, SRT14, SS15a, SKK10, SK12b, SS13b, SS13d, SS13e, SK14, SST14, SMMS15, SS14a, SS14b, SS15b, SS14c, SG11, TC14, TSS15, TTC17, YUSC14a, YUSC14b, YCL14, AHH16, ASG17, AN17a, AH17b, BS16a, BS16b, BP18c, BRAS17, BCP17, CSB19, EEE16, FGP16, HKH17, HLY16, JN18, KGH19, KJ17, Koy18, KZ19, LJW18, LTV17, MSB19a, MAG18, RM16, MA18, MF16, MT16, Moh19b, MKK17, MSK17, NLRD17, Pat16, PM16b, PT19, QKCH19, RSC16, Sen16b, Sen16a, Sen17a, Sen17b, Sen18b, SG17a, SG17b, SKT16, SSS16, SS16c, SKM16b, SHM16, SP16d, SG17e, SPS17, SKS17, SK18b, SSS18]. population [SMP18, SPY18, SG19, SS16g, TMZZ19, WXZ16, YSG17]. Populations [CF12, CF14, CHP14, CDZ14, Fle13, HX14, KT11, PGF11, RB10, Ryc14a, Sub14, TD14, ZL14a, AA17a, AB19, AAK17, CQ17, Gor17, Gun17, HD17, JT19, NKG16, NP17b, SS18, SG16e, TKJ17, WHZ17]. Portfolio [BCCH14, HF19, SLZ18]. portfolios [BNB18]. Portmanteau [BBC18, DY16]. Portugal [PMC13]. posed [AD17a, AD17b, MDR17]. Position [BSB12]. positional [Man17a]. positions [HT16b]. Positive [AD10, DMP11, De 11, LY16c, Mel13, MF10, XY12, XY13, ZC11, Bar19, DLZ17, Gor17, HB17a, LL19b, ZBA17]. Positive-Rule [XY12, XY13, LY16c]. Positively [LWY10, dXcY12, XY14, GP16a]. Possibilistic [WLY14a]. possible [LTZ19]. Possibly [PZ15, ZLF14, CDS19, FLY16, Miy17, ZL16g]. Post [MS15, SV10, TTC17, Woj11, WS12b, Nag17, Vis18, XHY16]. post-change [XHY16]. Post-Model-Selection [Woj11]. Post-Mortem [WS12b]. Post-Stratification [SV10, TTC17, Vis18]. Post-Stratified [MS15, Nag17]. Posterior [BP17, BC14, BC15, Gue13, HXL13, Miy10, Sag15, TL15a, Wan13b, WH13, Xin10, YC11b, ZLDT14, Abu18, CFA17, Das16a, JN18, MSEN18, Slj17, Su17b, Tsi18, ZY16]. posteriors [KC16]. poststratification [DAS16b]. Poststratified [Par15b]. Potency [Dji10, Che16a]. Potential [Chi11, Ko12, Wu13, Ko16a]. Poverty [CD15, Meh17, SNWL17]. Power [DLL13, DCG13, FZZ16, Faz11, Fu12c, GK19b, HD15, JK19, Kak10, LCG12, Lu13b, MN10, MRS12, Mur13, Pak14, PS11, Ree12, Sal15, Sto11, WS12a, WXZ13, XLL15a, Xu10, YK11, ZK15b, Ahm16, AE23d, BD16b, BMFS18, CRTdC17, ÇG18, CH17, CTAL17, Gen19, GG16b, HA16b, Hut19, IR19, JT16, Jon18, KM16a, KRMG19, LSW17, MFBG13, MFPG16, Mur16a, NORV19, NAMAK17, PWLO17, PLC17, RSG17, RS19, RN17, ZJLO17]. power-binomial [Gen19]. Power-Law [Ree12, WXZ13]. power-Lindley [IR19]. power-normal [BMFS18]. Power-Transformed [XLL15a, FZZ16, Lu13b]. powered [CFK17, LP18]. Powerful [CB12, Voi10, BMRSE17, PLC17]. Powers [Kak12]. pp [Mül12]. PPS [SMP18]. Practical [DG12a, ZK13b]. practically [Yam17b]. Practice [GdSC15b, AE23c, DMM16, OSX^{+}16]. Pre [NO12, YK11]. Pre-Test [NO12, YK11]. precautionary [GKP16]. precipitation [AEB17]. Precise [CS19c, FL18b, GWZ17, HSLQ17, HL15, KL11, LYG17, Lou19, LSX14, LSW16, LQ17b, LSL18, ST16c, SHLQ17, WW13c, XY10, YS16, WGW19, YC18]. Precision [Ng12, Sim14b, SSH10, CS16c, DZH^{+}19, WW17b, XM17b]. predator [KD16c]. predictability [CW19b]. Predicted [JTR15]. Predicting [SF16, HL18b, WER^{+}19]. Prediction [AB09, Ala14, AF10, BS12c, BA10, Bas14, Bin14, CB17, EDAAS12, Fau12, GB12, GBC12, Ham19, HL13b, LWR11, Men10b, MM11a, RA11, RGM15, SDJ15, SK10b, UC15, VB12, Zad15, AB17, BQC17, BAR16, BN18b, EDKAEN17, GAS17, HLY16, JN18, KTR18, MLD17, MAA17, OPK17, OPK18, Sha17, SBA17, STRD17, XZ16, YC18, ZLD16, ZH16]. Predictions [YYX14, GKP16]. Predictive [BS16b, BSSJ11, BC15, CMCSC12, CCSCME13, FU10, HXL13, KPS10, Kha12, Kha13, Mar15, Men10a, Men13a, ÖK15, Tra11, Tri17, AD19, CM17, CCMA18, DK17a, Ras17, Wu16b]. Predictor [AKL15, GR15, PC12, ZS12, BH17a, BCB19]. Predictors [BDKM12, Bea15, Bur09, HX11, LH11b, QZY11, SW14a, GST17, SJJ19, ZL16d]. predictors/estimators [SJJ19]. Preface [CNT14, CY12, DS10, Edi10a, EPRS14, EE15a, FvR12, Kol13, MSS11, RdC13, SS11a, SM13, Vas14c, VM14]. Preference [AA13, MSEN18]. Preferences [Far12, Fon14, GGFL11]. Preliminary [BAT15, CY14, CW19a, CBM14, LYX11, SS15a, YL15, AMP17, AW16, Cha18, KAT18, RPH17, XO19]. Premier [GDDC19]. Premium [WWW13, AHP19, FWLL19, GH19, NAN17b, YWX17]. premiums [WYMZ17]. Prescott [Nas18, Yam17b, Yam18a]. Presence [AB13, Bis12, BDM13, CZ15c, Chi12a, DG12b, FHS15, GB12, GSSCR12, HLL15, IT12, Kak12, KK11b, Mån12, MSS14b, QL11a, Sal15, SK13a, SSS14b, SN14, SS14c, SM14, TSS15, ZWZZ15, Zho10, ASG17, BR19, BIJ17, BMRSE17, BMQ16, BP18c, CSL17, CSB19, CM16c, DP18, Fan18, GFL19, GK19a, HJC17, KGH19, KJ17, KZ19, LP19, LWZD17, MSB19a, Man17a, NNZA16, NNZA18, NN18, NS18, NN16b, RWY16, Sen18a, SG17e, SSB17c, SSS17a]. present [ZL17]. Preservation [IRAB13, IAB16, WZ10, ZM13]. Preserved [Hou12]. Preserving [LCG12]. Prevalence [BAK13, Ken11]. Preventive [AEH11, CC10b, Che14, CC10d, SCC10, SCC12, Cha16, PM18]. Price [Bia12, Bia15, CCL14, LIPMF13, Bia17]. Prices [LIPMF13, BBS16]. Pricing [CR14, Hua14b, LSW17, QJWC18, QWWW19, ST13, WW10, WZJH19, YSK15, CMN^{+}18, FNM19, FSSW17, GDW18, HZ17b, HHC^{+}17, KYW16, LLLL18, LMC18, MDG19, XDW19, ZG17a, ZZ19c]. Principal [ASAK10, BYDM15, EZ14, KQ11, MK13, NKS15, OFS13, WY13, YS15a, Li18, PG16, Wu17b, YH16]. Principle [Mia11, WWW13, LS17e, MGLX14, PM16a, YWX17]. Principles [Fu12b, GM14, FWLL19, MGN16, WGW19]. Prior [DMST12, KG14a, MMV12, SS15a, SK10b, Xin15, YD14, AB17, CTAL17, JSC17, PK17, SUS19, YM17, YJ17, ZJLO17, ZRL19a]. priority [VM17, ZW17b]. Priors [AT15, GVGP11, Gri15, RS10, WS13, ZK13b, DE17b, TL16]. Pro [CC10d]. Probabilistic [DTO13, El 15, VBATK11, ZLDT14, GDDC19, KH19, WLZ17]. Probabilities [AJ11, AASR12, Far12, GÖ12, Gue13, GW15b, Kon11, ZYN11, Bic19, EKZ16, FL19, Geo17, Hay16, Hon17, Kyr16, LS16a, LWG18b, Rav19, RK16, Sel17a, Sun19, XHY16, XX18]. Probability [AA14, BS12a, BB14a, Can15, DTE13, FM11, FF13, GK11, Han11, HJ14, Kát11a, Lec11, LCG12, LBL^{+}10, LDY15, Mat11, PHW11, RMN16, SW14a, SO14a, TE11, Tan13, TP14, VYZ14, WXPL13, WLY14b, XLL13, Yan14, AA17b, ABTR19, BR17, BSW15, BR19, Bow17, CM16a, CC16a, CFA17, Cle18, Fro17b, GWL16, JR18b, LA17, LP19, LY16a, Li17b, LG16b, LGG16, LWG18a, Lu16a, RZHC16, SSSK18, SSS19b, TL16, VN18, YGL16, YWH17, ZK16]. Probability-weighted [RMN16]. Probit [BBB10, CT10, Fed13, Fed14, VMVD14, AXLJ17, Ker17, Roy16, YJSL19]. Probit-Beta-Normal [VMVD14]. Problem [BC15, BEZ11, GGS13, GVGP11, Gue14, KM14b, LH14a, MN13b, MM11a, SL14, TPTT14, VBATK11, WBH10, XT11, AD17a, AD17b, CO16a, CLZ16, CMS17, GG16a, GR18a, HA16a, JSC17, MWY18, Mar16, MS17b, PCL16, PTV18, RZHC16, SR17, Udo18, WRZ19, WZZ17]. Problems [CB10, DB13, DB15, DI10, KHM12, Kor11, MH11, Mon12, SP11b, Sto12, UC14, VSS12, YC11b, Far17b, FEP18, Kar17, KX18, LMZ17, MDR17, NAAC16, PB18, hXZ17, Yam18a]. Procedure [BK09, CW14, CD12, HD13b, IUT12, KXX11, MJ15, Sha10a, Sha10b, SJS15, TSS15, Tak10a, TS12, WH14, WX15, ZK13b, Che16b, EM18, FGGL18, LMZ17, MDR17, Pat16, RMN16, SS16c, SS18, WX17, WHZ17, Wu19]. Procedures [CHLP09, HH15, Koi10, LWWH14, Xin15, CK17c, JZZZ18, Kar17, Kou16, SH17a, SM18, SP16b, SSKP18]. Process [ARD14, AGR15, AAM14, AA15b, Ano13b, AYC^{+}12, BD11b, BQDH12, CCL14, CC15a, CC15b, DMS14, DMST12, Dos11, Fat13, GS11, GL15a, Gra11, GW15b, Hu14a, Hua15, HL15, JP13, Kac14, Kem11, KTW10, LS14, LPLW13, LH11c, MCY10, Min11, MS11b, MO14b, MS10d, MS10e, Neg09, NNA12, PS14, PHW11, Pul13, QL15, Rah12, RL11a, RNB12, RCdC14, SY10, SM09, She14a, SHH13, SW14b, Sim14a, THHL15, TKW12, TS17e, UU11, VYZ14, WXZ13, Web12, XLL13, Yen12, ZARH14, ZY10, ZWZ12, ZYN11, Zil11, AO18a, AR17a, AMV18, AESO19, AK19, AH18a, AAK16, Ali17, AABL17, BU17, BAL18b, Bal18a, BER16a, BBS16, BPS16b, BCK17, BAW19, COK17, CB17, CC16b, CL17b, CKT18, CDS19, CJ17, CY16, CWW16, DP18, DTWW17, DW19, El 19, Far17b, FN18, FCJ^{+}17, Fro17b]. process [GFS16, HL18a, Haq18, HUK19, HK19b, HJ14, Ho16, HBAJ17, JJ19, JBA18, KAS17, KAPA19, Ko16b, KT18, KD16c, KV17, LZY^{+}17, LW17b, LCG18, LS17b, Lu18, Mud16, NRP16, Nen18, NS17, OAOQ16, PLDS17, Pro18, QJWC18, Rah16, RBN19, SA16, Tra17, UK16, UB17, WCK16, YLW17, aYwSlDmS17, Yan17b, YCAJ18, YKL^{+}17, ZHW17, ZW17a, ZW17b, ZW18, ZW19b, ZL16g, ZG19b, vZvdM19]. process-waiting [KV17]. Processes [Aga11, AD10, AR14, BP15a, Ban13, BB12, BEZ11, BZ13, CC10a, CCZ15, CTQ11, CC11, Chr10, CLX14, DDGM13, Deb12, DO13, ES11b, FR10, Faz11, Fu11a, GZSS14, GGLV14, Gra10, Gri15, GL11, JRM10a, JWZ14, JT11, KP13, KB14, KOP13, KOP14, Li11, MR15, Mas15, MN13a, Mia11, MC11, MXD13, MZ15, MS10c, MS14c, MM11a, MSP12, NH15b, OP11, PS13, Par13a, PMC10, RZ13, RBB13, SA12, SA15a, SP11a, Sun11, SI13, TW13, TB15, TT11, cTtL14, Tun15, WTZ15, WST10, WH10, Yeh13, pZ13, ZWZ11, ZY13, dSE11, AP17, AG18, BD16a, BT18b, BK19, BD17a, BD17b, BG16b, CB17, CZ19b, CLW17, CS19c, CW19c, DKY16a, DKY16b, DYW18, DLZ17, Far17a, FPR17, Gra17, GPP19, Hac18, HZS17, KLMD16, Kap18, KD16a, KUWG19, Lii17, LPHC18, LZF17]. processes [Mal17, Man18, MG19, MSEN18, PSF16, PM18, PL17, RWY16, RC16, RPM18, SNWL17, SA19a, TN17, TMZZ19, WSW16, Wan19b, XX18, YSC17b, ZW19a, Zha16a, ZZ19d, ZSJ16]. Processing [Afs13]. producing [Has19]. Product [BZAZ15, BNE15, CA14, DMP11, Duf10, FR10, SS13c, Smi19, SYP13, TC14, TTC17, Wan13d, YUSC13, YUSC14b, AE17a, AE17b, AEHQ17, CXC19, CFK17, GSM16, GG16b, KPZ16, KZ19, LPDD18, LTS16, Mat18, RM16, MX19, NA17, SSS16, WAJ17]. Product-Limit [FR10]. Product-Type [TTC17]. Production [BQDH12, Fal14, Cox19, Kyr16]. production-inventory [Kyr16]. Products [Fu12b, KRT13, Li13b, SCSMN11, ZY13, DYW18, Wu17c, Zha16b]. Professional [AK10]. Professor [AK10, Edi10b]. Profile [Li17d, SN14, Sug13, WSK14, ZHL15, CH17, IKH19, PS18]. Profiles [ASMZ14, ES14, HX14, NKS15, NNI15, SN14, MATC18, SNAN17, SA17a, WT18]. Profit [AEH11]. Profitability [SYP13]. Prognostication [GMB^{+}13]. programing [BCB19]. Programming [Fal14, MGP11, NN11b, Smi13, UC14, Fon17, KAR16, MPG16, MGP17, Xia17a]. Progress [NM17]. Progressive [BHLK15, Bas14, CB11, CT14, JMEM14, JY14, MA14, PGF09, PGF11, RB10, Wan12, ZARH14, ZS15, AN17d, AAK17, CSL17, LSCW17, MK19, SK18a, STRD17, XZ16]. Progressively [BS12b, Cra14, RAB15, Rez17, SK15a, TTZ14a, BSL^{+}19, FGP16, GP16b, LS17c, WL17a, WSS17]. PROMETHEE [AM16]. promising [LL17f]. promotion [KDGV16].

Prone [ZZ11, LXT17, WJH15]. Proof [Kát11a, Nav16]. Propagating [New11b, Bic17]. propagation [RBK^{+}19]. propensity [GGtH^{+}19, LCW^{+}18]. Properties [AA10, ACO^{+}18, AFL19, AASR12, AE23d, BS15, BL13, Cha10, CC12, Cha15, CB14, CCZ15, CG11a, CM12, CARO17, CAPO17, DC13, El 15, Ery12, Fat13, Fu11a, GLZW12, GAK14, GS13, GCSB10, Gup15, Hod14, HGZG15, Hud14, Hun14, Imo16, IUT12, IS11, IRB13, Jon18, KM10, KKH17, Kim11, KRT11, KA15, KD17b, KN10, LK19b, LG10, LVBS10, MV14, MB15, MK13, MCX10, MML17, Mit13, MP10b, NCO15, NSV12, NAMAK17, NA19, PWLO17, PSP12, RLSL19, RCC12, RMNA17, RSAN17, SP11a, SU13, TCM^{+}16, TB12, TK14, Vas13, Vel14, VFP17b, WLF11, XWJ10, Yil12, ZLN11, ZL14a, AS19, BR19, BBS16, CNL19, CC16b, CBOR18, DDD17, FLY16, GMS18, JB18, JYX16, JXW17, JK19, KLMD16, KHY18, KSBN17, KH18, LL17a, LLD17, Lin19, LMH17, NSR17, Rao18, SG18a]. properties [SGGdML17, SG16d, SC19, SWZ16, TC17, TS17e, Too17, VFP17a, WH16c, Wu17b, XJW19, XHL18a, Xu16, YYL17, Zen17a, ZB17, ZWPB17, ZZ17d]. Property [Deb12, RR14, Sal15, SU14, WZ10, DY18, HYS17, Tra17, WGW19]. Proportion [KL10, LVVN11, RD10, JSSO18, Jun17, Lu18, MA18, SG19]. Proportional [AJGG14, CL15b, CW15a, DJ15, DSA16, HT10, HA14, HSHW14, IT12, LZ10, NSS18, RA11, RG15, SS12a, TWH10, TS12, XLL15b, Zha15a, AMP17, BBHN18, BBA16, BB19a, GP16b, GHL17, JR18b, KN18b, LP17, LCW^{+}18, PA18, Qiu17, Rez16, SAD16, SS16a, SSSK18, SSS19b, YWX17, ZYH17]. proportionality [FWZ^{+}17b]. Proportions [KZ15, LWWH14, MP12, NN11a, New11a, PJ10, SR11, Zha15b, BS16c, CS18, CGS16, JSS16, MP16, Sen17a, Sen17b, TK19, WY16]. proposal [Gol16]. proposing [TP19]. Proposition [Bia12]. protections [QJWC18]. protocol [MV16b, WH18]. Proxy [ÜY13]. Pseudo [Gar10, KM14a, KSBN17, KTBN17, KSDB18, MZ17, RSV10, VEKH15, KT17b, WXS18, ZX18]. Pseudo-Likelihood [KM14a, MZ17, KT17b, ZX18]. Pseudo-Likelihoods [RSV10]. Pseudo-maximum [KSDB18]. pseudo-observations [WXS18]. Pseudo-truncated [VEKH15]. pseudocounts [Oga16]. Pseudolikelihood [ZT13]. purchasing [PLC17]. pure [DYW18, OAOQ16]. pure-birth [OAOQ16]. PVC [HC12].

QAL [PDS09]. QML [DK13]. QMLE} [BG19c, AL15]. quadrant [YC17]. quadrant-dependent [YC17]. Quadratic [Afe13, Are14, FH13b, FN09, Kon10, LR10a, LWR11, PM12, RS13, RLL12, SK12c, Smi13, TL15b, BBC18, Fon17, KK17a, RK18, SK19a, ZZL16, Zha17b, ZZJ19, ZKW16]. quadratically [Fon17]. quadrivariate [YVCA17]. qualitative [SG19, ZLZ19]. Qualitatively [WBH10]. Quality [Ber12, Cui12, HD13b, HC12, PDS09, PVB13, SW14c, FY17a, FEP18, HME17, Ko16b, LPHC18, PMFK18, TS17e]. Quality-Based [SW14c]. quantifying [FEP18]. Quantile [AT19, CTT15, CZ15c, CT16, CLX14, DNL12, De 11, DZX13, FR10, Fau12, GZSS14, HS15d, JZQ15, Kay18, LK14, Lu13b, MCX10, MSN14, Pao10, Par15a, QL11a, Ryc14b, SY11, SS17a, SE15, SF12b, TTZ14b, WLF11, WL15, WWH15b, Yu12, ZWNS18, ZFD17, AM17b, ARJF19, BK18, BSGS16, BMR17, BLT16, BS15, CL17b, CA18, Ciu17, DGMC17, FTT16, FZZ16, GGO17, JQZ18, JHM19, LD18a, LY17, MG19, MMT17, NSS18, NI16, PLC17, Qiu18, SAD19, SYC19, SK19a, SC19, SLF18, SQCL19, SSH17, Sta17, Sun17c, SR18, TLT17, WH16a, WGG^{+}19, XTT17, YS19, YWL17, YWLW18, YD19b, ZZ18a, ZP17]. Quantile-based [Kay18, SS17a, BK18, BSGS16, BMR17, SC19, SR18]. Quantile-Copula [Fau12]. Quantiles [AB14, Bak11, BHLK15, KT11, RVD14, Ryc11, YWH14, BAB^{+}18, Hen17, JT19, KOA16, MST19, MRA16, SCYR19, Sze17, TKJ17, XY19, YZ17b]. Quantitative [Ber12, GMSD13, Zho10, Arn19, BSH17]. quantities [JZ16]. Quantity [Bia15, CC11, PVB13]. Quantity/Quality [PVB13]. Quantum [GV11]. quarter [Ham19]. quarter-plane [Ham19]. quasars [DBC16]. Quasi [GLZW12, GJ15, GL10b, PC16a, Phi18, Saj11, Wan13c, Wu17a, XML16, ZLN11, AD16, AD17c, BJC16, HA18, JXW17, LSL16, ST16a, WCW17, XJW17, XJW19, ZX16]. quasi- [ST16a]. quasi-associated [AD16, AD17c, HA18]. quasi-asymptotical [WCW17]. Quasi-Likelihood [GLZW12, GL10b, BJC16, JXW17, XJW17, XJW19]. Quasi-Maximum [ZLN11, PC16a, ZX16]. Quasi-Minimax [Wan13c, Wu17a]. Quasi-Ranges [Saj11]. quasi-stochastically [LSL16]. Quasi-sure [XML16]. Quaternion [ABLR15, LX10a]. query [LDL18]. query-agnostic [LDL18]. question [LPdDTH17, MA18, Sen17b, TS16, TS17d]. Questions [BBP14, TWTG14]. Queue [CM13a, Dra13, GMRN13, Kem11, SRK11, SU14, Vas14a, Bur19, GMR16, JX19, KC19a, KC19b, KY16, PBC17, SA19b, YL17, YL18c, Ye19, ZG19a]. Queueing [XZD10, BL19, PM16a, TSOT19, Vas16, ZXM16]. queues [AS19, SLOT17, YFV19]. queuing [DP18, KS18a, SK16e]. Quick [Gov11, BU17, BJU18, MB19]. Quickly [WXZ18]. Quotient [BNE15].

R [BAA^{+}18, GPÖ14, PBC17, TN13, Gue10, HBAJ17, KHY17]. R&D [Tsi10]. Race [GL15b]. radar [SM17]. Radial [FG11]. radiation [El 19]. Rahim [HME17, PKM17, PMKF17, PM18, PMFK18, RMM17]. raise [SGGdML17]. raising [GR17]. Random [Abd15, ADP10, AC15, AKK09, AK14c, BS12a, BP18b, (Ba14, BGT10, BCCH14, Cey11, CG11a, CW13, CW15a, CYL13, CA14, Con10, Dai16, DSO12, Don11, EMI14, FJ11, FWT13, Fle13, GZSS14, GK13, Guo14, GS15c, GE11, HS15b, HS13a, HEHM11, HH13, HS15d, HWYW14, HGZG15, HWW15, HP14, JA10, KR14, Ko13, Ko14e, KL11, Kor11, KRT11, KOP13, KOP14, KLM10, Kur10, LG10, LR10a, Li13a, LX15, Li15, LZ18, Lim11, LR10b, LS12, LSX14, MP15, Mas15, MWM13, Mon14, NRW15, ÖÇO14, Pul13, QKP12, Rat13, RB11, RMC11, RE10, SSA10, SY17a, SG10a, SWL14, SWLW14, Sho15, SMM10, SS13c, SS14b, Sun11, Sun12, Sun13, Sun15, TC14, TSS15, Tee14, Tho14, TCN15, VT15, WY11, WHVH13, WW13c, Wan13a, WLY14b, WS14, WDM14, WZY14, Wei12, WGG^{+}19, WJ11, XTG15, XY09]. Random [XY10, XM14, YUSC14b, Yan19, Yar11, Yar13, Yar14, YTX14, ZY10, ZWZ11, ZLD11, ZWPZ15, Zie13, AD17a, AEHQ17, AD16, AD17c, BSW15, BSL^{+}19, BZH19, BNZ17, BEY18, BD19, Bib16, BH17c, Bud16, CS18, CSB19, CLLW16, CF16, CKS17, CG17, CL17a, Che19a, CXC19, CS19b, CZW16, DP17, DLH^{+}18, DWM^{+}17, DTWW17, DRAS17, DW17, DS16b, DS17e, EG18, FL18a, FC18a, FPCP16, FWZ^{+}17a, FWW19, Fro17a, GHSZ19, GSM16, GMR16, GGO17, GHL17, GX18, Had17, Ham19, HF19, HSLQ17, HZZ19, JKL19, Kal17, KM18, Kay18, KKKÖ17, sK16c, Ko16c, Ko17b, Ko18, Kon16, Kon19, KVR17, Kun17b, KP18b, KN18c, LCG18, LPH16, LWWH16, LSYL17, LTS16, LTV17, LSZ16, LSW16, LQ17b, LSL18, LQYZ17, MSB19a, MAS16, MWW18, Men19, MX19, MKY16, MSK17, NC18b, NAC19, NMAA19, NHV17, QC16, Rao17a, RS19, Roo17, SB16a, aSaCN18]. random [SG16b, SAB18, SG17a, SG17b, SG17c, SH17b, SBA17, SWZ16, ST16c, SZV16, SS17b, SXW17, SYW16, SJY17, SY17b, SSB17c, SKS17, SK18b, SG19, SN18, SKK16, SS16h, SA17c, SHLQ17, SW19c, SJJ19, TWZ16, Tan18, Udo18, UZ19, VFP17a, VFP17b, WHV16, WL17b, WW19, WGW19, WN17a, WN17b, WV16, WG16b, Wu17d, WHY18, WXZ18, WP19, XTL^{+}17, XDWC18, XJW19, Xia17b, Xie16, XXMZ18, Yan18, YHW17, YZG17, YJS19, YYWH19, YQC17, YLS17, YC17, YW18, YZ18, ZG19a, ZK16, ZSL17, ZC17a, Zha17e, ZT17, Zha17a, ZWPB17, ZMN17, ZYH18, ZL16g, ZY16]. Random-effects [EMI14, SJJ19, XXMZ18, Zha17e]. random-type [Udo18]. Randomization [Ani11, OM13, YSS19]. Randomized [Arn11, ASN12, AS15, Chi10, DP14, GCR12, GMSD13, MJ15, MH15, RN13, Shi10, SKHL10, SHLK10, SK12d, TWTG14, Wu16a, XTT17, YW15, AAE17, AO18b, AO17b, Arn19, EM18, GZ17, HS18a, JSS16, JSSO18, KC19a, LPdDTH17, LHS19, Sen16b, Sen16a, Sen17a, Sen17b, Sen18b, SM18, ST16d, ST16f, ST16g, SG17f, SG17g, ST17a, SG17h, ST17b, SSSK18, SG19, SSS19b, TS16, TS17a, TS17b, TS17c, TS17d]. Randomly [GJ15, WHHW17, WHYW18, YXC11, YC17, LL17d, Roo17, TZW16, ZL17]. Range [BNE15, Dem14, GS15b, IS11, ZLF14, Ber16b, BS17, CZCM19, WH16b, Zan17]. range-based [CZCM19]. Ranges [Saj11]. Rank [CB10, CB12, CM12, FZ17b, Hou12, KAD14, Say14, SP15, Tia12, YDC15, Bin17, Che18, FZ18, KN17b, LVH19, ML19, OL19a, YGL16]. rank-estimates [ML19]. Ranked [AARN10, Bak11, DCL12, FDHH14, Fre14, GÖ12, HASA10, LCP^{+}14, MADSAS10, PL12, SAS11, VB12, VB13, ZAV14, CLZ19, EDKAEN17, EDT18, FF16, FF18, HBMAO16a, HBM16, HBMAO16b, Haq17, Koy18, LB16, MZ18, RM16, PT17, PT16, SM16]. Ranked-Set [Fre14, FF16, FF18]. Ranking [BG11b, CDZ14, Gia14, PL12, VB13, Yan16, ZAV14, Alv16]. ranks [LYR^{+}18, MC18]. Rao [GPÖ14]. Rare [FNDM12, SK15b, LHS19, QKCH19, ST17b, SSSK18, SSS19b, TS17b, WN16a]. Rasch [BBB14]. Rata [CC10d]. Rate [BP15b, BR13, CCZ15, CZ15a, COL15, Dom11, DMT^{+}15, ES11b, EdSMR14, GK11, KRB13, KOP13, LZPJ11, LH14b, LDY15, LB11, Mel13, Muk10, Nad10, NS15a, QWZ14, RPY^{+}13, RGI15, RB11, RR13, Ryc14a, SJS15, SWLW14, Sym19, TK14, WZY14, WX15, dXcY12, XY14, XLL15b, Yil12, AE17c, AD17b, AG18, BSW15, BKM18, BBA16, Bic17, CM16b, CLLW16, FNM19, FSSW17, GH19, GP16b, HP16, HB17a, HP17a, HB17b, HA16b, JZ17b, KJT17, Ko17b, LMdO18, LO17, MKSS17, NE17, NAN17b, NB17, PKM17, PK19, POS^{+}19, RG16, SAD16, SZV16, SLZ18, TPN18, VS19, WL17b, WH18, WP18, XX18, YJS19, YCL17, ZL17]. rate-making [NAN17b]. ratemaking [ZHW17]. Rates [CB12, Fu11b, PMC13, RvdM14, SBK13, VLTP14, XGZ15, XY10, Xin10, pZ13, ZZS13, Arv18, BBHN18, CE19, CF19, CL17b, Che17b, DLH^{+}18, DZL18, KYW16, LW17c, LL19c, LCA17, MCGMS17, SRN17, SV17]. rating [BKKP19]. Ratings [CG14b, BH17c]. Ratio [BSY12, Chi09, CM13b, EE15b, GAK14, HTN15, HS13a, KPZ16, Lu11, Mae10, MKK17, OJR15, ÖÇO14, PSS^{+}15, QL19, RS14, Rub10, SS13c, SAC13, TC14, TTC17, Tak14, TN15, WXPL13, YUSC13, YUSC14b, AE17a, AE17b, AEHQ17, ABK17, BS16a, BP18c, Che16c, CLL17, CE17c, DDP18, EEE16, Hua17a, KP18a, KZ19, LKK17, LWZ18, LTS16, MNM19, RM16, NKG16, mNK17, Par17b, RK18, SG17c, SSS16, SP17a, SPY18, SLSC18, SS16i, Tar17, Vis18, Wan18, XX17, YX17, YX18, YH19, ZB19]. Ratio- [MKK17, TTC17]. Ratio-Cum-Product [TC14, LTS16, RM16, SSS16]. Ratio-Product [YUSC13]. ratio-ratio-type [SPY18]. Rational [Cho10, Ko15a]. Rationally [sK12a, sK16c]. Ratios [AE13, Bur15, KRT13, MCY10, Ste11, Tho14, TP14, Bur17, SG17g, ST17a]. Rayleigh [BP17, IVVG17, ICBG18, JYX16, JY16, KTR18, KPS10, Kha12, KG14c, LL12, LG14, MAA17, MML17, PKJ14, PG18, Qom17]. RCEV [Çel19]. Re [LMCH12, ÇE17a, ST18]. re-estimated [ÇE17a]. Re-Estimation [LMCH12, ST18]. Real [HT13, KSDR15, aYwSlDmS17, GST17]. Real-time [aYwSlDmS17]. realization [ZYH18]. Really [YSHL11]. REBMIX [NF11b, NF11a]. Recapture [WBH10]. recipe [Key16]. Reciprocal [RR09, SS12e, SSR19]. reciprocity [LD17]. Recognition [Ano10b, TN19]. recommended [Hot16]. reconciling [WK19]. Record [AAR11, BAT15, AC16, BBA16, NK19, PT17, QK17, SAD16, SK17, SBA17, TD18, TP19, ZCCW18]. record-breaking [SAD16]. Records [AB14, AF10, EWRZ14, GLS12, KJD^{+}15, Ko11, AB19, BAB^{+}18, CM19, Jas16, Jas19, KQ17, MT14, MA14, Wan17b]. Recoveries [BCCH14]. Recovery [KvdVvZ13, Mat14, Mna17, UU11]. Recruitment [Ani11, LZY^{+}17]. Rectangular [AJ11]. rectifying [NN18]. Recurrence [NHN16, NA17, ZWZZ15]. Recurrent [Han12b, LSZ13, ZZS13, ZWZZ15, Fan18, GZ16, HHS18, OAOQ16, SLOT17, She18b]. Recursion [Mar15]. Recursive [BZAZ15, BLT16, LWY10, Nge11, Sla17]. Reduced [GMP11, Por11, Alm18, IK18, KD17b, SG16c]. Reduced-Bias [GMP11]. Reduction [FAS11, Gil12, LG14, YL10, ZYX12, Zhu10, ZYZ11, CS17, DLS17, DBG17, God16, HJW19, KNH17, SS16e, SA17b, XTS16, ZZ17c]. Redundancies [LYH11, ZL15, CZZZ18, YL18a]. Redundancy [LWZ13, YLL19, YL18d]. Referee [Ano10b]. Reference [AT15, GDDC19, Gri15, CM16b, CTAL17]. Refined [GR18a, BOT17, OOT19, OTA17]. Reflected [EB12, QNP12, LW18b, ZZ19d]. Reflections [CCZ15]. reformulation [ZZJ19]. Refusal [BBP14]. Regarding [MC14, RL09, NM17, NM18, SG17d]. Regime [DYW18, WW10, XLPL12, Cav17, FSSW17, GDW18, HZ17b, XDW19]. Regime-switching [DYW18, FSSW17, GDW18, HZ17b, XDW19]. regimes [OH19, WLC17, ZDDH17]. Region [AKL15, CC15b, Ko14a, UC15]. Regional [BHKHM13, Ko15a, sK12a, sK13b, Ko15b]. Regions [BL15, Can15, CGL13, HM15, sK12a, Pre11, XZD10, BKL19, Hua17a, sK16c, LH17c, MP17, PF19, QYLL17, YCH17]. Regression [Abd15, Aga11, AKL15, AK14b, AARN10, AAM14, APA14, ASMZ14, AJN15, AG11, BA10, BC12b, BBB10, BS15, BB12, CYE15, CO10, CY14, CS14a, CS15a, CKLS11, CG11a, CHH15, CTT15, CHLP09, CYL13, CYY14, CZ15c, CS14b, CA12, Cho10, Chv13, CR14, DNL12, Den13, DZ15, DG11, DZX13, DLP15, DI10, DD13, ER11, Fam15, FV11, FH13b, FP12, FH12, FG11, GSY10, GT15, Gil14, GL11, GR15, HS15b, HKLW12, Hol15, HL11, HS15d, IT12, IS11, JS13, KYS13, KMS13, KMSS14, Kha11, KCC14, KS15b, KCS12, KCY^{+}19, KE10, KS12b, KS12c, KLS10, KAD14, KP12b, LGV13, LK14, Lem14, LLC10, LY10, LX11, LYX11, LY11a, Lia11, LL14, LW15, LWR11, LG11, LGX12, Lop11, LM18b, Lu13b, ML12, MS11a, MSS14a, MSS14b, MKS15, MS14a, Mao15, MP14, MFS12, MH11]. Regression [MS12a, MT12, MM15, MX15, MW10, NBS13, NO12, NAMD14, NK15, Nos14, NIN11, ÖA15, ÖD10, Par13b, Par15a, QZY11, QQL14, RSV10, RVD14, RM10, RAG12, SE15, SGM11, She15b, SS15a, Smi13, SLSY14, SSM12, SZD12, SZL14, TOT15, THHL15, TTZ14b, Tsi10, Tsi14, UPA13a, UPA13b, ÜY13, Van16, VRC13, VV12, WDZ11, Wan13b, WW13a, WS13, WG15, WWT15, WH15a, WT15, WTZS15, WSlT15, WW15b, WXS18, WY13, zXyL13, Xu12, XY12, Xu13, XY13, YC10, YHA10b, YXC11, YC11a, Yan12, YY14, YS15b, YC11b, YW13, ZI12, ZI14, ZK13a, ZSK12, ZY14, ZZLL15, ZSY13, ZW09, ZS10, ZL14b, Zhu10, ZZSP14, ZLWZ10, ACO^{+}18, AB17, AM17b, AXLJ17, Alg19, AAO17, AW16, AEA17, ALR19, BD16b, BH17a, BD17a, BD17b, BG16a, BSS19, CYF^{+}18, CH19a, CBSN16, CW18, Cha18, CS16c, CCN17, CW16, Che16c]. regression [CKS17, CN17a, CT16, Che19b, CGC17, CA18, CE17c, CBOR18, DYXW19, DDP18, DDD17, DA17b, DGMC17, El 17, El 19, FAT17, FTT16, FZZ16, Fan18, FK19, FI17, FDL17, FWZ^{+}17b, FJS^{+}18, FRC16, GAS17, GGO17, Gro18, GK19c, HL18a, HBM16, He18, HW18a, HJW19, HWH18, HLY16, HWY17, Hua17b, HJY18, Ima17, Isl17, IKH19, JXW17, JQZ18, JHM19, KD17a, KKS19, Ker17, KY17b, KLKK17, KDG^{+}18, KH16, KL17b, KS18b, LZY^{+}17, LP19, LM18a, Li16, LY16b, LSL16, LX17, LCG18, Li18, LD18a, LX19, Lin16, LYR^{+}18, LL17d, LY16c, LY17, Liu17a, LWD^{+}18, LMdO18, LWHB16, MSK18, MBS^{+}19, MMN16, MMP18, MATC18, MN19, MG19, MFFMA17, MMP17, MD17b, Mur16b, NC19a, OW19, OE16, OOT19, PDdHNMR17, PY16a, PK18, PCS18, Pop16, PK17, Qin17, ROM16, RA16, Sab19, Sch17, SLF18, SNW19, SS16b, SSH17]. regression [Spa19, TLT17, TT17, TK17b, WXZ16, WHC16, WP16, WL17c, WH16c, Wie18, WB18, WF17, Wu17b, WAA18, WCF18, XQS19, XW16, XHL18a, XX17, XQ17b, XTT17, Xu16, XN19, Yam17a, Yan18, YS19, YGL16, YLG16a, YYY17, YWH17, YGM19, YJSL19, YWLW18, Yil18, YCL17, YD19b, ZB19, ZKS18, Zen17b, Zen17a, ZDZ17, ZZ18b, ZWNS18, ZZ18c, ZZ18d, ZDW17, ZP17, ZDS16, ZL16f, ZFD17, Zin18, dPD16, vZ17]. Regressions [GH13, LF10, WH15b, ZYZ11, ABES18, LP17, VSN16, ZKS18]. Regressive [MYM11, JDSS16]. regressive-spatial [JDSS16]. Regressor [LMR13]. Regressors [XWJ10, FP17b, SB16a]. Regret [NS15b, Xu13, Xu17a]. Regular [Bas14, CG14a, CJ15, DKY16a, Hay10, HLL15, Koi10, Lu14, LS15b, MM15, MW10, Tan10b, CZ19a, LMH17, Lu16b, MH16a, OB16, ZYLZ17, ZLYZ17, ZL16e]. Regularization [TPTT14, AD17a, ZZ19b]. Regularized [EY10, KK11a]. Regularly [HLX14]. Reinforced [DP14]. Reinsurance [ES11b, LLSM14, CPW19, FWLL19, HW18b, LRZ17, MWY18, YWX17, ZYH17]. reinsurance-investment [LRZ17]. reinsurer [FWLL19, ZYH17]. Rejection [LS11, Sco11]. rejections [MR17]. Relapse [KC12]. Related [Cha15, SWW^{+}12, TAY11, Wat11, AA19, Ayk19, BPV19, FFN^{+}19, Hom17, OL19a, PL17, VPMV16, WQ18, Yam18a]. Relating [CVC^{+}12]. Relation [BREA11, Lec11, GR17]. Relational [APM10]. Relations [BS12b, SZL14, ZC11, NA17, PHY19]. Relationship [BBG18, MCY14, Tsi10, Nag18, Uya17, ZMV^{+}17]. Relationships [Dom11, Tia12, Vij11, Ayk19]. Relative [Dji10, OMB17, RGI15, Tee12, Tsi14, WH14, WZ15, WWH15b, Zha15a, Che16a, Con16, DWW18, HN16, HW18b, Ko17a, Meh17, YYS^{+}18, ZZL16, dAdLA18]. relatives [AR17c]. Reliabilities [DLC16]. Reliability [AMGK13, AMGK15, BC12a, Ben14, Ber12, EI11, Ery12, Gup15, Gup16, Han12a, Hou12, JCY10, JC12, Kam14, KS14a, KN10, KN18b, LH14b, Lin14b, LSCW17, Mea15, MV16b, NSP12, NRW15, PSF16, Pao10, RSA19, RAK15, SWW^{+}12, SK15a, WCZP16, Wan17b, WYL12, WZT15, YGWQ17, ZZA12, ZM13, ARB19, CR16, CK17c, CM19, DMA17, DZH^{+}19, FK16, FN18, Han17, IS17, JZ17a, JSWL17, KS19, KG17, Moh19a, MIG17, NS16b, RAA17, RAM16, Sha18c, Sre18, TE18a, YX17, aYwSlDmS17, YYLM18, You17]. Reliability-Based [Ben14]. Reliable [DCL12, cTtL14]. Remainder [NNZA18]. Remarks [AK14a, BK17b, Mül12, YHA10b, KS16b]. removal [LSCW17]. removals [BSL^{+}19]. Removing [SÖ12]. reneging [KS18a]. Renewal [HKR15, KRH12, MS10e, SI13, AAK16, FS17, GCW19, GWZ17, Li17b, LGG16, LYG17, XX18]. renewal-reward [AAK16]. Renewing [CC10d, JHP10]. Renyi [AA10, AC16, DWX17, NY17, SK19b, Too17, JY14].

Renyi-type [DWX17]. Repair [CSC13, CDG12, DJ15, Fal14, KY10a, HWZ19, JX19, KC19a, KC19b, LCWG16, MAA17, QC19, RG16, YFV19, ZW17a, ZW17b, ZW17c, ZW18]. Repair-cost [CSC13]. Repair-Time [KY10a]. Repairable [BC12a, CGSM10, DCG13, DSC^{+}14, CVL19, Ery16a, LCWG16, QC19, VM17, WCZP16, ZW17b, ZW17c]. repairman [LCWG16]. Repeatability [DZ14]. Repeated [DS09, EMI14, ITG09, RL12, SAC13, WL14, ZLD16, ASR^{+}19, BAT^{+}18, CK17a, FZ17b, FZ18, HCW18, JKL19, KC19a, KC19b, LCZA17b, LCA17, NNvRS17, NVS19, RAST18, RRRA18, SLL17, XQ17a, YZFM17]. repeated-measures [FZ18]. repeats [ZCC16]. repetitive [AO18a, JBA18, YCAJ18]. Replacement [CSC13, CC10b, CC10d, KY10a, NN11b, SCC11, ZMN17, Cha16, Che19a, Moh19b, PYL17, PKM17, PM18, ZW17c]. Replicate [MLGZ14]. Replicated [UC15]. Replication [HB13, GL19]. Replications [NC14]. Reporting [Bic19]. Representation [ABLR15, CLX14, DO13, HP14, LN13b, MS10c, NN15, BCCR18, Car19, Tav16, XY19]. Representations [MS14c, Ric13, ZM13]. Reproducibility [DZ14]. Reproducing [Par15a, SN12]. Reproductive [CTW10, XWJ10]. required [DZH^{+}19]. Requirements [MH15]. reranked [Spe17]. Resampling [CS16b, JP13, LM13, Wan17a, WGG^{+}19]. Rescaled [WWH15a, BRAS17]. Research [JC12, KLS12, Wei14, ZW13, YH16]. Reserving [HQW15]. Reset [Par13a]. Residual [AC16, Bar10, BR12, CB14, CW15a, Fak15, HS12a, IRB13, JY14, KRB12, LM14, MB15, NRB14, RGM15, RG15, RSA13, RR13, Zha11, CB17, CNG17, FZZ16, GXB18, Kum17, KP18b, LMM18, MB17b, MBF17, MSR17, NP17a, PBA17, PBA18, Pro18, PE17, Qiu18, SHB17, SK19a, SY17a, Tav16, VS19, ZJZ17]. Residuals [CL15c, UDC12, ÇE17a, Çel17, Çel19, KL17a, KLKK17, LTCQ17, LM18b, NS19, WH16c, ZKS18]. Resolution [Hil15, HYK15b, RSV11, Tin10, YH11, KHY17]. resource [KAR16]. Resources [RE10]. Respect [AB14, HL13a, HX11, ZFL13, CS16a, She18a, SJY17, SS18]. Respecting [ZB12a]. Respondent [MV13]. respondents [RSC16]. Response [Aca15, AAG12a, Arn11, ASN12, AS15, AAG12b, BD11a, BO14, FHS15, GDS^{+}14, GMSD13, KK11b, LW13b, MMP14, MJ15, PP10, Par15b, RMP12, SK13a, SKK10, SK10a, SK14, SMPH15, SKHL10, SHLK10, SK12d, SZD12, TSS15, TWTG14, UC14, Wee12, WJH15, XTG15, YX15, ZZ11, ZS12, AAE17, ASG17, AO18b, AO17b, Arn19, AH17b, BQC17, CS17, CSB19, EM18, GAS17, JSS16, JSSO18, KJ17, KZ19, LP19, LPdDTH17, LHS19, LWZD17, LXT17, Lu16c, MSB19a, Mat18, MBN14, MSEN18, NC19b, OW19, PT19, Roy16, Roz18, Sen16b, Sen16a, Sen17a, Sen17b, Sen18b, SGA19, ST16d, SKM16a, ST16f, ST16g, SKM16b, SG17f, SG17g, ST17a, SSB17c, SSS17a, SG17h, SMK17, SKS17, ST17b, SKM17, SK18b, SSS18, SSKP18, SSSK18, SG19, SSS19b, TS16, TS17a, TS17b, TS17c, TS17d, Wu16a, WXZ18, XQ17a, YSG17, ZF18]. Responses [BBG14, BM15, BB14a, CDZ14, Fu12c, Kat11b, Kha12, LW13b, QL11b, SG14, THHL15, Wei12, AO18b, BM16, zF16, GX18, HXF18, KGH19, LLZ17, OD19, RSG17, SG16b, Sen16b, Sen16a, SG17e, TSG19, XTS16]. Restoration [OP11]. Restricted [AK14b, CHP14, HM15, LY10, LLY16, LGX12, NC14, OM13, Piz13, Pre11, SGM11, Šir14, WY13, XY11a, YC11a, AW16, AEA17, BCP17, GOLM19, Hu16, HWY17, KD17a, LY16b, Li18, Ren16, WXS18, Wu17b, WA17, WAA18]. Restriction [BAT11, CZ19c, FAT17]. Restrictions [Bru10, PC12, Tia12, YW11a, YY14, Kaç19, KAT18, SJJ19, TK17a]. restrictive [MDC16]. resubmitted [MB19]. Result [ABC12, Mel13, SB16a]. resulting [SGKB17]. Results [AB15, BHB15, BSCS11, CC10a, FJ11, Gue14, IAA15, LLV10, Li14b, Mah12a, MWM13, Pap13, RR13, SBR14, SS12e, WZ15, YL15, ZM13, AABL17, AER16, BNB18, BNB19, BCL16, BD17b, CF19, CB16, DAS16b, Had17, KSDB18, Kum17, KP18b, Li17c, LL17b, LL17f, NM18, OL19b, PC16b, Qiu18, SS16b, SI19, Tch19, Udo18, WHV16, WCW17, YL16, ZZ18e, dAdLA18]. retention [KS18a]. Retrial [Dra13, SLOT17]. Retrials [Dra13]. Retrospective [GLY17, XLZ^{+}12]. Returns [GW15b, LPS11, SN18]. Reversal [MS14b]. Reverse [Jac11]. Reversed [Dom11, IRB13, MB15, RGI15, BBHN18, HP16, HB17a, HP17a, HB17b, SRN17]. Reversible [BJ19, ÖD10, Sze17]. Review [ZL14a, ZN18]. Reviewed [GGL14]. Revised [Lu10, CKLC17, LWZ17]. Revisit [KCC14, SSS19b, KJ18]. Revisited [Kon15b, MO18, Spa19]. Revisiting [CDE10, MC19, PLC17]. revival [Fab16]. reward [AAK16]. Rewards [SMS14]. Reweighted [DTO11, DBA18]. Ridge [AK14b, Alk10, APA14, AJN15, BAT11, Gru12, Hol15, KMS13, KMSS14, KCC14, LY11a, LL14, LMR15, LG11, LGX12, LP17, MS11a, MSS14a, MSS14b, NO12, RAG12, SG15, TK17a, UPA13b, Xu12, XY13, YY14, ZSK12, Alg19, FAT17, FRC16, GR17, KD17a, LY16c, MSK18, RA16, WA17, Yam17a, ZKS18, Zin18]. Ridge-type [LL14]. Riemann [Dai15]. Riesz [HL13a, TZ17]. Right [DP10a, DP13a, DPP15b, Hua11, Par15a, She10a, She12a, YSHL11, ZLD11, CSC17, FGL18, GFL19, NB16, She16, sSW18, SSR19, Wan16d, WH16a, ZZ18a, ZJZ17]. Right-Censored [Hua11, Par15a, She12a, YSHL11, CSC17, FGL18, GFL19, She16, sSW18, ZZ18a, ZJZ17]. right-skewed [WH16a]. rises [Kon19]. Risk [BCCH14, CGW15, CPW19, Chi09, DMS14, ES11b, GJ15, GB15, HK15, HJ14, HZ17b, JRM10b, KM18, Kha11, KY10b, LD14, LDY15, LLQ13, LLSM14, PP12, PHW11, PK11, SW14b, SK10b, WW13c, WH14, WJY15, WZ15, XLL13, XGZ15, XLZ^{+}12, Yan14, Yan15, Zha15a, ZYN11, ZZL12, ATO19, BR19, BSGS16, CRTdC17, CCCT16, CXC19, CH19b, Con16, DLH^{+}18, DTT^{+}16, Ery18, FG17, FS17, FL17, FL18b, FL19, GWL16, GCW19, GWZ17, IF18, JN18, KOFA18, KL17b, Li17b, Li17c, LSY17, LG16b, LGG16, LYG17, LWG18a, LWG18b, LWH19, Lou19, LSL19, LO17, NCA17, NSR^{+}19, OMB17, QWWW19, She18a, SQCL19, TMZ^{+}17, Wan16b, WCW17, WRZ19, XX18, hXZ17, XZ17, XYL14, YZ17a, ZN18]. risk-adjusted [KOFA18]. risk-minimization [QWWW19]. Risk-minimizing [HZ17b]. Risks [Ko14b, SS12d, WY12, WW11, XT11, Yan14, Yan15, ANJ17, AN17d, BSL^{+}19, Bel17, CSL17, CK17b, FZZ16, LXL17, LL17b, LS16a, LS17c, LWG18a, LWG18b, SS16a, SDS17, SX18, SLSE19, SSD19, WXS18, WSS17, ZZ18a, dAdLA18]. Risky [LPS11, GWP19]. Ristić [CB16]. rival [Say17]. RJMCMC [DM10]. Robbins [AD17b, GFS16, RD16]. robot [FN18]. Robust [ÁF16, AMH14, BC12b, Bis12, BDM13, BDM15, CYE15, CL10b, CZ19a, DYXW19, DPA10, DPP15a, ES14, EP11, FQZ14, FH13b, GGLV14, GH13, GYL17, JN18, Kim15, KZW09, LMG13, LZZL17, LL18a, MEZ16, Mao17, MR16b, MBN14, MX15, SK18a, SA17a, SSS14a, SLZ18, SV11, TN15, TB13, WLS18, Wu13, XY11b, YLG16a, YLG16b, YZ18, YK11, ZGK17, ZYH17, Arv18, BMRSE17, BBR17, CSM16, FCSH16, GKP16, HJC17, KKS19, KP19, KH18, Man17a, MSEN18, OPK18, OH19, RAZ19, Ric17, SA16, SSY18, TL17, YL18b, YGM19, ZB19, ZX18]. Robustness [AB09, AAG12a, ADO13, CG12, Che16a, DTE13, HK13, Kha11, LH11b, VRC13, YDC15, CSM16, Kha16, Wan19a]. ROC [CM16d, CM17, DTE13, WNL18]. Role [BC14, CKD14, Hun14, Ram17, ST16f, SA14, VT17]. Romig [YF17]. Root [Kon15a, LS11, QTD10, FLY16, Hay17, MR16b, PLC17, Sen18a, WYY19, XP18, Zha16a, Zha17c]. rooted [HYS17, SY16, SY17c]. Roots [BB12, dBCR15, DPS11, ZS17b]. RoPLS [TB13]. Rosenblatt [YLW17]. Rosenthal [FTF17]. Ross [CCZ15, FNM19, RS14, SST14]. Rotatability [KP12b]. Rotatable [DPA10, DPP15a]. Rotation [SK10a, SSP^{+}12, SK12b, SK14, SMPH15, SMMS15, PM16b, SP16c, SKT16, SP16d, SSS17b, SP17c]. Rounded [Kac14]. routing [Lin14b, MV16b]. Row [VJV14, CLLW16, DJVV17, GPM18, VV17, Wan17d, WW19]. Row-column [VJV14, GPM18, Wan17d]. row-wise [CLLW16, WW19]. Rows [KLS12]. Rowwise [SWLW14, WHVH13, MWW18, SS17b, Wan17f, YJS19]. RRT [MA18]. Ruin [BS12a, GW15b, HJ14, Kon11, LDY15, PHW11, XLL13, Yan14, ZYN11, BSW15, BR19, Ery18, FL19, GWL16, Li17b, LGG16, LWG18a, LWG18b, Sun19, XX18, hXZ17, ZWY16]. Rule [CHP14, jK15, LR10a, SLZ11, ÜY13, XY12, XY13, BCP17, CKLC17, LY16c, OB16, VM17, Zha17d]. Rules [Hou12, KS14b, Lee13, LR10c, Mah12a, RL12, YHS13, GFM17, KCL^{+}16, KHC14, MR17, WCK16]. rules-the [KCL^{+}16]. Run [AMR14, BQDH12, Cox15, Lee13, LZH15, WCK16, BVJV17, FY17a, GFM17, LK17c, LK17b, RSA19]. Runs [HCS10, KS14b, Kon15b, KHC14, Mah12a, MP15, MZ14, SH11, Ste11, AMP18, CKLC17, KCL^{+}16, Kon16, Kon19]. Runs-rules [KHC14].

S [BT13, Had17]. s-increasing [Had17]. Saddlepoint [AE13, AE17b, AAE13, CG11b, AE17a]. Safe [BHB15, FTF17]. Safety [SCC11, LJC18]. Sahalia [JZ17b]. Sahalia-type [JZ17b]. Salary [DDGM13]. Salinity [RGM15]. Sample [AM15, BS12c, BDKM12, BEZ11, BSY12, Bur15, CY14, CG11a, CS12, CB10, CP14, CM13b, Din15, ET13, EDAAS12, Gue10, Him10, HL13b, HP10, HS12b, IP12, KL10, KPS10, Kha12, Kha13, KNK13, sK12a, Ko14a, Ko14d, Ko15a, Ko16b, KL12, Kum15, Kvå15, LMCH12, LG10, LF10, LL17f, LCZA17a, LCZA17b, LCA17, MS15, MH15, MS11c, NAD15, SG10b, SL14, Tar17, TS12, Tsu14, WWH15a, Xu10, YUSC13, YUSC14a, YLS10, YWH14, YMF15, YA10, Zil11, Ahm17, BOT17, BG19a, BC15, CBSN16, CC16c, CL17b, DDP18, DAS16b, DW18, FGP16, FLL17, Gau17, HN17b, Kar17, KM19b, sK16c, KPZ16, KX18, LPDD18, LTZ19, MN19, Mar16, MS17b, MAA17, MK19, NHN16, OL19a, OH16, PA18, Par17b, PFD19, PP17, Pou17, PK17, PT16, Sha17, SBA17, SR17, SP17b, SPS17, ST18]. sample [SW19a, Sze17, Tak14, TWTG14, Wan16c, Wan17a, WYY16, WW17b, WZZ17, Wu17b, XY19, Xu16, YYS^{+}18, Zan17, ZBHW18, ZSJ16]. sampler [Van17]. Samplers [JB15]. Samples [Bak11, GZSS14, LR10a, LQ13, LCP^{+}14, MZ15, NH15a, NGT13, PL12, QL11a, SK15a, SYP13, VB12, WJ11, XYL14, YCL14, ZPZ10, AJ19, BN18a, BKL19, DYL17, EDT18, FA19, GN17, HOK17, Kal17, KAW^{+}18, LQ17a, LB16, WdX17, YYY17]. Sampling [AC15, AARN10, AAR11, AS13, AYC^{+}12, BAK13, DM11, DCL12, ET13, ES11a, EF11, Fak15, Fre14, GÖ12, GB15, GS15c, HS13a, HASA10, HD13b, JFS14, Kaw13, KSG13, LCT12, LCLJ14, LC15, Mah12a, MS13, MADSAS10, NN13, NGT13, NAD15, OP11, ÖÇO14, RN13, SAS11, SG10a, SK13a, SRT14, SKK10, SS13c, SMPH15, SMMS15, SS14b, SS15c, Sub14, TC14, TL15a, TS12, VB13, Wys19, XLZ^{+}12, YUSC14b, ZAV14, AO18a, ANJ17, AHH16, ARJF19, ABK17, AO17b, AAJ16, ASeMJ17, AAJ17, BOT17, Bak17, BU17, BAL18b, BJU18, Bal18a, BAA^{+}18, BS16a, BS16b, BAW19, CY17b, COK17, CSB19, Cle18, CLZ19, EW16, EAE18, EDKAEN17, EDT18, FY17a, FF16, FF18, HBMAO16a, HBM16, HBMAO16b, Haq17, HBAJ17, IF18, JRF17, JBA18, JKJ16, JR18b, KSF17, KGH19, KS16b, KAJH18, Koy18]. sampling [KZ19, LW18a, LP19, LK17a, LK17d, LK19a, LDL18, LTS16, LTV17, MZ17, MDG19, MB19, MZ18, MSB19a, RM16, MF16, Moh19b, MRA16, MSK17, NNZA16, NNZA18, NS17, NN18, NS18, NSSS17, OOT19, OTA17, Ozg19b, PB17, PT17, Pou17, PT16, PM16b, PT19, PB18, QKCH19, RSC16, SKK18, SAD16, SSS19a, SCYR17, SCYR19, SFSH16, Sen16b, Sen16a, Sen17a, SG17a, SG17b, SG17c, Sha18a, SGA19, SH17b, SP16c, SKM16a, SSS16, SS16c, SM16, SP16b, SS16d, SKM16b, SHM16, SSS17b, SSB17c, SMK17, SKS17, SP17c, SKM17, SK18b, SSS18, SMP18, SSKP18, SSSK18, SG19, SSS19b, SS16g, SS16h, SS16i, TS17e, UB17, WT18, WHZ17, WAJ17, YF17, YCAJ18, ZWZwZ17]. Santos [AE23c]. SARMA [Ami18]. satisfaction [BCB19]. Satisfying [Yar14]. Saturated [WLL13, PYL17]. Saunders [AALS12, CLC^{+}16, CCOB16, KLJP16, LF10, MFFMA17, OMFB17, UM09, ZHZA18]. SAVE [Vel14]. SB [LMG13]. SB-Robust [LMG13]. SCAD [HZ18a, WSlT15]. SCAD-Penalized [WSlT15]. Scalar [CM13b, Sha19]. Scale [ADO13, CT14, FGTS12, GR11, JLV13, KT15, KP12a, MR15, Ng12, NY11, Pet10, PT15, ST10, Saj11, SV11, Wan12, WXZ13, WQ14, XFC11, YLS10, Bob19, CR16, DZL18, DL16, GCA18, GKP16, Gun17, HBMAO16a, Jon16, KK17a, KG18a, LW17a, LW18a, LKK17, LM18b, Mal17, MM17, MA14, MR16a, PDdHNMR17, RSA19, RPM18, SM16, TKP17]. scaled [OPK17, OPK18, WL17a]. Scales [Mar12, YTX14]. Scaling [HT10, GFS16]. Scan [CG16, Lin14a, RP12, SA15b, WG14, ZG17b, BT16, Cuc16]. Scanning [JB15]. Scatter [MCY10, YDC15]. scenario [CH19b, LWH19]. schedule [JX19]. scheduled [CC19]. scheduling [Che19a]. Schematic [PYL17]. Scheme [Kor11, LJ10, LCT12, Sag15, SKK10, SZ13, AAK17, BIJ17, CK19, GFM17, GP16b, KSF17, MK19, MRA16, SG17b, SSR19, UK16, WL17a, ZWZwZ17]. Schemes [AAR11, CR14, Sub14, HBM16, KS16b, SAD16, Sha19, Udo18, vZvdM19]. Schnabel [GG13]. Schur [Yar14]. Schwarz [GL10b, Sha10b]. Sciences [MA11, Mar16]. Scientific [Che10a]. Scope [Zhu10]. Score [Bea15, CHH15, DZ15, Fab16, KS15a, sK11, KSBM14b, KSBM14a, Lan14, LF10, MO10, NN11a, SK12d, YHA10a, BH17a, GGtH^{+}19, LCW^{+}18, WER^{+}19]. Score-Based [KS15a]. Scores [KAD14, OFS13, ZZ19a]. scoring [LMdO18, Zen17a, ZZ19a]. scrambled [ASG17, KGH19, PT19, SG17e, SSKP18]. Scrambling [GMSD13, HAZ16, SG17g, ST17a]. Screened [Kim12, jK15]. Screening [BAK13, GAK14, SNN12, Sty10, Web11, EIM18, She17, SLSC18]. Search [SP17c, TE11, LZ18, SSS17b]. Searching [Cha14, PM16b, TE11]. Seasonal [dBCR15, KCS12, LS09, SA15a]. Second [AAG12a, Aka10, AHKO16, AO17a, ADWR11, BBC13, CZ15b, DPA10, DPP15a, GMRN13, GD12, Han15, HM15, ITG09, Kak10, Kak12, Kát11a, MH16a, NC14, PP10, SCSMN11, Tak10b, TOT15, WL11, Yan15, ZLD11, CMN^{+}18, CS17, CLW17, CLL17, DPP17, GMR16, HYS17, KT17a, KD16a, LMH17, NE17, YSC17b]. Second-Hand [SCSMN11]. Second-Limit [Kát11a]. Second-Order [AAG12a, Aka10, ADWR11, BBC13, DPA10, DPP15a, GD12, Han15, ITG09, Kak10, Kak12, NC14, PP10, Tak10b, WL11, ZLD11, AHKO16, AO17a, CZ15b, HM15, MH16a, Yan15, CMN^{+}18, CS17, DPP17, HYS17, KT17a, LMH17, NE17]. secondary [dCA19, Haq17, LW17a]. Secretary [KM14b]. section [ZL16c]. sectional [Mao18]. Sections [DDÚF11]. Secure [Gup14]. Security [SI13]. See [SS13a]. Seemingly [RAG12, ZSK12, ZKS18]. Segmentation [UC14, WLS18]. Segmented [VV15]. Segments [GS15b]. Selected [AMV15, ARB19, KO10a, AN17a, BG19b, IS17, KG17, LWHB16, MAG18, MT16, NP17b]. Selecting [AM15, BAW19, BCP17, CHP14, KYS13, KS12b, SJS15, YCL14]. Selection [CT13, CT10, CHP14, Con12, DCG13, DLP15, FQZ14, Fu12c, GGLV14, GLP14, GHY11, HS12b, KJL11, Kim11, Kim15, KH16, KS12c, MJ15, MT12, NMJY15, OFS13, QNP12, RZ13, Say14, SA12, SJS15, ST13, SW14c, SK10b, Tak10a, Tra11, cTtL14, WS11, WLC11, WL15, Woj11, YX15, AXLJ17, AO18b, BCP17, CH18, CA18, DYXW19, DE17a, DE17b, DGMC17, EMH16, EM18, GYL17, HSW17, HZ18b, HJT19, KKS19, Kar17, KDG^{+}18, KH18, LPDD19, LPHC18, LS17e, MST19, NS19, OH16, OE16, PS18, Rej16, SI17, Tak18, WX17, WYY19, WXZ18, Xia17a, XQS19, YS19, YLG16a, YLG16b, YL18b, YWLW18, ZCHR19, ZZ19b, ZX18, ZZ18d, ZP17, ZGK17, ZL16f]. Selections [FM11]. Selective [MS10b, Mat11, ADKZ16]. Selector [WSC13, BAKZ18, Ten17]. Self [Gar12, MR15, MSP12, hCXxShY18, GR18b, Kal17, LPH16, LTC17, Mal17, WWWX19, Wu17c, Yu19, Zha16b, ZCC16]. Self-Exciting [MSP12, WWWX19]. self-intersection [Yu19]. self-normalized [hCXxShY18, Kal17, LPH16, Wu17c, Zha16b]. self-normalizing [GR18b]. self-overlapping [ZCC16]. Self-Similar [MR15, Mal17]. self-similarity [LTC17]. Self-Weighting [Gar12]. Semantic [ZLDT14]. Semi [AKK09, AK14c, BG11c, CJ15, DDGM13, DGM13, DMS14, DT12, GL15a, Hu14a, Hu14b, Hua17b, HJT19, KK11a, KR14, LMR13, Lim11, Muk10, Pap13, PV14, SMS14, UU11, Vas13, VLTP14, WCZP16, ZL12, ZZLL15, BG16a, CO16a, DPP17, Geo17, KKKÖ17, KD16c, OE16, XHL18a, Yeh17, ZSS18, ZZL19]. semi- [Yeh17]. Semi-folding [CJ15]. semi-functional [ZZL19]. Semi-Markov [DDGM13, DGM13, DMS14, DT12, GL15a, Lim11, PV14, SMS14, UU11, Vas13, VLTP14, WCZP16, DPP17, Geo17, KD16c]. Semi-Markovian [AKK09, AK14c, KKKÖ17]. Semi-Metric [LMR13]. Semi-non [Hua17b]. Semi-Nonparametric [BG11c]. Semi-Parametric [KR14, ZL12, Hu14b, BG16a, CO16a, OE16, XHL18a]. Semi-Sequential [Muk10]. Semi-Supervised [KK11a]. Semi-varying [Hu14a, HJT19, ZSS18]. semicontinuous [DR17, SGKB17]. semidefinite [DLZ17]. Semilinear [Fan15]. Semimartingales [LCLJ14, YLW17]. Semiparametric [dCA19, BJC16, CTW10, CTT15, CW15a, Den13, DLP15, FH13a, GSY10, HZL16, HJY18, LW13a, LWZ18, Nar13, RAG12, She10b, SS12d, SZD12, WK12, WW15b, WL15, YLS10, YX15, YDL14, ZBTL17, ZZ18c, ZT18, ZX10, ZZ10, ZXF17, CW16, DCZL17, DR17, HZ18b, LLWK18, WXZ16, WL17d, XQS19, XLS19, XKW19, ZCHR19, ZL16c]. Semivarying [Hu11a, Hu14b]. Semivarying-coefficient [Hu14b]. Sense [AB14]. Sensing [MN15]. Sensitive [AS13, GSSCR12, SK15b, TWTG14, ASS18, BSH17, GK19a, HAZ16, LHS19, MA18, PT19, Sen17b, SG17e, ST17b, SSKP18, SSSK18, SSS19b, SSS17c, TS17b, TS17c, YSS19]. Sensitivity [Ban13, Chi09, Chi12a, KHM12, RS10, WQ14, Zha14, ZT13, FKR16, HAZ16, LPDD18, LPDD19, LWWR18, MA18, RSA19]. Sensor [Gup14]. Separable [GS13, GS15a, LR14]. Separate [ÖÇO14, LTV17, SE17]. Sequence [EWRZ14, Guo14, LPLW13, WLHW14, YWH14, BNZ17, hCXxShY18, CZM17, DPZ17, HV17, LQHZ18, LD18b, LQW19, WWW17, zWDD19, ZQQY17, ZL16g]. Sequences [ASAK10, BR11, GM14, HW15, HWW15, LWY10, MP15, MZ14, Pap13, RM10, dXcY12, XY14, BVJV17, DWX17, Kon16, Kon19, LO17, LQYZ17, MR16a, NHV17, SJY17, TWZ16, WG16b, Wu17c, XY19, YZG17, Zha16b, ZMMX17]. Sequential [AW10, BD19, CS11, Koi10, Muk10, QTD15, RSV11, SCC12, SB11, WXPL13, ZW19a, Zha13, ZZL12, Fon17, HD17, Kar17, Kha16, LA17, Liu17b, Liu17c, LL19c, MLD17, RS16, Wu16d]. Serial [KCS12, LWHW11, LLHZ16, LTGK17a, LTGK17b]. serially [Hay17, LMC18]. Series [AP12, AT09, AMH14, Bin14, BR11, CZ15c, Dok12, Don11, El 15, hFmH14, Fu12c, GS15b, Igl13, JA10, Kha11, KRT11, KR15, LM13, LCG12, LYH11, LWZ13, LSC12, ML12, MO10, Mur13, Nas12, OCIS12, PS11, Pop12, SP11a, SSON11, TB15, TT11, TT17, UC14, WS12a, WH15b, WW15b, WILL13, YLL19, ZA14, ZK15a, ZZL12, Ahm16, BNB17, BNB19, BD16b, BB17a, BBS16, CNL19, CZZZ18, CH17, CTAL17, FLB16, FY17b, GS17, Gho17, GRSN18, GLY17, GICRR17, HA16b, HN17a, Ili16, JT16, JZZZ18, JM16, JKS18, KAS17, KAPA19, KAT18, LA17, LW17b, LL19a, LS17d, MHA19, MST19, MSL17, MAA17, Mur16a, NORV19, NRI17, NIN17, PB18, Raj18, RSG17, RS19, RN17, SAD19, SG16d, SK16b, STD18, Tri17, XZA18, YL18a, ZT18, cZzXgL18]. series/parallel [CZZZ18]. server [AS19, BL19, TSOT19]. Service [GMRN13, Kem11, GMR16, KC19a, KC19b, KY16]. services [ZG19a]. Set [AARN10, Bak11, DCL12, DDM10, DD13, Fre14, GÖ12, HASA10, LCP^{+}14, MN13b, MADSAS10, PL12, SAS11, SF12a, VB12, VB13, ZAV14, Bic17, CLZ19, EDKAEN17, EDT18, FF16, FF18, HBMAO16a, HBM16, HBMAO16b, Haq17, Koy18, LB16, MZ18, RM16, PT17, PT16, SM16]. Set-Up [DD13]. Sets [AA11b, Bic12, OFS13, TN13, Tor12, AGS18, Bic17, Kam19, YXZ17]. setting [Ben19, KP19, LLZ17, RG19]. Setup [Bis12, BDM13, DP13a, EI11, RR14, BL19]. Setups [RR13]. Seven [HH15]. Several [AJA14, GGS13, HS12b, LWWH14, MFSAS10, PT15, SASMF13, ST10, SKG12, Wan13d, Yam18a, Ahm17, ESAKK18, Gun17, LW17c, PK19, PT16, SKG18, Sin16, SHM16, SP16d, SSS18, Sou17, TKJ17, ZCCW18]. severity [Gol16]. Shannon [HYS18, MT14, SY16]. Shape [BNS11, CZ15a, KS10a, Bay16, KT17a, Kam19, MC17, Roz18, TPJ18]. Shaped [SSH10, AE17c, JC18, Kam19, MKSS17, NSR17, SK15a]. shapes [JD16]. Shapiro [Fly10]. Shared [HS15a, HS15b, HD15, HP17a, HB17b, HP17b, PBL19, HP16, HB16, HB17a]. sharing [SNN16, ZB17]. Sharp [KR17, Muk11, Ryc11, SASMF13, ZB12b]. Sharpe [ZVG19]. sharpening [HJW19, KN17a]. Sheet [Dai15, SSC18, XML16]. Shewhart [Bi15, HCS10, KHC14, Mah12a, Mah17, RC16, ZG19b]. Shewhart-Type [HCS10, KHC14, ZG19b]. Shift [SK10b, BRZR16, ZT17]. Shifted [Nas12]. shifts [WH16b]. Shiu [GH19, WL17b]. Shock [CR13, YWLY11, BMLY19, PMW17, PMKF17, RMM17, SMSF19, TE18a]. Shocks [CF12, SCC10, WZT15]. Short [BQDH12, FZ10c, LJ12, Web12, ZB12b, CCCT16, Gra17, Hay17]. Short-Cut [ZB12b]. Short-Run [BQDH12]. Short-Tailed [FZ10c, LJ12]. short-term [CCCT16]. Shortest [QK17, KQ17]. shortfall [NCA17, YZ17a]. Shot [Muk10, CS19c, GDW18]. shot-noise [GDW18]. Shrinkage [AKTS13, CK10, CM19, HWH18, HA14, HL13b, Jun17, NH15b, NAML16, Noo16, RA16, CS19a, Che16a, FM16, NO18, Qom17, XN19]. Shrunken [SF12a]. Shuffle [ZYH18]. shuffling [RZHC16]. Sided [Chv13, DW13, GAK14, GN15, Ko14d, KG15, PHW11, SE13, WXPL13, Yen12, AS18, AHKO16, ÇG18, CCZ15, CM16c, CFA17, Gen19, GG16b, GW15b, LWG18b, RC16, Tak14, Wu16d, YCAJ18, YW17b, Zha14, ZSJ16, ZK15b]. Sides [Abd15]. Sieve [JP13, Kom17]. Sigma [Rav19]. Sigmoid [YSD^{+}16]. Sigmoid-like [YSD^{+}16]. Sign [HCS10, KS10b, KHC14, LB16, HK19b, Lu18, PS17b]. Signal [Afs13, BH12, GGL14, MCY11, TN15, ABK17]. signal-to-noise [ABK17]. signaling [WH16b]. Signals [FM10a, PYB^{+}15]. signature [Moh19a]. Signatures [FY17b, EW16]. Signed [KG14b, KN17b, Say14, ZWZ12, BCK17, Che18]. Signed-rank [KN17b, Che18]. Significance [BG11b, DPS11, DPS12, Xu13, AdBPdBP18, Xu17a]. significant [Hot16]. Silva [AE23b]. Similar [Jor13, MR15, VN18, Mal17, SKK16]. Similarity [DI10, DL16, LTC17, NKG16]. Simple [AC15, AMH14, Bar17b, Bas14, CHH15, Chi11, ES14, HS13a, HYK15b, KXX11, MYM11, NKS15, Oll17, Par13b, PDS09, SG10a, SN14, SÖ12, Wan10a, Wit14, Yan12, BER16a, Gra17, HBM16, KD16c, LWWW16, LH17c, LS17c, MSK17, Nav16, PY16a, aSaCN18, SNAN17, SG17a, SG17b, SN18, VSN16, WYY19, ZW17c]. Simplicial [EPGTH15, KRT13]. Simulation [Afs13, Cah12, CR14, Hua14b, RM15, SS10b, SNN12, TN16, BOT17, HA18, OOT19, VPMV16, Zha17c]. Simulation/Regression [CR14]. Simultaneous [AKL15, ASS18, Chv13, HKNL15, KG14a, KJL11, LARS19, NC19b, PS17a, PGF09, Rub10, SKG18, SLSY14, SG16e, WLC17, Lin16, ML19, Sou17, Tu17, UK16]. simultaneously [Par17a]. Singh [Wan13d]. Single [AGR15, BZAZ15, CM13a, Hua11, Hua12, JFS14, JZQ15, KMC10, KMM12, LW15, LLY14, LWHW11, LWH12, LC15, TKW12, WHZ17, Wil09, Yar11, YHC14, AS19, BKL19, CK17b, CA18, DLC16, FTT16, HA18, HWZ19, HZL16, HZ19, LLHZ16, LY16a, LZZL17, LL18b, NA17, RSC16, SK16a, XHL18b, YLG16a, YCH17, YCZ17, YF17, ZF17, ZXF17]. single- [RSC16]. Single-Index [Hua11, Hua12, LW15, LWHW11, LWH12, JZQ15, YHC14, CA18, FTT16, HA18, HZL16, HZ19, LLHZ16, LY16a, LZZL17, LL18b, XHL18b, YLG16a, YCH17, YCZ17, ZF17, ZXF17]. Single-stage [WHZ17]. single-unit [DLC16]. Singly [Pao10, YSHL11]. Singular [AJN15, Kur10, LX10a, LY10, YW11a, AJT16, ARB19, AN17c, Bud16, DKY16b]. singular-generalized-hyperbolic [ARB19]. Singularities [Aoy10]. Sinh [Coo13, LVBS10]. Sinh-Normal [LVBS10]. Sinks [LH14b]. sinuous [TS16]. SIR [EKZ16]. SIS [FP17a]. Situations [BDKM12, CSB19, SK18b]. six [ZXSZ18, Rav19]. Size [ET13, Him10, Hua10, Hug09, HS12b, IP12, JFS14, sK12a, Ko14a, Ko14d, Ko15a, LMCH12, LZH15, MH15, NAD15, TIA11, TS12, Vas14a, YMF15, YA10, ZB12a, Ahm16, BAL18b, Bal18a, BM17, CBSN16, CC16c, FS17, FL18b, GWZ17, HN17b, IBA18, JR18b, KNC18, Kar17, sK16c, Ko16b, LTZ19, LL17f, LCZA17a, LCZA17b, LCA17, Moh19b, OL19a, Pou17, Ric17, Roo18, SMN18, SSSK18, SSS19b, ST18, Tar17, TK17b, Vas16, Wan17c, Xio16]. Size-Biased [ET13, Ahm16]. Size-Dependent [YMF15, FS17, GWZ17]. Sizes [Ano13b, AM15, BS12a, CS12, MS10d, PV14, ASR^{+}19, JKD17, LSY17, LMC18, LWG18b, OMB17, RRRA18, SBA17, ZSJ16]. Skeleton [YM18]. Skellam [AA17b, BCK15]. Skew [AAM14, AVCG10, AVCL13, AGVB10, BG11a, BBB10, Cah15, Can15, GCSB10, KSK15, KA15, Kur10, LSSS14, MM13, MH11, NAA13, PK11, Raz14, RMN16, SD14, VLD15, ZK13a, AM17b, AGS18, ATO19, AJT16, AEN16, BK17a, dSBCO18, DA17b, HL16, HEB16, HLP17, HST19, Hut19, KK17a, KKS18, LAF17, MN17, MFPG16, NY16, OHKK17, RGV16, RN17, Sab19, SSB17a, TN16, TWG18, Yil16, ZKW16, AT17]. Skew- [SD14, HLP17, HST19]. skew-bimodal [dSBCO18]. Skew-Curved-Symmetric [GCSB10]. Skew-Generalized [NAA13, OHKK17]. Skew-Normal [AVCL13, KSK15, MM13, Can15, Raz14, VLD15, AM17b, AJT16, KK17a, KKS18, TN16]. Skew-Normal-Cauchy [AGVB10]. Skew-Probit [BBB10]. Skew-Slash [ZK13a, RGV16]. Skew-Symmetric [Cah15, Kur10, AGS18, HEB16]. Skewed [LVBS10, Pun13, The19, Cos18, GP16a, SG18a, Shu19, WH16a]. skewed-normal [SG18a]. Skewness [KM15, MBS^{+}19, RPJ15, SSS14b, Zoi09, DYL17, LTZ19, NHN16]. skewnesses [MNM19]. SkSP [BAA^{+}18, HBAJ17]. SkSP-R [HBAJ17, BAA^{+}18]. Skull [SU11]. Slash [Gui15, OPSGB12, Pun13, ZK13a, MAAI17, RGV16, RGBG19, TWG18]. slash-elliptical [RGBG19]. Slashed [AGB17, IVVG17, IR19, ICBG18]. slashed-Rayleigh [ICBG18]. Sliced [CKLS11, CL15a, KLS10, CGC17]. sliding [TK17b]. slight [ZW17a]. SLLN [Ko17b, Yan19]. Slope [DPA10, DPP15a, KP12b]. Slope-Rotatability [KP12b]. Slope-Rotatable [DPA10, DPP15a]. Small [CS18, Kor11, LF10, Muk10, NNvRS17, PA18, PC12, PMC13, Rab13, Ryc11, TD14, Wee12, Zil11, DDP18, FP17b, LTZ19, MN19, NVS19, SYW16, SJY17, UK16, Wu17b, Yam17b]. Small-Sample [LF10, DDP18]. Smallest [Sch12, Bob19, BCP17]. Smirnov [BP18a]. Smooth [BCG11, DTO11, WH14, BKR19, Hua17b, ÖGGB16, WL19]. Smoothed [LCZ16, Zho17]. smoother [YJ19]. Smoothing [AP12, Bos15, Bru10, FHZ12, GS15b, Jeo12, LX11, Pie11, CH18, FJS^{+}18, LQ19, Nas18]. Smoothness [Yu19, GICRR17, NE17]. Sobol [LWWR18]. soccer [GDDC19]. Social [MA11, FKNR17, FK19, HCW18, Mar16, SSM16, TSG19]. society [CLMW19]. Soft [VRC12]. software [KAR16]. Sojourn [XZD10]. solar [El 19]. Solution [AG11, Bon14, CDZ14, GCR12, MS11b, ZS12, ZB12b, Bur19, GR18a, RG16, VSN16]. Solutions [Gia14, NY11, ZZ15, LYX18]. solve [MDR17]. Solving [NAMD14, UC14, VSS12]. Some [AA19, AB15, AK14a, APM10, ADP10, AA17b, AASR12, AABL17, AJT16, AG17, BZAZ15, BNB18, BSCS11, BAT^{+}18, Bea15, BCL16, BCDLA12, Cah15, CNL19, CC10a, CF19, Cha10, CC16b, CH18, CdSBB15, DAS16b, DSA16, DO13, Dji10, EZ14, FJ11, FM11, FWH^{+}15, GS13, Gra12, GY13, Gue14, GN12, Had17, HT13, Hud14, IRAB13, KLMD16, KHY18, KS16b, KRT11, KA15, KN10, KP18b, LWWH14, LLP10, Lem14, LLV10, LL17b, Lu14, LS15b, Lu15, Lu16b, MKS15, MC14, MV14, MS15, MC11, MWM13, MLGZ14, Muk11, NSV12, NSR17, NSL14, NM18, OQC17, Pap13, Pes15, PSP12, PT15, RAST18, RAAJ17, RAZ19, RCC12, SYC19, ST16a, Sch17, SBR14, SG10b, SC19, SG17d, SK10a, SS13e, SMMS15, SP16b, SMK17, SKS17, SSKP18, ŠS12c, SS15c, SS12e, SI19, TN10, Tav16]. Some [Tch19, TN13, Vin13, Voi10, WZ15, Wan17f, WCW17, Wei14, YHA10b, YW11b, YWH17, YL15, YL16, ZL14a, ZZ18e, ZC11, dAdLA18, BBG18, BMR17, BG16b, EdSMR14, HS19, HN16, HST19, HS18a, Jas19, KN18a, KG18b, KA18, MSK18, MC18, MP17, NCA16, NHAMG17, Rao18, SHANA18, UU17, VZ19, WHV16, Wu16c, ZF18, ZLYZ17, fZWD19]. Source [Dra13, Yar11, Vas16]. Sources [New11b, WBH10, PS17a]. Space [FP12, FN09, Ham12, LMR13, Mah12b, Par15a, TB15, Vas14b, Yeh15, Aga17a, CW19c, FAT17, Haj19, Rao17a]. Space-Time [TB15]. spaces [Ko18, STD18]. Spacings [AV15b, HV11, PT15, ST10]. Spanish [GDDC19]. Spanned [KQ11]. Spanos [Hod14]. Spare [CC10b]. Sparre [HJ14]. Sparse [AXLJ17, GBC12, MX15, YJSL19, PP17, WX19, XM17b, YA17]. Sparsistent [ZL16d]. sparsity [HTW17]. Spatial [AA17c, BRAS17, Cey11, Con12, CGL13, GSB12, GS13, GS15a, Gho15, HCW18, LIPMF13, PMC10, RGM15, RL12, SA15b, YDC15, ABES18, BPS16a, BT16, CS18, Con16, DA16, Far17a, FMPF16, JDSS16, LX19, PF19, SN17, Sab19, SA19a, SL18, XHL18a, Zha17e]. Spatial-Temporal [PMC10, AA17c]. spatially [Zha17e]. Spatio [DC10a]. Spatio-Temporal [DC10a]. Spatiotemporal [WW13a, WLW15, CY16]. SPC [HC12]. Special [Edi10b, KM15, MSD19, BG19a, Van17, VFP17b]. Specific [BBP14, Ko14a, He18, TK19, VV17]. Specification [CC15b, MN10, Yu12, CC16b, GR18b, LR19, Tch19]. specificity [LPDD18, LPDD19]. Specified [BCH12, BCS10, SCMR15, KS17, Mer16, YWZ18, Zha17e]. spectra [DBC16]. Spectral [HP10, BP18b, FM16, MHA19, WYY16, Xie16]. Spectrum [MR15]. Speed [QZY12, VFP17b]. Sphere [Sei13]. Spherical [KRT13, RP10, OPK18]. Spherically [KQ11, XY13, DE17b]. Sphericity [CM13b, Ahm17]. Spiked [Din15]. Spillover [Otr15]. spirit [HN16]. Splice [NN16a]. Spline [Bru10, Jeo12, LX11, CH18, FJS^{+}18, HNB17, JHM19, LWD^{+}18]. Splines [DM10, Li12a, KN17b, ZDS16]. Split [KMM12, WYZ15, LWWW16]. Split-Plot [KMM12, WYZ15, LWWW16]. Spot [ST13]. Spread [ZLD11]. spreadsheet [Cox19]. Spring [CG14a, Gra12]. SPSS [SLL17]. Spurious [VS10, WH15b, VSN16]. Square [FH13b, Gar10, HK11, Kha11, Oga15, SHH13, MD17a, RFM18]. Squared [BLN13, Fel10, GSA14, Mah12b, MS10b, Nad10, NVS19, CS19a, CN17b, KA18, LTCQ17, MAG18, Voi15, VPMV16]. Squared-Log [Mah12b]. Squares [AG11, CKD14, FP12, Hua14a, LHHC14, MA11, PYB^{+}15, QZY11, Rah12, TB13, WH15a, ZLZ15, AJ18b, CCZR17, LSL16, Nie17, OU16, OOT19, PCL16, SG18b, Spe17, Ten17, Yam18a, YWH17, ZWPB17, vZ17]. SR [Wu19]. SR-CUSUM [Wu19]. Stability [ARA16, AR13, RR13, BPS16a, LDS18, UZ19]. Stabilizing [NK15, GS19]. Stable [BH12, RP15, SU14, WST10, dSE11, Arv18, CFK17, DL16, GWK19, HB17a, Ike17, KG14b, SW19a, TN17, TMZZ19, Uya17, Van16, YSC17b]. Stage [AYC^{+}12, IUT12, KC12, LMCH12, NN11b, Tak10a, UU11, BS16c, BS19, Bib16, CSY19, Haq17, IF18, Kus18, LHS19, LJC18, LL19c, MSB19a, MLD17, PNPB19, SSS19a, ST16g, ST17b, SSSK18, TK17a, WHZ17, YL17, Ye19]. Stages [CT13]. Staggered [WK13]. Standard [BDM15, Liu11c, MB17a, MSW13, NN15, PJ10, Pao10, YA13, ÇG18, DZH^{+}19, FPPRG16, GCW19, Tar17, Yad16]. Standardization [WZ15]. standardized [CJ17]. Standby [BC12a, LWZ13, WZT15, CZZZ18, Ery16a, FK16, HN17a, LCWG16, MSR17, VM17, Wan16e, YVCA17, YYLM18, ZW17b]. Star [NC14, Sub14, SSH10, JD16, Kam19]. Star-Shaped [SSH10, Kam19]. Star-Type [Sub14]. Start [Sco11, NNZA16]. Start-Up [Sco11]. starting [LL19a, ZG17a]. State [DSC^{+}14, EI11, Ham12, KS14b, PV14, SY17a, SDC^{+}12, SS14a, Vas14a, Vas14b, ZHL15, BAL18b, DPP17, Haj19, HZ17a, KS19, LH14b, LAW^{+}19, SJY17, SS16f, Tat16, Vas16, YCAJ18, ZWZwZ17]. State-Space [Ham12]. Stated [BSB12]. states [CK19, QJKC17, ZCC16]. Stationary [CG11b, CLX14, GL15a, HJ14, HW15, Hug09, Mas15, MS14c, MM11a, NH15b, PBC17, XY09, zXyL13, ZWPZ15, BNZ17, BD17a, BD17b, hCXxShY18, Dvo17, FL18b, FL19, KT18, LGY17, LL17d, PC16a, PLC17, Rah16, Raj18, SN17, Wan19b, Wu17d, ZCC16]. Stationary-ergodicity [ZWPZ15]. Statistic [BDS11, BREA11, CM13b, Deb12, Fly10, GC12, KS10b, Lan14, LLP10, LL11, Li14b, MCY14, RAB15, RP12, SNN12, XHL18b, Xu10, Yam12, ZLD11, AAJ16, BT16, BAW19, Che18, Cuc16, FLY16, KAJH18, Lie16, RK16, TN19, Wan17e, WZX17, WT18]. Statistical [Ani11, ASNT15, AAK17, BSL^{+}19, BG11b, CSY19, CC11, Edi10b, FH12, HA17, HZ19, KAS17, Ko15b, LH17a, Liu17a, NKS15, NNI15, PBY^{+}14, PYB^{+}15, PF19, Qiu17, Roz18, SD14, SA10, SU13, Sti17, WL14, Wei12, WJH15, WSS17, XFWC19, ZSS18, ZS15, ZS10, ZCC16, AABL17, BS17, BK17b, CMS17, FKNR17, FCSH16, HME17, KSF17, KM19a, LK19a, MCGMS17, Nen18, PMKF17, PMFK18, RMM17, SKK18, Sel17b, SF16, SSSM16, Spa19, ZMMX17, ZZ17a]. statistically [SFSH16]. Statistician [Mon12]. Statistics [AA10, AB14, AH13, Ano23, AE23a, AE23b, AE23c, AE23d, BTK10, BS12b, BZAZ15, Bar10, BB14b, (Ba14, BC12c, Bur09, CNC14, CM12, Cra14, EB12, EDAAS12, FZ10b, Fu11a, Fu11b, Fu12a, Fu12b, GD12, GK13, HLX14, HWW15, Hui10, JP13, Jor13, KO10a, KJS19, KT15, LWL13, Lin14a, Mae10, MBS17, MCX10, MA14, Mül12, PS12, PJ14, RA11, Ryc14a, Ryc14b, Sei13, SDJ15, Shi10, SG17d, SA15b, SAC13, Ste11, SM14, TB12, TV11, TCN15, VT15, Voi10, WZ10, WG14, XZNL15, AMP17, AEHQ17, ARB19, AL16, AMP18, Are16, BNZ17, BEY18, BD19, Bar17c, BAR16, CS16a, ÇG18, CW19a, CG16, CZ19c, CCM19, CH19b, CCMA18, DYL17, EE14, Ery16b, GF17, Gor17, GR18a, HD17, HN17b, JPAB17, Kay16, KI19, KR17, KDN17, Len17, LWH19]. statistics [MSD19, Moh19a, NA17, Par17b, Rez17, RKMG19, RJN16, aSaCN18, Sch16, SBA17, TN16, VT17, VZ19, WDLZ16, Xie16, YMF15, ZWZ17, ZG17b]. Statistics-Theory [AE23a, KJS19]. Status [RPY^{+}13, She12b, ZTZ^{+}13, FDL17, NRN^{+}17, SSD19]. steadiness [KDGV16]. Steady [KS14b]. Steady-State [KS14b]. steepest [Kia17]. stein [CW18, DY18, FRC16, JN19, KNC18, KU17, LY16c, MT19, Tee10, TN13, TCN15, ÜY13, WZL10, XY12, XY13, Zha17d, ZTSL15]. Stein-Rule [ÜY13]. stein-type [CW18, XY12, XY13, DY18, LY16c]. Step [BP15a, Bas14, GB12, HN14, LvR12, TGXX12, Tsu14, WXPS12, XF12, Bel17, GH19, Hot16, HSSP16, Lin19, LS17c, LSCW17, SA16, SS16d]. Step-Stress [Bas14, WXPS12, XF12, HN14, TGXX12, Bel17, LS17c, LSCW17]. Stepanov [LYX18]. Stepanov-like [LYX18]. Stepwise [FBS10]. Stereological [Zie13]. Stieltjes [MH16a, SA17c]. Stirling [BM10]. Stochastic [AB11a, AK14b, AJ19, AJN15, BAT11, BC12c, Bia15, BCCH14, BPSV14, CC10a, Cha14, CY14, CL10b, CVL19, DB15, DMST12, Fan15, FLB16, FP17a, zF16, GGLV14, GV11, GW15a, GJM14, HW14, HQW15, JLV13, KT15, LLD17, LL19a, Lin14b, LWR11, Lu13a, MFS12, MS11b, MM11a, NN11b, PW11, PMC10, QWZ14, RSA13, Ric13, RCdC14, SN12, SW14c, SL16, Udo18, Wan17g, Web12, WY13, XWJ10, XGZ15, YC12, YW11a, YC11a, YY14, YTX14, ZZ15, ZB17, ZXD10, dSE11, AD17b, Ali17, AD16, AW16, BTC16, BBHN18, Bia17, CMN^{+}18, Car17, DMM16, DLH^{+}18, FSSW17, FRC16, JXW17, KD17a, Kaç19, KAT18, KDGV16, KT18, LRZ17, LDS18, Li18, LYX18, LL17e, LW18b, MDR17, MSD19, MN18, OD19, Rah18, RG16, SYC19, SW19b, SY17d, SLZ18, UZ19, Wan17f, zWDD19, WZJH19, WA17, XDW19, Xue17]. stochastic [ZLD16, ZG17a]. Stochastic-Flow [CL10b]. Stochastically [ZLD15, LSL16, Ren16]. stock [HHC^{+}17, RBK^{+}19]. stocks [MC19]. Stopped [AAE13]. Stopped-Sum [AAE13]. Stopping [MR17, LJC18]. storage [YGWQ17]. Strata [GFR12, RMP12]. Strategies [JB15, MS10b, CZZZ18, NAML16, Sen17a, TMZ^{+}17]. Strategy [Ken11, NN11b, GQC17, LRZ17, Sen16b, SSB17c, SK18b, WWWX19, XZ17, YSS19]. Stratification [SV10, TTC17, RK19, Vis18]. Stratified [AC15, BA10, BSY12, CKLS11, GS15c, HS13a, MZ18, MS15, NPC17, ÖÇO14, SG10a, SG10b, SS13c, SS14b, TC14, WDZ11, YUSC14b, BB19a, Cle18, KGH19, KJ17, Koy18, LHS19, LTS16, LTV17, MSK17, Nag17, NSSS17, Ozg19b, RK19, SCYR17, SCYR19, SG17a, SG17b, SG17c, Sha18a, SGA19, SH17b, ST16g, SG19, SS16h, TS16, TS17a, TS17c, TS17d]. Stratifying [SAS11]. stratum [FZS19]. Stream [WCL11, TK17b]. Strength [AMGK13, AMGK15, EI11, NRW15, RAK15, SK15a, BBA16, CK19, DMA17, MIG17, QJKC17, RAA17, Sha18c]. Stress [AMGK13, AMGK15, Bas14, EI11, hFmH14, MIG17, NRW15, QJKC17, RAK15, SK15a, WXPS12, XF12, BBA16, Bel17, CBSN16, CK19, DMA17, DW19, HN14, LS17c, LSCW17, RAA17, Sha18c, TGXX12, XK18, ZS15]. Stress-Strength [AMGK13, AMGK15, EI11, NRW15, RAK15, SK15a, MIG17, QJKC17, BBA16, CK19, DMA17, RAA17]. Strict [DMP11]. Strictly [XY09, ZZ18b]. strings [ZXSZ18]. Strong [AR14, BZH19, Bin17, BEZ11, BZ13, CS15a, CHZ17, CL17a, Che17b, FR10, GFA10, HH13, HWW15, Ko13, KB14, LPLW13, LMR15, SY11, SB16a, SY17c, Ste15, Sun13, Sun15, WJ11, dXcY12, XY14, XYL14, YWS14, YYY17, YYS^{+}18, ZL12, CZW16, Dan18, DWX17, GHSZ19, LX17, LLM16, Lou19, Men19, NHV17, Pen17, SZV16, SXZ17, SY16, TY17, WH17b, WP19, Yan16, YYW17, YWH17, Yan17b, Zha17a]. Strongly [KB14, TW13, BAT^{+}18, MPG16, RRRA18, Tan18, YWZ18]. Structural [AR14, Chi09, Chi12a, EY10, WILL13, XGZ11, AHZ17, BR19, CS16c, CN17a, DPZ17, DTT^{+}16, HS18b, KL17a, ZVG19]. Structure [BS12a, BG18, GHY11, GK19b, MBN14, Par13b, RL12, RVP15, SC10, SB12, SMM10, Web12, WW11, Yam12, ZWLI13, ZL14b, GPM18, HA18, MBG19, PY16a, PR17, WHHW17, WHYW18, Xu17b, YVCA17, YSC17a, YX18]. Structured [LR14, RL11b, SAvR12, WS12b, SÁHLvR17]. Structures [ASB13, ADO13, BBB14, CL15a, Ind11, MP10b, Oga11, RAP13, BKR19, FPCP16, GRSN18, KYW16, LL17a, PPMD19]. STSD [LZPJ11]. Student [LPS11, Cha18, CW19a, Maz19, CY14, Cos18, SF12b, TP14, ZX16]. Student- [CY14]. studentized [Çel19, Shi10]. Studies [DZ12, sK13b, KCS12, Liu11a, SD12, TWH10, WDZ11, YW15, CMN^{+}18, FTF17, MZ17, PAvS^{+}19]. Study [AZN15, BC12a, BNE15, BBP14, DCL12, HT13, HWW15, JCY10, JWZ14, KK11b, KCL^{+}16, MH15, RM15, SA14, WH13, XLZ^{+}12, YC14, YYX14, AG16, CSM16, GQC18, HS17, HST19, KZ19, KN16b, KN18b, OTA17, SSSM16, She17, SSS17a, SPY18, SA17b, SR18, TSOT19, VPMV16, Zha17c]. stylized [Ram17]. Sub [Ham13, SKK10, CHZ17, FAT17, FWW19, LTC17, WW19, YW18]. sub-fractional [LTC17]. Sub-Independence [Ham13]. sub-linear [CHZ17, FWW19, WW19, YW18]. Sub-Sampling [SKK10]. sub-space [FAT17]. subclasses [AN17a, PW16]. Subcritical [Rah12]. subdistribution [LXL17]. Subexponential [RP15, BSW15, Lou19, ZK16, ZC17a]. Subexponentiality [Kon11]. Subfamily [Afe13, GKP16]. Subfractional [SY14a, SY14b, Dai16]. Subgroup [CC15a]. subgroups [PS17a]. Subject [LK14, LJ13]. Subjected [SCC10, VM17]. Subjective [SUS19, VBAR15]. Subjects [ZWZZ15, Liu17c]. Sublinear [ZC14, CF16, CL17a, GHSZ19, HC16a]. Submodularity [YS15a]. subordinated [FKR16]. Subordinator [LPS11, GWK19]. subpopulation [Sou17]. subpopulations [Sou17, WW17b]. subsampling [Kyr17, RSC16]. Subset [CHP14, DGMC17, MJ15, BCP17, KC16, NP17b]. Subspace [KQ11, YGWQ17]. Subspaces [LGV13]. Subtangent [RR09, SS12e, SSR19]. successes [CW19c]. Successive [AS13, Ima17, NN13, SS12a, SMPH15, SMMS15, AHH16, BS16a, CSB19, GR17, PT19, SP16c, SKM16a, SS16c, Sin16, SP16b, SKM16b, SHM16, SSS17b, SSB17c, SMK17, SKS17, SP17c, SKM17, SS18, SK18b, SSS18, SSKP18, SS16i]. Sufficiency [Por11]. Sufficient [FAS11, SS16e, Tan10b, Zhu10, AP19, BBHN18, BMR17, SA17b]. Suicide [XLZ^{+}12]. Sum [AMR14, AAE13, Hua14a, QKP12, Rat13, SK12d, Wit14, CJ17, FL17, HS18a, Kha19a, Kha19b, LSZ16, SA17c, WZ16, ZC17a]. sum-to-zero [WZ16]. Summands [AK14c, ZC17a]. Summary [Wei14, RKMG19, WER^{+}19]. Sums [AE13, BCH15, CL15c, sDsWcCjK14, GJ15, Guo14, Ko14e, KL11, Li13b, LSX14, SSA10, SWLW14, Sun15, Tee14, WHVH13, WW13c, WLY14b, XM14, ZY13, ZLF14, CKS17, CZW16, DW17, FWW19, Fro17a, FS17, HF19, HSLQ17, HZZ19, Ko16c, Ko17b, LPH16, LSW16, LSL18, MX19, NMAA19, QC16, SAB18, SWZ16, ST16c, SXW17, SHLQ17, Spe17, Tan18, WHV16, WWW17, WHHW17, Wan17f, WHYW18, zWDD19, WV16, WG16b, Wu17c, XDWC18, Yan18, YJS19, YQC17, YC17, YW18, Zha16b, ZK16, ZMMX17, Zha17a]. Sup [XW16]. Sup-Hellinger [XW16]. Super [UC14, CSB19, WL19]. super-population [CSB19]. superadditive [MWW18]. Supercritical [Yar11]. superefficient [AESO19]. Superimposed [MS10e]. Superiorities [CW13]. Superiority [RL09, ŠS12c, Wu16c]. Superpopulation [AB09, BA10, HX11, SK12b, AB17]. Supersaturated [MGP14]. Superstructure [CC15b]. Supervised [KK11a, LLV10, PR17]. supplementary [AGS18]. Supplemented [KMM12]. Supplier [SW14c]. Suppliers [BS14]. Supply [JC12, SF16]. Support [Koi10, CC16a, SSH17, SA17b]. Supported [KOP14, Car17, PTV18]. Supremacy [Far12]. Sure [BR11, Li13b, XM14, BR17, LPH16, LO17, MGN16, MX19, zWDD19, Wu17c, Wu17d, XML16, Yan18, Zan17, Zha16b]. Surface [AAG12a, FHS15, PP10, ZS12, BQC17, CS17]. Surfaces [AAG12b, MBN14, UC14]. surgical [KOFA18]. surplus [Zha17c]. surplus-lag [Zha17c]. Surrogate [YL10, Roy16, YWLZ17]. Surrogates [Ko12]. Surveillance [FAS11, SS16e]. Survey [ASN12, AZN15, BBP14, DSSL15, Eid12, LB11, She15c, SS15c, AH17a, SSS16, SS16d, LS16b]. Survey-Weighted [LB11]. Surveys [Arn11, YUSC13, YUSC14a, Zad15, PHH18, Sen16b, Sen17a, SP17b, SPS17]. Survival [AB15, Abd15, CZ15a, DPB13, DM10, GMB^{+}13, HD13a, HS15a, HM12, JFS14, Ko10b, sK11, Ko12, sK13b, Ko14a, Ko14b, Ko14c, Ko15b, Ko15a, LSC13, MSSU14, Man11a, MC11, OTA17, PCS18, Ree12, SF16, She10a, SCdSP10, WTB11, ZSS15, AdBPdBP18, BB19a, BT16, BS19, CLZ18, GZ17, GW18, HB16, IKH19, Ko16a, KN16b, LL17c, Liu17c, LL17f, PA18, RASR17, SV17, Sha18b, She17, Tat16, VM17, WXS18, vSK16]. survivors [CCOB16]. Survivorship [SG10c]. Suspected [Kum15]. Sustainability [HH13]. sustained [VM17]. SVMs [Gue10, Gue13]. SVSSI [KSF17]. Swaps [LD14]. Swarm [CYE15, IES^{+}17, KDG^{+}18]. Swedish [LS11, MKS15]. swine [KD16a]. Swiss [JPAB17]. Switching [Gov11, Mah12a, NY11, Pie11, WW10, YTX14, ARA16, AH18b, BU17, BJU18, BG18, Cav17, Che16d, DYW18, FSSW17, FPR17, GDW18, HZ17b, MB19, TMZZ19, XDW19, ZDDH17]. Symbolic [DO13]. Symmetric [ADP10, AB11b, BL15, Cah15, CEK11, DC14, FZ10c, GCSB10, KG14b, KQ11, Kur10, LJ12, MS10a, PT15, Saj11, Tan13, Wil09, XY13, Yam12, AGS18, BMFS18, DE17b, HEB16, LRC17, Mod17, Roo17, aSaCN18, SN18, Uya17, YX18]. symmetric-asymmetric [BMFS18]. Symmetrical [VRC13, ZZC13, DAC17]. Symmetry [HT13, Li14b, SH11, BBR17, DS17b]. Synchronizations [Man11b]. Synergistic [PMA11]. Synergy [MS14a]. Synthetic [BP15a, KS10b, LK17d, MC14, ZSY13, LK17c, LK17b, LK19a, SG17d, WWZC18]. System [AALS12, AEH11, BC12a, CSC13, Che14, DZ12, DT12, EI11, hFmH14, Gov11, JC12, Kam14, KJL11, Lin14b, MTM13, QZZN10, Raq10, SCC10, TE18a, Vas14a, WYL12, XZD10, ZW13, BU17, BJU18, BL19, CC19, CZZZ18, Che19a, DP18, DLC16, Ery16a, GXB18, Han17, HWZ19, KAR16, KS18a, Kyr16, LCWG16, MB19, MAA17, MSR17, PBA18, SK16e, TSOT19, VM17, Vas16, WWZC18, WCZP16, WAJ17, YVCA17, YGWQ17, YL18a, YYLM18, You17, ZXM16, ZW17b, ZW17c, ZB17, ZH16]. Systematic [CKD14, KSG13, Sub14, CMS17, KS16b, NNZA16, NNZA18]. Systematically [TR16a]. Systemic [CH19b]. Systems [AA14, BHB15, BS14, CGSM10, DCG13, DSC^{+}14, Ery12, JCY10, KS14a, LYH11, LWZ13, MGF14, MS12b, NY11, NRB14, PV14, Rab13, RSA13, SDC^{+}12, Vas13, Vas14b, WZT15, YLL19, Yil12, Zha11, ZM13, ZZL12, ZL15, AHEA18, BBHN18, BNB17, BNB19, CNL19, CVL19, FLB16, FP17a, FK16, FN18, FY17b, HNB19, HN17a, JSWL17, Kom17, KDGV16, KS19, LL19a, NAN17b, NI16, PSF16, PBA17, PKM17, QC19, QCK19, SHB17, SNN16, Sym19, Tav16, Too17, UZ19, Wan16e, Wan18, YL18d].

t [DBA18, Sab19]. t-distribution [DBA18]. Tabapua [GL15b]. Table [GN10, SS10b, SS14a, MO18, SMN18, SS16f]. Tables [BF10, BSY12, CC14, DBC14, EPGTH15, FZ10a, SS11b, YSR11, Bar17b, LVH19, THV16]. Tabular [Jor13]. Tabulations [NCA17]. tackling [WF17]. Taguchi [BDS11, HD13b, PMKF17, PMFK18, TN15]. TAIEX [HL18b]. Tail [BZ13, DFS10, Fer14, FL17, FZ10c, GJ15, GBD10, GD12, GMP11, GWP19, HC16b, LPN14, MJH19, NY16, RGL12, SF12b, TP14, WLY14b, WW15c, Yan15, dCOSW13, JSZL17, KWM17, LG16b, PY16b, Sta17, WHYW18, ZK16, ZC17a]. Tailed [ADO13, BS12a, FZ10c, KL11, LJ12, Wan13a, WLY14b, WW15c, Ali17, DPZ17, FL19, GGDVP16, GCW19, GW15b, KQ17, LCZ16, LQ17b, QL18, Roo17, WHHW17]. Tailoring [Zoi09, FPZ16]. Tails [HLX14, LSX14, Yar13, DW17, GS17, HSLQ17, LSZ16, QK17, ST16c, SHLQ17, YC17]. Taking [Gue10, STD18]. Tandem [GMRN13, GMR16]. tapering [PCL16]. target [GDDC19]. targeted [Ko17a]. Tau [HL17]. Taxonomic [SZL14]. taxonomy [KM16b]. Taylor [WN17a]. technique [AAE17, AO17b, JSS16, MDG19, ST17a, SL18, SG19]. Techniques [Arn11, ASN12, AS15, TWTG14, AO18b, Arn19, HS18a, LPdDTH17, MSD19, SSKP18, Wan17e]. Technology [HM12]. temperature [Ayk19]. temperature-related [Ayk19]. Tempered [KUWG19, MRS12, SXZ17]. tempering [GSLZ16]. Temporal [DC10a, PMC10, ZA14, AA17c]. temporally [LW17b]. Ten [CDE10]. Term [EE14, GT15, LSC13, CCCT16, CE17b, CN17b, CCOB16, DAC17, KYW16, KKKÖ17, MR17, SBLC17, XO19]. term-by-term [CN17b]. Terminal [ZZS13, ZWZZ15, DDS16, Fan18, HDS16]. Termination [KNK13]. Terms [CHP14, RR13, Xu12, EQ17, OQ17, Tav16]. Terry [Yan16]. Test [Aca15, AR14, BF10, BT13, BC13, BAT15, BK09, BEKL13, Cag12, CY14, CM13b, DLL13, DPS11, DPS12, DI10, EE15b, hFmH14, GAK14, GSA14, HH15, Him10, JZQ15, JO10, KK14, KJD^{+}15, sK11, KSBM14a, LJ13, LYX11, LS11, LMM15, Lu10, Lu11, MSS14b, MO10, Mar12, Mar11, MDC16, NO12, PS12, RD10, RJ15, SH11, Say14, Sch12, Sco09, ST13, SS15a, SA15b, SS17c, Sur12, Tsa10, TJ13, VV12, Voi10, WDZ11, WXPS12, WXPL13, WQ14, WK13, WW15c, Wu11, YS12, YXC11, YMF15, YSR11, YK11, ZAV14, Zha15b, AMP17, AASRZE17, AL16, AW16, Arv18, AdBPdBP18, BP18a, BG19a, Bar17c, BB19a, BMRSE17, BDM15, Çel19, Cha18, CW19a, CK17b, CNG17, CQ17, CW19b, CZ19c, CBM14, Coi17, DW18, DO19, DW19, Dvo17, ESAKK18, FZS19, FPM19, FLY16, GG16a, GR18b]. test [HD17, Hay17, HOK17, HA17, Hot16, HL17, Hu16, HN17b, JRF17, JZZZ18, Kal19, KKS19, KAT18, KS19, LPC17, LB16, LVH19, LMM18, LWZ17, Mao17, MS17b, Mer16, MF16, MGG^{+}17, Nou17, Par17b, PFD19, RPH17, SDS17, Sen18a, SI17, SE17, Sou17, Tat16, TS17f, VV17, WH17a, Wu16d, XX17, XO19, YX18, YXZ17, YH19, YW17b, ZBTL17, ZBHW18, ZWP17, ZS17b, ZJC10]. Testing [AV18, Ahm17, Alm11, AGVGP^{+}11, BPS16a, BR12, BKR19, BD11b, Bon14, Che11, CGS16, DJ15, DCZL17, DP13b, ESAKK18, ER14, FZ18, Far12, GG10a, GFR12, Gia14, GVGP11, GMH16, HN14, HXL13, HH13, HK11, KHM12, KA17, Ken11, KG14c, KE10, KL12, Kum15, LLP10, LL11, Li14b, Li15, LTZ19, LL17c, LWHW11, LTGK17a, LTGK17b, Mån12, MMN16, MMP18, MHA19, Mao18, MN13b, Mod12, MM15, MH15, Nar13, NA13, NC18c, PMA14, PK17, Pro18, Rab13, RON12, RAB15, RL11b, RM14, SAKK15, SY17a, SLZ11, SKG12, SJ10, SE13, SR11, Sun17a, TD14, Tor12, XZD10, XF12, YSS18, Yam12, YHA10b, Yen12, CBSN16, CK17c, Che17a, CZ19c, DKY16a, DKY16b, FM16, FWZ^{+}17b, Isl17, IKL16, KAR16, Kha16, KLKK17, KR16, KT18, LLHZ16, LLY16, LMZ17, LWHB16, QL18, Roz18, SM18]. testing [SS18, Tch19, WX17, WH18, WL17d, Xio16, Zha17c, ZXW16]. Tests [AV15a, AJA14, AASR12, AMH14, BLN13, BL10, BF12, BC12c, dBCR15, Cey11, CY14, CB10, CB12, CL15c, DTO11, DZ15, DMRO14, DS17b, ET13, HME12, HS18b, JMEM14, KMC10, Kak10, Kak12, KM10, KNK13, KSBM14b, LLP10, LL11, LF10, Li12a, MGAf14, PSS^{+}15, PS11, Pes15, Sal15, SGR^{+}12, Sco11, SS11b, SB12, SL14, Sto11, Tak14, TGXX12, Tun15, VB13, VP10, Voi15, Wan10b, Wyl13, XFC11, Yan17a, YHA10a, ZPZ10, ZLD15, ZXD10, ZS15, AM17a, AABL17, BBC18, BBR17, BRY17, Che16c, DDP18, DY16, DL16, Far17a, Far17b, Gau17, GL19, GF17, JZ17a, JBA18, JSZL17, JK19, KL17a, KP18a, Kou16, KG18a, LM16, LSCW17, LR19, Mar16, MAS16, MR16b, Mur16b, NKG16, OL19a, ÖGGB16, Par17a, PLC17, PP17, Ric17, Sch17, She16, Sil18]. tests [SG16e, UU17, VPMV16, WK19, WBGS19, XK18, Zha16a]. Text [Iez12]. TGARCH [MC19]. th [DK17b, Jas16, Jas19, Pao10, RAB15, SYW16, SY16, SY17c]. th-order [SYW16, SY16, SY17c]. Their [AAG12a, AAR11, CHP14, Mae10, Men10a, OCIS12, SDJ15, YW11b, ZZL12, AJT16, Bos18, CSS16, CC16b, Dom11, JPAB17, JA10, JAA11, KK17a, Kur10, NN11a, SY17b, Sug13, WWW17, WHHW17, Yar14]. Theorem [BZ13, CYY14, GLS12, HWM12, JFA12, Kát11a, Li13b, MZ14, NR12, Sun15, WH15a, ZC14, hCXxShY18, DeR19, HZW17, Hu17, HYS18, Ko16c, LPH16, MGN16, MX19, SY16, SJY17, SN16, SNS19, Sze17, Wu17c, Wu17d, Xie16, XML16, Yam17a, Zan17, Zha16b]. Theorem-Based [HWM12]. Theorems [MC11, Sun13, CZCM19, Ike17, SYW16, Tav16, Wan17f, ZZ17a]. Theoretical [DG12a, Dut17, FWH^{+}15, GAS17, DAS16b, NM18]. Theory [Aca15, Ano23, AL15, AE23a, AE23b, AE23c, AE23d, Bic12, CLC^{+}16, DMM16, GJ15, GdSC15b, Kát11a, KJS19, MBS17, MAYH17, Mül12, Nag17, OSX^{+}16, Pes15, SG17d, SS14a, TN15, Vin13, ZY10, Arv18, HF19, IS17, KG17, LK19b, MSD19, PM16a, Sel17a, WCW17, XP18, ZL16g]. there [MT16]. Thinning [ZWZ12]. Third [AAG12b, Cav17, IUT12, KP12b, RAZ19, ZQZY18]. Third-Order [AAG12b, IUT12, KP12b, ZQZY18]. Thompson [Bak17, GÖ12, Sen17a]. Three [BNS11, Bi15, CY14, CM17, DGV11, FH12, GGS13, Jia12, KKKÖ17, LPDD19, MTM13, NBK14, NNA12, OQC14, Pou17, QZY12, RL11b, SAKK15, Tak10a, ZXSZ18, AJ18a, CK17a, CCZR17, GH19, GQC17, IF18, JC18, KVR17, LHS19, LOHW19, NB16, Pat16, RAM16, RRRA18, Sha18b, Yad16, YVCA17, ZZ19a]. Three-Decision [GGS13]. Three-Dimensional [BNS11, MTM13, AJ18a]. Three-factor [SAKK15]. Three-group [CM17]. Three-Level [RL11b, OQC14, Pou17, GQC17, LOHW19]. Three-Parameter [DGV11, NBK14, QZY12, CCZR17, JC18, NB16, RAM16, Sha18b]. Three-Part [FH12]. three-population [Pat16]. Three-Stage [Tak10a, IF18, LHS19]. three-step [GH19]. Three-term [KKKÖ17]. three-unit [YVCA17]. Three-way [SAKK15, ZZ19a]. Threshold [BGT10, CCL14, CSF12, DMRO14, KY10b, MSP12, NV15, Rab13, XLPL12, XLL15a, GK19b, LX19, LW17c, Mer16, MC19, QC19, VN17, WWWX19]. Thresholded [LX10b]. thresholding [CN17b, GP19]. thresholds [GEM^{+}18]. tied [ZZ19a]. ties [Spe17]. Tikhonov [AD17a, TPTT14]. Time [ADP10, AW10, AP12, Ano13b, AM15, AT09, AMH14, BLN13, BS12a, Bia15, Bin14, CC10b, CZ15c, CGSM10, DDGM13, DJ15, Den13, DT12, DN14, Dok12, Don11, EMI14, El 15, FMS15, GS11, GL15a, GL15b, GS15b, Gue14, HL11, HW18b, HQW15, Igl13, Kha11, KY10a, Kim15, sK11, Ko14a, Ko14c, Ko15b, Ko15a, KN10, LM13, Lim11, LY11b, LSC12, ML12, MO10, MV16a, MS10d, MH15, NY11, PS14, PHW11, Pop12, QNZ14, RL12, SM09, SG16d, She12a, SP11a, SS10b, SS14a, TW13, TB15, TT11, VYZ14, Vas14a, WH15b, WW15b, Wan19b, WILL13, XLL13, XZD10, YWO^{+}15, ZA14, ZK15a, ZLZ15, ZZSP14, BSW15, BR19, BB17a, BG19c, BBS16, BD17a, BD17b, BG16a, BL19, CSY19, CZ19b, CXC19, CT16, CH17, DE17a, DLZ17, Ery16a, Ery18, FPM19, FP17b, FL17, GWK19, GS17, GLY17]. time [GICRR17, HL18a, HJ14, HH19, HZ17a, HNB17, Ili16, JT19, JZZZ18, JM16, JKS18, Ko16a, KP18b, LW17b, Li17b, LL17c, LG16b, Liu17c, Liu17a, LL18b, LWG18a, LWG18b, LCZA17a, MHA19, Mal17, Man18, MST19, MSL17, NK19, NRI17, NIN17, OAOQ16, PAvS^{+}19, QC19, QCK19, Raj18, RLSL19, SDS17, SK16b, SF16, SV17, SJ19, SS16e, SW19a, TN17, Tav16, Tri17, UZ19, VM17, Vas16, WXS18, WWWX19, XZA18, aYwSlDmS17, Yan14, YTX14, Yu19, ZBA17, ZT18, ZH16, cZzXgL18]. time-averaged [LCZA17a]. time-changed [GWK19, SJ19]. Time-consistent [HW18b]. Time-Course [Kim15]. time-delay [UZ19]. Time-Dependent [Den13, GS11, She12a, YWO^{+}15, FL17, LG16b, SS16e]. time-depending [HL18a]. time-dynamic [LL18b]. Time-Homogeneous [Ano13b, MS10d]. time-modulated [Man18]. time-scale [Mal17]. Time-Series [CZ15c, JKS18, NRI17, XZA18]. Time-To-Event [Ko15a, Ko14a, EMI14, HNB17]. Time-Varying [Don11, MV16a, BG19c, HH19, ZH16, cZzXgL18]. Times [Bas14, Den13, LCLJ14, SS12a, WTB11, AHEA18, CC19, DDS16, Gra17, HHS18, KV17, LSY17, LCWG16, MAA17, NC18a, She18b, SSD19, YFV19]. Tissue [WS12b, AIIK17]. Tobit [KMSS14, MFBG13, MH11]. Toeplitz [LH17b]. Tolerance [Chv13, Ko14d, KLM10, KXX11, MS10b, KQ17, Ko16b, QK17]. tolerances [RWY16]. Too [Koz12]. Tool [AB13, Cox19]. Top [Ber12]. Top-event [Ber12]. Topics [KR14]. Topp [Bay16, Gen17, KI19, RSAN17]. Total [CG12, HS12b, FPM19, LJW18, RSC16, SMP18]. Touchard [MPD^{+}19]. Trace [AMZ15, XM17b]. tractable [Cos18]. Tracy [ST16g, Wan13d]. Trade [PVB13, SH17b]. Trade-Off [PVB13, SH17b]. traded [WQW16]. traditional [KS16b]. Traffic [CNC14, CM13a]. Training [LR10a]. Trait [Zho10]. Traits [Fed13]. trajectories [JJ19, SW19a]. Trajectory [NRS^{+}14]. Tranches [CR14]. Transect [ES11a, Eid14, EAE18]. Transform [HME12, Pri15, Ber18, LLLL18]. Transformation [FMS15, GE11, LSZ13, Mae10, MGAf14, Peg12, QNZ14, She15a, She15c, ZTZ^{+}13, ZZ10, HHS18, Jon16, Sel17a, She17, YSD^{+}16]. Transformations [FFA09, CH17, MP16, Zen17a]. Transformed [NB15, NA13, XLL15a, YUSC13, FZZ16, HG19, Lu13b, Mna17]. Transforms [BT13, FM10b, MH16a]. Transient [Bur19, KS18a, TSOT19, Vas14a, Vas16, DP18]. Transition [Rav19, SDC^{+}12, XHY16]. transitional [Che16d]. Transitions [SSM16, AG18]. Transmission [HM12]. Transmuted [GL15b, KY17a, KKH17, BKGD17, GFL19, MAYH17, NAYC17]. transmuted- [MAYH17, NAYC17]. Transport [AB11a]. Trapezoidal [LKK15, GSM16]. Treatment [BD11a, Bis12, CC10c, JFS14, Ko15a, SW14a, TWH10, YW15, BS19, CK17a, GPM18, He18, LL17c, LWZ18, MGP17, OH19, PPMD19, Ros19, WLC17]. Treatment-Control [Bis12, He18, Ros19]. Treatments [GGS13, LLZ17]. Tree [Ber12, GMB^{+}13, SRT14, CCCT16, Dan18, HYS17, Pen17, SY16, SY17b, SY17c, TY17]. tree-based [CCCT16]. tree-indexed [Pen17, SY17b]. Trees [Fle13, SA14, VRC12, HYS18]. Trend [GICRR17, Hil15, KCS12, SS14c, WW13b, WILL13, Wyl13, JZZZ18, NNZA16, NNZA18, NIN17, OL19a]. Trend-Free [WW13b]. Trends [Con10, Zha15b, ZK19]. Trial [sK12a, sK13b, Ko14a, Ko15b, Ko15a, RPY^{+}13, YW15, sK16c, Ko17a, Liu17b, Liu17c, LL17f]. Trials [Ani11, CLC^{+}14, Chi10, LMCH12, WTB11, AMP18, GZ17, LJC18, MLD17, MR17, PHH18, Wu16a, Yad16]. Triangle [CLGFB15]. Triangular [FFA09, ZK15b, HL16, NM17]. Trigonometric [Pop16]. trimmed [Bie16, Gri17]. Trimodal [LHZR14]. Trindade [AE23c]. tripartite [AAE17]. true [CW19c, JKL19, LN17]. Truncated [AT19, AB11b, Coo12, GN15, KRB12, Kit15, Mah12b, PJ10, PDdHNMR17, Par15a, PBS12, She10b, She12a, She13, She14b, WLF11, WH11, AGAM17, AHKO16, AO17a, BN17, FLB16, God16, He17, JBA18, Kal19, Kay18, Kun17a, Kun17b, LC19, NC18b, She16, sSW18, SK19b, VEKH15, WTZS15, Wan16c, Wan17b, Wu16d, YWL17, YWLW18, ZJZ17]. truncating [VPCG16]. Truncation [FJ11, Lia10, Lia11, SY11, She10a, AO17a, HL17, LLM16]. Truncations [GC12]. Tsallis [Kum17]. Tsui [MF16]. Tucker3 [Gal15]. tumor [She17]. Tuning [WSC13, ZP17]. tuple [LB16]. Turnbull [She14b]. Turtles [Far12]. Tutorial [ZSS15]. Tweedie [CFK17]. Twice [She15b, GMS18]. Twice-Censored [She15b]. twin [DP18, HB17a, HB17b]. twisting [MDG19]. Two [AW10, AASR12, AAR11, ASAN12, ASN12, AYC^{+}12, BS12c, BDKM12, BHLK15, BC12a, BF10, Bib16, BC15, BEZ11, BSY12, CNC14, CT11, CCZ15, CC10c, CJ15, CB10, CB12, CFA17, Chv13, CM18, CGSM10, CT14, DBC14, DM11, DW13, EDAAS12, EWRZ14, EE14, Ery12, FLL17, FLK12, GMRN13, GFG13, GdSC15b, GPÖ14, GW15b, GS15c, HKR15, Haq17, HMP12, Hay10, Him10, HLL15, Hua10, IUT12, JB15, KG14a, KC12, Kat11b, KPS10, KNK13, Ko14d, KR16, KG15, KSBM14b, KSBM14a, KLM10, KZ15, LMCH12, LY10, LL12, LvR12, LJC18, MM17, MS15, NN11b, NY11, Nos14, OFS13, OQC14, Pao10, Par13b, PGF11, PS11, PP17, Pop12, RB10, SCMR15, SK13a, SCSMN11, SBK13, SS10a, SK15a, SKK10, SK10a, SSP^{+}12, SK12b, SK14, SMPH15, SMMS15, SS16d, SHLK10, SR11, Tak14]. Two [TOT15, TGXX12, TWTG14, Tin10, Tsu14, VB13, WBH10, WL13, WLL13, WYZ15, WH10, WY13, XLPL12, Xu10, YLL19, YLX10, YC10, YLS10, YC11a, YSR11, Yil12, YTX14, YW17b, ZZP10, ZPZ10, ZLL11, ZS12, Zha14, ZFZ17, ZK15b, AE17a, AE17b, AE17c, AA17a, AHEA18, ASS18, AS18, AHKO16, AR17b, AN17a, AO17b, AB19, AEA17, AAK17, BS16a, BS16b, BNB18, BS16c, BS19, BMLY19, Bob19, CMN^{+}18, CSY19, Car17, ÇG18, CK17a, CW18, Cha18, CW19a, CK19, CZ19a, CM16c, CVL19, DS17b, DW18, EIM18, Ery18, FA19, FL17, GWL16, GSM16, Gen19, GMMA19, GMR16, GN16, GG16b, Hay16, Hot16, HSLQ17, HYS17, Hua17a, HWY17, HSSP16, HN17b, JSSO18, JT19, JKJ16, JYX16, JY16, KY17a, Kit15, KX18, KZ19, KA18, Kus18, LKK17, LY16b, LS17a, Li18, LPDD18, LX19]. two [LGG16, Liu17b, LOHW19, LCZA17b, LSW16, LLZ17, MMN16, MMP18, MNM19, MHA19, MSB19a, Mao18, Mar16, MS17b, MLD17, MK19, MRA16, MR17, MSK17, NLRD17, NKG16, mNK17, NP17b, NN18, Ozg19b, PSF16, PY16a, PPMD19, PK17, PM16b, PB18, QJKC17, RSC16, RAA17, RK18, SAD16, SSS19a, SG17a, SG17b, SG17c, Sha18a, SGA19, SA16, SR17, ST16c, SS15a, SP16c, SKM16a, SKT16, SS16c, ST16g, SKM16b, SP16d, SSB17c, SMK17, SKS17, SP17c, ST17b, SKM17, SK18b, SSSK18, SHLQ17, SSS17c, SG16e, TSOT19, TK17a, Uya17, Wan17a, Wan17c, Wan17d, WER^{+}19, WG16a, WAJ17, WZZ17, XHH17, XO19, Yam18b, aYwSlDmS17, YH16, YL17, YWZ18, Ye19, YLS17, ZG19a, ZLYZ17, ZL16e, ZS17a, ZG17b, ZZ18e, ZZ19e, vZvdM19]. Two- [OQC14, WYZ15, LOHW19, ZLYZ17]. Two-by-Two [KG14a]. Two-Component [JB15]. two-decks [JSSO18]. Two-Dimensional [SCSMN11, HKR15, Ery18, LX19, LGG16, ST16c, ZG17b]. Two-Factor [Hua10, CMN^{+}18, YWZ18, ZZ19e]. Two-Fold [Par13b, PY16a]. two-group [LCZA17b, Yam18b]. two-independent-sample [LPDD18]. Two-Item [BC12a]. Two-Level [CT11, HLL15, Hua10, Tin10, CJ15, WLL13, CZ19a, Wan17c, Wan17d, ZL16e, ZS17a, ZZ18e]. Two-Machines [NN11b]. Two-Occasion [SK10a, SSP^{+}12, SK12b, SK14, SMPH15, SMMS15, BS16a, PM16b, SKM16a, SKT16, SS16c, SKM16b, SP16d, SSB17c, SMK17, SP17c, SKM17]. Two-Parameter [CT14, GFG13, KPS10, Pao10, SK15a, TOT15, YC10, YC11a, BHLK15, TGXX12, AE17c, AA17a, AR17b, AEA17, Bob19, CW18, Cha18, CW19a, GMMA19, HWY17, KY17a, KX18, LY16b, WG16a, YH16, vZvdM19]. two-period [LS17a]. Two-Phase [DM11, Nos14, SKK10, SHLK10, GMRN13, AO17b, BS16b, GMR16, JKJ16, RSC16, SG17b, SG17c, SGA19, SKS17, SK18b, aYwSlDmS17]. Two-Piece [ASAN12, MM17, DS17b]. two-plan [WAJ17]. Two-Regime [XLPL12]. two-rooted [HYS17]. Two-Sample [BS12c, BDKM12, BEZ11, BSY12, CB10, EDAAS12, KNK13, Xu10, YLS10, BC15, FLL17, PP17, Tak14, TWTG14, DW18, KX18, Mar16, MS17b, SR17, Wan17a, WZZ17]. two-sampling [MRA16]. Two-Sequence [EWRZ14]. Two-Sided [Chv13, DW13, Ko14d, KG15, CCZ15, CFA17, GW15b, YW17b, Zha14, ZK15b, AS18, AHKO16, ÇG18, CM16c, Gen19, GG16b]. Two-Stage [AYC^{+}12, IUT12, KC12, LMCH12, Bib16, Haq17, LJC18, BS16c, CSY19, Kus18, MSB19a, MLD17, SSS19a, ST16g, ST17b, SSSK18, TK17a, YL17, Ye19]. Two-Step [Tsu14, SS16d, Hot16, HSSP16, SA16]. Two-Term [EE14]. Two-time [YTX14]. Two-Time-Scale [NY11]. Two-Treatment [CC10c]. two-unit [CVL19]. two-urn [PB18]. Two-Way [BF10, DBC14, KSBM14b, YSR11, LLZ17, Mao18]. Type [ARD14, AK13, Aoy10, BT13, BB14b, EE15b, EdSMR14, Gar10, GSA14, HK15, HCS10, Iqb13, LPLW13, SAB15, SK18a, SMM10, SST14, SAC13, Sto11, SYP13, Sub14, TC14, TTC17, Voi15, WTZ15, XY12, XY13, Yan19, AHAH17, Ali17, ANA17, AEA17, BN18a, BQC17, BP18c, ÇE17a, CO16b, CW18, CE17b, Che13, DT17, DWX17, DY18, Dvo17, Far17a, Fro17a, GGO17, Gri17, HHC16, INOC17, IES^{+}17, JZQ15, JZ17b, Kaç19, KS16a, Kim15, KDG^{+}18, KHC14, KS19, KZ19, Li18, LL14, LLM16, LCWG16, LY16c, LL19b, LTS16, MGF14, MDA16, MK19, MKK17, MZ14, PHY19, PM16b, Qom17, Rao17b, SG18a, SG16d, Sha19, SL16, SM16, SP17a, STRD17, SSS17a, SPY18, SS16i, Tan18, The19, Udo18, VPMV16, WX15, WSW16, WLS18, WHYW18, WF17, Wu18b, WCF18, YUSC14b, YYS19, YYWH19, Yil18]. type [ZMCM16, ZL16b, ZW17a, ZS17b, ZG19b, AA17a, AN17d, AAK17, BS12c, BS12b, BHLK15, Bas14, BREA11, CBSN16, CB10, CB12, CT14, Cra14, EDAAS12, FH12, GP16b, GC19, HN14, JBH12, JZTN17, Kha13, LJ13, MA14, Muk10, NB16, NC18c, Pao10, PGF11, RD10, RB10, RAB15, Rez17, RK16, Sha17, SBK13, TGXX12, TTZ14a, WL10, Wan12, Wan16c, WL17a, WK13, YLS10, ZS15, CK19]. Type-I [AN17d, AAK17, CBSN16, CT14, GP16b, GC19, HN14, NC18c, RD10, Sha17, TGXX12, WL17a, ZS15]. Type-II [SK18a, AHAH17, MK19, Qom17, AA17a, BS12c, BS12b, BHLK15, CB10, CB12, Cra14, FH12, MA14, NB16, PGF11, RB10, RAB15, Rez17, SBK13, TTZ14a, Wan12, Wan16c]. type-III [STRD17]. Types [CGSM10, ZLL11, BAF19, CS19a, El 17, GWL16, TPN18, TSOT19, XO19].

U.S. [Con12]. UEND [LSZ16]. Uhlenbeck [Tra17, WTZ15, YSC17b, ZW19a, ZZ19d]. ultra [WX19]. ultra-high [WX19]. ultrahigh [SLSC18]. umbrella [GG16a]. UMVU [Ghr19, SK16e]. Unavailable [BSSJ11]. Unbalanced [Bak11, YA13, TK19, Yad16]. Unbiased [BDKM12, ES11a, GD12, LYX11, LWR11, MB10, Neg09, NP15, QLZ14, RS14, SU14, Tsu14, XY11a, AR17b, Cha18, CW19a, HBMAO16a, HBM16, HWY17, Kaç19, Li18, LY16c, RG19, Sen18b, Wu16b, WAA18]. unbounded [CDS19, WWW17, WYY16]. Uncertain [KG14a, OAOQ16]. uncertainties [WCF18]. Uncertainty [Ala14, Ian14, TT11, Bic17, Sun19]. Unconditional [HB13]. uncorrelated [Sze17]. uncorrelatedness [Mao18]. Underdispersion [SM17]. Underlying [Che10a, DTE13, SBLC17]. undershoot [CWW16]. Understanding [SO14b, YC11b]. Unemployment [PMC13]. Unequal [AM15, SAJMF12, SAJ13, ASR^{+}19, EDT18, LP19, WK19]. unexpected [DHK17]. Unification [Chi12b]. Unified [FZ10a, JAA11, SO14a, AJT16, Has19, Kam16, Mah17, MN18, Mna17, OH16, Rez17, RCdCN16]. Uniform [ALR19, BSW15, Che11, Dem14, GCW19, GW15b, KOP14, LGG16, LSL19, MZ15, NR12, NFZ11, Rat13, RN13, Sco09, Tan10a, Tee10, TN13, Wan13a, Yan17b, ZK16, BCL16, CL17b, CQ17, Che17b, CWW16, DS17e, Gol16, KNC18, LL16, LOHW19, MAG18, PMW17, QYLL17, RK16, SN16, SM16, SN18, SNS19, TS19, WW17a]. Uniform-Laplace [Sco09]. Uniformity [WQ18, ADKZ16, EQ17, GQC18]. Uniformly [LG16b, MS14b, GC19, Man18, WdX17]. Unifying [APM10, ZLDT14, GL19]. unilateral [ABES18, BPS16a, CC16b]. Unimodal [SB14]. unimodality [Hom17]. union [LS17e]. union-intersection [LS17e]. Unique [GG10b, SK16b]. Uniqueness [MS11b]. Unit [dBCR15, DPS11, Hay17, Kon15a, LS11, Par15b, QTD10, WZT15, CVL19, DLC16, FLY16, GMMA19, Gun16, MMD18, MR16b, PLC17, Sen18a, VM17, XP18, YVCA17, Zha16a, Zha17c, ZS17b]. unit-gamma [MMD18]. unit-inverse [GMMA19]. Unit-Root [QTD10, PLC17, Zha16a]. Units [BJVV15, Sco11, KR17]. Univariate [Chv13, GMN13, NF11a, RR13, KSBN17, KTBN17, Lin16, WYY19]. Universal [FM12, Lu13a, Xin15, BAB^{+}18]. Universally [AIA14, FM17]. universe [Bak17]. University [MS12b, GRV15, Maz19]. Unknown [AAP11, CBB06, JS13, Kum15, PJ10, Pet10, Bic19, CFA17, KT18, KN17a, MBS17, RG19, WCK16, XHY16, ZY17a]. Unlabeled [Ind11]. Unmeasured [Chi09]. unobservable [FP17b, MSL17]. Unobserved [Tsi10, GDPR17]. unpublished [FTF17]. Unrecognized [sK13b, Ko15b, Ko15a]. Unrelated [RAG12, ZSK12, LPdDTH17, LHS19, MA18, Sen17b, SG17h, SSSK18, SSS17c, TS16, TS17d, ZKS18]. Unreliable [BS14, BL19, KC19a, KC19b, TSOT19]. Unreplicated [SAKK15]. unspecified [Pan19, XXMZ18]. unstable [LJW18]. untargeted [Ko17a]. Unusual [Muk11]. Unwanted [Bea13]. unweighted [LD17]. Upcrossings [VPMF14]. updated [LVH19]. Updating [FM10a]. upon [OL19a, Sco11]. Upper [AJ11, BC13, BCDLA12, CP14, GJ15, Gor17, KL12, LDY15, Raq10, SB14, Sun19, Tee12, EG18, GR18a, LG16b, SK17]. upper-bound [GR18a]. Upper-Tail [GJ15]. upside [AE17c, SSSM16]. upside-down [AE17c, SSSM16]. urn [PB18]. Urns [CW14]. Use [ASG17, ASN12, CJ15, DSC^{+}14, HXL13, LB11, PT15, RAB15, ST10, Saj11, SGR^{+}12, SK12b, SKM16b, SP16d, BSGS16, BPV19, Bow17, HKH17, LW17b, SGKB17, SHM16, ST17a, SSS17b, SKM17, Tak18, YSG17, ZW17b]. Used [AA14, Che14, Hol15, LMM15, MCY10, SA10, Bak17, MD17b, vZ17]. Useful [LH11b, Gen17, Yam17b]. user [LPDD19]. user-defined [LPDD19]. USGBs [KYW16]. Using [ARD14, AT15, APM10, AAP11, AC15, ABI12, AARN10, AP12, ADO13, AS15, AYC^{+}12, BA10, BDS11, BAK13, BEZ11, CCL14, CL10a, CC15a, Chi12a, CKD14, CGL13, Coo12, DS09, DB15, DM11, DMST12, DG11, ES14, EPGTH15, EE15b, FH12, FN09, Gar14, Gha19, GÖ12, GVGP11, GGFL11, Gro10, GR15, GE11, Han15, HASA10, HL11, HNV13, HLR11, IP12, JP13, JLL13, KP13, KS15a, KR14, KS10a, KK11b, KS10b, KRB13, Ko10b, Ko11, sK11, Ko14c, LIPMF13, LW17a, LSC12, MSSU14, MGP11, MCY11, MS15, Miy10, ML13, NAMD14, NSL14, NN11b, NRS^{+}14, NA13, ÖD10, Par13a, PK11, PG16, RSV10, RS10, SK13a, SGM11, SSS14a, SRT14, SV10, SKK10, SS13b, SS13d, SS13e, SST14, SNN12, SS14b, SS15b, SLL14, TSS15, TS12, TEF12, WXZ16, WCGZ12, WYL12, WCF18, XLZ^{+}12]. Using [YUSC14a, YHS13, YZFM17, AO18a, AHH16, AMP17, AHAH17, AAL17, ABTR19, ABES18, AEB17, ASeMJ17, BOT17, BBA16, BSH17, BQDH12, BBR17, BAW19, CE19, CSM16, CLZ19, D’A17a, DAC17, DYXW19, DDP18, DAS16b, Den17, Dut17, DS16b, DS17e, FK19, GFS16, GDW18, HS15a, HB16, HB17a, HBM16, Haq18, Har17, HP14, HLR16, Ima17, IBA18, JSSO18, JKJ16, JSC17, KM19a, KGH19, KAR16, KAJH18, KJ17, Ko16b, Kus18, LG18, Lie16, LS17e, LCZA17b, Maz19, RM16, MSK17, MADSAS10, NE17, PYL17, PK18, PLC17, PHH18, PM16b, PT19, RK19, SSD18, SK17, SG17a, SG17b, SF16, SA17a, Sha18c, SG18b, SX18, SM16, SP16b, SP17b, SP17c, ST17b, SSS18, SSKP18, SSSK18, SG19, SSS19b, SKK16, SS16g, SS16i, SSS17c, TS17a, TS17b, TS17c, TS17d, TP19, TR16b, Udo18, Van16, Vas14a, Vas16]. using [WNL18, WZJH19, WF17, Xio16, XM17a, XMP17, YSS19, ZA14, ZB19, ZLD16, ZW19b]. Usual [ÜY13, Arv18]. utility [RS17]. utilization [Sha18a]. Utilizations [Vin13]. Utilize [KG14a]. Utilizing [MH11, Pel17, SG17g].

vacation [HWZ19, KC19a, KY16, LCWG16, TSOT19, YL17]. Vacations [Kem11, BL19, LCWG16, Ye19]. vaccination [FP17a]. Valid [JP13, EW16]. Validation [VBATK11, hYhW14, LTGK17a, LTGK17b, MSB19b, MD17b, NRS^{+}14, Oga19, Ten17, XTS16, YWLZ17]. Validity [Gar14, Kvå17, NC18c]. Valuation [DW13, QWZ14, DYW18]. Value [ATO19, BC13, BGG11, Chr10, Cra14, Fer14, GD12, GMN13, MX15, WWL13, XYL14, AEB17, Ber16b, Bia17, BBS16, CM16c, Fel10, GF17, HLR16, Kvå17, LK19b, LL19b, NCA17, OGD17, RMN16, Roo18, SK17, YZ17a, ZXM16, ZN18]. Value-at-risk [ATO19, XYL14]. value-validity [Kvå17]. Valued [FZ10b, GSB12, Kac14, LMR13, MSP12, RNB12, SSA10, VP10, ZWZ11, BB17a, BCK17, JM16, KB13, KN18c, LPC17, NRP16, Quo17, Rao17a, Shu19, SW19c, YZ18]. Values [AAR11, BG11a, BAT15, BSSJ11, Bi15, FBS10, Gue10, GR15, HXL13, JWZ14, Kur10, LN13a, MFSAS10, MMP14, RL11a, SASMF13, SG14, SS15a, SHHH09, TW13, WL14, AC16, BBA16, BL15, CFA17, DHK17, HMT^{+}17, IBA18, LW17a, LAW^{+}19, NK19, QK17, RSG17, SNSA17, SK17, SBA17, STD18, TD18, TP19, ZXW16, ZCCW18]. Vandermonde [Aoy10, SKK16]. Vapnik [Gey18]. VAR [DP13b, Dvo17, FPR17, Raï15, Slj17, GMP11, LLSM14]. Variability [Hua15, KTW10, MCY10, SHH13]. Variable [AK14a, BC14, Cag12, CA18, DLP15, FQZ14, FWT13, GLP14, GSSCR12, Kim15, LCT12, Mah12a, NGT13, NAD15, OFS13, OE16, SS15a, WS11, WG14, WL15, Xia17a, XQS19, YX15, YWLW18, ZZ18d, AXLJ17, BAL18b, Bal18a, BMQ16, CLZ19, DYXW19, DE17b, DM18, EG18, FWZ^{+}17a, GAS17, GK19a, GYL17, HZ18b, HAZ16, JZTN17, KKS19, KJ17, KDG^{+}18, LK17a, MH16b, Pou17, PLDS17, RG16, RSG17, SKK18, SI17, SG17e, SKK16, VFP17a, VFP17b, WX17, WXZ18, YS19, YLG16a, YLG16b, YL18b, hYhW14, ZZ19b, ZGK17, ZYH18, ZL16f, HD13b]. Variables [AB09, AKL15, AYC^{+}12, (Ba14, BSSJ11, Bi15, Cag12, CA14, DM11, FLW12, FH13a, FHZ12, GdC12, GDS^{+}14, GCR12, Gil14, Guo14, GS15c, HH13, HD13b, HWYW14, HWW15, HP14, Ian12, Imo15, JA10, KYS13, KL11, KLM10, KS12b, KS12c, LWHW11, LSX14, MP15, MS15, MF14, MWM13, MT12, NAMD14, NRW15, QKP12, RON12, SSA10, SK13a, SDJ15, SGM11, SWL14, SWLW14, SHHH09, SK12d, SZD12, Sun11, Sun12, Sun13, Sun15, Tee14, Tho14, ÜY13, VT15, VS10, WY11, WHVH13, WW13c, WJH15, XLL15a, XTG15, XY09, XY10, XM14, Yan19, YX15, ZZ11, ZS12, ZW11, ZLD11, ZHL15, Zör14, AHH16, ASS18, AO18c, BZH19, BOT17, BU17, BD19, BH17a, CLLW16, CF16, CKS17, CG17, CL17a, CXC19, CS19b, CZW16, DeR19, DP17, DWM^{+}17, DTWW17, DRAS17, DW17, FL18a, FC18a, FWW19, FFN^{+}19]. variables [GHSZ19, GSM16, Had17, Hay16, HF19, HSLQ17, HZZ19, HZ19, Kay18, Kun17b, KN18c, LPH16, LWWH16, LSYL17, LXT17, LLWK18, LSZ16, LSW16, LQ17b, LSL18, MM19, MWW18, Men19, MX19, MSK17, NCA16, NC18b, NMAA19, NHV17, NRN^{+}17, NC19b, Ozg19b, Pes17, QC16, Roo17, SAB18, SG17a, SG17b, Sha18a, SWZ16, SZV16, SS17b, SXW17, SYW16, SJY17, SKM16b, SHM16, SP16d, SG17g, ST17a, SKM17, SSS18, SHLQ17, SW19c, SO17, TR16a, Uya17, WHV16, WW19, WGW19, WV16, WG16b, WAJ17, WP19, XTL^{+}17, XDWC18, Xia17b, XLS19, XHH17, XFWC19, Yan18, YHW17, YZG17, YW17a, YJS19, YYWH19, YQC17, YC17, YW18, ZL16b, ZK16, ZZ17a, ZZ17b, Zha17a, ZZJ19, ZZ19e, ZZL19]. Variance [AMZ15, Are14, AZN15, Bea13, BR13, BS15, Bru10, Bur15, CTQ11, CW13, CHP14, CP14, CA14, CLV14, DPA10, DZ12, DG12b, DMT^{+}15, Eid12, Fu12b, Gro10, GE11, GS19, Hu13, Kaz11, LMCH12, Li12b, MS15, MB15, Men10a, Men13b, Men13a, MB10, MS11c, NK15, OJR15, ÖÇO14, PZ15, Par13b, QTD15, Raï15, SGR^{+}12, SG10b, SB14, SL13, SP13, SS13d, SLSC18, SL15, TKW12, Tsa10, ÜY13, VV17, WT15, WX19, XY11a, YUSC13, YUSC14a, YS12, ZARH14, ZYX12, ZLF14, ZYZ11, AHH16, BRZR16, BOT17, BKR19, BCP17, CRTdC17, COK17, CFA17, CFK17, Dvo17, EG18, FLY16, Ghr19, GMH16, Huh16, JPAB17, Jas16, JZQ15, JZZZ18, KT17b, LRZ17, Liu17b, MGLX14, MGN16, MKK17, OPK17, OPK18, PK18, PS17a, QKCH19, Rah18, RPH17, Sen18a, Sen16a, Sen18b, SH17b, SKT16, SSB17c, SP17b]. variance [SPY18, Wan17e, WRZ19, XP18, Zha17d, ZRL19a, ZG17b, ZL16g]. Variance-balanced [VV17]. Variance-Covariance [AMZ15, SP13]. Variance-Stabilizing [NK15]. Variances [HEHM11, Him10, MO10, PV14, Rub10, SAJMF12, SAJ13, ZJC10, Ima18, Mat18, Par17a, RG19, Sin16, WXZ16, WDLZ16, WK19]. Variant [Sha10b, YFV19]. Variants [dSE11, BTC16]. Variate [ASNT15, BREA11, CLGFB15, IANT13, SST14, Zin12, AH17a, NRCG16, NAN17a, TZ17, ZKW16]. Variates [KRT13, AE17a, AE17b, GG16b, Zha17b]. Variation [AJA14, Gro11, Jaf15, KK11b, Kvå15, LS15b, Men10a, Sho15, TL15b, WW13c, ASG17, LMH17, LZF17, Lu16b, LSL18, MH16a, mNK17, SG17b, UK16, vZvdM17]. Variational [YOD10, Udo18]. Variations [LTC17]. Variograms [WWH15a]. Various [PS13, PV16, FJS^{+}18, OQC17]. VARs [OL19a]. Varying [Don11, FH13a, FHZ12, Hua12, HLX14, Lu13b, LSX14, Lu14, MO14b, Wei12, YX15, ZX10, ZZLL15, AV18, BG19c, DW17, FTT16, HH19, Hu14a, HJT19, JQZ18, JHM19, LLHZ16, LZS17, LLWK18, LL18a, LLF19, MJH19, MBS^{+}19, MV16a, RTT17, SHLQ17, WGW19, Wan19b, WX19, WXZ18, XQS19, XFWC19, YLG16a, YLG16b, YSC17a, ZSWA19, ZL16d, ZXF17, ZSS18, ZP17, ZH16, ZSJ16, cZzXgL18, ZY19]. Varying-Coefficient [FH13a, FHZ12, Hua12, Wei12, ZX10, Lu13b, FTT16, LZS17, LLWK18, LL18a, LLF19, RTT17, XFWC19, YLG16a, YSC17a, ZL16d, ZXF17]. vecd [Nag18]. vech [Nag18]. Vector [KK14, NV15, RZ13, SA12, Tak10b, VP10, WDM14, Bud16, CS16a, Cav17, KT17a, MF16, SSH17, Shu19, SA17b, SI19, Xia17b]. Vector-Valued [VP10, Shu19]. Vectors [BC13, LR10a, ESAKK18, Ko18, LCG18, NAC19, ST16c, WN17a, WN17b, YSS18]. verification [TSOT19]. Verma [BRY17]. Version [KA15, KR15, LJ12, Lu10, SBR14, Tor12, FNM19, LWZ17, MB17b, NMAA19, RMM17]. Versions [KLS10, Wan16a]. versus [ABI12, RS17, Wu16a]. very [Nav16]. Via [BCH15, CW15b, MO10, Mar15, NS15b, SDJ15, Tee10, AGA17b, BTC16, BK09, Cha14, CPW19, Con10, DGMC17, EdSMR14, Fau12, FF13, GP19, GMP11, GPÖ14, HA18, HS19, HB13, HH13, HSHW14, IF18, JSZL17, JN19, Jun17, KK11a, KDG^{+}18, Li12a, LD18a, LH17c, LYR^{+}18, LTC17, MS17a, Man11b, NM19, NS19, PS18, PY16b, QWWW19, Rao10, RZ13, SCC11, SUKK19, TN13, TCN15, WDZ11, Wan17a, WX17, WW17a, WLS18, WGG^{+}19, WW15c, WHY18, XLZ^{+}12, XM17b, YGM19, Zho10]. Victoria [BHKHM13]. Viewpoint [Par12]. VIF [HJC17]. VIF-based [HJC17]. Vine [HTW17, JSWL17]. Violation [BDM15, HYK15a]. Virtual [SCC10]. Visualization [HKV12]. Visualizing [JTR15]. Vitality [RASMN14]. Vol [Mül12]. volatilities [CMN^{+}18, MS17a]. Volatility [GJM14, JLL13, LCLJ14, LRZ17, LG18, ZG17a]. Volume [Ano18]. vs [Bon14, HQW15, Spa19]. VSI [HD13b, CY10, ZLT17]. Vulnerable [WW10].

Wadley [LH14a]. Wagner [KA17]. waiting [KV17]. Wald [Arv18, KKS19, Par17b, UDC12, Voi15, WF17, WCF18]. Wald-type [WF17, WCF18]. Walk [AKK09, AK14c, Wan13a, KKKÖ17, aSaCN18, SN18]. Walker [PCL16]. Walks [Yar11, Yar13, Yar14, Dai16]. Wallis [WLZ17]. Warahena [AE23d]. Warahena-Liyanage [AE23d]. Warde [ST16f]. warm [YVCA17]. Warner [Sen16a, Sen17b]. Warped [CW15b]. Warranty [CC10d, JHP10, SCSMN11, ZW19b]. water [ZH16]. Watson [BANC17, Fro17b, Li16, VYZ14]. Waugh [Yam17a]. Wavelet [Afs13, CS15a, CS15b, CF13, Che13, CS14b, HW18a, KOP13, KOP14, KAD14, LX10b, LS11, LPLW13, LX17, Mån12, Man18, NAMD14, QTD10, Rao17b, RMC11, WZY14, WZK17, WL19, ZL14b, cZzXgL18, BD17b, CCN17, CN17b, CDS19, GK19c, ZL17]. Wavelet-Based [CF13, Mån12, HW18a, LX17]. Wavelets [AP12, CW15b, MO10, Rao10, SA17a]. Way [BF10, DBC14, Eid12, KSBM14b, KSBM14a, SAJ13, YSR11, DO19, KS16b, LLZ17, Mao18, MGG^{+}17, SAJMF12, SAKK15, Su17b, WS14, YSS18, ZZ19a]. WCEV [Çel17]. Weak [CF14, CS19b, HV17, HS14, Jas19, SY17d, BB17b, NMAA19, Tch19]. Weakly [FM12, SEA15, LPH16]. Wealth [XZNL15]. Wear [AB13]. Web [Pap13]. Weibull [Ano23, Jon18, AHAH17, AFL15, Alm18, AFL13, AN17d, AAJ18, BJU18, BNB17, BNB19, BNS11, Bel17, BIJ17, CSL17, CK19, Coo12, CR13, hFmH14, FLB16, FA19, GC14, GHBM18, GBD10, GSA14, GW18, HP17a, HSMK13b, HASA10, IKH19, JB18, JS18, KSDR15, KKH17, KCY^{+}19, KN18a, KN19, LWL13, LKK17, LCO14, LS16a, LS17c, MBAJ19, NAMAK17, NB17, NMJY15, OBPL15, PP18, PGF11, PMKF17, PCS18, PC10b, QZY12, RAA17, RK19, RM14, SP16a, SX18, SM14, TCM^{+}16, TGXX12, UM09, VEKH15, WK13, WSS17, XZ16, YWLY11, Yeh12, Yeh13, Yeh15, YCL17, Yu12, YH19, vZ17]. Weibull-distributed [RK19]. Weibull-geometric [BNB17, NB17]. Weibull-Negative-Binomial [YCL17]. Weibull-Tail [GBD10]. Weighing [CG12, CG14a, Gra12, KS12a]. Weight [CG12, Sta17]. Weighted [AMGK15, BJ18, Cha10, sDsWcCjK14, DL16, ER11, GJ15, Guo14, HGZG15, Hua14a, HWM12, Iga16, IRB13, JQZ18, Ko14e, LD18a, LPDD18, LB11, MKS15, MZ10a, Mar11, Men10b, Men13a, NY17, QNZ14, SWLW14, SL13, SHH13, SP11b, Sun17c, Sun15, SS12e, TE11, TKW12, VSS12, WHVH13, WXPL13, WLY14a, WWW13, XM14, YD19b, ZLL11, ZC14, ZYX12, ZYZ11, AR17b, AP17, Ala17, AGA17b, AAJ16, BKM18, BBG18, CKS17, CCZR17, CZW16, Das17b, DW17, DL19, FWW19, GCA18, HZ18b, HZZ19, HN17b, IAB16, Kaç19, Ko16c, KPZ16, Kun17b, LD17, LY16a, Lu18, LD18b, LQW19, MFJ17, MB17b, MBF17, NMAA19, NO18, PCL16, PE17, QC16, RASR17, RLSL19, RMN16, Roo17, SAB18, SWZ16, SXW17, SNW19, SK19b, Ten17, WZ16, WHV16, WHHW17, WH17a, WT18, WHYW18, WV16, WG16b, WA17, XDWC18, XN19, Yan18, YGL16, YJS19]. weighted [YWLW18, YQC17, YKL^{+}17, YC17, YW18, Zen17b, ZK16, ZMMX17, Zha17a, ZX18]. Weighting [GZSS14, Gar12, IRAB13, LIPMF13, Lec11, SW14a, WP11, Zho09, ÇE17a, WGG^{+}19, ZY16]. weighting-based [WGG^{+}19]. weightings [Cle18]. Weights [Par15b, BPV19, LW18a, LPDD19, ST16f, SS16d, YJ19, vZ17]. well [LJC18]. where [PVS17]. Which [Cah15, DN14, DM18, Hua10, SSON11, GGDVP16]. white [CJ17, SW19a]. Whitney [Cuc16, DT17, HH15, WZX17]. Whittaker [YJ19, YD19a]. Whittaker-Henderson [YJ19]. Whittle [BK19, SP13, SMM10]. wide [WHYW18, YC17]. widely [CZW16, DWM^{+}17, LWWH16, LSYL17, QC16]. Wider [Zha15a]. Width [Liu11c, PR11]. Wiener [MS11b, SSC18, SM09, aYwSlDmS17, Zil11]. Wigner [MR15]. Wilcoxon [Che18, HH15]. Wild [MG15]. Wilk [Fly10]. Wilks [BREA11, GC12, TZ17]. Wilson [BSY12]. Win [Hay16]. Win-probabilities [Hay16]. Wind [QZY12]. Window [WG14, TK17b]. Winsorized [Bie16]. wise [CLLW16, WW19]. Wishart [LX10a]. within [Cox19, SN12]. Without [Fu12b, KK14, CHZ17, CS19b, FM17, Moh19b, TH17]. WLAD [YL18b]. WLAD-LASSO [YL18b]. WNOD [Yan18]. WOD [BZH19]. Wolfe [GG16a]. Wolfowitz [AD17c]. Wong [KHY18, PHY19]. word [MC17]. Work [Koz12]. Working [GLZW12, GHY11, HWZ19, JX19, YL18c, ZMN17]. works [Che19a]. worth [NP17b]. Wrap [OQC14, GQC17]. Wrap-around [OQC14, GQC17]. Wrapped [JJ18, LMG13]. Wright [Cah12, Pop12]. written [WQW16]. WUOD [LSL18, SHLQ17]. Wynn [AW10].

Xavier [AE23c]. XII [ACO^{+}18, BAT15, HME17, RAK15, SMSF19, Wat11].

yantram [Kha19b]. Years [CDE10, El 17]. Yield [CC11, Gra11, MN13a, WCL11, WT18]. yields [Kha19b, LS17b]. Youden [SG18b]. Youden-m [SG18b]. young [TE18b]. Yule [PCL16].

Zenga [NS16b, Por15]. Zero [DN14, FWH^{+}15, Fu12c, Gov11, GB15, Mon14, ÖD10, PS11, RBN19, YHA10b, YHA10a, ZI14, ZWZZ15, ZZLL15, BW17, BD16b, BG16a, CYF^{+}18, CTAL17, DDD17, FI17, LMdO18, Mur16a, RC16, SBRGD16, SG16c, SG16e, TSG19, WZ16]. Zero-Acceptance [GB15]. zero-distorted [SBRGD16]. Zero-Inflated [FWH^{+}15, Fu12c, Mon14, ÖD10, PS11, RBN19, YHA10b, ZI14, ZZLL15, BW17, BD16b, BG16a, CYF^{+}18, DDD17, FI17, LMdO18, RC16, SG16c, SG16e, TSG19]. Zero-Inflation [YHA10a]. zero-modified [CTAL17]. Zero-recurrence [ZWZZ15]. Zeros [DZ15]. ZIP [FNDM12]. Zografos [NCO15]. zone [EM18, Rav19]. Zygmund [Yan19, Ko13].

References

Abbasnejad:2010:REP

[AA10] M. Abbasnejad and N. R. Arghami. Renyi entropy properties of order statistics. Communications in Statistics: Theory and Methods, 40(1):40–52, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:2011:CNB

[AA11a] Rashid Ahmed and Munir Akhtar. Construction of neighborbalanced designs in linear blocks. Communications in Statistics: Theory and Methods, 40(17):3198–3205, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Almendra-Arao:2011:BCS

[AA11b] Félix Almendra-Arao. Barnard convex sets. Communications in Statistics: Theory and Methods, 40(14):2574–2582, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abbas:2013:MFI

[AA13] Nasir Abbas and Muhammad Aslam. Modeling the factors influencing the evaluations of items through mixture models for preference datasets. Communications in Statistics: Theory and Methods, 42(1):135–144, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asadi:2014:FPU

[AA14] Majid Asadi and Amir R. Asadi. On the failure probability of used coherent systems. Communications in Statistics: Theory and Methods, 43(10–12):2468–2475, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abouammoh:2015:DGD

[AA15a] A. M. Abouammoh and Najla S. Alhazzani. On discrete gamma distribution. Communications in Statistics: Theory and Methods, 44(14):3087–3098, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alvarez-Andrade:2015:CCH

[AA15b] Sergio Alvarez-Andrade. On complete convergence for the hybrid process. Communications in Statistics: Theory and Methods, 44(5):1044–1052, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdel-Aty:2017:ELI

[AA17a] Yahia Abdel-Aty. Exact likelihood inference for two populations from two-parameter exponential distributions under joint Type-II censoring. Communications in Statistics: Theory and Methods, 46(18):9026–9041, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1200093.

Akpoue:2017:SCM

[AA17b] Blache Paul Akpoue and Jean-François Angers. Some contributions on the multivariate Poisson–Skellam probability distribution. Communications in Statistics: Theory and Methods, 46 (1):49–68, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.983806.

Alawadhi:2017:STI

[AA17c] S. A. Alawadhi and F. A. Alawadhi. Spatial-temporal interpolation of non methane hydrocarbon levels in Kuwait. Communications in Statistics: Theory and Methods, 46(6):2764–2779, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.968731.

Abbasi:2019:SBR

[AA19] Somayeh Abbasi and Mohammasd Hossein Alamatsaz. Some bounds related to Harris family of distributions. Communications in Statistics: Theory and Methods, 48(16):4082–4095, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1487981.

Alvarez-Andrade:2017:SAR

[AABL17] Sergio Alvarez-Andrade, Salim Bouzebda, and Aimé Lachal. Some asymptotic results for the integrated empirical process with applications to statistical tests. Communications in Statistics: Theory and Methods, 46(7):3365–3392, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1060346.

Alhejaili:2013:SAS

[AAE13] Amal D. Alhejaili and Ehab F. Abd-Elfattah. Saddlepoint approximations for stopped-sum distributions. Communications in Statistics: Theory and Methods, 42(20):3735–3743, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adebola:2017:HTR

[AAE17] Femi B. Adebola, Adedamola A. Adediran, and Olusegun S. Ewemooje. Hybrid tripartite randomized response technique. Communications in Statistics: Theory and Methods, 46(23):11756–11763, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1277760.

Abouelmagd:2019:NAC

[AAE^{+}19] T. H. M. Abouelmagd, A. A. E. Ahmed, Enayat M. Abd Elrazik, Mahmoud M. Mansour, A-Hadi N. Ahmed, and Ahmed Z. Afify. A new aging concept derived from the increasing convex ordering. Communications in Statistics: Theory and Methods, 48(12):2904–2916, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473593.

Ahmad:2012:MAP

[AAG12a] Tanvir Ahmad, Munir Akhtar, and Steven G. Gilmour. Multilevel augmented pairs second-order response surface designs and their robustness to missing data. Communications in Statistics: Theory and Methods, 41(3):437–452, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arshad:2012:ABB

[AAG12b] Hafiz Muhammad Arshad, Munir Akhtar, and Steven G. Gilmour. Augmented Box–Behnken designs for fitting thirdorder response surfaces. Communications in Statistics: Theory and Methods, 41(23):4225–4239, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aslam:2016:MSP

[AAJ16] Muhammad Aslam, Muhammad Azam, and Chi-Hyuck Jun. Mixed sampling plan based on exponentially weighted moving average statistic. Communications in Statistics: Theory and Methods, 45(22):6709–6719, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
966841.

Aslam:2017:NSP

[AAJ17] M. Aslam, M. Azam, and C.-H. Jun. A new sampling plan under the exponential distribution. Communications in Statistics: Theory and Methods, 46(2):644–652, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.1002936.

Aslam:2018:HCC

[AAJ18] Muhammad Aslam, Muhammad Azam, and Chi-Hyuck Jun. A HEWMA-CUSUM control chart for the Weibull distribution. Communications in Statistics: Theory and Methods, 47(24): 5973–5985, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1404100.

Aliyev:2016:AAR

[AAK16] Rovshan Aliyev, Ozlem Ardic, and Tahir Khaniyev. Asymptotic approach for a renewal-reward process with a general interference of chance. Communications in Statistics: Theory and Methods, 45(14):4237–4248, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.917679.

Ashour:2017:SIT

[AAK17] S. K. Ashour and O. E. Abo-Kasem. Statistical inference for two exponential populations under joint progressive Type-I censored scheme. Communications in Statistics: Theory and Methods, 46(7):3479–3488, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1065329.

Altaf:2017:DMP

[AAL17] Saima Altaf, Muhammad Aslam, and Madiha Liaqat. Decisionmaking for paired comparison using the extended amended Davidson model. Communications in Statistics: Theory and Methods, 46(4):1766–1778, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1026998.

Athayde:2012:ABS

[AALS12] Emilia Athayde, Cecilia Azevedo, Víctor Leiva, and Antonio Sanhueza. About Birnbaum–Saunders distributions based on the Johnson system. Communications in Statistics: Theory and Methods, 41(11):2061–2079, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alodat:2014:SGP

[AAM14] M. T. Alodat and E. Y. Al-Momani. Skew Gaussian process for nonlinear regression. Communications in Statistics: Theory and Methods, 43(23):4936–4961, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alzaid:2017:BRM

[AAO17] Abdulhamid A. Alzaid, Fatimah E. Almuhayfith, and Maha A. Omair. Bivariate regression models based on compound Poisson distribution. Communications in Statistics: Theory and Methods, 46(15):7375–7389, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1152483.

Ahmed:2011:IUH

[AAP11] Munir Ahmed, Muhammad Aslam, and G. R. Pasha. Inference under heteroscedasticity of unknown form using an adaptive estimator. Communications in Statistics: Theory and Methods, 40(24):4431–4457, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amini:2011:FIR

[AAR11] Morteza Amini, Jafar Ahmadi, and M. Razmkhah. Fisher information in record values and their concomitants: a comparison of two sampling schemes. Communications in Statistics: Theory and Methods, 40(7):1298–1314, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alodat:2010:IAR

[AARN10] M. T. Alodat, M. Y. Al-Rawwash, and I. M. Nawajah. Inference about the regression parameters using median-ranked set sampling. Communications in Statistics: Theory and Methods, 39(14):2604–2616, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Almendra-Arao:2012:SPN

[AASR12] Félix Almendra-Arao and David Sotres-Ramos. Some properties of non inferiority tests for two independent probabilities. Communications in Statistics: Theory and Methods, 41(9):

1636–1646, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Almendra-Arao:2017:EBT

[AASRZE17] Félix Almendra-Arao, David Sotres-Ramos, and Magin Zuñiga-Estrada. Extending the Barnard’s test to non-inferiority. Communications in Statistics: Theory and Methods, 46(13):6293– 6302, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2013.875577.

Aggarwal:2009:RBP

[AB09] Priyanka Aggarwal and Ashok K. Bansal. Robustness of Bayes prediction under error-in-variables superpopulation model. Communications in Statistics: Theory and Methods, 39(2):327– 339, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Afanasyeva:2011:SMT

[AB11a] Larisa Afanasyeva and Ekaterina Bulinskaya. Stochastic models of transport flows. Communications in Statistics: Theory and Methods, 40(16):2830–2846, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ato:2011:CSL

[AB11b] José Candel Ato and Milenko Bernadic. The characterization of the symmetric Lorenz curves by doubly truncated mean function. Communications in Statistics: Theory and Methods, 40 (18):3269–3280, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anis:2013:APT

[AB13] M. Z. Anis and Ishita Basak. On C_{pk} assessment in the presence of tool wear. Communications in Statistics: Theory and Methods, 42(10):1759–1773, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmadi:2014:NRP

[AB14] Mosayeb Ahmadi and G. R. Mohtashami Borzadaran. On the nearness of records to population quantiles with respect to order statistics in the sense of Pitman closeness. Communications in Statistics: Theory and Methods, 43(20):4357–4370, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abbasnejad:2015:SRD

[AB15] M. Abbasnejad and G. R. Mohtashami Borzadaran. Some results on dynamic generalized survival entropy. Communications in Statistics: Theory and Methods, 44(8):1653–1668, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:2017:BPP

[AB17] Priyanka Aggarwal and Ashok K. Bansal. Bayes prediction of Poisson regression superpopulation mean with a non gamma prior. Communications in Statistics: Theory and Methods, 46(11):5531–5543, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1104355.

Arshad:2019:ECL

[AB19] Mohd. Arshad and Ayman Baklizi. Estimation of common location parameter of two exponential populations based on records. Communications in Statistics: Theory and Methods, 48(6):1545–1552, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1435805.

Atlagh:2012:CRC

[ABC12] M. Atlagh, M. Broniatowski, and G. Celant. A convergence result for a class of asymptotically linear estimators. Communications in Statistics: Theory and Methods, 41(16–17):2922– 2931, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/03610926.2011.593286.

Abdushukurov:2015:ESF

[Abd15] A. A. Abdushukurov. Estimation of survival function in Cox regression model under random censoring from both sides. Communications in Statistics: Theory and Methods, 44(3):533–553, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arbia:2018:FSR

[ABES18] Giuseppe Arbia, Marco Bee, Giuseppe Espa, and Flavio Santi. Fitting spatial regressions to large datasets using unilateral approximations. Communications in Statistics: Theory and Methods, 47(1):222–238, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1301476.

Aljawadi:2012:NVP

[ABI12] Bader A. I. Aljawadi, Mohd Rizam A. Bakar, and Noor Akma Ibrahim. Nonparametric versus parametric estimation of the cure fraction using interval censored data. Communications in Statistics: Theory and Methods, 41(23):4251–4275, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albatineh:2017:NCI

[ABK17] Ahmed N. Albatineh, Ibrahimou Boubakari, and B. M. Golam Kibria. New confidence interval estimator of the signal-to-noise ratio based on asymptotic sampling distribution. Communications in Statistics: Theory and Methods, 46(2):574–590, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.1000498.

Arashi:2015:IRQ

[ABLR15] M. Arashi, A. Bekker, M. T. Loots, and J. J. J. Roux. Integral representation of quaternion elliptical density and its applications. Communications in Statistics: Theory and Methods, 44 (4):778–789, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andreopoulos:2019:MMU

[ABTR19] Panagiotis Andreopoulos, G. Fragkiskos Bersimis, Alexandra Tragaki, and Antonis Rovolis. Mortality modeling using probability distributions. Application in Greek mortality data. Communications in Statistics: Theory and Methods, 48(1):127–140, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1501485.

Abuzaid:2018:HCD

[Abu18] Ali H. Abuzaid. A half circular distribution for modeling the posterior corneal curvature. Communications in Statistics: Theory and Methods, 47(13):3118–3124, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1348521.

Ahmed:2011:DBA

[AC11] S. Ejaz Ahmed and S. Chitsaz. Data-based adaptive estimation in an investment model. Communications in Statistics: Theory and Methods, 40(19–20):3540–3554, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aladag:2015:IEP

[AC15] Sibel Aladag and Hulya Cingi. Improvement in estimating the population median in simple random sampling and stratified random sampling using auxiliary information. Communications in Statistics: Theory and Methods, 44(5):1013–1032, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asha:2016:RRE

[AC16] P. S. Asha and Manoj Chacko. Residual Renyi entropy of k-record values. Communications in Statistics: Theory and Methods, 45(16):4874–4885, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
932806.

Acar:2015:IAB

[Aca15] Tülin Acar. An investigation of agreement between the item difficulty coefficient calculated in accordance with classical test theory and item response theory with Bland–Altman method. Communications in Statistics: Theory and Methods, 44(21): 4614–4621, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Afify:2018:FPB

[ACO^{+}18] Ahmed Z. Afify, Gauss M. Cordeiro, Edwin M. M. Ortega, Haitham M. Yousof, and Nadeem Shafique Butt. The four-parameter Burr XII distribution: Properties, regression model, and applications. Communications in Statistics: Theory and Methods, 47(11):2605–2624, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1231821.

Andel:2010:BPN

[AD10] Jirí Andel and Sárka Doslá. Bernoulli processes with nonpositive correlations. Communications in Statistics: Theory and Methods, 39(17):3191–3201, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arab:2016:CSA

[AD16] Idir Arab and Abdelnasser Dahmani. Consistency of stochastic approximation algorithm with quasi-associated random errors. Communications in Statistics: Theory and Methods, 45(23): 6883–6890, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.968737.

Aiane:2017:CTR

[AD17a] Nabila Aiane and Abdelnasser Dahmani. Consistency of the Tikhonov’s regularization in an ill-posed problem with αmixing random data. Communications in Statistics: Theory and Methods, 46(11):5633–5642, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1107586.

Aiane:2017:RCR

[AD17b] Nabila Aiane and Abdelnasser Dahmani. On the rate of convergence of the Robbins–Monro’s algorithm in a linear stochastic ill-posed problem with α-mixing data. Communications in Statistics: Theory and Methods, 46(13):6694–6703, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1133825.

Arab:2017:KWA

[AD17c] Idir Arab and Abdelnasser Dahmani. Kiefer–Wolfowitz algorithm under quasi-associated random errors. Communications in Statistics: Theory and Methods, 46(6):2686–2695, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1044668.

Aminzadeh:2019:BPM

[AD19] M. S. Aminzadeh and M. Deng. Bayesian predictive modeling for inverse gamma-Pareto composite distribution. Communications in Statistics: Theory and Methods, 48(8):1938–1954, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1440595.

Assane:2018:BST

[AdBPdBP18] Cachimo Combo Assane, Basilio de Bragança Pereira, and Carlos Alberto de Bragança Pereira. Bayesian significance test for discriminating between survival distributions. Communications in Statistics: Theory and Methods, 47(24):6095–6107, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1406117.

Androulakis:2016:MUE

[ADKZ16] E. Androulakis, K. Drosou, C. Koukouvinos, and Y.-D. Zhou. Measures of uniformity in experimental designs: a selective overview. Communications in Statistics: Theory and Methods, 45(13):3782–3806, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
966843.

Andrade:2013:RBM

[ADO13] J. A. A. Andrade, C. C. Y. Dorea, and C. E. Guevara Otiniano. On the robustness of Bayesian modelling of location and scale structures using heavy-tailed distributions. Communications in Statistics: Theory and Methods, 42(8):1502–1514, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmadi:2010:BEB

[ADP10] Jafar Ahmadi, Mahdi Doostparast, and Ahmad Parsian. Bayes estimation based on random censored data for some life time models under symmetric and asymmetric loss functions. Communications in Statistics: Theory and Methods, 39(17):3058– 3071, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ardakani:2011:ESO

[ADWR11] Mostafa K. Ardakani, Debashis Das, Shaun S. Wulff, and Timothy J. Robinson. Estimation in second-order models with errors in the factor levels. Communications in Statistics: Theory and Methods, 40(9):1573–1590, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abd-Elfattah:2013:SAD

[AE13] Ehab F. Abd-Elfattah. Saddlepoint approximations to the density and the distribution functions of linear combinations of ratios of partial sums. Communications in Statistics: Theory and Methods, 42(23):4185–4198, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abd-Elfattah:2017:DDF

[AE17a] Ehab F. Abd-Elfattah. Densities and distribution functions of the ratio of two linear functions and the product of gamma variates: A saddlepoint approach. Communications in Statistics: Theory and Methods, 46(15):7327–7336, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1148736.

Abd-Elfattah:2017:SDD

[AE17b] Ehab F. Abd-Elfattah. Saddlepoint density and distribution functions for the ratio of two linear functions and the product of generalized gamma variates. Communications in Statistics: Theory and Methods, 46(23):11745–11755, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1277758.

Abd-Elrahman:2017:NTP

[AE17c] Ayman M. Abd-Elrahman. A new two-parameter lifetime distribution with decreasing, increasing or upside-down bathtub-shaped failure rate. Communications in Statistics: Theory and Methods, 46(18):8865–8880, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1193198.

Azimi:2023:CGK

[AE23a] Reza Azimi and Mahdy Esmailian. Correction to: Ghosh, I. K., and Hamedani, G. G. (2018). The Gamma-Kumaraswamy distribution: an alternative to Gamma distribution. Communications in Statistics-Theory and Methods, 47, 2056–2072. Communications in Statistics: Theory and Methods, 52(18):??, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2021.1999476. See [GH18].

Azimi:2023:CGS

[AE23b] Reza Azimi and Mahdy Esmailian. Correction to: Gomes, A. E., da-Silva, C., Cordeiro, G.M. (2015). The exponentiated G Poisson model. Communications in Statistics — Theory and Methods, 44:20, 4217–4240. Communications in Statistics: Theory and Methods, 52(16):??, 2023. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2021.2018464. See [GDSC15a].

Azimi:2023:COJ

[AE23c] Reza Azimi and Mahdy Esmailian. Correction to: Oliveira, J., J. Santos, C. Xavier, D. Trindade, and G. M. Cordeiro. 2016. The McDonald half-logistic distribution: Theory and practice. Communications in Statistics — Theory and methods 45 (7):2005–2022. Communications in Statistics: Theory and Methods, 52(17):??, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2021.

1931336. See [OSX^{+}16].

Azimi:2023:CPM

[AE23d] Reza Azimi and Mahdy Esmailian. Correction to: Pararai, M., Warahena-Liyanage, G., and Oluyede, B. (2017). Exponentiated power Lindley Poisson distribution: Properties and applications. Communications in Statistics — Theory and Methods, 46, 4726–4755. Communications in Statistics: Theory and Methods, 52(18):??, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2021.

2009871. See [PWLO17].

Asar:2017:DRT

[AEA17] Yasin Asar, Murat Erisoglu, and Mohammad Arashi. Developing a restricted two-parameter Liu-type estimator: a comparison of restricted estimators in the binary logistic regression model. Communications in Statistics: Theory and Methods, 46(14):6864–6873, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1137597.

Ashoori:2017:MMP

[AEB17] Farnoosh Ashoori, Malihe Ebrahimpour, and Abolghasem Bozorgnia. Modeling of maximum precipitation using maximal generalized extreme value distribution. Communications in Statistics: Theory and Methods, 46(6):3025–3033, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1034325.

Attia:2011:OPM

[AEH11] A. F. Attia, E. D. Abou Elela, and H. A. Hosham. The optimal preventive maintenance policy for the multistate system profit. Communications in Statistics: Theory and Methods, 40(23): 4189–4199, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alawady:2017:ARE

[AEHQ17] M. A. Alawady, A. M. Elsawah, Jianwei Hu, and Hong Qin. Asymptotic random extremal ratio and product based on generalized order statistics and its dual. Communications in Statistics: Theory and Methods, 46(18):8881–8896, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1197255.

Asgharzadeh:2016:BSL

[AEN16] A. Asgharzadeh, L. Esmaeili, and S. Nadarajah. Balakrishnan skew logistic distribution. Communications in Statistics: Theory and Methods, 45(2):444–464, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asha:2016:FRD

[AER16] G. Asha, I. Elbatal, and C. J. Rejeesh. Further results on discrete mean past lifetime. Communications in Statistics: Theory and Methods, 45(4):1081–1098, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.857415.

Alazemi:2019:ESE

[AESO19] Fares Alazemi, Khalifa Es-Sebaiy, and Youssef Ouknine. Efficient and superefficient estimators of filtered Poisson process intensities. Communications in Statistics: Theory and Methods, 48(7):1682–1692, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1438622.

Asgharzadeh:2010:EPE

[AF10] A. Asgharzadeh and A. Fallah. Estimation and prediction for exponentiated family of distributions based on records. Communications in Statistics: Theory and Methods, 40(1):68–83, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alvarez:2016:REA

[ÁF16] Enrique E. Álvarez and Julieta Ferrario. Robust estimation in the additive hazards model. Communications in Statistics: Theory and Methods, 45(4):906–921, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.853790.

Afendras:2013:MBI

[Afe13] G. Afendras. Moment-based inference for Pearson’s quadratic q subfamily of distributions. Communications in Statistics: Theory and Methods, 42(12):2271–2280, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alzaatreh:2013:WPD

[AFL13] Ayman Alzaatreh, Felix Famoye, and Carl Lee. Weibull–Pareto distribution and its applications. Communications in Statistics: Theory and Methods, 42(9):1673–1691, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aljarrah:2015:NWP

[AFL15] Mohammad A. Aljarrah, Felix Famoye, and Carl Lee. A new Weibull–Pareto distribution. Communications in Statistics: Theory and Methods, 44(19):4077–4095, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic). See correction [Ano23].

Aljarrah:2019:NGN

[AFL19] Mohammad A. Aljarrah, Felix Famoye, and Carl Lee. A new generalized normal distribution: Properties and applications. Communications in Statistics: Theory and Methods, 48(18): 4474–4491, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1483509.

Afshari:2013:FWA

[Afs13] M. Afshari. A fast wavelet algorithm for analyzing onedimensional signal processing and asymptotic distribution of wavelet coefficients with numerical example and simulation. Communications in Statistics: Theory and Methods, 42(22): 4156–4169, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Atilgan:2011:LMS

[AG11] Y. Kayhan Atilgan and Suleyman Gunay. Least median of squares solution of multiple linear regression models through the origin. Communications in Statistics: Theory and Methods, 40(22):4125–4137, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alzaatreh:2016:SGP

[AG16] Ayman Alzaatreh and Indranil Ghosh. A study of the gamma-Pareto (IV) distribution and its applications. Communications in Statistics: Theory and Methods, 45(3):636–654, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2017:SAB

[AG17] Barry C. Arnold and Indranil Ghosh. Some alternative bivariate Kumaraswamy models. Communications in Statistics: Theory and Methods, 46(18):9335–9354, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1208244.

Azais:2018:NCJ

[AG18] Romain Azaïs and Alexandre Genadot. A new characterization of the jump rate for piecewise-deterministic Markov processes with discrete transitions. Communications in Statistics: Theory and Methods, 47(8):1812–1829, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1327072.

Aga:2011:EEL

[Aga11] Mosisa Aga. Edgeworth expansion for linear regression processes with long-memory errors. Communications in Statistics: Theory and Methods, 40(4):663–673, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agahi:2017:MCI

[Aga17a] Hamzeh Agahi. Multivariate Chebyshev inequality in generalized measure space based on Choquet integral. Communications in Statistics: Theory and Methods, 46(19):9625–9628, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1213293.

Arnold:2017:CBM

[AGA17b] Barry C. Arnold, Indranil Ghosh, and Ayman Alzaatreh. Construction of bivariate and multivariate weighted distributions via conditioning. Communications in Statistics: Theory and Methods, 46(18):8897–8912, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1197256.

Ahsanullah:2017:CLD

[AGAM17] M. Ahsanullah, M. E. Ghitany, and D. K. Al-Mutairi. Characterization of Lindley distribution by truncated moments. Communications in Statistics: Theory and Methods, 46(12):6222– 6227, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1124117.

Astorga:2017:SGE

[AGB17] Juan M. Astorga, Héctor W. Gómez, and Heleno Bolfarine. Slashed generalized exponential distribution. Communications in Statistics: Theory and Methods, 46(5):2091–2102, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1032426.

Adnaik:2015:SAC

[AGR15] S. B. Adnaik, M. P. Gadre, and R. N. Rattihalli. Single attribute control charts for a Markovian-dependent process. Communications in Statistics: Theory and Methods, 44(17):3723– 3737, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alanazi:2018:CAE

[AGS18] T. Alanazi, S. D. Georgiou, and S. Stylianou. Construction and analysis of edge designs from skew-symmetric supplementary difference sets. Communications in Statistics: Theory and Methods, 47(20):5064–5076, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1386310.

Arrue:2010:SNC

[AGVB10] Jaime Arrue, Héctor W. Gómez, Héctor Varela, and Heleno Bolfarine. On the skew-normal-Cauchy distribution. Communications in Statistics: Theory and Methods, 40(1):15–27, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ausin:2011:BAM

[AGVGP^{+}11] M. C. Ausín, M. A. Gómez-Villegas, B. González-Pérez, M. T. Rodríguez-Bernal, I. Salazar, and L. Sanz. Bayesian analysis of multiple hypothesis testing with applications to microarray experiments. Communications in Statistics: Theory and Methods, 40(13):2276–2291, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahsanullah:2013:CCD

[AH13] M. Ahsanullah and G. G. Hamedani. Characterizations of continuous distributions based on conditional expectation of generalized order statistics. Communications in Statistics: Theory and Methods, 42(19):3608–3613, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderlucci:2014:CCD

[AH14] Laura Anderlucci and Christian Hennig. The clustering of categorical data: a comparison of a model-based and a distancebased approach. Communications in Statistics: Theory and Methods, 43(4):704–721, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2017:BMM

[AH17a] Barry C. Arnold and G. G. Hamedani. Bivariate, multivariate, and matrix variate normal characterizations: a brief survey II. Communications in Statistics: Theory and Methods, 46(24):11949–11971, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1285923.

Azeem:2017:JIM

[AH17b] Muhammad Azeem and Muhammad Hanif. Joint influence of measurement error and non response on estimation of population mean. Communications in Statistics: Theory and Methods, 46(4):1679–1693, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1026992.

Ali:2018:MGC

[AH18a] Rizwan Ali and Abdul Haq. A mixed GWMA-CUSUM control chart for monitoring the process mean. Communications in Statistics: Theory and Methods, 47(15):3779–3801, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361994.

Aliat:2018:MSP

[AH18b] Billel Aliat and Fayçal Hamdi. On Markov-switching periodic ARMA models. Communications in Statistics: Theory and Methods, 47(2):344–364, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1303734.

Al-Hussaini:2017:BIU

[AHAH17] Essam K. Al-Hussaini and Alaa H. Abdel-Hamid. Bayes inference using half-logistic generated Weibull model based on type-II censoring. Communications in Statistics: Theory and Methods, 46(5):2103–2122, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1032427.

Abouelmagd:2018:TEW

[AHEA18] T. H. M. Abouelmagd, M. S. Hamed, Abd El Hadi N. Ebraheim, and Ahmed Z. Afify. Toward the evaluation of P (X_{(t)} > Y_{(t)}) when both X_{(t)} and Y_{(t)} are inactivity times of two systems. Communications in Statistics: Theory and Methods, 47(14):3293–3304, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1353624.

Ahmad:2016:EFP

[AHH16] Zahoor Ahmad, Ijaz Hussain, and Muhammad Hanif. Estimation of finite population variance in successive sampling using multi-auxiliary variables. Communications in Statistics: Theory and Methods, 45(3):553–565, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akahira:2016:SOA

[AHKO16] M. Akahira, S. Hashimoto, K. Koike, and N. Ohyauchi. Secondorder asymptotic comparison of the MLE and MCLE for a twosided truncated exponential family of distributions. Communications in Statistics: Theory and Methods, 45(19):5637–5659, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.948202.

Ahmad:2016:BEP

[Ahm16] Peer Bilal Ahmad. On the Bayes estimators of the parameters of size-biased generalized power series distributions. Communications in Statistics: Theory and Methods, 45(12):3612–3624, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.904353.

Ahmad:2017:THS

[Ahm17] M. Rauf Ahmad. Testing homogeneity of several covariance matrices and multi-sample sphericity for high-dimensional data under non-normality. Communications in Statistics: Theory and Methods, 46(8):3738–3753, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1073310.

Amato:2019:DPB

[AHP19] Valeria D. Amato, Steven Haberman, and Gabriella Piscopo. The dependency premium based on a multifactor model for dependent mortality data. Communications in Statistics: Theory and Methods, 48(1):50–61, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1366523.

Ahmed:2017:DSC

[AHZ17] Mumtaz Ahmed, Gulfam Haider, and Asad Zaman. Detecting structural change with heteroskedasticity. Communications in Statistics: Theory and Methods, 46(21):10446–10455, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1235200.

Ai:2017:KLE

[Ai17] Xiaohui Ai. Karhunen–Loeve expansion for the additive detrended Brownian motion. Communications in Statistics: Theory and Methods, 46(16):8210–8216, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1177079.

Ahmed:2014:UOD

[AIA14] Rashid Ahmed, Ijaz Iqbal, and Munir Akhtar. Universally optimal designs balanced for neighbor effects. Communications in Statistics: Theory and Methods, 43(16):3371–3379, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Astuti:2017:BMM

[AIIK17] Ani Budi Astuti, Nur Iriawan, Irhamah, and Heri Kuswanto. Bayesian mixture model averaging for identifying the different gene expressions of chickpea (Cicer arietinum) plant tissue. Communications in Statistics: Theory and Methods, 46(21):10564–10581, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1239112.

Ahmadi-Javid:2011:LUB

[AJ11] Amir Ahmadi-Javid. Lower and upper bounds for rectangular probabilities. Communications in Statistics: Theory and Methods, 40(8):1352–1362, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andreassen:2018:TDN

[AJ18a] Camilla Mondrup Andreassen and Jens Ledet Jensen. On the three-dimensional negative binomial distribution. Communications in Statistics: Theory and Methods, 47(10):2314–2326, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.864766.

Ankargren:2018:LSM

[AJ18b] Sebastian Ankargren and Shaobo Jin. On the least-squares model averaging interval estimator. Communications in Statistics: Theory and Methods, 47(1):118–132, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1300272.

Amiri:2019:SOM

[AJ19] Mehdi Amiri and Ahad Jamalizadeh. Stochastic ordering of medians in samples from normal distributions. Communications in Statistics: Theory and Methods, 48(13):3413–3420, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1494835.

Al-Jarallah:2014:NTC

[AJA14] Reem Al-Jarallah and Emad-Eldin A. A. Aly. Nonparametric tests for comparing several coefficients of variation. Communications in Statistics: Theory and Methods, 43(17):3602–3613, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Jarallah:2014:PHM

[AJGG14] R. A. Al-Jarallah, M. E. Ghitany, and Ramesh C. Gupta. A proportional hazard Marshall–Olkin extended family of distributions and its application to Gompertz distribution. Communications in Statistics: Theory and Methods, 43(20):4428–4443, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arashi:2015:SRR

[AJN15] M. Arashi, M. Janfada, and M. Norouzirad. Singular ridge regression with stochastic constraints. Communications in Statistics: Theory and Methods, 44(6):1281–1292, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amiri:2016:SMS

[AJT16] Mehdi Amiri, Ahad Jamalizadeh, and Mina Towhidi. Some multivariate singular unified skew-normal distributions and their application. Communications in Statistics: Theory and Methods, 45(8):2159–2171, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
875575.

Abeynayake:2011:NBB

[AJV11] N. R. Abeynayake, Seema Jaggi, and Cini Varghese. Neighborbalanced bipartite block designs. Communications in Statistics: Theory and Methods, 40(22):4041–4052, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aoshima:2010:PCC

[AK10] Makoto Aoshima and Ken-Ichi Koike. The professional career and contributions of Professor Masafumi Akahira. Communications in Statistics: Theory and Methods, 39(8–9):1324–1342, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alheety:2013:MLT

[AK13] Mustafa Ismaeel Alheety and B. M. Golam Kibria. Modified Liu-type estimator based on (r − k) class estimator. Communications in Statistics: Theory and Methods, 42(2):304–319, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agresti:2014:SRL

[AK14a] Alan Agresti and Maria Kateri. Some remarks on latent variable models in categorical data analysis. Communications in Statistics: Theory and Methods, 43(4):801–814, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alheety:2014:GSR

[AK14b] M. I. Alheety and B. M. Golam Kibria. A generalized stochastic restricted ridge regression estimator. Communications in Statistics: Theory and Methods, 43(20):4415–4427, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aliyev:2014:MSM

[AK14c] Rovshan Aliyev and Tahir Khaniyev. On the moments of a semi-Markovian random walk with Gaussian distribution of summands. Communications in Statistics: Theory and Methods, 43(1):90–104, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alevizakos:2019:ACI

[AK19] Vasileios Alevizakos and Christos Koukouvinos. An asymptotic confidence interval for the process capability index C_{pm}. Communications in Statistics: Theory and Methods, 48(20): 5138–5144, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1508713.

Akahira:2010:FSO

[Aka10] Masafumi Akahira. The first- and second-order large-deviation efficiency for an exponential family and certain curved exponential models. Communications in Statistics: Theory and Methods, 39(8–9):1387–1403, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akici:2011:LE

[Aki11] Fatih Akici. Letter to the Editor. Communications in Statistics: Theory and Methods, 40(23):4322–4323, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aliyev:2009:AEM

[AKK09] Rovshan Aliyev, Tahir Khaniyev, and Tulay Kesemen. Asymptotic expansions for the moments of the semi-Markovian random walk with gamma distributed interference of chance. Communications in Statistics: Theory and Methods, 39(1):130–143, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ah-Kine:2015:OSC

[AKL15] P. Ah-Kine and W. Liu. Optimal simultaneous confidence bands in multiple linear regression with predictor variables constrained in an ellipsoidal region. Communications in Statistics: Theory and Methods, 44(3):441–452, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arashi:2013:SEU

[AKTS13] M. Arashi, Shahjahan Khan, S. M. M. Tabatabaey, and H. Soleimani. Shrinkage estimation under multivariate elliptic models. Communications in Statistics: Theory and Methods, 42(11):2084–2103, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arvanitis:2015:LTQ

[AL15] Stelios Arvanitis and Alexandros Louka. Limit theory for the QMLE of the GQARCH(1, 1) model. Communications in Statistics: Theory and Methods, 44(17):3549–3575, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:2016:NFC

[AL16] Gordon Anderson and Teng Wah Leo. A note on a family of criteria for evaluating test statistics. Communications in Statistics: Theory and Methods, 45(11):3138–3144, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.897137.

Alam:2014:LPG

[Ala14] Md. Moudud Alam. Likelihood prediction for generalized linear mixed models under covariate uncertainty. Communications in Statistics: Theory and Methods, 43(2):219–234, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alavi:2017:GCF

[Ala17] S. M. R. Alavi. A generalized class of form-invariant bivariate weighted distributions. Communications in Statistics: Theory and Methods, 46(5):2193–2201, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1035395.

Alexandrovich:2016:PML

[Ale16] Grigory Alexandrovich. Penalized maximum likelihood estimation for Gaussian hidden Markov models. Communications in Statistics: Theory and Methods, 45(20):6133–6148, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.957855.

Algamal:2019:PRE

[Alg19] Zakariya Yahya Algamal. Performance of ridge estimator in inverse Gaussian regression model. Communications in Statistics: Theory and Methods, 48(15):3836–3849, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1481977.

Aliyev:2017:SPH

[Ali17] Rovshan Aliyev. On a stochastic process with a heavytailed distributed component describing inventory model type of (s, S). Communications in Statistics: Theory and Methods, 46(5):2571–2579, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
1002932.

Alkhamisi:2010:REL

[Alk10] M. A. Alkhamisi. Ridge estimation in linear models with autocorrelated errors. Communications in Statistics: Theory and Methods, 39(14):2630–2644, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Almasri:2011:NAT

[Alm11] Abdullah Almasri. A new approach for testing periodicity. Communications in Statistics: Theory and Methods, 40(7):

1196–1217, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Almalki:2018:RNM

[Alm18] Saad J. Almalki. A reduced new modified Weibull distribution. Communications in Statistics: Theory and Methods, 47(10): 2297–2313, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2013.857416.

Attouch:2019:LLE

[ALR19] Mohammed Attouch, Ali Laksaci, and Fatima Rafaa. On the local linear estimate for functional regression: Uniform in bandwidth consistency. Communications in Statistics: Theory and Methods, 48(8):1836–1853, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1440308.

Alvo:2016:BGL

[Alv16] Mayer Alvo. Bridging the gap: a likelihood function approach for the analysis of ranking data. Communications in Statistics: Theory and Methods, 45(19):5835–5847, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.950751.

Arshad:2015:SEP

[AM15] Mohd. Arshad and Neeraj Misra. Selecting the exponential population having the larger guarantee time with unequal sample sizes. Communications in Statistics: Theory and Methods, 44(19):4144–4171, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albuquerque:2016:PIT

[AM16] Pedro Henrique Melo Albuquerque and Mariana Rosa Montenegro. PROMETHEE IV through kernel density estimation. Communications in Statistics: Theory and Methods, 45(18): 5355–5362, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.942432.

Afhami:2017:EBG

[AM17a]  Bahareh Afhami and Mohsen Madadi.  Entropy-based

goodness-of-fit tests for the Pareto I distribution. Communications in Statistics: Theory and Methods, 46(8):3649–3666, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1069350.

Aghamohammadi:2017:BQR

[AM17b] A. Aghamohammadi and M. R. Meshkani. Bayesian quantile regression for skew-normal linear mixed models. Communications in Statistics: Theory and Methods, 46(22):10953–10972, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1257713.

Al-Mutairi:2013:ISS

[AMGK13] D. K. Al-Mutairi, M. E. Ghitany, and Debasis Kundu. Inferences on stress-strength reliability from Lindley distributions. Communications in Statistics: Theory and Methods, 42(8):1443–1463, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Mutairi:2015:ISS

[AMGK15] D. K. Al-Mutairi, M. E. Ghitany, and Debasis Kundu. Inferences on stress-strength reliability from weighted Lindley distributions. Communications in Statistics: Theory and Methods, 44(19):4096–4113, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Awosoga:2014:SRT

[AMH14] Oluwagbohunmi A. Awosoga, Joseph W. Mckean, and Bradley E. Huitema. Simple robust tests for autocorrelated errors in time series design intervention models. Communications in Statistics: Theory and Methods, 43(13):2629–2641, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amin:2018:BID

[Ami18] Ayman A. Amin. Bayesian inference for double SARMA models. Communications in Statistics: Theory and Methods, 47(21):5333–5345, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1390132.

Ahmadi:2017:CPT

[AMP17] Jafar Ahmadi, Elham Mirfarah, and Ahmad Parsian. Comparison of preliminary test estimators based on generalized order statistics from proportional hazard family using Pitman measure of closeness. Communications in Statistics: Theory and Methods, 46(7):3200–3216, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1060336.

Arapis:2018:DSD

[AMP18] Anastasios N. Arapis, Frosso S. Makri, and Zaharias M. Psillakis. Distributions of statistics describing concentration of runs in non homogeneous Markov-dependent trials. Communications in Statistics: Theory and Methods, 47(9):2238–2250, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1337144.

Acosta-Mejia:2014:CRS

[AMR14] Cesar A. Acosta-Mejia and Luis Rincon. The continuous run sum chart. Communications in Statistics: Theory and Methods, 43(20):4371–4383, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Mosawi:2015:EPS

[AMV15] Riyadh R. Al-Mosawi and P. Vellaisamy. Estimation of the parameter of the selected binomial population. Communications in Statistics: Theory and Methods, 44(22):4796–4811, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

AlMasry:2018:GMM

[AMV18] Z. Al Masry, S. Mercier, and G. Verdier. Generalized method of moments for an extended gamma process. Communications in Statistics: Theory and Methods, 47(15):3687–3714, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361988.

Arab:2015:TVC

[AMZ15] Taher Ben Arab, Afif Masmoudi, and Mourad Zribi. Trace of the variance-covariance matrix in natural exponential families. Communications in Statistics: Theory and Methods, 44(6): 1241–1254, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amini:2017:EPD

[AN17a] Morteza Amini and Nader Nematollahi. Estimation of the parameter of a dynamically selected population for two subclasses of the exponential family. Communications in Statistics: Theory and Methods, 46(7):3217–3234, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1060337.

Arashi:2017:GED

[AN17b] M. Arashi and S. Nadarajah. Generalized elliptical distributions. Communications in Statistics: Theory and Methods, 46(13):6412–6432, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1129415.

Arashi:2017:SEM

[AN17c] M. Arashi and S. Nadarajah. On singular elliptical models. Communications in Statistics: Theory and Methods, 46(1):247– 258, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.990103.

Ashour:2017:IWD

[AN17d] S. K. Ashour and M. Nassar. Inference for Weibull distribution under adaptive Type-I progressive hybrid censored competing risks data. Communications in Statistics: Theory and Methods, 46(10):4756–4773, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1083111.

Arashi:2017:DLT

[ANA17] M. Arashi, Saralees Nadarajah, and Fikri Akdeniz. The distribution of the Liu-type estimator of the biasing parameter in elliptically contoured models. Communications in Statistics: Theory and Methods, 46(8):3829–3837, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1073315.

Anisimov:2011:SMC

[Ani11] Vladimir V. Anisimov. Statistical modeling of clinical trials (recruitment and randomization). Communications in Statistics: Theory and Methods, 40(19–20):3684–3699, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adibfar:2017:OAS

[ANJ17] Saeed Adibfar, Mohammad Saber Fallah Nezhad, and Roya Jafari. Optimal acceptance sampling policy considering Bayesian risks. Communications in Statistics: Theory and Methods, 46(11):5228–5237, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1099670.

Anonymous:2010:C

[Ano10a] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 39(16):3009, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2010:RR

[Ano10b] Anonymous. Referee recognition. Communications in Statistics: Theory and Methods, 39(8–9):1343, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2011:Cc

[Ano11a] Anonymous. Corrigenda. Communications in Statistics: Theory and Methods, 40(15):2820, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2011:Ca

[Ano11b] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 40(15):2818, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2011:Cb

[Ano11c] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 40(15):2819, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2011:Cd

[Ano11d] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 40(21):3935, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2011:EBE

[Ano11e] Anonymous. Editorial board EOV. Communications in Statistics: Theory and Methods, 40(24):ebi, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2011:E

[Ano11f] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 40(23):4324, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2012:Ca

[Ano12a] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 41(7):1313, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2012:Cb

[Ano12b] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 41(10):1885, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2012:Cc

[Ano12c] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 41(22):4222, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2012:EBE

[Ano12d] Anonymous. Editorial board EOV. Communications in Statistics: Theory and Methods, 41(24):ebi, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2012:E

[Ano12e] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 41(22):4223, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2013:C

[Ano13a] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 42(9):1712, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2013:CDF

[Ano13b] Anonymous. Corrigendum for: The distribution of family sizes under a time-homogeneous birth and death process. Communications in Statistics: Theory and Methods, 42(3):566–567, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [MS10d].

Anonymous:2013:Eb

[Ano13c] Anonymous. Errata. Communications in Statistics: Theory and Methods, 42(6):1141, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2013:Ea

[Ano13d] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 42(3):568, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2013:Ec

[Ano13e] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 42(18):3420, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2014:C

[Ano14] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 43(23):5114, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2015:C

[Ano15a] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 44(8):1762, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2015:Ea

[Ano15b] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 44(8):1760, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2015:Eb

[Ano15c] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 44(8):1761, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:2016:Ea

[Ano16a] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 45(10):3095, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1166871.

Anonymous:2016:Eb

[Ano16b] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 45(19):5871, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1184861.

Anonymous:2017:EEB

[Ano17a] Anonymous. EOV: Editorial Board. Communications in Statistics: Theory and Methods, 46(24):??, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1400512.

Anonymous:2017:E

[Ano17b] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 46(8):4150, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1259841.

Anonymous:2018:EEV

[Ano18] Anonymous. EOV: End of volume editorial board. Communications in Statistics: Theory and Methods, 47(24):??, 2018.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1475544.

Anonymous:2019:Cd

[Ano19a] Anonymous. Correction. Communications in Statistics: Theory and Methods, 48(23):5922, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2019.
1527071.

Anonymous:2019:C

[Ano19b] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 48(2):434, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1494939.

Anonymous:2019:Ca

[Ano19c] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 48(13):3421, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1505820.

Anonymous:2019:Cb

[Ano19d] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 48(15):3941–3942, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1481981.

Anonymous:2019:Cc

[Ano19e] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 48(20):5191, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1494907.

Anonymous:2019:E

[Ano19f] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 48(8):2074, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1444121.

Anonymous:2023:CAM

[Ano23] Anonymous. Correction to: Aljarrah, M.A., Famoye, F. and Lee, C. (2015). A New Weibull–Pareto distribution. Communications in Statistics — Theory and Methods, 44:19, 4077–4095. Communications in Statistics: Theory and Methods, 52(16):??, 2023. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2020.1822408. See [AFL15].

Asgharzadeh:2017:GIL

[ANS17] A. Asgharzadeh, S. Nadarajah, and F. Sharafi. Generalized inverse Lindley distribution with application to Danish fire insurance data. Communications in Statistics: Theory and Methods, 46(10):5001–5021, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1096394.

Alzaid:2012:EBD

[AO12] Abdulhamid A. Alzaid and Maha A. Omair. An extended binomial distribution with applications. Communications in Statistics: Theory and Methods, 41(19):3511–3527, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akahira:2017:SOA

[AO17a] M. Akahira and N. Ohyauchi. Second-order asymptotic loss of the MLE of a truncation parameter for a truncated exponential family of distributions. Communications in Statistics: Theory and Methods, 46(12):6085–6097, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1117638.

Arnab:2017:ORR

[AO17b] Raghunath Arnab and J. O. Olaomi. Optional randomized response technique in two-phase sampling. Communications in Statistics: Theory and Methods, 46(18):8973–8986, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1197257.

Adeoti:2018:CIB

[AO18a] Olatunde A. Adeoti and John O. Olaomi. Capability indexbased control chart for monitoring process mean using repetitive sampling. Communications in Statistics: Theory and Methods, 47(2):493–507, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1307401.

Ardah:2018:MSR

[AO18b] Husam I. Ardah and Evrim Oral. Model selection in randomized response techniques for binary responses. Communications in Statistics: Theory and Methods, 47(14):3305–3323, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1353626.

Ayodeji:2018:DCC

[AO18c] Idowu O. Ayodeji and Titilola O. Obilade. Development of a canonical correlation model involving non linearity and asymmetric variables. Communications in Statistics: Theory and Methods, 47(2):399–414, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1303731.

Aksoy:2017:PLD

[AOÖ17] Ümit Aksoy, Sofiya Ostrovska, and Ahmet Yasar Özban. Polynomial logistic distribution associated with a cubic polynomial. Communications in Statistics: Theory and Methods, 46(14): 7007–7019, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1140781.

Aoyagi:2010:BLC

[Aoy10] Miki Aoyagi. A Bayesian learning coefficient of generalization error and Vandermonde matrix-type singularities. Communications in Statistics: Theory and Methods, 39(15):2667–2687, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aminghafari:2012:NTS

[AP12] Mina Aminghafari and Jean-Michel Poggi. Nonstationary time series forecasting using wavelets and kernel smoothing. Communications in Statistics: Theory and Methods, 41(3):485–499, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akhundjanov:2017:EWM

[AP17] Sherzod B. Akhundjanov and Francis G. Pascual. Exponentially weighted moving average charts for correlated multivariate Poisson processes. Communications in Statistics: Theory and Methods, 46(10):4977–5000, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1096392.

Andrich:2019:MOA

[AP19] David Andrich and Pender Pedler. Modelling ordinal assessments: fit is not sufficient. Communications in Statistics: Theory and Methods, 48(12):2932–2947, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473595.

Amanullah:2014:IEM

[APA14] M. Amanullah, G. R. Pasha, and M. Aslam. Influence in the elliptical modified ridge regression. Communications in Statistics: Theory and Methods, 43(21):4650–4667, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ah-Pine:2010:USA

[APM10] Julien Ah-Pine and Jean-François Marcotorchino. Unifying some association criteria between partitions by using relational matrices. Communications in Statistics: Theory and Methods, 39(3):531–542, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alizadeh:2013:HMM

[AR13] S. H. Alizadeh and S. Rezakhah. Hidden Markov mixture autoregressive models: Stability and moments. Communications in Statistics: Theory and Methods, 42(6):1087–1104, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Azizzadeh:2014:CTD

[AR14] F. Azizzadeh and S. Rezakhah. The CUSUM test for detecting structural changes in strong mixing processes. Communications in Statistics: Theory and Methods, 43(17):3733–3750, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ajadi:2017:MME

[AR17a] Jimoh Olawale Ajadi and Muhammad Riaz. Mixed multivariate EWMA–CUSUM control charts for an improved process monitoring. Communications in Statistics: Theory and Methods, 46(14):6980–6993, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1139132.

Akdeniz:2017:EGD

[AR17b] Fikri Akdeniz and Mahdi Roozbeh. Efficiency of the generalized-difference-based weighted mixed almost unbiased two-parameter estimator in partially linear model. Communications in Statistics: Theory and Methods, 46(24):12259–12280, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1295075.

Amini:2017:LEC

[AR17c] Zahra Amini and Hossein Rabbani. Letter to the editor: Correction to “The Normal-Laplace distribution and its relatives”. Communications in Statistics: Theory and Methods, 46(4): 2076–2078, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1040510.

Alemohammad:2016:MSC

[ARA16] N. Alemohammad, S. Rezakhah, and S. H. Alizadeh. Markov switching component GARCH model: Stability and forecasting. Communications in Statistics: Theory and Methods, 45(15):4332–4348, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
841934.

Amiri:2019:SSG

[ARB19] Mehdi Amiri, Roohollah Roozegar, and Yousef Behbahani. Selected singular-generalized-hyperbolic distributions with applications to order statistics and reliability. Communications in Statistics: Theory and Methods, 48(12):3122–3135, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473608.

Abbas:2014:ETC

[ARD14] Nasir Abbas, Muhammad Riaz, and Ronald J. M. M. Does. An EWMA-type control chart for monitoring the process mean using auxiliary information. Communications in Statistics: Theory and Methods, 43(16):3485–3498, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arendacka:2014:IQF

[Are14] Barbora Arendacká. Independent quadratic forms in 3variance-component models. Communications in Statistics: Theory and Methods, 43(5):975–988, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arevalillo:2016:CFE

[Are16] Jorge M. Arevalillo. The Cornish–Fisher expansion for a class of statistics in first order autoregression. Communications in Statistics: Theory and Methods, 45(11):3196–3205, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901366.

Arfi:2013:NEH

[Arf13] Mounir Arfi. Nonparametric estimation for the hazard function. Communications in Statistics: Theory and Methods, 42(14): 2543–2550, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akbari:2019:NEQ

[ARJF19] Mahboubeh Akbari, Majid Rezaei, Sarah Jomhoori, and Vahid Fakoor. Nonparametric estimators for quantile density function under length-biased sampling. Communications in Statistics: Theory and Methods, 48(19):4918–4935, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1549245.

Arnab:2011:AER

[Arn11] Raghunath Arnab. Alternative estimators for randomized response techniques in multi-character surveys. Communications in Statistics: Theory and Methods, 40(10):1839–1848, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:2019:ORR

[Arn19] Raghunath Arnab. Optional randomized response techniques for quantitative characteristics. Communications in Statistics: Theory and Methods, 48(16):4154–4170, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1489554.

Arvanitis:2018:NSL

[Arv18] Stelios Arvanitis. A note on stable limit theory for the OLSE with non usual rates and the heteroskedasticity robust Wald test. Communications in Statistics: Theory and Methods, 47 (1):28–41, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1300277.

Arnab:2013:EMS

[AS13] Raghunath Arnab and Sarjinder Singh. Estimation of mean of sensitive characteristics for successive sampling. Communications in Statistics: Theory and Methods, 42(14):2499–2524, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:2015:RRT

[AS15] Raghunath Arnab and D. K. Shangodoyin. Randomized response techniques using maximum likelihood estimator. Communications in Statistics: Theory and Methods, 44(16):3340– 3349, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Sharadqah:2017:NPF

[AS17] Ali Al-Sharadqah. A new perspective in functional EIV linear model: Part I. Communications in Statistics: Theory and Methods, 46(14):7039–7062, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1143009.

Ai:2018:KLE

[AS18] Xiaohui Ai and Yang Sun. Karhunen-Loeve expansion for the additive two-sided Brownian motion. Communications in Statistics: Theory and Methods, 47(13):3085–3091, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1346809.

Acharya:2019:APM

[AS19] Sarat Kumar Acharya and Saroja Kumar Singh. Asymptotic properties of maximum likelihood estimators from single server queues: a martingale approach. Communications in Statistics: Theory and Methods, 48(14):3549–3557, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1477958.

Sama-ae:2018:HNA

[aSaCN18] Al ameen Sama-ae, Nattakarn Chaidee, and Kritsana Neammanee. Half-normal approximation for statistics of symmetric simple random walk. Communications in Statistics: Theory and Methods, 47(4):779–792, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1139129.

Alqallaf:2010:CAS

[ASAK10] F. Alqallaf, A. R. Soltani, and N. Al-Kandari. Continuity and analysis of sequences of principal components. Communications in Statistics: Theory and Methods, 39(19):3558–3567, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ardalan:2012:TPN

[ASAN12] A. Ardalan, S. M. Sadooghi-Alvandi, and A. R. Nematollahi. The two-piece normal-Laplace distribution. Communications in Statistics: Theory and Methods, 41(20):3759–3785, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amini:2013:AND

[ASB13] M. Amini, H. R. Nili Sani, and A. Bozorgnia. Aspects of negative dependence structures. Communications in Statistics: Theory and Methods, 42(5):907–917, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aslam:2017:DSP

[ASeMJ17] Muhammad Aslam, Saqlain Latif Satti, Mitwali Abd el Moemen, and Chi-Hyuck Jun. Design of sampling plan using auxiliary information. Communications in Statistics: Theory and Methods, 46(8):3772–3781, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1073311.

Ahmed:2017:USR

[ASG17] Shakeel Ahmed, Javid Shabbir, and Sat Gupta. Use of scrambled response model in estimating the finite population mean in presence of non response when coefficient of variation is known. Communications in Statistics: Theory and Methods, 46(17): 8435–8449, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1183782.

Amiri:2014:DAM

[ASMZ14] Amirhossein Amiri, Abbas Saghaei, Mohammad Mohseni, and Yaser Zerehsaz. Diagnosis aids in multivariate multiple linear regression profiles monitoring. Communications in Statistics: Theory and Methods, 43(14):3057–3079, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:2012:UTD

[ASN12] Raghunath Arnab, Sarjinder Singh, and D. North. Use of two decks of cards in randomized response techniques for complex survey designs. Communications in Statistics: Theory and Methods, 41(16–17):3198–3210, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2012.682634.

Arashi:2015:BSI

[ASNT15] M. Arashi, A. K. Md. Ehsanes Saleh, Daya K. Nagar, and S. M. M. Tabatabaey. Bayesian statistical inference for Laplacian class of matrix variate elliptically contoured models. Communications in Statistics: Theory and Methods, 44(13):2774– 2787, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:2019:MCB

[ASR^{+}19] Rashid Ahmed, Farrukh Shehzad, Muhammad Rajab, Muhammad Daniyal, and M. H. Tahir. Minimal circular balanced repeated measurement designs in unequal period sizes. Communications in Statistics: Theory and Methods, 48(21):5223–5232, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1508719.

Ahmed:2018:SEM

[ASS18] Segun Ahmed, Stephen A. Sedory, and Sarjinder Singh. Simultaneous estimation of means of two sensitive variables. Communications in Statistics: Theory and Methods, 47(2):324–343, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1303733.

Artes:2009:AMT

[AT09] Rinaldo Artes and Clelia M. C. Toloi. An autoregressive model for time series of circular data. Communications in Statistics: Theory and Methods, 39(1):186–194, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adam:2011:MPD

[AT11] M. B. Adam and J. A. Tawn. Modification of Pickands’ dependence function for ordered bivariate extreme distribution. Communications in Statistics: Theory and Methods, 40(9):1687– 1700, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abbas:2015:AFD

[AT15] Kamran Abbas and Yincai Tang. Analysis of Frechet distribution using reference priors. Communications in Statistics: Theory and Methods, 44(14):2945–2956, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alavi:2017:SBN

[AT17] Sayed Mohammad Reza Alavi and Mahtab Tarhani. On a Skew Bimodal Normal–Normal distribution fitted to the Old-Faithful geyser data. Communications in Statistics: Theory and Methods, 46(15):7301–7312, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1148731.

Adjoudj:2019:CQE

[AT19] Latifa Adjoudj and Abdelkader Tatachak. Conditional quantile estimation for truncated and associated data. Communications in Statistics: Theory and Methods, 48(18):4598–4641, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1498895.

Atkinson:2012:OED

[Atk12] Anthony C. Atkinson. Optimum experimental designs for choosing between competitive and non competitive models of enzyme inhibition. Communications in Statistics: Theory and Methods, 41(13–14):2283–2296, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Altun:2019:VRE

[ATO19] Emrah Altun, Huseyin Tatlidil, and Gamze Ozel. Value-at-risk estimation with new skew extension of generalized normal distribution. Communications in Statistics: Theory and Methods, 48(14):3663–3681, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481970.

Ahmad:2015:TCM

[AV15a] M. Rauf Ahmad and Dietrich Von Rosen. Tests of covariance matrices for high dimensional multivariate data under non normality. Communications in Statistics: Theory and Methods, 44(7):1387–1398, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:2015:OS

[AV15b] Barry C. Arnold and Jose A. Villasenor. Orthogonal spacings. Communications in Statistics: Theory and Methods, 44(19):

3998–4006, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahkim:2018:TCV

[AV18] Mohamed Ahkim and Anneleen Verhasselt. Testing for constancy in varying coefficient models. Communications in Statistics: Theory and Methods, 47(4):890–911, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1300271.

Arellano-Valle:2010:EES

[AVCG10] Reinaldo B. Arellano-Valle, Milton A. Cortés, and Héctor W. Gómez. An extension of the epsilon-skew-normal distribution. Communications in Statistics: Theory and Methods, 39(5):912– 922, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arellano-Valle:2013:CPD

[AVCL13] Reinaldo B. Arellano-Valle, Luis M. Castro, and Rosangela H. Loschi. Change point detection in the skew-normal model parameters. Communications in Statistics: Theory and Methods, 42(4):603–618, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

AlLabadi:2010:MWS

[AW10] L. Al Labadi and Z. Wang. Modified Wynn’s sequential algorithm for constructing D-optimal designs: Adding two points at a time. Communications in Statistics: Theory and Methods, 39(15):2818–2828, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See remarks [Mül12].

Arumairajan:2016:GPT

[AW16] S. Arumairajan and P. Wijekoon. Generalized preliminary test stochastic restricted estimator in the linear regression model. Communications in Statistics: Theory and Methods, 45(20): 6061–6086, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.957850.

Aijun:2017:SBV

[AXLJ17] Yang Aijun, Jiang Xuejun, Xiang Liming, and Lin Jinguan. Sparse Bayesian variable selection in multinomial probit regression model with application to high-dimensional data classification. Communications in Statistics: Theory and Methods, 46(12):6137–6150, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1122056.

Aslam:2012:TSV

[AYC^{+}12] Muhammad Aslam, Ching-Ho Yen, Chia-Hao Chang, Chi-Hyuck Jun, Munir Ahmad, and Mujahid Rasool. Two-stage variables acceptance sampling plans using process loss functions. Communications in Statistics: Theory and Methods, 41 (20):3633–3647, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aykroyd:2019:BMT

[Ayk19] Robert G. Aykroyd. Bayesian modeling of temperature-related mortality with latent functional relationships. Communications in Statistics: Theory and Methods, 48(1):3–14, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1421223.

Yan:2017:RTR

[aYwSlDmS17] Wei an Yan, Bao wei Song, Gui lin Duan, and Yi min Shi. Real-time reliability evaluation of two-phase Wiener degradation process. Communications in Statistics: Theory and Methods, 46(1):176–188, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
988262.

Arnab:2015:VEC

[AZN15] Raghunath Arnab, T. Zewotir, and D. North. Variance estimation from complex survey designs: a case study of household income and expenditure survey design 2002/03, Botswana. Communications in Statistics: Theory and Methods, 44(1):63–79, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:2010:BPS

[BA10] Ashok K. Bansal and Priyanka Aggarwal. Bayes prediction for a stratified regression superpopulation model using balanced loss function. Communications in Statistics: Theory and Methods, 39(15):2789–2799, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bayramoglu:2014:CIR

[(Ba14] Ismihan Bayramoglu (Bairamov). On conditionally independent random variables, copula and order statistics. Communications in Statistics: Theory and Methods, 43(10–12):2105– 2117, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balamurali:2018:ODS

[BAA^{+}18] Saminathan Balamurali, Liaquat Ahmad, Muhammad Aslam, Jaffer Hussain, and Chi-Hyuck Jun. Optimal designing of an SkSP-r double sampling plan. Communications in Statistics: Theory and Methods, 47(17):4329–4337, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1373819.

Belaghi:2018:DFC

[BAB^{+}18] R. Arabi Belaghi, M. Noori Asl, H. Bevrani, William Volterman, and N. Balakrishnan. On the distribution-free confidence intervals and universal bounds for quantiles based on joint records. Communications in Statistics: Theory and Methods, 47(12):2969–2978, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1343845.

Bekker:2019:NBG

[BAF19] A. Bekker, M. Arashi, and J. T. Ferreira. New bivariate gamma types with MIMO application. Communications in Statistics: Theory and Methods, 48(3):596–615, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1417428.

Baklizi:2011:ELI

[Bak11] Ayman Baklizi. Empirical likelihood inference for population quantiles with unbalanced ranked set samples. Communications in Statistics: Theory and Methods, 40(23):4179–4188, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Birkner:2013:EFH

[BAK13] Thomas Birkner, Inmaculada B. Aban, and Charles R. Katholi. Evaluation of a frequentist hierarchical model to estimate prevalence when sampling from a large geographic area using pool screening. Communications in Statistics: Theory and Methods, 42(19):3571–3595, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bak:2017:EHT

[Bak17] Tomasz Bak. An extension of Horvitz–Thompson estimator used in adaptive cluster sampling to continuous universe. Communications in Statistics: Theory and Methods, 46(19):9777– 9786, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1218028.

Belaid:2018:BAB

[BAKZ18] Nawal Belaid, Smail Adjabi, Célestin C. Kokonendji, and Nabil Zougab. Bayesian adaptive bandwidth selector for multivariate discrete kernel estimator. Communications in Statistics: Theory and Methods, 47(12):2988–3001, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1346807.

Balanzario:2015:BLM

[Bal15] Eugenio P. Balanzario. Benford’s law for mixtures. Communications in Statistics: Theory and Methods, 44(4):698–709, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balamurali:2018:ODN

[Bal18a] S. Balamurali. Optimal designing of a new mixed variable lotsize chain sampling plan based on the process capability index. Communications in Statistics: Theory and Methods, 47(10): 2490–2503, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1339089.

Balamurali:2018:DNM

[BAL18b] S. Balamurali, Muhammad Aslam, and Ahmad Liaquat. Determination of a new mixed variable lot-size multiple dependent state sampling plan based on the process capability index. Communications in Statistics: Theory and Methods, 47 (3):615–627, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1309435.

Banerjee:2013:SDL

[Ban13] Anurag N. Banerjee. The sensitivity of detrended long-memory processes. Communications in Statistics: Theory and Methods, 42(20):3770–3780, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barros:2017:DIO

[BANC17] C. M. Barros, G. J. A. Amaral, A. D. C. Nascimento, and A. H. M. A. Cysneiros. Detecting influential observations in Watson data. Communications in Statistics: Theory and Methods, 46(14):6882–6898, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1139130.

Baratpour:2010:CBC

[Bar10] S. Baratpour. Characterizations based on cumulative residual entropy of first-order statistics. Communications in Statistics: Theory and Methods, 39(20):3645–3651, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basiri:2016:CAN

[BAR16] Elham Basiri, Jafar Ahmadi, and Mohammad Z. Raqab. Comparison among non parametric prediction intervals of order statistics. Communications in Statistics: Theory and Methods, 45(9):2699–2713, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
887117.

Barrett:2017:NEC

[Bar17a] Bruce Barrett. A note on exact calculation of the non central hypergeometric distribution. Communications in Statistics: Theory and Methods, 46(13):6737–6741, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1134573.

Barrett:2017:SEI

[Bar17b] Bruce Barrett. Simple, exact inference for 2 × 2 tables. Communications in Statistics: Theory and Methods, 46(1):221–233, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.990101.

Bartlett:2017:NIC

[Bar17c] Thomas E. Bartlett. Network inference and community detection, based on covariance matrices, correlations, and test statistics from arbitrary distributions. Communications in Statistics: Theory and Methods, 46(18):9150–9165, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1205624.

Barbiero:2019:BGD

[Bar19] Alessandro Barbiero. A bivariate geometric distribution allowing for positive or negative correlation. Communications in Statistics: Theory and Methods, 48(11):2842–2861, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473428.

Basak:2014:PTF

[Bas14] Indrani Basak. Prediction of times to failure of censored items for a simple step-stress model with regular and progressive Type I censoring from the exponential distribution. Communications in Statistics: Theory and Methods, 43(10–12):2322– 2341, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bashtian:2011:REU

[BAT11] M. Hassanzadeh Bashtian, M. Arashi, and S. M. M. Tabatabaey. Ridge estimation under the stochastic restriction. Communications in Statistics: Theory and Methods, 40 (21):3711–3747, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Belaghi:2015:CPT

[BAT15] R. Arabi Belaghi, M. Arashi, and S. M. M. Tabatabaey. On the construction of preliminary test estimator based on record values for the Burr XII model. Communications in Statistics: Theory and Methods, 44(1):1–23, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bashir:2018:SEC

[BAT^{+}18] Zahid Bashir, Rashid Ahmed, M. H. Tahir, Syed Shakir Ali Ghazali, and Farrukh Shehzad. Some extensions of circular balanced and circular strongly balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 47(9):2183–2194, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1337140.

Butt:2019:SBP

[BAW19] Kalsoom Afzal Butt, Muhammad Aslam, and Fu-Kwun Wang. Selecting better process based on difference statistic using double sampling plan. Communications in Statistics: Theory and Methods, 48(11):2641–2656, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472785.

Bayoud:2016:AME

[Bay16] Husam Awni Bayoud. Admissible minimax estimators for the shape parameter of Topp–Leone distribution. Communications in Statistics: Theory and Methods, 45(1):71–82, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blanke:2012:CAP

[BB12] Delphine Blanke and Denis Bosq. Comparing ARMA processes with roots of modulus 1 and polynomial regression. Communications in Statistics: Theory and Methods, 41(16–17):2851– 2863, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/03610926.2011.617482.

Bashiri:2014:OPI

[BB14a] Mahdi Bashiri and Mohammad Hasan Bakhtiarifar. An optimality probability index for correlated multiple responses. Communications in Statistics: Theory and Methods, 43(20): 4324–4336, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bayramoglu:2014:BLH

[BB14b] K. Bayramoglu and I. Bayramoglu (Bairamov). Baker–Lin– Huang type bivariate distributions based on order statistics. Communications in Statistics: Theory and Methods, 43(10–12): 1992–2006, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bentarzi:2017:PIV

[BB17a] Mohamed Bentarzi and Wissam Bentarzi. Periodic integervalued bilinear time series model. Communications in Statistics: Theory and Methods, 46(3):1184–1201, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1014107.

Boualam:2017:HEU

[BB17b] Karima Boualam and Youcef Berkoun. Hill’s estimator under weak dependence. Communications in Statistics: Theory and Methods, 46(18):9218–9229, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1205615.

BenElouefi:2019:GFT

[BB19a] Rim Ben Elouefi and Rim Ben Elouefi. A goodness-of-fit test for the stratified proportional hazards model for survival data. Communications in Statistics: Theory and Methods, 48(17): 4206–4220, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1489965.

Bhati:2019:NID

[BB19b] Deepesh Bhati and Hassan S. Bakouch. A new infinitely divisible discrete distribution with applications to count data modeling. Communications in Statistics: Theory and Methods, 48(6): 1401–1416, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1433847.

Basirat:2016:ESS

[BBA16] M. Basirat, S. Baratpour, and Jafar Ahmadi. On estimation of stress-strength parameter using record values from proportional hazard rate models. Communications in Statistics: Theory and Methods, 45(19):5787–5801, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.948727.

Bazan:2010:FSP

[BBB10] Jorge L. Bazán, Heleno Bolfarine, and Márcia D. Branco. A framework for skew-probit links in binary regression. Communications in Statistics: Theory and Methods, 39(4):678–697, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bertoli-Barsotti:2014:IGS

[BBB14] Lucio Bertoli-Barsotti and Silvia Bacci. Identifying Guttman structures in incomplete Rasch datasets. Communications in Statistics: Theory and Methods, 43(3):470–497, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barroso:2013:SOC

[BBC13] Lúcia P. Barroso, Denise A. Botter, and Gauss M. Cordeiro. Second-order covariance matrix formula for heteroskedastic generalized linear models. Communications in Statistics: Theory and Methods, 42(9):1618–1627, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baragona:2018:PTB

[BBC18] Roberto Baragona, Francesco Battaglia, and Domenico Cucina. Portmanteau tests based on quadratic forms in the autocorrelations. Communications in Statistics: Theory and Methods, 47(17):4355–4374, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1380829.

Bacci:2014:CML

[BBG14] Silvia Bacci, Francesco Bartolucci, and Michela Gnaldi. A class of multidimensional latent class IRT models for ordinal polytomous item responses. Communications in Statistics: Theory and Methods, 43(4):787–800, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Behdani:2018:RBW

[BBG18] Zahra Behdani, Gholam Reza Mohtashami Borzadaran, and Bahram Sadeghpour Gildeh. Relationship between the weighted distributions and some inequality measures. Communications in Statistics: Theory and Methods, 47(22):5573–5589, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1397171.

Balakrishnan:2018:NSC

[BBHN18] Narayanaswamy Balakrishnan, Ghobad Barmalzan, Abedin Haidari, and Amir T. Payandeh Najafabadi. Necessary and sufficient conditions for stochastic orders between (n−r+1)-out-of-n systems in proportional hazard (reversed hazard) rates model. Communications in Statistics: Theory and Methods, 47(23): 5854–5866, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1406517.

Bertoli-Barsotti:2014:RAS

[BBP14] Lucio Bertoli-Barsotti and Antonio Punzo. Refusal to answer specific questions in a survey: a case study. Communications in Statistics: Theory and Methods, 43(4):826–838, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bianco:2017:CTE

[BBR17] Ana M. Bianco, Graciela Boente, and Isabel M. Rodrigues. Conditional tests for elliptical symmetry using robust estimators. Communications in Statistics: Theory and Methods, 46(4):1744–1765, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1026997.

Bodnar:2016:MAE

[BBS16] Rostyslav Bodnar, Taras Bodnar, and Wolfgang Schmid. Multivariate autoregressive extreme value process and its application for modeling the time series properties of the extreme daily asset prices. Communications in Statistics: Theory and Methods, 45(12):3421–3440, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
791370.

Bidram:2013:NGE

[BBT13] Hamid Bidram, Javad Behboodian, and Mina Towhidi. A new generalized exponential geometric distribution. Communications in Statistics: Theory and Methods, 42(3):528–542, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bao:2012:SRT

[BC12a] Xinzhuo Bao and Lirong Cui. A study on reliability for a twoitem cold standby Markov repairable system with neglected failures. Communications in Statistics: Theory and Methods, 41(21):3988–3999, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bashir:2012:RML

[BC12b] Shaheena Bashir and E. M. Carter. Robust mixture of linear regression models. Communications in Statistics: Theory and Methods, 41(18):3371–3388, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berrendero:2012:TSO

[BC12c] José R. Berrendero and Javier Cárcamo. Tests for stochastic orders and mean order statistics. Communications in Statistics: Theory and Methods, 41(8):1497–1509, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bianconcini:2012:MLG

[BC12d] S. Bianconcini and S. Cagnone. Multivariate latent growth models for mixed data with covariate effects. Communications in Statistics: Theory and Methods, 41(16–17):3079–3093, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/03610926.2011.609955.

Bazyari:2013:UBV

[BC13] Abouzar Bazyari and Rahim Chinipardaz. Upper bound for p-value of the test of multivariate normal ordered mean vectors against all alternatives. Communications in Statistics: Theory and Methods, 42(10):1748–1758, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bianconcini:2014:RPD

[BC14] Silvia Bianconcini and Silvia Cagnone. The role of posterior densities in latent variable models for ordinal data. Communications in Statistics: Theory and Methods, 43(4):681–692, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blomstedt:2015:PPC

[BC15] Paul Blomstedt and Jukka Corander. Posterior predictive comparisons for the two-sample problem. Communications in Statistics: Theory and Methods, 44(2):376–389, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brusco:2019:IDA

[BCB19] Michael J. Brusco, J. Dennis Cradit, and Susan Brudvig. An integrated dominance analysis and dynamic programing approach for measuring predictor importance for customer satisfaction. Communications in Statistics: Theory and Methods, 48(21):5290–5307, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1510004.

Bielecki:2014:BDM

[BCCH14] Tomasz R. Bielecki, Areski Cousin, Stéphane Crépey, and Alexander Herbertsson. A bottom-up dynamic model of portfolio credit risk with stochastic intensities and random recoveries. Communications in Statistics: Theory and Methods, 43(7): 1362–1389, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bolivar-Cime:2018:BDM

[BCCR18] A. Bolivar-Cime and L. M. Cordova-Rodriguez. Binary discrimination methods for high-dimensional data with a geometric representation. Communications in Statistics: Theory and Methods, 47(11):2720–2740, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1342838.

Brombin:2012:MCC

[BCD12] Chiara Brombin, Massimo Crippa, and Clelia Di Serio. Modeling cancer cells growth. Communications in Statistics: Theory and Methods, 41(16–17):3043–3059, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2012.685547.

Broniatowski:2012:UBE

[BCDLA12] Michel Broniatowski, Giorgio Celant, Marco Di Battista, and Samuela Leoni-Aubin. Upper bounds for the error in some interpolation and extrapolation designs. Communications in Statistics: Theory and Methods, 41(16–17):3002–3019, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2011.607532.

Baur:2011:SCE

[BCG11] Benedict Baur, Florian Conrad, and Martin Grothaus. Smooth contractive embeddings and application to Feynman formula for parabolic equations on smooth bounded domains. Communications in Statistics: Theory and Methods, 40(19–20):3452– 3464, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baek:2012:BCP

[BCH12] J. S. Baek, M. S. Choi, and S. Y. Hwang. A broad class of partially specified autoregressions on multi-casting data. Communications in Statistics: Theory and Methods, 41(1):178–193, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berenhaut:2015:AED

[BCH15] Kenneth S. Berenhaut, James W. Chernesky, Jr., and Ross P. Hilton. Asymptotic expansions for i.i.d. sums via lower-order convolutions. Communications in Statistics: Theory and Methods, 44(11):2330–2350, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bulla:2015:BSD

[BCK15] Jan Bulla, Christophe Chesneau, and Maher Kachour. On the bivariate Skellam distribution. Communications in Statistics: Theory and Methods, 44(21):4552–4567, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bulla:2017:BFO

[BCK17] Jan Bulla, Christophe Chesneau, and Maher Kachour. A bivariate first-order signed integer-valued autoregressive process. Communications in Statistics: Theory and Methods, 46(13): 6590–6604, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1132322.

Bouzebda:2016:SUC

[BCL16] Salim Bouzebda, Khalid Chokri, and Djamal Louani. Some uniform consistency results in the partially linear additive model components estimation. Communications in Statistics: Theory and Methods, 45(5):1278–1310, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.861491.

Buzaianu:2017:SNP

[BCP17] Elena M. Buzaianu, Pinyuen Chen, and S. Panchapakesan. Selecting the normal population with the smallest variance: a restricted subset selection rule. Communications in Statistics: Theory and Methods, 46(16):7887–7901, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1165849.

Balakrishnan:2010:BCD

[BCS10] N. Balakrishnan, Enrique Castillo, and José María Sarabia. Bivariate continuous distributions with specified conditional hazard functions. Communications in Statistics: Theory and Methods, 39(14):2473–2484, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bandyopadhyay:2011:MTA

[BD11a] Uttam Bandyopadhyay and Saurav De. On multi-treatment adaptive allocation design for dichotomous response. Communications in Statistics: Theory and Methods, 40(22):4104–4124, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berkoun:2011:TIL

[BD11b] Youcef Berkoun and Hamid Douki. Testing independence in linear process with non-normal innovations. Communications in Statistics: Theory and Methods, 40(13):2377–2384, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ba:2016:MDE

[BD16a] Demba Bocar Ba and Ali Souleymane Dabye. On minimum L^{p}-distance estimation for inhomogeneous Poisson processes. Communications in Statistics: Theory and Methods, 45(21): 6461–6470, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.927503.

Barriga:2016:BID

[BD16b] Gladys D. Cacsire Barriga and Dipak K. Dey. Bayesian inference and diagnostics in zero-inflated generalized power series regression model. Communications in Statistics: Theory and Methods, 45(22):6553–6568, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
919397.

Bouzebda:2017:ARM

[BD17a] Salim Bouzebda and Sultana Didi. Additive regression model for stationary and ergodic continuous time processes. Communications in Statistics: Theory and Methods, 46(5):2454–2493, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1048882.

Bouzebda:2017:MWD

[BD17b] Salim Bouzebda and Sultana Didi. Multivariate wavelet density and regression estimators for stationary and ergodic discrete time processes: Asymptotic results. Communications in Statistics: Theory and Methods, 46(3):1367–1406, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019144.

Baratnia:2019:SOS

[BD19] Mohammad Baratnia and Mahdi Doostparast. Sequential order statistics from dependent random variables. Communications in Statistics: Theory and Methods, 48(18):4569–4580, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1496257.

Balakrishnan:2012:PCC

[BDKM12] N. Balakrishnan, Katherine F. Davies, Jerome P. Keating, and Robert L. Mason. Pitman closeness comparison of best linear unbiased and invariant predictors for exponential distribution in one- and two-sample situations. Communications in Statistics: Theory and Methods, 41(1):1–15, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Biswas:2013:DRA

[BDM13] Amrita Biswas, Premadhis Das, and Nripes Kumar Mandal. Designs robust against the presence of an aberration in a diallel cross design setup. Communications in Statistics: Theory and Methods, 42(15):2810–2817, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Biswas:2015:DRA

[BDM15] A. Biswas, P. Das, and N. K. Mandal. Designs robust against violation of normality assumption in the standard F -test. Communications in Statistics: Theory and Methods, 44(14):2898– 2910, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beh:2011:CAC

[BDS11] Eric J. Beh, Luigi D’ambra, and Biagio Simonetti. Correspondence analysis of cumulative frequencies using a decomposition of Taguchi’s statistic. Communications in Statistics: Theory and Methods, 40(9):1620–1632, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beauducel:2013:NUV

[Bea13] André Beauducel. A note on unwanted variance in exploratory factor models. Communications in Statistics: Theory and Methods, 42(3):561–565, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beauducel:2015:NEI

[Bea15] André Beauducel. A note on equality and inequality of some factor score predictors. Communications in Statistics: Theory and Methods, 44(17):3653–3657, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bouzebda:2013:NEE

[BEKL13] Salim Bouzebda, Issam Elhattab, Amor Keziou, and Tewfik Lounis. New entropy estimator with an application to test of normality. Communications in Statistics: Theory and Methods, 42(12):2245–2270, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beltrami:2017:WLE

[Bel17] Joleen Beltrami. Weibull lagged effect step-stress model with competing risks. Communications in Statistics: Theory and Methods, 46(11):5419–5442, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1102283.

Benavides:2014:GMR

[Ben14] Efrén M. Benavides. General model for reliability-based engineering design. Communications in Statistics: Theory and Methods, 43(10–12):2342–2356, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Benhaddou:2019:BDM

[Ben19] Rida Benhaddou. Blind deconvolution model in periodic setting with fractional Gaussian noise. Communications in Statistics: Theory and Methods, 48(3):438–449, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1417431.

Berni:2012:QRT

[Ber12] Rossella Berni. Quality and reliability in top-event estimation: Quantitative fault tree analysis in case of dependent events. Communications in Statistics: Theory and Methods, 41(16–17): 3138–3149, 2012. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/03610926.2011.621574.

Bean:2016:FIP

[BER16a] Nigel G. Bean, Ali Eshragh, and Joshua V. Ross. Fisher information for a partially observable simple birth process. Communications in Statistics: Theory and Methods, 45(24):7161–7183, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.978024.

Beran:2016:ELR

[Ber16b] Jan Beran. On the effect of long-range dependence on extreme value copula estimation with fixed marginals. Communications in Statistics: Theory and Methods, 45(19):5590–5618, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.948198.

Bernardoff:2018:MOL

[Ber18] Philippe Bernardoff. Marshall-Olkin Laplace transform copulas of multivariate gamma distributions. Communications in Statistics: Theory and Methods, 47(3):655–670, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1310241.

Barakat:2018:GOS

[BEY18] H. M. Barakat, M. A. Abd Elgawad, and Ting Yan. Generalized-order statistics with random indices. Communications in Statistics: Theory and Methods, 47(12):2859–2868, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1342837.

Bouzebda:2011:MTS

[BEZ11] Salim Bouzebda, Nour-Eddin El Faouzi, and Tarek Zari. On the multivariate two-sample problem using strong approximations of empirical copula processes. Communications in Statistics: Theory and Methods, 40(8):1490–1509, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barhdadi:2010:ETA

[BF10] Amina Barhdadi and Sorana Froda. An exact test for additivity in two-way tables under biadditive modelling. Communications in Statistics: Theory and Methods, 39(11):1960–1978, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basso:2012:EMP

[BF12] Dario Basso and Livio Finos. Exact multivariate permutation tests for fixed effects in mixed-models. Communications in Statistics: Theory and Methods, 41(16–17):2991–3001, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/03610926.2011.627103.

Baghfalaki:2011:EEA

[BG11a] T. Baghfalaki and M. Ganjali. An EM estimation approach for analyzing bivariate skew normal data with non monotone missing values. Communications in Statistics: Theory and Methods, 40(9):1671–1686, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bargagliotti:2011:SSR

[BG11b] Anna E. Bargagliotti and Raymond N. Greenwell. Statistical significance of ranking paradoxes. Communications in Statistics: Theory and Methods, 40(5):916–928, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boucher:2011:SNA

[BG11c] Jean-Philippe Boucher and Montserrat Guillen. A seminonparametric approach to model panel count data. Communications in Statistics: Theory and Methods, 40(4):622–634, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Braekers:2011:FMK

[BG11d] Roel Braekers and Auguste Gaddah. Flexible modeling in the Koziol–Green model by a copula function. Communications in Statistics: Theory and Methods, 40(7):1218–1235, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Braekers:2016:SPC

[BG16a] Roel Braekers and Yves Grouwels. A semi-parametric Cox’s regression model for zero-inflated left-censored time to event data. Communications in Statistics: Theory and Methods, 45(7):1969–1988, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
870207.

Bulinskaya:2016:ABP

[BG16b] Ekaterina Bulinskaya and Alexander Gromov. Asymptotic behavior of the processes describing some insurance models. Communications in Statistics: Theory and Methods, 45(6):1778– 1793, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.985840.

Bowman:2017:LEE

[BG17] Dale Bowman and E. Olusegun George. Likelihood estimation for exchangeable multinomial data. Communications in Statistics: Theory and Methods, 46(6):2882–2892, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1053934.

Bibi:2018:MSB

[BG18] Abdelouahab Bibi and Ahmed Ghezal. Markov-switching BILINEAR–GARCH models: Structure and estimation. Communications in Statistics: Theory and Methods, 47(2):307–323, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1303732.

Baringhaus:2019:HTS

[BG19a] Ludwig Baringhaus and Daniel Gaigall. On Hotelling’s T^{2} test in a special paired sample case. Communications in Statistics: Theory and Methods, 48(2):257–267, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1408828.

Bhattacharjee:2019:BCE

[BG19b] Abhishek Bhattacharjee and Malay Ghosh. Bias corrected empirical Bayes confidence intervals for the selected mean. Communications in Statistics: Theory and Methods, 48(3):583–595, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1417427.

Bibi:2019:QPT

[BG19c] Abdelouahab Bibi and Ahmed Ghezal. QMLE of periodic time-varying bilinear-GARCH models. Communications in Statistics: Theory and Methods, 48(13):3291–3310, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1476703.

Beirlant:2011:EBM

[BGG11] Jan Beirlant, Ségolen Geffray, and Armelle Guillou. Estimation of the bias of the maximum likelihood estimators in an extreme value context. Communications in Statistics: Theory and Methods, 40(22):3959–3971, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beirlant:2010:PTM

[BGT10] Jan Beirlant, Armelle Guillou, and Gwladys Toulemonde. Peaks-over-threshold modeling under random censoring. Communications in Statistics: Theory and Methods, 39(7):1158– 1179, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishna:2012:SAM

[BH12] N. Balakrishna and G. Hareesh. Stable autoregressive models and signal estimation. Communications in Statistics: Theory and Methods, 41(11):1969–1988, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beauducel:2017:DRF

[BH17a] André Beauducel and Norbert Hilger. The determinacy of the regression factor score predictor based on continuous parameter estimates from categorical variables. Communications in Statistics: Theory and Methods, 46(7):3417–3425, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1062106.

Bender:2017:ICF

[BH17b] Shaun Bender and Daniel F. Heitjan. Ignorability conditions for frequentist non parametric analysis of conditional distributions with incomplete data. Communications in Statistics: Theory and Methods, 46(11):5252–5264, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1099673.

Bou-Hamad:2017:BCR

[BH17c] Imad Bou-Hamad. Bayesian credit ratings: a random forest alternative approach. Communications in Statistics: Theory and Methods, 46(15):7289–7300, 2017. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1148730.

Balakrishnan:2015:IOR

[BHB15] N. Balakrishnan, Abedin Haidari, and Ghobad Barmalzan. Improved ordering results for fail-safe systems with exponential components. Communications in Statistics: Theory and Methods, 44(10):2010–2023, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:2013:FIC

[BHKHM13] Nicholas Brown, Rachael Hamilton-Keene, Philippa Hartney, and Terence Mills. Forecasting the incidence of cancer in regional Victoria, Australia. Communications in Statistics: Theory and Methods, 42(16):2935–2943, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2015:CIQ

[BHLK15] N. Balakrishnan, A. J. Hayter, W. Liu, and S. Kiatsupaibul. Confidence intervals for quantiles of a two-parameter exponential distribution under progressive Type-II censoring. Communications in Statistics: Theory and Methods, 44(14):3001–3010, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bi:2015:NFV

[Bi15] Henry H. Bi. A note on the factor values of three common Shewhart variables control charts. Communications in Statistics: Theory and Methods, 44(13):2655–2673, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bialek:2012:PGF

[Bia12] Jacek Bialek. Proposition of a general formula for price indices. Communications in Statistics: Theory and Methods, 41(5):943– 952, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bialek:2015:GDP

[Bia15] Jacek Bialek. Generalization of the divisia price and quantity indices in a stochastic model with continuous time. Communications in Statistics: Theory and Methods, 44(2):309–328, 2015.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bialek:2017:BEV

[Bia17] Jacek Bialek. Bounds for the expected value of the stochastic Divisia’s price index. Communications in Statistics: Theory and Methods, 46(14):7134–7146, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1148737.

Bibi:2016:TSG

[Bib16] Abdelouahab Bibi. Two-stage generalized moment method approach for bidimensional random coefficient autoregressive models. Communications in Statistics: Theory and Methods, 45(14):4268–4284, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
919401.

Bickel:2012:CFD

[Bic12] David R. Bickel. Coherent frequentism: a decision theory based on confidence sets. Communications in Statistics: Theory and Methods, 41(8):1478–1496, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bickel:2017:CDA

[Bic17] David R. Bickel. Confidence distributions applied to propagating uncertainty to inference based on estimating the local false discovery rate: a fiducial continuum from confidence sets to empirical Bayes set estimates as the number of comparisons increases. Communications in Statistics: Theory and Methods, 46(21):10788–10799, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1248781.

Bickel:2018:NFM

[Bic18] David R. Bickel. A note on fiducial model averaging as an alternative to checking Bayesian and frequentist models. Communications in Statistics: Theory and Methods, 47(13):3125–3137, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1348522.

Bickel:2019:RBF

[Bic19] David R. Bickel. Reporting Bayes factors or probabilities to decision makers of unknown loss functions. Communications in Statistics: Theory and Methods, 48(9):2163–2174, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1459713.

Bieniek:2016:CBT

[Bie16] Mariusz Bieniek. Comparison of the bias of trimmed and Winsorized means. Communications in Statistics: Theory and Methods, 45(22):6641–6650, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
963620.

Bennett:2017:AGW

[BIJ17] Nathan Bennett, Srikanth K. Iyer, and S. Rao Jammalamadaka. Analysis of Gamma and Weibull lifetime data under a general censoring scheme and in the presence of covariates. Communications in Statistics: Theory and Methods, 46(5): 2277–2289, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1041981.

Bingham:2014:MPF

[Bin14] N. H. Bingham. Modelling and prediction of financial time series. Communications in Statistics: Theory and Methods, 43(7):1351–1361, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bindele:2017:SCG

[Bin17] Huybrechts F. Bindele. Strong consistency of the general rank estimator. Communications in Statistics: Theory and Methods, 46(2):532–539, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
999091.

Barczy:2010:IOM

[BIP^{+}10] Mátyás Barczy, Márton Ispány, Gyula Pap, Manuel Scotto, and Maria Eduarda Silva. Innovational outliers in INAR(1) models. Communications in Statistics: Theory and Methods, 39(18):

3343–3362, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Biswas:2012:BDR

[Bis12] Amrita Biswas. Block designs robust against the presence of an aberration in a treatment-control setup. Communications in Statistics: Theory and Methods, 41(5):920–933, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhati:2018:WGD

[BJ18] Deepesh Bhati and Savitri Joshi. Weighted geometric distribution with new characterizations of geometric distribution. Communications in Statistics: Theory and Methods, 47(6):1510– 1527, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1321765.

Baghfalaki:2019:RJM

[BJ19] T. Baghfalaki and E. Farahani Jalali. Reversible jump MCMC to identify dropout mechanism in longitudinal data. Communications in Statistics: Theory and Methods, 48(11):2717–2733, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1472790.

Bravo:2016:SQL

[BJC16] Francesco Bravo and David T. Jacho-Chávez. Semiparametric quasi-likelihood estimation with missing data. Communications in Statistics: Theory and Methods, 45(5):1345–1369, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.863928.

Balamurali:2018:EDQ

[BJU18] S. Balamurali, P. Jeyadurga, and M. Usha. Economic design of quick switching sampling system for assuring Weibull distributed mean life. Communications in Statistics: Theory and Methods, 47(2):385–398, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1303738.

Bhowmik:2015:OBD

[BJVV15] Arpan Bhowmik, Seema Jaggi, Cini Varghese, and Eldho Varghese. Optimal block designs with interference effects from neighboring units under a non additive model. Communications in Statistics: Theory and Methods, 44(10):2092–2103, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhowmik:2017:OBD

[BJVV17] Arpan Bhowmik, Seema Jaggi, Eldho Varghese, and Cini Varghese. Optimal block designs with non additive mixed effects interference. Communications in Statistics: Theory and Methods, 46(24):12103–12112, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1291978.

Bouzebda:2009:NTP

[BK09] Salim Bouzebda and Amor Keziou. A new test procedure of independence in copula models via χ^{2}-divergence. Communications in Statistics: Theory and Methods, 39(1):1–20, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Babayi:2017:ISN

[BK17a] Salman Babayi and Esmaile Khorram. Inference of R = P [X < Y] for skew normal distribution. Communications in Statistics: Theory and Methods, 46(20):10315–10326, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1235191.

Budzynski:2017:RDS

[BK17b] Robert J. Budzyński and Witold Kondracki. Remarks on the L_{1} distance in statistical data analysis. Communications in Statistics: Theory and Methods, 46(19):9355–9363, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.799691.

Baratpour:2018:QBG

[BK18] S. Baratpour and A. H. Khammar. A quantile-based generalized dynamic cumulative measure of entropy. Communications in Statistics: Theory and Methods, 47(13):3104–3117, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1348520.

Bibi:2019:WEM

[BK19] Abdelouahab Bibi and Karima Kimouche. Whittle estimation in multivariate CCC–GARCH processes. Communications in Statistics: Theory and Methods, 48(15):3921–3940, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1484140.

Bhati:2017:NCM

[BKGD17] Deepesh Bhati, Pooja Kumawat, and E. Gómez-Déniz. A new count model generated from mixed Poisson transmuted exponential family with an application to health care data. Communications in Statistics: Theory and Methods, 46(22):11060– 11076, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1257712.

Boreiko:2019:IHM

[BKKP19] D. V. Boreiko, S. Y. Kaniovski, Y. M. Kaniovski, and G. Ch. Pflug. Identification of hidden Markov chains governing dependent credit-rating migrations. Communications in Statistics: Theory and Methods, 48(1):75–87, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1342841.

Bedbur:2019:CRP

[BKL19] Stefan Bedbur, Udo Kamps, and Jens M. Lennartz. Confidence regions for Pareto parameters from a single and independent samples. Communications in Statistics: Theory and Methods, 48(13):3341–3359, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1476706.

Balakrishnan:2018:WPD

[BKM18] N. Balakrishnan, M. V. Koutras, and F. S. Milienos. A weighted Poisson distribution and its application to cure rate models. Communications in Statistics: Theory and Methods, 47(17):

4297–4310, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1373817.

Bakouch:2016:EPD

[BKN16] Hassan S. Bakouch, Maher Kachour, and Saralees Nadarajah. An extended Poisson distribution. Communications in Statistics: Theory and Methods, 45(22):6746–6764, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.967587.

Brouste:2011:FDP

[BKP11] Alexandre Brouste, Marina Kleptsyna, and Alexandre Popier. Fractional diffusion with partial observations. Communications in Statistics: Theory and Methods, 40(19–20):3479–3491, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

BenHajria:2019:TAB

[BKR19] Raja Ben Hajria, Salah Khardani, and Hamdi Raïssi. Testing for abrupt breaks in variance structures with smooth changes. Communications in Statistics: Theory and Methods, 48(21): 5195–5212, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1510002.

Bartroff:2010:MTM

[BL10] Jay Bartroff and Tze Leung Lai. Multistage tests of multiple hypotheses. Communications in Statistics: Theory and Methods, 39(8–9):1597–1607, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bibi:2013:EAP

[BL13] Abdelouahab Bibi and Ines Lescheb. Estimation and asymptotic properties in periodic GARCH(1, 1) models. Communications in Statistics: Theory and Methods, 42(19):3497–3513, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beh:2015:CRA

[BL15] Eric J. Beh and Rosaria Lombardo. Confidence regions and approximate p-values for classical and non symmetric correspondence analysis. Communications in Statistics: Theory and Methods, 44(1):95–114, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bu:2019:MGC

[BL19] Qihui Bu and Liwei Liu. An M/G/1 clearing queueing system with setup time and multiple vacations for an unreliable server. Communications in Statistics: Theory and Methods, 48(11): 2810–2826, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1473426.

Bagdonavicius:2013:CSG

[BLN13] Vilijandas B. Bagdonavicius, Ruta J. Levuliene, and Mikhail S. Nikulin. Chi-squared goodness-of-fit tests for parametric accelerated failure time models. Communications in Statistics: Theory and Methods, 42(15):2768–2785, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Benziadi:2016:RKE

[BLT16] Fatima Benziadi, Ali Laksaci, and Fethallah Tebboune. Recursive kernel estimate of the conditional quantile for functional ergodic data. Communications in Statistics: Theory and Methods, 45(11):3097–3113, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
901364.

Bhattacharjee:2010:SFE

[BM10] Debanjan Bhattacharjee and Nitis Mukhopadhyay. Stirling’s formula and its extensions: Heuristic approaches. Communications in Statistics: Theory and Methods, 39(6):1046–1053, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bentarzi:2011:MMP

[BM11] M. Bentarzi and M. Merzougui. Moments of mixture periodic autoregressive models. Communications in Statistics: Theory and Methods, 40(22):3937–3947, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bandyopadhyay:2015:ACD

[BM15] Uttam Bandyopadhyay and Shirsendu Mukherjee. Adaptive crossover design for normal responses. Communications in Statistics: Theory and Methods, 44(7):1466–1482, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boente:2016:EAM

[BM16] Graciela Boente and Alejandra M. Martínez. Estimating additive models with missing responses. Communications in Statistics: Theory and Methods, 45(2):413–429, 2016. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beder:2017:NMS

[BM17] Jay H. Beder and Margaret Ann McComack. A note on the minimum size of an orthogonal array. Communications in Statistics: Theory and Methods, 46(8):3690–3697, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1071391.

Bolfarine:2018:BSA

[BMFS18] Heleno Bolfarine, Guillermo Martínez-Flórez, and Hugo S. Salinas. Bimodal symmetric-asymmetric power-normal families. Communications in Statistics: Theory and Methods, 47(2):259– 276, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2013.765475.

Bian:2019:LDT

[BMLY19] Lina Bian, Ming Ma, Hua Liu, and Jian Hua Ye. Lifetime distribution of two discrete censored δ shock models. Communications in Statistics: Theory and Methods, 48(14):3451–3463, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1477961.

Bessegato:2016:OCA

[BMQ16] Lupércio F. Bessegato, Lucas S. Mota, and Roberto C. Quinino. Online control by attributes in the presence of classification errors with variable inspection interval. Communications in Statistics: Theory and Methods, 45(11):3283–3301, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901376.

Belzunce:2017:SCC

[BMR17] Félix Belzunce and Carolina Martínez-Riquelme. On sufficient conditions for the comparison of some quantile-based measures. Communications in Statistics: Theory and Methods, 46(13): 6512–6527, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1129421.

Bera:2017:NRM

[BMRSE17] Anil K. Bera, Gabriel Montes-Rojas, and Walter Sosa-Escudero. A new robust and most powerful test in the presence of local misspecification. Communications in Statistics: Theory and Methods, 46(16):8187–8198, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1177077.

Bordbar:2016:MEG

[BN16] F. Bordbar and A. R. Nematollahi. The modified exponentialgeometric distribution. Communications in Statistics: Theory and Methods, 45(1):173–181, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baek:2017:ECM

[BN17] Jong-Il Baek and Si-Li Niu. Estimation of conditional mode with truncated, censored, and dependent data. Communications in Statistics: Theory and Methods, 46(12):6000–6016, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1116575.

Babilua:2018:DBK

[BN18a] P. K. Babilua and E. A. Nadaraya. On deviations between kernel-type estimators of a distribution density in p ≥ 2 independent samples. Communications in Statistics: Theory and Methods, 47(2):475–492, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1307404.

Bejleri:2018:BFP

[BN18b] Valbona Bejleri and Balgobin Nandram. Bayesian and frequentist prediction limits for the Poisson distribution. Communications in Statistics: Theory and Methods, 47(17):4254–4271, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1373814.

Barmalzan:2017:OSP

[BNB17] Ghobad Barmalzan, Amir T. Payandeh Najafabadi, and Narayanaswamy Balakrishnan. Orderings for series and parallel systems comprising heterogeneous exponentiated Weibullgeometric components. Communications in Statistics: Theory and Methods, 46(20):9869–9880, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1222432.

Barmalzan:2018:SNR

[BNB18] Ghobad Barmalzan, Amir. T. Payandeh Najafabadi, and Narayanaswamy Balakrishnan. Some new results on aggregate claim amounts from two heterogeneous Marshall-Olkin extended exponential portfolios. Communications in Statistics: Theory and Methods, 47(11):2779–2794, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1343844.

Barmalzan:2019:ORS

[BNB19] Ghobad Barmalzan, Amir T. Payandeh Najafabadi, and Narayanaswamy Balakrishnan. Ordering results for series and parallel systems comprising heterogeneous exponentiated Weibull components. Communications in Statistics: Theory and Methods, 48(3):660–675, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1417433.

Barakat:2015:ADG

[BNE15] H. M. Barakat, E. M. Nigm, and A. M. Elsawah. Asymptotic distributions of the generalized range, midrange, extremal quotient, and extremal product, with a comparison study. Communications in Statistics: Theory and Methods, 44(5):900–913, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bartkute-Norkuniene:2011:CES

[BNS11] Vaida Bartkute-Norkuniene and Leonidas Sakalauskas. Consistent estimator of the shape parameter of three-dimensional Weibull distribution. Communications in Statistics: Theory and Methods, 40(16):2985–2996, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barakat:2017:LDO

[BNZ17] H. M. Barakat, E. M. Nigm, and E. O. Abo Zaid. Limit distributions of order statistics with random indices in a stationary Gaussian sequence. Communications in Statistics: Theory and Methods, 46(14):7099–7113, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1148732.

Bhar:2014:OMR

[BO14] Lalmohan Bhar and Sankalpa Ojha. Outliers in multi-response experiments. Communications in Statistics: Theory and Methods, 43(13):2782–2798, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhar:2017:IMB

[BO17] Lalmohan Bhar and Sankalpa Ojha. Influence measures in blocked designs of experiments with correlated errors. Communications in Statistics: Theory and Methods, 46(5):2411–2434, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1045079.

Bobotas:2019:ESS

[Bob19] Panayiotis Bobotas. Estimation of the smallest scale parameter of two-parameter exponential distributions. Communications in Statistics: Theory and Methods, 48(11):2748–2765, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1472792.

Bonnini:2014:THC

[Bon14] Stefano Bonnini. Testing for heterogeneity with categorical data: Permutation solution vs. bootstrap method. Communications in Statistics: Theory and Methods, 43(4):906–917, 2014.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bosq:2015:MAE

[Bos15] Denis Bosq. Models associated with extended exponential smoothing. Communications in Statistics: Theory and Methods, 44(3):468–475, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bosq:2018:DIJ

[Bos18] D. Bosq. Detecting instants of jumps and estimating their intensity in the context of p derivatives with continuous or discrete data. Communications in Statistics: Theory and Methods, 47(13):3234–3251, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1353622.

Baiche:2017:LSV

[BOT17] Leila Baiche and Megdouda Ourbih-Tari. Large-sample variance of simulation using refined descriptive sampling: Case of independent variables. Communications in Statistics: Theory and Methods, 46(1):510–519, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
997362.

Bowater:2017:FCP

[Bow17] Russell J. Bowater. A formulation of the concept of probability based on the use of experimental devices. Communications in Statistics: Theory and Methods, 46(10):4774–4790, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1085564.

Bajirao:2015:MSC

[BP15a] Adnaik Sachin Bajirao and Gadre Mukund Parasharam. Modified synthetic control chart for one-step Markov-dependent processes. Communications in Statistics: Theory and Methods, 44 (5):942–952, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2015:LIF

[BP15b] N. Balakrishnan and Suvra Pal. Likelihood inference for flexible cure rate models with gamma lifetimes. Communications in Statistics: Theory and Methods, 44(19):4007–4048, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bakouch:2016:LFO

[BP16] Hassan S. Bakouch and Bozidar V. Popović. Lindley first-order autoregressive model with applications. Communications in Statistics: Theory and Methods, 45(17):4988–5006, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.935429.

Barot:2017:PAC

[BP17] D. R. Barot and M. N. Patel. Posterior analysis of the compound Rayleigh distribution under balanced loss functions for censored data. Communications in Statistics: Theory and Methods, 46(3):1317–1336, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1019140.

Banerjee:2018:KST

[BP18a] Buddhananda Banerjee and Biswabrata Pradhan. Kolmogorov– Smirnov test for life test data with hybrid censoring. Communications in Statistics: Theory and Methods, 47(11):2590–2604, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1205616.

Baxevani:2018:RSM

[BP18b] Anastassia Baxevani and Krzysztof Podgórski. Random spectral measure for non Gaussian moving averages. Communications in Statistics: Theory and Methods, 47(2):448–462, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1303737.

Bhushan:2018:ORT

[BP18c] Shashi Bhushan and Abhay Pratap Pandey. Optimality of ratio type estimation methods for population mean in the presence of missing data. Communications in Statistics: Theory and Methods, 47(11):2576–2589, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1167906.

Baran:2016:TSS

[BPS16a] Sándor Baran, Gyula Pap, and Kinga Sikolya. Testing stability in a spatial unilateral autoregressive model. Communications in Statistics: Theory and Methods, 45(4):933–949, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.853792.

Bordes:2016:PIP

[BPS16b] Laurent Bordes, Christian Paroissin, and Ali Salami. Parametric inference in a perturbed gamma degradation process. Communications in Statistics: Theory and Methods, 45(9):2730– 2747, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.892133.

Bonnini:2012:PIC

[BPSS12] Stefano Bonnini, Domenico Piccolo, Luigi Salmaso, and Francesca Solmi. Permutation inference for a class of mixture models. Communications in Statistics: Theory and Methods, 41(16–17):2879–2895, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
590915.

Bonnini:2014:PAS

[BPSV14] Stefano Bonnini, Nicola Prodi, Luigi Salmaso, and Chiara Visentin. Permutation approaches for stochastic ordering. Communications in Statistics: Theory and Methods, 43(10–12): 2227–2235, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bersimis:2019:UCW

[BPV19] Fragkiskos G. Bersimis, Demosthenes Panagiotakos, and Malvina Vamvakari. The use of components’ weights improves the diagnostic accuracy of a health-related index. Communications in Statistics: Theory and Methods, 48(1):141–164, 2019.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1388401.

Balakrishnan:2017:MBT

[BQC17] Narayanaswamy Balakrishnan, Hong Qin, and Kashinath Chatterjee. Multivariate Bayesian U-type asymmetric designs for non parametric response surface prediction under correlated errors. Communications in Statistics: Theory and Methods, 46(9):4226–4239, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1080843.

Bessegato:2012:LPC

[BQDH12] Lupércio F. Bessegato, Roberto C. Quinino, Luiz H. Duczmal, and Linda Lee Ho. On-line process control using attributes with misclassification errors: an economical design for shortrun production. Communications in Statistics: Theory and Methods, 41(10):1813–1832, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buldygin:2011:ASC

[BR11] Valerii V. Buldygin and Marina K. Runovska. Almost sure convergence of the series of Gaussian Markov sequences. Communications in Statistics: Theory and Methods, 40(19–20):3407– 3424, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baratpour:2012:TGF

[BR12] S. Baratpour and A. Habibi Rad. Testing goodness-of-fit for exponential distribution based on cumulative residual entropy. Communications in Statistics: Theory and Methods, 41(8): 1387–1396, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beyene:2013:NME

[BR13] Semhar Beyene and V. Ramakrishnan. A numerical method for estimating the variance of age at maximum growth rate in growth models. Communications in Statistics: Theory and Methods, 42(8):1464–1475, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Badaoui:2017:ASC

[BR17] Mohammed Badaoui and Noureddine Rhomari. L_{2}(R^{d})- almost sure convergence for multivariate probability density estimate from dependent observations. Communications in Statistics: Theory and Methods, 46(3):1306–1316, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019139.

Bazyari:2019:FTR

[BR19] Abouzar Bazyari and Rasool Roozegar. Finite time ruin probability and structural density properties in the presence of dependence in insurance risk model. Communications in Statistics: Theory and Methods, 48(5):1284–1304, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1429628.

Biswas:2017:SER

[BRAS17] Ankur Biswas, Anil Rai, Tauqueer Ahmad, and Prachi Misra Sahoo. Spatial estimation and rescaled spatial bootstrap approach for finite population. Communications in Statistics: Theory and Methods, 46(1):373–388, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
995820.

Bekker:2011:BVB

[BREA11] A. Bekker, J. J. J. Roux, R. Ehlers, and M. Arashi. Bimatrix variate beta Type IV distribution: Relation to Wilks’s statistic and bimatrix variate Kummer–beta Type IV distribution. Communications in Statistics: Theory and Methods, 40(23): 4165–4178, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brumback:2010:BRL

[Bru10] Babette A. Brumback. On the built-in restrictions in linear mixed models, with application to smoothing spline analysis of variance. Communications in Statistics: Theory and Methods, 39(4):579–591, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bitaraf:2017:GFT

[BRY17] M. Bitaraf, M. Rezaei, and F. Yousefzadeh. Goodness-of-fit tests based on Verma Kullback–Leibler information. Communications in Statistics: Theory and Methods, 46(24):12152–12164, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1291970.

Babadi:2016:VSM

[BRZR16] B. Babadi, A. Rasekh, K. Zare, and A. A. Rasekhi. A variance shift model for detection of outliers in the linear mixed measurement error models. Communications in Statistics: Theory and Methods, 45(24):7350–7366, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.980517.

Bai:2012:ABR

[BS12a] Xiaodong Bai and Lixin Song. Asymptotic behavior of random time ruin probability under heavy-tailed claim sizes and dependence structure. Communications in Statistics: Theory and Methods, 41(10):1721–1732, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2012:RMP

[BS12b] N. Balakrishnan and H. M. Saleh. Relations for moments of progressively Type-II censored order statistics from log-logistic distribution with applications to inference. Communications in Statistics: Theory and Methods, 41(5):880–906, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2012:OTS

[BS12c] N. Balakrishnan and A. R. Shafay. One- and two-sample Bayesian prediction intervals based on Type-II hybrid censored data. Communications in Statistics: Theory and Methods, 41(9):1511–1531, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bulinskaya:2014:ISU

[BS14] Ekaterina Bulinskaya and Kseniya Shakhgildyan. Inventory systems with unreliable suppliers. Communications in Statistics: Theory and Methods, 43(7):1563–1574, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bianchi:2015:APV

[BS15] Annamaria Bianchi and Nicola Salvati. Asymptotic properties and variance estimators of the M-quantile regression coefficients estimators. Communications in Statistics: Theory and Methods, 44(11):2416–2429, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bandyopadhyay:2016:ERP

[BS16a] A. Bandyopadhyay and G. N. Singh. Estimation of ratio of population means in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 45(14):4136–4155, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.917185.

Bandyopadhyay:2016:PEP

[BS16b] A. Bandyopadhyay and G. N. Singh. Predictive estimation of population mean in two-phase sampling. Communications in Statistics: Theory and Methods, 45(14):4249–4267, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.919396.

Batool:2016:TSD

[BS16c] Fatima Batool and Javid Shabbir. A two-stage design for multivariate estimation of proportions. Communications in Statistics: Theory and Methods, 45(18):5412–5426, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.942435.

Beran:2017:NPS

[BS17] Jan Beran and Nadja Schumm. On non parametric statistical inference for densities under long-range dependence. Communications in Statistics: Theory and Methods, 46(22):11296–11314, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1263740.

Bhattacharya:2019:MTT

[BS19] Rahul Bhattacharya and Madhumita Shome. A multitreatment two stage adaptive allocation for survival outcomes.

Communications in Statistics: Theory and Methods, 48(8): 2001–2013, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1440599.

Bush:2012:ODS

[BSB12] Stephen Bush, Deborah J. Street, and Leonie Burgess. Optimal designs for stated choice experiments that incorporate position effects. Communications in Statistics: Theory and Methods, 41(10):1771–1795, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barreto-Souza:2011:SRB

[BSCS11] Wagner Barreto-Souza, Gauss M. Cordeiro, and Alexandre B. Simas. Some results for beta fréchet distribution. Communications in Statistics: Theory and Methods, 40(5):798–811, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Belles-Sampera:2016:UFQ

[BSGS16] Jaume Belles-Sampera, Montserrat Guillén, and Miguel Santolino. The use of flexible quantile-based measures in risk assessment. Communications in Statistics: Theory and Methods, 45(6):1670–1681, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
938829.

Batool:2017:ESQ

[BSH17] Fatima Batool, Javid Shabbir, and Zawar Hussain. On the estimation of a sensitive quantitative mean using blank cards. Communications in Statistics: Theory and Methods, 46(6): 3070–3079, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1053944.

Bai:2019:SAC

[BSL^{+}19] Xuchao Bai, Yimin Shi, Yiming Liu, Hon Keung Tony Ng, and Bin Liu. Statistical analysis of competing risks model from Marshall–Olkin extended Chen distribution under adaptive progressively interval censoring with random removals. Communications in Statistics: Theory and Methods, 48(15): 3683–3702, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1481973.

Brusco:2019:MIR

[BSS19] Michael J. Brusco, Douglas Steinley, and Jordan Stevens. K-medoids inverse regression. Communications in Statistics: Theory and Methods, 48(20):4999–5011, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1504076.

Bhattacharjee:2011:PIU

[BSSJ11] S. K. Bhattacharjee, Ahmed Shamiri, Md. Sabiruzzaman, and S. Rao Jammalamadaka. Predictive influence of unavailable values of future explanatory variables in a linear model. Communications in Statistics: Theory and Methods, 40(24):4458–4466, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bai:2015:UAR

[BSW15] Xiaodong Bai, Lixin Song, and Yuebao Wang. Uniform asymptotics for random time ruin probability with subexponential claims and constant interest rate. Communications in Statistics: Theory and Methods, 44(14):2976–2983, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:2012:WCI

[BSY12] B. M. Brown, Thomas Suesse, and Von Bing Yap. Wilson confidence intervals for the two-sample log-odds-ratio in stratified 2 × 2 contingency tables. Communications in Statistics: Theory and Methods, 41(18):3355–3370, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baringhaus:2013:KTT

[BT13] L. Baringhaus and F. Taherizadeh. A K–S type test for exponentiality based on empirical Hankel transforms. Communications in Statistics: Theory and Methods, 42(20):3781–3792, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhatt:2016:SSS

[BT16] Vijaya Bhatt and Neeraj Tiwari. A spatial scan statistic for survival data based on generalized life distribution. Communications in Statistics: Theory and Methods, 45(19):5730–5744, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.948207.

Bee:2018:CFB

[BT18a] M. Bee and L. Trapin. A characteristic function-based approach to approximate maximum likelihood estimation. Communications in Statistics: Theory and Methods, 47(13):3138–3160, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1348523.

Beran:2018:NDE

[BT18b] Jan Beran and Klaus Telkmann. On nonparametric density estimation for multivariate linear long-memory processes. Communications in Statistics: Theory and Methods, 47(22):5460– 5473, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1395048.

Baey:2016:NLM

[BTC16] Charlotte Baey, Samis Trevezas, and Paul-Henry Cournède. A non linear mixed effects model of plant growth and estimation via stochastic variants of the EM algorithm. Communications in Statistics: Theory and Methods, 45(6):1643–1669, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.930909.

Balakrishnan:2010:DFC

[BTK10] N. Balakrishnan, I. S. Triantafyllou, and M. V. Koutras. A distribution-free control chart based on order statistics. Communications in Statistics: Theory and Methods, 39(20):3652– 3677, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balamurali:2017:DVQ

[BU17] S. Balamurali and M. Usha. Designing of variables quick switching sampling system by considering process loss functions. Communications in Statistics: Theory and Methods, 46(5): 2299–2314, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1041983.

Budny:2016:EMC

[Bud16] Katarzyna Budny. An extension of the multivariate Chebyshev’s inequality to a random vector with a singular covariance matrix. Communications in Statistics: Theory and Methods, 45(17):5220–5223, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
941499.

Burkschat:2009:LEP

[Bur09] M. Burkschat. Linear estimators and predictors based on generalized order statistics from generalized Pareto distributions. Communications in Statistics: Theory and Methods, 39(2):311– 326, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burch:2015:ILS

[Bur15] Brent D. Burch. Improved large-sample confidence intervals for ratios of variance components in nonnormal distributions. Communications in Statistics: Theory and Methods, 44(2):349– 362, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burgess:2017:ECA

[Bur17] Stephen Burgess. Estimating and contextualizing the attenuation of odds ratios due to non collapsibility. Communications in Statistics: Theory and Methods, 46(2):786–804, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1006778.

Bura:2019:TSM

[Bur19] Gulab Singh Bura. Transient solution of an M/M/∞ queue with catastrophes. Communications in Statistics: Theory and Methods, 48(14):3439–3450, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1477960.

Bhowmik:2017:MCR

[BVJV17] Arpan Bhowmik, Eldho Varghese, Seema Jaggi, and Cini Varghese. Minimally changed run sequences in factorial experiments. Communications in Statistics: Theory and Methods, 46(15):7444–7459, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1152490.

Baetschmann:2017:DHM

[BW17] Gregori Baetschmann and Rainer Winkelmann. A dynamic hurdle model for zero-inflated count data. Communications in Statistics: Theory and Methods, 46(14):7174–7187, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1146766.

Boumaza:2015:IPC

[BYDM15] Rachid Boumaza, Smail Yousfi, and Sabine Demotes-Mainard. Interpreting the principal component analysis of multivariate density functions. Communications in Statistics: Theory and Methods, 44(16):3321–3339, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bouzebda:2013:SIT

[BZ13] Salim Bouzebda and Tarek Zari. A strong invariance theorem of the tail empirical copula processes. Communications in Statistics: Theory and Methods, 42(1):11–27, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:2015:RAS

[BZAZ15] N. Balakrishnan, Xiaojun Zhu, and Bander Al-Zahrani. A recursive algorithm for the single and product moments of order statistics from the exponential-geometric distribution and some estimation methods. Communications in Statistics: Theory and Methods, 44(17):3576–3598, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bai:2019:SLL

[BZH19] Xiaodong Bai, Qingjian Zhou, and Zhiqiang Hua. Strong law of large numbers and complete convergence for non-identically distributed WOD random variables. Communications in Statistics: Theory and Methods, 48(11):2687–2699, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472787.

Chitsaz:2012:IER

[CA12] S. Chitsaz and S. Ejaz Ahmed. An improved estimation in regression parameter matrix in multivariate regression model. Communications in Statistics: Theory and Methods, 41(13–14): 2305–2320, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coelho:2014:ENE

[CA14] Carlos A. Coelho and Barry C. Arnold. On the exact and near-exact distributions of the product of generalized gamma random variables and the generalized variance. Communications in Statistics: Theory and Methods, 43(10–12):2007–2033, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christou:2018:VSH

[CA18] Eliana Christou and Michael G. Akritas. Variable selection in heteroscedastic single-index quantile regression. Communications in Statistics: Theory and Methods, 47(24):6019–6033, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1405271.

Cagnone:2012:NGF

[Cag12] S. Cagnone. A note on goodness-of-fit test in latent variable models with categorical variables. Communications in Statistics: Theory and Methods, 41(16–17):2983–2990, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2011.622424.

Cahoy:2012:ESW

[Cah12] Dexter O. Cahoy. Estimation and simulation for the M-Wright function. Communications in Statistics: Theory and Methods, 41(8):1466–1477, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cahoy:2015:SSS

[Cah15] Dexter O. Cahoy. Some skew-symmetric distributions which include the bimodal ones. Communications in Statistics: Theory and Methods, 44(3):554–563, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Canale:2015:NRG

[Can15] Antonio Canale. A note on regions of given probability of the extended skew-normal distribution. Communications in Statistics: Theory and Methods, 44(12):2507–2516, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2017:OLL

[CAPO17] Gauss M. Cordeiro, Morad Alizadeh, Rodrigo R. Pescim, and Edwin M. M. Ortega. The odd log-logistic generalized halfnormal lifetime distribution: Properties and applications. Communications in Statistics: Theory and Methods, 46(9):4195– 4214, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1080841.

Carley:2017:EDS

[Car17] Holly Carley. Examples of doubly stochastic measures supported on the graphs of two functions. Communications in Statistics: Theory and Methods, 46(17):8628–8630, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1186189.

Caruana:2019:NPE

[Car19] Mark Anthony Caruana. Non parametric estimation of the measure associated with the lévy-Khintchine canonical representation. Communications in Statistics: Theory and Methods, 48(1):100–111, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1401083.

Cordeiro:2017:GOH

[CARO17] Gauss M. Cordeiro, Morad Alizadeh, Thiago G. Ramires, and Edwin M. M. Ortega. The generalized odd half-Cauchy family of distributions: Properties and applications. Communications in Statistics: Theory and Methods, 46(11):5685–5705, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1109665.

Cavicchioli:2017:TFM

[Cav17] Maddalena Cavicchioli. Third and fourth moments of vector autoregressions with regime switching. Communications in Statistics: Theory and Methods, 46(9):4181–4194, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1080840.

Chikkagoudar:2010:RTT

[CB10] M. S. Chikkagoudar and B. S. Biradar. Rank tests for twosample problems based on multiple Type-II censored data. Communications in Statistics: Theory and Methods, 39(18): 3203–3221, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cramer:2011:MMI

[CB11] Erhard Cramer and Christain Bagh. Minimum and maximum information censoring plans in progressive censoring. Communications in Statistics: Theory and Methods, 40(14):2511–2527, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chikkagoudar:2012:LMP

[CB12] M. S. Chikkagoudar and B. S. Biradar. Locally most powerful rank tests for comparison of two failure rates based on multiple Type-II censored data. Communications in Statistics: Theory and Methods, 41(23):4315–4331, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chamany:2014:DDI

[CB14] A. Chamany and S. Baratpour. A dynamic discrimination information based on cumulative residual entropy and its properties. Communications in Statistics: Theory and Methods, 43(6): 1041–1049, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2016:NRR

[CB16] Gauss M. Cordeiro and Marcelo Bourguignon. New results on the Ristić–Balakrishnan family of distributions. Communications in Statistics: Theory and Methods, 45(23):6969–6988, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.972573.

Cha:2017:PRF

[CB17] Ji Hwan Cha and F. G. Badía. Prediction of the residual failure processes based on the process history. Communications in Statistics: Theory and Methods, 46(13):6336–6357, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1122059.

Cordeiro:2006:CMF

[CBB06] Gauss M. Cordeiro, Lúcia P. Barroso, and Denise A. Botter. Covariance matrix formula for generalized linear models with unknown dispersion. Communications in Statistics: Theory and Methods, 35(1):113–120, 2006. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See corrigendum [MBS17].

Coetsee:2014:PTB

[CBM14] J. Coetsee, A. Bekker, and S. Millard. Preliminary test and Bayes estimation of a location parameter under BLINEX loss. Communications in Statistics: Theory and Methods, 43(17): 3641–3660, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2018:GMP

[CBOR18] Gauss M. Cordeiro, Marcelo Bourguignon, Edwin M. M. Ortega, and Thiago G. Ramires. General mathematical properties, regression and applications of the log-gamma-generated family. Communications in Statistics: Theory and Methods, 47(5):1050–1070, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1316403.

Chan:2016:OSS

[CBSN16] Ping Shing Chan, Narayanaswamy Balakrishnan, Hon Yiu So, and Hon Keung Tony Ng. Optimal sample size allocation for multi-level stress testing with exponential regression under Type-I censoring. Communications in Statistics: Theory and Methods, 45(6):1831–1852, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1078474.

Cerqueti:2010:SNA

[CC10a] Roy Cerqueti and Mauro Costantini. Some nonparametric asymptotic results for a class of stochastic processes. Communications in Statistics: Theory and Methods, 39(14):2552–2560, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2010:OSO

[CC10b] Jih-An Chen and Yu-Hung Chien. Optimal spare ordering time for preventive replacement to maximize the cost effectiveness. Communications in Statistics: Theory and Methods, 39(13): 2404–2421, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2010:GCC

[CC10c] Vivian Yi-Ju Chen and Vernon M. Chinchilli. A generalized correlation coefficient for the general two-treatment crossover design. Communications in Statistics: Theory and Methods, 39(11):2063–2080, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chien:2010:OAP

[CC10d] Yu-Hung Chien and Mingchih Chen. An optimal age for preventive replacement under fully renewing combination free replacement and Pro rata warranty. Communications in Statistics: Theory and Methods, 39(13):2422–2439, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2011:SDC

[CC11] Pei-Wen Chen and Chuen-Sheng Cheng. On statistical design of the cumulative quantity control chart for monitoring high yield processes. Communications in Statistics: Theory and Methods, 40(11):1911–1928, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborty:2012:DGD

[CC12] Subrata Chakraborty and Dhrubajyoti Chakravarty. Discrete gamma distributions: Properties and parameter estimations. Communications in Statistics: Theory and Methods, 41(18): 3301–3324, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cazzaro:2014:MNI

[CC14] Manuela Cazzaro and Roberto Colombi. Marginal nested interactions for contingency tables. Communications in Statistics: Theory and Methods, 43(13):2799–2814, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:2015:DPC

[CC15a] Moutushi Chatterjee and Ashis Kumar Chakraborty. Distributions and process capability control charts for C_{P U} and C_{P L} using subgroup information. Communications in Statistics: Theory and Methods, 44(20):4333–4353, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:2015:SPC

[CC15b] Moutushi Chatterjee and Ashis Kumar Chakraborty. A superstructure of process capability indices for circular specification region. Communications in Statistics: Theory and Methods, 44(6):1158–1181, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborty:2016:NDP

[CC16a] Subrata Chakraborty and Dhrubajyoti Chakravarty. A new discrete probability distribution with integer support on (−∞, ∞). Communications in Statistics: Theory and Methods, 45(2):492– 505, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:2016:SPC

[CC16b] Moutushi Chatterjee and Ashis Kumar Chakraborty. Some process capability indices for unilateral specification limits — their properties and the process capability control charts. Communications in Statistics: Theory and Methods, 45(24):7130–7160, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.978023.

Chen:2016:ECM

[CC16c] Zhengjia Chen and Xinjia Chen. Exact calculation of minimum sample size for estimating a Poisson parameter. Communications in Statistics: Theory and Methods, 45(16):4692–4715, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.927497.

Chang:2019:OCD

[CC19] Chin-Chih Chang and Yen-Luan Chen. Optimization of continuous and discrete scheduled times for a cumulative damage system with age-dependent maintenance. Communications in Statistics: Theory and Methods, 48(17):4261–4277, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1491992.

Chang:2016:EDT

[CCCT16] Yung-Chia Chang, Kuei-Hu Chang, Heng-Hsuan Chu, and Lee-Ing Tong. Establishing decision tree-based short-term default credit risk assessment models. Communications in Statistics: Theory and Methods, 45(23):6803–6815, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.968730.

Chen:2010:CFF

[CCH^{+}10] Chun-Ying Chen, Pei-Ju Chou, Jeff Yu-Shun Hsu, Wisely Po-Hong Liu, Yuh-Dauh Lyuu, and Chuan-Ju Wang. A closedform formula for an option with discrete and continuous barriers. Communications in Statistics: Theory and Methods, 40 (2):345–357, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2013:NIF

[CCL^{+}13] Feng Chen, Qiang Cheng, Hong Liu, Wenli Xu, and Song Wang. A new inference framework for dependency networks. Communications in Statistics: Theory and Methods, 42(1):56–75, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:2014:MEP

[CCL14] Jennifer S. K. Chan, S. T. Boris Choy, and Connie P. Y. Lam. Modeling electricity price using a threshold conditional autoregressive geometric process jump model. Communications in Statistics: Theory and Methods, 43(10–12):2505–2515, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2019:PEG

[CCM19] Haiqing Chen, Weihu Cheng, and Jin Mingzhong. Parameter estimation for generalized logistic distribution by estimating equations based on the order statistics. Communications in Statistics: Theory and Methods, 48(6):1506–1516, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1433854.

Coolen:2018:NPP

[CCMA18] Frank P. A. Coolen, Tahani Coolen-Maturi, and Hana N. Alqifari. Non-parametric predictive inference for future order statistics. Communications in Statistics: Theory and Methods, 47(10):2527–2548, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1342834.

Chaubey:2017:LWE

[CCN17] Yogendra P. Chaubey, Christophe Chesneau, and Fabien Navarro. Linear wavelet estimation of the derivatives of a regression function based on biased data. Communications in Statistics: Theory and Methods, 46(19):9541–9556, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1213287.

Cordeiro:2016:MLT

[CCOB16] Gauss M. Cordeiro, Vicente G. Cancho, Edwin M. M. Ortega, and Gladys D. C. Barriga. A model with long-term survivors: negative binomial Birnbaum–Saunders. Communications in Statistics: Theory and Methods, 45(5):1370–1387, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.863929.

Coolen:2013:NPI

[CCSCME13] F. P. A. Coolen, P. Coolen-Schrijner, T. Coolen-Maturi, and F. F. Elkhafifi. Nonparametric predictive inference for ordinal data. Communications in Statistics: Theory and Methods, 42 (19):3478–3496, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cui:2017:ALG

[CCY17] Yan Cui, Xia Chen, and Li Yan. Adaptive Lasso for generalized linear models with a diverging number of parameters. Communications in Statistics: Theory and Methods, 46(23):11826– 11842, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1285926.

Chang:2015:PCI

[CCZ15] Liu Chang, Shouting Chen, and Ailin Zhu. Properties of the Cox–Ingersoll–Ross interest rate processes with two-sided reflections. Communications in Statistics: Theory and Methods, 44(4):657–670, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2017:PET

[CCZR17] Haiqing Chen, Weihu Cheng, Leilei Zhu, and Yaohua Rong. Parameter estimation for three-parameter generalized Pareto distribution by weighted non linear least squares. Communications in Statistics: Theory and Methods, 46(23):11440–11449, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1202286.

Costa:2012:PBP

[CD12] Michele Costa and Luca De Angelis. A permutation based procedure for classification assessment. Communications in Statistics: Theory and Methods, 41(16–17):3126–3137, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2011.608475.

Costa:2015:DLM

[CD15] Michele Costa and Luca De Angelis. A dynamic latent model for poverty measurement. Communications in Statistics: Theory and Methods, 44(23):5037–5048, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2017:AED

[CD17] Shouquan Chen and Lingling Du. Asymptotic expansions of density of normalized extremes from logarithmic general error distribution. Communications in Statistics: Theory and Methods, 46(7):3459–3478, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1062109.

Chavez-Demoulin:2010:RET

[CDE10] Valérie Chavez-Demoulin and Paul Embrechts. Revisiting the edge, ten years on. Communications in Statistics: Theory and Methods, 39(8–9):1674–1688, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Corset:2012:BAA

[CDG12] Franck Corset, Laurent Doyen, and Olivier Gaudoin. Bayesian analysis of ARA imperfect repair models. Communications in Statistics: Theory and Methods, 41(21):3915–3941, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chesneau:2019:AWE

[CDS19] Christophe Chesneau, Hassan Doosti, and Lewi Stone. Adaptive wavelet estimation of a function from an m-dependent process with possibly unbounded m. Communications in Statistics: Theory and Methods, 48(5):1123–1135, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1423700.

Cordeiro:2015:MSK

[CdSBB15] Gauss M. Cordeiro and Rejane dos S. B. Bager. Moments for some Kumaraswamy generalized distributions. Communications in Statistics: Theory and Methods, 44(13):2720–2737, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Corain:2014:PCB

[CDZ14] Livio Corain, Valeria De Giuli, and Roberto Zecchin. A permutation and combination-based solution on the ranking of multivariate populations in the case of ordered categorical responses with application to the evaluation of the indoor environment. Communications in Statistics: Theory and Methods, 43(4):879– 890, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Celik:2017:AMC

[ÇE17a] Resit Çelik and Aydin Erar. An alternative method correcting BDR type of heteroskedasticity by the weighting re-estimated absolute residuals. Communications in Statistics: Theory and Methods, 46(23):11514–11538, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1271429.

Chen:2017:MLT

[CE17b] Ai-Chun Chen and Takeshi Emura. A modified Liu-type estimator with an intercept term under mixture experiments. Communications in Statistics: Theory and Methods, 46(13): 6645–6667, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1132327.

Clement:2017:ERE

[CE17c] Etebong P. Clement and Ekaette I. Enang. On the efficiency of ratio estimator over the regression estimator. Communications in Statistics: Theory and Methods, 46(11):5357–5367, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1100741.

Camilleri:2019:MMR

[CE19] Liberato Camilleri and Kathleen England. Modeling mortality rates in Malta using GEE models. Communications in Statistics: Theory and Methods, 48(1):15–24, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1386313.

Chatterjee:2011:LBM

[CEK11] K. Chatterjee, H. Evangelaras, and C. Koukouvinos. A lower bound to the measure of optimality for main effect plans in the symmetric factorial experiments. Communications in Statistics: Theory and Methods, 40(13):2358–2365, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Celik:2017:CDO

[Çel17] Resit Çelik. Correcting double outward box distributed residuals by WCEV. Communications in Statistics: Theory and Methods, 46(19):9566–9590, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1213289.

Celik:2019:RHT

[Çel19] Resit Çelik. RCEV heteroscedasticity test based on the studentized residuals. Communications in Statistics: Theory and Methods, 48(13):3258–3268, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1475566.

Ceyhan:2011:SCT

[Cey11] Elvan Ceyhan. Spatial clustering tests based on the domination number of a new random digraph family. Communications in Statistics: Theory and Methods, 40(8):1363–1395, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cha:2012:SBH

[CF12] Ji Hwan Cha and Maxim Finkelstein. Shocks as burn-in in heterogeneous populations. Communications in Statistics: Theory and Methods, 41(2):325–340, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chesneau:2013:WBD

[CF13] Christophe Chesneau and Jalal Fadili. Wavelet-based density estimation in a heteroscedastic convolution model. Communications in Statistics: Theory and Methods, 42(17):3085–3099, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cha:2014:BEW

[CF14] Ji Hwan Cha and Maxim Finkelstein. Burn-in for eliminating weak items in heterogeneous populations. Communications in Statistics: Theory and Methods, 43(24):5115–5129, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2016:LDN

[CF16] Zengjing Chen and Xinwei Feng. Large deviation for negatively dependent random variables under sublinear expectation. Communications in Statistics: Theory and Methods, 45(2):400–412, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cha:2019:SRD

[CF19] Ji Hwan Cha and Maxim Finkelstein. Some results on discrete mixture failure rates. Communications in Statistics: Theory and Methods, 48(15):3884–3898, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1482343.

Chinipardaz:2017:DVP

[CFA17] Rahim Chinipardaz, Behzad Falahifard, and Mohammad Reza Akhoond. The discrepancy of P -values and posterior probability: Two-sided hypothesis of variance of normal distribution with unknown mean. Communications in Statistics: Theory and Methods, 46(11):5544–5555, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1105978.

Cuenin:2017:GVP

[CFK17] Johann Cuenin, Adrien Faivre, and Célestin C. Kokonendji. On generalized variance of product of powered components and multiple stable Tweedie models. Communications in Statistics: Theory and Methods, 46(14):7225–7237, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1146770.

Chen:2011:FSP

[CG11a] Qian Chen and David E. Giles. Finite-sample properties of the maximum likelihood estimator for the Poisson regression model with random covariates. Communications in Statistics: Theory and Methods, 40(6):1000–1014, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2011:SAD

[CG11b] Qian Chen and David E. Giles. A saddlepoint approximation to the distribution of the half-life estimator in a stationary autoregressive model. Communications in Statistics: Theory and Methods, 40(21):3903–3916, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ceranka:2012:ROC

[CG12] Bronisław Ceranka and Małgorzata Graczyk. Robustness of optimal chemical balance weighing designs for estimation of total weight. Communications in Statistics: Theory and Methods, 41 (13–14):2297–2304, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ceranka:2014:ROS

[CG14a] Bronisław Ceranka and Małgorzata Graczyk. Regular Eoptimal spring balance weighing designs with correlated errors. Communications in Statistics: Theory and Methods, 43(5):947– 953, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cerchiello:2014:BCR

[CG14b] Paola Cerchiello and Paolo Guidici. Bayesian credit ratings. Communications in Statistics: Theory and Methods, 43(4):867– 878, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborty:2015:EGD

[CG15] Subrata Chakraborty and Rameshwar D. Gupta. Exponentiated geometric distribution: Another generalization of geometric distribution. Communications in Statistics: Theory and Methods, 44(6):1143–1157, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2016:SSM

[CG16] Jie Chen and Joseph Glaz. Scan statistics for monitoring data modeled by a negative binomial distribution. Communications in Statistics: Theory and Methods, 45(6):1632–1642, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.923460.

Chen:2017:EDC

[CG17] Xu Chen and Chang Guisong. Exact distribution of the convolution of negative binomial random variables. Communications in Statistics: Theory and Methods, 46(6):2851–2856, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1053931.

Cetinkaya:2018:MOS

[ÇG18]  Çagatay Çetinkaya and Ali I. Genç. Moments of order statis-

tics of the standard two-sided power distribution. Communications in Statistics: Theory and Methods, 47(17):4311–4328, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1373818.

Ceranka:2019:RDO

[CG19] B. Ceranka and M. Graczyk. Recent developments in Doptimal designs. Communications in Statistics: Theory and Methods, 48(6):1470–1480, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1433851.

Chiancone:2017:CSI

[CGC17] Alessandro Chiancone, Stéphane Girard, and Jocelyn Chanussot. Collaborative sliced inverse regression. Communications in Statistics: Theory and Methods, 46(12):6035–6053, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1116578.

Cook:2013:SCD

[CGL13] Andrea J. Cook, Diane R. Gold, and Yi Li. Spatial cluster detection for longitudinal outcomes using administrative regions. Communications in Statistics: Theory and Methods, 42(12): 2105–2117, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2016:THM

[CGS16] Jie Chen, Joseph Glaz, and Cristina P. Sison. Testing homogeneity of the multinomial proportions. Communications in Statistics: Theory and Methods, 45(6):1747–1777, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.966839.

Colosimo:2010:OMT

[CGSM10] Enrico A. Colosimo, Gustavo L. Gilardoni, Wagner B. Santos, and Sergio B. Motta. Optimal maintenance time for repairable systems under two types of failures. Communications in Statistics: Theory and Methods, 39(7):1289–1298, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cai:2015:MDE

[CGW15] Yujie Cai, Fuqing Gao, and Shaochen Wang. Moderate deviations for estimators of financial risk under an asymmetric Laplace law. Communications in Statistics: Theory and Methods, 44(2):329–339, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chung:2017:PGI

[CH17] Sunah Chung and S. Y. Hwang. A profile godambe information of power transformations for ARCH time series. Communications in Statistics: Theory and Methods, 46(14):6899–6908, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1139133.

Chen:2018:SCS

[CH18] Chun-Shu Chen and Yi-Tsz Huang. Some characteristics on the selection of spline smoothing parameter. Communications in Statistics: Theory and Methods, 47(6):1307–1317, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1317808.

Cao:2019:LAE

[CH19a] MingXiang Cao and DaoJiang He. Linearly admissible estimators on linear functions of regression coefficient under balanced loss function. Communications in Statistics: Theory and Methods, 48(11):2700–2706, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472788.

Chen:2019:SRS

[CH19b] Yanhong Chen and Yijun Hu. Systemic risk statistics with scenario analysis. Communications in Statistics: Theory and Methods, 48(14):3558–3569, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1477959.

Chakraborty:2010:SDP

[Cha10] Subrata Chakraborty. On some distributional properties of the family of weighted generalized Poisson distribution. Communications in Statistics: Theory and Methods, 39(15):2767–2788, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2014:MCL

[Cha14] Sheng-Mao Chang. Maximizing complex likelihoods via directed stochastic searching algorithm. Communications in Statistics: Theory and Methods, 43(20):4281–4296, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborty:2015:NDD

[Cha15] Subrata Chakraborty. A new discrete distribution related to generalized gamma distribution and its properties. Communications in Statistics: Theory and Methods, 44(8):1691–1705, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2016:OPP

[Cha16] C. C. Chang. Optimal preventive policies for imperfect maintenance models with replacement first and last. Communications in Statistics: Theory and Methods, 45(17):5123–5136, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.936942.

Chang:2018:PTA

[Cha18] Xinfeng Chang. On preliminary test almost unbiased twoparameter estimator in linear regression model with student’s t errors. Communications in Statistics: Theory and Methods, 47 (3):583–600, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1309433.

Chechile:2010:NBP

[Che10a] Richard A. Chechile. A novel Bayesian parameter mapping method for estimating the parameters of an underlying scientific model. Communications in Statistics: Theory and Methods, 39(7):1190–1201, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2010:IM

[Che10b] Jinwen Chen. Inference for mixtures. Communications in Statistics: Theory and Methods, 39(7):1247–1257, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2011:EBT

[Che11] Lee-Shen Chen. Empirical Bayes testing for uniform distributions: Non identical components case. Communications in Statistics: Theory and Methods, 40(17):3181–3197, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chesneau:2013:WED

[Che13] Christophe Chesneau. Wavelet estimation of a density in a GARCH-type model. Communications in Statistics: Theory and Methods, 42(1):98–117, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2014:OAP

[Che14] Yen-Luan Chen. Optimal age policy for a used system with imperfect preventive maintenance and cumulative damage model. Communications in Statistics: Theory and Methods, 43(19): 4062–4073, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2016:RSE

[Che16a] Ding-Geng (Din) Chen. Robustness of the shrinkage estimator for the relative potency in the combination of multivariate bioassays. Communications in Statistics: Theory and Methods, 45(18):5380–5391, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
815205.

Chen:2016:NCC

[Che16b] John Tuhao Chen. A nonparametric coherent confidence procedure. Communications in Statistics: Theory and Methods, 45(11):3397–3409, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
901378.

Chen:2016:ELR

[Che16c] Li Chen. Empirical likelihood ratio tests for the linear regression model with inequality constraints. Communications in Statistics: Theory and Methods, 45(20):5933–5945, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.953690.

Cheng:2016:TMS

[Che16d] J. Cheng. A transitional Markov switching autoregressive model. Communications in Statistics: Theory and Methods, 45(10):2785–2800, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
894065.

Chen:2017:EBT

[Che17a] Lee-Shen Chen. Empirical Bayes testing for guarantee lifetime: Non identical components case. Communications in Statistics: Theory and Methods, 46(2):683–705, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.1002939.

Cheng:2017:SUC

[Che17b] Fuxia Cheng. Strong uniform consistency rates of kernel estimators of cumulative distribution functions. Communications in Statistics: Theory and Methods, 46(14):6803–6807, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1136417.

Chechile:2018:BAW

[Che18] Richard A. Chechile. A Bayesian analysis for the Wilcoxon signed-rank statistic. Communications in Statistics: Theory and Methods, 47(21):5241–5254, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1388402.

Chen:2019:OSR

[Che19a] Yen-Luan Chen. Optimal scheduling replacement policies for a system with multiple random works. Communications in Statistics: Theory and Methods, 48(3):676–688, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1417434.

Cheng:2019:CED

[Che19b] Fuxia Cheng. The L_{p} consistency of error density estimator in censored linear regression. Communications in Statistics: Theory and Methods, 48(7):1579–1584, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1438621.

Chen:2015:SCS

[CHH15] Jian Chen, John J. Hanfelt, and Yijian Huang. A simple corrected score for logistic regression with errors-in-covariates. Communications in Statistics: Theory and Methods, 44(10): 2024–2036, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiba:2009:SAU

[Chi09] Yasutaka Chiba. Sensitivity analysis of unmeasured confounding for the causal risk ratio by applying marginal structural models. Communications in Statistics: Theory and Methods, 39(1):65–76, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiba:2010:AEC

[Chi10] Yasutaka Chiba. An approach for estimating causal effects in randomized trials with noncompliance. Communications in Statistics: Theory and Methods, 39(12):2146–2156, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiba:2011:SFE

[Chi11] Yasutaka Chiba. Simple formulae for evaluating the potential impact of confounding bias. Communications in Statistics: Theory and Methods, 40(23):4278–4288, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiba:2012:MCS

[Chi12a] Yasutaka Chiba. Monte-Carlo sensitivity analysis for controlled direct effects using marginal structural models in the presence of confounded mediators. Communications in Statistics: Theory and Methods, 41(10):1739–1749, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Childs:2012:UDM

[Chi12b] Aaron Childs. Unification of Dirichlet methodology. Communications in Statistics: Theory and Methods, 41(9):1647–1662, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2009:MDP

[CHLP09] Shuenn-Ren Cheng, Deng-Yuan Huang, Hung-Yun Lu, and S. Panchapakesan. Multiple decision procedures for inference in regression models. Communications in Statistics: Theory and Methods, 39(2):241–248, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Che:2016:IAP

[CHM16] Suisui Che, Kai Huang, and Jie Mi. Inference about parameters in binomial-Poisson distribution with additional incomplete data. Communications in Statistics: Theory and Methods, 45(11):3206–3222, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
901367.

Cho:2010:RRM

[Cho10] Jaerin Cho. Rational regression models with continuous Chebyshev design. Communications in Statistics: Theory and Methods, 39(5):855–873, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2014:RSS

[CHP14] Pinyuen Chen, Lifang Hsu, and S. Panchapakesan. A restricted subset selection rule for selecting at least one of the t best normal populations in terms of their means when their common variance is known, case II. Communications in Statistics: Theory and Methods, 43(10–12):2250–2259, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christian:2010:LAV

[Chr10] Farinetto Christian. On least absolute value estimators for Poisson processes. Communications in Statistics: Theory and Methods, 39(14):2645–2658, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chvostekova:2013:STS

[Chv13] M. Chvosteková. Simultaneous two-sided tolerance intervals for a univariate linear regression model. Communications in Statistics: Theory and Methods, 42(7):1145–1152, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2017:SLLa

[CHZ17] Zengjing Chen, Cheng Hu, and Gaofeng Zong. Strong laws of large numbers for sub-linear expectation without independence. Communications in Statistics: Theory and Methods, 46(15): 7529–7545, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1154157.

Ciuperca:2017:ECP

[Ciu17] Gabriela Ciuperca. Estimation in a change-point non linear quantile model. Communications in Statistics: Theory and Methods, 46(12):6017–6034, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1116576.

Chen:2015:USF

[CJ15] Yibing Oliver Chen and Mike Jacroux. On the use of semifolding in regular blocked two-level factorial designs. Communications in Statistics: Theory and Methods, 44(12):2473–2506, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cho:2017:CDM

[CJ17] Sungwhan Cho and Jin Jiang. Change detection in the mean of a white Gaussian process by the backward standardized sum. Communications in Statistics: Theory and Methods, 46(17): 8401–8418, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1179761.

Cho:2019:IIB

[CJ19] Sang-Hoon Cho and Richard A. Johnson. Inference for Izawa’s bivariate gamma distribution. Communications in Statistics: Theory and Methods, 48(17):4350–4366, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1494288.

Chaturvedi:2010:EMD

[CK10] Anoop Chaturvedi and Suchita Kesarwani. Effect of misspecifying the disturbance covariance matrix on a family of shrinkage estimators. Communications in Statistics: Theory and Methods, 40(1):53–67, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chalikias:2017:OTT

[CK17a] Miltiadis Chalikias and Stratis Kounias. Optimal two treatment repeated measurement designs for three periods. Communications in Statistics: Theory and Methods, 46(1):200–209, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.988264.

Chao:2017:NTS

[CK17b] Cao Chao and Masahito Kobayashi. A new test for single against competing risks models in duration analysis. Communications in Statistics: Theory and Methods, 46(21):10672–10684, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1242742.

Chaturvedi:2017:ETP

[CK17c] Ajit Chaturvedi and Taruna Kumari. Estimation and testing procedures for the reliability functions of a general class of distributions. Communications in Statistics: Theory and Methods, 46(22):11370–11382, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1267765.

Chaturvedi:2019:ECS

[CK19] Ajit Chaturvedi and Taruna Kumari. Estimation and comparison of the stress-strength models with more than two states under Weibull distribution and type II censoring scheme. Communications in Statistics: Theory and Methods, 48(3):537–548, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1414264.

Clare:2014:EMC

[CKD14] John F. Clare, Annette Koo, and Robert B. Davies. Evaluation of measurement comparisons using generalized least squares: the role of participants’ estimates of systematic error. Communications in Statistics: Theory and Methods, 43(20):4297–4307, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chong:2017:GRR

[CKLC17] Zhi Lin Chong, Michael B. C. Khoo, Ming Ha Lee, and Chung-Ho Chen. Group runs revised m-of-k runs rule control chart. Communications in Statistics: Theory and Methods, 46(14): 6916–6935, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1140779.

Chavent:2011:SIR

[CKLS11] Marie Chavent, Vanessa Kuentz, Benoit Liquet, and Jérôme Saracco. A sliced inverse regression approach for a stratified population. Communications in Statistics: Theory and Methods, 40(21):3857–3878, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2017:CCW

[CKS17] Pingyan Chen, Ningning Kong, and Soo Hak Sung. Complete convergence for weighted sums of i.i.d. random variables with applications in regression estimation and EV model. Communications in Statistics: Theory and Methods, 46(7):3599–3613, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1066817.

Chernoyarov:2018:MCP

[CKT18] O. V. Chernoyarov, Yu. A. Kutoyants, and A. Top. On multiple change-point estimation for Poisson process. Communications in Statistics: Theory and Methods, 47(5):1215–1233, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1317810.

Chang:2010:BEB

[CL10a] Shui-Ching Chang and Tze Fen Li. Bayes empirical Bayes classification of components using masked data. Communications in Statistics: Theory and Methods, 39(13):2312–2320, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2010:AAR

[CL10b] Shin-Guang Chen and Yi-Kuei Lin. An approximate algorithm for the robust design in a stochastic-flow network. Communications in Statistics: Theory and Methods, 39(13):2440–2454, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2010:NEF

[CL10c] Xi Chen and Lu Lin. Nonparametric estimation for FBSDEs models with applications in finance. Communications in Statistics: Theory and Methods, 39(14):2492–2514, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calinski:2014:CIG

[CL14] Tadeusz Caliński and Agnieszka Lacka. On combining information in generally balanced nested block designs. Communications in Statistics: Theory and Methods, 43(5):954–974, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2015:NLH

[CL15a] Hao Chen and Min-Qian Liu. Nested latin hypercube designs with sliced structures. Communications in Statistics: Theory and Methods, 44(22):4721–4733, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2015:LIL

[CL15b] Yurong Chen and Luqin Liu. Laws of the iterated logarithm and a moderate deviation of MLE for the proportional hazards model with incomplete information. Communications in Statistics: Theory and Methods, 44(22):4696–4708, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christensen:2015:LFT

[CL15c] Ronald Christensen and Yong Lin. Lack-of-fit tests based on partial sums of residuals. Communications in Statistics: Theory and Methods, 44(13):2862–2880, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castellares:2016:MOE

[CL16] Fredy Castellares and Artur J. Lemonte. On the Marshall– Olkin extended distributions. Communications in Statistics: Theory and Methods, 45(15):4537–4555, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.922986.

Chen:2017:SLLb

[CL17a] Xiaoyan Chen and Fang Liu. Strong laws of large numbers for negatively dependent random variables under sublinear expectations. Communications in Statistics: Theory and Methods, 46(24):12387–12400, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1300274.

Chen:2017:MCR

[CL17b] Youyou Chen and Tailong Li. Moment convergence rates for the uniform empirical process and the uniform sample quantile process. Communications in Statistics: Theory and Methods, 46(18):9086–9091, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1205614.

Chen:2014:EBM

[CLC^{+}14] Yong Chen, Sheng Luo, Haitao Chu, Xiao Su, and Lei Nie. An empirical Bayes method for multivariate meta-analysis with an application in clinical trials. Communications in Statistics: Theory and Methods, 43(16):3536–3551, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:2016:EBS

[CLC^{+}16] Gauss M. Cordeiro, Maria Do Carmo S. Lima, Audrey H. M. A. Cysneiros, Marcelino A. R. Pascoa, Rodrigo R. Pescim, and Edwin M. M. Ortega. An extended birnbaum-Saunders distribution: Theory, estimation, and applications. Communications in Statistics: Theory and Methods, 45(8):2268–2297, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.879182.

Clement:2018:NCW

[Cle18] Etebong P. Clement. A note on calibration weightings for stratified double sampling with equal probability. Communications in Statistics: Theory and Methods, 47(12):2835–2847, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1342836.

Caro-Lopera:2015:GPT

[CLGFB15] Francisco J. Caro-Lopera, Graciela González-Farías, and N. Balakrishnan. The generalized Pascal triangle and the matrix variate Jensen-logistic distribution. Communications in Statistics: Theory and Methods, 44(13):2738–2752, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2017:ELR

[CLL17] Conghua Cheng, Yiming Liu, and Zhi Liu. Empirical likelihood ratio under infinite second moment. Communications in Statistics: Theory and Methods, 46(14):6909–6915, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1139135.

Chen:2016:CRA

[CLLW16] Zhiyong Chen, Tingting Liu, Jimin Ling, and Xuejun Wang. On the convergence rate for arrays of row-wise NOD random variables. Communications in Statistics: Theory and Methods, 45(5):1215–1223, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
786784.

Chan:2019:MIS

[CLMW19] Ka Ching Chan, Christopher T. Lenard, Terence M. Mills, and Ruth F. G. Williams. Measuring inequality in society. Communications in Statistics: Theory and Methods, 48(1):88–99, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1501486.

Coeurjolly:2014:VEF

[CLV14] Jean-François Coeurjolly, Kichun Lee, and Brani Vidakovic. Variance estimation for fractional Brownian motions with fixed Hurst parameters. Communications in Statistics: Theory and Methods, 43(8):1845–1858, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2017:ELC

[CLW17] Conghua Cheng, Zhi Liu, and Yi Wan. Empirical likelihood for compound Poisson processes under infinite second moment. Communications in Statistics: Theory and Methods, 46(17): 8618–8627, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1185122.

Chu:2014:BRL

[CLX14] Huang Chu and Zhang Li-Xin. Bahadur representation of linear kernel quantile estimator for stationary processes. Communications in Statistics: Theory and Methods, 43(22):4669–4678, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2016:GEP

[CLZ16] Baicheng Chen, Hua Liang, and Yong Zhou. GMM estimation in partial linear models with endogenous covariates causing an over-identified problem. Communications in Statistics: Theory and Methods, 45(11):3168–3184, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901363.

Chen:2018:ISE

[CLZ18] Shuai Chen, Wenbin Lu, and Hongwei Zhao. An improved survival estimator for censored medical costs with a kernel approach. Communications in Statistics: Theory and Methods, 47(23):5702–5716, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1400059.

Cui:2019:MER

[CLZ19] Chang Cui, Tao Li, and Lei Zhang. Mean estimate in ranked set sampling using a length-biased concomitant variable. Communications in Statistics: Theory and Methods, 48(12):2917–2931, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1473594.

Christoph:2012:APG

[CM12] Gerd Christoph and Sergey V. Malov. Asymptotic properties of generalized multivariate rank statistics. Communications in Statistics: Theory and Methods, 41(12):2133–2159, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chowdhury:2013:ETI

[CM13a] Shovan Chowdhury and S. P. Mukherjee. Estimation of traffic intensity based on queue length in a single M/M/1 queue. Communications in Statistics: Theory and Methods, 42(13): 2376–2390, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coelho:2013:MSB

[CM13b] Carlos A. Coelho and Filipe J. Marques. The multi-sample block-scalar sphericity test: Exact and near-exact distributions for its likelihood ratio test statistic. Communications in Statistics: Theory and Methods, 42(7):1153–1175, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cha:2016:DMP

[CM16a] Ji Hwan Cha and Jie Mi. Dynamic mixing probability measures of mixtures. Communications in Statistics: Theory and Methods, 45(16):4824–4839, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
930913.

Chatterjee:2016:IEO

[CM16b] Kiranmoy Chatterjee and Diganta Mukherjee. An improved estimator of omission rate for census count: With particular reference to India. Communications in Statistics: Theory and Methods, 45(4):1047–1062, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.854911.

Chinipardaz:2016:IVB

[CM16c] Rahim Chinipardaz and Khadijeh Mehri. Irreconcilability of P -value and Bayesian measure in two-sided hypothesis: Pareto distribution with the presence of nuisance parameter. Communications in Statistics: Theory and Methods, 45(8):2438–2449, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.881495.

Chrzanowski:2016:CNP

[CM16d] Michal Chrzanowski and Ryszard Magiera. Consistency of non parametric estimators of the area under the ROC curve. Communications in Statistics: Theory and Methods, 45(1):132–141, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coolen-Maturi:2017:TGR

[CM17] Tahani Coolen-Maturi. Three-group ROC predictive analysis for ordinal outcomes. Communications in Statistics: Theory and Methods, 46(19):9476–9493, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1212074.

Cockriel:2018:TMG

[CM18] William M. Cockriel and James B. McDonald. Two multivariate generalized beta families. Communications in Statistics: Theory and Methods, 47(23):5688–5701, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1400058.

Chaturvedi:2019:SER

[CM19] Ajit Chaturvedi and Ananya Malhotra. Shrinkage estimators of the reliability characteristics of a family of lifetime distributions based on records. Communications in Statistics: Theory and Methods, 48(21):5245–5264, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1510001.

Campos:2010:IMD

[CMCCO10] D. Campos, C. E. Martínez, A. Contreras-Cristán, and F. O’Reilly. Inferences for mixtures of distributions for centrally censored data with partial identification. Communications in Statistics: Theory and Methods, 39(12):2241–2263, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coolen-Maturi:2012:NPM

[CMCSC12] Tahani Coolen-Maturi, Pauline Coolen-Schrijner, and Frank P. A. Coolen. Nonparametric predictive multiple comparisons of lifetime data. Communications in Statistics: Theory and Methods, 41(22):4164–4181, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Canhanga:2018:ANS

[CMN^{+}18] Betuel Canhanga, Anatoliy Malyarenko, Ying Ni, Milica Rancić, and Sergei Silvestrov. Analytical and numerical studies on the second-order asymptotic expansion method for European option pricing under two-factor stochastic volatilities. Communications in Statistics: Theory and Methods, 47(6): 1328–1349, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1318923.

Conti:2017:HFI

[CMS17] Pier Luigi Conti, Daniela Marella, and Mauro Scanu. How far from identifiability? a systematic overview of the statistical matching problem in a non parametric framework. Communications in Statistics: Theory and Methods, 46(2):967–994, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1010005.

Chuntee:2015:BNA

[CN15] Wichairat Chuntee and Kritsana Neammanee. Bounds on normal approximations for the number of descents and inversions. Communications in Statistics: Theory and Methods, 44(11): 2310–2329, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2017:ECP

[CN17a] Fuqi Chen and Sévérien Nkurunziza. On estimation of the change points in multivariate regression models with structural changes. Communications in Statistics: Theory and Methods, 46(14):7157–7173, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1143510.

Chesneau:2017:PMS

[CN17b] Christophe Chesneau and Fabien Navarro. On the pointwise mean squared error of a multidimensional term-by-term thresholding wavelet estimator. Communications in Statistics: Theory and Methods, 46(11):5643–5655, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1107587.

Chuntee:2017:EBN

[CN17c] Wichairat Chuntee and Kritsana Neammanee. Exponential bounds for normal approximation of the number of descents and inversions. Communications in Statistics: Theory and Methods, 46(3):1218–1229, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1014109.

Castillo:2014:TAS

[CNC14] Enrique Castillo, MarÍa Nogal, and Aida Calviño. Two applications of statistics to traffic models. Communications in Statistics: Theory and Methods, 43(10–12):2118–2134, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2017:JEL

[CNG17] Ying-Ju Chen, Wei Ning, and Arjun K. Gupta. Jackknife empirical likelihood test for mean residual life functions. Communications in Statistics: Theory and Methods, 46(7):3111–3122, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1054943.

Cai:2019:SOP

[CNL19] Nanlian Cai, Wenqing Ni, and Chen Li. Some ordering properties of series and parallel systems with dependent component lifetimes. Communications in Statistics: Theory and Methods, 48(19):4764–4779, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1478099.

Chan:2014:P

[CNT14] P. S. Chan, H. K. T. Ng, and M. L. Tang. Preface. Communications in Statistics: Theory and Methods, 43(10–12):1991, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2010:ORM

[CO10] Jing Chang and David J. Olive. OLS for 1D regression models. Communications in Statistics: Theory and Methods, 39(10): 1869–1882, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Casanova:2016:BSP

[CO16a] María Paz Casanova and Yasna Orellana. A Bayesian semiparametric approach to the ordinal calibration problem. Communications in Statistics: Theory and Methods, 45(22):6596– 6610, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.963617.

Chakraborty:2016:CPT

[CO16b] S. Chakraborty and S. H. Ong. A COM–Poisson-type generalization of the negative binomial distribution. Communications in Statistics: Theory and Methods, 45(14):4117–4135, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.917184.

Coin:2017:GFT

[Coi17] Daniele Coin. A goodness-of-fit test for generalized error distribution. Communications in Statistics: Theory and Methods, 46(23):11485–11499, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1271426.

Castagliola:2017:DSC

[COK17] Philippe Castagliola, Pedro Carlos Oprime, and Michael B. C. Khoo. The double sampling S^{2} chart with estimated process variance. Communications in Statistics: Theory and Methods, 46(7):3556–3573, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1066814.

Cordeiro:2015:PGL

[COL15] Gauss M. Cordeiro, Edwin Ortega, and Artur Lemonte. The Poisson generalized linear failure rate model. Communications in Statistics: Theory and Methods, 44(10):2037–2058, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Congdon:2010:ADA

[Con10] Peter Congdon. Assessing differential area mortality trends via Bayesian random effects. Communications in Statistics: Theory and Methods, 39(12):2205–2230, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Congdon:2012:SHF

[Con12] Peter Congdon. Spatial health factors with selection among multiple causes: Lung cancer in U.S. counties. Communications in Statistics: Theory and Methods, 41(11):1933–1953, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Congdon:2016:LJC

[Con16] Peter Congdon. A local join counts methodology for spatial clustering in disease from relative risk models. Communications in Statistics: Theory and Methods, 45(10):3059–3075, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.894071.

Cooray:2012:AGC

[Coo12] Kahadawala Cooray. Analyzing grouped, censored, and truncated data using the odd Weibull family. Communications in Statistics: Theory and Methods, 41(15):2661–2680, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cooray:2013:ESC

[Coo13] Kahadawala Cooray. Exponentiated sinh Cauchy distribution with applications. Communications in Statistics: Theory and Methods, 42(21):3838–3852, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cooray:2019:NEF

[Coo19] Kahadawala Cooray. A new extension of the FGM copula for negative association. Communications in Statistics: Theory and Methods, 48(8):1902–1919, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1440312.

Costa:2018:ATA

[Cos18] Liborio I. Costa. An accurate, tractable, and analytically integrable polynomial expansion of the skewed Student’s tdistribution. Communications in Statistics: Theory and Methods, 47(1):239–246, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1301478.

Cox:2015:EAR

[Cox15] M. A. A. Cox. Exact average run lengths for monitoring Poisson counts. Communications in Statistics: Theory and Methods, 44(22):4757–4771, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cox:2019:TFP

[Cox19] M. A. A. Cox. A tool to facilitate the production of control charts within a spreadsheet. Communications in Statistics: Theory and Methods, 48(20):5050–5057, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1504969.

Ciginas:2014:NUB

[CP14] Andrius Ciginas and Dalius Pumputis. A note on the upper bound to variance of the sample extreme from a finite population. Communications in Statistics: Theory and Methods, 43(8):1793–1799, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2019:RMI

[CPW19] Fenge Chen, Xingchun Peng, and Wenyuan Wang. Risk minimization for an insurer with investment and reinsurance via g-expectation. Communications in Statistics: Theory and Methods, 48(20):5012–5035, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1504077.

Chen:2017:PHT

[CQ17] Xu Chen and Wang Qi. The parametric hypothesis test of multiple uniform populations. Communications in Statistics: Theory and Methods, 46(6):2681–2685, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1042274.

Costa:2013:EDC

[CR13] Antonio F. B. Costa and M. A. Rahim. Economic design of and R charts under Weibull shock models. Communications in Statistics: Theory and Methods, 42(21):3902–3925, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crepey:2014:SRP

[CR14] Stéphane Crépey and Abdallh Rahal. Simulation/regression pricing schemes for CVA computations on CDO tranches. Communications in Statistics: Theory and Methods, 43(7):1390– 1408, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chandrasekar:2016:EEP

[CR16] B. Chandrasekar and S. Amala Revathy. Equivariant estimation for the parameters of location-scale multivariate exponential models and its application in reliability analysis. Communications in Statistics: Theory and Methods, 45(18):5550–5559, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.948195.

Cramer:2014:EVA

[Cra14] Erhard Cramer. Extreme value analysis for progressively Type-II censored order statistics. Communications in Statistics: Theory and Methods, 43(10–12):2135–2155, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chandra:2013:GPD

[CRG13] Nimai Kumar Chandra, Dilip Roy, and Tirthankar Ghosh. A generalized Poisson distribution. Communications in Statistics: Theory and Methods, 42(15):2786–2797, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calsavara:2017:FMP

[CRTdC17] Vinicius Fernando Calsavara, Agatha Sacramento Rodrigues, Vera Lúcia Damasceno Tomazella, and Mário de Castro. Frailty models power variance function with cure fraction and latent risk factors negative binomial. Communications in Statistics: Theory and Methods, 46(19):9763–9776, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1218029.

Christen:2011:ASD

[CS11] J. Andrés Christen and Bruno Sansó. Advances in the sequential design of computer experiments based on active learning. Communications in Statistics: Theory and Methods, 40(24): 4467–4483, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2012:ESS

[CS12] Yi-Chun Chen and Wheyming Tina Song. The effects of sample sizes in Phases I and II on p control chart performance. Communications in Statistics: Theory and Methods, 41(22): 4047–4068, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaturvedi:2014:BER

[CS14a] Anoop Chaturvedi and Shalabh. Bayesian estimation of regression coefficients under extended balanced loss function. Communications in Statistics: Theory and Methods, 43(20):4253– 4264, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chesneau:2014:NWR

[CS14b] Christophe Chesneau and Esmaeil Shirazi. Nonparametric wavelet regression based on biased data. Communications in Statistics: Theory and Methods, 43(13):2642–2658, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:2015:MNL

[CS15a] Yogendra P. Chaubey and Esmaeil Shirazi. On MISE of a non linear wavelet estimator of the regression function based on biased data under strong mixing. Communications in Statistics: Theory and Methods, 44(5):885–899, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:2015:WED

[CS15b] Yogendra P. Chaubey and Esmaeil Shirazi. On wavelet estimation of the derivatives of a density based on biased data. Communications in Statistics: Theory and Methods, 44(21): 4491–4506, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cao:2016:LAE

[CS16a] Ming-Xiang Cao and Guang-Jun Shena. Linearly admissible estimators of mean vector with respect to balanced loss function in multivariate statistics. Communications in Statistics: Theory and Methods, 45(4):1063–1069, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.854912.

Chakraborty:2016:RME

[CS16b] Hrishikesh Chakraborty and Pranab K. Sen. Resampling method to estimate intra-cluster correlation for clustered binary data. Communications in Statistics: Theory and Methods, 45(8):2368–2377, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
870202.

Chaturvedi:2016:BAL

[CS16c] Anoop Chaturvedi and Arvind Shrivastava. Bayesian analysis of a linear model involving structural changes in either regression parameters or disturbances precision. Communications in Statistics: Theory and Methods, 45(2):307–320, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charyulu:2017:NRD

[CS17] N. Ch. Bhatra Charyulu and Sk. Ameen Saheb. Note on reduction of dimensionality for second-order response surface design model. Communications in Statistics: Theory and Methods, 46(7):3520–3525, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1065332.

Chandra:2018:SAE

[CS18] Hukum Chandra and Nicola Salvati. Small area estimation of proportions under a spatial dependent aggregated level random effects model. Communications in Statistics: Theory and Methods, 47(5):1234–1255, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1317806.

Chang:2019:NTS

[CS19a] Yuan-Tsung Chang and Nobuo Shinozaki. New types of shrinkage estimators of Poisson means under the normalized squared error loss. Communications in Statistics: Theory and Methods, 48(5):1108–1122, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1423699.

Chen:2019:WLL

[CS19b] Pingyan Chen and Soo Hak Sung. Weak laws of large numbers for nonnegative independent random variables without finite means. Communications in Statistics: Theory and Methods, 48(21):5401–5413, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1513145.

Cheng:2019:PDC

[CS19c] Zailei Cheng and Youngsoo Seol. Precise deviations for Cox processes with a shot noise intensity. Communications in Statistics: Theory and Methods, 48(23):5850–5861, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1522351.

Chatterjee:2019:EPM

[CSB19] A. Chatterjee, G. N. Singh, and A. Bandyopadhyay. Estimation of population mean in successive sampling under superpopulation model in presence of random non response situations. Communications in Statistics: Theory and Methods, 48(15):3850–3863, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481978.

Chang:2013:BOR

[CSC13] Chin-Chih Chang, Shey-Huei Sheu, and Yen-Luan Chen. A bivariate optimal replacement policy for a system with agedependent minimal repair and cumulative repair-cost limit. Communications in Statistics: Theory and Methods, 42(22): 4108–4126, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2017:ACM

[CSC17] Yu-Mei Chang, Pao-Sheng Shen, and Chun-Shu Chen. Adaptive-Cox model averaging for right-censored data. Communications in Statistics: Theory and Methods, 46(19):9364– 9376, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1208237.

Chen:2012:ELT

[CSF12] Yanhong Chen, Lixin Song, and Jinghai Feng. Empirical likelihood for threshold autoregressive models. Communications in Statistics: Theory and Methods, 41(6):970–982, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cai:2017:AID

[CSL17] Jing Cai, Yimin Shi, and Bin Liu. Analysis of incomplete data in the presence of dependent competing risks from Marshall–Olkin bivariate Weibull distribution under progressive hybrid censoring. Communications in Statistics: Theory and Methods, 46(13):6497–6511, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1129420.

Chalaki:2016:CSE

[CSM16] Kamyar Chalaki, Abbas Saghaei, and M. B. Moghadam. A comparison study of effectiveness and robustness of robust economic designs of T^{2} chart using genetic algorithm. Communications in Statistics: Theory and Methods, 45(11):3383–3396, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.904352.

Chaidee:2016:PAT

[CSS16] Nattakarn Chaidee, Tippawan Santiwipanont, and Songkiat Sumetkijakan. Patched approximations and their convergence. Communications in Statistics: Theory and Methods, 45(9): 2654–2664, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.887112.

Cao:2019:SIA

[CSY19] Yongxiu Cao, Yueyong Shi, and Jichang Yu. Statistical inference for the accelerated failure time model under two-stage generalized case-cohort design. Communications in Statistics: Theory and Methods, 48(24):6063–6079, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1528363.

Chen:2010:PLM

[CT10] Guo Chen and Hiroki Tsurumi. Probit and logit model selection. Communications in Statistics: Theory and Methods, 40 (1):159–175, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2011:OTL

[CT11] Fu-Kai Chang and Chao-Ping Ting. Optimal two-level fractional factorial designs for location main effects with dispersion factors. Communications in Statistics: Theory and Methods, 40(11):2035–2043, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chacon:2013:DBC

[CT13] J. E. Chacón and C. Tenreiro. Data-based choice of the number of pilot stages for plug-in bandwidth selection. Communications in Statistics: Theory and Methods, 42(12):2200–2214, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cramer:2014:CSM

[CT14] Erhard Cramer and Miriam Tamm. On a correction of the scale MLE for a two-parameter exponential distribution under progressive Type-I censoring. Communications in Statistics: Theory and Methods, 43(20):4401–4414, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2016:QRM

[CT16] Jung-Yu Cheng and Shinn-Jia Tzeng. Quantile regression methods for censored gap time data. Communications in Statistics: Theory and Methods, 45(17):5154–5165, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.941491.

Conceicao:2017:BZM

[CTAL17] Katiane S. Conceição, Vera Tomazella, Marinho G. Andrade, and Francisco Louzada. Biparametric zero-modified power series distributions: Bayesian analysis under a reference prior approach. Communications in Statistics: Theory and Methods, 46(21):10518–10536, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1236960.

Chen:2011:MVC

[CTQ11] Zhanshou Chen, Zheng Tian, and Ruibing Qin. Monitoring variance change in infinite order moving average processes and nonstationary autoregressive processes. Communications in Statistics: Theory and Methods, 40(7):1254–1270, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2015:SHC

[CTT15] Yanliang Chen, Man-Lai Tang, and Maozai Tian. Semiparametric hierarchical composite quantile regression. Communications in Statistics: Theory and Methods, 44(5):996–1012, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:2014:OSM

[cTtL14] Chih chun Tsai and Chien tai Lin. Optimal selection of the most reliable design based on gamma degradation processes. Communications in Statistics: Theory and Methods, 43(10–12): 2419–2428, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2010:IDS

[CTW10] Xue-Dong Chen, Nian-Sheng Tang, and Xue-Ren Wang. Influence diagnostics of semiparametric nonlinear reproductive dispersion models. Communications in Statistics: Theory and Methods, 39(17):3021–3040, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cucala:2016:MWS

[Cuc16] Lionel Cucala. A Mann–Whitney scan statistic for continuous data. Communications in Statistics: Theory and Methods, 45 (2):321–329, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cui:2012:AFC

[Cui12] Lixin Cui. Applying fuzzy comprehensive evaluation method to evaluate quality in crisis and emergency management. Communications in Statistics: Theory and Methods, 41(21):3942–3959, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cuadras:2012:DBM

[CVC^{+}12] Carles M. Cuadras, Silvia Valero, Daniel Cuadras, Philippe Salembier, and Jocelyn Chanussot. Distance-based measures of association with applications in relating hyperspectral images. Communications in Statistics: Theory and Methods, 41(13–14): 2342–2355, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Corujo:2019:SCT

[CVL19] Josué M. Corujo, José E. Valdés, and Juan C. Laria. Stochastic comparisons of two-unit repairable systems. Communications in Statistics: Theory and Methods, 48(23):5820–5838, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1522349.

Chen:2013:SEB

[CW13] Ling Chen and Laisheng Wei. The superiorities of empirical Bayes estimation of variance components in random effects model. Communications in Statistics: Theory and Methods, 42(22):4017–4033, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chandrasekhar:2014:MPE

[CW14] Rameela Chandrasekhar and Gregory E. Wilding. A modification of a percentile estimation procedure based on generalized Pólya urns. Communications in Statistics: Theory and Methods, 43(14):2951–2957, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2015:SPM

[CW15a] Xiaolin Chen and Qihua Wang. Semiparametric proportional mean residual life model with censoring indicators missing at random. Communications in Statistics: Theory and Methods, 44(24):5161–5188, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chesneau:2015:ECD

[CW15b] Christophe Chesneau and Thomas Willer. Estimation of a cumulative distribution function under interval censoring “case 1” via warped wavelets. Communications in Statistics: Theory and Methods, 44(17):3680–3702, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2016:ELI

[CW16] Wei Chen and Dehui Wang. Empirical likelihood inference for a semiparametric hazards regression model. Communications in Statistics: Theory and Methods, 45(11):3236–3248, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.901368.

Chang:2018:PST

[CW18] Xinfeng Chang and Jibo Wu. Performance of the stein-type two-parameter estimator in multiple linear regression model. Communications in Statistics: Theory and Methods, 47(8): 1935–1952, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1332220.

Chang:2019:PTA

[CW19a] Xinfeng Chang and Hexiang Wang. Preliminary test almost unbiased two-parameter estimators with student’s t errors and conflicting test statistics. Communications in Statistics: Theory and Methods, 48(18):4449–4473, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1502782.

Chen:2019:ELT

[CW19b] Kun Chen and Man Wang. On empirical likelihood test for predictability. Communications in Statistics: Theory and Methods, 48(10):2499–2508, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1465092.

Cui:2019:CML

[CW19c] Yunwei Cui and Xiaoyin Wang. Conditional maximum likelihood estimation in negative binomial INGARCH processes with known number of successes when the true parameter is at the boundary of the parameter space. Communications in Statistics: Theory and Methods, 48(13):3388–3401, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1476710.

Cui:2016:ULA

[CWW16] Zhaolei Cui, Yuebao Wang, and Kaiyong Wang. The uniform local asymptotics for a Lévy process and its overshoot and undershoot. Communications in Statistics: Theory and Methods, 45(4):1156–1181, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
861492.

Cao:2017:LAE

[CX17] Mingxiang Cao and Xingzhong Xu. Linearly admissible estimators of the common mean parameter under balanced loss function. Communications in Statistics: Theory and Methods, 46(17):8307–8317, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
786783.

Chen:2019:PDD

[CXC19] Jikun Chen, Hui Xu, and Fengyang Cheng. The product distribution of dependent random variables with applications to a discrete-time risk model. Communications in Statistics: Theory and Methods, 48(13):3325–3340, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1476705.

Chen:2011:EGC

[CXS11] Min Chen, Yang Xie, and Michael Story. An exponentialgamma convolution model for background correction of Illumina BeadArray data. Communications in Statistics: Theory and Methods, 40(17):3055–3069, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2010:EDV

[CY10] F. L. Chen and C. H. Yeh. Economic design of VSI x-bar control charts for non normally distributed data under Gamma(λ, 2) failure models. Communications in Statistics: Theory and Methods, 39(10):1743–1760, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cui:2012:P

[CY12] Lirong Cui and Guoxiao Yang. Preface. Communications in Statistics: Theory and Methods, 41(21):3813, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:2014:PTE

[CY14] Xinfeng Chang and Hu Yang. Preliminary test estimators induced by three large sample tests for stochastic constraints in a regression model with multivariate Student-t error. Communications in Statistics: Theory and Methods, 43(17):3629–3640, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cronie:2016:DOI

[CY16] Ottmar Cronie and Jun Yu. The discretely observed immigration–death process: Likelihood inference and spatiotemporal applications. Communications in Statistics: Theory and Methods, 45(18):5279–5298, 2016. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
942433.

Cang:2017:EBA

[CY17a] Yuquan Cang and Yang Yang. On extremal behavior of aggregation of largest claims. Communications in Statistics: Theory and Methods, 46(2):917–926, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1006787.

Cao:2017:OGC

[CY17b] Yongxiu Cao and Jichang Yu. Optimal generalized case-cohort sampling design under the additive hazard model. Communications in Statistics: Theory and Methods, 46(9):4484–4493, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1085563.

Cagcag:2015:NRR

[CYE15] Ozge Cagcag, Ufuk Yolcu, and Erol Egrioglu. A new robust regression method based on particle swarm optimization. Communications in Statistics: Theory and Methods, 44(6):1270– 1280, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cancho:2018:EID

[CYF^{+}18] Vicente G. Cancho, Bao Yiqi, Jose A. Fiorucci, Gladys D. C. Barriga, and Dipak K. Dey. Estimation and influence diagnostics for zero-inflated hyper-Poisson regression model: full Bayesian analysis. Communications in Statistics: Theory and Methods, 47(11):2741–2759, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1342839.

Cheng:2013:ODR

[CYL13] Jing Cheng, Rong-Xian Yue, and Xin Liu. Optimal designs for random coefficient regression models with heteroscedastic errors. Communications in Statistics: Theory and Methods, 42(15):2798–2809, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2014:EGC

[CYY14] Fuxia Cheng, Jigao Yan, and Lijian Yang. Extended Glivenko– Cantelli theorem in nonparametric regression. Communications in Statistics: Theory and Methods, 43(17):3720–3725, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:2015:ECF

[CZ15a] Yogendra P. Chaubey and Rui Zhang. An extension of Chen’s family of survival distributions with bathtub shape or increasing hazard rate function. Communications in Statistics: Theory and Methods, 44(19):4049–4064, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2015:LLE

[CZ15b] Yingyu Chen and Lixin Zhang. Local linear estimation of second-order jump-diffusion model. Communications in Statistics: Theory and Methods, 44(18):3903–3920, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:2015:QRM

[CZ15c] Yebin Cheng and Dawit Zerom. A quantile regression model for time-series data in the presence of additive components. Communications in Statistics: Theory and Methods, 44(20): 4354–4379, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:2019:RTL

[CZ19a] Chao Chen and Run-Chu Zhang. Robust two-level regular fractional factorial designs. Communications in Statistics: Theory and Methods, 48(19):4794–4803, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1490431.

Chen:2019:CAD

[CZ19b] Chen Chen and Panpan Zhang. Characterizations of asymptotic distributions of continuous-time Pólya processes. Communications in Statistics: Theory and Methods, 48(21):5308–5321, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1510005.

Chen:2019:NTS

[CZ19c] Ziqi Chen and Yan Zhou. New test statistics for hypothesis testing of parameters in conditional moment restriction models. Communications in Statistics: Theory and Methods, 48(10): 2521–2528, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1472775.

Cai:2019:CLT

[CZCM19] Jingwei Cai, Quanxin Zhu, Ping Chen, and Xia Mei. Central limit theorems of range-based estimators for diffusion models. Communications in Statistics: Theory and Methods, 48(24): 5969–5984, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1523432.

Chen:2017:CCM

[CZM17] Ying-Xia Chen, Shui-Li Zhang, and Fu-Qiang Ma. On the complete convergence for martingale difference sequence. Communications in Statistics: Theory and Methods, 46(15):7603–7611, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1157188.

Cheng:2016:SCW

[CZW16] Dongya Cheng, Wenhai Zhang, and Yuebao Wang. On the strong convergence of weighted sums of widely dependent random variables. Communications in Statistics: Theory and Methods, 45(21):6447–6460, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
922988.

Zhou:2018:WET

[cZzXgL18] Xing cai Zhou, Ying zhi Xu, and Jin guan Lin. Wavelet estimation in time-varying coefficient time series models with measurement errors. Communications in Statistics: Theory and Methods, 47(10):2504–2519, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1339801.

Chen:2018:ASS

[CZZZ18] Jianbin Chen, Yiying Zhang, Peng Zhao, and Shan Zhou. Allocation strategies of standby redundancies in series/parallel system. Communications in Statistics: Theory and Methods, 47 (3):708–724, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1313984.

Drinkuth:2016:AIG

[DA16] Carsten Drinkuth and Matthias Arnold. Asymptotics of improved generalized moment estimators for spatial autoregressive error models. Communications in Statistics: Theory and Methods, 45(7):1939–1952, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
870203.

DAmbra:2017:CCA

[D’A17a] Antonello D’Ambra. Cumulative correspondence analysis using orthogonal polynomials. Communications in Statistics: Theory and Methods, 46(6):2942–2954, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1053938.

Dogru:2017:PEM

[DA17b] Fatma Zehra Dogru and Olcay Arslan. Parameter estimation for mixtures of skew Laplace normal distributions and application in mixture regression modeling. Communications in Statistics: Theory and Methods, 46(21):10879–10896, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1252400.

DAmbra:2017:DME

[DAC17] Antonello D’Ambra, Pietro Amenta, and Anna Crisci. Decomposition of the main effects and interaction term by using orthogonal polynomials in multiple non symmetrical correspondence analysis. Communications in Statistics: Theory and Methods, 46(20):10179–10188, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1231817.

deAndrade:2018:SRM

[dAdLA18] Bernardo Borba de Andrade and Joanlise Marco de Leon Andrade. Some results for maximum likelihood estimation of adjusted relative risks. Communications in Statistics: Theory and Methods, 47(23):5750–5769, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1402045.

Dai:2015:AMR

[Dai15] Hongshuai Dai. Approximation to multifractional Riemann– Liouville Brownian sheet. Communications in Statistics: Theory and Methods, 44(7):1399–1410, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dai:2016:RWS

[Dai16] Hongshuai Dai. Random walks and subfractional Brownian motion. Communications in Statistics: Theory and Methods, 45(10):2834–2841, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
889163.

Dimitrakopoulou:2012:BEE

[DAL12] Theodora Dimitrakopoulou, Konstantinos Adamidis, and Sotirios Loukas. Bivariate extended exponential-geometric distributions. Communications in Statistics: Theory and Methods, 41(7):1129–1150, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dang:2018:SLL

[Dan18] Hui Dang. The strong law of large numbers for non homogeneous M-bifurcating Markov chains indexed by a Mbranch Cayley tree. Communications in Statistics: Theory and Methods, 47(9):2110–2125, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1335417.

Dasgupta:2016:HDP

[Das16a] Shibasish Dasgupta. High-dimensional posterior consistency of the Bayesian lasso. Communications in Statistics: Theory and Methods, 45(22):6700–6708, 2016. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
966840.

Dastbaravarde:2016:STR

[DAS16b] Ali Dastbaravarde, Nasser Reza Arghami, and Majid Sarmad. Some theoretical results concerning non parametric estimation by using a judgment poststratification sample. Communications in Statistics: Theory and Methods, 45(8):2181–2203, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.878355.

Das:2017:GGD

[Das17a] Shubhabrata Das. On generalized geometric distributions and improved estimation of batting average in cricket. Communications in Statistics: Theory and Methods, 46(6):2736–2750, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1048888.

Das:2017:WGE

[Das17b] Suchismita Das. On weighted generalized entropy. Communications in Statistics: Theory and Methods, 46(12):5707–5727, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.960583.

Dahmani:2013:EIC

[DB13] Abdelnasser Dahmani and Karima Belaide. Exponential inequalities in calibration problems with Gaussians errors. Communications in Statistics: Theory and Methods, 42(19):3596– 3607, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahmani:2015:EIS

[DB15] Abdelnasser Dahmani and Karima Belaide. Exponential inequalities in stochastic inverse problems using an iterative method. Communications in Statistics: Theory and Methods, 44(11):2385–2397, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dogru:2018:DRE

[DBA18] Fatma Zehra Dogru, Y. Murat Bulut, and Olcay Arslan. Doubly reweighted estimators for the parameters of the multivariate t-distribution. Communications in Statistics: Theory and Methods, 47(19):4751–4771, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1445861.

DAmbra:2014:CCA

[DBC14] Luigi D’Ambra, Eric J. Beh, and Ida Camminatiello. Cumulative correspondence analysis of two-way ordinal contingency tables. Communications in Statistics: Theory and Methods, 43(6):1099–1113, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

De:2016:CLN

[DBC16] Tuli De, Didier Fraix Burnet, and Asis Kumar Chattopadhyay. Clustering large number of extragalactic spectra of galaxies and quasars through canopies. Communications in Statistics: Theory and Methods, 45(9):2638–2653, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.848286.

Castro:2015:AFT

[dBCR15] Tomás del Barrio Castro and Andreu Sansó Rossello. On augmented franses tests for seasonal unit roots. Communications in Statistics: Theory and Methods, 44(24):5204–5212, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dortet-Bernadet:2017:MMD

[DBG17] Jean-Luc Dortet-Bernadet and Laurent Gardes. A mixture model for dimension reduction. Communications in Statistics: Theory and Methods, 46(21):10768–10787, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1248576.

deBrito:2016:MGM

[dBSCD16] Edleide de Brito, Giovana Oliveira Silva, Gauss M. Cordeiro, and Clarice Garcia B. Demétrio. The McDonald Gumbel model. Communications in Statistics: Theory and Methods, 45(11):3367–3382, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
904351.

Demattei:2010:MST

[DC10a] C. Demattei and L. Cucala. Multiple spatio-temporal cluster detection for case event data: an ordering-based approach. Communications in Statistics: Theory and Methods, 40(2):358– 372, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diawara:2010:MBE

[DC10b] Norou Diawara and Mark Carpenter. Mixture of bivariate exponential distributions. Communications in Statistics: Theory and Methods, 39(15):2711–2720, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Domma:2013:BDD

[DC13] Filippo Domma and Francesca Condino. The beta-Dagum distribution: Definition and properties. Communications in Statistics: Theory and Methods, 42(22):4070–4090, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DAmbra:2014:CED

[DC14] Antonello D’Ambra and Anna Crisci. The confidence ellipses in decomposition multiple non symmetric correspondence analysis. Communications in Statistics: Theory and Methods, 43(6): 1209–1221, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz:2017:MGC

[DC17] Walter Díaz and Carles M. Cuadras. On a multivariate generalization of the covariance. Communications in Statistics: Theory and Methods, 46(9):4660–4669, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1056368.

Amorim:2019:SES

[dCA19] Gustavo Guimarães de Castro Amorim. Semiparametric estimator for a secondary analysis with correlated outcomes. Communications in Statistics: Theory and Methods, 48(19):4703– 4711, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1146768.

DiasDeOliveira:2013:PLS

[DCG13] Maristela Dias De Oliveira, Enrico A. Colosimo, and Gustavo L. Gilardoni. Power law selection model for repairable systems. Communications in Statistics: Theory and Methods, 42(4):570– 578, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:2012:FSR

[DCL12] Xiaofang Dong, Lirong Cui, and Fangyu Liu. A further study on reliable life estimation under ranked set sampling. Communications in Statistics: Theory and Methods, 41(21):3888–3902, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deCarvalho:2013:ELE

[dCOSW13] Miguel de Carvalho, Boris Oumow, Johan Segers, and Michal Warchol. A Euclidean likelihood estimator for bivariate tail dependence. Communications in Statistics: Theory and Methods, 42(7):1176–1192, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Di:2017:THS

[DCZL17] Chong-Zhi Di, Kwun Chuen Gary Chan, Cheng Zheng, and Kung-Yee Liang. Testing homogeneity in semiparametric mixture case-control models. Communications in Statistics: Theory and Methods, 46(18):9092–9100, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1205612.

Dutta:2013:ODE

[DD13] Ganesh Dutta and Premadhis Das. Optimum design for estimation of regression parameters in multi-factor set-up. Communications in Statistics: Theory and Methods, 42(24):4431–4443, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diallo:2017:APM

[DDD17] Alpha Oumar Diallo, Aliou Diop, and Jean-François Dupuy. Asymptotic properties of the maximum-likelihood estimator in zero-inflated binomial regression. Communications in Statistics: Theory and Methods, 46(20):9930–9948, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1222437.

DAmico:2013:IEG

[DDGM13] Guglielmo D’Amico, Giuseppe Di Biase, Fulvio Gismondi, and Raimondo Manca. The input evaluation of generalized Bernoulli processes for salary lines construction by means of continuous time generalized non-homogeneous semi-Markov processes. Communications in Statistics: Theory and Methods, 42(16):2889–2901, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutta:2010:ODC

[DDM10] Ganesh Dutta, Premadhis Das, and Nripes K. Mandal. Doptimal designs for covariate parameters in block design set up. Communications in Statistics: Theory and Methods, 39 (19):3434–3443, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutta:2014:ODC

[DDM14] Ganesh Dutta, Premadhis Das, and Nripes K. Mandal. Doptimal designs for covariate models. Communications in Statistics: Theory and Methods, 43(1):165–174, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:2018:CLR

[DDP18] Ujjwal Das, Subhra Sankar Dhar, and Vivek Pradhan. Corrected likelihood-ratio tests in logistic regression using smallsample data. Communications in Statistics: Theory and Methods, 47(17):4272–4285, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1373815.

Du:2016:JME

[DDS16] Ting Du, Jieli Ding, and Liuquan Sun. Joint modeling and estimation for longitudinal data with informative observation and terminal event times. Communications in Statistics: Theory and Methods, 45(22):6521–6539, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.960589.

DeBaets:2011:CCG

[DDÚF11] Bernard De Baets, Hans De Meyer, and Manuel Úbeda-Flores. Constructing copulas with given diagonal and opposite diagonal sections. Communications in Statistics: Theory and Methods, 40(5):828–843, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLaHorra:2011:PFD

[De 11] Julián De La Horra. The positive false discovery quantile. Communications in Statistics: Theory and Methods, 40(15): 2629–2638, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:2017:CSA

[DE17a] Ujjwal Das and Nader Ebrahimi. Covariate selection for accelerated failure time data. Communications in Statistics: Theory and Methods, 46(8):4051–4064, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1078475.

DeKock:2017:BVS

[DE17b] Michiel B. De Kock and Hans C. Eggers. Bayesian variable selection with spherically symmetric priors. Communications in Statistics: Theory and Methods, 46(9):4250–4263, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1081945.

Debrabant:2012:PPG

[Deb12] Birgit Debrabant. Point processes with a generalized order statistic property. Communications in Statistics: Theory and Methods, 41(19):3449–3467, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demirtas:2014:GBU

[Dem14] Hakan Demirtas. Generating bivariate uniform data with a full range of correlations and connections to bivariate binary data. Communications in Statistics: Theory and Methods, 43(17): 3574–3579, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:2013:SRA

[Den13] Shirong Deng. Semiparametric regression analysis of panel count data with time-dependent covariates and informative observation and censoring times. Communications in Statistics: Theory and Methods, 42(22):4170–4183, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:2017:MBF

[Den17] Yihao Deng. Modeling binary familial data using Gaussian copula. Communications in Statistics: Theory and Methods, 46(20):10097–10102, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1228971.

DeRiggi:2019:CLT

[DeR19] Dennis DeRiggi. A central limit theorem for correlated variables with limited normal or gamma distributions. Communications in Statistics: Theory and Methods, 48(21):5213–5222, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1536212.

Donevska:2016:CCM

[DFH16] Sandra Donevska, Eva Fiserová, and Karel Hron. Calibration of compositional measurements. Communications in Statistics: Theory and Methods, 45(22):6773–6788, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.839039.

Durante:2010:DCC

[DFS10] Fabrizio Durante, Rachele Foschi, and Peter Sarkoci. Distorted copulas: Constructions and tail dependence. Communications in Statistics: Theory and Methods, 39(12):2288–2301, 2010.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Desmet:2011:CFU

[DG11] L. Desmet and I. Gijbels. Curve fitting under jump and peak irregularities using local linear regression. Communications in Statistics: Theory and Methods, 40(22):4001–4020, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DiLascio:2012:CMD

[DG12a] F. Marta L. Di Lascio and Simone Giannerini. Clustering microarray data: Theoretical and practical issues. Communications in Statistics: Theory and Methods, 41(16–17):3211–3232, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/03610926.2011.654036.

Diana:2012:FPV

[DG12b] Giancarlo Diana and Marco Giordan. Finite population variance estimation in presence of measurement errors. Communications in Statistics: Theory and Methods, 41(23):4302–4314, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DAmico:2013:SMD

[DGM13] Guglielmo D’Amico, Montserrat Guillen, and Raimondo Manca. Semi-Markov disability insurance models. Communications in Statistics: Theory and Methods, 42(16):2872–2888, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunder:2017:SSQ

[DGMC17] Emre Dünder, Serpil Gümüstekin, Naci Murat, and Mehmet Ali Cengiz. Subset selection in quantile regression analysis via alternative Bayesian information criteria and heuristic optimization. Communications in Statistics: Theory and Methods, 46(22):11091–11098, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1257718.

Diaz-Garcia:2011:MTP

[DGV11] José A. Díaz-García and J. Fernando Vera. Multivariate three-parameter log-elliptical distributions. Communications in Statistics: Theory and Methods, 40(11):1964–1976, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dai:2017:EUV

[DHK17] D. Dai, T. Holgersson, and P. Karlsson. Expected and unexpected values of individual Mahalanobis distances. Communications in Statistics: Theory and Methods, 46(18):8999–9006, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1200096.

Daryabari:2017:EME

[DHKM17] S. Abedin Daryabari, S. Mohammad Hashemian, Ali Keyvandarian, and Shekary A. Maryam. The effects of measurement error on the MAX EWMAMS control chart. Communications in Statistics: Theory and Methods, 46(12):5766–5778, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1112911.

DAmato:2016:MMM

[DHP^{+}16] Valeria D’Amato, Steven Haberman, Gabriella Piscopo, Maria Russolillo, and Lorenzo Trapani. Multiple mortality modeling in Poisson Lee–Carter framework. Communications in Statistics: Theory and Methods, 45(6):1723–1732, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.960580.

Durio:2010:CDR

[DI10] Alessandra Durio and Ennio Davide Isaia. Clusters detection in regression problems: a similarity test between estimated models. Communications in Statistics: Theory and Methods, 39(3): 508–516, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ding:2015:CSE

[Din15] Xue Ding. Convergence of sample eigenvectors of spiked population model. Communications in Statistics: Theory and Methods, 44(18):3825–3840, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Durante:2010:NCB

[DJ10] Fabrizio Durante and Piotr Jaworski. A new characterization of bivariate copulas. Communications in Statistics: Theory and Methods, 39(16):2901–2912, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dauxois:2015:TED

[DJ15] J. Y. Dauxois and S. Jomhoori. Testing an exponential delay time model against an intensity proportional repair alert model. Communications in Statistics: Theory and Methods, 44(19): 4065–4076, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Djira:2010:RPE

[Dji10] Gemechis Dilba Djira. Relative potency estimation in parallelline assays — method comparison and some extensions. Communications in Statistics: Theory and Methods, 39(7):1180– 1189, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Djordjevic:2017:EIM

[Djo17] Miodrag S. Djordjević. An extension on INAR models with discrete Laplace marginal distributions. Communications in Statistics: Theory and Methods, 46(12):5896–5913, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1115071.

Datta:2017:GCR

[DJVV17] Anindita Datta, Seema Jaggi, Eldho Varghese, and Cini Varghese. Generalized confounded row–column designs. Communications in Statistics: Theory and Methods, 46(12):6213–6221, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1122062.

Dey:2009:DBL

[DK09] Arabin Kumar Dey and Debasis Kundu. Discriminating between the log-normal and log-logistic distributions. Communications in Statistics: Theory and Methods, 39(2):280–292, 2009.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demos:2013:EMA

[DK13] Antonis Demos and Dimitra Kyriakopoulou. Edgeworth and moment approximations: The case of MM and QML estimators for the MA(1) models. Communications in Statistics: Theory and Methods, 42(10):1713–1747, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dawoud:2017:EPP

[DK17a] Issam Dawoud and Selahattin Kaçiranlar. Evaluation of the predictive performance of the r–k and r–d class estimators. Communications in Statistics: Theory and Methods, 46(8): 4031–4050, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1076482.

Dawoud:2017:OMA

[DK17b] Issam Dawoud and Selahattin Kaçiranlar. An optimal k of k-th MA-ARIMA models under a class of ARIMA model. Communications in Statistics: Theory and Methods, 46(12):5754–5765, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1112910.

Dachian:2016:HTPa

[DKY16a] S. Dachian, Yu. A. Kutoyants, and L. Yang. On hypothesis testing for Poisson processes: Regular case. Communications in Statistics: Theory and Methods, 45(23):6816–6832, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.968733.

Dachian:2016:HTPb

[DKY16b] S. Dachian, Yu. A. Kutoyants, and L. Yang. On hypothesis testing for Poisson processes. Singular cases. Communications in Statistics: Theory and Methods, 45(23):6833–6859, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.968734.

Dudek:2011:BAD

[DL11] Anna Dudek and Jacek Leśkow. A bootstrap algorithm for data from a periodic multiplicative intensity function. Communications in Statistics: Theory and Methods, 40(8):1468–1489, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dorea:2016:WST

[DL16] Chang C. Y. Dorea and Luciene P. Lopes. Weighted similarity tests for location-scale families of stable distributions. Communications in Statistics: Theory and Methods, 45(16):4912–4922, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.932809.

Du:2019:WJM

[DL19] Ruofei Du and Ji-Hyun Lee. A weighted jackknife method for clustered data. Communications in Statistics: Theory and Methods, 48(8):1963–1980, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1440597.

Dine:2017:MFM

[DLB17] Abdessamad Dine, Denis Larocque, and François Bellavance. Multivariate forests with missing mixed outcomes. Communications in Statistics: Theory and Methods, 46(23):11500–11513, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1271427.

Du:2016:RSU

[DLC16] Shijia Du, Cong Lin, and Lirong Cui. Reliabilities of a singleunit system with multi-phased missions. Communications in Statistics: Theory and Methods, 45(9):2524–2537, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.882953.

Deng:2018:DIR

[DLH^{+}18] Yingchun Deng, Juan Liu, Ya Huang, Man Li, and Jieming Zhou. On a discrete interaction risk model with delayed claims and stochastic incomes under random discount rates. Communications in Statistics: Theory and Methods, 47(23):5867–5883, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1406518.

Desgagne:2013:TNA

[DLL13] Alain Desgagné, Pierre Lafaye de Micheaux, and Alexandre Leblanc. Test of normality against generalized exponential power alternatives. Communications in Statistics: Theory and Methods, 42(1):164–190, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Du:2015:VSS

[DLP15] Jiang Du, Gaorong Li, and Heng Peng. Variable selection for semiparametric partially linear covariate-adjusted regression models. Communications in Statistics: Theory and Methods, 44(13):2809–2826, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DaVeiga:2017:EEC

[DLS17] Sébastien Da Veiga, Jean-Michel Loubes, and Maikol Solís. Efficient estimation of conditional covariance matrices for dimension reduction. Communications in Statistics: Theory and Methods, 46(9):4403–4424, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1083109.

Du:2017:PEM

[DLZ17] Xiu-Li Du, Jin-Guan Lin, and Xiu-Qing Zhou. Parameter estimation for multivariate diffusion processes with the time inhomogeneously positive semidefinite diffusion matrix. Communications in Statistics: Theory and Methods, 46(22):11010–11025, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1257721.

Denis:2010:FKS

[DM10] M. Denis and N. Molinari. Free knot splines with RJMCMC in survival data analysis. Communications in Statistics: Theory and Methods, 39(14):2617–2629, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dash:2011:ICE

[DM11] P. R. Dash and G. Mishra. An improved class of estimators in two-phase sampling using two auxiliary variables. Communications in Statistics: Theory and Methods, 40(24):4347–4352, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dossar:2018:CAL

[DM18] Petan Dossar and Mounir Mesbah. Cumulative or adjacent logits: Which choice for an ordinal logistic latent variable model? Communications in Statistics: Theory and Methods, 47(11): 2563–2575, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1060342.

Dey:2017:ERM

[DMA17] Sanku Dey, Josmar Mazucheli, and M. Z. Anis. Estimation of reliability of multicomponent stress-strength for a Kumaraswamy distribution. Communications in Statistics: Theory and Methods, 46(4):1560–1572, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1022457.

DAmico:2016:TPS

[DMM16] Guglielmo D’Amico, Raimondo Manca, and Sally McClean. Theory and practice of stochastic models and data analysis. Communications in Statistics: Theory and Methods, 45(6): 1577–1579, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1107333.

DeIaco:2011:SPD

[DMP11] S. De Iaco, D. E. Myers, and D. Posa. Strict positive definiteness of a product of covariance functions. Communications in Statistics: Theory and Methods, 40(24):4400–4408, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dentler:2014:ETN

[DMRO14] Alexander Dentler, Gabriel Montes-Rojas, and Jose Olmo. Endogeneity in threshold nonlinearity tests. Communications in Statistics: Theory and Methods, 43(1):105–114, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DAmico:2014:BSM

[DMS14] Guglielmo D’Amico, Raimondo Manca, and Giovanni Salvi. Bivariate semi-Markov process for counterparty credit risk. Communications in Statistics: Theory and Methods, 43(7):1503– 1522, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:2012:BMM

[DMST12] Haisong Deng, Yizhong Ma, Wenze Shao, and Yiliu Tu. A Bayesian meta-modeling approach for Gaussian stochastic process models using a non informative prior. Communications in Statistics: Theory and Methods, 41(5):829–850, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:2015:ECP

[DMT^{+}15] C. Dong, B. Miao, C. Tan, D. Wei, and Y. Wu. An estimate of a change point in variance of measurement errors and its convergence rate. Communications in Statistics: Theory and Methods, 44(4):790–797, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dixit:2014:EPM

[DN14] V. U. Dixit and S. P. Nabar. Estimation of parameters of a mixed failure time distribution which is a mixture of degenerate (degenerate at zero) and I.G. distribution. Communications in Statistics: Theory and Methods, 43(23):5039–5051, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dabo-Niang:2012:NQR

[DNL12] Sophie Dabo-Niang and Ali Laksaci. Nonparametric quantile regression estimation for functional dependent data. Communications in Statistics: Theory and Methods, 41(7):1254–1268, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DiNardo:2013:SAS

[DO13] E. Di Nardo and I. Oliva. On some applications of a symbolic representation of non centered lévy processes. Communications in Statistics: Theory and Methods, 42(21):3974–3988, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Don:2019:BAO

[DO19] Hasthika S. Rupasinghe Arachchige Don and David J. Olive. Bootstrapping analogs of the one way MANOVA test. Communications in Statistics: Theory and Methods, 48(22):5546–5558, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1515363.

Dokuchaev:2012:DDT

[Dok12] Nikolai Dokuchaev. On detecting the dependence of time series. Communications in Statistics: Theory and Methods, 41(5):934– 942, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Domma:2011:BRH

[Dom11] Filippo Domma. Bivariate reversed hazard rate, notions, and measures of dependence and their relationships. Communications in Statistics: Theory and Methods, 40(6):989–999, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Donchev:2011:RST

[Don11] Doncho S. Donchev. Random series with time-varying discounting. Communications in Statistics: Theory and Methods, 40 (16):2866–2878, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dosla:2011:EPC

[Dos11] Sárka Doslá. Estimation of parameters of a clipped MA(1) process. Communications in Statistics: Theory and Methods, 40(14):2437–2454, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davarzani:2010:BID

[DP10a] Nasser Davarzani and Ahmad Parsian. Bayesian inference in dependent right censoring. Communications in Statistics: Theory and Methods, 39(7):1270–1288, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diana:2010:IEP

[DP10b] Giancarlo Diana and Pier Francesco Perri. Improved estimators of the population mean for missing data. Communications in Statistics: Theory and Methods, 39(18):3245–3251, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davarzani:2013:IUR

[DP13a] Nasser Davarzani and Ahmad Parsian. Inference under right censoring in a discrete setup. Communications in Statistics: Theory and Methods, 42(13):2362–2375, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dvorak:2013:TCA

[DP13b] Marek Dvorák and Zuzana Prásková. On testing changes in autoregressive parameters of a VAR model. Communications in Statistics: Theory and Methods, 42(7):1208–1226, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:2014:RRB

[DP14] Rabindra Nath Das and Jeong-Soo Park. A reinforced randomized block design with correlated errors. Communications in Statistics: Theory and Methods, 43(1):191–209, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dehua:2017:CCM

[DP17] Qiu Dehua and Chen Pingyan. Complete and complete moment convergence for i.i.d. random variables under exponential moment conditions. Communications in Statistics: Theory and Methods, 46(9):4510–4519, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1085566.

Dash:2018:NTQ

[DP18] Priyaranjan Dash and Bijoy Kumar Pradhan. A note on a transient queuing system of a linear birth-immigration process in presence of twin births. Communications in Statistics: Theory and Methods, 47(15):3605–3615, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361983.

Das:2010:RSO

[DPA10] Rabindra Nath Das, Sung H. Park, and Manohar Aggarwal. Robust second-order slope-rotatable designs with maximum directional variance. Communications in Statistics: Theory and Methods, 39(5):803–814, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dai:2013:MSE

[DPB13] Hongsheng Dai, Jianxin Pan, and Yanchun Bao. Modelling survival events with longitudinal covariates measured with error. Communications in Statistics: Theory and Methods, 42(21): 3819–3837, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

dePaula:2016:GLR

[dPD16] Marcelo de Paula and Carlos Alberto Ribeiro Diniz. Generalized linear regression models incorporating original outcome distributions. Communications in Statistics: Theory and Methods, 45(19):5762–5786, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
948726.

Das:2015:MRS

[DPP15a] Rabindra Nath Das, Partha Pal, and Sung H. Park. Modified robust second-order slope-rotatable designs. Communications in Statistics: Theory and Methods, 44(1):80–94, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davarzani:2015:DRC

[DPP15b] Nasser Davarzani, Ahmad Parsian, and Ralf Peeters. Dependent right censorship in the MOMW distribution. Communications in Statistics: Theory and Methods, 44(11):2222–2242, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DAmico:2017:LDS

[DPP17] Guglielmo D’Amico, Filippo Petroni, and Flavio Prattico. On the limit distribution of a second-order semi-Markov chain in state and duration. Communications in Statistics: Theory and Methods, 46(12):5994–5999, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1116574.

Diniz:2011:URB

[DPS11] M. Diniz, C. A. B. Pereira, and J. M. Stern. Unit roots: Bayesian significance test. Communications in Statistics: Theory and Methods, 40(23):4200–4213, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diniz:2012:CBS

[DPS12] M. Diniz, C. A. B. Pereira, and J. M. Stern. Cointegration: Bayesian significance test. Communications in Statistics: Theory and Methods, 41(19):3562–3574, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dan:2017:DES

[DPZ17] Wang Dan, Guo Pengjiang, and Xia Zhiming. Detection and estimation of structural change in heavy-tailed sequence. Communications in Statistics: Theory and Methods, 46(2):815–827, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1006780.

Dreassi:2017:BSM

[DR17] Emanuela Dreassi and Emilia Rocco. A Bayesian semiparametric model for non negative semicontinuous data. Communications in Statistics: Theory and Methods, 46(10):5133–5146, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1096389.

Dragieva:2013:FSR

[Dra13] Velika I. Dragieva. A finite source retrial queue: Number of retrials. Communications in Statistics: Theory and Methods, 42 (5):812–829, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dolati:2017:EDD

[DRAS17] Ali Dolati, Rasool Roozegar, Najmeh Ahmadi, and Zohreh Shishebor. The effect of dependence on distribution of the functions of random variables. Communications in Statistics: Theory and Methods, 46(21):10704–10717, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1242740.

Druilhet:2017:FPL

[Dru17] P. Druilhet. On the Flatland paradox and limiting arguments. Communications in Statistics: Theory and Methods, 46(24):12281–12289, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1295157.

Dagne:2009:BAR

[DS09] Getachew A. Dagne and James Snyder. Bayesian analysis of repeated events using event-dependent frailty models: an application to behavioral observation data. Communications in Statistics: Theory and Methods, 39(2):293–310, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dohi:2010:P

[DS10] Tadashi Dohi and Shey-Huei Sheu. Preface. Communications in Statistics: Theory and Methods, 39(13):2311, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Durante:2014:NCB

[DS14] Fabrizio Durante and Juan Fernández Sánchez. A note on the compatibility of bivariate copulas. Communications in Statistics: Theory and Methods, 43(9):1918–1923, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deniz:2016:DDI

[DS16a] Emilio Gómez Déniz and José María Sarabia. A discrete distribution including the Poisson. Communications in Statistics: Theory and Methods, 45(8):2311–2322, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.879896.

Dutta:2016:CMD

[DS16b] Santanu Dutta and Koushik Saha. Consistency of multivariate density estimators using random bandwidths. Communications in Statistics: Theory and Methods, 45(2):252–266, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:2017:MDO

[DS17a] Premadhis Das and Bikas K. Sinha. Mixture designs with orthogonal blocks. Communications in Statistics: Theory and Methods, 46(16):8155–8165, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1175629.

Dixit:2017:TST

[DS17b] V. U. Dixit and Leela Subramanian. Tests for symmetry of two-piece normal distribution. Communications in Statistics: Theory and Methods, 46(4):1807–1820, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030419.

Dutta:2017:NOCa

[DS17c] Ganesh Dutta and Bikas Kumar Sinha. Nearly optimal covariate designs — Part I. Communications in Statistics: Theory and Methods, 46(19):9691–9702, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1217020.

Dutta:2017:NOCb

[DS17d] Ganesh Dutta and Bikas Kumar Sinha. Nearly optimal covariate designs — Part II. Communications in Statistics: Theory and Methods, 46(19):9703–9725, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1217021.

Dutta:2017:PUC

[DS17e] Santanu Dutta and Koushik Saha. Pointwise and uniform convergence of multivariate kernel density estimators using random bandwidths. Communications in Statistics: Theory and Methods, 46(6):2708–2723, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1048886.

Dewan:2016:PHM

[DSA16] Isha Dewan, K. K. Sudheesh, and P. Anisha. Proportional hazards model for discrete data: Some new developments. Communications in Statistics: Theory and Methods, 45(21):6481–6493, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.957864.

Braga:2018:NSB

[dSBCO18] Altemir da Silva Braga, Gauss M. Cordeiro, and Edwin M. M. Ortega. A new skew-bimodal distribution with applications. Communications in Statistics: Theory and Methods, 47(12): 2950–2968, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1343851.

Dui:2014:UIM

[DSC^{+}14] Hongyan Dui, Shubin Si, Zhiqiang Cai, Shudong Sun, and Yingfeng Zhang. On the use of the importance measure for multi-state repairable k-out-of-n: G systems. Communications in Statistics: Theory and Methods, 43(13):2766–2781, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

daSilva:2011:DSD

[dSE11] José Luís da Silva and Mohamed Erraoui. The α-dependence of stochastic differential equations driven by variants of α-stable processes. Communications in Statistics: Theory and Methods, 40(19–20):3465–3478, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demidenko:2012:REC

[DSO12] Eugene Demidenko, James Sargent, and Tracy Onega. Random effects coefficient of determination for mixed and meta-analysis models. Communications in Statistics: Theory and Methods, 41 (6):953–969, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dykes:2015:CEP

[DSSL15] Lee Dykes, Sarjinder Singh, Stephen A. Sedory, and Vincent Louis. Calibrated estimators of population mean for a Mail survey design. Communications in Statistics: Theory and Methods, 44(16):3403–3427, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dimitriou:2012:ASM

[DT12] V. A. Dimitriou and N. Tsantas. The augmented semi-Markov system in continuous time. Communications in Statistics: Theory and Methods, 41(1):88–107, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeNeve:2017:MWT

[DT17] Jan De Neve and Olivier Thas. A Mann–Whitney type effect measure of interaction for factorial designs. Communications in Statistics: Theory and Methods, 46(22):11243–11260, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1263739.

Devlin:2013:RAR

[DTE13] Sean M. Devlin, Elizabeth G. Thomas, and Scott S. Emerson. Robustness of approaches to ROC curve modeling under misspecification of the underlying probability model. Communications in Statistics: Theory and Methods, 42(20):3655–3664, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeBoeck:2011:RST

[DTO11] B. De Boeck, O. Thas, and J. P. Ottoy. Reweighted smooth tests of goodness of fit. Communications in Statistics: Theory and Methods, 40(9):1651–1670, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeNeve:2013:GFM

[DTO13] Jan De Neve, Olivier Thas, and Jean-Pierre Ottoy. Goodnessof-fit methods for probabilistic index models. Communications in Statistics: Theory and Methods, 42(7):1193–1207, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ding:2016:NDL

[DTT^{+}16] Qian Ding, Qiaohui Tong, Qiaoling Tong, Xuecheng Zou, and Hengqing Tong. A new definite linear algorithm for the structural equation model with application in fire risk evaluation. Communications in Statistics: Theory and Methods, 45(23):

7075–7085, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.974824.

Ding:2017:CMC

[DTWW17] Yang Ding, Xuefei Tang, Hui Wang, and Xuejun Wang. Complete moment convergence of moving-average process generated by a class of random variables. Communications in Statistics: Theory and Methods, 46(22):10903–10913, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1252401.

Dufresne:2010:BPD

[Duf10] Daniel Dufresne. The beta product distribution with complex parameters. Communications in Statistics: Theory and Methods, 39(5):837–854, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutang:2017:TMM

[Dut17] Christophe Dutang. Theoretical L-moments and T L-moments using combinatorial identities and finite operators. Communications in Statistics: Theory and Methods, 46(8):3801–3828, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1073313.

Dvorak:2017:DET

[Dvo17] Marek Dvorák. Darling–Erdős-type test for change detection in parameters and variance for stationary VAR models. Communications in Statistics: Theory and Methods, 46(1):465–484, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.995828.

Dong:2013:FVL

[DW13] Yinghui Dong and Guojing Wang. Fair valuation of life insurance contracts under a two-sided jump diffusion model. Communications in Statistics: Theory and Methods, 42(21):3926– 3948, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:2017:MDR

[DW17] Yinghua Dong and Dingcheng Wang. Moderate deviations for the random weighted sums of END random variables with consistently varying tails. Communications in Statistics: Theory and Methods, 46(20):10116–10134, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1231820.

Don:2018:BAT

[DW18] H. S. Rupasinghe Arachchige Don and L. C. R. Pelawa Watagoda. Bootstrapping analogs of the two-sample Hotelling’s T^{2} test. Communications in Statistics: Theory and Methods, 47(9):2172–2182, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1337146.

Duan:2019:ODC

[DW19] Fengjun Duan and Guanjun Wang. Optimal design for constant-stress accelerated degradation test based on gamma process. Communications in Statistics: Theory and Methods, 48(9):2229–2253, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1459718.

Ding:2017:CCC

[DWM^{+}17] Yang Ding, Yi Wu, Songlin Ma, Xinran Tao, and Xuejun Wang. Complete convergence and complete moment convergence for widely orthant-dependent random variables. Communications in Statistics: Theory and Methods, 46(16):8278–8294, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1177085.

Ding:2018:REB

[DWW18] Hao Ding, Zhanfeng Wang, and Yaohua Wu. A relative errorbased estimation with an increasing number of parameters. Communications in Statistics: Theory and Methods, 47(1):196– 209, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1301474.

Deng:2017:HRT

[DWX17] Xin Deng, Xuejun Wang, and Fengxi Xia. Hajek–Renyi-type inequality and strong law of large numbers for END sequences. Communications in Statistics: Theory and Methods, 46(2):672– 682, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.1002938.

Xing:2012:NCR

[dXcY12] Guo dong Xing and Shan chao Yang. A note on the convergence rate of strong law of large numbers for positively associated sequences. Communications in Statistics: Theory and Methods, 41(6):1116–1125, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeGooijer:2016:NPP

[DY16] Jan G. De Gooijer and Ao Yuan. Non parametric portmanteau tests for detecting non linearities in high dimensions. Communications in Statistics: Theory and Methods, 45(2):385–399, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:2018:PMG

[DY18] Xiang Deng and Jing Yao. On the property of multivariate generalized hyperbolic distribution and the Stein-type inequality. Communications in Statistics: Theory and Methods, 47(21): 5346–5356, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1390134.

Ding:2017:SEO

[DYL17] Weiyong Ding, Jiaxin Yang, and Xiaoliang Ling. On the skewness of extreme order statistics from heterogeneous samples. Communications in Statistics: Theory and Methods, 46(5): 2315–2331, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1041984.

Dong:2018:RSP

[DYW18] Yinghui Dong, Kam Chuen Yuen, and Guojing Wang. Regimeswitching pure jump processes and applications in the valuation of mortality-linked products. Communications in Statistics: Theory and Methods, 47(6):1372–1391, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1319483.

Dai:2019:RVS

[DYXW19] Lin Dai, Junhui Yin, Zhengfen Xie, and Liucang Wu. Robust variable selection in finite mixture of regression models using the t distribution. Communications in Statistics: Theory and Methods, 48(21):5370–5386, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1513143.

Deldossi:2012:CIV

[DZ12] Laura Deldossi and Diego Zappa. Confidence intervals for variance components in measurement system capability studies. Communications in Statistics: Theory and Methods, 41(16–17): 2932–2943, 2012. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/03610926.2011.589956.

Deldossi:2014:NAE

[DZ14] Laura Deldossi and Diego Zappa. A novel approach to evaluate repeatability and reproducibility for ordinal data. Communications in Statistics: Theory and Methods, 43(4):851–866, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:2015:STB

[DZ15] Dianliang Deng and Yu Zhang. Score tests for both extra zeros and extra ones in binomial mixed regression models. Communications in Statistics: Theory and Methods, 44(14):2881–2897, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:2019:CSN

[DZH^{+}19] Yuge Dong, Haimeng Zhang, Liangguo He, Can Wang, and Minghui Wang. The computation of standard normal distribution integral in any required precision based on reliability method. Communications in Statistics: Theory and Methods, 48(6):1517–1528, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1433855.

Deng:2018:FDR

[DZL18] Lu Deng, Xuemin Zi, and Zhonghua Li. False discovery rates for large-scale model checking under certain dependence. Communications in Statistics: Theory and Methods, 47(1):64–79, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1300279.

Du:2013:FIC

[DZX13] Jiang Du, Zhongzhan Zhang, and Tianfa Xie. Focused information criterion and model averaging in quantile regression. Communications in Statistics: Theory and Methods, 42(20): 3716–3734, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eidous:2018:BCH

[EAE18] Omar Eidous and Fahid Al-Eibood. A bias-corrected histogram estimator for line transect sampling. Communications in Statistics: Theory and Methods, 47(15):3675–3686, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361987.

El-Baset:2012:JMG

[EB12] A. Ahmad Abd El-Baset. Joint moment generating functions of nonadjacent dual generalized order statistics from reflected generalized Pareto distributions. Communications in Statistics: Theory and Methods, 41(15):2762–2772, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Din:2012:TSB

[EDAAS12] M. M. Mohie El-Din, Y. Abdel-Aty, and A. R. Shafay. Twosample Bayesian prediction intervals of generalized order statistics based on multiply Type II censored data. Communications in Statistics: Theory and Methods, 41(3):381–392, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Editor:2010:P

[Edi10a] Christos H. Skiadas Guest Editor. Preface. Communications in Statistics: Theory and Methods, 39(3):383–384, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Editor:2010:PSI

[Edi10b] Makoto Aoshima Organizing Editor. A path to the special issue on recent advances in statistical inference — in honor of Professor Masafumi Akahira. Communications in Statistics: Theory and Methods, 39(8–9):1321–1323, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Din:2017:BIP

[EDKAEN17] Mostafa M. Mohie El-Din, Mohamed S. Kotb, Ehab F. Abd-Elfattah, and Haidy A. Newer. Bayesian inference and prediction of the Pareto distribution based on ordered ranked set sampling. Communications in Statistics: Theory and Methods, 46(13):6264–6279, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1124118.

Esquivel:2014:RCS

[EdSMR14] Manuel L. Esquível, João Lita da Silva, João Tiago Mexia, and Luís Ramos. The rate of convergence of some asymptotic expansions for distribution approximations via an Esséen type estimate. Communications in Statistics: Theory and Methods, 43(2):266–290, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eskandarzadeh:2018:MIM

[EDT18] Maryam Eskandarzadeh, Antonio Di Crescenzo, and Saeid Tahmasebi. Measures of information for maximum ranked set sampling with unequal samples. Communications in Statistics: Theory and Methods, 47(19):4692–4709, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1445857.

Ennis:2011:CBM

[EE11] John M. Ennis and Daniel M. Ennis. Confidence bounds for multiplicative comparisons. Communications in Statistics:

Theory and Methods, 40(17):3049–3054, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elfadaly:2014:TTE

[EE14] Fadlalla G. Elfadaly and Sanaa M. El Gayar. Two-term Edgeworth expansions for the classes of U- and V -statistics. Communications in Statistics: Theory and Methods, 43(22):4679–4706, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Editors:2015:P

[EE15a] Syed N. U. A. Kirmani Guest Editors and Ram C. Tripathi Guest Editors. Preface. Communications in Statistics: Theory and Methods, 44(19):3997, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

ElBarmi:2015:CCI

[EE15b] Hammou El Barmi and Lahcen El Bermi. On comparing cumulative incidence functions using an empirical likelihood ratio type test. Communications in Statistics: Theory and Methods, 44(23):4940–4952, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Emami:2016:IDC

[EE16] Hadi Emami and Mostafa Emami. Influence diagnostics in constrained general linear models. Communications in Statistics: Theory and Methods, 45(18):5331–5340, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.942430.

Etuk:2016:MCR

[EEE16] Sylvanus Itoro Etuk, Ekaette Inyang Enang, and Emmanuel John Ekpenyong. A modified class of ratio estimators for population mean. Communications in Statistics: Theory and Methods, 45(4):837–849, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
851238.

Eleftheriou:2011:ECC

[EF11] Mavroudis Eleftheriou and Nicolas Farmakis. Expected cost for continuous sampling plans. Communications in Statistics:

Theory and Methods, 40(16):2969–2984, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Egozcue:2018:VUB

[EG18] Martín Egozcue and Luis Fuentes García. The variance upper bound for a mixed random variable. Communications in Statistics: Theory and Methods, 47(22):5391–5395, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1390136.

Eryilmaz:2011:REM

[EI11] Serkan Eryilmaz and Funda Isçioglu. Reliability evaluation for a multi-state system under stress-strength setup. Communications in Statistics: Theory and Methods, 40(3):547–558, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eideh:2012:FVC

[Eid12] Abdulhakeem A. H. Eideh. Fitting variance components model and fixed effects model for one-way analysis of variance to complex survey data. Communications in Statistics: Theory and Methods, 41(16–17):3278–3300, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2012.692425.

Eidous:2014:HMP

[Eid14] Omar Eidous. Histogram model with plausible parametric detection function for line transect data. Communications in Statistics: Theory and Methods, 43(1):1–12, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebrahimzadeh:2013:CDI

[EIJK13] Mahdi Ebrahimzadeh, Noor Akma Ibrahim, Abdul Aziz Jemain, and Adem Kilicman. Claim dependence induced by common effects in hierarchical credibility models. Communications in Statistics: Theory and Methods, 42(18):3373–3400, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ekhosuehi:2018:MAC

[EIM18] Virtue U. Ekhosuehi, Daniel O. Iruegbukpe, and Julian I. Mbegbu. Minimum aberration criterion for screening experiments at two levels from an entropy-based perspective. Communications in Statistics: Theory and Methods, 47(8):1868– 1881, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1328513.

ElMaroufy:2016:FOP

[EKZ16] Hamid El Maroufy, Driss Kiouach, and Taib Ziad. Final outcome probabilities for SIR epidemic model. Communications in Statistics: Theory and Methods, 45(8):2426–2437, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.881494.

ElGhini:2015:PPP

[El 15] Ahmed El Ghini. Probabilistic properties of parametric dual and inverse time series models generated by ARMA models. Communications in Statistics: Theory and Methods, 44(21): 4651–4661, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

ElGenidy:2017:MNL

[El 17] Mohammed Mohammed El Genidy. Multiple non linear regression model for the maximum number of migratory bird types during migration years. Communications in Statistics: Theory and Methods, 46(16):7969–7975, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1167909.

ElGenidy:2019:MNR

[El 19] Mohammed El Genidy. Multiple nonlinear regression of the Markovian arrival process for estimating the daily global solar radiation. Communications in Statistics: Theory and Methods, 48(22):5427–5444, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1517890.

Ellis:2011:WCC

[Ell11] Steven P. Ellis. When a constant classifier is as good as any linear classifier. Communications in Statistics: Theory and Methods, 40(21):3800–3811, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

ElAnbari:2013:AGE

[EM13] Mohammed El Anbari and Abdallah Mkhadri. The adaptive gril estimator with a diverging number of parameters. Communications in Statistics: Theory and Methods, 42(14):2634–2660, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Monsef:2016:NLD

[EM16] M. M. E. Abd El-Monsef. A new Lindley distribution with location parameter. Communications in Statistics: Theory and Methods, 45(17):5204–5219, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
941496.

Elkadry:2018:ACR

[EM18] Alaa Elkadry and Gary C. McDonald. Analyzing continuous randomized response data with an indifference-zone selection procedure. Communications in Statistics: Theory and Methods, 47(14):3508–3522, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1359298.

ElMatouat:2016:CIC

[EMH16] Abdelaziz El Matouat, Freedath Djibril Moussa, and Hassania Hamzaoui. Consistency of information criteria for model selection with missing data. Communications in Statistics: Theory and Methods, 45(23):6900–6914, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.972568.

El-Monsef:2017:ELD

[EMHK17] M. M. E. Abd El-Monsef, W. A. Hassanein, and N. M. Kilany. Erlang–Lindley distribution. Communications in Statistics: Theory and Methods, 46(19):9494–9506, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1212069.

Efendi:2014:MCG

[EMI14] Achmad Efendi, Geert Molenberghs, and Samuel Iddi. A marginalized combined gamma frailty and normal randomeffects model for repeated, overdispersed, time-to-event outcomes. Communications in Statistics: Theory and Methods, 43(22):4806–4828, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eljabri:2017:KGD

[EN17] S. Eljabri and S. Nadarajah. The Kumaraswamy GEV distribution. Communications in Statistics: Theory and Methods, 46(20):10203–10235, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1231815.

Eshkaftaki:2011:RBE

[EP11] Abdollah Bayati Eshkaftaki and Ahmad Parsian. Robust Bayes estimation. Communications in Statistics: Theory and Methods, 40(5):929–941, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Enea:2017:GBD

[EP17] Marco Enea and Antonella Plaia. A gradient-based deletion diagnostic measure for generalized linear mixed models. Communications in Statistics: Theory and Methods, 46(4):1972–1982, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1030956.

Egozcue:2015:ICT

[EPGTH15] Juan José Egozcue, Vera Pawlowsky-Glahn, Matthias Templ, and Karel Hron. Independence in contingency tables using simplicial geometry. Communications in Statistics: Theory and Methods, 44(18):3978–3996, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Editors:2014:P

[EPRS14] Paola Monari Guest Editors, Domenico Piccolo, Carla Rampichini, and Luigi Salmaso. Preface. Communications in Statistics: Theory and Methods, 43(4):663, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elsawah:2017:NLO

[EQ17] A. M. Elsawah and Hong Qin. A new look on optimal foldover plans in terms of uniformity criteria. Communications in Statistics: Theory and Methods, 46(4):1621–1635, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1024862.

ElMethni:2011:LWA

[ER11] Mohamed El Methni and Mustapha Rachdi. Local weighted average estimation of the regression operator for functional data. Communications in Statistics: Theory and Methods, 40(17): 3141–3153, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Emami:2014:IDT

[ER14] Hadi Emami and Abdolrahman Rasekh. Influence diagnostics on testing linear hypothesis in linear models with correlated errors. Communications in Statistics: Theory and Methods, 43(6):1050–1060, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eryilmaz:2012:RPS

[Ery12] Serkan Eryilmaz. Reliability properties of systems with two exchangeable log-logistic components. Communications in Statistics: Theory and Methods, 41(18):3416–3427, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eryilmaz:2016:DTC

[Ery16a] Serkan Eryilmaz. Discrete time cold standby repairable system: Combinatorial analysis. Communications in Statistics: Theory and Methods, 45(24):7399–7405, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.776689.

Eryilmaz:2016:ACO

[Ery16b] Serkan Eryilmaz. On an application of concomitants of order statistics. Communications in Statistics: Theory and Methods, 45(19):5628–5636, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
948201.

Eryilmaz:2017:NCB

[Ery17] Serkan Eryilmaz. A new class of bivariate lifetime distributions. Communications in Statistics: Theory and Methods, 46(24):12324–12335, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1295078.

Eryilmaz:2018:FTR

[Ery18] Serkan Eryilmaz. On the first time of ruin in two-dimensional discrete time risk model with dependent claim occurrences. Communications in Statistics: Theory and Methods, 47(9): 2251–2258, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1337145.

Eidous:2011:AUK

[ES11a] Omar Eidous and M. K. Shakhatreh. Asymptotic unbiased kernel estimator for line transect sampling. Communications in Statistics: Theory and Methods, 40(24):4353–4363, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ekheden:2011:CER

[ES11b] Erland Ekheden and Dmitrii Silvestrov. Coupling and explicit rate of convergence in Cramér–Lundberg approximation for reinsurance risk processes. Communications in Statistics: Theory and Methods, 40(19–20):3524–3539, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ebadi:2014:REP

[ES14] Mohsen Ebadi and Hamid Shahriari. Robust estimation of parameters in simple linear profiles using M-estimators. Communications in Statistics: Theory and Methods, 43(20):4308–4323, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eftekhar:2018:TES

[ESAKK18] Sana Eftekhar, Mohammad Sadooghi-Alvandi, and Mahmood Kharrati-Kopaei. Testing the equality of several multivariate normal mean vectors under heteroscedasticity: a fiducial approach and an approximate test. Communications in Statistics: Theory and Methods, 47(7):1747–1766, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1324984.

Economou:2013:STD

[ET13] P. Economou and G. Tzavelas. Sample tests for detection of size-biased sampling mechanism. Communications in Statistics: Theory and Methods, 42(18):3280–3295, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eakin:2016:SDE

[EW16] Mark E. Eakin and Mary M. Whiteside. The sampling distribution of the W estimator of the number of valid signatures on a petition. Communications in Statistics: Theory and Methods, 45(5):1224–1240, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
837183.

ElBarmi:2017:EPO

[EW17] Hammou El Barmi and Rongning Wu. On estimation of peakedness-ordered distributions. Communications in Statistics: Theory and Methods, 46(10):4855–4869, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1089291.

El-Wahish:2014:PCK

[EWRZ14] Huda A. El-Wahish, Mohammad Z. Raqab, and Ahmad A. Zghoul. Pitman closeness of kth records based on a twosequence case. Communications in Statistics: Theory and Methods, 43(23):5091–5104, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eguchi:2010:MRL

[EY10] Shinto Eguchi and Koichi Yoshioka. Maximum regularized likelihood estimator of finite mixtures with a structural model. Communications in Statistics: Theory and Methods, 39(8–9): 1498–1510, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Einbeck:2014:SAL

[EZ14] Jochen Einbeck and Mohammad A. Zayed. Some asymptotics for localized principal components and curves. Communications in Statistics: Theory and Methods, 43(8):1736–1749, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feroze:2019:ABA

[FA19] Navid Feroze and Muhammad Aslam. Approximate Bayesian analysis of doubly censored samples from mixture of two Weibull distributions. Communications in Statistics: Theory and Methods, 48(11):2862–2878, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473430.

Fabian:2016:SFD

[Fab16] Zdenek Fabián. Score function of distribution and revival of the moment method. Communications in Statistics: Theory and Methods, 45(4):1118–1136, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.857688.

Fakoor:2015:NMR

[Fak15] V. Fakoor. On the nonparametric mean residual life estimator in length-biased sampling. Communications in Statistics: Theory and Methods, 44(3):512–519, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fallahnezhad:2014:FHD

[Fal14] Mohammad Saber Fallahnezhad. A finite horizon dynamic programming model for production and repair decisions. Communications in Statistics: Theory and Methods, 43(15):3302–3313, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Famoye:2015:MGP

[Fam15] Felix Famoye. A multivariate generalized Poisson regression model. Communications in Statistics: Theory and Methods, 44 (3):497–511, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2015:NAS

[Fan15] Xi-Liang Fan. Non autonomous semilinear stochastic evolution equations. Communications in Statistics: Theory and Methods, 44(9):1806–1818, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

FangYuan:2018:AMR

[Fan18] Kang FangYuan. An additive marginal regression model for clustered recurrent event in the presence of a terminal event. Communications in Statistics: Theory and Methods, 47(8): 1901–1912, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1332221.

Farchione:2011:FBI

[Far11] Davide Farchione. Frequentist and Bayesian interval estimators for the normal mean. Communications in Statistics: Theory and Methods, 40(15):2680–2694, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farcomeni:2012:TSI

[Far12] Alessio Farcomeni. Testing supremacy or inferiority of multinomial cell probabilities with application to biting preferences of loggerhead marine turtles. Communications in Statistics: Theory and Methods, 41(1):34–45, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farinetto:2017:HTD

[Far17a] Christian Farinetto. On hypothesis tests for disk-type intensities of spatial Poisson processes. Communications in Statistics: Theory and Methods, 46(9):4353–4368, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1082590.

Farinetto:2017:HTM

[Far17b] Christian Farinetto. On hypothesis tests in misspecified changepoint problems for a Poisson process. Communications in Statistics: Theory and Methods, 46(20):10103–10115, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1231819.

Frisen:2011:SRM

[FAS11] Marianne Frisén, Eva Andersson, and Linus Schiöler. Sufficient reduction in multivariate surveillance. Communications in Statistics: Theory and Methods, 40(10):1821–1838, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fathallah:2013:APL

[Fat13] Hamdi Fathallah. Asymptotic properties of the LS-estimator of a Gaussian autoregressive process by an averaging method. Communications in Statistics: Theory and Methods, 42(17): 3148–3173, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fallah:2017:RRE

[FAT17] R. Fallah, M. Arashi, and S. M. M. Tabatabaey. On the ridge regression estimator with sub-space restriction. Communications in Statistics: Theory and Methods, 46(23):11854–11865, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1285928.

Faugeras:2012:PQC

[Fau12] Olivier P. Faugeras. Prediction via the quantile-copula conditional density estimator. Communications in Statistics: Theory and Methods, 41(1):16–33, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fazli:2011:PLI

[Faz11] Khosrow Fazli. Power loss for inhomogeneous Poisson processes. Communications in Statistics: Theory and Methods, 40 (4):648–662, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finos:2010:ASV

[FBS10] Livio Finos, Chiara Brombin, and Luigi Salmaso. Adjusting stepwise p-values in generalized linear models. Communications in Statistics: Theory and Methods, 39(10):1832–1846, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fazekas:2018:CMP

[FC18a] István Fazekas and Alexey Chuprunov. The conditional maximum of Poisson random variables. Communications in Statistics: Theory and Methods, 47(16):3857–3870, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1364388.

Feng:2018:OIL

[FC18b] Yanqin Feng and Yurong Chen. Oracle inequalities for the lasso in the additive hazards model with interval-censored data. Communications in Statistics: Theory and Methods, 47(12): 2927–2949, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1343850.

Feng:2017:EPP

[FCJ^{+}17] Yarong Feng, Xing Chen, Liyi Jia, Xiruo Song, and Hosam M. Mahmoud. Estimating the pólya process. Communications in Statistics: Theory and Methods, 46(19):9397–9406, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1208242.

Faraz:2016:RES

[FCSH16] Alireza Faraz, Kamyar Chalaki, Erwin M. Saniga, and Cedric Heuchenne. The robust economic statistical design of the Hotelling’s T^{2} chart. Communications in Statistics: Theory and Methods, 45(23):6989–7001, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.972574.

Francis:2014:MML

[FDHH14] Brian Francis, Regina Dittrich, Reinhold Hatzinger, and Les Humphreys. A mixture model for longitudinal partially ranked data. Communications in Statistics: Theory and Methods, 43 (4):722–734, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feng:2017:AHR

[FDL17] Yanqin Feng, Yuan Dong, and Yang Li. Additive hazards regression of current status data with auxiliary covariates. Communications in Statistics: Theory and Methods, 46(21):10657– 10671, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1242737.

Feddag:2013:CLE

[Fed13] M.-L. Feddag. Composite likelihood estimation for multivariate probit latent traits models. Communications in Statistics: Theory and Methods, 42(14):2551–2566, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feddag:2014:GEE

[Fed14] M-L. Feddag. Generalized estimating equations to binary probit model. Communications in Statistics: Theory and Methods, 43(19):3997–4010, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fellows:2010:MMV

[Fel10] Ian Fellows. The minimaxity of the mid p-value under linear and squared loss functions. Communications in Statistics: Theory and Methods, 40(2):244–254, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feldman:2018:MQE

[FEP18] Michael Feldman, Adir Even, and Yisrael Parmet. A methodology for quantifying the effect of missing data on decision quality in classification problems. Communications in Statistics: Theory and Methods, 47(11):2643–2663, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1277752.

Ferreira:2014:BTD

[Fer14] Helena Ferreira. Bivariate tail dependence and the generation of multivariate extreme value distributions. Communications in Statistics: Theory and Methods, 43(24):5318–5325, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Filus:2013:MMP

[FF13] Jerzy K. Filus and Lidia Z. Filus. A method for multivariate probability distributions construction via parameter dependence. Communications in Statistics: Theory and Methods, 42 (4):716–721, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feeman:2016:EBG

[FF16] Timothy G. Feeman and Jesse Frey. Efficiency bounds for a generalization of ranked-set sampling. Communications in Statistics: Theory and Methods, 45(3):739–756, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frey:2018:FME

[FF18] Jesse Frey and Timothy G. Feeman. Finding the maximum efficiency for multistage ranked-set sampling. Communications in Statistics: Theory and Methods, 47(17):4131–4141, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1367816.

Ferreira:2019:ADB

[FF19] Helena Ferreira and Marta Ferreira. Asymptotic dependence of bivariate maxima. Communications in Statistics: Theory and Methods, 48(13):3269–3279, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1475568.

Filus:2009:FMD

[FFA09] Jerzy K. Filus, Lidia Z. Filus, and Barry C. Arnold. Families of multivariate distributions involving “Triangular” transformations. Communications in Statistics: Theory and Methods, 39 (1):107–116, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferreira:2019:CRI

[FFN^{+}19] Dário Ferreira, Sandra S. Ferreira, Célia Nunes, Miguel Fonseca, and João T. Mexia. Chisquared and related inducing pivot variables: an application to orthogonal mixed models. Communications in Statistics: Theory and Methods, 48(22):5445–5466, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.770532.

Fokoue:2011:OED

[FG11] Ernest Fokoué and Prem Goel. An optimal experimental design perspective on radial basis function regression. Communications in Statistics: Theory and Methods, 40(7):1184–1195, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Figini:2017:CRA

[FG17] Silvia Figini and Paolo Giudici. Credit risk assessment with Bayesian model averaging. Communications in Statistics: Theory and Methods, 46(19):9507–9517, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1212070.

Fernandez:2018:CBP

[FGGL18] M. Fernández, Jesús E. García, and V. A. González-López. A copula-based partition Markov procedure. Communications in Statistics: Theory and Methods, 47(14):3408–3417, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1359291.

Fu:2018:PMM

[FGL18] Qiang Fu, Xin Guo, and Kenneth C. Land. A Poissonmultinomial mixture approach to grouped and right-censored counts. Communications in Statistics: Theory and Methods, 47 (2):427–447, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1303736.

Fallah:2016:IBP

[FGP16] Lida Fallah, Leila Golparvar, and Ahmad Parsian. Inference based on progressively censored sample from Pareto population. Communications in Statistics: Theory and Methods, 45(1):9– 24, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2012:MAC

[FGTS12] Lianyan Fu, Wei Gao, Man-Lai Tang, and Ning-Zhong Shi. On modelling agreement and category distinguishability on an ordinal scale. Communications in Statistics: Theory and Methods, 41(24):4413–4426, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fiserova:2012:SIO

[FH12] Eva Fiserová and Karel Hron. Statistical inference in orthogonal regression for three-part compositional data using a linear model with Type-II constraints. Communications in Statistics: Theory and Methods, 41(13–14):2367–2385, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2013:ELS

[FH13a] Guo-Liang Fan and Zhen-Sheng Huang. Empirical likelihood for semiparametric varying-coefficient heteroscedastic partially linear errors-in-variables models. Communications in Statistics: Theory and Methods, 42(13):2437–2450, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Filova:2013:CRE

[FH13b] Lenka Filová and Radoslav Harman. Criterion-robust experimental designs for the quadratic regression on a square and a cube. Communications in Statistics: Theory and Methods, 42(11):2044–2055, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:2015:ERS

[FHS15] Juntao Fang, Zhen He, and Linxi Song. Evaluation of response surface designs in presence of errors in factor levels. Communications in Statistics: Theory and Methods, 44(18):3769–3781, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feng:2012:EVC

[FHZ12] Jingyan Feng, Zhensheng Huang, and Riquan Zhang. Estimation on varying-coefficient partially linear model with different smoothing variables. Communications in Statistics: Theory and Methods, 41(3):516–529, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Faroughi:2017:BZI

[FI17] Pouya Faroughi and Noriszura Ismail. Bivariate zero-inflated generalized Poisson regression model with flexible covariance. Communications in Statistics: Theory and Methods, 46(15):

7769–7785, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1165846.

Fakoor:2011:SAR

[FJ11] Vahid Fakoor and Sarah Jomhoori. Some asymptotic results of kernel density estimators under random left-truncation and dependent data. Communications in Statistics: Theory and Methods, 40(23):4249–4262, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fernandes:2018:CSE

[FJS^{+}18] Adji Achmad Rinaldo Fernandes, Paul Janssen, Umu Sa’adah, Solimun, Achmad Effendi, Nurjannah, and Luthfatul Amaliana. Comparison of spline estimator at various levels of autocorrelation in smoothing spline non parametric regression for longitudinal data. Communications in Statistics: Theory and Methods, 47(21):5265–5285, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1388404.

Fan:2017:ELH

[FJW17] Guo-Liang Fan, Zhi-Qiang Jiang, and Jiang-Feng Wang. Empirical likelihood for high-dimensional partially linear model with martingale difference errors. Communications in Statistics: Theory and Methods, 46(22):11228–11242, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1260739.

Fathizadeh:2016:AAR

[FK16] M. Fathizadeh and K. Khorshidian. An alternative approach to reliability analysis of cold standby systems. Communications in Statistics: Theory and Methods, 45(21):6471–6480, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.944660.

Farahani:2019:PMS

[FK19] Ebrahim Mazrae Farahani and Reza Baradaran Kazemzadeh. Phase I monitoring of social network with baseline periods using Poisson regression. Communications in Statistics: Theory and Methods, 48(2):311–331, 2019. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1408836.

Farahani:2017:SAS

[FKNR17] Ebrahim Mazrae Farahani, Reza Baradaran Kazemzadeh, Rassoul Noorossana, and Ghazaleh Rahimian. A statistical approach to social network monitoring. Communications in Statistics: Theory and Methods, 46(22):11272–11288, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1263741.

Fort:2016:NSA

[FKR16] Jean-Claude Fort, Thierry Klein, and Nabil Rachdi. New sensitivity analysis subordinated to a contrast. Communications in Statistics: Theory and Methods, 45(15):4349–4364, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901369.

Fu:2017:TBS

[FL17] Ke-Ang Fu and Jie Li. Tail behavior for the sum of two correlated classes of discounted aggregate claims in a time-dependent risk model. Communications in Statistics: Theory and Methods, 46(5):2559–2570, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.1000499.

Fang:2018:OEI

[FL18a] Rui Fang and Xiaohu Li. Ordering extremes of interdependent random variables. Communications in Statistics: Theory and Methods, 47(17):4187–4201, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1371754.

Fu:2018:PLD

[FL18b] Ke-Ang Fu and Jie Li. Precise large deviations of aggregate claims in a risk model with size dependence and non stationary arrivals. Communications in Statistics: Theory and Methods, 47(3):698–707, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1310244.

Fu:2019:ARP

[FL19] Ke-Ang Fu and Jie Li. Asymptotic ruin probabilities for a bidimensional risk model with heavy-tailed claims and nonstationary arrivals. Communications in Statistics: Theory and Methods, 48(24):6169–6178, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1529242.

Fang:2016:SCS

[FLB16] Longxiang Fang, Jie Ling, and N. Balakrishnan. Stochastic comparisons of series and parallel systems with independent heterogeneous lower-truncated Weibull components. Communications in Statistics: Theory and Methods, 45(2):540–551, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Flesia:2013:NDR

[Fle13] Ana Georgina Flesia. A note on distinguishing random trees populations. Communications in Statistics: Theory and Methods, 42(2):239–251, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feigin:2012:ACL

[FLK12] P. D. Feigin, Ya. P. Lumelskii, and A. V. Kryvoshaev. Approximate confidence limits for the difference between parameters of two Pólya distributions. Communications in Statistics: Theory and Methods, 41(8):1428–1454, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:2017:TSH

[FLL17] Jianglin Fang, Wanrong Liu, and Xuewen Lu. Two-sample high-dimensional empirical likelihood. Communications in Statistics: Theory and Methods, 46(13):6323–6335, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1115072.

Fan:2012:ELH

[FLW12] Guo-Liang Fan, Han-Ying Liang, and Jiang-Feng Wang. Empirical likelihood for a heteroscedastic partial linear errors-in-variables model. Communications in Statistics: Theory and Methods, 41(1):108–127, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2011:ELH

[FLX11] Guo-Liang Fan, Han-Ying Liang, and Hong-Xia Xu. Empirical likelihood for a heteroscedastic partial linear model. Communications in Statistics: Theory and Methods, 40(8):1396–1417, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Flynn:2010:CMS

[Fly10] Michael R. Flynn. Constrained maximization of the Shapiro– Wilk W statistic to estimate parameters of the Johnson S_{B} distribution. Communications in Statistics: Theory and Methods, 39(20):3635–3644, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2016:APB

[FLY16] Ke-Ang Fu, Jie Li, and Xiao-Rong Yang. Asymptotic properties of the bootstrap unit root test statistic under possibly infinite variance. Communications in Statistics: Theory and Methods, 45(11):3158–3167, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
901362.

Farzinnia:2010:BUC

[FM10a] Neda Farzinnia and Kevin F. McCardle. Bayesian updating with confounded signals. Communications in Statistics: Theory and Methods, 39(6):956–972, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fitzsimmons:2010:JFL

[FM10b] Patrick Fitzsimmons and Tucker McElroy. On joint Fourier– Laplace transforms. Communications in Statistics: Theory and Methods, 39(10):1883–1885, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fei:2011:SAP

[FM11] Lin Fei and Jie Mi. Some aspects of probability of correct selections. Communications in Statistics: Theory and Methods, 40(24):4484–4495, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Filipiak:2012:UOC

[FM12] Katarzyna Filipiak and Augustyn Markiewicz. On universal optimality of circular weakly neighbor balanced designs under an interference model. Communications in Statistics: Theory and Methods, 41(13–14):2356–2366, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fonseca:2013:DMM

[FM13] Miguel Fonseca and João Tiago Mexia. Delta method, moment convergence, and inference. Communications in Statistics: Theory and Methods, 42(12):2165–2171, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Filipiak:2014:OCB

[FM14] Katarzyna Filipiak and Augustyn Markiewicz. On the optimality of circular block designs under a mixed interference model. Communications in Statistics: Theory and Methods, 43(21): 4534–4545, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Farne:2016:DES

[FM16] Matteo Farné and Angela Montanari. Different estimators of the spectral matrix: an empirical comparison testing a new shrinkage estimator. Communications in Statistics: Theory and Methods, 45(2):354–364, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Filipiak:2017:UOD

[FM17] Katarzyna Filipiak and Augustyn Markiewicz. Universally optimal designs under interference models with and without block effects. Communications in Statistics: Theory and Methods, 46(3):1127–1143, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1011786.

Fonseca:2016:DMS

[FMPF16] C. Fonseca, A. P. Martins, L. Pereira, and H. Ferreira. Dependence matrices for spatial extreme events. Communications in Statistics: Theory and Methods, 45(21):6321–6341, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.781649.

Feng:2015:EAI

[FMS15] Yanqin Feng, Ling Ma, and Jianguo Sun. Empirical analysis of interval-censored failure time data with linear transformation models. Communications in Statistics: Theory and Methods, 44(24):5120–5135, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fornius:2009:UCD

[FN09] Ellinor Fackle Fornius and Hans Nyquist. Using the canonical design space to obtain c-optimal designs for the quadratic logistic model. Communications in Statistics: Theory and Methods, 39(1):144–157, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fazlollahtabar:2018:MBP

[FN18] Hamed Fazlollahtabar and Seyed Taghi Akhavan Niaki. Modified branching process for the reliability analysis of complex systems: Multiple-robot systems. Communications in Statistics: Theory and Methods, 47(7):1641–1652, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1324985.

Fatahi:2012:CBB

[FNDM12] Amir Afshin Fatahi, Rassoul Noorossana, Pershang Dokouhaki, and Babak Farhang Moghaddam. Copula-based bivariate ZIP control chart for monitoring rare events. Communications in Statistics: Theory and Methods, 41(15):2699–2716, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fallah:2019:FVC

[FNM19] Somayeh Fallah, Ali Reza Najafi, and Farshid Mehrdoust. A fractional version of the Cox–Ingersoll–Ross interest rate model and pricing double barrier option with Hurst index H ∈ (2/3, 1). Communications in Statistics: Theory and Methods, 48(9):2254–2266, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1464580.

Fontana:2014:FFD

[Fon14] Roberto Fontana. Fractional factorial designs for model-based evaluation of customer preferences. Communications in Statistics: Theory and Methods, 43(4):693–703, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fontana:2017:GMA

[Fon17] Roberto Fontana. Generalized minimum aberration mixed-level orthogonal arrays: a general approach based on sequential integer quadratically constrained quadratic programming. Communications in Statistics: Theory and Methods, 46(9):4275–4284, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1081947.

Fiorin:2012:LSC

[FP12] Silvano Fiorin and Francesca Parpinel. Least squares consistent estimates for arbitrary regression functions over an abstract space. Communications in Statistics: Theory and Methods, 41(16–17):2959–2982, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
647216.

Farnoosh:2017:SDE

[FP17a] Rahman Farnoosh and Mahmood Parsamanesh. Stochastic differential equation systems for an SIS epidemic model with vaccination and immigration. Communications in Statistics: Theory and Methods, 46(17):8723–8736, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1189571.

Forchini:2017:MFD

[FP17b] Giovanni Forchini and Bin Peng. Modified first-difference estimator in a panel data model with unobservable factors both in the errors and the regressors when the time dimension is small. Communications in Statistics: Theory and Methods, 46(24):12226–12239, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1295156.

Fei:2016:MCS

[FPCP16] Yu Fei, Yating Pan, Yin Chen, and Jianxin Pan. Modeling of covariance structures of random effects and random errors in linear mixed models. Communications in Statistics: Theory and Methods, 45(9):2748–2769, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1089290.

Fani:2019:NAO

[FPM19] S. Fani, M. A. Pasha, and M. Bameni Moghadam. A new approach to optimal design of control charts based on change point estimation and total time on test plot. Communications in Statistics: Theory and Methods, 48(22):5571–5584, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1515956.

Fernandez-Ponce:2016:NMA

[FPPRG16] J. M. Fernández-Ponce, F. Pellery, and M. R. Rodríguez-Griñolo. New multivariate aging notions based on the corrected orthant and the standard construction. Communications in Statistics: Theory and Methods, 45(10):2958–2974, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.894063.

Fiorentini:2017:MDM

[FPR17] Gabriele Fiorentini, Christophe Planas, and Alessandro Rossi. Marginal distribution of Markov-switching VAR processes. Communications in Statistics: Theory and Methods, 46(13): 6605–6623, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1132324.

Faliva:2016:OPT

[FPZ16] M. Faliva, V. Potì, and M. G. Zoia. Orthogonal polynomials for tailoring density functions to excess kurtosis, asymmetry, and dependence. Communications in Statistics: Theory and Methods, 45(1):49–62, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:2014:RVS

[FQZ14] Yali Fan, Guoyou Qin, and Zhong Yi Zhu. Robust variable selection in linear mixed models. Communications in Statistics: Theory and Methods, 43(21):4566–4581, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fakoor:2010:SGA

[FR10] V. Fakoor and N. Nakhaei Rad. Strong Gaussian approximations of product-limit and quantile processes for strong mixing and censored data. Communications in Statistics: Theory and Methods, 39(12):2271–2279, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Firinguetti:2016:NSR

[FRC16] Luis Firinguetti, Hernán Rubio, and Yogendra P. Chaubey. A non stochastic ridge regression estimator and comparison with the James–Stein estimator. Communications in Statistics: Theory and Methods, 45(8):2298–2310, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.879892.

Frey:2014:NFI

[Fre14] Jesse Frey. A note on Fisher information and imperfect rankedset sampling. Communications in Statistics: Theory and Methods, 43(13):2726–2733, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frolov:2017:ETI

[Fro17a] Andrei N. Frolov. On Esséen type inequalities for combinatorial random sums. Communications in Statistics: Theory and Methods, 46(12):5932–5940, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1115074.

From:2017:NPC

[Fro17b] Steven G. From. Non parametric confidence intervals for the extinction probability of a Galton–Watson branching process. Communications in Statistics: Theory and Methods, 46(3): 1202–1217, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1014108.

Fu:2017:MDS

[FS17] Ke-Ang Fu and Xinmei Shen. Moderate deviations for sums of dependent claims in a size-dependent renewal risk model. Communications in Statistics: Theory and Methods, 46(7):3235– 3243, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1060338.

Fan:2017:FAO

[FSSW17] Kun Fan, Yang Shen, Tak Kuen Siu, and Rongming Wang. An FFT approach for option pricing under a regime-switching stochastic interest rate model. Communications in Statistics: Theory and Methods, 46(11):5292–5310, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1100740.

Ferrani:2016:KDM

[FST16] Yacine Ferrani, Elias Ould Saïd, and Abdelkader Tatachak. On kernel density and mode estimates for associated and censored data. Communications in Statistics: Theory and Methods, 45(7):1853–1862, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
867996.

Fragkos:2017:EDE

[FTF17] Konstantinos C. Fragkos, Michail Tsagris, and Christos C. Frangos. Exploring the distribution for the estimator of Rosenthal’s ‘fail-safe’ number of unpublished studies in metaanalysis. Communications in Statistics: Theory and Methods, 46(11):5672–5684, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1109664.

Fan:2016:CQR

[FTT16] Yan Fan, Manlai Tang, and Maozai Tian. Composite quantile regression for varying-coefficient single-index models. Communications in Statistics: Theory and Methods, 45(10):3027–3047, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.894069.

Figini:2010:MAP

[FU10] Silvia Figini and Pierpaolo Uberti. Model assessment for predictive classification models. Communications in Statistics: Theory and Methods, 39(18):3238–3244, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2011:APS

[Fu11a] Ke-Ang Fu. Asymptotic properties of the R/S statistics for linear processes. Communications in Statistics: Theory and Methods, 40(18):3207–3217, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2011:EMC

[Fu11b] Ke-Ang Fu. Exact moment convergence rates of U-statistics. Communications in Statistics: Theory and Methods, 40(6): 1030–1040, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2012:ASN

[Fu12a] Ke-Ang Fu. An application of U-statistics to nonparametric functional data analysis. Communications in Statistics: Theory and Methods, 41(9):1532–1542, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2012:IPP

[Fu12b] Ke-Ang Fu. Invariance principles for products of U-statistics without variance. Communications in Statistics: Theory and Methods, 41(4):674–683, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2012:MSZ

[Fu12c] Ying-Zi Fu. Model selection of zero-inflated generalized power series distribution with missing responses. Communications in Statistics: Theory and Methods, 41(6):1013–1028, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fernandez:2011:GRM

[FV11] Antonio Fernández and Manuel Vázquez. A generalized regression methodology for bivariate heteroscedastic data. Communications in Statistics: Theory and Methods, 40(4):598–621, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Filipiak:2012:P

[FvR12] Katarzyna Filipiak and Dietrich von Rosen. Preface. Communications in Statistics: Theory and Methods, 41(13–14):2281– 2282, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feng:2015:STC

[FWH^{+}15] Changyong Feng, Hongyue Wang, Yu Han, Yinglin Xia, Naiji Lu, and Xin M. Tu. Some theoretical comparisons of negative binomial and zero-inflated Poisson distributions. Communications in Statistics: Theory and Methods, 44(15):3266–3277, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:2019:POR

[FWLL19] Ying Fang, Xia Wang, Hongli Liu, and Tong Li. Paretooptimal reinsurance for both the insurer and the reinsurer with general premium principles. Communications in Statistics: Theory and Methods, 48(24):6134–6154, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1528364.

Feng:2013:GMN

[FWT13] Changyong Feng, Hongyue Wang, and Xin M. Tu. Geometric mean of nonnegative random variable. Communications in Statistics: Theory and Methods, 42(15):2714–2717, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feng:2019:CCW

[FWW19] Feng-Xiang Feng, Ding-Cheng Wang, and Qun-Ying Wu. Complete convergence for weighted sums of negatively dependent random variables under the sub-linear expectations. Communications in Statistics: Theory and Methods, 48(6):1351–1366, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1429632.

Feng:2017:GDG

[FWZ^{+}17a] Changyong Feng, Hongyue Wang, Yun Zhang, Yu Han, Yuefeng Liang, and Xin M. Tu. Generalized definition of the geometric mean of a non negative random variable. Communications in Statistics: Theory and Methods, 46(7):3614–3620, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1066818.

Feng:2017:TPC

[FWZ^{+}17b] Changyong Feng, Hongyue Wang, Yun Zhang, Yu Han, Yuefeng Liang, and Xin M. Tu. On testing proportionality in the Cox regression model by Andersen’s plot. Communications in Statistics: Theory and Methods, 46(7):3489–3500, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1065330.

FallahNezhad:2017:OME

[FY17a] Mohammad Saber FallahNezhad and Ahmad Ahmadi Yazdi. An optimization model for economic design of sampling plans based on conforming run length considering outgoing quality. Communications in Statistics: Theory and Methods, 46(5): 2202–2211, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1035396.

Franko:2017:SSP

[FY17b] C. Franko and F. Yalcin. Signatures of series and parallel systems consisting of non disjoint modules. Communications in Statistics: Theory and Methods, 46(23):11425–11439, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1080839.

Fioruci:2016:EPL

[FYLC16] José Augusto Fioruci, Bao Yiqi, Francisco Louzada, and Vicente G. Cancho. The exponential Poisson logarithmic distribution. Communications in Statistics: Theory and Methods, 45(9):2556–2575, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
887106.

Feng:2010:UAL

[FZ10a] Yanqin Feng and Guoxin Zuo. A unified approach to likelihood inference on linear inequality constraints in contingency tables. Communications in Statistics: Theory and Methods, 39(19): 3399–3420, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:2010:CLB

[FZ10b] Ke-Ang Fu and Li-Xin Zhang. Characterization of LIL behavior for non-degenerate B-valued U-statistics. Communications in Statistics: Theory and Methods, 39(7):1258–1269, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fuming:2010:TBE

[FZ10c] Lin Fuming and Peng Zuoxiang. Tail behavior and extremes of short-tailed symmetric distribution. Communications in Statistics: Theory and Methods, 39(15):2811–2817, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fakoor:2017:DKD

[FZ17a] Vahid Fakoor and Raheleh Zamini. L_{p} distance for kernel density estimator in length-biased data. Communications in Statistics: Theory and Methods, 46(18):9247–9264, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1205621.

Fan:2017:RRM

[FZ17b] Chunpeng Fan and Donghui Zhang. Rank repeated measures analysis of covariance. Communications in Statistics: Theory and Methods, 46(3):1158–1183, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1014106.

Fan:2018:TFC

[FZ18] Chunpeng Fan and Donghui Zhang. Testing factor-covariate interaction in rank repeated-measures analysis of covariance models. Communications in Statistics: Theory and Methods, 47(11):2760–2778, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1342840.

Fan:2019:NTS

[FZS19] Xiaodong Fan, Shishun Zhao, and Jianguo Sun. Nonparametric test for stratum effects in the Cox model with interval-censored data. Communications in Statistics: Theory and Methods, 48 (2):300–310, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1408832.

Zhang:2019:CSN

[fZWD19] Tian fang Zhang, Guobin Wu, and Aloke Dey. Construction of some new families of nested orthogonal arrays. Communications in Statistics: Theory and Methods, 48(3):774–779, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1435816.

Fan:2016:PTL

[FZZ16] Caiyun Fan, Feipeng Zhang, and Yong Zhou. Powertransformed linear regression on quantile residual life for censored competing risks data. Communications in Statistics: Theory and Methods, 45(20):5884–5905, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.950873.

Ghosh:2018:NCG

[GA18] Indranil Ghosh and Ayman Alzaatreh. A new class of generalized logistic distribution. Communications in Statistics: Theory and Methods, 47(9):2043–2055, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.835420.

Gao:2014:POS

[GAK14] Hongjiang Gao, Inmaculada B. Aban, and Charles R. Katholi. Properties of a one-sided likelihood ratio test as applied to pool screening. Communications in Statistics: Theory and Methods, 43(21):4546–4565, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gallo:2015:TMC

[Gal15] Michele Gallo. Tucker3 model for compositional data. Communications in Statistics: Theory and Methods, 44(21):4441–4453, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garg:2010:PNM

[Gar10] Davinder Kumar Garg. Pseudo new modified latin square (NML_{m}(m)) type PBIB designs. Communications in Statistics: Theory and Methods, 39(19):3485–3491, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia:2012:SWM

[Gar12] Edel Garcia. The self-weighting model. Communications in Statistics: Theory and Methods, 41(8):1421–1427, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gargoum:2014:UHD

[Gar14] Ali S. Gargoum. On using Hellinger distance in checking the validity of dynamic approximations. Communications in Statistics: Theory and Methods, 43(19):4179–4186, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gangsei:2017:TEP

[GAS17] Lars Erik Gangsei, Trygve Almøy, and Solve Sæbø. Theoretical evaluation of prediction error in linear regression with a bivariate response variable containing missing data. Communications in Statistics: Theory and Methods, 46(20):9921–9929, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1222434.

Gaur:2017:CSD

[Gau17] Anil Gaur. A class of k-sample distribution-free tests for location against ordered alternatives. Communications in Statistics: Theory and Methods, 46(5):2343–2353, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1041986.

Guo:2012:AOS

[GB12] Jiin-Huarng Guo and Lynne Billard. Assessing one-step-ahead prediction error based on the median for first-order autoregressive models in the presence of outliers. Communications in Statistics: Theory and Methods, 41(15):2738–2749, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindaraju:2015:CRC

[GB15] K. Govindaraju and M. Bebbington. The consumer’s risk and costs in zero-acceptance number sampling. Communications in Statistics: Theory and Methods, 44(14):2933–2944, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghapani:2018:MLE

[GB18] F. Ghapani and B. Babadi. Mixed Liu estimator in linear measurement error models. Communications in Statistics: Theory and Methods, 47(7):1561–1570, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321768.

Gupta:2012:ENM

[GBC12] Arindam Gupta, Sabyasachi Bhattacharya, and Asis Kumar Chattyopadhyay. Exploring new models for population prediction in detecting demographic phase change for sparse census data. Communications in Statistics: Theory and Methods, 41(7):1171–1193, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goegebeur:2010:GKE

[GBD10] Yuri Goegebeur, Jan Beirlant, and Tertius De Wet. Generalized kernel estimators for the Weibull-tail coefficient. Communications in Statistics: Theory and Methods, 39(20):3695–3716, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grilo:2012:FNE

[GC12] Luís M. Grilo and Carlos A. Coelho. A family of near-exact distributions based on truncations of the exact distribution for the generalized Wilks Lambda statistic. Communications in Statistics: Theory and Methods, 41(13–14):2321–2341, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garrido:2014:HWN

[GC14] Liliana Garrido and Edilberto C. Cuervo. Heteroscedastic Weibull–normal mixture models: a Bayesian approach. Communications in Statistics: Theory and Methods, 43(2):249–265, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gorny:2019:THC

[GC19] Julian Górny and Erhard Cramer. Type-I hybrid censoring of uniformly distributed lifetimes. Communications in Statistics: Theory and Methods, 48(2):412–433, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1414255.

Ghorbanpour:2018:FIE

[GCA18] S. Ghorbanpour, R. Chinipardaz, and S. M. R. Alavi. Fisher information in exponential-weighted location-scale distributions. Communications in Statistics: Theory and Methods, 47(9): 2213–2226, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1337147.

Giancristofaro:2012:MRC

[GCR12] Rosa Arboretti Giancristofaro, Livio Corain, and Susanna Ragazzi. The multivariate randomized complete block design: a novel permutation solution in case of ordered categorical variables. Communications in Statistics: Theory and Methods, 41(16–17):3094–3109, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
601836.

Gomez:2010:PIS

[GCSB10] Héctor W. Gómez, Luis M. Castro, Hugo S. Salinas, and Heleno Bolfarine. Properties and inference on the skew-curved-symmetric family of distributions. Communications in Statistics: Theory and Methods, 39(5):884–898, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Geng:2019:UAN

[GCW19] Bingzhen Geng, Cen Chen, and Shijie Wang. Uniform asymptotics for a non standard renewal risk model with CLWD heavy-tailed claims. Communications in Statistics: Theory and Methods, 48(16):4051–4066, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1484484.

Giordan:2011:CMC

[GD11] Marco Giordan and Giancarlo Diana. A clustering method for categorical ordinal data. Communications in Statistics: Theory and Methods, 40(7):1315–1334, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goegebeur:2012:LES

[GD12] Yuri Goegebeur and Tertius De Wet. Local estimation of the second-order parameter in extreme value statistics and local unbiased estimation of the tail index. Communications in Statistics: Theory and Methods, 41(19):3575–3607, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Galea:2012:IAH

[GdC12] Manuel Galea and Mário de Castro. Influence assessment in an heteroscedastic errors-in-variables model. Communications in Statistics: Theory and Methods, 41(8):1350–1363, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Deniz:2016:PCL

[GDCO16] E. Gómez-Déniz and E. Calderín-Ojeda. The Poisson-conjugate Lindley mixture distribution. Communications in Statistics: Theory and Methods, 45(10):2857–2872, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.892134.

Gomez-Deniz:2013:GGI

[GDCOS13] Emilio Gómez-Déniz, Enrique Calderín-Ojeda, and José María Sarabia. Gamma-generalized inverse Gaussian class of distributions with applications. Communications in Statistics: Theory and Methods, 42(6):919–933, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Deniz:2019:RPG

[GDDC19] Emilio Gómez-Déniz and Nancy Dávila-Cárdenes. Reference points for getting a target in a soccer competition. a probabilistic model with applications to the Spanish and English Premier League. Communications in Statistics: Theory and Methods, 48(9):2075–2087, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1465083.

Gomez-Deniz:2017:BGG

[GDGG17] E. Gómez-Déniz, M. E. Ghitany, and Ramesh C. Gupta. A bivariate generalized geometric distribution with applications. Communications in Statistics: Theory and Methods, 46(11): 5453–5465, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1102285.

Gomez-Deniz:2017:MIG

[GDPR17] Emilio Gómez-Déniz and Jorge V. Pérez-Rodríguez. Mixture inverse Gaussian for unobserved heterogeneity in the autoregressive conditional duration model. Communications in Statistics: Theory and Methods, 46(18):9007–9025, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1200094.

Georgescu:2014:AMA

[GDS^{+}14] Vera Georgescu, Nicolas Desassis, Samuel Soubeyrand, André Kretzschmar, and Rachid Senoussi. An automated MCEM algorithm for hierarchical models with multivariate and multitype response variables. Communications in Statistics: Theory and Methods, 43(17):3698–3719, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomes:2015:EPM

[GDSC15a] Antonio E. Gomes, Cibele Q. Da-Silva, and Gauss M. Cordeiro. The exponentiated G Poisson model. Communications in Statistics: Theory and Methods, 44(20):4217–4240, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See correction [AE23b].

Gomes:2015:TEB

[GdSC15b] Antonio E. Gomes, Cibele Q. da Silva, and Gauss M. Cordeiro. Two extended Burr models: Theory and practice. Communications in Statistics: Theory and Methods, 44(8):1706–1734, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2018:BCP

[GDW18] Jie Guo, Yinghui Dong, and Guojing Wang. Basket CDS pricing with default intensities using a regime-switching shotnoise model. Communications in Statistics: Theory and Methods, 47(18):4443–4458, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1376087.

Gurka:2011:EVC

[GE11] Matthew J. Gurka and Lloyd J. Edwards. Estimating variance components and random effects using the Box–Cox transformation in the linear mixed model. Communications in Statistics: Theory and Methods, 40(3):515–531, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghalibaf:2010:ADC

[GEA^{+}10] M. Bolbolian Ghalibaf, N. Eghbal, M. Amini, H. A. Azarnoosh, and A. Bozorgnia. Aspects of dependence in Cuadras–Auge family. Communications in Statistics: Theory and Methods, 39(11):2094–2107, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gebregziabher:2018:JMM

[GEM^{+}18] Mulugeta Gebregziabher, Mark A. Eckert, Lois J. Matthews, Abeba A. Teklehaimanot, and Judy R. Dubno. Joint modeling of multivariate hearing thresholds measured longitudinally at multiple frequencies. Communications in Statistics: Theory and Methods, 47(22):5418–5434, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1395045.

Genc:2017:ACB

[Gen17] Ali I. Genç. An absolutely continuous bivariate Topp–Leone distribution: a useful model on a bounded domain. Communications in Statistics: Theory and Methods, 46(19):9726–9742, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1218026.

Genc:2019:TSP

[Gen19] Ali I. Genç. The two-sided power-binomial distribution for modeling binary count data. Communications in Statistics: Theory and Methods, 48(10):2548–2561, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472778.

Georgiadis:2017:FHP

[Geo17] Stylianos Georgiadis. First hitting probabilities for semi-Markov chains and estimation. Communications in Statistics: Theory and Methods, 46(5):2435–2446, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1045080.

Gey:2018:VCD

[Gey18] Servane Gey. Vapnik–Chervonenkis dimension of axis-parallel cuts. Communications in Statistics: Theory and Methods, 47(9):2291–2296, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1339088.

Gontscharuk:2017:AGF

[GF17] Veronika Gontscharuk and Helmut Finner. Asymptotics of goodness-of-fit tests based on minimum p-value statistics. Communications in Statistics: Theory and Methods, 46(5):2332– 2342, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1041985.

Ghalibaf:2010:ABL

[GFA10] M. Bolbolian Ghalibaf, V. Fakoor, and H. A. Azarnoosh. Asymptotic behaviors of the Lorenz curve for censored data under strong mixing. Communications in Statistics: Theory and Methods, 40(2):323–332, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giles:2013:BML

[GFG13] David E. Giles, Hui Feng, and Ryan T. Godwin. On the bias of the maximum likelihood estimator for the two-parameter Lomax distribution. Communications in Statistics: Theory and Methods, 42(11):1934–1950, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giles:2016:BCM

[GFG16] David E. Giles, Hui Feng, and Ryan T. Godwin. Biascorrected maximum likelihood estimation of the parameters of the generalized Pareto distribution. Communications in Statistics: Theory and Methods, 45(8):2465–2483, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.887104.

Granzotto:2019:LBI

[GFL19] Daniele C. T. Granzotto, Paulo H. Ferreira, and Francisco Louzada. Likelihood-based inference for the transmuted loglogistic model in the presence of right-censored data. Communications in Statistics: Theory and Methods, 48(7):1798–1813, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1440313.

Golbafian:2017:NES

[GFM17] V. Golbafian, M. S. Fallahnezhad, and Y. Zare Mehrjerdi. A new economic scheme for CCC charts with run rules based on average number of inspected items. Communications in Statistics: Theory and Methods, 46(24):12023–12044, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1291967.

Giancristofaro:2012:CNT

[GFR12] Rosa Arboretti Giancristofaro, Roberto Fontana, and Susanna Ragazzi. Construction and nonparametric testing of orthogonal arrays through algebraic strata and inequivalent permutation matrices. Communications in Statistics: Theory and Methods, 41(16–17):3162–3178, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2011.579380.

Garthwaite:2016:AOS

[GFS16] P. H. Garthwaite, Y. Fan, and S. A. Sisson. Adaptive optimal scaling of Metropolis–Hastings algorithms using the Robbins–Monro process. Communications in Statistics: Theory and Methods, 45(17):5098–5111, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.936562.

Galea:2010:ETE

[GG10a] Manuel Galea and Patricia Giménez. Estimation and testing in elliptical functional measurement error models. Communications in Statistics: Theory and Methods, 39(11):2031–2045, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guyon:2010:EDB

[GG10b] Frédéric Guyon and Alain Guénoche. An evolutionary distance based on maximal unique matches. Communications in Statistics: Theory and Methods, 39(3):385–397, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goudie:2013:MLE

[GG13] I. B. J. Goudie and R. Gormley. Maximum likelihood estimates for the Schnabel census with plants. Communications in Statistics: Theory and Methods, 42(20):3704–3715, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gokpinar:2016:MMW

[GG16a] Esra Gökpinar and Fikri Gökpinar. A modified Mack–Wolfe test for the umbrella alternative problem. Communications in Statistics: Theory and Methods, 45(24):7226–7241, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.975821.

Gunduz:2016:DPP

[GG16b] Selim Gündüz and Ali I. Genç. Distribution of the product of a pair of independent two-sided power variates. Communications in Statistics: Theory and Methods, 45(20):6149–6164, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.957861.

Ganji:2018:DDN

[GG18] M. Ganji and F. Gharari. The discrete delta and nabla Mittag-Leffler distributions. Communications in Statistics: Theory and Methods, 47(18):4568–4589, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1377254.

Garcia:2016:MOF

[GGDVP16] V. J. García, E. Gómez-Déniz, and F. J. Vázquez-Polo. A Marshall–Olkin family of heavy-tailed distributions which includes the lognormal one. Communications in Statistics: Theory and Methods, 45(7):2023–2044, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.873132.

Gonzalez:2011:IPC

[GGFL11] Luis G. González, Graciela González-Farías, and M. Vittoria Levati. Identifying preferences for conditional cooperation using individual beliefs. Communications in Statistics: Theory and Methods, 40(17):3099–3118, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia:2014:MAS

[GGL14] Jesús E. García and V. A. González-López. Modeling of acoustic signal energies with a generalized Frank copula. A linguistic conjecture is reviewed. Communications in Statistics: Theory and Methods, 43(10–12):2034–2044, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia:2014:RMS

[GGLV14] Jesús E. García, V. A. González-López, and M. L. L. Viola. Robust model selection for stochastic processes. Communications in Statistics: Theory and Methods, 43(10–12):2516–2526, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goegebeur:2017:LMT

[GGO17] Yuri Goegebeur, Armelle Guillou, and Michael Osmann. A local moment type estimator for an extreme quantile in regression with random covariates. Communications in Statistics: Theory and Methods, 46(1):319–343, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.991039.

Gill:2013:MTD

[GGS13] Amar Nath Gill, Anju Goyal, and Parminder Singh. On a multiple three-decision problem for comparing several treatments with the best control. Communications in Statistics: Theory and Methods, 42(19):3432–3454, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunn:2019:MEM

[GGtH^{+}19] Laura H. Gunn, Henryk Gzyl, Enrique ter Horst, Miller Janny Ariza, and German Molina. Maximum entropy in the mean methods in propensity score matching for interval and noisy data. Communications in Statistics: Theory and Methods, 48(18):4581–4597, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1497656.

Ge:2013:EGE

[GH13] Yonghui Ge and Xiaohui Han. The extent of Gross errors eliminated by robust multiple linear regressions. Communications in Statistics: Theory and Methods, 42(23):4210–4221, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:2018:GKD

[GH18] Indranil Ghosh and G. G. Hamedani. The gamma-Kumaraswamy distribution: an alternative to gamma distribution. Communications in Statistics: Theory and Methods, 47(9):2056–2072, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1122055. See correction [AE23a].

Gao:2019:GSF

[GH19] Zhongqin Gao and Jingmin He. The Gerber–Shiu function for the compound Poisson Omega model with a three-step premium rate. Communications in Statistics: Theory and Methods, 48(24):6019–6037, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1524488.

Ghapani:2019:ULE

[Gha19] Fatemeh Ghapani. Using Liu estimator for detection of influential observations in linear measurement error models. Communications in Statistics: Theory and Methods, 48(19):4748–4763, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1475567.

Ghosh:2018:MWP

[GHBM18] I. Ghosh, G. G. Hamedani, N. Bansal, and M. Maadooliat. On the mixtures of Weibull and Pareto (IV) distribution: an alternative to Pareto distribution. Communications in Statistics: Theory and Methods, 47(9):2073–2084, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1171353.

Guo:2017:EUC

[GHL17] Lisha Guo, X. Joan Hu, and Yanyan Liu. Estimation under Cox proportional hazards model with covariates missing not at random. Communications in Statistics: Theory and Methods, 46(18):8952–8972, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1197252.

Ghorbel:2009:IDD

[Gho09] M. Ghorbel. On the inverted Dirichlet distribution. Communications in Statistics: Theory and Methods, 39(1):21–37, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:2015:CSG

[Gho15] Sucharita Ghosh. Computation of spatial Gini coefficients. Communications in Statistics: Theory and Methods, 44(22): 4709–4720, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:2017:EMD

[Gho17] Sucharita Ghosh. On estimating the marginal distribution of a detrended series with long memory. Communications in Statistics: Theory and Methods, 46(23):11539–11557, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1275698.

Ghribi:2019:UEG

[Ghr19] Abdelaziz Ghribi. On the UMVU estimator for the generalized variance of the multinomial family. Communications in Statistics: Theory and Methods, 48(24):5943–5952, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1520888.

Gao:2019:SLL

[GHSZ19] Miaomiao Gao, Feng Hu, Jingbo Sun, and Zhaojun Zong. A strong law of large number for negatively dependent and non identical distributed random variables in the framework of sublinear expectation. Communications in Statistics: Theory and Methods, 48(20):5058–5073, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1508708.

Gosho:2011:CSW

[GHY11] Masahiko Gosho, Chikuma Hamada, and Isao Yoshimura. Criterion for the selection of a working correlation structure in the generalized estimating equation approach for longitudinal balanced data. Communications in Statistics: Theory and Methods, 40(21):3839–3856, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giancristofaro:2014:PSM

[Gia14] Rosa Arboretti Giancristofaro. Permutation solutions for multivariate ranking and testing with applications. Communications in Statistics: Theory and Methods, 43(4):891–905, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guerrero:2017:TEM

[GICRR17] Victor M. Guerrero, Alejandro Islas-Camargo, and L. Leticia Ramirez-Ramirez. Trend estimation of multivariate time series with controlled smoothness. Communications in Statistics: Theory and Methods, 46(13):6704–6726, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1133826.

Giles:2012:BRM

[Gil12] David E. Giles. Bias reduction for the maximum likelihood estimators of the parameters in the half-logistic distribution. Communications in Statistics: Theory and Methods, 41(2):212– 222, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gillard:2014:MME

[Gil14] Jonathan Gillard. Method of moments estimation in linear regression with errors in both variables. Communications in Statistics: Theory and Methods, 43(15):3208–3222, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gzyl:2014:FMM

[GIT14] Henryk Gzyl, Pier Luigi Novi Inverardi, and Aldo Tagliani. Fractional moments and maximum entropy: Geometric meaning. Communications in Statistics: Theory and Methods, 43 (17):3596–3601, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:2015:RWS

[GJ15] Qingwu Gao and Na Jin. Randomly weighted sums of pairwise quasi upper-tail independent increments with application to risk theory. Communications in Statistics: Theory and Methods, 44(18):3885–3902, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Graja:2014:IES

[GJM14] Asma Graja, Aida Jarraya, and Afif Masmoudi. Implicit estimation for the stochastic volatility model. Communications in Statistics: Theory and Methods, 43(6):1061–1076, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gurler:2011:AHR

[GK11] Ülkü Gürler and Paul Kvam. Adjusted hazard rate estimator based on a known censoring probability. Communications in Statistics: Theory and Methods, 40(24):4409–4416, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goroncy:2013:RCC

[GK13] Agnieszka Goroncy and Udo Kamps. Random convex combinations of m-generalized order statistics. Communications in Statistics: Theory and Methods, 42(17):3076–3084, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grover:2019:CGA

[GK19a] Lovleen Kumar Grover and Amanpreet Kaur. Corrigendum to: Gupta et al. (2012). Estimation of the mean of a sensitive variable in the presence of auxiliary information. CSTM, 41(13–14):2394–2404. Communications in Statistics: Theory and Methods, 48(12):3162–3163, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1475570. See [GSSCR12].

Guerbyenne:2019:PPT

[GK19b] Hafida Guerbyenne and Abderrahim Kessira. Power periodic threshold GARCH model: Structure and estimation. Communications in Statistics: Theory and Methods, 48(19):4834–4860, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1496258.

Guo:2019:NLW

[GK19c] Huijun Guo and Junke Kou. Non linear wavelet estimation of regression derivatives based on biased data. Communications in Statistics: Theory and Methods, 48(13):3219–3235, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1473613.

Golparvar:2016:BRB

[GKP16] Leila Golparvar, Ali Karimnezhad, and Ahmad Parsian. Bayes and robust Bayes predictions in a subfamily of scale parameters under a precautionary loss function. Communications in Statistics: Theory and Methods, 45(13):3970–3992, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.915041.

Gorecki:2014:FDC

[GKW14] Tomasz Górecki, Mirosław Krzyśko, and Łukasz Waszak. Functional discriminant coordinates. Communications in Statistics: Theory and Methods, 43(5):1013–1025, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gathy:2010:MPD

[GL10a] Maude Gathy and Claude Lefèvre. On Markov–Pólya distribution and the Katz family of distributions. Communications in Statistics: Theory and Methods, 40(2):267–278, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2010:SMP

[GL10b] Guangbao Guo and Shaoling Lin. Schwarz method for penalized quasi-likelihood in generalized additive models. Communications in Statistics: Theory and Methods, 39(10):1847–1854, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gu:2011:FDR

[GL11] Wentao Gu and Soyeon Lee. On fixed design regression for general linear processes. Communications in Statistics: Theory and Methods, 40(23):4233–4248, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Georgiadis:2015:NES

[GL15a] Stylianos Georgiadis and Nikolaos Limnios. Nonparametric estimation of the stationary distribution of a discrete-time semi-Markov process. Communications in Statistics: Theory and Methods, 44(7):1319–1337, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Granzotto:2015:TLL

[GL15b] Daniele Cristina Tita Granzotto and Francisco Louzada. The transmuted log-logistic distribution: Modeling, inference, and an application to a polled tabapua race time up to first calving data. Communications in Statistics: Theory and Methods, 44 (16):3387–3402, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Girardin:2019:EBG

[GL19] Valérie Girardin and Justine Lequesne. Entropy-based goodness-of-fit tests — a unifying framework: Application to DNA replication. Communications in Statistics: Theory and Methods, 48(1):62–74, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1401084.

Giordano:2014:IVS

[GLP14] Francesco Giordano, Michele La Rocca, and Cira Perna. Input variable selection in neural network models. Communications in Statistics: Theory and Methods, 43(4):735–750, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gouet:2012:CLT

[GLS12] Raúl Gouet, F. Javier López, and Gerardo Sanz. Central limit theorem for the number of near-records. Communications in Statistics: Theory and Methods, 41(2):309–324, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gombay:2017:RCD

[GLY17] Edit Gombay, Fuxiao Li, and Hao Yu. Retrospective change detection in categorical time series. Communications in Statistics: Theory and Methods, 46(14):6831–6845, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1137595.

Gao:2012:APM

[GLZW12] Qi-Bing Gao, Jin-Guan Lin, Chun-Hua Zhu, and Yao-Hua Wu. Asymptotic properties of maximum quasi-likelihood estimates in generalized linear models with “Working” covariance matrix and adaptive designs. Communications in Statistics: Theory and Methods, 41(19):3544–3561, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giuliano:2014:LDP

[GM14] Rita Giuliano and Claudio Macci. Large deviation principles for sequences of maxima and minima. Communications in Statistics: Theory and Methods, 43(6):1077–1098, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garg:2013:EMD

[GMB^{+}13] Lalit Garg, Sally McClean, Maria Barton, Brian Meenan, and Ken Fullerton. An extended mixture distribution survival tree for patient pathway prognostication. Communications in Statistics: Theory and Methods, 42(16):2912–2934, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Graiche:2016:TCV

[GMH16] F. Graiche, D. Merabet, and D. Hamadouche. Testing change in the variance with epidemic alternatives. Communications in Statistics: Theory and Methods, 45(13):3822–3837, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.907024.

Ghitany:2019:UIG

[GMMA19] M. E. Ghitany, J. Mazucheli, A. F. B. Menezes, and F. Alqallaf. The unit-inverse Gaussian distribution: a new alternative to two-parameter distributions on the unit interval. Communications in Statistics: Theory and Methods, 48(14):3423–3438, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1476717.

Gomes:2013:GJB

[GMN13] M. Ivette Gomes, M. João Martins, and M. Manuela Neves. Generalized jackknife-based estimators for univariate extremevalue modeling. Communications in Statistics: Theory and Methods, 42(7):1227–1245, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomes:2011:ARB

[GMP11] M. Ivette Gomes, Sandra Mendonça, and Dinis Pestana. Adaptive reduced-bias tail index and VaR estimation via the bootstrap methodology. Communications in Statistics: Theory and Methods, 40(16):2946–2968, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorbani-Mandolakani:2016:MBE

[GMR16] Mehrzad Ghorbani-Mandolakani and Mohammad Reza Salehi Rad. ML and Bayes estimation in a two-phase tandem queue with a second optional service and random feedback. Communications in Statistics: Theory and Methods, 45(9):2576–2591, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.887107.

Ghorbani-Mandolakani:2013:EPT

[GMRN13] Mehrzad Ghorbani-Mandolakani, Mohammad Reza Salehi Rad, and Nader Nematollahi. Estimation of the parameters of a twophase tandem queue with a second optional service. Communications in Statistics: Theory and Methods, 42(17):3031–3052, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guessoum:2018:APK

[GMS18] Zohra Guessoum, Mohamed-Amine Mansouri, and Elias Ould Saïd. Asymptotic properties of the kernel mode estimator under twice censorship model. Communications in Statistics: Theory and Methods, 47(9):2195–2212, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1337143.

Gupta:2013:GSQ

[GMSD13] Sat Gupta, Samridhi Mehta, Javid Shabbir, and B. K. Dass. Generalized scrambling in quantitative optional randomized response models. Communications in Statistics: Theory and Methods, 42(22):4034–4042, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govaert:2010:LBM

[GN10] Gérard Govaert and Mohamed Nadif. Latent block model for contingency table. Communications in Statistics: Theory and Methods, 39(3):416–425, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2012:SBB

[GN12] Arjun K. Gupta and Daya K. Nagar. Some bimatrix beta distributions. Communications in Statistics: Theory and Methods, 41(5):869–879, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:2015:IHT

[GN15] Indranil Ghosh and Saralees Nadarajah. Inference for a hidden truncated (both-sided) bivariate Pareto (II) distribution. Communications in Statistics: Theory and Methods, 44(10): 2136–2150, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:2016:AAC

[GN16] Indranil Ghosh and Saralees Nadarajah. An alternative approach for compatibility of two discrete conditional distributions. Communications in Statistics: Theory and Methods, 45(15):4416–4432, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
921302.

Ghosh:2017:BIK

[GN17] Indranil Ghosh and Saralees Nadarajah. On the Bayesian inference of Kumaraswamy distributions based on censored samples. Communications in Statistics: Theory and Methods, 46(17): 8760–8777, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1193197.

Gokpinar:2012:HTE

[GÖ12] Fikri Gökpinar and Yaprak Arzu Özdemir. A Horvitz– Thompson estimator of the population mean using inclusion probabilities of ranked set sampling. Communications in Statistics: Theory and Methods, 41(6):1029–1039, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Godwin:2016:BRM

[God16] Ryan T. Godwin. Bias reduction for the maximum likelihood estimator of the doubly-truncated Poisson distribution. Communications in Statistics: Theory and Methods, 45(7):1887– 1901, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2013.867999.

Golia:2016:PCS

[Gol16] Silvia Golia. A proposal for categorizing the severity of non uniform differential item functioning: The polytomous case. Communications in Statistics: Theory and Methods, 45(2):236– 251, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grajales:2019:DRE

[GOLM19] Luis Fernando Grajales, Raydonal Ospina, Luis A. López, and Oscar O. Melo. A doubly restricted exponential dispersion model. Communications in Statistics: Theory and Methods, 48(11):2827–2841, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473427.

Goroncy:2017:UNP

[Gor17] Agnieszka Goroncy. Upper non positive bounds on expectations of generalized order statistics from DD and DDA populations. Communications in Statistics: Theory and Methods, 46(24):11972–11987, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1288250.

Govindaraju:2011:ZAN

[Gov11] K. Govindaraju. Zero acceptance number chained quick switching system. Communications in Statistics: Theory and Methods, 40(12):2104–2116, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gijbels:2012:FMD

[GP12] I. Gijbels and I. Prosdocimi. Flexible mean and dispersion function estimation in extended generalized additive models. Communications in Statistics: Theory and Methods, 41(16–17): 3259–3277, 2012. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/03610926.2012.654881.

Ghosh:2016:NBC

[GP16a] Santu Ghosh and Alan M. Polansky. New bootstrap confidence intervals for means of positively skewed distributions. Communications in Statistics: Theory and Methods, 45(23):6915–6927, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.972569.

Golparvar:2016:IPH

[GP16b] Leila Golparvar and Ahmad Parsian. Inference on proportional hazard rate model parameter under Type-I progressively hybrid censoring scheme. Communications in Statistics: Theory and Methods, 45(24):7258–7274, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.978020.

Giurcanu:2019:OGE

[GP19] Mihai Giurcanu and Brett Presnell. Oracle GMM estimation for misspecified models via thresholding. Communications in Statistics: Theory and Methods, 48(11):2657–2686, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472786.

Gopinath:2018:IRC

[GPM18] Pratheesh P. Gopinath, Rajender Parsad, and Baidya Nath Mandal. Incomplete row-column designs with factorial treatment structure for estimating main effects with full efficiency. Communications in Statistics: Theory and Methods, 47(18): 4493–4502, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1376091.

Guler:2014:BTL

[GPÖ14] Nesrın Güler, Simo Puntanen, and Halım Özdemir. On the BLUEs in two linear models via C. R. Rao’s Pandora’s Box. Communications in Statistics: Theory and Methods, 43(5):921– 931, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guida:2019:BAN

[GPP19] Maurizio Guida, Fabio Postiglione, and Gianpaolo Pulcini. A Bayesian approach for non-homogeneous gamma degradation processes. Communications in Statistics: Theory and Methods, 48(7):1748–1765, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1440306.

Gou:2017:NES

[GQC17] Tingxun Gou, Hong Qin, and Kashinath Chatterjee. A new extension strategy on three-level factorials under wrap-around L_{2}-discrepancy. Communications in Statistics: Theory and Methods, 46(18):9063–9074, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1202284.

Gou:2018:SUP

[GQC18] Tingxun Gou, Hong Qin, and Kashinath Chatterjee. A study of uniformity pattern for extended designs. Communications in Statistics: Theory and Methods, 47(12):2869–2880, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1343843.

Gideon:2011:LSE

[GR11] Rudy Gideon and Adele Marie Rothan. Location and scale estimation with correlation coefficients. Communications in Statistics: Theory and Methods, 40(9):1561–1572, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guria:2015:MVL

[GR15] Sibnarayan Guria and Sugata Sen Roy. Missing values in linear regression: Imputations using an error-contaminated linear predictor. Communications in Statistics: Theory and Methods, 44(8):1735–1744, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia:2017:SRE

[GR17] J. García and D. E. Ramirez. The successive raising estimator and its relation with the ridge estimator. Communications in Statistics: Theory and Methods, 46(22):11123–11142, 2017.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1260738.

Goroncy:2018:RSU

[GR18a] Agnieszka Goroncy and Tomasz Rychlik. Refined solution to upper-bound problem for the expectations of order statistics from decreasing density on the average distributions. Communications in Statistics: Theory and Methods, 47(16):4029–4041, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1367813.

Groenvik:2018:SNA

[GR18b] Henriette Groenvik and Yeonwoo Rho. A self-normalizing approach to the specification test of mixed-frequency models. Communications in Statistics: Theory and Methods, 47(8): 1913–1922, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1332222.

Grau:2010:NCI

[Gra10] Daniel Grau. New capability indices for non normal processes. Communications in Statistics: Theory and Methods, 39(16): 2913–2929, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grau:2011:PYC

[Gra11] Daniel Grau. Process yield and capability indices. Communications in Statistics: Theory and Methods, 40(15):2751–2771, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Graczyk:2012:OSB

[Gra12] Małgorzata Graczyk. A-optimal spring balance weighing designs under some conditions. Communications in Statistics: Theory and Methods, 41(13–14):2386–2393, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grabchak:2017:SCM

[Gra17] Michael Grabchak. A simple condition for the multivariate CLT and the attraction to the Gaussian of Lévy processes at long and short times. Communications in Statistics: Theory and Methods, 46(1):446–456, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
995826.

Grientschnig:2015:RPP

[Gri15] Dieter Grientschnig. Reference priors for Poisson processes involving nuisance parameters. Communications in Statistics: Theory and Methods, 44(23):4893–4911, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gribkova:2017:CTM

[Gri17] Nadezhda Gribkova. Cramér-type moderate deviations for intermediate trimmed means. Communications in Statistics: Theory and Methods, 46(23):11918–11932, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1285930.

Grover:2010:CNI

[Gro10] Lovleen Kumar Grover. A correction note on improvement in variance estimation using auxiliary information. Communications in Statistics: Theory and Methods, 39(5):753–764, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Groeneveld:2011:IFC

[Gro11] Richard A. Groeneveld. Influence functions for the coefficient of variation, its inverse, and CV comparisons. Communications in Statistics: Theory and Methods, 40(23):4139–4150, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Groenitz:2018:LRA

[Gro18] Heiko Groenitz. Logistic regression analyses for indirect data. Communications in Statistics: Theory and Methods, 47(16): 3838–3856, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1364387.

Goldoust:2018:LDM

[GRSN18] M. Goldoust, S. Rezaei, Y. Si, and S. Nadarajah. A lifetime distribution motivated by parallel and series structures. Communications in Statistics: Theory and Methods, 47(13):3052–3072, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1346802.

Gruber:2012:LRE

[Gru12] Marvin H. J. Gruber. Liu and ridge estimators — a comparison. Communications in Statistics: Theory and Methods, 41(20): 3739–3749, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grilli:2015:BMM

[GRV15] Leonardo Grilli, Carla Rampichini, and Roberta Varriale. Binomial mixture modeling of University credits. Communications in Statistics: Theory and Methods, 44(22):4866–4879, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gani:2011:IDP

[GS11] J. Gani and R. J. Swift. An immigration–death process with a time-dependent delay. Communications in Statistics: Theory and Methods, 40(7):1236–1240, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghodsi:2013:SPN

[GS13] Alireza Ghodsi and Mahendran Shitan. Some properties of the normalized periodogram of a Fractionally Integrated Separable Spatial ARMA (FISSARMA) model. Communications in Statistics: Theory and Methods, 42(8):1515–1530, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghodsi:2015:AFF

[GS15a] Alireza Ghodsi and Mahendran Shitan. Autocovariance function of the fractionally integrated separable spatial ARMA (FISSARMA) models. Communications in Statistics: Theory and Methods, 44(5):933–941, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guerrero:2015:STS

[GS15b] Victor M. Guerrero and Eliud Silva. Smoothing a time series by segments of the data range. Communications in Statistics:

Theory and Methods, 44(21):4568–4585, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2015:EFP

[GS15c] Sat Gupta and Javid Shabbir. Estimation of finite population mean in stratified random sampling with two auxiliary variables under double sampling design. Communications in Statistics: Theory and Methods, 44(13):2798–2808, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garthoff:2017:MMC

[GS17] Robert Garthoff and Wolfgang Schmid. Monitoring means and covariances of multivariate non linear time series with heavy tails. Communications in Statistics: Theory and Methods, 46(21):10394–10415, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1085567.

Gustafsson:2019:VSF

[GS19] Oskar Gustafsson and Pär Stockhammar. Variance stabilizing filters. Communications in Statistics: Theory and Methods, 48(24):6155–6168, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1528369.

Goual:2014:CST

[GSA14] Hafida Goual and Nacira Seddik-Ameur. Chi-squared type test for the AFT-generalized inverse Weibull distribution. Communications in Statistics: Theory and Methods, 43(13):2605–2617, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghodsi:2012:FOS

[GSB12] Alireza Ghodsi, Mahendran Shitan, and Hassan S. Bakouch. A first-order spatial integer-valued autoregressive SINAR(1, 1) model. Communications in Statistics: Theory and Methods, 41(15):2773–2787, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2016:PTD

[GSLZ16] Guangbao Guo, Wei Shao, Lu Lin, and Xuehu Zhu. Parallel tempering for dynamic generalized linear models. Communications in Statistics: Theory and Methods, 45(21):6299–6310, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.960586.

Garg:2016:DPT

[GSM16] Mridula Garg, Ajay Sharma, and Pratibha Manohar. The distribution of the product of two independent generalized trapezoidal random variables. Communications in Statistics: Theory and Methods, 45(21):6369–6384, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.882954.

Gupta:2012:EMS

[GSSCR12] Sat Gupta, Javid Shabbir, Rita Sousa, and Pedro Corte-Real. Estimation of the mean of a sensitive variable in the presence of auxiliary information. Communications in Statistics: Theory and Methods, 41(13–14):2394–2404, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See corrigendum [GK19a].

Gan:2017:EPU

[GST17] Shengjun Gan, Yuqin Sun, and Yongge Tian. Equivalence of predictors under real and over-parameterized linear models. Communications in Statistics: Theory and Methods, 46(11): 5368–5383, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1100742.

Gannoun:2010:SMR

[GSY10] Ali Gannoun, Jerome Saracco, and Keming Yu. On semiparametric mode regression estimation. Communications in Statistics: Theory and Methods, 39(7):1141–1157, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:2015:BAC

[GT15] Cheng Gao and Hiroki Tsurumi. Bayesian analysis of the censored regression model with an AEPD error term. Communications in Statistics: Theory and Methods, 44(5):953–971, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gou:2017:EPF

[GT17] Jiangtao Gou and Ajit C. Tamhane. Estimation of a parametric function associated with the lognormal distribution. Communications in Statistics: Theory and Methods, 46(16):8134–8154, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1175628.

George:2019:KMO

[GT19] Roshini George and S. Thobias. Kumaraswamy Marshall-Olkin exponential distribution. Communications in Statistics: Theory and Methods, 48(8):1920–1937, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1440594.

Garg:2017:MLD

[GTW17] Vikram V. Garg, Luis Tenorio, and Karen Willcox. Minimum local distance density estimation. Communications in Statistics: Theory and Methods, 46(1):148–164, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.988260.

Guermeur:2010:SCC

[Gue10] Yann Guermeur. Sample complexity of classifiers taking values in R^{Q}, application to multi-class SVMs. Communications in Statistics: Theory and Methods, 39(3):543–557, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guermeur:2013:CMC

[Gue13] Yann Guermeur. Combining multi-class SVMs with linear ensemble methods that estimate the class posterior probabilities. Communications in Statistics: Theory and Methods, 42(16): 3011–3030, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guerry:2014:SRE

[Gue14] Marie-Anne Guerry. Some results on the embeddable problem for discrete-time Markov models in manpower planning. Communications in Statistics: Theory and Methods, 43(7):1575– 1584, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gui:2015:AHN

[Gui15] Wenhao Gui. An alpha half normal slash distribution for analyzing non negative data. Communications in Statistics: Theory and Methods, 44(22):4783–4795, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gunasekera:2016:BIO

[Gun16] Sumith Gunasekera. Bayesian inference for the offered optical network unit load. Communications in Statistics: Theory and Methods, 45(10):2890–2919, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
892136.

Gunasekera:2017:BIC

[Gun17] Sumith Gunasekera. Bayesian inference for a common scale parameter of several Pareto populations. Communications in Statistics: Theory and Methods, 46(6):2780–2800, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1048890.

Guo:2014:ECC

[Guo14] Ming Le Guo. Equivalent conditions of complete moment convergence of weighted sums for φ-mixing sequence of random variables. Communications in Statistics: Theory and Methods, 43(10–12):2527–2539, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2014:SCS

[Gup14] Bhupendra Gupta. On secure communication in a sensor network. Communications in Statistics: Theory and Methods, 43 (21):4642–4649, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2015:PRF

[Gup15] Pushpa L. Gupta. Properties of reliability functions of discrete distributions. Communications in Statistics: Theory and Methods, 44(19):4114–4131, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2016:RCF

[Gup16] Ramesh C. Gupta. Reliability characteristics of Farlie– Gumbel–Morgenstern family of bivariate distributions. Communications in Statistics: Theory and Methods, 45(8):2342– 2353, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2013.828075.

Gliklikh:2011:DMQ

[GV11] Yuri E. Gliklikh and Natalia V. Vinokurova. On the description of motion of a quantum particle in the classical gauge field in the language of stochastic mechanics. Communications in Statistics: Theory and Methods, 40(19–20):3630–3640, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Villegas:2011:BAH

[GVGP11] Miguel A. Gómez-Villegas and Beatriz González-Pérez. A Bayesian analysis for the homogeneity testing problem using ϵ-contaminated priors. Communications in Statistics: Theory and Methods, 40(6):1049–1062, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gou:2015:MCC

[GW15a] Jianwei Gou and Jinde Wang. Multiple comparisons with control under stochastic ordering: Controlling FDR. Communications in Statistics: Theory and Methods, 44(21):4507–4520, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2015:UAE

[GW15b] Fenglong Guo and Dingcheng Wang. Uniform asymptotic estimates for ruin probabilities with exponential lévy process investment returns and two-sided linear heavy-tailed claims. Communications in Statistics: Theory and Methods, 44(20): 4278–4306, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:2018:ASD

[GW18] Ramesh C. Gupta and Muhammad Waleed. Analysis of survival data by a Weibull–Bessel distribution. Communications in Statistics: Theory and Methods, 47(4):980–995, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1316402.

Gunasekera:2019:GCL

[GW19] Sumith Gunasekera and Danush K. Wijekularathna. Generalized confidence limits for the performance index of the exponentially distributed lifetime. Communications in Statistics: Theory and Methods, 48(3):755–773, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1435810.

Gajda:2019:FLS

[GWK19] Janusz Gajda, Agnieszka Wylomanska, and Arun Kumar. Fractional Lévy stable motion time-changed by gamma subordinator. Communications in Statistics: Theory and Methods, 48(24):5953–5968, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1523430.

Gao:2016:PRC

[GWL16] Jianwei Gao, Liyuan Wu, and Huihui Liu. On the probability of ruin in a continuous risk model with two types of delayed claims. Communications in Statistics: Theory and Methods, 45(13):3734–3750, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
909939.

Guo:2019:TAD

[GWP19] Fenglong Guo, Dingcheng Wang, and Jiangyan Peng. Tail asymptotic of discounted aggregate claims with compound dependence under risky investment. Communications in Statistics: Theory and Methods, 48(4):810–830, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1417437.

Guo:2017:PLD

[GWZ17] Hui Guo, Shijie Wang, and Chuanwei Zhang. Precise large deviations of aggregate claims in a compound sizedependent renewal risk model. Communications in Statistics: Theory and Methods, 46(3):1107–1116, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1010011.

Guo:2018:CBE

[GX18] Donglin Guo and Liugen Xue. CBPS-based estimation for linear models with responses missing at random. Communications in Statistics: Theory and Methods, 47(17):4160–4169, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1371752.

Goliforushani:2018:MRL

[GXB18] S. Goliforushani, M. Xie, and N. Balakrishnan. On the mean residual life of a generalized k-out-of-n system. Communications in Statistics: Theory and Methods, 47(10):2362–2372, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.988266.

Gu:2013:SNP

[GY13] Bei-Qing Gu and Rong-Xian Yue. Some notes on parameter estimation for generalized exponential distribution. Communications in Statistics: Theory and Methods, 42(10):1787–1805, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:2017:RVS

[GYL17] Chaohui Guo, Hu Yang, and Jing Lv. Robust variable selection for generalized linear models with a diverging number of parameters. Communications in Statistics: Theory and Methods, 46(6):2967–2981, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1053940.

Gao:2016:CIR

[GZ16] Xiang Gao and Ming Zheng. Causal inference for recurrent events data with all-or-none compliance. Communications in Statistics: Theory and Methods, 45(24):7306–7325, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.980515.

Gao:2017:ECE

[GZ17] Xiang Gao and Ming Zheng. Estimating the causal effects in randomized trials for survival data with a cure fraction and non compliance. Communications in Statistics: Theory and Methods, 46(8):4065–4087, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1076481.

Gao:2014:RWE

[GZSS14] Shesheng Gao, Yongmin Zhong, Chunmeng Sang, and Bijan Shirinzadeh. Random weighting estimation for quantile processes and negatively associated samples. Communications in Statistics: Theory and Methods, 43(3):656–662, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hossain:2014:PSE

[HA14] Shakhawat Hossain and S. Ejaz Ahmed. Penalized and shrinkage estimation in the Cox proportional hazards model. Communications in Statistics: Theory and Methods, 43(5):1026–1040, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Halima:2016:EIL

[HA16a] Zerouati Halima and Dahmani Abdelnasser. Exponential inequalities in linear calibration problem. Communications in Statistics: Theory and Methods, 45(18):5251–5262, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.941497.

Harandi:2016:GLF

[HA16b] S. Shams Harandi and M. H. Alamatsaz. Generalized linear failure rate power series distribution. Communications in Statistics: Theory and Methods, 45(8):2204–2227, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.878357.

Hesamian:2017:STB

[HA17] Gholamreza Hesamian and Mohamad Ghasem Akbari. Statistical test based on intuitionistic fuzzy hypotheses. Communications in Statistics: Theory and Methods, 46(18):9324–9334, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1208241.

Hadjila:2018:ESC

[HA18] Tabti Hadjila and Ait Saidi Ahmed. Estimation and simulation of conditional hazard function in the quasi-associated framework when the observations are linked via a functional single-index structure. Communications in Statistics: Theory and Methods, 47(4):816–838, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1213294.

Hackmann:2018:KLE

[Hac18] Daniel Hackmann. Karhunen-loève expansions of lévy processes. Communications in Statistics: Theory and Methods, 47(23):5675–5687, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1400057.

Hadjikyriakou:2017:SRD

[Had17] Milto Hadjikyriakou. Some results on dependent random variables and a connection with the multivariate s-increasing convex order. Communications in Statistics: Theory and Methods, 46(24):12087–12102, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1291977.

Hajrajabi:2019:ANS

[Haj19] A. Hajrajabi. Analysis of nonlinear state space model with dependent measurement noises. Communications in Statistics: Theory and Methods, 48(9):2088–2101, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1459706.

Hamdan:2010:MDF

[Ham10] Hasan Hamdan. Modeling data with flat densities. Communications in Statistics: Theory and Methods, 39(4):722–728, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamdi:2012:CEF

[Ham12] Fayçal Hamdi. Computing the exact Fisher information matrix of periodic state-space models. Communications in Statistics: Theory and Methods, 41(22):4182–4199, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamedani:2013:SIE

[Ham13] G. G. Hamedani. Sub-independence: an expository perspective. Communications in Statistics: Theory and Methods, 42(20): 3615–3638, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamaz:2019:PRF

[Ham19] Abdelghani Hamaz. Prediction of random fields with incomplete quarter-plane past. Communications in Statistics: Theory and Methods, 48(11):2707–2716, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472789.

Han:2011:BEH

[Han11] Ming Han. E-Bayesian estimation and hierarchical Bayesian estimation of failure probability. Communications in Statistics: Theory and Methods, 40(18):3303–3314, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2012:BCL

[Han12a] Ming Han. The M-Bayesian credible limits of the reliability derived from binomial distribution. Communications in Statistics: Theory and Methods, 41(21):3814–3830, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanif:2012:LLE

[Han12b] Muhammad Hanif. Local linear estimation of recurrent jumpdiffusion models. Communications in Statistics: Theory and Methods, 41(22):4142–4163, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanif:2015:NPE

[Han15] Muhammad Hanif. Non parametric estimation of second-order diffusion equation by using asymmetric kernels. Communications in Statistics: Theory and Methods, 44(9):1896–1910, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2017:BHB

[Han17] Ming Han. The E-Bayesian and hierarchical Bayesian estimations for the system reliability parameter. Communications in Statistics: Theory and Methods, 46(4):1606–1620, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1024861.

Han:2019:BEE

[Han19] Ming Han. E-Bayesian estimation of the exponentiated distribution family parameter under LINEX loss function. Communications in Statistics: Theory and Methods, 48(3):648–659, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1417432.

Hazen:2017:AMB

[HAP17] Gordon Hazen, Daniel Apley, and Neehar Parikh. ANOVA models for Brownian motion. Communications in Statistics: Theory and Methods, 46(15):7642–7660, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1158834.

Haq:2017:TSC

[Haq17] Abdul Haq. Two-stage cluster sampling with hybrid ranked set sampling in the secondary sampling frame. Communications in Statistics: Theory and Methods, 46(17):8450–8467, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1183783.

Haq:2018:NAE

[Haq18] Abdul Haq. A new adaptive EWMA control chart using auxiliary information for monitoring the process mean. Communications in Statistics: Theory and Methods, 47(19):4840–4858, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1448417.

Haridas:2017:MCD

[Har17] Haritha N. Haridas. Modeling censored data using modified lambda family. Communications in Statistics: Theory and Methods, 46(8):3838–3847, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1073316.

Haselimashhadi:2019:UCP

[Has19] Hamed Haselimashhadi. A unified class of penalties with the capability of producing a differentiable alternative to l_{1} norm penalty. Communications in Statistics: Theory and Methods, 48(22):5530–5545, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1515362.

Helu:2010:BEW

[HASA10] Amal Helu, Muhammad Abu-Salih, and Osama Alkam. Bayes estimation of Weibull distribution parameters using ranked set sampling. Communications in Statistics: Theory and Methods, 39(14):2533–2551, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayashi:2010:TNR

[Hay10] Masahito Hayashi. Two non-regular extensions of large deviation bound. Communications in Statistics: Theory and Methods, 39(8–9):1404–1423, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayter:2016:WPC

[Hay16] A. J. Hayter. Win-probabilities for comparing two Poisson variables. Communications in Statistics: Theory and Methods, 45(20):5966–5976, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
953693.

Hayakawa:2017:URT

[Hay17] Kazuhiko Hayakawa. Unit root test for short panels with serially correlated errors. Communications in Statistics: Theory and Methods, 46(8):3891–3900, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1076471.

Hussain:2016:MSE

[HAZ16] Zawar Hussain and Bander Al-Zahrani. Mean and sensitivity estimation of a sensitive variable through additive scrambling. Communications in Statistics: Theory and Methods, 45(1):182– 193, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:2013:ECU

[HB13] Zhulin He and Babette A. Brumback. An equivalence of conditional and unconditional maximum likelihood estimators via infinite replication of observations. Communications in Statistics: Theory and Methods, 42(18):3267–3279, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:2016:MBS

[HB16] David D. Hanagal and Susmita M. Bhambure. Modeling bivariate survival data using shared inverse Gaussian frailty model. Communications in Statistics: Theory and Methods, 45(17): 4969–4987, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.901380.

Hanagal:2017:MAT

[HB17a] David D. Hanagal and Susmita M. Bhambure. Modeling Australian twin data using shared positive stable frailty models based on reversed hazard rate. Communications in Statistics: Theory and Methods, 46(8):3754–3771, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1071395.

Hanagal:2017:SGF

[HB17b] David D. Hanagal and Susmita M. Bhambure. Shared gamma frailty models based on reversed hazard rate for modeling Australian twin data. Communications in Statistics: Theory and Methods, 46(12):5812–5826, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1116581.

Hussain:2017:SRS

[HBAJ17] Jaffer Hussain, S. Balamurali, Muhammad Aslam, and Chi-Hyuck Jun. SkSP-R sampling plan based on process capability index. Communications in Statistics: Theory and Methods, 46(6):2955–2966, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1053939.

Haq:2016:IBL

[HBM16] Abdul Haq, Jennifer Brown, and Elena Moltchanova. Improved best linear unbiased estimators for the simple linear regression model using double ranked set sampling schemes. Communications in Statistics: Theory and Methods, 45(12):3541–3561, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.904350.

Haq:2016:BLU

[HBMAO16a] Abdul Haq, Jennifer Brown, Elena Moltchanova, and Amer Ibrahim Al-Omari. Best linear unbiased and invariant estimation in location-scale families based on double-ranked set sampling. Communications in Statistics: Theory and Methods, 45(1):25– 48, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:2016:PDR

[HBMAO16b] Abdul Haq, Jennifer Brown, Elena Moltchanova, and Amer Ibrahim Al-Omari. Paired double-ranked set sampling. Communications in Statistics: Theory and Methods, 45(10):2873–2889, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.892135.

Huang:2010:EDM

[HC10] Ching-Chun Huang and Fei-Long Chen. Economic design of max charts. Communications in Statistics: Theory and Methods, 39(16):2961–2976, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2012:QCA

[HC12] Yaoguang Hu and Xi Cheng. A quality control approach of PVC gloves based on the integration of SPC and EPC. Communications in Statistics: Theory and Methods, 41(24):4470–4483, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2016:GLL

[HC16a] Feng Hu and Zengjing Chen. General laws of large numbers under sublinear expectations. Communications in Statistics: Theory and Methods, 45(14):4215–4229, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.917677.

Huang:2016:TBG

[HC16b] Jianwen Huang and Shouquan Chen. Tail behavior of the generalized Maxwell distribution. Communications in Statistics: Theory and Methods, 45(14):4230–4236, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.917678.

Human:2010:NST

[HCS10] S. W. Human, S. Chakraborti, and C. F. Smit. Nonparametric Shewhart-type sign control charts based on runs. Communications in Statistics: Theory and Methods, 39(11):2046–2062, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2018:SAR

[HCW18] Danyang Huang, Xiangyu Chang, and Hansheng Wang. Spatial autoregression with repeated measurements for social networks. Communications in Statistics: Theory and Methods, 47(15): 3715–3727, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1361989.

Cai:2018:SNC

[hCXxShY18] Guang hui Cai, Lin Xiang, Xin xing Su, and Xue hai Ying. A self-normalized central limit theorem for a ρmixing stationary sequence. Communications in Statistics: Theory and Methods, 47(8):1785–1791, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1277754.

Hu:2019:CDE

[HCZT19] XueLong Hu, Philippe Castagliola, XiaoJian Zhou, and AnAn Tang. Conditional design of the EWMA median chart with estimated parameters. Communications in Statistics: Theory and Methods, 48(8):1871–1889, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1440310.

Hanagal:2013:MIG

[HD13a] David D. Hanagal and Alok D. Dabade. Modeling of inverse Gaussian frailty model for bivariate survival data. Communications in Statistics: Theory and Methods, 42(20):3744–3769, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ho:2013:VSI

[HD13b] Linda Lee Ho and Roberto Da Costa Quinino. The Variable Sampling Interval (VSI) in an analysis of Taguchi’s on-line quality monitoring procedure for variables. Communications in Statistics: Theory and Methods, 42(6):974–987, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:2015:CSF

[HD15] David D. Hanagal and Alok D. Dabade. Comparison of shared frailty models for kidney infection data under exponential power baseline distribution. Communications in Statistics: Theory and Methods, 44(23):5091–5108, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hashempour:2017:BIM

[HD17] Majid Hashempour and Mahdi Doostparast. Bayesian inference on multiply sequential order statistics from heterogeneous exponential populations with GLR test for homogeneity. Communications in Statistics: Theory and Methods, 46(16):8086–8100, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1175625.

He:2016:JML

[HDS16] Sui He, Ting Du, and Liuquan Sun. Joint modeling of longitudinal data with a dependent terminal event. Communications in Statistics: Theory and Methods, 45(3):813–835, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:2017:BMC

[He17] Chaobing He. Bayesian multiple change-point estimation for exponential distribution with truncated and censored data. Communications in Statistics: Theory and Methods, 46(12): 5827–5839, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1161797.

He:2018:CLR

[He18] Zhulin He. Conditional logistic regression in cluster-specific 1: m matched treatment-control designs. Communications in Statistics: Theory and Methods, 47(9):2134–2145, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1337141.

Hassan:2016:BSS

[HEB16] M. Y. Hassan and M. Y. El-Bassiouni. Bimodal skewsymmetric normal distribution. Communications in Statistics: Theory and Methods, 45(5):1527–1541, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.882950.

Herrador:2011:FHM

[HEHM11] M. Herrador, M. D. Esteban, T. Hobza, and D. Morales. A Fay– Herriot model with different random effect variances. Communications in Statistics: Theory and Methods, 40(5):785–797, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henchiri:2017:NPL

[Hen17] Yousri Henchiri. Non parametric learning approach to estimate conditional quantiles in the dependent functional data case. Communications in Statistics: Theory and Methods, 46(9): 4369–4387, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1082592.

Herbertsson:2014:PEC

[HF14] Alexander Herbertsson and Rüdiger Frey. Parameter estimation in credit models under incomplete information. Communications in Statistics: Theory and Methods, 43(7):1409–1436, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Houser:2019:CSD

[HF19] Madelyn Houser and Pak-Wing Fok. Convergence of sums of dependent Bernoulli random variables: an application from portfolio theory. Communications in Statistics: Theory and Methods, 48(23):5673–5681, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1517891.

Fan:2014:CSA

[hFmH14] Tsai hung Fan and Tsung ming Hsu. Constant stress accelerated life test on a multiple-component series system under Weibull lifetime distributions. Communications in Statistics: Theory and Methods, 43(10–12):2370–2383, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

H:2019:ETL

[HG19] Rimsha H. and Dais George. q-Esscher transformed Laplace distribution. Communications in Statistics: Theory and Methods, 48(7):1563–1578, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1435812.

Hu:2015:APR

[HGZG15] Gaoge Hu, Shesheng Gao, Yongmin Zhong, and Chengfan Gu. Asymptotic properties of random weighted empirical distribution function. Communications in Statistics: Theory and Methods, 44(18):3812–3824, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herzberg:2013:TFS

[HH13] Frederik Herzberg and Angélique Herzberg. Testing fiscal sustainability via strong convergence of Gaussian random variables. Communications in Statistics: Theory and Methods, 42 (18):3235–3242, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hao:2015:APW

[HH15] Li Hao and Daniel Houser. Adaptive procedures for the Wilcoxon–Mann–Whitney test: Seven decades of advances. Communications in Statistics: Theory and Methods, 44(9): 1939–1957, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamad:2016:NBP

[HH16] Naqvi Hamad and Muhammad Hanif. Non binary partially neighbor balanced designs for circular blocks. Communications in Statistics: Theory and Methods, 45(20):5961–5965, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.953692.

Hayakawa:2019:ETV

[HH19] Kazuhiko Hayakawa and Jie Hou. Estimation of time-varying coefficient dynamic panel data models. Communications in Statistics: Theory and Methods, 48(13):3311–3324, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1476704.

Huang:2016:TEC

[HHC16] Lele Huang, Tao Hu, and Hengjian Cui. The T -type estimate of a class of partially non linear models. Communications in Statistics: Theory and Methods, 45(4):976–999, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.844253.

Huang:2017:AMC

[HHC^{+}17] Jui-Chieh Huang, Wen-Tso Huang, Pei-Tzu Chu, Wen-Yi Lee, Hsin-Ping Pai, Chih-Chen Chuang, and Ya-Wen Wu. Applying a Markov chain for the stock pricing of a novel forecasting model. Communications in Statistics: Theory and Methods, 46(9):4388–4402, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1083108.

Han:2018:CPL

[HHS18] Miao Han, Dongxiao Han, and Liuquan Sun. A class of partially linear transformation models for recurrent gap times. Communications in Statistics: Theory and Methods, 47(3):739–766, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1313986.

He:2014:CCC

[HHW14] Shuguang He, Zhen He, and G. Alan Wang. CUSUM control charts for multivariate Poisson distribution. Communications in Statistics: Theory and Methods, 43(6):1192–1208, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hilow:2015:MCL

[Hil15] Hisham Hilow. Minimum cost linear trend free 2^{n−(n−k)} designs of resolution IV. Communications in Statistics: Theory and Methods, 44(15):3158–3172, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Himeno:2010:TEM

[Him10] Tetsuto Himeno. A test for the equality of means of two groups with different variances when the sample size and the dimension are large. Communications in Statistics: Theory and Methods, 39(11):2081–2093, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hashorva:2014:AFT

[HJ14] Enkelejd Hashorva and Lanpeng Ji. Asymptotics of the finitetime ruin probability for the Sparre Andersen risk model perturbed by an inflated stationary chi-process. Communications in Statistics: Theory and Methods, 43(10–12):2540–2548, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2017:VBA

[HJC17] Chien-Chia Liäm Huang, Yow-Jen Jou, and Hsun-Jung Cho. VIF-based adaptive matrix perturbation method for heteroskedasticity-robust covariance estimators in the presence of multicollinearity. Communications in Statistics: Theory and Methods, 46(7):3255–3263, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1060340.

Huang:2019:CMS

[HJT19] Lei Huang, Hui Jiang, and Haitao Tian. The consistency of model selection for dynamic semi-varying coefficient models with autocorrelated errors. Communications in Statistics: Theory and Methods, 48(3):549–558, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1414265.

He:2019:HOD

[HJW19] Xuyang He, Yuexiang Jiang, and Jiazhen Wang. High-order data sharpening with dependent errors for regression bias reduction. Communications in Statistics: Theory and Methods, 48(23):5748–5755, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1520885.

Huang:2018:SHM

[HJY18] Mian Huang, Qinghua Ji, and Weixin Yao. Semiparametric hidden Markov model with non-parametric regression. Communications in Statistics: Theory and Methods, 47(21):5196–5204, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1388398.

Hudgens:2011:OCS

[HK11] Michael G. Hudgens and Hae-Young Kim. Optimal configuration of a square array group testing algorithm. Communications in Statistics: Theory and Methods, 40(3):436–448, 2011.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hrabakova:2013:CRM

[HK13] J. Hrabáková and V. Kus. The consistency and robustness of modified Cramér–von Mises and Kolmogorov–Cramér estimators. Communications in Statistics: Theory and Methods, 42 (20):3665–3677, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hashimoto:2015:BTI

[HK15] Shintaro Hashimoto and Ken-Ichi Koike. Bhattacharyya-type information inequality for the Bayes risk. Communications in Statistics: Theory and Methods, 44(24):5213–5224, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamad:2019:HME

[HK19a] Farag Hamad and Nezamoddin N. Kachouie. A hybrid method to estimate the full parametric hazard model. Communications in Statistics: Theory and Methods, 48(22):5477–5491, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1513149.

Haq:2019:NNP

[HK19b] Abdul Haq and Michael B. C. Khoo. A new non-parametric multivariate EWMA sign control chart for monitoring process dispersion. Communications in Statistics: Theory and Methods, 48(15):3703–3716, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1476708.

Hussain:2019:NIM

[HK19c] Zawar Hussain and Akbar Ali Khan. A new index for measuring evenness. Communications in Statistics: Theory and Methods, 48(2):354–367, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1410715.

Haq:2017:NEF

[HKH17] Abdul Haq, Manzoor Khan, and Zawar Hussain. A new estimator of finite population mean based on the dual use of the auxiliary information. Communications in Statistics: Theory and Methods, 46(9):4425–4436, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1083112.

Hayter:2012:IPM

[HKLW12] A. J. Hayter, S. Kiatsupaibul, W. Liu, and H. P. Wynn. An independence point method of confidence band construction for multiple linear regression models. Communications in Statistics: Theory and Methods, 41(22):4132–4141, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayter:2015:SIC

[HKNL15] A. J. Hayter, S. Kiatsupaibul, P. Napalai, and W. Liu. Simultaneous inferences on the cumulative distribution function of a normal distribution. Communications in Statistics: Theory and Methods, 44(24):5136–5145, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hadji:2015:TDR

[HKR15] Ehsan Moghimi Hadji, Nirmal Singh Kambo, and Alagar Rangan. Two-dimensional renewal function approximation. Communications in Statistics: Theory and Methods, 44(15):3107– 3124, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Horova:2012:VBM

[HKV12] Ivana Horová, Jan Kolácek, and Kamila Vopatová. Visualization and bandwidth matrix choice. Communications in Statistics: Theory and Methods, 41(4):759–777, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hong:2011:IGN

[HL11] Zhaoping Hong and Heng Lian. Inference of genetic networks from time course expression data using functional regression with lasso penalty. Communications in Statistics: Theory and Methods, 40(10):1768–1779, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hassairi:2013:MRD

[HL13a] Abdelhamid Hassairi and Mahdi Louati. Mixture of the Riesz distribution with respect to a multivariate Poisson. Communications in Statistics: Theory and Methods, 42(6):1124–1140, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huber:2013:SEP

[HL13b] Nina Huber and Hannes Leeb. Shrinkage estimators for prediction out-of-sample: Conditional performance. Communications in Statistics: Theory and Methods, 42(7):1246–1264, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hui:2015:PAC

[HL15] Jiang Hui and Yu Lei. Precise asymptotics in complete moment convergence of parameter estimator in the Gaussian autoregressive process. Communications in Statistics: Theory and Methods, 44(7):1483–1496, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hashorva:2016:MSE

[HL16] Enkelejd Hashorva and Chengxiu Ling. Maxima of skew elliptical triangular arrays. Communications in Statistics: Theory and Methods, 45(12):3692–3705, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
906613.

Hsieh:2017:ETC

[HL17] Jin-Jian Hsieh and Zih-Jyun Li. Estimation and test of conditional Kendall’s Tau under bivariate left-truncation data. Communications in Statistics: Theory and Methods, 46(13):6635– 6644, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1132326.

Hang:2018:NPR

[HL18a] Yin Hang and Shu Liu. Non parametric regression analysis for longitudinal data with time-depending autoregressive error process. Communications in Statistics: Theory and Methods, 47(18):4503–4533, 2018. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1377251.

Huang:2018:EAM

[HL18b] Wen-Tso Huang and Cheng-Chang Lu. An enhanced absorbing Markov chain model for predicting TAIEX index futures. Communications in Statistics: Theory and Methods, 47(1):133–146, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1300281.

Hsiao:2015:MAR

[HLL15] Ya-Chun Hsiao, Chen-Tuo Liao, and Li-Yu D. Liu. Minimum aberration regular two-level designs in the presence of dispersion factors. Communications in Statistics: Theory and Methods, 44(13):2764–2773, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hou:2017:HOE

[HLP17] Jingyao Hou, Xin Liao, and Zuoxiang Peng. Higher-order expansions of extremes from mixed skew-t distribution. Communications in Statistics: Theory and Methods, 46(20):9949–9971, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1228972.

Husler:2011:EGP

[HLR11] Jürg Hüsler, Deyuan Li, and Mathias Raschke. Estimation for the generalized Pareto distribution using maximum likelihood and goodness of fit. Communications in Statistics: Theory and Methods, 40(14):2500–2510, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Husler:2016:EVI

[HLR16] Jürg Hüsler, Deyuan Li, and Mathias Raschke. Extreme value index estimator using maximum likelihood and moment estimation. Communications in Statistics: Theory and Methods, 45(12):3625–3636, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
861495.

Huang:2014:AOS

[HLX14] Yu Huang, Yongling Li, and Maochao Xu. Analysis of order statistics from distributions with regularly varying tails. Communications in Statistics: Theory and Methods, 43(13):2702– 2713, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2016:LMP

[HLY16] Guikai Hu, Qingguo Li, and Shenghua Yu. Linear minimax prediction of finite population regression coefficient under a balanced loss function. Communications in Statistics: Theory and Methods, 45(24):7197–7209, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.978945.

Higuchi:2012:MMO

[HM12] Isao Higuchi and Toshio Mikami. Maxima and minima of overall survival functions with fixed marginal distributions and transmission of technology. Communications in Statistics: Theory and Methods, 41(1):46–61, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huda:2015:DMS

[HM15] S. Huda and R. M’Hallah. D-minimax second-order designs over hypercubes for extrapolation and restricted interpolation regions. Communications in Statistics: Theory and Methods, 44(21):4600–4613, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:2012:GFT

[HME12] Norbert Henze, Simos G. Meintanis, and Bruno Ebner. Goodness-of-fit tests for the gamma distribution based on the empirical Laplace transform. Communications in Statistics: Theory and Methods, 41(9):1543–1556, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Heydari:2017:EBR

[HME17] A. A. Heydari, M. B. Moghadam, and F. Eskandari. An extension of Banerjee and Rahim model in economic and economic statistical designs for multivariate quality characteristics under Burr XII distribution. Communications in Statistics: Theory and Methods, 46(16):7855–7871, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1140782.

Hauke:2012:CBU

[HMP12] Jan Hauke, Augustyn Markiewicz, and Simo Puntanen. Comparing the BLUEs under two linear models. Communications in Statistics: Theory and Methods, 41(13–14):2405–2418, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:2017:NPM

[HMT^{+}17] Fanyin He, Sati Mazumdar, Gong Tang, Triptish Bhatia, Stewart J. Anderson, Mary Amanda Dew, Robert Krafty, Vishwajit Nimgaonkar, Smita Deshpande, Martica Hall, and Charles F. Reynolds III. Non-parametric MANOVA approaches for nonnormal multivariate outcomes with missing values. Communications in Statistics: Theory and Methods, 46(14):7188–7200, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1146767.

Han:2014:ACB

[HN14] David Han and H. K. T. Ng. Asymptotic comparison between constant-stress testing and step-stress testing for Type-I censored data from exponential distribution. Communications in Statistics: Theory and Methods, 43(10–12):2384–2394, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hazra:2016:SGO

[HN16] Nil Kamal Hazra and Asok K. Nanda. On some generalized orderings: In the spirit of relative ageing. Communications in Statistics: Theory and Methods, 45(20):6165–6181, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.957862.

Hazra:2017:GSA

[HN17a] Nil Kamal Hazra and Asok K. Nanda. General standby allocation in series and parallel systems. Communications in Statistics: Theory and Methods, 46(19):9842–9858, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1222436.

Hyodo:2017:OSL

[HN17b] Masashi Hyodo and Takahiro Nishiyama. A one-sample location test based on weighted averaging of two test statistics when the dimension and the sample size are large. Communications in Statistics: Theory and Methods, 46(7):3526–3541, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1066812.

Huong:2017:PSJ

[HNB17] Pham Thi Thu Huong, Darfiana Nur, and Alan Branford. Penalized spline joint models for longitudinal and time-to-event data. Communications in Statistics: Theory and Methods, 46(20):10294–10314, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1235195.

Haidari:2019:CBP

[HNB19] Abedin Haidari, Amir T. Payandeh Najafabadi, and Narayanaswamy Balakrishnan. Comparisons between parallel systems with exponentiated generalized gamma components. Communications in Statistics: Theory and Methods, 48(6):1316–1332, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1429630.

Hosseini-Nasab:2013:PBC

[HNV13] S. Mohammad E. Hosseini-Nasab and Masoumeh Vakili. Percentile-t bootstrap confidence intervals for period using periodogram. Communications in Statistics: Theory and Methods, 42(17):3053–3064, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ho:2016:FPP

[Ho16] Chih-Hsiang Ho. Forecasting a point process with an ARIMA model. Communications in Statistics: Theory and Methods, 45(17):5066–5076, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
936560.

Haghbin:2017:NGO

[HOAH17] Hossein Haghbin, Gamze Ozel, Morad Alizadeh, and G. G. Hamedani. A new generalized odd log-logistic family of distributions. Communications in Statistics: Theory and Methods, 46(20):9897–9920, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1222428.

Hodoshima:2014:PLF

[Hod14] Jiro Hodoshima. On the properties of the likelihood function of Spanos’ conditional t heteroskedastic model. Communications in Statistics: Theory and Methods, 43(3):578–598, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayes:2017:DBB

[HOK17] Kevin Hayes, Kevin O’Brien, and Anthony Kinsella. A decomposition of the Bradley–Blackwood paired-samples omnibus test. Communications in Statistics: Theory and Methods, 46(20):9892–9896, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1222439.

Holgersson:2015:NCU

[Hol15] H. E. T. Holgersson. A note on a commonly used ridge regression Monte Carlo design. Communications in Statistics: Theory and Methods, 44(10):2176–2179, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Homei:2017:CAR

[Hom17] Hazhir Homei. Characterizations of arcsin and related distributions based on a new generalized unimodality. Communications in Statistics: Theory and Methods, 46(2):1024–1030, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1006788.

Hong:2017:ECP

[Hon17] Liang Hong. Estimators of contingent probabilities and means with actuarial applications. Communications in Statistics: Theory and Methods, 46(2):927–941, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1010003.

Hothorn:2016:TSA

[Hot16] Ludwig A. Hothorn. The two-step approach — a significant ANOVA F -test before Dunnett’s comparisons against a control — is not recommended. Communications in Statistics: Theory and Methods, 45(11):3332–3343, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.902225.

Hou:2012:RPA

[Hou12] Fujun Hou. Rank preserved aggregation rules and application to reliability allocation. Communications in Statistics: Theory and Methods, 41(21):3831–3845, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hui:2010:LSD

[HP10] Jun Hui and Guangming Pan. Limiting spectral distribution for large sample covariance matrices with m-dependent elements. Communications in Statistics: Theory and Methods, 39(6):935– 941, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hui:2014:RHR

[HP14] Stefen Hui and C. J. Park. The representation of hypergeometric random variables using independent Bernoulli random variables. Communications in Statistics: Theory and Methods, 43(19):4103–4108, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:2016:GSF

[HP16] David D. Hanagal and Arvind Pandey. Gamma shared frailty model based on reversed hazard rate. Communications in Statistics: Theory and Methods, 45(7):2071–2088, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.870204.

Hanagal:2017:SFM

[HP17a] David D. Hanagal and Arvind Pandey. Shared frailty models based on reversed hazard rate for modified inverse Weibull distribution as baseline distribution. Communications in Statistics: Theory and Methods, 46(1):234–246, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.990102.

Hanagal:2017:SIG

[HP17b] David D. Hanagal and Arvind Pandey. Shared inverse Gaussian frailty models based on additive hazards. Communications in Statistics: Theory and Methods, 46(22):11143–11162, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1260740.

Huang:2015:SLR

[HQW15] Jinlong Huang, Chunjuan Qiu, and Xianyi Wu. Stochastic loss reserving in discrete time: Individual vs. aggregate data models. Communications in Statistics: Theory and Methods, 44 (10):2180–2206, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2017:OME

[HR17] Alan Huang and Paul J. Rathouz. Orthogonality of the mean and error distribution in generalized linear models. Communications in Statistics: Theory and Methods, 46(7):3290–3296, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.851241.

Huang:2011:EGL

[HS11] Yujie Huang and Lixin Song. M-estimator of a generalized linear model with measurement errors. Communications in Statistics: Theory and Methods, 40(3):532–546, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2012:CDB

[HS12a] Wen-Jang Huang and Nan-Cheng Su. Characterizations of distributions based on moments of residual life. Communications in Statistics: Theory and Methods, 41(15):2750–2761, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hyodo:2012:MSC

[HS12b] Masashi Hyodo and Takashi Seo. A model selection criterion for discriminant analysis of several groups when the dimension is larger than the total sample size. Communications in Statistics: Theory and Methods, 41(13–14):2419–2436, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:2013:IFR

[HS13a] Abdul Haq and Javid Shabbir. Improved family of ratio estimators in simple and stratified random sampling. Communications in Statistics: Theory and Methods, 42(5):782–799, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2013:HDA

[HS13b] Liwen Huang and Lianta Su. Hierarchical discriminant analysis and its application. Communications in Statistics: Theory and Methods, 42(11):1951–1957, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:2014:BBK

[HS14] Eunju Hwang and Dong Wan Shin. Block bootstrapping for kernel density estimators under ψ-weak dependence. Communications in Statistics: Theory and Methods, 43(17):3751–3761, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:2015:ABS

[HS15a] David D. Hanagal and Richa Sharma. Analysis of bivariate survival data using shared inverse Gaussian frailty model. Communications in Statistics: Theory and Methods, 44(7):1351–1380, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:2015:BIM

[HS15b] David D. Hanagal and Richa Sharma. Bayesian inference in Marshall–Olkin bivariate exponential shared gamma frailty regression model under random censoring. Communications in Statistics: Theory and Methods, 44(1):24–47, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:2015:CFM

[HS15c] David D. Hanagal and Richa Sharma. Comparison of frailty models for acute leukemia data under Gompertz baseline distribution. Communications in Statistics: Theory and Methods, 44(7):1338–1350, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Horrigue:2015:NPR

[HS15d] Walid Horrigue and Elias Ould Saïd. Non parametric regression quantile estimation for dependent functional data under random censorship: Asymptotic normality. Communications in Statistics: Theory and Methods, 44(20):4307–4332, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2017:SGN

[HS17] Wen-Jang Huang and Nan-Cheng Su. A study of generalized normal distributions. Communications in Statistics: Theory and Methods, 46(11):5612–5632, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1107585.

Hussain:2018:SNR

[HS18a] Zawar Hussain and Naila Shabbir. On some new randomized item sum techniques. Communications in Statistics: Theory and Methods, 47(20):4977–4990, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1383428.

Hwang:2018:TSB

[HS18b] Eunju Hwang and Dong Wan Shin. Tests for structural breaks in memory parameters of long-memory heterogeneous autoregressive models. Communications in Statistics: Theory and Methods, 47(21):5378–5389, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1408827.

Hassanein:2019:CSC

[HS19] W. A. Hassanein and M. M. Seyam. Construction of some compound criteria via A-optimality. Communications in Statistics: Theory and Methods, 48(22):5559–5570, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1515364.

Hou:2014:PEL

[HSHW14] Wen Hou, Lixin Song, Xiangyan Hou, and Xiaoguang Wang. Penalized empirical likelihood via bridge estimator in Cox’s proportional hazard model. Communications in Statistics: Theory and Methods, 43(2):426–440, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hua:2017:PLD

[HSLQ17] Zhiqiang Hua, Lixin Song, Dawei Lu, and Xiaomeng Qi. Precise large deviations for the difference of two sums of END random variables with heavy tails. Communications in Statistics: Theory and Methods, 46(2):736–746, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1004094.

Hanook:2013:LE

[HSMK13a] Sharoon Hanook, Muhammad Qaiser Shahbaz, Muhammad Mohsin, and B. M. Golam Kibria. Letter to the Editor. Communications in Statistics: Theory and Methods, 42(18):3371– 3372, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanook:2013:NBI

[HSMK13b] Sharoon Hanook, Muhammad Qaiser Shahbaz, Muhammad Mohsin, and B. M. Golam Kibria. A note on beta inverse-Weibull distribution. Communications in Statistics: Theory and Methods, 42(2):320–335, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hyodo:2016:ECM

[HSSP16] Masashi Hyodo, Nobumichi Shutoh, Takashi Seo, and Tatjana Pavlenko. Estimation of the covariance matrix with two-step monotone missing data. Communications in Statistics: Theory and Methods, 45(7):1910–1922, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.868085.

Huang:2019:SSS

[HST19] Wen-Jang Huang, Nan-Cheng Su, and Hui-Yi Teng. On some study of skew-t distributions. Communications in Statistics: Theory and Methods, 48(19):4712–4729, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2012.700369.

Huang:2017:NAM

[HSW17] Wenxue Huang, Yong Shi, and Xiaogang Wang. A nominal association matrix with feature selection for categorical data. Communications in Statistics: Theory and Methods, 46(16): 7798–7819, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.930911.

Hara:2010:LAI

[HT10] Hisayuki Hara and Akimichi Takemura. A localization approach to improve iterative proportional scaling in Gaussian graphical models. Communications in Statistics: Theory and Methods, 39(8–9):1643–1654, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2013:SSD

[HT13] Wen-Jang Huang and Hui-Yi Teng. A study of some different concepts of symmetry on the real line. Communications in Statistics: Theory and Methods, 42(6):1042–1055, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayakawa:2016:EEP

[HT16a] Jumpei Hayakawa and Akimichi Takemura. Estimation of exponential–polynomial distribution by holonomic gradient descent. Communications in Statistics: Theory and Methods, 45(23):6860–6882, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
968735.

Hosseini:2016:OLO

[HT16b] Reza Hosseini and Akimichi Takemura. An objective look at obtaining the plotting positions for QQ-plots. Communications in Statistics: Theory and Methods, 45(16):4716–4728, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.927498.

Han:2015:DDD

[HTN15] Dong Han, Fugee Tsung, and Xianghui Ning. Detection and diagnosis of distribution changes of degree ratio in complex networks. Communications in Statistics: Theory and Methods, 44(9):1911–1938, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:2017:VCM

[HTW17] Dezhao Han, Ken Seng Tan, and Chengguo Weng. Vine copula models with GLM and sparsity. Communications in Statistics: Theory and Methods, 46(13):6358–6381, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1122061.

Hu:2011:BEL

[Hu11a] Xuemei Hu. The block empirical likelihood method of the semivarying coefficient model with application to longitudinal data. Communications in Statistics: Theory and Methods, 40(8):1342–1351, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2011:LFC

[Hu11b] Yu-Pin Hu. Likelihood function and canonical correlation analysis of the Peña–Box model. Communications in Statistics: Theory and Methods, 40(8):1453–1467, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2013:NEV

[Hu13] Yuao Hu. Nonparametric estimation of variance function for functional data under mixing conditions. Communications in Statistics: Theory and Methods, 42(10):1774–1786, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2014:EES

[Hu14a] Xuemei Hu. Efficient estimation for semi-varying coefficient model with an invertible linear process error. Communications in Statistics: Theory and Methods, 43(15):3117–3134, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2014:SPE

[Hu14b] Xuemei Hu. Semi-parametric efficient inference for heteroscedastic semivarying-coefficient models. Communications in Statistics: Theory and Methods, 43(18):3927–3942, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2016:AHT

[Hu16] Xiaomi Hu. On an ad hoc test for order restricted multivariate normal means. Communications in Statistics: Theory and Methods, 45(5):1501–1507, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
864771.

Hu:2017:MCT

[Hu17] Feng Hu. The modulus of continuity theorem for G-Brownian motion. Communications in Statistics: Theory and Methods, 46(7):3586–3598, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1066816.

Huang:2010:TLD

[Hua10] Pi-Hsiang Huang. Two-level designs in blocks of size two in which all main effects and two-factor interactions are estimable. Communications in Statistics: Theory and Methods, 39(14): 2485–2491, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2011:ELP

[Hua11] Zhensheng Huang. Empirical likelihood for a partially linear single-index measurement error model with right-censored data. Communications in Statistics: Theory and Methods, 40(6):1015–1029, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2012:INP

[Hua12] Zhensheng Huang. Inference for nonparametric parts in singleindex varying-coefficient model. Communications in Statistics: Theory and Methods, 41(7):1214–1227, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2014:SSG

[Hua14a] Chi-Jui Huang. A sum of squares generally weighted moving average control chart. Communications in Statistics: Theory and Methods, 43(23):5052–5071, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2014:MEM

[Hua14b] Shih-Feng Huang. A modified empirical martingale simulation for financial derivative pricing. Communications in Statistics: Theory and Methods, 43(2):328–342, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2015:NGC

[Hua15] Chi-Jui Huang. A new GWMA control chart for monitoring process mean and variability. Communications in Statistics: Theory and Methods, 44(18):3841–3856, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2017:ANP

[Hua17a] Li-Fei Huang. Approximated non parametric confidence regions for the ratio of two percentiles. Communications in Statistics: Theory and Methods, 46(8):4004–4015, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1076479.

Huang:2017:SNP

[Hua17b] Xianzheng Huang. Semi-non parametric smooth isotonic regression. Communications in Statistics: Theory and Methods, 46(20):10071–10087, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1228963.

Hudecova:2014:SPA

[Hud14] Sárka Hudecová. On some properties of autopersistence functions and autopersistence graphs. Communications in Statistics: Theory and Methods, 43(17):3661–3673, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huggins:2009:NES

[Hug09] Richard Huggins. A note on estimating the size of an open population of stationary size. Communications in Statistics: Theory and Methods, 39(2):363–371, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huh:2016:ECP

[Huh16] Jib Huh. Estimation of a change point in the variance function based on the χ^{2}-distribution. Communications in Statistics: Theory and Methods, 45(17):4937–4968, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.930912.

Hui:2010:BEB

[Hui10] Jiang Hui. Berry–Esséen bound for a class of normalized Lstatistics. Communications in Statistics: Theory and Methods, 39(17):3095–3107, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:2019:EEC

[HUK19] Abdul Haq, Zain Ul Abidin, and Michael B. C. Khoo. An enhanced EWMA-t control chart for monitoring the process mean. Communications in Statistics: Theory and Methods, 48(6):1333–1350, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1429631.

Hunter:2014:RKC

[Hun14] Jeffrey J. Hunter. The role of Kemeny’s constant in properties of Markov chains. Communications in Statistics: Theory and Methods, 43(7):1309–1321, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hutson:2019:ASE

[Hut19] Alan D. Hutson. An alternative skew exponential power distribution formulation. Communications in Statistics: Theory and Methods, 48(12):3005–3024, 2019. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473600.

Hamedani:2011:DCS

[HV11] G. G. Hamedani and Hans Volkmer. On the distributions of certain spacings. Communications in Statistics: Theory and Methods, 40(3):510–514, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamedani:2017:WCB

[HV17] G. G. Hamedani and Hans Volkmer. Weak convergence of bivariate sequence to bivariate normal. Communications in Statistics: Theory and Methods, 46(3):1337–1341, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019141.

Hsu:2011:GGB

[HW11] Yu-Sheng Hsu and Cheng-Hsun Wu. A generalization of geometric Brownian motion with applications. Communications in Statistics: Theory and Methods, 40(12):2081–2103, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2014:SCO

[HW14] Shaoyong Hu and Rongming Wang. Stochastic comparisons and optimal allocation for policy limits and deductibles. Communications in Statistics: Theory and Methods, 43(1):151–164, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hashorva:2015:LLM

[HW15] Enkelejd Hashorva and Zhichao Weng. Limit laws for maxima of contracted stationary Gaussian sequences. Communications in Statistics: Theory and Methods, 44(21):4641–4650, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:2018:WBE

[HW18a] Yunzhu He and Dongsheng Wu. Wavelet-based estimators of multivariable mean regression function with long-memory data. Communications in Statistics: Theory and Methods, 47(10):

2389–2406, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1112913.

Hu:2018:TCI

[HW18b] Duni Hu and Hailong Wang. Time-consistent investment and reinsurance under relative performance concerns. Communications in Statistics: Theory and Methods, 47(7):1693–1717, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1324987.

Hoque:2018:SER

[HWH18] Zahirul Hoque, Jacek Wesolowski, and Shahadut Hossain. Shrinkage estimator of regression model under asymmetric loss. Communications in Statistics: Theory and Methods, 47(22): 5547–5557, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1397169.

Huang:2018:HOE

[HWLP18] Jianwen Huang, Jianjun Wang, Guowang Luo, and Hao Pu. Higher order expansion for moments of extreme for generalized Maxwell distribution. Communications in Statistics: Theory and Methods, 47(14):3441–3452, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1359293.

Hutson:2012:ITB

[HWM12] Alan D. Hutson, Dongliang Wang, and Jeffrey C. Miecznikowski. An inversion theorem-based kernel density estimator for a weighted average and difference of weighted averages with applications. Communications in Statistics: Theory and Methods, 41(7):1299–1312, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2015:FSM

[HWW15] Haiwu Huang, Dingcheng Wang, and Qunying Wu. Further study on the Marcinkiewicz strong laws for linear statistics of ρ^{∗}-mixing sequences of random variables. Communications in Statistics: Theory and Methods, 44(1):125–134, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2017:RAU

[HWY17] Hua Huang, Jibo Wu, and Wende Yi. On the restricted almost unbiased two-parameter estimator in linear regression model. Communications in Statistics: Theory and Methods, 46(4): 1668–1678, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1026991.

Hu:2014:NIM

[HWYW14] Shuhe Hu, Xinghui Wang, Wenzhi Yang, and Xuejun Wang. A note on the inverse moment for the non negative random variables. Communications in Statistics: Theory and Methods, 43(8):1750–1757, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:2019:PAM

[HWZ19] Gang He, Wenqing Wu, and Yuanyuan Zhang. Performance analysis of machine repair system with single working vacation. Communications in Statistics: Theory and Methods, 48(22): 5602–5620, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1515958.

He:2011:ALP

[HX11] Daojiang He and Xingzhong Xu. Admissibility of linear predictors in the superpopulation model with respect to inequality constraints under matrix loss function. Communications in Statistics: Theory and Methods, 40(21):3789–3799, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harrar:2014:AMP

[HX14] Solomon W. Harrar and Jin Xu. Analyzing mean profiles of nonnormal populations. Communications in Statistics: Theory and Methods, 43(17):3553–3573, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:2018:ELI

[HXF18] Yuping Hu, Liugen Xue, and Sanying Feng. Empirical likelihood inference for partial functional linear model with missing responses. Communications in Statistics: Theory and Methods, 47(19):4673–4691, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1445856.

He:2013:UPP

[HXL13] Daojiang He, Xingzhong Xu, and Xuhua Liu. The use of posterior predictive p-values in testing goodness-of-fit. Communications in Statistics: Theory and Methods, 42(23):4287–4297, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xie:2017:DBR

[hXZ17] Jie hua Xie and Wei Zou. Dividend barrier and ruin problems for a risk model with delayed claims. Communications in Statistics: Theory and Methods, 46(14):7063–7084, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1143010.

Yu:2014:ELF

[hYhW14] Shi hang Yu and De hui Wang. Empirical likelihood for firstorder autoregressive error-in-variable of models with validation data. Communications in Statistics: Theory and Methods, 43(8):1800–1823, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hojo:2015:EVN

[HYK15a] Shintaro Hojo, Michio Yamamoto, and Yutaka Kano. Effect of violation of the normal assumption on MI and ML estimators in the analysis of incomplete data. Communications in Statistics: Theory and Methods, 44(15):3234–3250, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hyodo:2015:ECB

[HYK15b] Y. Hyodo, H. Yumiba, and M. Kuwada. Existence conditions for balanced fractional 2^{m} factorial designs of resolution 2l + 1 derived from simple arrays. Communications in Statistics: Theory and Methods, 44(12):2564–2570, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:2017:AEP

[HYS17] Huilin Huang, Weiguo Yang, and Zhiyan Shi. Asymptotic equipartition property for second-order circular Markov chains indexed by a two-rooted Cayley tree. Communications in Statistics: Theory and Methods, 46(3):1275–1289, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1014113.

Huang:2018:SMT

[HYS18] Huilin Huang, Weiguo Yang, and Zhiyan Shi. The Shannon– McMillan theorem for Markov chains in Markovian environments indexed by homogeneous trees. Communications in Statistics: Theory and Methods, 47(21):5286–5297, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1388405.

Hong:2017:NPT

[HZ17a] Wenming Hong and Ke Zhou. A note on the passage time of finite-state Markov chains. Communications in Statistics: Theory and Methods, 46(1):438–445, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.995825.

Hu:2017:RMP

[HZ17b] Shaoyong Hu and Ailin Zhu. Risk-minimizing pricing and hedging foreign currency options under regime-switching jumpdiffusion models. Communications in Statistics: Theory and Methods, 46(4):1821–1842, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1030420.

Hu:2018:MEM

[HZ18a] Jianhua Hu and NengHui Zhu. M-estimation and model identification based on double SCAD penalization. Communications in Statistics: Theory and Methods, 47(23):5639–5661, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1400055.

Huang:2018:OWE

[HZ18b] Jiting Huang and Peixin Zhao. Orthogonal weighted empirical likelihood-based variable selection for semiparametric instrumental variable models. Communications in Statistics: Theory and Methods, 47(18):4375–4388, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1373821.

Huang:2019:SEP

[HZ19] Zhensheng Huang and Xin Zhao. Statistical estimation for a partially linear single-index model with errors in all variables. Communications in Statistics: Theory and Methods, 48(5): 1136–1148, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1425446.

Huang:2016:SIP

[HZL16] Zhensheng Huang, Riquan Zhang, and Yazhao Lv. Semiparametric inference on partially linear single-index model. Communications in Statistics: Theory and Methods, 45(14):4068–4085, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.915045.

Haghbin:2017:IAB

[HZS17] Hossein Haghbin, Atefeh Zamani, and Zohreh Shishebor. Inference on the asymptotic behavior of covariance operator of first-order periodically correlated autoregressive Hilbertian processes. Communications in Statistics: Theory and Methods, 46 (2):761–769, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1005099.

Hu:2017:FTN

[HZW17] Feng Hu, Zhaojun Zong, and Helin Wu. Fubini theorem for non additive measures in the framework of g-expectation. Communications in Statistics: Theory and Methods, 46(23):11776– 11785, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1285921.

Huang:2019:ECC

[HZZ19] Haiwu Huang, Hang Zou, and Qingxia Zhang. Equivalent conditions of the complete convergence for weighted sums of NSD random variables. Communications in Statistics: Theory and Methods, 48(18):4675–4689, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1500601.

Izadkhah:2015:FRG

[IAA15] S. Izadkhah, H. Ahmadzade, and M. Amini. Further results for a general family of bivariate copulas. Communications in Statistics: Theory and Methods, 44(15):3146–3157, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Izadkhah:2016:PDC

[IAB16] S. Izadkhah, M. Amini, and G. R. Mohtashami Borzadaran. Preservation of dependence concepts under bivariate weighted distributions. Communications in Statistics: Theory and Methods, 45(15):4589–4599, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
923459.

Iriarte:2017:GMD

[IABG17] Yuri A. Iriarte, Juan M. Astorga, Heleno Bolfarine, and Héctor W. Gómez. Gamma–Maxwell distribution. Communications in Statistics: Theory and Methods, 46(9):4264–4274, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1081946.

Iannario:2012:HCM

[Ian12] Maria Iannario. Hierarchical CUB models for ordinal variables. Communications in Statistics: Theory and Methods, 41(16–17): 3110–3125, 2012. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/03610926.2011.645987.

Iannario:2014:MUO

[Ian14] Maria Iannario. Modelling uncertainty and overdispersion in ordinal data. Communications in Statistics: Theory and Methods, 43(4):771–786, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iranmanesh:2013:IMV

[IANT13] Anis Iranmanesh, M. Arashi, D. K. Nagar, and S. M. M. Tabatabaey. On inverted matrix variate gamma distribution. Communications in Statistics: Theory and Methods, 42(1):28– 41, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iqbal:2018:CND

[IBA18] I. Iqbal, P. F. Buzdar, and I. Ahmed. Construction of neighbor designs of block size 4 for all values of v >= 5 by using the method of addition. Communications in Statistics: Theory and Methods, 47(10):2482–2489, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1301472.

Iriarte:2018:MSR

[ICBG18] Yuri A. Iriarte, Nabor O. Castillo, Heleno Bolfarine, and Héctor W. Gómez. Modified slashed-Rayleigh distribution. Communications in Statistics: Theory and Methods, 47(13): 3220–3233, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1353621.

Inan:2017:PSO

[IES^{+}17] Deniz Inan, Erol Egrioglu, Busenur Sarica, Oykum Esra Askin, and Mujgan Tez. Particle swarm optimization based Liu-type estimator. Communications in Statistics: Theory and Methods, 46(22):11358–11369, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1267759.

Iezzi:2012:CMT

[Iez12] Domenica Fioredistella Iezzi. Centrality measures for text clustering. Communications in Statistics: Theory and Methods, 41(16–17):3179–3197, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
633729.

Isogai:2018:BRE

[IF18] Eiichi Isogai and Andreas Futschik. Bounded risk estimation of a linear combination of location parameters in negative exponential distributions via three-stage sampling. Communications in Statistics: Theory and Methods, 47(15):3728–3733, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1361990.

Igarashi:2016:WLN

[Iga16] Gaku Igarashi. Weighted log-normal kernel density estimation. Communications in Statistics: Theory and Methods, 45(22): 6670–6687, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.963623.

Iglesias:2013:BBB

[Igl13] Emma M. Iglesias. The block-block bootstrap for time series. Communications in Statistics: Theory and Methods, 42(14): 2584–2600, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Igarashi:2018:LBR

[IK18] Gaku Igarashi and Yoshihide Kakizawa. Limiting bias-reduced amoroso kernel density estimators for non-negative data. Communications in Statistics: Theory and Methods, 47(20):4905– 4937, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1380832.

Ikeda:2017:NML

[Ike17] Shin S. Ikeda. A note on mixingale limit theorems and stable convergence in law. Communications in Statistics: Theory and Methods, 46(19):9377–9387, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1208238.

Islam:2019:MPL

[IKH19] Md. Mazharul Islam, Md. Hasinur Rahaman Khan, and Tamanna Hawlader. Modified profile likelihood estimation for the Weibull regression models in survival analysis. Communications in Statistics: Theory and Methods, 48(9):2329–2343, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1472784.

Izadi:2016:TMI

[IKL16] Muhyiddin Izadi, Baha-Eldin Khaledi, and Chin-Diew Lai. On testing more IFRA ordering — II. Communications in Statistics: Theory and Methods, 45(15):4556–4568, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.922987.

Ilic:2016:GTS

[Ili16] Ana V. Miletić Ilić. A geometric time series model with a new dependent Bernoulli counting series. Communications in Statistics: Theory and Methods, 45(21):6400–6415, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.895840.

Imada:2017:SCB

[Ima17] T. Imada. Successive comparisons between ordered normal means using isotonic regression estimators. Communications in Statistics: Theory and Methods, 46(7):3186–3199, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1056370.

Imada:2018:MCC

[Ima18] T. Imada. Multiple comparison for checking differences among normal variances. Communications in Statistics: Theory and Methods, 47(18):4402–4414, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1376083.

Imoto:2015:CBN

[Imo15] Tomoaki Imoto. Convolution of binomial and negative binomial variables. Communications in Statistics: Theory and Methods, 44(23):5005–5022, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Imoto:2016:PLD

[Imo16] Tomoaki Imoto. Properties of Lagrangian distributions. Communications in Statistics: Theory and Methods, 45(3):712–721, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Indlekofer:2011:LUC

[Ind11] Karl-Heinz Indlekofer. On labeled and unlabeled combinatorial structures. Communications in Statistics: Theory and Methods, 40(19–20):3641–3653, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Imoto:2017:MCM

[INOC17] T. Imoto, C. M. Ng, S. H. Ong, and S. Chakraborty. A modified Conway–Maxwell–Poisson type binomial distribution and its applications. Communications in Statistics: Theory and Methods, 46(24):12210–12225, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1291974.

Imoto:2013:LNC

[IO13] Tomoaki Imoto and Seng Huat Ong. A Lagrangian non central negative binomial distribution of the first kind. Communications in Statistics: Theory and Methods, 42(3):466–477, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Islam:2012:BSS

[IP12] A. F. M. Saiful Islam and L. I. Pettit. Bayesian sample size determination using LINEX loss and linear cost. Communications in Statistics: Theory and Methods, 41(2):223–240, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ieva:2013:DMM

[IP13] Francesca Ieva and Anna M. Paganoni. Depth measures for multivariate functional data. Communications in Statistics: Theory and Methods, 42(7):1265–1276, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iqbal:2013:DCG

[Iqb13] Farhat Iqbal. Diagnostic checking for GARCH-type models. Communications in Statistics: Theory and Methods, 42(6):934– 953, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iriarte:2019:SPL

[IR19] Yuri A. Iriarte and Mario A. Rojas. Slashed power-Lindley distribution. Communications in Statistics: Theory and Methods, 48(7):1709–1720, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1438626.

Izadkhah:2013:GAP

[IRAB13] S. Izadkhah, A. H. Rezaei, M. Amini, and G. R. Mohtashami Borzadaran. A general approach for preservation of some aging classes under weighting. Communications in Statistics: Theory and Methods, 42(10):1899–1909, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Izadkhah:2013:PRM

[IRB13] S. Izadkhah, A. H. Rezaei Roknabadi, and G. R. Mohtashami Borzadaran. On properties of reversed mean residual life order for weighted distributions. Communications in Statistics: Theory and Methods, 42(5):838–851, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ivanov:2011:APK

[IS11] Alexander V. Ivanov and Irina N. Savich. Asymptotic properties of Koenker–Bassett estimator in regression model with long-range dependence. Communications in Statistics: Theory and Methods, 40(19–20):3555–3568, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwinska:2016:CED

[IS16] Maria Iwińska and Magdalena Szymkowiak. Characterizations of the exponential distribution by Pascal compound. Communications in Statistics: Theory and Methods, 45(1):63–70, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwinska:2017:CDT

[IS17] Maria Iwińska and Magdalena Szymkowiak. Characterizations of distributions through selected functions of reliability theory. Communications in Statistics: Theory and Methods, 46(1):69– 74, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.985837.

Islam:2017:EHT

[Isl17] M. Qamarul Islam. Estimation and hypothesis testing in multivariate linear regression models under non normality. Communications in Statistics: Theory and Methods, 46(17):8521–8543, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1183789.

Inan:2012:NEC

[IT12] Deniz Inan and Müjgan Tez. A new estimator for Cox proportional hazard regression model in presence of collinearity. Communications in Statistics: Theory and Methods, 41(13–14): 2437–2444, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iqbal:2009:CFS

[ITG09] Ijaz Iqbal, M. H. Tahir, and Syed Skahir Ali Ghazali. Circular first- and second-order balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 39 (2):228–240, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isogai:2012:TSP

[IUT12] Eiichi Isogai, Chikara Uno, and Daisuke Takeuchi. Two-stage procedure having exact third-order properties for a normal mean. Communications in Statistics: Theory and Methods, 41(23):4332–4347, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iriarte:2017:SGR

[IVVG17] Yuri A. Iriarte, F. Vilca, Héctor Varela, and Héctor W. Gómez. Slashed generalized Rayleigh distribution. Communications in Statistics: Theory and Methods, 46(10):4686–4699, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1066811.

Jazi:2010:OCL

[JA10] Mansour Aghababaei Jazi and Mohammad Hossein Alamatsaz. Ordering comparison of logarithmic series random variables with their mixtures. Communications in Statistics: Theory and Methods, 39(18):3252–3263, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jazi:2011:UAO

[JAA11] Mansour Aghababaei Jazi, Mohammad Hossein Alamatsaz, and Somayyeh Abbasi. A unified approach to ordering comparison of GPS distributions with their mixtures. Communications in Statistics: Theory and Methods, 40(14):2591–2604, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jacroux:2011:RFB

[Jac11] Mike Jacroux. Reverse foldovers for blocked 2^{m−k} fractional factorial designs. Communications in Statistics: Theory and Methods, 40(15):2799–2808, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jafari:2013:LOD

[Jaf13] Habib Jafari. Locally D-optimal design for a logit model in discrete choice experiment. Communications in Statistics: Theory and Methods, 42(5):867–882, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jafari:2015:ICV

[Jaf15] Ali Akbar Jafari. Inferences on the coefficients of variation in a multivariate normal population. Communications in Statistics: Theory and Methods, 44(12):2630–2643, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jasinski:2016:MVR

[Jas16] Krzysztof Jasiński. On maximum variance of k-th records. Communications in Statistics: Theory and Methods, 45(8): 2392–2401, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2013.879895.

Jasinski:2019:WRG

[Jas19] Krzysztof Jasiński. Weak k-th records from geometric distribution and some characterizations. Communications in Statistics: Theory and Methods, 48(24):6179–6187, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1529245.

Jaworski:2014:CCD

[Jaw14] Piotr Jaworski. On the characterization of copulas by differential equations. Communications in Statistics: Theory and Methods, 43(16):3402–3428, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:2015:GES

[JB15] Alicia A. Johnson and Owen Burbank. Geometric ergodicity and scanning strategies for two-component Gibbs samplers. Communications in Statistics: Theory and Methods, 44(15): 3125–3145, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jayakumar:2018:DWG

[JB18] K. Jayakumar and M. Girish Babu. Discrete Weibull geometric distribution and its properties. Communications in Statistics: Theory and Methods, 47(7):1767–1783, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1327074.

Jeyadurga:2018:DAN

[JBA18] P. Jeyadurga, S. Balamurali, and M. Aslam. Design of an attribute np control chart for process monitoring based on repetitive group sampling under truncated life tests. Communications in Statistics: Theory and Methods, 47(24):5934–5955, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1404098.

Jacobs:2012:BKB

[JBH12] R. Jacobs, A. Bekker, and S. W. Human. On the bivariate Kummer–beta Type IV distribution. Communications in Statistics: Theory and Methods, 41(18):3339–3354, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jia:2012:RRS

[JC12] Xujie Jia and Lirong Cui. Reliability research of k-out-of-n: G supply chain system based on copula. Communications in Statistics: Theory and Methods, 41(21):4023–4033, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jung:2018:BIT

[JC18] Myoungjin Jung and Younshik Chung. Bayesian inference of three-parameter bathtub-shaped lifetime distribution. Communications in Statistics: Theory and Methods, 47(17):4229–4241, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1371757.

Jiang:2017:NCD

[JCN17] Xiao Jiang, Jeffrey Chu, and Saralees Nadarajah. New classes of discrete bivariate distributions with application to football data. Communications in Statistics: Theory and Methods, 46(16):8069–8085, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1171358.

Jia:2010:SRC

[JCY10] Xujie Jia, Lirong Cui, and Jinqiu Yan. A study on the reliability of consecutive k-out-of-n: G systems based on copula. Communications in Statistics: Theory and Methods, 39(13): 2455–2472, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:2016:IPC

[JD16] Edmund Jones and Vanessa Didelez. Inequalities on partial correlations in Gaussian graphical models containing star shapes. Communications in Statistics: Theory and Methods, 45(20): 5990–5996, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.953696.

Jin:2016:DOM

[JDSS16] Libin Jin, Xiaowen Dai, Anqi Shi, and Lei Shi. Detection of outliers in mixed regressive-spatial autoregressive models. Communications in Statistics: Theory and Methods, 45(17):5179–5192, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.941493.

Jeon:2012:CLD

[Jeo12] Yongho Jeon. A characterization of the log-density smoothing spline ANOVA model. Communications in Statistics: Theory and Methods, 41(12):2081–2087, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jomhoori:2012:CLT

[JFA12] Sarah Jomhoori, Vahid Fakoor, and Hasanali Azarnoosh. Central limit theorem for ISE of kernel density estimators in censored dependent model. Communications in Statistics: Theory and Methods, 41(8):1334–1349, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeske:2014:MSS

[JFS14] Daniel R. Jeske, James M. Flegal, and Stephen R. Spindler. Minimum size survival analysis sampling plans for comparing multiple treatment groups to a single control group. Communications in Statistics: Theory and Methods, 43(13):2689–2701, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:2019:BSE

[JHM19] Jun Jin, Chenyan Hao, and Tiefeng Ma. B-spline estimation for partially linear varying coefficient composite quantile regression models. Communications in Statistics: Theory and Methods, 48(21):5322–5335, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1510006.

Jung:2010:BAM

[JHP10] Ki Mun Jung, Sung Sil Han, and Dong Ho Park. A Bayesian approach to maintenance policy based on cost and downtime after non-renewing warranty. Communications in Statistics:

Theory and Methods, 39(13):2321–2332, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:2012:TNL

[Jia12] Renyan Jiang. Three new life distribution models for modeling field failure data. Communications in Statistics: Theory and Methods, 41(21):3974–3987, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jacroux:2016:OBM

[JJ16] Mike Jacroux and Tom Jacroux. On the E-optimality of blocked main effects plans when n ≡ 1 (mod 4). Communications in Statistics: Theory and Methods, 45(19):5584–5589, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.948197.

Joshi:2018:WLD

[JJ18] Savitri Joshi and K. K. Jose. Wrapped Lindley distribution. Communications in Statistics: Theory and Methods, 47(5): 1013–1021, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1280168.

Jaffa:2019:LBA

[JJ19] Miran A. Jaffa and Ayad A. Jaffa. A likelihood based approach for joint modeling of longitudinal trajectories and informative censoring process. Communications in Statistics: Theory and Methods, 48(12):2982–3004, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473599.

Kim:2015:BLC

[jK15] Hea jung Kim. Best linear classification rule for multivariate interval screened data. Communications in Statistics: Theory and Methods, 44(8):1669–1690, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:2019:PPM

[JK19] Peiyun Jiang and Eiji Kurozumi. Power properties of the modified CUSUM tests. Communications in Statistics: Theory and Methods, 48(12):2962–2981, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473598.

Jacroux:2017:OBM

[JKD17] Mike Jacroux and Bonni Kealy-Dichone. On the E-optimality of blocked main effects plans in blocks of different sizes. Communications in Statistics: Theory and Methods, 46(5):2132– 2138, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1033427.

Jhajj:2016:EFE

[JKJ16] H. S. Jhajj, Harpreet Kaur, and Puneet Jhajj. Efficient family of estimators of median using two-phase sampling design. Communications in Statistics: Theory and Methods, 45(15): 4325–4331, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.911912.

Jo:2019:APE

[JKL19] Seongil Jo, Jeongseon Kim, and Woojoo Lee. An analysis pipeline for estimating true intake from repeated measurements with random errors. Communications in Statistics: Theory and Methods, 48(5):1239–1254, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1429624.

Jowaheer:2018:BTS

[JKS18] V. Jowaheer, N. Mamode Khan, and Y. Sunecher. A BI-NAR(1) time-series model with cross-correlated COM–Poisson innovations. Communications in Statistics: Theory and Methods, 47(5):1133–1154, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1316400.

Jing:2013:EIC

[JLL13] Bing-Yi Jing, Cui-Xia Li, and Zhi Liu. On estimating the integrated co-volatility using noisy high-frequency data with jumps. Communications in Statistics: Theory and Methods, 42(21): 3889–3901, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joannides:2013:GSS

[JLV13] Marc Joannides and Irène Larramendy-Valverde. On geometry and scale of a stochastic chemostat. Communications in Statistics: Theory and Methods, 42(16):2902–2911, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jose:2016:NIV

[JM16] K. K. Jose and K. D. Mariyamma. A note on an integer valued time series model with Poisson-negative binomial marginal distribution. Communications in Statistics: Theory and Methods, 45(1):123–131, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jaheen:2014:BIC

[JMEM14] Z. F. Jaheen, H. M. Moustafa, and G. H. Abd El-Monem. Bayes inference in constant partially accelerated life tests for the generalized exponential distribution with progressive censoring. Communications in Statistics: Theory and Methods, 43(14):2973–2988, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jafaraghaie:2018:RBP

[JN18] Razieh Jafaraghaie and Nader Nematollahi. Robust Bayesian prediction under a general linear-exponential posterior risk function and its application in finite population. Communications in Statistics: Theory and Methods, 47(13):3269–3292, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1383425.

Jongpreechaharn:2019:NAC

[JN19] Suporn Jongpreechaharn and Kritsana Neammanee. Normal approximation for call function via Stein’s method. Communications in Statistics: Theory and Methods, 48(14):3498–3517, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1476716.

Jureckova:2010:EPI

[JO10] Jana Jurecková and Marek Omelka. Estimator of the Pareto index based on nonparametric test. Communications in Statistics: Theory and Methods, 39(8–9):1536–1551, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:2016:BTS

[Jon16] M. C. Jones. On bivariate transformation of scale distributions. Communications in Statistics: Theory and Methods, 45(3):577– 588, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:2018:LEC

[Jon18] M. C. Jones. Letter to the Editor concerning “A new method for generating distributions with an application to exponential distribution” and “Alpha power Weibull distribution: Properties and applications”. Communications in Statistics: Theory and Methods, 47(20):5096, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1386314. See [MK17, NAMAK17].

Jorgensen:2013:FCC

[Jor13] Murray A. Jorgensen. Forming clusters from census areas with similar tabular statistics. Communications in Statistics: Theory and Methods, 42(12):2136–2151, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jentsch:2013:VRH

[JP13] Carsten Jentsch and Dimitris N. Politis. Valid resampling of higher-order statistics using the linear process bootstrap and autoregressive sieve bootstrap. Communications in Statistics: Theory and Methods, 42(7):1277–1293, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jamalizadeh:2017:OST

[JPAB17] Ahad Jamalizadeh, Reza Pourmousa, Mehdi Amiri, and N. Balakrishnan. Order statistics and their concomitants from multivariate normal mean-variance mixture distributions with application to Swiss Markets Data. Communications in Statistics: Theory and Methods, 46(22):10991–11009, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1257719.

Jiang:2018:WCQ

[JQZ18] Rong Jiang, Wei-Min Qian, and Zhan-Gong Zhou. Weighted composite quantile regression for partially linear varying coefficient models. Communications in Statistics: Theory and Methods, 47(16):3987–4005, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1366522.

Jensen:2018:DEF

[JR18a] D. R. Jensen and D. E. Ramirez. Designs enhancing Fisher information. Communications in Statistics: Theory and Methods, 47(20):4895–4904, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1380831.

Joshi:2018:MPP

[JR18b] Kalyan Joshi and M. B. Rajarshi. Modified probability proportional to size sampling. Communications in Statistics: Theory and Methods, 47(4):805–815, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1139131.

Jahanshahi:2017:GFT

[JRF17] S. M. A. Jahanshahi, A. Habibi Rad, and V. Fakoor. Goodnessof-fit test under length-biased sampling. Communications in Statistics: Theory and Methods, 46(15):7580–7592, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1157187.

Jayakumar:2010:GBM

[JRM10a] K. Jayakumar, Miroslav M. Ristić, and Davis Antony Mundassery. A generalization to bivariate Mittag-Leffler and bivariate discrete Mittag-Leffler autoregressive processes. Communications in Statistics: Theory and Methods, 39(6):942–955, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jokiel-Rokita:2010:MRI

[JRM10b] Alicja Jokiel-Rokita and Ryszard Magiera. Minimum risk invariant estimators of a continuous cumulative distribution function. Communications in Statistics: Theory and Methods, 39 (18):3332–3342, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johannes:2013:AGI

[JS13] Jan Johannes and Maik Schwarz. Adaptive Gaussian inverse regression with partially unknown operator. Communications in Statistics: Theory and Methods, 42(7):1343–1362, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jayakumar:2018:GDW

[JS18] K. Jayakumar and K. K. Sankaran. A generalization of discrete Weibull distribution. Communications in Statistics: Theory and Methods, 47(24):6064–6078, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1406115.

Jung:2017:BCP

[JSC17] Myoungjin Jung, Seongho Song, and Younshik Chung. Bayesian change-point problem using Bayes factor with hierarchical prior distribution. Communications in Statistics: Theory and Methods, 46(3):1352–1366, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1019143.

Javed:2016:EFR

[JSS16] M. Javed, S. S. Sidhu, and Sarjinder Singh. An extension of Franklin’s randomized response technique for multiproportions. Communications in Statistics: Theory and Methods, 45(16):4637–4662, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
927493.

Jayaraj:2018:NEP

[JSSO18] Augustus Jayaraj, Stephen A. Sedory, Sarjinder Singh, and Oluseun Odumade. A new estimator of proportion with a linear function using data from two-decks randomized response model. Communications in Statistics: Theory and Methods, 47(6):1475–1490, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1321764.

Jia:2017:VCC

[JSWL17] Xujie Jia, Jingyuan Shen, Liying Wang, and Zhongping Li. Vine copula constructions of higher-dimensional dependent reliability systems. Communications in Statistics: Theory and Methods, 46(18):9126–9136, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1205620.

Jiang:2017:NPT

[JSZL17] Yuexiang Jiang, Haoze Sun, Yi Zhang, and Huaigang Long. Non-parametric tests for the tail equivalence via empirical likelihood. Communications in Statistics: Theory and Methods, 46(21):10640–10656, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1242736.

Jose:2011:GLD

[JT11] K. K. Jose and Manu Mariam Thomas. Generalized Laplacian distributions and autoregressive processes. Communications in Statistics: Theory and Methods, 40(23):4263–4277, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jafari:2016:GPS

[JT16] A. A. Jafari and S. Tahmasebi. Gompertz-power series distributions. Communications in Statistics: Theory and Methods, 45(13):3761–3781, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
911904.

Jena:2019:EOQ

[JT19] Adarsha Kumar Jena and Manas Ranjan Tripathy. Estimating ordered quantiles of two exponential populations with a common minimum guarantee time. Communications in Statistics: Theory and Methods, 48(14):3570–3585, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1478100.

Juutilainen:2015:VPO

[JTR15] I. Juutilainen, S. Tamminen, and J. Röning. Visualizing predicted and observed densities jointly with beanplot. Communications in Statistics: Theory and Methods, 44(2):340–348, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ju:2015:MBB

[Ju15] Hyunsu Ju. Moving block bootstrap for analyzing longitudinal data. Communications in Statistics: Theory and Methods, 44(6):1130–1142, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jung:2017:SEP

[Jun17] Yoonsuh Jung. Shrinkage estimation of proportion via logit penalty. Communications in Statistics: Theory and Methods, 46(5):2447–2453, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1048881.

Jia:2014:SMV

[JWZ14] Boting Jia, Dehui Wang, and Haixiang Zhang. A study for missing values in PINAR(1)_{T} processes. Communications in Statistics: Theory and Methods, 43(22):4780–4789, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:2019:CAQ

[JX19] Tao Jiang and Baogui Xin. Computational analysis of the queue with working breakdowns and delaying repair under a Bernoulli-schedule-controlled policy. Communications in Statistics: Theory and Methods, 48(4):926–941, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1422756.

Jiang:2017:APM

[JXW17] Xuejun Jiang, Tian Xia, and Xueren Wang. Asymptotic properties of maximum quasi-likelihood estimator in quasilikelihood non linear models with stochastic regression. Communications in Statistics: Theory and Methods, 46(13):6229– 6239, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1129422.

Jomhoori:2014:ERR

[JY14] Sarah Jomhoori and Fatemeh Yousefzadeh. On estimating the residual Rényi entropy under progressive censoring. Communications in Statistics: Theory and Methods, 43(10–12):2395– 2405, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:2016:MDM

[JY16] Hui Jiang and Qingshan Yang. Moderate deviations for the moment estimators in Rayleigh distribution with two parameters. Communications in Statistics: Theory and Methods, 45 (2):330–339, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:2016:APM

[JYX16] Hui Jiang, Xin Yun, and Chen Xie. Asymptotic properties for the moment estimators in Rayleigh distribution with two parameters. Communications in Statistics: Theory and Methods, 45(16):4760–4770, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
927500.

Jiang:2019:ODM

[JYZ19] Hong-Yan Jiang, Rong-Xian Yue, and Xiao-Dong Zhou. Optimal designs for multivariate logistic mixed models with longitudinal data. Communications in Statistics: Theory and Methods, 48(4):850–864, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1419263.

Jiang:2016:EAQ

[JZ16] Ting Jiang and XiaoJian Zhou. Estimation of actuarial quantities at fractional ages with kriging. Communications in Statistics: Theory and Methods, 45(23):6943–6954, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.972571.

Jayasinghe:2017:NPH

[JZ17a] Chathuri L. Jayasinghe and Panlop Zeephongsekul. Non parametric hypothesis tests for comparing reliability functions. Communications in Statistics: Theory and Methods, 46(8): 3698–3717, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1071392.

Jin:2017:EGA

[JZ17b] Xinghu Jin and Zhenzhong Zhang. Ergodicity of generalized Ait–Sahalia-type interest rate model. Communications in Statistics: Theory and Methods, 46(16):8199–8209, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1177078.

Jiang:2015:GAV

[JZQ15] Rong Jiang, Zhan-Gong Zhou, and Wei-Min Qian. Generalized analysis-of-variance-type test for the single-index quantile model. Communications in Statistics: Theory and Methods, 44(13):2842–2861, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:2017:TCL

[JZTN17] Jiao Jin, Liang Zhu, Xingwei Tong, and Kirsten K. Ness. t-Type corrected-loss estimation for error-in-variable model. Communications in Statistics: Theory and Methods, 46(2):616–627, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.1002934.

Jin:2018:BPV

[JZZZ18] Hao Jin, Si Zhang, Jinsuo Zhang, and Shougang Zhang. Bootstrap procedures for variance breaks test in time series with a changing trend. Communications in Statistics: Theory and Methods, 47(18):4609–4627, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1377256.

Kumar:2015:EVS

[KA15] C. Satheesh Kumar and M. R. Anusree. On an extended version of skew generalized normal distribution and some of its properties. Communications in Statistics: Theory and Methods, 44(3):573–586, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaur:2017:TWL

[KA17] Kirandeep Kaur and Umme Afifa. Testing Wagner’s Law in India: a cointegration and causality analysis. Communications in Statistics: Theory and Methods, 46(17):8510–8520, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1183788.

Kurnaz:2018:MMS

[KA18] Fatma Sevinç Kurnaz and Kadri Ulas Akay. Matrix mean squared error comparisons of some biased estimators with two biasing parameters. Communications in Statistics: Theory and Methods, 47(8):2022–2035, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1335415.

Kachour:2014:RIV

[Kac14] Maher Kachour. On the rounded integer-valued autoregressive process. Communications in Statistics: Theory and Methods, 43(2):355–376, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaciranlar:2019:CWM

[Kaç19] Selahattin Kaçiranlar. Comments on “On the weighted mixed Liu-type estimator under unbiased stochastic restrictions”. Communications in Statistics: Theory and Methods, 48(3):435– 437, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1438624.

Kwessi:2014:ERR

[KAD14] E. Kwessi, A. Abebe, and G. De Souza. Efficient rank regression with wavelet estimated scores. Communications in Statistics: Theory and Methods, 43(18):3986–3996, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2018:DAS

[KAJH18] Nasrullah Khan, Muhammad Aslam, Chi-Hyuck Jun, and Jaffer Hussain. Design of acceptance sampling plan using a modified EWMA statistic. Communications in Statistics: Theory and Methods, 47(12):2881–2891, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1343846.

Kakizawa:2010:SOP

[Kak10] Yoshihide Kakizawa. Second-order power comparison of tests. Communications in Statistics: Theory and Methods, 39(8–9): 1424–1436, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kakizawa:2012:SOP

[Kak12] Yoshihide Kakizawa. Second-order powers of a class of tests in the presence of a nuisance parameter. Communications in Statistics: Theory and Methods, 41(20):3676–3691, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalemkerian:2017:IFD

[Kal17] Juan Kalemkerian. An integral formula for the distribution of self-normalized Gaussian random samples. Communications in Statistics: Theory and Methods, 46(10):4671–4685, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1060335.

Kalemkerian:2019:TCM

[Kal19] Juan Kalemkerian. A truncated Cramér–von Mises test of normality. Communications in Statistics: Theory and Methods, 48(16):3956–3975, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1465093.

Kamalja:2014:BRI

[Kam14] K. K. Kamalja. Birnbaum reliability importance for (n, f, k) and ⟨n, k, f⟩ system. Communications in Statistics: Theory and Methods, 43(10–12):2406–2418, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kamiya:2016:UAM

[Kam16] Hidehiko Kamiya. A unified approach to marginal equivalence in the general framework of group invariance. Communications in Statistics: Theory and Methods, 45(23):6955–6968, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.972572.

Kamiya:2019:NES

[Kam19] Hidehiko Kamiya. A note on estimation of the shape of density level sets of star-shaped distributions. Communications in Statistics: Theory and Methods, 48(12):3092–3104, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473606.

Kapelko:2018:MDP

[Kap18] Rafał Kapelko. On the moment distance of Poisson processes. Communications in Statistics: Theory and Methods, 47(24): 6052–6063, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1406114.

Kara:2019:PES

[KAPA19] Mahmut Kara, Ömer Altindag, Mustafa Hilmi Pekalp, and Halil Aydogdu. Parameter estimation in α-series process with lognormal distribution. Communications in Statistics: Theory and Methods, 48(20):4976–4998, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1504075.

Khan:2016:DOA

[KAR16] M. G. M. Khan, N. Ahmad, and L. S. Rafi. Determining the optimal allocation of testing resource for modular software system using dynamic programming. Communications in Statistics: Theory and Methods, 45(3):670–694, 2016. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kareev:2017:LBE

[Kar17] Iskander Kareev. Lower bounds for expected sample size of sequential procedures for the multinomial selection problems. Communications in Statistics: Theory and Methods, 46(19): 9787–9794, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1222429.

Kara:2017:SIS

[KAS17] Mahmut Kara, Halil Aydogdu, and Birdal Senoglu. Statistical inference for α-series process with gamma distribution. Communications in Statistics: Theory and Methods, 46(13): 6727–6736, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1134571.

Katai:2011:PGS

[Kát11a] Imre Kátai. A proof of the generalized second-limit theorem in the theory of probability. Communications in Statistics: Theory and Methods, 40(19–20):3654–3662, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kateri:2011:CTO

[Kat11b] Maria Kateri. On the comparison of two ordinal responses. Communications in Statistics: Theory and Methods, 40(21): 3748–3763, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karbalaee:2018:PAP

[KAT18] M. H. Karbalaee, M. Arashi, and S. M. M. Tabatabaey. Performance analysis of the preliminary test estimator with series of stochastic restrictions. Communications in Statistics: Theory and Methods, 47(1):1–17, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1300275.

Kawai:2013:FIF

[Kaw13] Reiichiro Kawai. Fisher information for fractional Brownian motion under high-frequency discrete sampling. Communications in Statistics: Theory and Methods, 42(9):1628–1636, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kang:2018:ABF

[KAW^{+}18] Suling Kang, Rashid Abbasi, Huiling Wang, Lixiang Xu, Xiaofeng Wang, and Yuanyan Tang. Asymptotic behaviour of frequency polygons under ϕ-mixing samples. Communications in Statistics: Theory and Methods, 47(21):5369–5377, 2018.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1406516.

Kayal:2016:CBG

[Kay16] Suchandan Kayal. Characterization based on generalized entropy of order statistics. Communications in Statistics: Theory and Methods, 45(15):4628–4636, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
927491.

Kayal:2018:QBC

[Kay18] Suchandan Kayal. Quantile-based Chernoff distance for truncated random variables. Communications in Statistics: Theory and Methods, 47(20):4938–4957, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1383426.

Kaziska:2011:FAV

[Kaz11] David M. Kaziska. Functional analysis of variance, discriminant analysis, and clustering in a manifold of elastic curves. Communications in Statistics: Theory and Methods, 40(14):2487–2499, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2013:DMH

[KB13] Jaejik Kim and L. Billard. Dissimilarity measures for histogram-valued observations. Communications in Statistics: Theory and Methods, 42(2):283–303, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kontorovich:2014:SLL

[KB14] A. Kontorovich and A. Brockwell. A strong law of large numbers for strongly mixing processes. Communications in Statistics: Theory and Methods, 43(18):3777–3796, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2017:OIM

[KB17] Raju Kumar and Lal Mohan Bhar. Outliers in incomplete multiresponse experiments. Communications in Statistics: Theory and Methods, 46(21):10426–10445, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1235199.

Kao:2012:DMC

[KC12] Lie-Jane Kao and Li-Shya Chen. The dynamic model for cancer relapse based on two-stage model of carcinogenesis. Communications in Statistics: Theory and Methods, 41(10):1701–1720, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kohonen:2016:CEC

[KC16] Jukka Kohonen and Jukka Corander. Computing exact clustering posteriors with subset convolution. Communications in Statistics: Theory and Methods, 45(10):3048–3058, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.894070.

Kalita:2019:AUM

[KC19a] Chandi Ram Kalita and Gautam Choudhury. Analysis of an unreliable M^{X} /\frac{G_{1}}{G_{2}} /1/ repeated service queue with delayed repair under randomized vacation policy. Communications in Statistics: Theory and Methods, 48(21):5336–5369, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1513142.

Kalita:2019:PUQ

[KC19b] Chandi Ram Kalita and Gautam Choudhury. The N-policy for an unreliable queue (M^{x} /\frac{G_{1}}{G_{2}} /1) queue with optional repeated service and delayed repair. Communications in Statistics: Theory and Methods, 48(15):3717–3745, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1481967.

Khurana:2014:JRR

[KCC14] Mansi Khurana, Yogendra P. Chaubey, and Shalini Chandra. Jackknifing the ridge regression estimator: a revisit. Communications in Statistics: Theory and Methods, 43(24):5249–5262, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koutrouvelis:2010:CPA

[KCK10] Ioannis A. Koutrouvelis, George C. Canavos, and Athanasios G. Kallioras. Cumulant plots for assessing the gamma distribution. Communications in Statistics: Theory and Methods, 39(4):626– 641, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khoo:2016:SEC

[KCL^{+}16] Michael B. C. Khoo, Philippe Castagliola, J. Y. Liew, W. L. Teoh, and Petros E. Maravelakis. A study on EWMA charts with runs rules-the Markov chain approach. Communications in Statistics: Theory and Methods, 45(14):4156–4180, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.917187.

Kong:2012:ELR

[KCS12] Maiying Kong, Alex Cambon, and Michael J. Smith. Extended logistic regression model for studies with interrupted events, seasonal trend, and serial correlation. Communications in Statistics: Theory and Methods, 41(19):3528–3543, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korkmaz:2019:WMO

[KCY^{+}19] Mustafa Ç. Korkmaz, Gauss M. Cordeiro, Haitham M. Yousof, Rodrigo R. Pescim, Ahmed Z. Afify, and Saralees Nadarajah. The Weibull Marshall–Olkin family: Regression model and application to censored data. Communications in Statistics: Theory and Methods, 48(16):4171–4194, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1490430.

Kashikar:2016:ESO

[KD16a] Akanksha S. Kashikar and S. R. Deshmukh. Estimation in second order branching processes with application to swine flu data. Communications in Statistics: Theory and Methods, 45(4):1031–1046, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
853796.

Koutras:2016:GPL

[KD16b] Vasileios M. Koutras and Konstantinos Drakos. A generalized polynomial-logistic distribution and applications. Communications in Statistics: Theory and Methods, 45(18):5478–5494, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.944666.

Kyriakidis:2016:SMD

[KD16c] E. G. Kyriakidis and T. D. Dimitrakos. A semi-Markov decision model for the optimal control of a simple immigration-birth-death process through the introduction of a predator. Communications in Statistics: Theory and Methods, 45(13):3860–3873, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.911905.

Kaciranlar:2017:CGS

[KD17a] Selahattin Kaçiranlar and Issam Dawoud. Comment on a generalized stochastic restricted ridge regression estimator. Communications in Statistics: Theory and Methods, 46(10):4957–4960, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1096387.

Kumar:2017:ERK

[KD17b] C. Satheesh Kumar and S. H. S. Dharmaja. The exponentiated reduced Kies distribution: Properties and applications. Communications in Statistics: Theory and Methods, 46(17): 8778–8790, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1193199.

Koc:2018:PSO

[KDG^{+}18] Haydar Koç, Emre Dünder, Serpil Gümüstekin, Tuba Koç, and Mehmet Ali Cengiz. Particle swarm optimization-based variable selection in Poisson regression analysis via information complexity-type criteria. Communications in Statistics: Theory and Methods, 47(21):5298–5306, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1390129.

Komarudin:2016:BDP

[KDGV16] Komarudin, Tim De Feyter, Marie-Anne Guerry, and Greet Vanden Berghe. Balancing desirability and promotion steadiness in partially stochastic manpower planning systems. Communications in Statistics: Theory and Methods, 45(6):1805– 1818, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.1001495.

Koutras:2014:PLD

[KDK14] Vasileios M. Koutras, Konstantinos Drakos, and Markos V. Koutras. A polynomial logistic distribution and its applications in finance. Communications in Statistics: Theory and Methods, 43(10–12):2045–2065, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2017:EED

[KDN17] Devendra Kumar, Sanku Dey, and Saralees Nadarajah. Extended exponential distribution based on order statistics. Communications in Statistics: Theory and Methods, 46(18):9166– 9184, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1205625.

Kubokawa:2010:TLH

[KE10] Tatsuya Kubokawa and Nyambaa Erdembat. On testing linear hypothesis in a nested error regression model. Communications in Statistics: Theory and Methods, 39(8–9):1552–1562, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kempa:2011:ADP

[Kem11] Wojciech M. Kempa. Analysis of departure process in batch arrival queue with multiple vacations and exhaustive service. Communications in Statistics: Theory and Methods, 40(16): 2856–2865, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kennedy:2011:DEP

[Ken11] Nyongesa L. Kennedy. Dual estimation of prevalence and disease incidence in pool-testing strategy. Communications in Statistics: Theory and Methods, 40(18):3218–3229, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kerns:2017:CBL

[Ker17] Lucy Kerns. Confidence bands for the logistic and probit regression models over intervals. Communications in Statistics: Theory and Methods, 46(8):3878–3890, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1073319.

Key:2016:RBC

[Key16] Eric Key. A recipe for bivariate copulas. Communications in Statistics: Theory and Methods, 45(21):6416–6420, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.915040.

Kabaila:2014:SCI

[KG14a] Paul Kabaila and Khageswor Giri. Simultaneous confidence intervals for the population cell means, for two-by-two factorial data, that utilize uncertain prior information. Communications in Statistics: Theory and Methods, 43(19):4074–4087, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kodia:2014:ECS

[KG14b] B. Kodia and B. Garel. Estimation and comparison of signed symmetric covariation coefficient and generalized association parameter for alpha-stable dependence modeling. Communications in Statistics: Theory and Methods, 43(24):5156–5174, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuang:2014:HTR

[KG14c] Nenghui Kuang and Fuqing Gao. Hypothesis testing in a Rayleigh diffusion model. Communications in Statistics: Theory and Methods, 43(9):1907–1917, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Korkmaz:2015:TSG

[KG15] Mustafa Ç. Korkmaz and Ali I. Genç. Two-sided generalized exponential distribution. Communications in Statistics: Theory and Methods, 44(23):5049–5070, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundu:2017:IIE

[KG17] Chanchal Kundu and Amit Ghosh. Inequalities involving expectations of selected functions in reliability theory to characterize distributions. Communications in Statistics: Theory and Methods, 46(17):8468–8478, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1183784.

Kumar:2018:GCN

[KG18a] Narinder Kumar and Manish Goyal. A general class of non parametric tests for comparing scale parameters. Communications in Statistics: Theory and Methods, 47(24):5956–5972, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1404099.

Kundu:2018:SDG

[KG18b] Chanchal Kundu and Amit Ghosh. On some dynamic generalized information measures for bivariate lifetimes. Communications in Statistics: Theory and Methods, 47(22):5497–5513, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1395051.

Khalil:2019:EFP

[KGH19] Sadia Khalil, Sat Gupta, and Muhammad Hanif. Estimation of finite population mean in stratified sampling using scrambled responses in the presence of measurement errors. Communications in Statistics: Theory and Methods, 48(6):1553–1561, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1435817.

Kolacek:2016:SBK

[KH16] Jan Kolácek and Ivanka Horová. Selection of bandwidth for kernel regression. Communications in Statistics: Theory and Methods, 45(5):1487–1500, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
864770.

Kurata:2018:RGA

[KH18] Sumito Kurata and Etsuo Hamada. A robust generalization and asymptotic properties of the model selection criterion family. Communications in Statistics: Theory and Methods, 47(3): 532–547, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1307405.

Kurata:2019:DPM

[KH19] Sumito Kurata and Etsuo Hamada. A discrete probabilistic model for analyzing pairwise comparison matrices. Communications in Statistics: Theory and Methods, 48(15):3801–3815, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1481975.

Kharin:2011:RMS

[Kha11] Yuriy Kharin. Robustness of the mean square risk in forecasting of regression time series. Communications in Statistics: Theory and Methods, 40(16):2893–2906, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2012:IPI

[Kha12] Hafiz M. R. Khan. Improved predictive inference for responses based on location parameter median from the Rayleigh model given a doubly censored sample. Communications in Statistics: Theory and Methods, 41(24):4427–4438, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2013:PIH

[Kha13] Hafiz M. R. Khan. Predictive inference from the half-normal model given a Type II censored sample. Communications in Statistics: Theory and Methods, 42(1):42–55, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kharin:2016:PRE

[Kha16] Alexey Kharin. Performance and robustness evaluation in sequential hypotheses testing. Communications in Statistics: Theory and Methods, 45(6):1693–1709, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.944659.

Khattree:2019:NNC

[Kha19a] Ravindra Khattree. A note on the nonexistence of the constant block-sum balanced incomplete block designs. Communications in Statistics: Theory and Methods, 48(20):5165–5168, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1508715.

Khattree:2019:PYY

[Kha19b] Ravindra Khattree. The Parshvanath yantram yields a constant-sum partially balanced incomplete block design. Communications in Statistics: Theory and Methods, 48(4):841–849, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1417439.

Kritzinger:2014:IST

[KHC14] P. Kritzinger, S. W. Human, and S. Chakraborti. Improved Shewhart-type runs-rules nonparametric sign charts. Communications in Statistics: Theory and Methods, 43(22):4723–4748, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kachiashvili:2012:SAC

[KHM12] K. J. Kachiashvili, M. A. Hashmi, and A. Mueed. Sensitivity analysis of classical and conditional Bayesian problems of many hypotheses testing. Communications in Statistics: Theory and Methods, 41(4):591–605, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuwada:2017:ECB

[KHY17] Masahide Kuwada, Yoshifumi Hyodo, and Hiromu Yumiba. Existence conditions for balanced fractional 3^{m} factorial designs of resolution R(00, 10, 01, 20, 11). Communications in Statistics: Theory and Methods, 46(2):942–966, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1010004.

Khalili:2018:SPL

[KHY18] M. Khalili, A. Habibirad, and F. Yousefzadeh. Some properties of Lin-Wong divergence on the past lifetime data. Communications in Statistics: Theory and Methods, 47(14):3464–3476, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1359295.

Khan:2019:MDG

[KI19] M. J. S. Khan and S. Iqrar. On moments of dual generalized order statistics from Topp–Leone distribution. Communications in Statistics: Theory and Methods, 48(3):479–492, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1414260.

Kiani:2017:AED

[Kia17] M. Kiani. The augmentation of existing data for improving the path of steepest ascent. Communications in Statistics: Theory and Methods, 46(5):2506–2518, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2011.577547.

Kim:2011:CMB

[Kim11] Hea-Jung Kim. A class of multivariate bilateral selection t distributions and its properties. Communications in Statistics: Theory and Methods, 40(12):2136–2154, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2012:SPC

[Kim12] Hea-Jung Kim. On a screened Pearson correlation and its application to interdependence methods. Communications in Statistics: Theory and Methods, 41(19):3497–3510, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2015:LTR

[Kim15] Ji Young Kim. A lasso-type robust variable selection for timecourse microarray data. Communications in Statistics: Theory and Methods, 44(7):1411–1425, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kiraci:2012:FEL

[Kir12] Arzdar Kiraci. Formulas for the exact LMS and LQS estimators. Communications in Statistics: Theory and Methods, 41 (3):530–543, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kittaneh:2015:MDB

[Kit15] Omar Abd Al-Rahman Ibrahim Kittaneh. A measure of discrimination between two double truncated distributions. Communications in Statistics: Theory and Methods, 44(9):1797– 1805, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khare:2017:CEP

[KJ17] B. B. Khare and P. S. Jha. Classes of estimators for population mean using auxiliary variable in stratified population in the presence of non response. Communications in Statistics: Theory and Methods, 46(13):6579–6589, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1130841.

Kuo:2018:RDL

[KJ18] Kun-Lin Kuo and Thomas J. Jiang. A revisit of the distribution of linear combinations of Dirichlet components. Communications in Statistics: Theory and Methods, 47(3):509–520, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1307402.

Krishna:2013:MOF

[KJAR13] E. Krishna, K. K. Jose, T. Alice, and Miroslav M. Ristić. The Marshall–Olkin fréchet distribution. Communications in Statistics: Theory and Methods, 42(22):4091–4107, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khraibani:2015:NPE

[KJD^{+}15] Zaher Khraibani, Christine Jacob, Christian Ducrot, Myriam Charras-Garrido, and Carole Sala. A non parametric exact test based on the number of records for an early detection of emerging events: Illustration in epidemiology. Communications in Statistics: Theory and Methods, 44(4):726–749, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keerativibool:2011:MSS

[KJL11] Warangkhana Keerativibool, Jirawan Jitthavech, and Vichit Lorchirachoonkul. Model selection in a system of simultaneous equations model. Communications in Statistics: Theory and Methods, 40(3):373–393, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:2019:CST

[KJS19] N. Mamode Khan, V. Jowaheer, and Y. Sunecher. Communication in statistics-theory and methods improved GQL estimation method for the generalised BINMA(1) model. Communications in Statistics: Theory and Methods, 48(3):709–725, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1435807.

Kazemi:2017:EGL

[KJT17] M. R. Kazemi, A. A. Jafari, and S. Tahmasebi. An extension of the generalized linear failure rate distribution. Communications in Statistics: Theory and Methods, 46(16):7916–7933, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1167905.

Kawano:2011:SSL

[KK11a] Shuichi Kawano and Sadanori Konishi. Semi-supervised logistic discrimination via regularized Gaussian basis expansions. Communications in Statistics: Theory and Methods, 40(13): 2412–2423, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khare:2011:EPM

[KK11b] B. B. Khare and Sanjay Kumar. Estimation of population mean using known coefficient of variation of the study character in the presence of non response. Communications in Statistics: Theory and Methods, 40(11):2044–2058, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katayama:2014:NTH

[KK14] Shota Katayama and Yutaka Kano. A new test on highdimensional mean vector without any assumption on population covariance matrix. Communications in Statistics: Theory and Methods, 43(24):5290–5304, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2017:MSM

[KK17a] Hyoung-Moon Kim and Chiwhan Kim. Moments of scale mixtures of skew-normal distributions and their quadratic forms. Communications in Statistics: Theory and Methods, 46(3): 1117–1126, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1011339.

Kundu:2017:BED

[KK17b] Amarjit Kundu and Chanchal Kundu. Bivariate extension of (dynamic) cumulative past entropy. Communications in Statistics: Theory and Methods, 46(9):4163–4180, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1080838.

Kundu:2018:BEG

[KK18] Amarjit Kundu and Chanchal Kundu. Bivariate extension of generalized cumulative past entropy. Communications in Statistics: Theory and Methods, 47(8):1962–1977, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1335412.

Khan:2017:TWD

[KKH17] Muhammad Shuaib Khan, Robert King, and Irene Lena Hudson. Transmuted Weibull distribution: Properties and estimation. Communications in Statistics: Theory and Methods, 46(11):5394–5418, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1100744.

Kesemen:2017:TTA

[KKKÖ17] Tülay Kesemen, Zafer Küçük, Tahir Khaniyev, and Çisem Öçal. Three-term asymptotic expansion: a semi-Markovian random walk with a generalized beta distributed interference of chance. Communications in Statistics: Theory and Methods, 46(3):1445–1455, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1019148.

Kollo:2018:ANE

[KKS18] Tõnu Kollo, Meelis Käärik, and Anne Selart. Asymptotic normality of estimators for parameters of a multivariate skew-normal distribution. Communications in Statistics: Theory and Methods, 47(15):3640–3655, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361985.

Kamble:2019:CRV

[KKS19] T. S. Kamble, D. N. Kashid, and D. M. Sakate. Consistent and robust variable selection in regression based on Wald test. Communications in Statistics: Theory and Methods, 48(8):1981– 2000, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1440598.

Kang:2010:NCI

[KL10] Seung-Ho Kang and Yonghee Lee. New confidence intervals for the proportion of interest in one-sample correlated binary data. Communications in Statistics: Theory and Methods, 39 (16):2865–2876, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Konstantinides:2011:PLD

[KL11] Dimitrios G. Konstantinides and Fotis Loukissas. Precise large deviations for sums of negatively dependent random variables with common long-tailed distributions. Communications in Statistics: Theory and Methods, 40(19–20):3663–3671, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2012:TUO

[KL12] Nirpeksh Kumar and S. Lalitha. Testing for upper outliers in gamma sample. Communications in Statistics: Theory and Methods, 41(5):820–828, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2017:LFT

[KL17a] Jaehee Kim and Geung-Hee Lee. Local Fourier tests for structural change based on residuals. Communications in Statistics: Theory and Methods, 46(1):133–147, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.988259.

Kou:2017:NPR

[KL17b] Junke Kou and Youming Liu. Non parametric regression estimations over L^{p} risk based on biased data. Communications in Statistics: Theory and Methods, 46(5):2375–2395, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1044670.

Khosravi:2016:NBS

[KLJP16] Mohsen Khosravi, Víctor Leiva, Ahad Jamalizadeh, and Emilio Porcu. On a nonlinear Birnbaum–Saunders model based on a bivariate construction and its characteristics. Communications in Statistics: Theory and Methods, 45(3):772–793, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kiesse:2014:DNK

[KLK14] Tristan Senga Kiessé, Tristan Lorino, and Hussein Khraibani. Discrete nonparametric kernel and parametric methods for the modeling of pavement deterioration. Communications in Statistics: Theory and Methods, 43(6):1164–1178, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2017:OBH

[KLKK17] Yongku Kim, Woo Dong Lee, Sang Gil Kang, and Dal Ho Kim. Objective Bayesian hypothesis testing in regression models with first-order autoregressive residuals. Communications in Statistics: Theory and Methods, 46(12):5872–5887, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1112915.

Krishnamoorthy:2010:TID

[KLM10] K. Krishnamoorthy, Xiaodong Lian, and Sumona Mondal. Tolerance intervals for the distribution of the difference between two independent normal random variables. Communications in Statistics: Theory and Methods, 40(1):117–129, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kacem:2016:SMP

[KLMD16] Manel Kacem, Stéphane Loisel, and Véronique Maume-Deschamps. Some mixing properties of conditionally independent processes. Communications in Statistics: Theory and Methods, 45(5):1241–1259, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
851235.

Kuentz:2010:BVS

[KLS10] Vanessa Kuentz, Benoît Liquet, and Jérôme Saracco. Bagging versions of sliced inverse regression. Communications in Statistics: Theory and Methods, 39(11):1985–1996, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kozlowska:2012:BDN

[KLS12] Maria Kozlowska, Agnieszka Lacka, and Anna Skorupska. Block design with nested rows and columns for research on food acceptability limitation. Communications in Statistics: Theory and Methods, 41(13–14):2456–2464, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karagrigoriou:2010:DPG

[KM10] A. Karagrigoriou and K. Mattheou. On distributional properties and goodness-of-fit tests for generalized measures of divergence. Communications in Statistics: Theory and Methods, 39 (3):472–482, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalema:2014:PLM

[KM14a] George Kalema and Geert Molenberghs. Pseudo-likelihood methodology for hierarchical count data. Communications in Statistics: Theory and Methods, 43(22):4790–4805, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuchta:2014:SPM

[KM14b] Małgorzata Kuchta and Michal Morayne. A secretary problem with many lives. Communications in Statistics: Theory and Methods, 43(1):210–218, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kerman:2015:SKB

[KM15] Sean C. Kerman and James B. McDonald. Skewness–kurtosis bounds for EGB1, EGB2, and special cases. Communications in Statistics: Theory and Methods, 44(18):3857–3864, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kawamura:2016:PMA

[KM16a] Toshihiko Kawamura and Hitoshi Motoyama. Performance measures for asymmetric power loss functions. Communications in Statistics: Theory and Methods, 45(5):1260–1269, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.854914.

Kenward:2016:TMO

[KM16b] Michael G. Kenward and Geert Molenberghs. A taxonomy of mixing and outcome distributions based on conjugacy and bridging. Communications in Statistics: Theory and Methods, 45(7):1953–1968, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
870205.

Karmakar:2018:REE

[KM18] Bikram Karmakar and Indranil Mukhopadhyay. Risk efficient estimation of fully dependent random coefficient autoregressive models of general order. Communications in Statistics: Theory and Methods, 47(17):4242–4253, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1371758.

Katebi:2019:OSE

[KM19a] Mehdi Katebi and M. Bameni Moghadam. Optimal statistical, economic and economic statistical designs of attribute np control charts using a full adaptive approach. Communications in Statistics: Theory and Methods, 48(18):4528–4549, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1494837.

Kazemi:2019:NLD

[KM19b] Mohammad Reza Kazemi and Mohammad Reza Mahmoudi. A note on limiting distribution of the sample auto-covariance function for the first-order autoregressive (AR(1)) model. Communications in Statistics: Theory and Methods, 48(15):3876– 3883, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1481980.

Kairalla:2010:CIP

[KMC10] John A. Kairalla, Keith E. Muller, and Christopher S. Coffey. Combining an internal pilot with an interim analysis for single degree of freedom tests. Communications in Statistics: Theory and Methods, 39(20):3717–3738, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuriki:2012:ISP

[KMM12] Shinji Kuriki, Iwona Mejza, and Stanisław Mejza. Incomplete split-plot designs supplemented by a single control. Communications in Statistics: Theory and Methods, 41(13–14):2490– 2502, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khalaf:2013:MRR

[KMS13] G. Khalaf, Kristofer Månsson, and Ghazi Shukur. Modified ridge regression estimators. Communications in Statistics: Theory and Methods, 42(8):1476–1487, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khalaf:2014:TRR

[KMSS14] G. Khalaf, K. Månsson, P. Sjölander, and G. Shukur. A tobit ridge regression estimator. Communications in Statistics: Theory and Methods, 43(1):131–140, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundu:2010:SRP

[KN10] Chanchal Kundu and Asok K. Nanda. Some reliability properties of the inactivity time. Communications in Statistics: Theory and Methods, 39(5):899–911, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2016:GHP

[KN16a] C. Satheesh Kumar and B. Unnikrishnan Nair. On generalized hyper-Poisson distribution of order k and its application. Communications in Statistics: Theory and Methods, 45(22): 6720–6731, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.966842.

Kundu:2016:SDS

[KN16b] Amarjit Kundu and Asok K. Nanda. On study of dynamic survival and cumulative past entropies. Communications in Statistics: Theory and Methods, 45(1):104–122, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kudo:2017:DSU

[KN17a] Masaki Kudo and Kanta Naito. Data sharpening on unknown manifold. Communications in Statistics: Theory and Methods, 46(23):11721–11744, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1277756.

Kwessi:2017:SRA

[KN17b] Eddy Kwessi and Brice M. Nguelifack. Signed-rank analysis of a partial linear model with B-splines estimated monotone non parametric function. Communications in Statistics: Theory and Methods, 46(10):4843–4854, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1089289.

Kumar:2018:SAF

[KN18a] C. Satheesh Kumar and Subha R. Nair. On some aspects of a flexible class of additive Weibull distribution. Communications in Statistics: Theory and Methods, 47(5):1028–1049, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1316399.

Kundu:2018:RSP

[KN18b] Pradip Kundu and Asok K. Nanda. Reliability study of proportional odds family of discrete distributions. Communications in Statistics: Theory and Methods, 47(5):1091–1103, 2018.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1316397.

Kwong:2018:EFI

[KN18c] Hok Shing Kwong and Saralees Nadarajah. Expectation formulas for integer valued multivariate random variables. Communications in Statistics: Theory and Methods, 47(22):5514–5518, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1395052.

Kundu:2019:BDW

[KN19] Debasis Kundu and Vahid Nekoukhou. On bivariate discrete Weibull distribution. Communications in Statistics: Theory and Methods, 48(14):3464–3481, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1476712.

Kamjornkittikoon:2018:NUE

[KNC18] Kamonrat Kamjornkittikoon, Kritsana Neammanee, and Nattakarn Chaidee. Non uniform exponential bounds on normal approximation by Stein’s method and monotone size bias couplings. Communications in Statistics: Theory and Methods, 47(5):1117–1132, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1316404.

Kurosawa:2017:BRM

[KNH17] Takeshi Kurosawa, Kohei Noguchi, and Fumiaki Honda. Bias reduction of the maximum-likelihood estimator for a conditional Gaussian MA(1) model. Communications in Statistics: Theory and Methods, 46(17):8588–8602, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1185119.

Keyvandarian:2017:ACE

[KNHM17] Ali Keyvandarian, R. Noorossana, S. Mohammad Hashemian, and Shekary A. Maryam. Adaptive c-chart with estimated parameter. Communications in Statistics: Theory and Methods, 46(1):87–103, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
985842.

Kim:2013:NTS

[KNK13] Yang-Jin Kim, Chung Mo Nam, and Jinheum Kim. Nonparametric two-sample tests of longitudinal data with termination events. Communications in Statistics: Theory and Methods, 42(6):1056–1069, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaluszka:2010:BMM

[KO10a] Marek Kaluszka and Andrzej Okolewski. Bounds for moments of the maximum of concomitants of selected order statistics with application. Communications in Statistics: Theory and Methods, 39(15):2753–2766, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2010:UFM

[Ko10b] Feng-Shou Ko. Using frailty models to identify the longitudinal biomarkers in survival analysis. Communications in Statistics: Theory and Methods, 39(18):3222–3237, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2011:DOL

[Ko11] Feng-Shou Ko. Detecting outliers in large medical records: Using a frontier model. Communications in Statistics: Theory and Methods, 40(15):2772–2780, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2012:IPL

[Ko12] Feng-Shou Ko. The identification of potential longitudinal biomarkers and measurements of effectiveness for biomarkers as surrogates in multivariate survival data. Communications in Statistics: Theory and Methods, 41(15):2681–2698, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2013:AMZ

[Ko13] Mi-Hwa Ko. On the asymmetric Marcinkiewicz–Zygmund strong law of large numbers for linear random fields. Communications in Statistics: Theory and Methods, 42(8):1553–1562, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2014:ADS

[Ko14a] Feng-Shou Ko. An approach to determine sample size and to allocate sample size for a specific region in a multiregional trial for survival (time-to-event) data under accelerated failure time model. Communications in Statistics: Theory and Methods, 43(23):4925–4935, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2014:ILB

[Ko14b] Feng-Shou Ko. Identification of longitudinal biomarkers in survival analysis for competing risks data. Communications in Statistics: Theory and Methods, 43(16):3329–3342, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2014:MIB

[Ko14c] Feng-Shou Ko. The method to identify a biomarker and to evaluate its efficiency for survival by using the joint model of the accelerate failure time and longitudinal data. Communications in Statistics: Theory and Methods, 43(1):72–89, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2014:SSD

[Ko14d] Feng-Shou Ko. Sample size determination for a two-sided multiple components tolerance interval. Communications in Statistics: Theory and Methods, 43(24):5241–5248, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2014:CCW

[Ko14e] Mi-Hwa Ko. On complete convergence for weighted sums of NA random fields. Communications in Statistics: Theory and Methods, 43(17):3726–3732, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2015:SSD

[Ko15a] Feng-Shou Ko. Sample size determination and rational partition for a multi-regional trial for survival (time-to-event) data with unrecognized heterogeneity that interacts with treatment.

Communications in Statistics: Theory and Methods, 44(3):486– 496, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2015:SIA

[Ko15b] Feng-Shou Ko. A statistical issue about a phase III multiregional trial for the survival data under accelerated failure time model with unrecognized heterogeneity. Communications in Statistics: Theory and Methods, 44(12):2620–2629, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2016:IPL

[Ko16a] Feng-Shou Ko. Identification of potential longitudinal biomarkers under the accelerated failure time model in multivariate survival data. Communications in Statistics: Theory and Methods, 45(3):655–669, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2016:SSD

[Ko16b] Feng-Shou Ko. Sample size determination for quality control process using a multiple components tolerance interval. Communications in Statistics: Theory and Methods, 45(12):3668– 3674, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.904358.

Ko:2016:CLT

[Ko16c] Mi-Hwa Ko. A central limit theorem for weighted sums of associated random field. Communications in Statistics: Theory and Methods, 45(1):1–8, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2017:ERE

[Ko17a] Feng-Shou Ko. Evaluate the relative efficiency of a targeted clinical trial design to an untargeted design under the issue of cost. Communications in Statistics: Theory and Methods, 46(24):12336–12344, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1295080.

Ko:2017:CRS

[Ko17b] Mi-Hwa Ko. On convergence rate in the SLLN for maximums of moving-average sums of ALNQD random fields. Communications in Statistics: Theory and Methods, 46(2):609–615, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.1000678.

Ko:2018:CCC

[Ko18] Mi-Hwa Ko. Complete convergence for coordinatewise asymptotically negatively associated random vectors in Hilbert spaces. Communications in Statistics: Theory and Methods, 47 (3):671–680, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1310242.

Knefati:2016:LLD

[KOA16] Muhammad Anas Knefati, Abderrahim Oulidi, and Belkacem Abdous. Local linear double and asymmetric kernel estimation of conditional quantiles. Communications in Statistics: Theory and Methods, 45(12):3473–3488, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.889923.

Khosravi:2018:PRA

[KOFA18] Ramezan Khosravi, Mohammad Saleh Owlia, Mohammad Saber Fallahnezhad, and Amirhossein Amiri. Phase I risk-adjusted control charts for surgical data with ordinal outcomes. Communications in Statistics: Theory and Methods, 47(18):4422–4432, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1376085.

Koike:2010:SEP

[Koi10] Ken-Ichi Koike. Sequential estimation procedures for end points of support in a non-regular distribution. Communications in Statistics: Theory and Methods, 39(8–9):1585–1596, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kolev:2013:P

[Kol13] Nikolai Kolev. Preface. Communications in Statistics: Theory and Methods, 42(4):569, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Komarova:2017:ESE

[Kom17] Tatiana Komarova. Extremum sieve estimation in k-out-of-n systems. Communications in Statistics: Theory and Methods, 46(10):4915–4931, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1091081.

Konno:2010:EMC

[Kon10] Yoshihiko Konno. Estimation of multivariate complex normal covariance matrices under an invariant quadratic loss. Communications in Statistics: Theory and Methods, 39(8–9):1490– 1497, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Konstantinides:2011:ESR

[Kon11] Dimitrios G. Konstantinides. Extremal subexponentiality in ruin probabilities. Communications in Statistics: Theory and Methods, 40(16):2907–2918, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kong:2015:EMD

[Kon15a] Xin-Bing Kong. M-estimation for moderate deviations from a unit root. Communications in Statistics: Theory and Methods, 44(3):476–485, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kong:2015:DRR

[Kon15b] Yong Kong. Distributions of runs revisited. Communications in Statistics: Theory and Methods, 44(22):4663–4678, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kong:2016:NAE

[Kon16] Yong Kong. Number of appearances of events in random sequences: a new approach to non-overlapping runs. Communications in Statistics: Theory and Methods, 45(22):6765–6772, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.968728.

Kong:2019:JDR

[Kon19] Yong Kong. Joint distribution of rises, falls, and number of runs in random sequences. Communications in Statistics: Theory and Methods, 48(3):493–499, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1414261.

Kozachenko:2013:CRW

[KOP13] Yuriy Kozachenko, Andriy Olenko, and Olga Polosmak. Convergence rate of wavelet expansions of Gaussian random processes. Communications in Statistics: Theory and Methods, 42(21):3853–3872, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kozachenko:2014:UCC

[KOP14] Yuriy Kozachenko, Andriy Olenko, and Olga Polosmak. Uniform convergence of compactly supported wavelet expansions of Gaussian random processes. Communications in Statistics: Theory and Methods, 43(10–12):2549–2562, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koroliuk:2011:LDP

[Kor11] Volodymyr S. Koroliuk. Large deviation problems for Markov random evolution with independent increments in the scheme of asymptotically small diffusion. Communications in Statistics: Theory and Methods, 40(19–20):3385–3395, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kostrzewski:2014:BIJ

[Kos14] Maciej Kostrzewski. Bayesian inference for the jump-diffusion model with M jumps. Communications in Statistics: Theory and Methods, 43(18):3955–3985, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koutrouvelis:2016:GPG

[Kou16] Ioannis A. Koutrouvelis. Graphical procedures and goodnessof-fit tests for the compound Poisson–exponential distribution. Communications in Statistics: Theory and Methods, 45(8):

2450–2464, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.881496.

Koyuncu:2018:CEP

[Koy18] Nursel Koyuncu. Calibration estimator of population mean under stratified ranked set sampling design. Communications in Statistics: Theory and Methods, 47(23):5845–5853, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1402051.

Kozak:2012:DMC

[Koz12] Marcin Kozak. “A Dendrite Method for Cluster Analysis” by Caliński and Harabasz: a classical work that is far too often incorrectly cited. Communications in Statistics: Theory and Methods, 41(12):2279–2280, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kravchuk:2012:HLS

[KP12a] O. Y. Kravchuk and P. K. Pollett. Hodges–Lehmann scale estimator for Cauchy distribution. Communications in Statistics: Theory and Methods, 41(20):3621–3632, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kwon:2012:CSR

[KP12b] Soon S. Kwon and Sung H. Park. Conditions of sloperotatability for third-order polynomial regression models. Communications in Statistics: Theory and Methods, 41(23):4348– 4359, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karlis:2013:FBP

[KP13] Dimitris Karlis and Xanthi Pedeli. Flexible bivariate INAR(1) processes using copulas. Communications in Statistics: Theory and Methods, 42(4):723–740, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2018:LRT

[KP18a] Ilmun Kim and Sangun Park. Likelihood ratio tests for multivariate normality. Communications in Statistics: Theory and Methods, 47(8):1923–1934, 2018. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1332218.

Kundu:2018:SRR

[KP18b] Chanchal Kundu and Arijit Patra. Some results on residual life and inactivity time at random time. Communications in Statistics: Theory and Methods, 47(2):372–384, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1303735.

Karimnezhad:2019:BRB

[KP19] Ali Karimnezhad and Ahmad Parsian. Bayesian and robust Bayesian analysis in a general setting. Communications in Statistics: Theory and Methods, 48(15):3899–3920, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1482344.

Khan:2010:PIT

[KPS10] Hafiz M. R. Khan, Serge B. Provost, and Ashima Singh. Predictive inference from a two-parameter Rayleigh life model given a doubly censored sample. Communications in Statistics: Theory and Methods, 39(7):1237–1246, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishnamoorthy:2016:MLS

[KPZ16] K. Krishnamoorthy, Jie Peng, and Dan Zhang. Modified large sample confidence intervals for Poisson distributions: Ratio, weighted average, and product of means. Communications in Statistics: Theory and Methods, 45(1):83–97, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kurata:2011:LSS

[KQ11] Hiroshi Kurata and Dingxi Qiu. Linear subspace spanned by principal points of a mixture of spherically symmetric distributions. Communications in Statistics: Theory and Methods, 40(15):2737–2750, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kiapour:2017:ETS

[KQ17] A. Kiapour and M. Naghizadeh Qomi. Equal-tailed and shortest Bayesian tolerance intervals based on exponential k-records.

Communications in Statistics: Theory and Methods, 46(8): 3949–3956, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1076476.

Kazemi:2014:TDP

[KR14] Iraj Kazemi and Reyhaneh Rikhtehgaran. Topics on dynamic panel data models with random effects using semi-parametric Bayesian approach. Communications in Statistics: Theory and Methods, 43(8):1630–1648, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2015:MVL

[KR15] C. Satheesh Kumar and A. Riyaz. A modified version of logarithmic series distribution and its applications. Communications in Statistics: Theory and Methods, 44(14):3011–3021, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kohansal:2016:TNE

[KR16] Akram Kohansal and Saeid Rezakhah. Two new estimators of entropy for testing normality. Communications in Statistics: Theory and Methods, 45(18):5392–5411, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.942436.

Kozyra:2017:SBE

[KR17] Paweł Marcin Kozyra and Tomasz Rychlik. Sharp bounds on the expectations of L-statistics expressed in the Gini mean difference units. Communications in Statistics: Theory and Methods, 46(6):2921–2941, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1053937.

Khorashadizadeh:2012:CLD

[KRB12] M. Khorashadizadeh, A. H. Rezaei Roknabadi, and G. R. Mohtashami Borzadaran. Characterizations of lifetime distributions based on doubly truncated mean residual life and mean past to failure. Communications in Statistics: Theory and Methods, 41(6):1105–1115, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khorashadizadeh:2013:CLD

[KRB13] M. Khorashadizadeh, A. H. Rezaei Roknabadi, and G. R. Mohtashami Borzadaran. Characterization of life distributions using log-odds rate in discrete aging. Communications in Statistics: Theory and Methods, 42(1):76–87, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krebs:2018:BIE

[Kre18] Johannes T. N. Krebs. A Bernstein inequality for exponentially growing graphs. Communications in Statistics: Theory and Methods, 47(20):5097–5106, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1386317.

Kambo:2012:MBA

[KRH12] Nirmal S. Kambo, Alagar Rangan, and Ehsan Moghimi Hadji. Moments-based approximation to the renewal function. Communications in Statistics: Theory and Methods, 41(5):851–868, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kreidler:2019:CPG

[KRMG19] S. M. Kreidler, B. M. Ringham, K. E. Muller, and D. H. Glueck. Calculating power for the general linear multivariate model with one or more Gaussian covariates. Communications in Statistics: Theory and Methods, 48(6):1435–1448, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1433849.

Kozachenko:2011:PSR

[KRT11] Yury V. Kozachenko, Iryna V. Rozora, and Yevgen V. Turchyn. Properties of some random series. Communications in Statistics: Theory and Methods, 40(19–20):3672–3683, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kalke:2013:LCP

[KRT13] S. Kalke, W.-D. Richter, and F. Thauer. Linear combinations, products and ratios of simplicial or spherical variates. Communications in Statistics: Theory and Methods, 42(3):505–527, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaziska:2010:JGC

[KS10a] David Kaziska and Anuj Srivastava. Joint gait-cadence analysis for human identification using an elastic shape framework. Communications in Statistics: Theory and Methods, 39(10): 1817–1831, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khilare:2010:NSC

[KS10b] S. K. Khilare and D. T. Shirke. A nonparametric synthetic control chart using sign statistic. Communications in Statistics: Theory and Methods, 39(18):3282–3293, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katulska:2012:OCB

[KS12a] Krystyna Katulska and Lukasz Smaga. D-optimal chemical balance weighing designs with N ≡ 0 (mod 4) and 3 objects. Communications in Statistics: Theory and Methods, 41(13–14): 2445–2455, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubokawa:2012:AIC

[KS12b] Tatsuya Kubokawa and Muni S. Srivastava. Akaike information criterion for selecting variables in the nested error regression model. Communications in Statistics: Theory and Methods, 41(15):2626–2642, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubokawa:2012:SVM

[KS12c] Tatsuya Kubokawa and Muni S. Srivastava. Selection of variables in multivariate regression models for large dimensions. Communications in Statistics: Theory and Methods, 41(13–14): 2465–2489, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kamalja:2014:RS

[KS14a] K. K. Kamalja and R. L. Shinde. On the reliability of (n, f, k) and ⟨n, f, k⟩ systems. Communications in Statistics: Theory and Methods, 43(8):1649–1665, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khilare:2014:SSP

[KS14b] S. K. Khilare and D. T. Shirke. The steady-state performance of cumulative count of a conforming control chart based on runs rules. Communications in Statistics: Theory and Methods, 43 (15):3135–3147, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kasza:2015:CSB

[KS15a] Jessica Kasza and Patty Solomon. Comparing score-based methods for estimating Bayesian networks using the Kullback– Leibler divergence. Communications in Statistics: Theory and Methods, 44(1):135–152, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2015:ICL

[KS15b] Sungsu Kim and Ashis SenGupta. Inverse circular-linear/ linear-circular regression. Communications in Statistics: Theory and Methods, 44(22):4772–4782, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kallberg:2016:EET

[KS16a] David Källberg and Oleg Seleznjev. Estimation of entropytype integral functionals. Communications in Statistics: Theory and Methods, 45(4):887–905, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
853789.

Khan:2016:SRT

[KS16b] Zaheen Khan and Javid Shabbir. Some remarks on the traditional way of circular and diagonal circular systematic sampling schemes. Communications in Statistics: Theory and Methods, 45(7):2105–2117, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
873133.

Kayal:2017:GKI

[KS17] S. Kayal and S. M. Sunoj. Generalized Kerridge’s inaccuracy measure for conditionally specified models. Communications in Statistics: Theory and Methods, 46(16):8257–8268, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1177083.

Kumar:2018:TAM

[KS18a] Rakesh Kumar and Sapana Sharma. Transient analysis of an M/M/c queuing system with balking and retention of reneging customers. Communications in Statistics: Theory and Methods, 47(6):1318–1327, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1319485.

Kurosawa:2018:ADR

[KS18b] Takeshi Kurosawa and Nobuhiro Suzuki. Asymptotic distribution of regression correlation coefficient for Poisson regression model. Communications in Statistics: Theory and Methods, 47 (1):166–180, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1300285.

Kulkarni:2019:COR

[KS19] Leena Kulkarni and Sanjeev Sabnis. Construction of optimal reliability test plans for multi-state coherent systems of type 1. Communications in Statistics: Theory and Methods, 48(24): 6080–6097, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1528366.

Kouassi:2014:JST

[KSBM14a] Eugene Kouassi, Joel Sango, J. M. Bosson Brou, and Mbodja Mougoué. A joint score test for heteroscedasticity in the two way error components model. Communications in Statistics: Theory and Methods, 43(13):2734–2751, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kouassi:2014:CST

[KSBM14b] Eugene Kouassi, Joel Sango, Jm Bosson Brou, and Mbodja Mougoué. Conditional score tests for heteroscedasticity in the two-way error components model. Communications in Statistics: Theory and Methods, 43(18):3812–3835, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kouassi:2017:PMLa

[KSBN17] Eugene Kouassi, Patrice Takam Soh, Jean Marcelin Bosson Brou, and Emile Herve Ndoumbe. Pseudo maximum-likelihood estimation of the univariate GARCH(1, 1) and asymptotic properties. Communications in Statistics: Theory and Methods, 46(20):10253–10271, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1231824.

Kouassi:2018:AMP

[KSDB18] Eugene Kouassi, Patrice Takam Soh, Morvan N. Donfack, and Jean Marcelin B. Bosson. The ARCH(2) model: Pseudomaximum estimation and asymptotic results under dependent innovations. Communications in Statistics: Theory and Methods, 47(9):2146–2171, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1337142.

Khan:2015:IEW

[KSDR15] Hafiz M. R. Khan, Anshul Saxena, Sankalp Das, and Elizabeth Ross. Inference from the exponentiated Weibull model with applications to real data. Communications in Statistics: Theory and Methods, 44(22):4679–4695, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katebi:2017:EES

[KSF17] Mehdi Katebi, Asghar Seif, and Alireza Faraz. Economic and economic-statistical designs of the T^{2} control charts with SVSSI sampling scheme. Communications in Statistics: Theory and Methods, 46(20):10149–10165, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1231823.

Khan:2013:NSD

[KSG13] Zaheen Khan, Javid Shabbir, and Sat Gupta. A new sampling design for systematic sampling. Communications in Statistics: Theory and Methods, 42(18):3359–3370, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaarik:2015:PMS

[KSK15] Meelis Käärik, Anne Selart, and Ene Käärik. On parametrization of multivariate skew-normal distribution. Communications in Statistics: Theory and Methods, 44(9):1869–1885, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2011:EQN

[KT11] Somesh Kumar and Manas Ranjan Tripathy. Estimating quantiles of normal populations with a common mean. Communications in Statistics: Theory and Methods, 40(15):2719–2736, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kochar:2015:SCL

[KT15] Subhash C. Kochar and Nuria Torrado. On stochastic comparisons of largest order statistics in the scale model. Communications in Statistics: Theory and Methods, 44(19):4132–4143, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kamitamari:2017:SOA

[KT17a] Eita Kamitamari and Hidekazu Tanaka. On second-order admissibilities of the estimators of gamma shape vector. Communications in Statistics: Theory and Methods, 46(1):485–496, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.997359.

Kwon:2017:AIM

[KT17b] Amy M. Kwon and Gong Tang. Application of an imputation method for variance estimation under pseudo-likelihood when missing data are NMAR. Communications in Statistics: Theory and Methods, 46(14):6959–6966, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1143008.

Kozachenko:2018:CCT

[KT18] Yury V. Kozachenko and Viktor B. Troshki. Construction of a criterion for testing hypothesis about covariance function of a stationary Gaussian stochastic process with unknown mean.

Communications in Statistics: Theory and Methods, 47(18): 4556–4567, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1377253.

Kouassi:2017:PMLb

[KTBN17] Eugene Kouassi, Patrice Soh Takam, Jean Marcelin Bosson Brou, and Emile Herve Ndoumbe. Pseudo maximum likelihood estimation of the univariate GARCH(2, 2) and asymptotic normality under dependent innovations. Communications in Statistics: Theory and Methods, 46(23):11558–11574, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1275694.

Kayal:2018:EPI

[KTR18] Tanmay Kayal, Yogesh Mani Tripathi, and Manoj Kumar Rastogi. Estimation and prediction for an inverted exponentiated Rayleigh distribution under hybrid censoring. Communications in Statistics: Theory and Methods, 47(7):1615–1640, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1322702.

Khoo:2010:MPM

[KTW10] Michael B. C. Khoo, S. Y. Teh, and Zhang Wu. Monitoring process mean and variability with one double EWMA chart. Communications in Statistics: Theory and Methods, 39(20): 3678–3694, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2017:PSO

[KU17] A. N. Kumar and N. S. Upadhye. On perturbations of Stein operator. Communications in Statistics: Theory and Methods, 46(18):9284–9302, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1206937.

Kumar:2015:TSO

[Kum15] Nirpeksh Kumar. Testing of suspected observations in an exponential sample with unknown origin. Communications in Statistics: Theory and Methods, 44(17):3668–3679, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2017:CRB

[Kum17] Vikas Kumar. Characterization results based on dynamic Tsallis cumulative residual entropy. Communications in Statistics: Theory and Methods, 46(17):8343–8354, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1179757.

Kundu:2017:CDD

[Kun17a] Chanchal Kundu. Chernoff distance for doubly truncated distributions. Communications in Statistics: Theory and Methods, 46(21):10594–10606, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1239109.

Kundu:2017:WMI

[Kun17b] Chanchal Kundu. On weighted measure of inaccuracy for doubly truncated random variables. Communications in Statistics: Theory and Methods, 46(7):3135–3147, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1056365.

Kuriki:2010:DLS

[Kur10] Satoshi Kuriki. Distributions of the largest singular values of skew-symmetric random matrices and their applications to paired comparisons. Communications in Statistics: Theory and Methods, 39(8–9):1522–1535, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kushary:2018:FCB

[Kus18] Debashis Kushary. Finding confidence bound using twostage data. Communications in Statistics: Theory and Methods, 47(13):3043–3051, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1062110.

Kumar:2019:TML

[KUWG19] A. Kumar, N. S. Upadhye, A. Wyłomańska, and J. Gajda. Tempered Mittag-Leffler lévy processes. Communications in Statistics: Theory and Methods, 48(2):396–411, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1410719.

Klaassen:2010:OLC

[KV10] Chris Klaassen and Shanti Venetiaan. Optimizing lengths of confidence intervals: Fourth-order efficiency in location models. Communications in Statistics: Theory and Methods, 39(8–9): 1437–1448, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kyriakoussis:2017:HPA

[KV17] Andreas Kyriakoussis and Malvina Vamvakari. Heine process as a q-analog of the Poisson process-waiting and interarrival times. Communications in Statistics: Theory and Methods, 46(8): 4088–4102, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1078476.

Kvaalseth:2015:BSV

[Kvå15] Tarald O. Kvålseth. Bounds on sample variation measures based on majorization. Communications in Statistics: Theory and Methods, 44(16):3375–3386, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kvaalseth:2017:AMO

[Kvå17] Tarald O. Kvålseth. An alternative measure of ordinal association as a value-validity correction of the Goodman–Kruskal gamma. Communications in Statistics: Theory and Methods, 46(21):10582–10593, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1239114.

Kvaalseth:2018:ACC

[Kvå18] Tarald O. Kvålseth. An alternative to Cramér’s coefficient of association. Communications in Statistics: Theory and Methods, 47(23):5662–5674, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1400056.

Knapik:2013:BRI

[KvdVvZ13] B. T. Knapik, A. W. van der Vaart, and J. H. van Zanten. Bayesian recovery of the initial condition for the heat equation. Communications in Statistics: Theory and Methods, 42 (7):1294–1313, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumari:2017:ABM

[KVR17] J. Vasantha Kumari, R. Vasudeva, and S. Ravi. On the asymptotic behavior of maxima and near-maxima of random observations from three parameter lognormal distribution. Communications in Statistics: Theory and Methods, 46(17):8737–8747, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1189572.

Kollo:2011:EHD

[KVV11] Tõnu Kollo, Tatjana Von Rosen, and Dietrich Von Rosen. Estimation in high-dimensional analysis and multivariate linear models. Communications in Statistics: Theory and Methods, 40(7):1241–1253, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabaila:2017:PMA

[KWM17] Paul Kabaila, A. H. Welsh, and Rheanna Mainzer. The performance of model averaged tail area confidence intervals. Communications in Statistics: Theory and Methods, 46(21):10718– 10732, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1242741.

Krishnamoorthy:2018:CIT

[KX18] K. Krishnamoorthy and Yanping Xia. Confidence intervals for a two-parameter exponential distribution: One- and twosample problems. Communications in Statistics: Theory and Methods, 47(4):935–952, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1313983.

Krishnamoorthy:2011:SAP

[KXX11] K. Krishnamoorthy, Yanping Xia, and Fang Xie. A simple approximate procedure for constructing binomial and Poisson tolerance intervals. Communications in Statistics: Theory and Methods, 40(12):2243–2258, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:2010:RTL

[KY10a] Ho Gyun Kim and Won Young Yun. A repair-time limit replacement policy with estimation error. Communications in Statistics: Theory and Methods, 39(13):2365–2378, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kwan:2010:DIR

[KY10b] Isaac K. M. Kwan and Hailiang Yang. Dependent insurance risk model: Deterministic threshold. Communications in Statistics: Theory and Methods, 39(5):765–776, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ke:2016:ABA

[KY16] Jau-Chuan Ke and Dong-Yuh Yang. Analysis of a batch arrival queue with vacation policy and exceptional service. Communications in Statistics: Theory and Methods, 45(12):3637–3657, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.878356.

Kemaloglu:2017:TTP

[KY17a] Sibel Acik Kemaloglu and Mehmet Yilmaz. Transmuted twoparameter Lindley distribution. Communications in Statistics: Theory and Methods, 46(23):11866–11879, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1285933.

Khardani:2017:NLP

[KY17b] Salah Khardani and Anne Françoise Yao. Non linear parametric mode regression. Communications in Statistics: Theory and Methods, 46(6):3006–3024, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
1002940.

Kyriakidis:2016:EPP

[Kyr16] E. G. Kyriakidis. Equilibrium probabilities for a productioninventory system maintained by a control-limit policy. Communications in Statistics: Theory and Methods, 45(1):194–200, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kyriacou:2017:OSI

[Kyr17] Maria Kyriacou. Overlapping subsampling and invariance to initial conditions. Communications in Statistics: Theory and Methods, 46(2):540–553, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
999093.

Kamo:2013:BCA

[KYS13] Ken-Ichi Kamo, Hirokazu Yanagihara, and Kenichi Satoh. Biascorrected AIC for selecting variables in Poisson regression models. Communications in Statistics: Theory and Methods, 42 (11):1911–1921, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kariya:2016:EEB

[KYW16] Takeaki Kariya, Yoshiro Yamamura, and Zhu Wang. Empirically effective bond pricing model for USGBs and analysis on term structures of implied interest rates in financial crisis. Communications in Statistics: Theory and Methods, 45(6): 1580–1606, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.901377.

Krishnamoorthy:2015:AFC

[KZ15] K. Krishnamoorthy and Dan Zhang. Approximate and fiducial confidence intervals for the difference between two binomial proportions. Communications in Statistics: Theory and Methods, 44(8):1745–1759, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:2019:ITP

[KZ19] Kamlesh Kumar and Sayed Mohammed Zeeshan. Improved two phase sampling exponential ratio and product type estimators for population mean of study character in the presence of non response. Communications in Statistics: Theory and Methods, 48(9):2305–2319, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1465082.

Kotlyarova:2009:REB

[KZW09] Yulia Kotlyarova and Victoria Zinde-Walsh. Robust estimation in binary choice models. Communications in Statistics: Theory and Methods, 39(2):266–279, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lagha:2017:NPS

[LA17] Karima Lagha and Smail Adjabi. Non parametric sequential estimation of the probability density function by orthogonal series. Communications in Statistics: Theory and Methods, 46(12):5941–5955, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1115075.

Louzada:2017:BAS

[LAF17] Francisco Louzada, Anderson Ara, and Guilherme Fernandes. The bivariate alpha-skew-normal distribution. Communications in Statistics: Theory and Methods, 46(14):7147–7156, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1024865.

Lang:2014:PSS

[Lan14] Joseph B. Lang. The Pearson score statistic for multinomial-Poisson models. Communications in Statistics: Theory and Methods, 43(21):4471–4491, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lisawadi:2019:SEC

[LARS19] Supranee Lisawadi, S. Ejaz Ahmed, Orawan Reangsephet, and Muhammad Kashif Ali Shah. Simultaneous estimation of Cronbach’s alpha coefficients. Communications in Statistics: Theory and Methods, 48(13):3236–3257, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473882.

Lou:2019:EMS

[LAW^{+}19] Wenjie Lou, Erin L. Abner, Lijie Wan, David W. Fardo, Richard Lipton, Mindy Katz, and Richard J. Kryscio. Estimation of multi-state models with missing covariate values based on observed data likelihood. Communications in Statistics: Theory and Methods, 48(23):5733–5747, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1520884.

Loughin:2011:ULR

[LB11] Thomas M. Loughin and Christopher R. Bilder. On the use of a log-rate model for survey-weighted categorical data. Communications in Statistics: Theory and Methods, 40(15):2661–2669, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2016:STB

[LB16] Tao Li and Narayanaswamy Balakrishnan. Sign test based on k-tuple general ranked set samples. Communications in Statistics: Theory and Methods, 45(13):3874–3883, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.911906.

Li:2010:DMA

[LBL^{+}10] Shubiao Li, Dennis Black, Carl Lee, Felix Famoye, and Sung Li. Dependence models arising from the Lagrangian probability distributions. Communications in Statistics: Theory and Methods, 39(10):1729–1742, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2015:NDA

[LC15] Fangyu Liu and Lirong Cui. A new design on attributes single sampling plans. Communications in Statistics: Theory and Methods, 44(16):3350–3362, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2017:AML

[LC17] Kuo-Chin Lin and Yi-Ju Chen. Assessment of modeling longitudinal binary data based on graphical methods. Communications in Statistics: Theory and Methods, 46(7):3426–3437, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1062107.

Leoni:2019:PTM

[LC19] Roberto Campos Leoni and Antonio Fernando Branco Costa. The performance of the truncated mixed control chart. Communications in Statistics: Theory and Methods, 48(17):4294– 4301, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1494284.

Lou:2017:SSEb

[LCA17] Ying Lou, Jing Cao, and Chul Ahn. Sample size estimation for comparing rates of change in K-group repeated count outcomes. Communications in Statistics: Theory and Methods, 46(22):11204–11213, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1260744.

Levy:2012:MPF

[LCG12] Martin S. Levy, James J. Cochran, and Saeed Golnabi. A moment preserving finitization across the power series family of probability distributions. Communications in Statistics: Theory and Methods, 41(4):653–664, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2018:ERV

[LCG18] Chen Li, George Casella, and Malay Ghosh. Estimation of regression vectors in linear mixed models with Dirichlet process random effects. Communications in Statistics: Theory and Methods, 47(16):3935–3954, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1366519.

Li:2014:IVI

[LCLJ14] Cui-Xia Li, Jin-Yuan Chen, Zhi Liu, and Bing-Yi Jing. On integrated volatility of itô semimartingales when sampling times are endogenous. Communications in Statistics: Theory and Methods, 43(24):5263–5275, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lemonte:2014:AWD

[LCO14] Artur J. Lemonte, Gauss M. Cordeiro, and Edwin M. M. Ortega. On the additive Weibull distribution. Communications in Statistics: Theory and Methods, 43(10–12):2066–2080, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lim:2014:KDE

[LCP^{+}14] Johan Lim, Min Chen, Sangun Park, Xinlei Wang, and Lynne Stokes. Kernel density estimator from ranked set samples. Communications in Statistics: Theory and Methods, 43(10–12): 2156–2168, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2012:DCC

[LCT12] Pei-Hsi Lee, Yi-Chia Chang, and Chau-Chen Torng. A design of s control charts with a combined double sampling and variable sampling interval scheme. Communications in Statistics: Theory and Methods, 41(1):153–165, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2018:IPH

[LCW^{+}18] Bo Lu, Dingjiao Cai, Luheng Wang, Xingwei Tong, and Huiyun Xiang. Inference for proportional hazard model with propensity score. Communications in Statistics: Theory and Methods, 47(12):2908–2918, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1343849.

Liu:2016:CSR

[LCWG16] Baoliang Liu, Lirong Cui, Yanqing Wen, and Furi Guo. A cold standby repairable system with the repairman having multiple vacations and operational, repair, and vacation times following phase-type distributions. Communications in Statistics: Theory and Methods, 45(4):850–858, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.851239.

Li:2016:SEL

[LCZ16] Jinyu Li, Xingtong Chen, and Song Zhu. Smoothed empirical likelihood for GARCH models with heavy-tailed errors.

Communications in Statistics: Theory and Methods, 45(24): 7275–7293, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.978947.

Lou:2017:SSC

[LCZA17a] Ying Lou, Jing Cao, Song Zhang, and Chul Ahn. Sample size calculations for time-averaged difference of longitudinal binary outcomes. Communications in Statistics: Theory and Methods, 46(1):344–353, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
991040.

Lou:2017:SSEa

[LCZA17b] Ying Lou, Jing Cao, Song Zhang, and Chul Ahn. Sample size estimation for a two-group comparison of repeated count outcomes using GEE. Communications in Statistics: Theory and Methods, 46(14):6743–6753, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1134572.

Li:2013:MBG

[LD13] Jing Li and Sunil K. Dhar. Modeling with bivariate geometric distributions. Communications in Statistics: Theory and Methods, 42(2):252–266, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2014:MCC

[LD14] Xue Liang and Yinghui Dong. A Markov chain copula model for credit default swaps with bilateral counterparty risk. Communications in Statistics: Theory and Methods, 43(3):498–514, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ledolter:2017:RCW

[LD17] Johannes Ledolter and Franklin Dexter. On the reciprocity of connections in weighted and unweighted networks. Communications in Statistics: Theory and Methods, 46(12):5728–5737, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1085570.

Li:2018:WCQ

[LD18a] Yunxia Li and Jiali Ding. Weighted composite quantile regression method via empirical likelihood for non linear models. Communications in Statistics: Theory and Methods, 47(17): 4286–4296, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1373816.

Luo:2018:ADW

[LD18b] Jing Luo and Shan Duan. Affiliation discrete weighted networks with an increasing degree sequence. Communications in Statistics: Theory and Methods, 47(24):6079–6094, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1406116.

Liu:2018:AQA

[LDL18] Wenzhao Liu, Yanlei Diao, and Anna Liu. An analysis of queryagnostic sampling for interactive data exploration. Communications in Statistics: Theory and Methods, 47(16):3820–3837, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1363231.

Labib:2017:CFE

[LdM17] Richard Labib and Simon de Montigny. Closed-form evaluation of integrals involving the gamma function. Communications in Statistics: Theory and Methods, 46(17):8328–8342, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1179756.

Li:2018:ESS

[LDS18] Xiaofei Li, Deng Ding, and Di Sang. Exponential stability of stochastic cellular neural networks with mixed delays. Communications in Statistics: Theory and Methods, 47(19):4881–4894, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1459710.

Lin:2015:MUB

[LDY15] Xu Lin, Zhu Dongjin, and Zhou Yanru. Minimizing upper bound of ruin probability under discrete risk model with Markov chain interest rate. Communications in Statistics: Theory and Methods, 44(4):810–822, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lechner:2011:NRI

[Lec11] Michael Lechner. A note on the relation of inverse-probability-weighting and matching estimators. Communications in Statistics: Theory and Methods, 40(4):674–683, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2013:AHC

[Lee13] Ming Ha Lee. Adaptive Hotelling’s T^{2} control charts with run rules. Communications in Statistics: Theory and Methods, 42 (5):883–897, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lefebvre:2010:NGD

[Lef10] Mario Lefebvre. Nearly Gaussian distributions and application. Communications in Statistics: Theory and Methods, 39(5):823– 836, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lemonte:2014:NNA

[Lem14] Artur J. Lemonte. Non null asymptotic distributions of some criteria in censored exponential regression. Communications in Statistics: Theory and Methods, 43(15):3314–3328, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lengyel:2017:OSL

[Len17] T. Lengyel. Order statistics and the length of the best-n-of- (2n − 1) competition. Communications in Statistics: Theory and Methods, 46(1):406–414, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
995822.

Lemonte:2010:SSC

[LF10] Artur J. Lemonte and Silvia L. P. Ferrari. Small-sample corrections for score tests in Birnbaum–Saunders regressions. Communications in Statistics: Theory and Methods, 40(2):232–243, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lucini:2017:CDO

[LF17] M. Magdalena Lucini and Alejandro C. Frery. Comments on “Detecting Outliers in Gamma Distribution” by M. Jabbari Nooghabi et al. (2010). Communications in Statistics: Theory and Methods, 46(11):5223–5227, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1099669. See [NNN10].

Lehrke:2010:LSP

[LG10] S. G. Lehrke and J. K. Ghorai. Large sample properties of ML estimator of the parameters of multivariate O-U random fields. Communications in Statistics: Theory and Methods, 39 (4):738–752, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2011:LRR

[LG11] Xu-Qing Liu and Feng Gao. Linearized ridge regression estimator in linear regression. Communications in Statistics: Theory and Methods, 40(12):2182–2192, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ling:2014:BRM

[LG14] Xiao Ling and David E. Giles. Bias reduction for the maximum likelihood estimator of the parameters of the generalized Rayleigh family of distributions. Communications in Statistics: Theory and Methods, 43(8):1778–1792, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2016:OIC

[LG16a] Xiaoqing Liang and Junyi Guo. Optimal investment, consumption, and life insurance in an incomplete market. Communications in Statistics: Theory and Methods, 45(13):3884–3903, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.911907.

Liu:2016:UAB

[LG16b] Xijun Liu and Qingwu Gao. Uniformly asymptotic behavior for the tail probability of discounted aggregate claims in the time-dependent risk model with upper tail asymptotically independent claims. Communications in Statistics: Theory and Methods, 45(18):5341–5354, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.942431.

Li:2018:EIV

[LG18] Cuixia Li and Erlin Guo. Estimation of the integrated volatility using noisy high-frequency data with jumps and endogeneity. Communications in Statistics: Theory and Methods, 47(3):521– 531, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1307403.

Liu:2016:UAR

[LGG16] Xijun Liu, Qingwu Gao, and Ermin Guo. Uniform asymptotics for ruin probability of a two-dimensional dependent renewal risk model. Communications in Statistics: Theory and Methods, 45(20):6045–6060, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
955117.

Li:2018:MCA

[LGLX18] Dingfang Li, Yifan Gui, Yifan Li, and Lihua Xiong. A method for constructing asymmetric pair-copula and its application. Communications in Statistics: Theory and Methods, 47(17): 4202–4214, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1371755.

Lang:2013:ISR

[LGV13] Patrick Lang, Ann Gironella, and Rieken Venema. Invariant subspaces and regression. Communications in Statistics: Theory and Methods, 42(3):491–504, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2012:LRR

[LGX12] Xu-Qing Liu, Feng Gao, and Jian-Wen Xu. Linearized restricted ridge regression estimator in linear regression. Communications in Statistics: Theory and Methods, 41(24):4503–4514, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leon-Gonzalez:2017:BIF

[LGY17] Roberto Leon-Gonzalez and Fuyu Yang. Bayesian inference and forecasting in the stationary bilinear model. Communications in Statistics: Theory and Methods, 46(20):10327–10347, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1235193.

Lee:2011:AOE

[LH11a] Cheng-Hung Lee and Leng-Cheng Hwang. Asymptotic optimal estimation of Poisson mean under LINEX loss function. Communications in Statistics: Theory and Methods, 40(23): 4308–4321, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2011:RCP

[LH11b] Chen-Yu Li and Yu-Pin Hu. Robustness comparison of the Peña–Box model and the factor model to extract useful predictors. Communications in Statistics: Theory and Methods, 40(9):1633–1650, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2011:APM

[LH11c] Jixue Liu and Shuyuan He. Autoregressive process with measurement errors. Communications in Statistics: Theory and Methods, 40(6):1041–1048, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2013:NCM

[LH13] Yimei Li and Daniel F. Heitjan. A note on the complementary mixture Pareto II distribution. Communications in Statistics: Theory and Methods, 42(2):201–213, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leask:2014:BPD

[LH14a] Kerry L. Leask and Linda M. Haines. The beta-Poisson distribution in Wadley’s problem. Communications in Statistics: Theory and Methods, 43(23):4962–4971, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2014:REM

[LH14b] Yi-Kuei Lin and Cheng-Fu Huang. Reliability evaluation of a multi-state network with multiple sinks under individual accuracy rate constraint. Communications in Statistics: Theory and Methods, 43(21):4519–4533, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2017:SMC

[LH17a] Hong Li and Donald Hedeker. Statistical methods for continuous outcomes in partially clustered designs. Communications in Statistics: Theory and Methods, 46(8):3915–3933, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1076474.

Li:2017:TLC

[LH17b] Jiyanglin Li and Ze-Chun Hu. Toeplitz lemma, complete convergence, and complete moment convergence. Communications in Statistics: Theory and Methods, 46(4):1731–1743, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1026996.

Lin:2017:SMD

[LH17c] Chen-Yen Lin and Susan Halabi. A simple method for deriving the confidence regions for the penalized Cox’s model via the minimand perturbation. Communications in Statistics: Theory and Methods, 46(10):4791–4808, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1085568.

Liau:2014:HGL

[LHHC14] Pen-Hwang Liau, Pi-Hsiang Huang, Jui-Jung Ho, and Yen-Hung Chen. Hyper-graeco-latin squares and fractional factorial designs. Communications in Statistics: Theory and Methods, 43(10–12):2286–2296, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2019:NST

[LHS19] Gi-Sung Lee, Ki-Hak Hong, and Chang-Kyoon Son. A new stratified three-stage unrelated randomized response model for estimating a rare sensitive attribute based on the Poisson distribution. Communications in Statistics: Theory and Methods, 48(7):1585–1610, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1438625.

Li:2014:FTD

[LHZR14] Xiaowu Li, Jingyan Hao, Jie Zhou, and Dasen Ren. On a family of trimodal distributions. Communications in Statistics: Theory and Methods, 43(14):2886–2896, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2011:NEP

[Li11] Yuqiang Li. A note on estimation of parameters for branching processes with immigration. Communications in Statistics: Theory and Methods, 40(18):3337–3348, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2012:LFT

[Li12a] Chin-Shang Li. Lack-of-fit tests for generalized linear models via splines. Communications in Statistics: Theory and Methods, 41(23):4240–4250, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2012:CEV

[Li12b] Zaixing Li. A comparison of error variance estimates in nonparametric mixed models. Communications in Statistics: Theory and Methods, 41(4):778–790, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2013:KEA

[Li13a] Jiexiang Li. Kernel estimators for additive models with dependent observations from random fields. Communications in Statistics: Theory and Methods, 42(5):771–781, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2013:EAS

[Li13b] Yunxia Li. An extension of the almost sure central limit theorem for products of sums under association. Communications in Statistics: Theory and Methods, 42(3):478–490, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2014:AND

[Li14a] Jie Li. Asymptotics of the L_{p}-norms of density estimators in the nonlinear autoregressive models. Communications in Statistics:

Theory and Methods, 43(22):4845–4855, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2014:MDR

[Li14b] Lina Li. Moderate deviations results for a symmetry testing statistic based on the kernel density estimator for directional data. Communications in Statistics: Theory and Methods, 43 (14):3007–3018, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2015:TRE

[Li15] Zaixing Li. Testing for random effects in growth curve models. Communications in Statistics: Theory and Methods, 44(3):564– 572, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2016:ABN

[Li16] Jiexiang Li. On asymptotic behavior of Nadaraya–Watson regression estimator. Communications in Statistics: Theory and Methods, 45(19):5751–5761, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
948209.

Li:2017:KEM

[Li17a] Dewang Li. Kernel estimations for multivariate density functional with bootstrap. Communications in Statistics: Theory and Methods, 46(9):4631–4641, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1012392.

Li:2017:ITR

[Li17b] Jinzhu Li. The infinite-time ruin probability for a bidimensional renewal risk model with constant force of interest and dependent claims. Communications in Statistics: Theory and Methods, 46(4):1959–1971, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030428.

Li:2017:NCR

[Li17c] Jinzhu Li. A note on a by-claim risk model: Asymptotic results. Communications in Statistics: Theory and Methods, 46(22):11289–11295, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1263743.

Li:2017:PML

[Li17d] Zaixing Li. Profile maximal likelihood estimation for non linear mixed models with longitudinal data. Communications in Statistics: Theory and Methods, 46(9):4449–4463, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1085561.

Li:2018:ETC

[Li18] Yalian Li. Efficiency of two classes of stochastic restricted almost unbiased type principal component estimators in linear regression model. Communications in Statistics: Theory and Methods, 47(4):793–804, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1137598.

Li:2019:CIG

[Li19] Yong Li. Comments on “Improvement of generalized differencebased mixed Liu estimator in partially linear model”. Communications in Statistics: Theory and Methods, 48(21):5193–5194, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1508721. See [Wu18a].

Liang:2010:ELC

[Lia10] Han-Ying Liang. Empirical likelihood for conditional density under left truncation and α-mixing condition. Communications in Statistics: Theory and Methods, 39(4):707–721, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2011:ANR

[Lia11] Han-Ying Liang. Asymptotic normality for regression function estimate under truncation and α-mixing conditions. Communications in Statistics: Theory and Methods, 40(11):1999–2021, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liebscher:2016:ADU

[Lie16] Eckhard Liebscher. Approximation of distributions by using the Anderson Darling statistic. Communications in Statistics: Theory and Methods, 45(22):6732–6745, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.966844.

Liitiainen:2017:AEF

[Lii17] Elia Liitiäinen. Asymptotics for Euclidean functionals of mixing processes. Communications in Statistics: Theory and Methods, 46(23):11786–11800, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1285925.

Limnios:2011:DTS

[Lim11] Nikolaos Limnios. Discrete-time semi-Markov random evolutions — average and diffusion approximation of difference equations and additive functionals. Communications in Statistics: Theory and Methods, 40(19–20):3396–3406, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2014:AMS

[Lin14a] Chien-Tai Lin. Approximations for multiple scan statistics on the circle. Communications in Statistics: Theory and Methods, 43(10–12):2476–2488, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2014:SRA

[Lin14b] Yi-Kuei Lin. System reliability assessment through p minimal paths in stochastic case with backup-routing. Communications in Statistics: Theory and Methods, 43(3):455–469, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2016:CSC

[Lin16] Shan Lin. Comparison of simultaneous confidence bands for univariate polynomial regression over an interval. Communications in Statistics: Theory and Methods, 45(3):589–596, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Linke:2019:APO

[Lin19] Yuliana Linke. Asymptotic properties of one-step Mestimators. Communications in Statistics: Theory and Methods, 48(16):4096–4118, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1487982.

Larraz-Iribas:2013:WEP

[LIPMF13] Beatriz Larraz-Iribas, Jose M. Pavía-Miralles, and Guido Ferrari. Weighting elementary prices in consumer price index construction using spatial autocorrelation. Communications in Statistics: Theory and Methods, 42(24):4460–4475, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2011:EBC

[Liu11a] Fang Liu. Estimation bias in complete-case analysis in crossover studies with missing data. Communications in Statistics: Theory and Methods, 40(5):812–827, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2011:ANP

[Liu11b] Qiang Liu. Asymptotic normality for the partially linear EV models with longitudinal data. Communications in Statistics: Theory and Methods, 40(7):1149–1158, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2011:ECS

[Liu11c] Xiaofeng Liu. The effect of a covariate on standard error and confidence interval width. Communications in Statistics: Theory and Methods, 40(3):449–456, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2017:SIP

[Liu17a] Tian Liu. Statistical inference of partially linear panel data regression models with fixed individual and time effects. Communications in Statistics: Theory and Methods, 46(15):7267–7288, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1116577.

Liu:2017:CVE

[Liu17b] Yanning Liu. Covariance and variance evaluations of two estimators for drug-placebo difference in a trial with sequential parallel design. Communications in Statistics: Theory and Methods, 46(18):9194–9217, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1205613.

Liu:2017:PDS

[Liu17c] Yanning Liu. Planning the duration of a survival group sequential trial with a fixed follow-up time for all subjects. Communications in Statistics: Theory and Methods, 46(15):7460–7478, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1152488.

Liu:2019:AEC

[Liu19] Jin Liu. Averaging estimation for conditional covariance models. Communications in Statistics: Theory and Methods, 48(16):3992–4007, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1483511.

Lee:2010:NCS

[LJ10] Sang-Ho Lee and Chi-Hyuck Jun. A new control scheme always better than X-bar chart. Communications in Statistics: Theory and Methods, 39(19):3492–3503, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2012:GVS

[LJ12] Fuming Lin and Yingying Jiang. A general version of the shorttailed symmetric distribution. Communications in Statistics: Theory and Methods, 41(12):2088–2095, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lesch:2013:NEG

[LJ13] Scott M. Lesch and Daniel R. Jeske. A new exponential GOF test for data subject to multiply Type II censoring. Communications in Statistics: Theory and Methods, 42(24):4392–4410, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2018:TSP

[LJC18] Shixin Li, Hua Jin, and Wanyi Chen. Two-stage phase II trials with early stopping for effectiveness and safety as well as ineffectiveness or harm. Communications in Statistics: Theory and Methods, 47(22):5626–5638, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1400054.

Li:2018:LMD

[LJW18] Shimin Li, Hui Jiang, and Shaochen Wang. Large and moderate deviations for the total population in the nearly unstable INAR(1) model. Communications in Statistics: Theory and Methods, 47(7):1718–1730, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1324988.

Lee:2014:QRL

[LK14] Minjae Lee and Lan Kong. Quantile regression for longitudinal biomarker data subject to left censoring and dropouts. Communications in Statistics: Theory and Methods, 43(21):4628–4641, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2017:CDS

[LK17a] Ming Ha Lee and Michael B. C. Khoo. Combined double sampling and variable sampling interval np chart. Communications in Statistics: Theory and Methods, 46(23):11892–11917, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1285924.

Lee:2017:ODS

[LK17b] Ming Ha Lee and Michael B. C. Khoo. Optimal design of synthetic np control chart based on median run length. Communications in Statistics: Theory and Methods, 46(17):8544–8556, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1183790.

Lee:2017:ODM

[LK17c] Ming Ha Lee and Michael B. C. Khoo. Optimal designs of multivariate synthetic |S| control chart based on median run length. Communications in Statistics: Theory and Methods, 46(6): 3034–3053, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1048884.

Lee:2017:SDS

[LK17d] Ming Ha Lee and Michael B. C. Khoo. Synthetic double sampling s chart. Communications in Statistics: Theory and Methods, 46(12):5914–5931, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1115073.

Lee:2019:ESD

[LK19a] Ming Ha Lee and Michael B. C. Khoo. Economic-statistical design of synthetic double sampling T^{2} chart. Communications in Statistics: Theory and Methods, 48(23):5862–5876, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1523428.

Lee:2019:EGP

[LK19b] Seyoon Lee and Joseph H. T. Kim. Exponentiated generalized Pareto distribution: Properties and applications towards extreme value theory. Communications in Statistics: Theory and Methods, 48(8):2014–2038, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1441418.

Lawrence:2015:OTD

[LKK15] Jim Lawrence, Raghu Kacker, and Rüdiger Kessel. Obtaining a trapezoidal distribution. Communications in Statistics: Theory and Methods, 44(21):4586–4599, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2017:OBI

[LKK17] Woo Dong Lee, Sang Gil Kang, and Yongku Kim. Objective Bayesian inference for the ratio of the scale parameters of two Weibull distributions. Communications in Statistics: Theory and Methods, 46(10):4943–4956, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1091477.

Lauriski-Karriker:2015:IEC

[LKN15] Tonya Lauriski-Karriker and William Navidi. Improving the efficiency of the case-crossover design. Communications in Statistics: Theory and Methods, 44(12):2600–2619, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lemeshko:2011:MSD

[LL11] B. Yu. Lemeshko and S. B. Lemeshko. Models of statistic distributions of nonparametric goodness-of-fit tests in composite hypotheses testing for double exponential law cases. Communications in Statistics: Theory and Methods, 40(16):2879–2892, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2012:MEP

[LL12] Yaowu Li and Meihua Li. Moment estimation of the parameters in Rayleigh distribution with two parameters. Communications in Statistics: Theory and Methods, 41(15):2643–2660, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2014:GDR

[LL14] Fei-Bao Liang and Yi-Xin Lan. A generalized diagonal ridgetype estimator in linear regression. Communications in Statistics: Theory and Methods, 43(6):1145–1163, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2016:CUD

[LL16] Yan Liu and Min-Qian Liu. Construction of uniform designs for mixture experiments with complex constraints. Communications in Statistics: Theory and Methods, 45(8):2172–2180, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.875576.

Li:2017:AOP

[LL17a] Chen Li and Xiaohu Li. Aging and ordering properties of multivariate lifetimes with Archimedean dependence structures.

Communications in Statistics: Theory and Methods, 46(2):874– 891, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1006783.

Li:2017:SNR

[LL17b] Chen Li and Xiaohu Li. Some new results on allocation of coverage limits and deductibles to mutually independent risks. Communications in Statistics: Theory and Methods, 46(8):3934– 3948, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1076475.

Lim:2017:TET

[LL17c] Johan Lim and Sungim Lee. Testing the effect of treatment on survival time with an immediate intermediate event. Communications in Statistics: Theory and Methods, 46(8):3718–3727, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1071393.

Ling:2017:KRE

[LL17d] Nengxiang Ling and Yang Liu. The kernel regression estimation for randomly censored functional stationary ergodic data. Communications in Statistics: Theory and Methods, 46(17): 8557–8574, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1185117.

Liu:2017:MEA

[LL17e] Weiguo Liu and Jiaowan Luo. Modified Euler approximation of stochastic differential equation driven by Brownian motion and fractional Brownian motion. Communications in Statistics: Theory and Methods, 46(15):7427–7443, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1152487.

Liu:2017:SSI

[LL17f] Yanning Liu and Pilar Lim. Sample size increase during a survival trial when interim results are promising. Communications in Statistics: Theory and Methods, 46(14):6846–6863, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1137596.

Liu:2018:REM

[LL18a] Shu Liu and Heng Lian. Robust estimation and model identification for longitudinal data varying-coefficient model. Communications in Statistics: Theory and Methods, 47(11):2701–2719, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1342835.

Liu:2018:EET

[LL18b] Shu Liu and Liangyuan Liu. Efficient estimation for timedynamic longitudinal single-index model. Communications in Statistics: Theory and Methods, 47(15):3656–3674, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361986.

Li:2019:SCP

[LL19a] Chen Li and Xiaohu Li. Stochastic comparisons of parallel and series systems of dependent components equipped with starting devices. Communications in Statistics: Theory and Methods, 48 (3):694–708, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1435806.

Liu:2019:LIN

[LL19b] Chuandi Liu and Chengxiu Ling. A location-invariant nonpositive moment-type estimator of the extreme value index. Communications in Statistics: Theory and Methods, 48(5): 1166–1176, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1425448.

Liu:2019:OGS

[LL19c] Yanning Liu and Dayong Li. Optimal group sequential designs constrained on both overall and stage one error rates. Communications in Statistics: Theory and Methods, 48(20):4959–4975, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1440314.

Li:2010:CAP

[LLC10] Feng Li, Lu Lin, and Xia Cui. Covariate-adjusted partially linear regression models. Communications in Statistics: Theory and Methods, 39(6):1054–1074, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2017:SPM

[LLD17] Ping Li, Xiaoliang Ling, and Weiyong Ding. Stochastic properties of the mixed accelerated hazard models. Communications in Statistics: Theory and Methods, 46(13):6433–6445, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1129416.

Lu:2019:LLE

[LLF19] Yiqiang Lu, Feng Li, and Sanying Feng. Local linear estimation for covariate-adjusted varying-coefficient models. Communications in Statistics: Theory and Methods, 48(15):3816–3835, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1481976.

Li:2016:ELB

[LLHZ16] Jianbo Li, Yuan Li, Zhensheng Huang, and Riquan Zhang. Empirical likelihood-based serial correlation testing in partially varying coefficient single-index models. Communications in Statistics: Theory and Methods, 45(15):4471–4485, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.921306.

Li:2018:EOP

[LLLL18] Wenhan Li, Lixia Liu, Guiwen Lv, and Cuixiang Li. Exchange option pricing in jump-diffusion models based on Esscher transform. Communications in Statistics: Theory and Methods, 47(19):4661–4672, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1444180.

Liang:2016:BET

[LLM16] Han-Ying Liang, Deli Li, and Tianxuan Miao. Berry–Esséen type bound of conditional mode estimation under truncation and strong mixing assumptions. Communications in Statistics: Theory and Methods, 45(17):5077–5097, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.936561.

Lemeshko:2010:SDM

[LLP10] B. Yu Lemeshko, S. B. Lemeshko, and S. N. Postovalov. Statistic distribution models for some nonparametric goodness-of-fit tests in testing composite hypotheses. Communications in Statistics: Theory and Methods, 39(3):460–471, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2013:LPM

[LLQ13] Wanrong Liu, Xuewen Lu, and Yongcheng Qi. On local polynomial modelling of the additive risk model. Communications in Statistics: Theory and Methods, 42(11):1958–1981, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2014:ORU

[LLSM14] Zhi-Yi Lu, Le-Ping Liu, Qing-Jie Shen, and Li-Li Meng. Optimal reinsurance under VaR and CTE risk measures when ceded loss function is concave. Communications in Statistics: Theory and Methods, 43(15):3223–3247, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lenca:2010:CCE

[LLV10] P. Lenca, S. Lallich, and B. Vaillant. Construction of an off-centered entropy for the supervised learning of imbalanced classes: Some first results. Communications in Statistics: Theory and Methods, 39(3):493–507, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2018:ELB

[LLWK18] Feng Liu, Chunyan Li, Pengfei Wang, and Xinmei Kang. Empirical likelihood based diagnostics for heteroscedasticity in semiparametric varying-coefficient partially linear errorsin-variables models. Communications in Statistics: Theory and Methods, 47(22):5485–5496, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1395050.

Ling:2014:CDE

[LLY14] Nengxiang Ling, Zhihuan Li, and Wenzhi Yang. Conditional density estimation in the single functional index model for αmixing functional data. Communications in Statistics: Theory and Methods, 43(3):441–454, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2016:RET

[LLY16] Wenxue Li, Tingting Li, and Hu Yang. Restricted estimation and testing of hypothesis in linear measurement errors models. Communications in Statistics: Theory and Methods, 45(18): 5318–5330, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.942429.

Lu:2017:PCT

[LLZ17] Tong-Yu Lu, Yueqiong Lin, and Junjiang Zhong. Pairwise comparisons of treatments with ordinal responses in a twoway setting. Communications in Statistics: Theory and Methods, 46(21):10549–10563, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1239111.

Leskow:2013:RMT

[LM13] Jacek Leśkow and Mariola Molenda. Resampling methods for time series level crossings. Communications in Statistics: Theory and Methods, 42(23):4298–4322, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lorenzo:2014:IAD

[LM14] Edgardo Lorenzo and Hari Mukerjee. Improved asymptotics of a decreasing mean residual life estimator. Communications in Statistics: Theory and Methods, 43(21):4514–4518, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lenart:2016:GFT

[LM16] Adam Lenart and Trifon I. Missov. Goodness-of-fit tests for the Gompertz distribution. Communications in Statistics: Theory and Methods, 45(10):2920–2937, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.892323.

Leulmi:2018:LLE

[LM18a] Sara Leulmi and Fatiha Messaci. Local linear estimation of a generalized regression function with functional dependent data. Communications in Statistics: Theory and Methods, 47(23): 5795–5811, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1402048.

Louzada:2018:LLS

[LM18b] Francisco Louzada and Vitor A. A. Marchi. A log-location-scale lifetime distribution with censored data: Regression modeling, residuals, and global influence. Communications in Statistics: Theory and Methods, 47(11):2549–2562, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1021016.

Lin:2018:AJD

[LMC18] Xenos Chang-Shuo Lin, Daniel Wei-Chung Miao, and Wan-Ling Chao. Analysis of a jump-diffusion option pricing model with serially correlated jump sizes. Communications in Statistics: Theory and Methods, 47(4):953–979, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1315731.

Lai:2012:SSR

[LMCH12] Dejian Lai, Lemuel A. Moyé, Kuang-Chao Chang, and Robert J. Hardy. Sample size re-estimation based on two-stage analysis of variance: Interim analysis of clinical trials. Communications in Statistics: Theory and Methods, 41(22):4124–4131, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Louzada:2018:ZIN

[LMdO18] Francisco Louzada, Fernando F. Moreira, and Mauro Ribeiro de Oliveira. A zero-inflated non default rate regression model for credit scoring data. Communications in Statistics: Theory and Methods, 47(12):3002–3021, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1346803.

Laha:2013:SRE

[LMG13] Arnab Kumar Laha, K. C. Mahesh, and Dilip Kumar Ghosh. SB-robust estimators of the parameters of the wrapped normal distribution. Communications in Statistics: Theory and Methods, 42(4):660–672, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2017:CPS

[LMH17] Qing Liu, Tiantian Mao, and Taizhong Hu. Closure properties of the second-order regular variation under convolutions. Communications in Statistics: Theory and Methods, 46(1):104–119, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.985843.

Lorenzo:2015:NTN

[LMM15] Edgardo Lorenzo, Ganesh Malla, and Hari Mukerjee. A new test for new better than used in expectation lifetimes. Communications in Statistics: Theory and Methods, 44(23):4927–4939, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lorenzo:2018:NTD

[LMM18] Edgardo Lorenzo, Ganesh Malla, and Hari Mukerjee. A new test for decreasing mean residual lifetimes. Communications in Statistics: Theory and Methods, 47(12):2805–2812, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.985841.

Laksaci:2013:KCD

[LMR13] Ali Laksaci, Fethi Madani, and Mustapha Rachdi. Kernel conditional density estimation when the regressor is valued in a semi-metric space. Communications in Statistics: Theory and Methods, 42(19):3544–3570, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

LitaDaSilva:2015:SCR

[LMR15] João Lita Da Silva, João Tiago Mexia, and Luís Pedro Ramos. On the strong consistency of ridge estimates. Communications in Statistics: Theory and Methods, 44(3):617–626, 2015.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loertscher:2017:GNC

[LMT17] Simon Loertscher, Ellen V. Muir, and Peter G. Taylor. A general non-central hypergeometric distribution. Communications in Statistics: Theory and Methods, 46(9):4579–4598, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1088033.

Li:2017:GHP

[LMZ17] Huajiang Li, Yi Ma, and Hong Zhou. Generalized Holm’s procedure for multiple hypotheses testing problems. Communications in Statistics: Theory and Methods, 46(15):7503–7510, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1154158.

Ling:2013:CVO

[LN13a] Yan Ling and Paul I. Nelson. Consistency of p-values obtained by averaging over nuisance parameters. Communications in Statistics: Theory and Methods, 42(5):852–866, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopuhaa:2013:ALR

[LN13b] Hendrik P. Lopuhaä and Gabriela F. Nane. An asymptotic linear representation for the breslow estimator. Communications in Statistics: Theory and Methods, 42(7):1314–1324, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2017:TML

[LN17] Rui Li and Saralees Nadarajah. The true maximum-likelihood estimators for the generalized Gaussian distribution with p = 3, 4, 5. Communications in Statistics: Theory and Methods, 46(18):8821–8835, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1185121.

Luo:2017:ASC

[LO17] Zhongde Luo and Shide Ou. The almost sure convergence rate of the estimator of optimized certainty equivalent risk measure under α-mixing sequences. Communications in Statistics: Theory and Methods, 46(16):8166–8177, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1175630.

Liu:2019:LDM

[LOHW19] Jiaqi Liu, Zujun Ou, Liuping Hu, and Kang Wang. Lee discrepancy on mixed two- and three-level uniform augmented designs. Communications in Statistics: Theory and Methods, 48(10): 2409–2424, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1465087.

Lopez:2011:NEM

[Lop11] Olivier Lopez. Nonparametric estimation of the multivariate distribution function in a censored regression model with applications. Communications in Statistics: Theory and Methods, 40(15):2639–2660, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loukissas:2019:PLD

[Lou19] Fotios Loukissas. Precise large deviations for strong subexponential distributions and applications on a multi risk model. Communications in Statistics: Theory and Methods, 48(20): 5175–5190, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1508717.

Le:2010:DDM

[LP10] Sébastien Lê and Jérôme Pagès. DMFA: Dual multiple factor analysis. Communications in Statistics: Theory and Methods, 39(3):483–492, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2017:REG

[LP17] Guanghan Liu and Steven Piantadosi. Ridge estimation in generalized linear models and proportional hazards regressions. Communications in Statistics: Theory and Methods, 46(23):11466–11479, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1267767.

Li:2018:MCP

[LP18] Tingting Li and Zuoxiang Peng. Moment convergence of powered normal extremes. Communications in Statistics: Theory and Methods, 47(14):3453–3463, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1359294.

Lawson:2019:NGR

[LP19] Nuanpan Lawson and Chugiat Ponkaew. New generalized regression estimator in the presence of non response under unequal probability sampling. Communications in Statistics: Theory and Methods, 48(10):2483–2498, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1465091.

Lee:2017:FDT

[LPC17] Sangyeol Lee, Siyun Park, and Cathy W. S. Chen. On Fisher’s dispersion test for integer-valued autoregressive Poisson models with applications. Communications in Statistics: Theory and Methods, 46(20):9985–9994, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1228970.

Li:2018:WPI

[LPDD18] Dan-Ling Li, Jun-Xiang Peng, Chong-Yang Duan, and Ju-Min Deng. Weighted product index and its two-independent-sample comparison based on weighted sensitivity and specificity. Communications in Statistics: Theory and Methods, 47(22):5445– 5459, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1395047.

Li:2019:TOC

[LPDD19] Dan-Ling Li, Jun-Xiang Peng, Chong-Yang Duan, and Ju-Min Deng. Three optimal cut-point selection criteria based on sensitivity and specificity with user-defined weights. Communications in Statistics: Theory and Methods, 48(3):742–754, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1435809.

Lee:2017:IEM

[LPdDTH17] Shen-Ming Lee, Ter-Chao Peng, Jean de Dieu Tapsoba, and Shu-Hui Hsieh. Improved estimation methods for unrelated question randomized response techniques. Communications in Statistics: Theory and Methods, 46(16):8101–8112, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1175626.

Lin:2016:ASC

[LPH16] Zhengyan Lin, Tian-Xiao Pang, and Kyo-Shin Hwang. An almost sure central limit theorem for self-normalized partial sums of weakly dependent random variables. Communications in Statistics: Theory and Methods, 45(12):3411–3420, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.776688.

Lin:2018:GSP

[LPHC18] Chen-Ju Lin, W. L. Pearn, J. Y. Huang, and Y. H. Chen. Group selection for processes with multiple quality characteristics. Communications in Statistics: Theory and Methods, 47(16): 3923–3934, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1364392.

Li:2013:BET

[LPLW13] Yongming Li, Junhao Peng, Yufang Li, and Chengdong Wei. A Berry–Esséen type bound of wavelet estimator under linear process errors based on a strong mixing sequence. Communications in Statistics: Theory and Methods, 42(22):4146–4155, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liao:2014:TBL

[LPN14] Xin Liao, Zuoxiang Peng, and Saralees Nadarajah. Tail behavior and limit distribution of maximum of logarithmic general error distribution. Communications in Statistics: Theory and Methods, 43(24):5276–5289, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leonenko:2011:SSM

[LPS11] N. N. Leonenko, S. Petherick, and A. Sikorskii. The student subordinator model with dependence for risky asset returns. Communications in Statistics: Theory and Methods, 40(19–20): 3509–3523, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2013:ELC

[LQ13] Yufang Li and Yongsong Qin. Empirical likelihood confidence intervals for distribution functions under negatively associated samples. Communications in Statistics: Theory and Methods, 42(23):4357–4364, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lei:2017:EBE

[LQ17a] Qingzhu Lei and Yongsong Qin. Empirical Bayes estimation in continuous one-parameter exponential families under associated samples. Communications in Statistics: Theory and Methods, 46(7):3621–3630, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1066820.

Lu:2017:PLD

[LQ17b] Dawei Lu and Xiaomeng Qi. Precise large deviations for ϕmixing and UND random variables with long-tailed distributions. Communications in Statistics: Theory and Methods, 46(11):5443–5452, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1102284.

Luo:2019:JEL

[LQ19] Shan Luo and Gengsheng Qin. Jackknife empirical likelihoodbased inferences for Lorenz curve with kernel smoothing. Communications in Statistics: Theory and Methods, 48(3):559–582, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1417426.

Luo:2018:ADA

[LQHZ18] Jing Luo, Hong Qin, Jianwei Hu, and Yong Zhang. Asymptotic distributions in affiliation networks with an increasing sequence. Communications in Statistics: Theory and Methods, 47 (2):365–371, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1303741.

Luo:2019:ADA

[LQW19] Jing Luo, Hong Qin, and Zhenghong Wang. Asymptotic distribution in affiliation finite discrete weighted networks with an increasing degree sequence. Communications in Statistics: Theory and Methods, 48(17):4195–4205, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1487984.

Luo:2017:NAD

[LQYZ17] Jing Luo, Hong Qin, Ting Yan, and Laala Zeyneb. A note on asymptotic distributions in directed exponential random graph models with bi-degree sequences. Communications in Statistics: Theory and Methods, 46(18):8852–8864, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1193202.

Leiva:2010:QCR

[LR10a] Ricardo Leiva and Anuradha Roy. A quadratic classification rule with equicorrelated training vectors for non random samples. Communications in Statistics: Theory and Methods, 40 (2):213–231, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2010:AEL

[LR10b] Xu-Qing Liu and Jian-Ying Rong. Admissible estimation for linear combination of fixed and random effects in general mixed linear models. Communications in Statistics: Theory and Methods, 39(12):2192–2204, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Louden:2010:CRU

[LR10c] Christopher Louden and Anuradha Roy. Classification rules under autoregressive and general circulant covariance. Communications in Statistics: Theory and Methods, 39(18):3294–3315, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leiva:2014:CHO

[LR14] Ricardo Leiva and Anuradha Roy. Classification of higher-order data with separable covariance and structured multiplicative or additive mean models. Communications in Statistics: Theory and Methods, 43(5):989–1012, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2019:CIM

[LR19] Yun Liu and Yeonwoo Rho. On the choice of instruments in mixed frequency specification tests. Communications in Statistics: Theory and Methods, 48(24):6098–6118, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1528367.

Li:2017:EOM

[LRC17] Yanyan Li, Damaraju Raghavarao, and Inna Chervoneva. Extensions of D-optimal minimal designs for symmetric mixture models. Communications in Statistics: Theory and Methods, 46(5):2542–2558, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
988258.

Li:2017:EEL

[LRZ17] Danping Li, Ximin Rong, and Hui Zhao. Equilibrium excessof-loss reinsurance-investment strategy for a mean-variance insurer under stochastic volatility model. Communications in Statistics: Theory and Methods, 46(19):9459–9475, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1212071.

Lee:2009:GEM

[LS09] O. Lee and D. W. Shin. Geometric ergodicity and moment conditions for a seasonal GARCH model with periodic coefficients. Communications in Statistics: Theory and Methods, 39 (1):38–51, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2011:WIR

[LS11] Yushu Li and Ghazi Shukur. Wavelet improvement of the overrejection of unit root test under GARCH errors: an application to Swedish immigration data. Communications in Statistics: Theory and Methods, 40(13):2385–2396, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2012:ABB

[LS12] Dawei Lu and Lixin Song. The asymptotic behavior of a Brownian motion with a drift from a random domain. Communications in Statistics: Theory and Methods, 41(1):62–75, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leu:2014:NHP

[LS14] Szu-Yun Leu and Pranab K. Sen. Non-homogeneous Poisson process model for genetic crossover interference. Communications in Statistics: Theory and Methods, 43(1):44–71, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2015:BHM

[LS15a] Qie Li and Junfeng Shang. A Bayesian hierarchical model for multiple comparisons in mixed models. Communications in Statistics: Theory and Methods, 44(23):5071–5090, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2015:SAFa

[LS15b] Dawei Lu and Lixin Song. Some asymptotic formulas for a Brownian motion with a regular variation from a parabolic domain. Communications in Statistics: Theory and Methods, 44(9):1763–1778, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:2016:IMP

[LS16a] Ye Liang and Dongchu Sun. Identifiability of masking probabilities in competing risks models with emphasis on Weibull models. Communications in Statistics: Theory and Methods, 45(7): 2143–2157, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2013.878358.

Link:2016:MPD

[LS16b] William A. Link and John R. Sauer. Modeling participation duration, with application to the North American Breeding Bird Survey. Communications in Statistics: Theory and Methods, 45(21):6311–6320, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
957854.

Li:2017:OBE

[LS17a] Dandan Li and Siva Sivaganesan. An objective Bayesian estimation in a two-period crossover design. Communications in Statistics: Theory and Methods, 46(15):7409–7426, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1152486.

Lin:2017:CMP

[LS17b] Chen-Ju Lin and Yi-Ling Shen. Comparing multiple process overall yields from multiple manufacturing lines. Communications in Statistics: Theory and Methods, 46(23):11764–11775, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1280169.

Liu:2017:ISS

[LS17c] Fen Liu and Yimin Shi. Inference for a simple stepstress model with progressively censored competing risks data from Weibull distribution. Communications in Statistics: Theory and Methods, 46(14):7238–7255, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1147585.

Liu:2017:LMD

[LS17d] Jian Liu and Lei Shang. Large and moderate deviations for modified Engel series. Communications in Statistics: Theory and Methods, 46(16):8295–8306, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1177086.

Lorestani:2017:MSU

[LS17e] Hamid Lorestani and Abdolreza Sayyareh. Model selection using union-intersection principle for non nested models. Communications in Statistics: Theory and Methods, 46(4):1636–1649, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1024863.

Lu:2012:ODT

[LSC12] Jun Lu, Lei Shi, and Fei Chen. Outlier detection in time series models using local influence method. Communications in Statistics: Theory and Methods, 41(12):2202–2220, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Louzada:2013:FLT

[LSC13] F. Louzada, A. K. Suzuki, and V. G. Cancho. The FGM longterm bivariate survival copula model: Modeling, Bayesian estimation, and case influence diagnostics. Communications in Statistics: Theory and Methods, 42(4):673–691, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2017:RAM

[LSCW17] Bin Liu, Yimin Shi, Jing Cai, and Ruibing Wang. Reliability analysis of masked data in adaptive step-stress partially accelerated lifetime tests with progressive removal. Communications in Statistics: Theory and Methods, 46(12):6174–6191, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1122058.

Li:2010:CAA

[LSD10] Bao-Hong Li, Zhen-Qiu Sun, and Shi-Fu Dong. Correspondence analysis and its application in oncology. Communications in Statistics: Theory and Methods, 39(7):1229–1236, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2016:QSC

[LSL16] Siming Li, Yao Sheng, and Yong Li. The quasi-stochastically constrained least squares method for ill linear regression. Communications in Statistics: Theory and Methods, 45(2):217–225, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2018:PLD

[LSL18] Dawei Lu, Lixin Song, and Henan Li. Precise large deviations for sums of WUOD and ϕ-mixing random variables with dominated variation. Communications in Statistics: Theory and Methods, 47(19):4791–4807, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1445864.

Lu:2019:UAD

[LSL19] Dawei Lu, Lixin Song, and Fuqi Li. Uniform asymptotics for discounted aggregate claims in dependent multi-risk model. Communications in Statistics: Theory and Methods, 48(4):781– 793, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1417435.

Li:2014:DCC

[LSSS14] Chung-I Li, Nan-Cheng Su, Pei-Fang Su, and Yu Shyr. The design of, and R control charts for, skew normal distributed data. Communications in Statistics: Theory and Methods, 43 (23):4908–4924, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2016:PLD

[LSW16] Dawei Lu, Lixin Song, and Xiaohu Wang. Precise large deviation for the difference of two sums of random variables. Communications in Statistics: Theory and Methods, 45(2):291–306, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liao:2017:PPO

[LSW17] Juan Liao, Huisheng Shu, and Chao Wei. Pricing power options with a generalized jump diffusion. Communications in Statistics: Theory and Methods, 46(22):11026–11046, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1257138.

Lu:2014:PLD

[LSX14] Dawei Lu, Lixin Song, and Ying Xu. Precise large deviations for sums of independent random variables with consistently varying tails. Communications in Statistics: Theory and Methods, 43(1):28–43, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2017:PDR

[LSY17] Zhong Li, Kristina P. Sendova, and Chen Yang. On a perturbed dual risk model with dependence between intergain times and gain sizes. Communications in Statistics: Theory and Methods, 46(21):10507–10517, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1236959.

Liu:2017:CMC

[LSYL17] Xi Liu, Yan Shen, Jie Yang, and Yamei Lu. Complete moment convergence of widely orthant dependent random variables. Communications in Statistics: Theory and Methods, 46(14): 7256–7265, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1148728.

Liu:2013:ATM

[LSZ13] Yutao Liu, Liuquan Sun, and Yong Zhou. Additive transformation models for recurrent events. Communications in Statistics: Theory and Methods, 42(22):4043–4055, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2016:LDS

[LSZ16] Dawei Lu, Lixin Song, and Tao Zhang. Large deviations for sum of UEND and ϕ-mixing random variables with heavy tails. Communications in Statistics: Theory and Methods, 45(7): 2118–2129, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2013.873134.

Liu:2017:VES

[LTC17] Junfeng Liu, Donglei Tang, and Yuquan Cang. Variations and estimators for self-similarity parameter of sub-fractional Brownian motion via Malliavin calculus. Communications in Statistics: Theory and Methods, 46(7):3276–3289, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.819923.

Li:2017:MDC

[LTCQ17] Fuxiao Li, Zheng Tian, Zhanshou Chen, and Peiyan Qi. Monitoring distributional changes of squared residuals in GARCH models. Communications in Statistics: Theory and Methods, 46 (1):354–372, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.995819.

Liu:2017:TSCa

[LTGK17a] Feng Liu, Xiangyong Tan, Sitong Guo, and Xinmei Kang. Testing for serial correlation in linear model with validation data. Communications in Statistics: Theory and Methods, 46(12): 6074–6084, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1116580.

Liu:2017:TSCb

[LTGK17b] Feng Liu, Xiangyong Tan, Sitong Guo, and Xinmei Kang. Testing serial correlation in partially linear models with validation data. Communications in Statistics: Theory and Methods, 46(19):9795–9806, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1222431.

Lone:2016:GRP

[LTS16] Hilal A. Lone, Rajesh Tailor, and Housila P. Singh. Generalized ratio-cum-product type exponential estimator in stratified random sampling. Communications in Statistics: Theory and Methods, 45(11):3302–3309, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
901379.

Lone:2017:ESC

[LTV17] Hilal A. Lone, Rajesh Tailor, and Med Ram Verma. Efficient separate class of estimators of population mean in stratified random sampling. Communications in Statistics: Theory and Methods, 46(2):554–573, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
999094.

Liang:2019:THD

[LTZ19] Jiajuan Liang, Man-Lai Tang, and Xuejing Zhao. Testing highdimensional normality based on classical skewness and kurtosis with a possible small sample size. Communications in Statistics: Theory and Methods, 48(23):5719–5732, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1520882.

Lu:2010:RVM

[Lu10] Yunqing Lu. A revised version of McNemar’s test for paired binary data. Communications in Statistics: Theory and Methods, 39(19):3525–3539, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2011:CCO

[Lu11] Yunqing Lu. Considering the concordant observations in likelihood ratio test for paired binary data. Communications in Statistics: Theory and Methods, 40(23):4214–4232, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2013:UDS

[Lu13a] Chang-Yu Lu. Universal domination and stochastic domination — an improved lower bound for the dimensionality. Communications in Statistics: Theory and Methods, 42(23):4276–4286, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2013:EPT

[Lu13b] Yiqiang Lu. Estimation for the power-transformed varyingcoefficient quantile regression model. Communications in Statistics: Theory and Methods, 42(14):2617–2633, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2014:SAF

[Lu14] Dawei Lu. Some asymptotic formulas for a Brownian motion from the maximum and minimum domains with regular varying boundary. Communications in Statistics: Theory and Methods, 43(18):3848–3865, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2015:SAFb

[Lu15] Dawei Lu. Some asymptotic formulas for a Brownian motion from the maximum and minimum complicated domains. Communications in Statistics: Theory and Methods, 44(15):3192– 3217, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2016:NEM

[Lu16a] Dawei Lu. A note on the estimates of multivariate Gaussian probability. Communications in Statistics: Theory and Methods, 45(5):1459–1465, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
863936.

Lu:2016:SAF

[Lu16b] Dawei Lu. Some asymptotic formulas of a Brownian motion with regular variation from the maximum and minimum complicated domains. Communications in Statistics: Theory and Methods, 45(22):6569–6595, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.960590.

Lu:2016:AFI

[Lu16c] Shin-Li Lu. Applying fast initial response features on GWMA control charts for monitoring autocorrelation data. Communications in Statistics: Theory and Methods, 45(11):3344–3356, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.904348.

Lu:2018:NPD

[Lu18] Shin-Li Lu. Non parametric double generally weighted moving average sign charts based on process proportion. Communications in Statistics: Theory and Methods, 47(11):2684–2700, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1342832.

Leiva:2010:SSN

[LVBS10] Víctor Leiva, Filidor Vilca, N. Balakrishnan, and Antonio Sanhueza. A skewed sinh-normal distribution and its properties and application to air pollution. Communications in Statistics: Theory and Methods, 39(3):426–443, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopez-Vazquez:2019:EUT

[LVH19] Carlos López-Vázquez and Esther Hochsztain. Extended and updated tables for the Friedman rank test. Communications in Statistics: Theory and Methods, 48(2):268–281, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1408829.

Li:2012:MLE

[LvR12] Ying Li and Dietrich von Rosen. Maximum likelihood estimators in a two step model for PLS. Communications in Statistics: Theory and Methods, 41(13–14):2503–2511, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lumelskii:2011:ACL

[LVVN11] Ya. Lumelskii, V. Voinov, E. Voinov, and M. Nikulin. Approximate confidence limits for a proportion of the Pólya distribution. Communications in Statistics: Theory and Methods, 40(9):1601–1619, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2013:SEA

[LW13a] Daniel Li and Liqun Wang. A semiparametric estimation approach for linear mixed models. Communications in Statistics: Theory and Methods, 42(11):1982–1997, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2013:RAD

[LW13b] Xuan Li and Xikui Wang. Response adaptive designs with misclassified responses. Communications in Statistics: Theory and Methods, 42(11):2071–2083, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2014:ELP

[LW14] Zheng-Yan Lin and Ran Wang. Empirical likelihood for partial linear models under negatively associated errors. Communications in Statistics: Theory and Methods, 43(19):4088–4102, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2015:ELS

[LW15] Zheng-Yan Lin and Ran Wang. Empirical likelihood for singleindex regression models under negatively associated errors.

Communications in Statistics: Theory and Methods, 44(9): 1854–1868, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Laukaityte:2017:UPV

[LW17a] Inga Laukaityte and Marie Wiberg. Using plausible values in secondary analysis in large-scale assessments. Communications in Statistics: Theory and Methods, 46(22):11341–11357, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1267764.

Lee:2017:UTA

[LW17b] Bu Hyoung Lee and William W. S. Wei. The use of temporally aggregated data on detecting a mean change of a time series process. Communications in Statistics: Theory and Methods, 46(12):5851–5871, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1091082.

Lin:2017:DCT

[LW17c] S. H. Lin and R. S. Wang. Discussion on the common threshold of several exponential distributions with different rates. Communications in Statistics: Theory and Methods, 46(12):5779– 5794, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1112912.

Laukaityte:2018:ISW

[LW18a] Inga Laukaityte and Marie Wiberg. Importance of sampling weights in multilevel modeling of international largescale assessment data. Communications in Statistics: Theory and Methods, 47(20):4991–5012, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1383429.

Lv:2018:IHR

[LW18b] Siyu Lv and Zhen Wu. Infinite horizon reflected backward stochastic differential equations with Markov chains. Communications in Statistics: Theory and Methods, 47(14):3360–3376, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1353629.

Liu:2018:JDC

[LWD^{+}18] Guo-Xiang Liu, Meng-Meng Wang, Xiu-Li Du, Jin-Guan Lin, and Qi-Bing Gao. Jump-detection and curve estimation methods for discontinuous regression functions based on the piecewise B-spline function. Communications in Statistics: Theory and Methods, 47(23):5729–5749, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1400061.

Liu:2018:FTR

[LWG18a] Rongfei Liu, Dingcheng Wang, and Fenglong Guo. The finite-time ruin probability of a discrete-time risk model with GARCH discounted factors and dependent risks. Communications in Statistics: Theory and Methods, 47(17):4170–4186, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1371753.

Liu:2018:RPD

[LWG18b] Rongfei Liu, Dingcheng Wang, and Fenglong Guo. The ruin probabilities of a discrete time risk model with one-sided linear claim sizes and dependent risks. Communications in Statistics: Theory and Methods, 47(7):1529–1550, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1202281.

Liu:2012:EPL

[LWH12] X. H. Liu, Z. Z. Wang, and X. M. Hu. Estimation in partially linear single-index models with missing covariates. Communications in Statistics: Theory and Methods, 41(18):3428–3447, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2019:MCR

[LWH19] Wei Liu, Linxiao Wei, and Yijun Hu. Multivariate convex risk statistics with scenario analysis. Communications in Statistics: Theory and Methods, 48(22):5585–5601, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1515957.

Lu:2016:BHT

[LWHB16] Yonggang Lu, Peter H. Westfall, Gang Han, and Marilyn M. Bui. Bayesian hypothesis testing for selected regression coefficients. Communications in Statistics: Theory and Methods, 45(23):7011–7026, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
972745.

Liu:2011:TSC

[LWHW11] Xiaohui Liu, Zhizhong Wang, Xuemei Hu, and Guofu Wang. Testing serial correlation in partially linear single-index errorsin-variables models. Communications in Statistics: Theory and Methods, 40(14):2554–2573, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2013:CMW

[LWL13] Wen-Chuan Lee, Jong-Wuu Wu, and Chun-Te Li. Characterization of mixtures of Weibull and Pareto distributions based on conditional expectation of order statistics. Communications in Statistics: Theory and Methods, 42(3):413–428, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2011:BQU

[LWR11] Xu-Qing Liu, Yan-Dong Wu, and Jian-Ying Rong. Best quadratic unbiased prediction in a general linear model with stochastic regression coefficients. Communications in Statistics: Theory and Methods, 40(8):1418–1433, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:2014:SMD

[LWWH14] Ren-Fen Lee, Hsu-Sheng Wang, Wen-Hung Wu, and Deng-Yuan Huang. Some multiple decision procedures for comparing several population proportions. Communications in Statistics: Theory and Methods, 43(10–12):2236–2249, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2016:EIW

[LWWH16] Tingting Liu, Xuejun Wang, Xinghui Wang, and Shuhe Hu. On the exponential inequalities for widely orthantdependent random variables. Communications in Statistics: Theory and Methods, 45(19):5848–5856, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.950752.

Lu:2018:ESS

[LWWR18] Rong Lu, Danxin Wang, Min Wang, and Grzegorz A. Rempala. Estimation of Sobol’s sensitivity indices under generalized linear models. Communications in Statistics: Theory and Methods, 47(21):5163–5195, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1388397.

Lin:2016:SMO

[LWWW16] Chun-Yi Lin, P. C. Wang, Meijie Wang, and Ming-Hao Wang. A simple method to obtain minimum aberration split-plot designs. Communications in Statistics: Theory and Methods, 45(10):2988–2997, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
894066.

Li:2010:RKD

[LWY10] Yongming Li, Chengdong Wei, and Shanchao Yang. The recursive kernel distribution function estimator based on negatively and positively associated sequences. Communications in Statistics: Theory and Methods, 39(20):3585–3595, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2013:AGS

[LWZ13] Xiaohu Li, Yudan Wu, and Zhengchang Zhang. On the allocation of general standby redundancy to series and parallel systems. Communications in Statistics: Theory and Methods, 42(22):4056–4069, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:2017:NRV

[LWZ17] Yunqing Lu, Min Wang, and Gengsheng Zhang. A new revised version of McNemar’s test for paired binary data. Communications in Statistics: Theory and Methods, 46(20):10010–10024, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1228962.

Lin:2018:SET

[LWZ18] Cunjie Lin, Wenhua Wei, and Yong Zhou. Semiparametric estimation of treatment effect with density ratio model. Communications in Statistics: Theory and Methods, 47(14):3338–3359, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1353628.

Li:2017:MRE

[LWZD17] Yongjin Li, Qihua Wang, Liping Zhu, and Xiaobo Ding. Mean response estimation with missing response in the presence of high-dimensional covariates. Communications in Statistics: Theory and Methods, 46(2):628–643, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.1002935.

Li:2010:DFS

[LX10a] Fei Li and Yifeng Xue. The density functions of the singular quaternion normal matrix and the singular quaternion Wishart matrix. Communications in Statistics: Theory and Methods, 39(18):3316–3331, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2010:NBT

[LX10b] Linyuan Li and Yimin Xiao. A note on block-thresholded wavelet estimators with correlated noise. Communications in Statistics: Theory and Methods, 39(7):1111–1128, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2011:CNR

[LX11] Na Li and Xingzhong Xu. Comparison of nonparametric regression curves by spline smoothing. Communications in Statistics: Theory and Methods, 40(22):3972–3987, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2013:FEB

[LX13] Yonghui Liu and Chenqi Xia. Fundamental equations of BLUE and BLUP in the multivariate linear model with applications.

Communications in Statistics: Theory and Methods, 42(3):398– 412, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2015:EIG

[LX15] Wanbin Li and Liugen Xue. Efficient inference in a generalized partially linear model with random effect for longitudinal data. Communications in Statistics: Theory and Methods, 44(2):241– 260, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2017:WBE

[LX17] Linyuan Li and Yimin Xiao. Wavelet-based estimation of regression function with strong mixing errors under fixed design. Communications in Statistics: Theory and Methods, 46(10): 4824–4842, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1089288.

Liang:2019:DJL

[LX19] Xuexue Liang and Zhiming Xia. Detection of jump location curve in spatial linear regression model with twodimensional threshold. Communications in Statistics: Theory and Methods, 48(4):876–888, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1422186.

Li:2017:ASH

[LXL17] Wanxing Li, Xiaoming Xue, and Yonghong Long. An additive subdistribution hazard model for competing risks data. Communications in Statistics: Theory and Methods, 46(23):11667– 11687, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1277759.

Liu:2017:GEL

[LXT17] Juanfang Liu, Liugen Xue, and Ruiqin Tian. Generalized empirical likelihood inference in partially linear model for longitudinal data with missing response variables and errorprone covariates. Communications in Statistics: Theory and Methods, 46(19):9743–9762, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1218027.

Li:2010:TKR

[LY10] Yalian Li and Hu Yang. Two kinds of restricted estimators in singular linear regression. Communications in Statistics: Theory and Methods, 39(14):2594–2603, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2011:PJM

[LY11a] Yalian Li and Hu Yang. On the performance of the jackknifed modified ridge estimator in the linear regression model with correlated or heteroscedastic errors. Communications in Statistics: Theory and Methods, 40(15):2695–2708, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:2011:EGE

[LY11b] Yanyan Liu and Zhongshang Yuan. Estimated generalized estimating equation for correlated failure time data with auxiliary covariates. Communications in Statistics: Theory and Methods, 40(22):4086–4103, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2016:IPW

[LY16a] Tingting Li and Hu Yang. Inverse probability weighted estimators for single-index models with missing covariates. Communications in Statistics: Theory and Methods, 45(5):1199–1214, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2012.705208.

Li:2016:MTP

[LY16b] Yalian Li and Hu Yang. More on the two-parameter estimation in the restricted regression. Communications in Statistics: Theory and Methods, 45(24):7184–7196, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.978944.

Liu:2016:PRS

[LY16c] Chaolin Liu and Hu Yang. Positive-rule Stein-type almost unbiased ridge estimator in linear regression model. Communications in Statistics: Theory and Methods, 45(8):2228–2255, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.879180.

Liu:2017:PCQ

[LY17] Huilan Liu and Hu Yang. Penalized composite quantile estimation for censored regression model with a diverging number of parameters. Communications in Statistics: Theory and Methods, 46(13):6558–6578, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1130840.

Liu:2017:PLD

[LYG17] Xijun Liu, Changjun Yu, and Qingwu Gao. Precise large deviations of aggregate claim amount in a dependent renewal risk model. Communications in Statistics: Theory and Methods, 46(5):2354–2363, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1044666.

Li:2011:ARS

[LYH11] Xiaohu Li, Rongfang Yan, and Xiaoxiao Hu. On the allocation of redundancies in series and parallel systems. Communications in Statistics: Theory and Methods, 40(6):959–968, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Linder:2018:IFI

[LYR^{+}18] Daniel F. Linder, Jingjing Yin, Haresh Rochani, Hani Samawi, and Sanjay Sethi. Increased Fisher’s information for parameters of association in count regression via extreme ranks. Communications in Statistics: Theory and Methods, 47(5):1181–1203, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1316859.

Li:2011:MPT

[LYX11] Yalian Li, Hu Yang, and Jianwen Xu. More on the preliminary test estimator in almost unbiased Liu regression. Communications in Statistics: Theory and Methods, 40(13):2292–2304, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2018:SLA

[LYX18] Zhi Li, Litan Yan, and Liping Xu. Stepanov-like almost automorphic solutions for stochastic differential equations with lévy noise. Communications in Statistics: Theory and Methods, 47(6):1350–1371, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1319482.

Li:2010:MPO

[LZ10] Xiaohu Li and Peng Zhao. On the mixture of proportional odds models. Communications in Statistics: Theory and Methods, 40 (2):333–344, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ling:2014:GMM

[LZ14] Xiaoliang Ling and Peng Zhao. On general multivariate mixture models. Communications in Statistics: Theory and Methods, 43(20):4223–4240, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:2018:RSA

[LZ18] Guanghui Li and Chongqi Zhang. Random search algorithm for optimal mixture experimental design. Communications in Statistics: Theory and Methods, 47(6):1413–1422, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321122.

Liu:2017:MVG

[LZF17] Guangying Liu, Lixin Zhang, and Xinian Fang. Multipower variation from generalized difference for fractional integral processes with jumps. Communications in Statistics: Theory and Methods, 46(19):9662–9678, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1217019.

Luo:2015:AOA

[LZH15] Chun Luo, Yingshan Zhang, and Sihui He. Asymmetrical orthogonal arrays with run size 100. Communications in Statistics: Theory and Methods, 44(6):1222–1240, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2011:RCS

[LZPJ11] Fuming Lin, Xinhua Zhang, Zuoxiang Peng, and Yingying Jiang. On the rate of convergence of STSD extremes. Communications in Statistics: Theory and Methods, 40(10):1795–1806, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:2017:LEV

[LZS17] Hongmei Lin, Riquan Zhang, and Jianhong Shi. Local estimation for varying-coefficient models with longitudinal data. Communications in Statistics: Theory and Methods, 46(15): 7511–7528, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1154156.

Lai:2017:CML

[LZY^{+}17] Dejian Lai, Qiang Zhang, Jose-Miguel Yamal, Paula T. Einhorn, and Barry R. Davis. Conditional moving linear regression: Modeling the recruitment process for ALLHAT. Communications in Statistics: Theory and Methods, 46(18):8943–8951, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1197251.

Li:2017:REP

[LZZL17] Zean Li, Weihua Zhao, Riquan Zhang, and Jicai Liu. Robust estimation for partially linear single-index model. Communications in Statistics: Theory and Methods, 46(11):5342–5356, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1100739.

Mateos-Aparicio:2011:PLS

[MA11] Gregoria Mateos-Aparicio. Partial least squares (PLS) methods: Origins, evolution, and application to social sciences. Communications in Statistics: Theory and Methods, 40(13): 2305–2317, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mirfarah:2014:CRP

[MA14] Elham Mirfarah and Jafar Ahmadi. Closeness of k-records to progressive Type-II censored order statistics for location-scale families. Communications in Statistics: Theory and Methods, 43(10–12):2169–2182, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mesfioui:2017:MLD

[MA17] Mhamed Mesfioui and Abdulrahman M. Abouammoh. On a multivariate Lindley distribution. Communications in Statistics: Theory and Methods, 46(16):8027–8045, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1171355.

Mehta:2018:BES

[MA18] Samridhi Mehta and Priyanka Aggarwal. Bayesian estimation of sensitivity level and population proportion of a sensitive characteristic in a binary optional unrelated question RRT model. Communications in Statistics: Theory and Methods, 47(16):4021–4028, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1367812.

MirMostafaee:2017:BPM

[MAA17] S. M. T. K. MirMostafaee, Morteza Amini, and A. Asgharzadeh. Bayesian prediction of minimal repair times of a series system based on hybrid censored sample of components’ lifetimes under Rayleigh distribution. Communications in Statistics: Theory and Methods, 46(4):1788–1806, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030418.

Moradi:2017:NFM

[MAAI17] Z. Ghayour Moradi, M. Arashi, O. Arslan, and Anis Iranmanesh. A new family of multivariate slash distributions. Communications in Statistics: Theory and Methods, 46(7):3264– 3275, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2013.781644.

Muttlak:2010:EUR

[MADSAS10] H. A. Muttlak, W. A. Abu-Dayyeh, M. F. Saleh, and E. Al-Sawi. Estimating P (Y < X) using ranked set sampling in case of the exponential distribution. Communications in Statistics: Theory and Methods, 39(10):1855–1868, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maesono:2010:EEN

[Mae10] Yoshihiko Maesono. Edgeworth expansion and normalizing transformation of ratio statistics and their application. Communications in Statistics: Theory and Methods, 39(8–9):1344– 1358, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meena:2018:EPS

[MAG18] K. R. Meena, Mohd. Arshad, and Aditi Kar Gangopadhyay. Estimating the parameter of selected uniform population under the squared log error loss function. Communications in Statistics: Theory and Methods, 47(7):1679–1692, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1324986.

Mahadik:2012:ERV

[Mah12a] Shashibhushan B. Mahadik. Exact results for variable sampling interval Shewhart control charts with runs rules for switching between sampling interval lengths. Communications in Statistics: Theory and Methods, 41(24):4453–4469, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mahmoudi:2012:AME

[Mah12b] Eisa Mahmoudi. Admissible and minimax estimators of θ^{r} with truncated parameter space under squared-log error loss function. Communications in Statistics: Theory and Methods, 41(7):1242–1253, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mahadik:2017:UAA

[Mah17] Shashibhushan B. Mahadik. A unified approach to adaptive Shewhart control charts. Communications in Statistics: Theory and Methods, 46(20):10272–10293, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1235192.

Maleki:2017:OCE

[Mal17] Yasaman Maleki. Optimal covariance estimation of discretetime locally self-similar processes in time-scale and ambiguity domains. Communications in Statistics: Theory and Methods, 46(10):4700–4712, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1069352.

Manda:2011:NFM

[Man11a] Samuel O. M. Manda. A nonparametric frailty model for clustered survival data. Communications in Statistics: Theory and Methods, 40(5):863–875, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Manita:2011:BPI

[Man11b] Anatoly Manita. Brownian particles interacting via synchronizations. Communications in Statistics: Theory and Methods, 40(19–20):3440–3451, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maansson:2012:WBA

[Mån12] Kristofer Månsson. A wavelet-based approach of testing for Granger causality in the presence of GARCH effects. Communications in Statistics: Theory and Methods, 41(4):717–728, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2017:BDR

[Man17a] N. K. Mandal. Block designs robust against the presence of positional effects. Communications in Statistics: Theory and Methods, 46(5):2290–2298, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1041982.

Mandal:2017:MNO

[Man17b] Nripes Kumar Mandal. Measures of non orthogonality in block designs. Communications in Statistics: Theory and Methods, 46(5):2364–2369, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1044667.

Mansouri:2018:WAU

[Man18] Behzad Mansouri. Wavelet analysis of uniformly timemodulated processes. Communications in Statistics: Theory and Methods, 47(8):2036–2041, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1335416.

Mao:2015:EPE

[Mao15] Guangyu Mao. Efficient penalized estimation for linear regression model. Communications in Statistics: Theory and Methods, 44(7):1436–1449, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mao:2017:RTI

[Mao17] Guangyu Mao. Robust test for independence in high dimensions. Communications in Statistics: Theory and Methods, 46(20):10036–10050, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1228965.

Mao:2018:TEC

[Mao18] Guangyu Mao. Testing for error cross-sectional uncorrelatedness in a two-way error components panel data model. Communications in Statistics: Theory and Methods, 47(19):4808–4839, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1446087.

Martynov:2011:WCM

[Mar11] Gennadi Martynov. Weighted Cramér-von Mises test with estimated parameters. Communications in Statistics: Theory and Methods, 40(19–20):3569–3586, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marozzi:2012:MHP

[Mar12] Marco Marozzi. A modified Hall–Padmanabhan test for the homogeneity of scales. Communications in Statistics: Theory and Methods, 41(16–17):3068–3078, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
606483.

Martin:2015:AON

[Mar15] Ryan Martin. Asymptotically optimal nonparametric empirical Bayes via predictive recursion. Communications in Statistics: Theory and Methods, 44(2):286–299, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marozzi:2016:CDF

[Mar16] Marco Marozzi. Combination of distribution-free tests for the general two-sample problem with application to the social sciences. Communications in Statistics: Theory and Methods, 45(21):6421–6435, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
919398.

Masry:2015:LOS

[Mas15] Elias Masry. On linear operations on stationary and non stationary random processes. Communications in Statistics: Theory and Methods, 44(10):2104–2106, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meintanis:2016:DTD

[MAS16] Simos G. Meintanis, James S. Allison, and Leonard Santana. Diagnostic tests for the distribution of random effects in multivariate mixed effects models. Communications in Statistics: Theory and Methods, 45(1):201–215, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matsuura:2011:OPP

[Mat11] Shun Matsuura. Optimal partitioning of probability distributions under general convex loss functions in selective assembly. Communications in Statistics: Theory and Methods, 40(9): 1545–1560, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matteucci:2014:IPR

[Mat14] Mariagiulia Matteucci. An investigation of parameter recovery in MCMC estimation for the additive IRT model. Communications in Statistics: Theory and Methods, 43(4):751–770, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matsuura:2018:CCR

[Mat18] Shun Matsuura. Correlation coefficient of response variances in compound noise and product array experiments. Communications in Statistics: Theory and Methods, 47(3):725–738, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1313985.

Maleki:2018:PMC

[MATC18] M. R. Maleki, A. Amiri, A. R. Taheriyoun, and P. Castagliola. Phase I monitoring and change point estimation of autocorrelated Poisson regression profiles. Communications in Statistics: Theory and Methods, 47(24):5885–5903, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1402052.

Merovci:2017:ETF

[MAYH17] Faton Merovci, Morad Alizadeh, Haitham M. Yousof, and G. G. Hamedani. The exponentiated transmuted-G family of distributions: Theory and applications. Communications in Statistics: Theory and Methods, 46(21):10800–10822, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1248782.

Mazzoleni:2019:ASP

[Maz19] Marcella Mazzoleni. The analysis of student paths at the University using the multivariate joint models. Communications in Statistics: Theory and Methods, 48(1):25–39, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1390683.

Mukhopadhyay:2010:NMV

[MB10] Nitis Mukhopadhyay and Debanjan Bhattacharjee. A note on minimum variance unbiased estimation. Communications in Statistics: Theory and Methods, 39(8–9):1466–1476, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minkova:2014:BPA

[MB14] Leda D. Minkova and N. Balakrishnan. On a bivariate Pólya– Aeppli distribution. Communications in Statistics: Theory and Methods, 43(23):5026–5038, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathew:2015:CDL

[MB15] Deemat C. Mathew and M. I. Beg. Characterizing discrete life distributions by properties of reversed variance residual life. Communications in Statistics: Theory and Methods, 44(13): 2753–2763, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maqsood:2017:SEL

[MB17a] Arfa Maqsood and S. M. Aqil Burney. Standard errors for the Laspeyres index number with autocorrelated error models. Communications in Statistics: Theory and Methods, 46(21):10607–10616, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1239113.

Mirali:2017:DVW

[MB17b] M. Mirali and S. Baratpour. Dynamic version of weighted cumulative residual entropy. Communications in Statistics: Theory and Methods, 46(22):11047–11059, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1257711.

Mahalingam:2019:EDQ

[MB19] Usha Mahalingam and Saminathan Balamurali. Economic design of quick switching sampling system for resubmitted lots. Communications in Statistics: Theory and Methods, 48(16): 4019–4033, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1484141.

Mirzaei:2019:NGW

[MBAJ19] S. Mirzaei, G. R. Mohtashami Borzadaran, M. Amini, and H. Jabbari. A new generalized Weibull distribution in income economic inequality curves. Communications in Statistics: Theory and Methods, 48(4):889–908, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1422754.

Mirali:2017:WCR

[MBF17] M. Mirali, S. Baratpour, and V. Fakoor. On weighted cumulative residual entropy. Communications in Statistics: Theory and Methods, 46(6):2857–2869, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1053932.

Mortezanejad:2019:ESC

[MBG19] Seyedeh Azadeh Fallah Mortezanejad, Gholamreza Mohtashami Borzadaran, and Bahram Sadeghpour Gildeh. An entropic structure in capability indices. Communications in Statistics: Theory and Methods, 48(23):5911–5921, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1524910.

Moslemi:2014:REM

[MBN14] Amir Moslemi, Mahdi Bashiri, and Seyed Taghi Akhavan Niaki. Robust estimation of multi-response surfaces considering correlation structure. Communications in Statistics: Theory and Methods, 43(22):4749–4765, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Magalhaes:2017:CCM

[MBS17] Tiago M. Magalhães, Denise A. Botter, and Mônica C. Sandoval. Corrigendum to: “Covariance matrix formula for generalized linear models with unknown dispersion” by G. M. Cordeiro, L. P. Barroso, and D. A. Botter [Communications in Statistics — Theory and Methods (2006) 35(1), 113–120]. Communications in Statistics: Theory and Methods, 46(21):10702– 10704, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1242745. See [CBB06].

Magalhaes:2019:SML

[MBS^{+}19] Tiago M. Magalhães, Denise A. Botter, Mônica C. Sandoval, Gustavo H. A. Pereira, and Gauss M. Cordeiro. Skewness of maximum likelihood estimators in the varying dispersion beta regression model. Communications in Statistics: Theory and Methods, 48(17):4250–4260, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1490768.

Marbac:2017:MBC

[MBV17] Matthieu Marbac, Christophe Biernacki, and Vincent Vandewalle. Model-based clustering of Gaussian copulas for mixed data. Communications in Statistics: Theory and Methods, 46(23):11635–11656, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1277753.

Miao:2011:SLT

[MC11] Yu Miao and Ying-Xia Chen. Some limit theorems of survival function estimator for m-dependent processes. Communications in Statistics: Theory and Methods, 40(12):2193–2204, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Machado:2014:SCR

[MC14] Marcela A. G. Machado and Antonio F. B. Costa. Some comments regarding the synthetic chart. Communications in Statistics: Theory and Methods, 43(14):2897–2906, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See comments [SG17d].

Matheson:2017:SHF

[MC17] Matthew B. Matheson and Christopher Cox. The shape of the hazard function: Does the generalized gamma have the last word? Communications in Statistics: Theory and Methods, 46(23):11657–11666, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1277757.

Makinde:2018:SCB

[MC18] Olusola Makinde and Biman Chakraborty. On some classifiers based on multivariate ranks. Communications in Statistics: Theory and Methods, 47(16):3955–3969, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1366520.

Munir:2019:RCE

[MC19] Qaiser Munir and Kok Sook Ching. Revisiting calendar effects in Malaysian finance stocks market: Evidence from threshold GARCH (TGARCH) model. Communications in Statistics: Theory and Methods, 48(6):1377–1400, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1433846.

Monsalve-Cobis:2017:SII

[MCGMS17] A. E. Monsalve-Cobis, W. González-Manteiga, and W. Stute. The statistical impact of inflation on interest rates. Communications in Statistics: Theory and Methods, 46(14):6754–6763, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1130842.

Miao:2010:APD

[MCX10] Yu Miao, Ying-Xia Chen, and Shou-Fang Xu. Asymptotic properties of the deviation between order statistics and p-quantile. Communications in Statistics: Theory and Methods, 40(1):8– 14, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:2010:DSR

[MCY10] Robert L. Mason, Youn-Min Chou, and John C. Young. Decomposition of scatter ratios used in monitoring multivariate process variability. Communications in Statistics: Theory and Methods, 39(12):2128–2145, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:2011:DIM

[MCY11] Robert L. Mason, Youn-Min Chou, and John C. Young. Detection and interpretation of a multivariate signal using combined charts. Communications in Statistics: Theory and Methods, 40 (5):942–957, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:2014:RBS

[MCY14] Robert L. Mason, Youn-Min Chou, and John C. Young. The relationship between the T^{2} statistic and the influence function. Communications in Statistics: Theory and Methods, 43(13): 2844–2857, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2017:BIL

[MD17a] B. N. Mandal and Sukanta Dash. On balanced incomplete Latin square designs. Communications in Statistics: Theory and Methods, 46(13):6258–6263, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1124119.

Morin:2017:CVW

[MD17b] Kristi Morin and John L. Davis. Cross-validation: What is it and how is it used in regression? Communications in Statistics: Theory and Methods, 46(11):5238–5251, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1099672.

Mirhosseini:2016:BKT

[MDA16] S. M. Mirhosseini, A. Dolati, and M. Amini. On a bivariate Kumaraswamy type exponential distribution. Communications in Statistics: Theory and Methods, 45(18):5461–5477, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.944664.

Merzougui:2016:TPR

[MDC16] M. Merzougui, H. Dridi, and A. Chadli. Test for periodicity in restrictive EXPAR models. Communications in Statistics: Theory and Methods, 45(9):2770–2783, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.887110.

Ma:2019:EET

[MDG19] Junmei Ma, Kun Du, and Guiding Gu. An efficient exponential twisting importance sampling technique for pricing financial derivatives. Communications in Statistics: Theory and Methods, 48(2):203–219, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1530788.

Maouche:2017:SPS

[MDR17] Fouad Maouche, Abdelnasser Dahmani, and Nadji Rahmania. A stochastic procedure to solve linear ill-posed problems. Communications in Statistics: Theory and Methods, 46(3):1519– 1531, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1019153.

Mead:2015:GIG

[Mea15] M. E. Mead. Generalized inverse gamma distribution and its application in reliability. Communications in Statistics: Theory and Methods, 44(7):1426–1435, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mahmoud:2015:FIM

[MEG15] Mahmoud Riad Mahmoud and Amani Shaheen Abd El-Ghafour. Fisher information matrix for the generalized Feller– Pareto distribution. Communications in Statistics: Theory and Methods, 44(20):4396–4407, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehdi:2017:PCC

[Meh17] Tahsin Mehdi. Poverty comparisons with common relative poverty lines. Communications in Statistics: Theory and Methods, 46(4):2029–2036, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1040504.

Melnykov:2013:PFD

[Mel13] Igor Melnykov. A positive false discovery rate convergence result. Communications in Statistics: Theory and Methods, 42 (23):4239–4246, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:2010:CCV

[Men10a] Ulrich Menzefricke. Control charts for the variance and coefficient of variation based on their predictive distribution. Communications in Statistics: Theory and Methods, 39(16):2930– 2941, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:2010:MEW

[Men10b] Ulrich Menzefricke. Multivariate exponentially weighted moving average charts for a mean based on its prediction distribution. Communications in Statistics: Theory and Methods, 39(16):2942–2960, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:2013:CEW

[Men13a] Ulrich Menzefricke. Combined exponentially weighted moving average charts for the mean and variance based on the predictive distribution. Communications in Statistics: Theory and Methods, 42(22):4003–4016, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:2013:CCM

[Men13b] Ulrich Menzefricke. Control charts for the mean and variance based on changepoint methodology. Communications in Statistics: Theory and Methods, 42(6):988–1007, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meng:2019:GSC

[Men19] Yanjiao Meng. General strong convergence for negatively associated random variables. Communications in Statistics: Theory and Methods, 48(9):2206–2217, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1459716.

Merzougui:2016:OTD

[Mer16] M. Merzougui. Optimal test for PAR(1) dependence against PSETAR(2, 1, 1) models with specified threshold. Communications in Statistics: Theory and Methods, 45(4):872–886, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.853788.

Meyer:2013:BNC

[Mey13] Christian Meyer. The bivariate normal copula. Communications in Statistics: Theory and Methods, 42(13):2402–2422, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mann:2016:RDE

[MEZ16] Rena K. Mann, Roderick Edwards, and Julie Zhou. Robust designs for experiments with blocks. Communications in Statistics: Theory and Methods, 45(18):5363–5379, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.942434.

Minh:2010:NFP

[MF10] Do Le (Paul) Minh and Nicholas R. Farnum. A new fixed point iteration to find percentage points for distributions on the positive axis. Communications in Statistics: Theory and Methods, 40(1):1–7, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Melo:2014:ALI

[MF14] Tatiane F. N. Melo and Silvia L. P. Ferrari. Adjusted likelihood inference in an elliptical multivariate errors-in-variables model. Communications in Statistics: Theory and Methods, 43(24): 5226–5240, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mingoti:2016:HTT

[MF16] Sueli Aparecida Mingoti and Graziele Umbelina Alves Ferreira. Hayter and Tsui’s test with double sampling for the vector mean of multivariate normal population. Communications in Statistics: Theory and Methods, 45(15):4365–4377, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.909940.

Manzi:2019:BBE

[MF19] Giancarlo Manzi and Martin Forster. Biases in bias elicitation. Communications in Statistics: Theory and Methods, 48(18): 4656–4674, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1500598.

Martinez-Florez:2013:APT

[MFBG13] Guillermo Martínez-Flórez, Heleno Bolfarine, and Héctor W. Gómez. The alpha-power tobit model. Communications in Statistics: Theory and Methods, 42(4):633–643, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martinez-Florez:2017:MLB

[MFFMA17] Guillermo Martínez-Flórez, Rafael Bráz Azevedo Farias, and Germán Moreno-Arenas. Multivariate log-Birnbaum–Saunders regression models. Communications in Statistics: Theory and Methods, 46(20):10166–10178, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1231818.

Mahdavi:2017:BWE

[MFJ17] Abbas Mahdavi, Malek Fathizadeh, and Ahad Jamalizadeh. On the bivariate weighted exponential distribution based on the generalized exponential distribution. Communications in Statistics: Theory and Methods, 46(8):3641–3648, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1069349.

Martinez-Florez:2016:ILA

[MFPG16] Guillermo Martínez-Flórez, Mario Pacheco, and Ramón Giraldo. Inference in log-alpha-power and log-skew-normal multivariate models. Communications in Statistics: Theory and Methods, 45(15):4397–4415, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
921301.

Martins-Filho:2012:ANL

[MFS12] Carlos Martins-Filho and Paulo Saraiva. On asymptotic normality of the local polynomial regression estimator with stochastic bandwidths. Communications in Statistics: Theory and Methods, 41(6):1052–1068, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mardani-Fard:2010:BBD

[MFSAS10] H. A. Mardani-Fard, S. M. Sadooghi-Alvandi, and Z. Shishebor. Bounds on bivariate distribution functions with given margins and known values at several points. Communications in Statistics: Theory and Methods, 39(20):3596–3621, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Modugno:2015:WBM

[MG15] Lucia Modugno and Simone Giannerini. The wild bootstrap for multilevel models. Communications in Statistics: Theory and Methods, 44(22):4812–4825, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mao:2019:MDQ

[MG19] Mingzhi Mao and Wanli Guo. Moderate deviations for quantile regression processes. Communications in Statistics: Theory and Methods, 48(12):2879–2892, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473429.

Meintanis:2014:CGF

[MGAf14] Simos G. Meintanis, Ma Dolores JimÉnez Gamero, and V. Alba-fernÁndez. A class of goodness-of-fit tests based on transformation. Communications in Statistics: Theory and Methods, 43(8):1708–1735, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McClean:2014:CMC

[MGF14] Sally McClean, Lalit Garg, and Ken Fullerton. Costing mixed Coxian phase-type systems with Poisson arrivals. Communications in Statistics: Theory and Methods, 43(7):1437–1452, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mutlu:2017:NCA

[MGG^{+}17] Hakan Tahiri Mutlu, Fikri Gökpinar, Esra Gökpinar, Hasan Hüseyin Gül, and Gamze Güven. A new computational approach test for one-way ANOVA under heteroscedasticity. Communications in Statistics: Theory and Methods, 46(16):8236–8256, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1177082.

Miao:2014:MDP

[MGLX14] Yu Miao, Weiqiang Geng, Nan Li, and Qing Xiao. Moderate deviation principle for the error variance estimator in linear models. Communications in Statistics: Theory and Methods, 43(19):4215–4222, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Melo:2013:CMM

[MGM13] Sandra E. Melo, Betty J. Garzón, and Oscar O. Melo. Cell means model for balanced factorial designs with nested mixed factors. Communications in Statistics: Theory and Methods, 42(11):2009–2024, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miao:2016:IPA

[MGN16] Yu Miao, Wei-Qiang Geng, and Saralees Nadarajah. Invariance principles and almost sure central limit theorem for the error variance estimator in linear models. Communications in Statistics: Theory and Methods, 45(11):3223–3235, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.895841.

Mandal:2011:CPD

[MGP11] B. N. Mandal, V. K. Gupta, and Rajender Parsad. Construction of polygonal designs using linear integer programming. Communications in Statistics: Theory and Methods, 40(10): 1787–1794, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2014:CEM

[MGP14] B. N. Mandal, V. K. Gupta, and Rajender Parsad. Construction of efficient multi-level k-circulant supersaturated designs. Communications in Statistics: Theory and Methods, 43(3):599– 615, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2017:BTI

[MGP17] B. N. Mandal, V. K. Gupta, and Rajender Parsad. Balanced treatment incomplete block designs through integer programming. Communications in Statistics: Theory and Methods, 46(8):3728–3737, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1071394.

Mashtare:2011:UFE

[MH11] Terry L. Mashtare, Jr. and Alan D. Hutson. Utilizing the flexibility of the epsilon-skew-normal distribution for tobit regression problems. Communications in Statistics: Theory and Methods, 40(3):408–423, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moser:2015:SSR

[MH15] Barry Kurt Moser and Susan Halabi. Sample size requirements and study duration for testing main effects and interactions in completely randomized factorial designs when time to event is the outcome. Communications in Statistics: Theory and Methods, 44(2):275–285, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mao:2016:SOR

[MH16a] Tiantian Mao and Lei Hua. Second-order regular variation inherited from Laplace–Stieltjes transforms. Communications in Statistics: Theory and Methods, 45(15):4569–4588, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.922990.

Maraun:2016:NIF

[MH16b] Michael Maraun and Moritz Heene. A note on the implications of factorial invariance for common factor variable equivalence. Communications in Statistics: Theory and Methods, 45(14): 4181–4193, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.917186.

Mahmoudi:2019:TDB

[MHA19] Mohammad Reza Mahmoudi, Mohammad Hossein Heydari, and Zakieh Avazzadeh. Testing the difference between spectral densities of two independent periodically correlated (cyclostationary) time series models. Communications in Statistics: Theory and Methods, 48(9):2320–2328, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472776.

Mousa:2016:NNP

[MHF16] A. M. Mousa, M. Kh. Hassan, and A. Fathi. A new non parametric estimator for Pdf based on inverse gamma distribution. Communications in Statistics: Theory and Methods, 45(23): 7002–7010, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.972575.

Miao:2011:DLD

[Mia11] Yu Miao. The discounted large deviation principle for autoregressive processes. Communications in Statistics: Theory and Methods, 40(14):2455–2461, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mokhlis:2017:SSR

[MIG17] Nahed A. Mokhlis, Emad J. Ibrahim, and Dina M. Gharieb. Stress-strength reliability with general form distributions. Communications in Statistics: Theory and Methods, 46(3): 1230–1246, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1014110.

Minkova:2010:PAD

[Min10] Leda D. Minkova. Pólya–Aeppli distribution of order k. Communications in Statistics: Theory and Methods, 39(3):408–415, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minkova:2011:PPA

[Min11] Leda D. Minkova. I-Pólya process and applications. Communications in Statistics: Theory and Methods, 40(16):2847–2855, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minimol:2017:GDC

[Min17] S. Minimol. On generalized dynamic cumulative past entropy measure. Communications in Statistics: Theory and Methods, 46(6):2816–2822, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1053928.

Mitnik:2013:NPK

[Mit13] Pablo A. Mitnik. New properties of the Kumaraswamy distribution. Communications in Statistics: Theory and Methods, 42 (5):741–755, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miyata:2010:AMP

[Miy10] Yoichi Miyata. Approximate marginal posterior distributions using asymptotic modes. Communications in Statistics: Theory and Methods, 39(7):1129–1140, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miyata:2017:LAB

[Miy17] Yoichi Miyata. Laplace approximations and Bayesian information criteria in possibly misspecified models. Communications in Statistics: Theory and Methods, 46(24):12240–12258, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1295079.

McDonald:2015:OCS

[MJ15] Gary C. McDonald and Feiyi Jia. Operating characteristics of a subset selection procedure applied to a hybrid randomized response model. Communications in Statistics: Theory and Methods, 44(4):750–777, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ma:2019:TIV

[MJH19] Yaolan Ma, Yuexiang Jiang, and Wei Huang. Tail index varying coefficient model. Communications in Statistics: Theory and Methods, 48(2):235–256, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1406519.

Matsuura:2013:DPD

[MK13] Shun Matsuura and Hiroshi Kurata. Definition and properties of m-dimensional n-principal points. Communications in Statistics: Theory and Methods, 42(2):267–282, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mahdavi:2017:NMG

[MK17] Abbas Mahdavi and Debasis Kundu. A new method for generating distributions with an application to exponential distribution. Communications in Statistics: Theory and Methods, 46(13):6543–6557, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.

1130839. See comment [Jon18].

Mondal:2019:NTS

[MK19] Shuvashree Mondal and Debasis Kundu. A new two sample type-II progressive censoring scheme. Communications in Statistics: Theory and Methods, 48(10):2602–2618, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472781.

Michiels:2011:NMC

[MKD11] Frederik Michiels, Inge Koch, and Ann De Schepper. A new method for the construction of bivariate Archimedean copulas based on the λ function. Communications in Statistics: Theory and Methods, 40(15):2670–2679, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mostofi:2017:RDT

[MKK17] Amin Ghalamfarsa Mostofi and Mahmood Kharrati-Kopaei. Ratio- and difference-type estimators for estimating finite population variance. Communications in Statistics: Theory and Methods, 46(7):3542–3555, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1066813.

Maansson:2015:PSW

[MKS15] Kristofer Månsson, B. M. Golam Kibria, and Ghazi Shukur. Performance of some weighted Liu estimators for logit regression model: an application to Swedish accident data. Communications in Statistics: Theory and Methods, 44(2):363–375, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maurya:2017:NCD

[MKSS17] S. K. Maurya, A. Kaushik, S. K. Singh, and U. Singh. A new class of distribution having decreasing, increasing, and bathtub-shaped failure rate. Communications in Statistics: Theory and Methods, 46(20):10359–10372, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1235196.

Mohamed:2016:CEF

[MKY16] I. Mohamed, K. Khalid, and M. S. Yahya. Combined estimating function for random coefficient models with correlated errors. Communications in Statistics: Theory and Methods, 45(4):967– 975, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2013.853794.

Ma:2012:CAR

[ML12] Yunyan Ma and Yihui Luan. Covariate-adjusted regression for time series. Communications in Statistics: Theory and Methods, 41(3):422–436, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monfared:2013:BEC

[ML13] M. Esmail Dehghan Monfared and Fazlollah Lak. Bayesian estimation of the change point using X^{−} control chart. Communications in Statistics: Theory and Methods, 42(8):1572–1582, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mansouri:2019:SCI

[ML19] Hossein Mansouri and Bo Li. On simultaneous confidence intervals based on rank-estimates with application to analysis of gene expression data. Communications in Statistics: Theory and Methods, 48(17):4339–4349, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1494287.

Merabet:2017:BPT

[MLD17] Hayet Merabet, Ahlam Labdaoui, and Pierre Druilhet. Bayesian prediction for two-stage sequential analysis in clinical trials. Communications in Statistics: Theory and Methods, 46(19):9807–9816, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1222438.

Miao:2014:SLB

[MLGZ14] Yu Miao, Nan Li, Weiqiang Geng, and Yaohua Zhang. Some limit behaviors for linear EV model with replicate observations. Communications in Statistics: Theory and Methods, 43(15): 3170–3185, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moklyachuk:2011:MPP

[MM11a] Mikhail Moklyachuk and Aleksandr Masyutka. Minimax prediction problem for multidimensional stationary stochastic processes. Communications in Statistics: Theory and Methods, 40 (19–20):3700–3710, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mu:2011:GMI

[MM11b] Jian-Yong Mu and Yu Miao. Generalizing the Marshall’s inequality. Communications in Statistics: Theory and Methods, 40(15):2809–2817, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mameli:2013:GSN

[MM13] Valentina Mameli and Monica Musio. A generalization of the skew-normal distribution: The beta skew-normal. Communications in Statistics: Theory and Methods, 42(12):2229–2244, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mohdeb:2015:TLR

[MM15] Zaher Mohdeb and Abdelkader Mokkadem. Testing linear regression models in non regular case. Communications in Statistics: Theory and Methods, 44(21):4476–4490, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maleki:2017:TPL

[MM17] Mohsen Maleki and Mohammad Reza Mahmoudi. Two-piece location-scale distributions based on scale mixtures of normal family. Communications in Statistics: Theory and Methods, 46(24):12356–12369, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1295160.

Magnus:2019:ENM

[MM19] Gideon Magnus and Jan R. Magnus. The estimation of normal mixtures with latent variables. Communications in Statistics: Theory and Methods, 48(5):1255–1269, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1429625.

Mazucheli:2018:IML

[MMD18] Josmar Mazucheli, André Felipe Berdusco Menezes, and Sanku Dey. Improved maximum-likelihood estimators for the parameters of the unit-gamma distribution. Communications in Statistics: Theory and Methods, 47(15):3767–3778, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361993.

MirMostafaee:2017:MOE

[MML17] S. M. T. K. MirMostafaee, M. Mahdizadeh, and Artur J. Lemonte. The Marshall–Olkin extended generalized Rayleigh distribution: Properties and applications. Communications in Statistics: Theory and Methods, 46(2):653–671, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.1002937.

Mahmoudi:2016:TDB

[MMN16] Mohammad Reza Mahmoudi, Marziyeh Mahmoudi, and Elaheh Nahavandi. Testing the difference between two independent regression models. Communications in Statistics: Theory and Methods, 45(21):6284–6289, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.960584.

Mattei:2014:ILC

[MMP14] Alessandra Mattei, Fabrizia Mealli, and Barbara Pacini. Identification of local causal effects with missing outcome values and an instrument for non response. Communications in Statistics:

Theory and Methods, 43(4):815–825, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mojirsheibani:2017:DRE

[MMP17] Majid Mojirsheibani, Kevin Manley, and William Pouliot. On density and regression estimation with incomplete data. Communications in Statistics: Theory and Methods, 46(23):11688– 11711, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1277751.

Mahmoudi:2018:TET

[MMP18] Mohammad Reza Mahmoudi, Mohsen Maleki, and Abbas Pak. Testing the equality of two independent regression models. Communications in Statistics: Theory and Methods, 47(12): 2919–2926, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1343847.

Mauler:2017:PAQ

[MMT17] David J. Mauler, James B. McDonald, and Logan C. Tatham. Partially adaptive quantile estimators. Communications in Statistics: Theory and Methods, 46(11):5327–5341, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1096391.

Matteucci:2012:PDI

[MMV12] Mariagiulia Matteucci, Stefania Mignani, and Bernard P. Veldkamp. Prior distributions for item parameters in IRT models. Communications in Statistics: Theory and Methods, 41(16–17): 2944–2958, 2012. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/03610926.2011.639973.

Main:2010:CSE

[MN10] Paloma Main and Hilario Navarro. Conditional specification with exponential power distributions. Communications in Statistics: Theory and Methods, 39(12):2231–2240, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mavroudis:2013:ECC

[MN13a] Eleftheriou Mavroudis and Farmakis Nicolas. EWMA control charts for monitoring high yield processes. Communications in Statistics: Theory and Methods, 42(20):3639–3654, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meeden:2013:HTF

[MN13b] Glen Meeden and Siamak Noorbaloochi. Hypotheses testing as a fuzzy set estimation problem. Communications in Statistics: Theory and Methods, 42(10):1806–1820, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moore:2015:CSB

[MN15] Brian Moore and Balasubramaniam Natarajan. A compressive sensing based analysis of anomalies in generalized linear models. Communications in Statistics: Theory and Methods, 44(13):2705–2719, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maleki:2017:BAE

[MN17] M. Maleki and A. R. Nematollahi. Bayesian approach to epsilon-skew-normal family. Communications in Statistics: Theory and Methods, 46(15):7546–7561, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1157186.

Misra:2018:UAS

[MN18] Neeraj Misra and Sameen Naqvi. A unified approach to stochastic comparisons of multivariate mixture models. Communications in Statistics: Theory and Methods, 47(19):4724–4740, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1445859.

Mansour:2019:ESS

[MN19] S. Mehr Mansour and M. Niaparast. The effect of small sample on optimal designs for logistic regression models. Communications in Statistics: Theory and Methods, 48(12):2893–2903, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1473592.

Mnatsakanov:2017:RFT

[Mna17] Robert M. Mnatsakanov. Recovery of functions from transformed moments: a unified approach. Communications in Statistics: Theory and Methods, 46(7):3174–3185, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1056369.

Nam:2017:IRT

[mNK17] Jun mo Nam and Deukwoo Kwon. Inference on the ratio of two coefficients of variation of two lognormal distributions. Communications in Statistics: Theory and Methods, 46(17):8575–8587, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1185118.

Mahmoudi:2019:RTI

[MNM19] Mohammad Reza Mahmoudi, Roya Nasirzadeh, and Mahmoud Mohammadi. On the ratio of two independent skewnesses. Communications in Statistics: Theory and Methods, 48(7): 1721–1727, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1438627.

Mahalingam:2010:STH

[MO10] Nalini Mahalingam and Alwell Oyet. Score test for homogeneity of variances in longitudinal time series via wavelets. Communications in Statistics: Theory and Methods, 39(12):2157–2171, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minkova:2014:NMB

[MO14a] Leda D. Minkova and Edward Omey. A new Markov binomial distribution. Communications in Statistics: Theory and Methods, 43(13):2674–2688, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mitov:2014:BPI

[MO14b] Kosto V. Mitov and Edward Omey. A branching process with immigration in varying environments. Communications in Statistics: Theory and Methods, 43(24):5211–5225, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malov:2018:LTE

[MO18] Sergey V. Malov and Stephen J. O’Brien. Life table estimator revisited. Communications in Statistics: Theory and Methods, 47(9):2126–2133, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1335418.

Mendoza:2016:ELL

[MOC16] Natalie V. R. Mendoza, Edwin M. M. Ortega, and Gauss M. Cordeiro. The exponentiated-log-logistic geometric distribution: Dual activation. Communications in Statistics: Theory and Methods, 45(13):3838–3859, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
909937.

Modarres:2012:TEE

[Mod12] Reza Modarres. Testing the equality and exchangeability of bivariate distributions. Communications in Statistics: Theory and Methods, 41(20):3750–3758, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Modarres:2017:EDF

[Mod17] Reza Modarres. Estimating the distribution function of symmetric pairs. Communications in Statistics: Theory and Methods, 46(4):1843–1854, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1030421.

Mohammadi:2019:JRS

[Moh19a] Leila Mohammadi. The joint reliability signature of order statistics. Communications in Statistics: Theory and Methods, 48(19):4730–4747, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473596.

Mohammadi:2019:CIC

[Moh19b] Mohammad Mohammadi. Confidence intervals for the closed population size under inverse sampling without replacement. Communications in Statistics: Theory and Methods, 48(14):

3518–3529, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1476718.

Monari:2012:FPS

[Mon12] Paola Monari. Fortunato Pesarin: a statistician for complex problems. Communications in Statistics: Theory and Methods, 41(16–17):2833–2835, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2012.
694734.

Monod:2014:REM

[Mon14] Anthea Monod. Random effects modeling and the zero-inflated Poisson distribution. Communications in Statistics: Theory and Methods, 43(4):664–680, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:2010:NML

[MP10a] N. Martin and L. Pardo. A new measure of leverage cells in multinomial loglinear models. Communications in Statistics: Theory and Methods, 39(3):517–530, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mulero:2010:BAP

[MP10b] Julio Mulero and Franco Pellerey. Bivariate aging properties under Archimedean dependence structures. Communications in Statistics: Theory and Methods, 39(17):3108–3121, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandal:2012:ODE

[MP12] Nripes Kumar Mandal and Manisha Pal. Optimum designs for estimation of optimum amount and proportions in a mixtureamount model. Communications in Statistics: Theory and Methods, 41(11):1989–1999, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

MartIn:2014:NIM

[MP14] Nirian MartÍn and Leandro Pardo. New influence measures in polytomous logistic regression models based on phi-divergence measures. Communications in Statistics: Theory and Methods, 43(10–12):2311–2321, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Makri:2015:ORO

[MP15] Frosso S. Makri and Zaharias M. Psillakis. On l l-overlapping runs of ones of length k in sequences of independent binary random variables. Communications in Statistics: Theory and Methods, 44(18):3865–3884, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malik:2016:GET

[MP16] Waqas Ahmed Malik and Hans-Peter Piepho. On generalized exponential transformations for proportions. Communications in Statistics: Theory and Methods, 45(19):5857–5870, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.950753.

Mandal:2017:OMD

[MP17] Nripes Kumar Mandal and Manisha Pal. Optimum mixture designs in some constrained experimental regions. Communications in Statistics: Theory and Methods, 46(9):4240–4249, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1080844.

Matsushita:2019:TD

[MPD^{+}19] Raul Matsushita, Donald Pianto, Bernardo B. De Andrade, Andre Cançado, and Sergio Da Silva. The Touchard distribution. Communications in Statistics: Theory and Methods, 48(8): 2049–2059, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1444177.

Mandal:2016:CCB

[MPG16] B. N. Mandal, Rajender Parsad, and V. K. Gupta. Cyclic circular balanced and strongly balanced crossover designs through integer programming. Communications in Statistics: Theory and Methods, 45(4):859–871, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
853787.

Maleki:2015:SIW

[MR15] Y. Maleki and S. Rezakhah. The scale invariant Wigner spectrum estimation of Gaussian locally self-similar processes. Communications in Statistics: Theory and Methods, 44(23): 4983–5004, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Modarresi:2016:CDS

[MR16a] N. Modarresi and S. Rezakhah. Characterization of discrete scale invariant Markov sequences. Communications in Statistics: Theory and Methods, 45(18):5263–5278, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.942427.

Moreno:2016:RUR

[MR16b] Marta Moreno and Juan Romo. Robust unit root tests with autoregressive errors. Communications in Statistics: Theory and Methods, 45(20):5997–6021, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
955114.

Mubasher:2017:SRL

[MR17] Mohamed Mubasher and Howard E. Rockette. Stopping rules for long-term clinical trials based on two consecutive rejections of the null hypothesis. Communications in Statistics: Theory and Methods, 46(3):1342–1351, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019142.

Morabbi:2016:PEL

[MRA16] H. Morabbi, M. Razmkhah, and Jafar Ahmadi. Point estimation of lower quantiles based on two-sampling scheme. Communications in Statistics: Theory and Methods, 45(13):4018–4034, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.915044.

Meerschaert:2012:PEE

[MRS12] Mark M. Meerschaert, Parthanil Roy, and Qin Shao. Parameter estimation for exponentially tempered power law distributions.

Communications in Statistics: Theory and Methods, 41(10): 1839–1856, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maiboroda:2010:GEE

[MS10a] Rostyslav Maiboroda and Olena Sugakova. Generalized estimating equations for symmetric distributions observed with admixture. Communications in Statistics: Theory and Methods, 40(1):96–116, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matsuura:2010:OBS

[MS10b] Shun Matsuura and Nobuo Shinozaki. Optimal binning strategies under squared error loss in selective assembly with a tolerance constraint. Communications in Statistics: Theory and Methods, 39(4):592–605, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mohammadpour:2010:FMA

[MS10c] M. Mohammadpour and A. R. Soltani. Forward moving average representation in multivariate MA(1) processes. Communications in Statistics: Theory and Methods, 39(4):729–737, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moschopoulos:2010:DFS

[MS10d] Panagis Moschopoulos and Max Shpak. The distribution of family sizes under a time-homogeneous birth and death process. Communications in Statistics: Theory and Methods, 39(10): 1761–1775, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See corrigendum [Ano13b].

Mukhopadhyay:2010:BAS

[MS10e] Chiranjit Mukhopadhyay and Mathews P. Samuel. Bayesian analysis of a superimposed renewal process. Communications in Statistics: Theory and Methods, 40(2):279–303, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maansson:2011:RPL

[MS11a] Kristofer Månsson and Ghazi Shukur. On ridge parameters in logistic regression. Communications in Statistics: Theory and Methods, 40(18):3366–3381, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mishura:2011:EUS

[MS11b] Yulia S. Mishura and Georgiy M. Shevchenko. Existence and uniqueness of the solution of stochastic differential equation involving Wiener process and fractional Brownian motion with Hurst index H > 1/2. Communications in Statistics: Theory and Methods, 40(19–20):3492–3508, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:2011:CBS

[MS11c] Nitis Mukhopadhyay and Mun S. Son. On the covariance between the sample mean and variance. Communications in Statistics: Theory and Methods, 40(7):1142–1148, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Melas:2012:FAB

[MS12a] Viatcheslav B. Melas and Petr Shpilev. On the functional approach to Bayesian efficient designs for nonlinear regression models. Communications in Statistics: Theory and Methods, 41(16–17):2908–2921, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
615972.

Monari:2012:FMA

[MS12b] P. Monari and L. Stracqualursi. A forecast model to assess the critical points in university systems. Communications in Statistics: Theory and Methods, 41(16–17):2896–2907, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/03610926.2011.615439.

Mai:2013:SEH

[MS13] Jan-Frederik Mai and Matthias Scherer. Sampling exchangeable and hierarchical Marshall–Olkin distributions. Communications in Statistics: Theory and Methods, 42(4):619–632, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maciag:2014:LSR

[MS14a] Krystyna Maciag and Czesław Stepniak. Lack of synergy regression models. Communications in Statistics: Theory and Methods, 43(2):388–391, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McAssey:2014:UON

[MS14b] Michael P. McAssey and Francisco J. Samaniego. On uniformly optimal networks: a reversal of fortune? Communications in Statistics: Theory and Methods, 43(10–12):2452–2467, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mohammadpour:2014:FMA

[MS14c] M. Mohammadpour and A. R. Soltani. Forward moving average representations for MA processes of finite order: Multivariate stationary and periodically correlated. Communications in Statistics: Theory and Methods, 43(1):141–150, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Masood:2015:SFE

[MS15] Saadia Masood and Javid Shabbir. On some families of estimators of variance of post-stratified sample mean using two auxiliary variables. Communications in Statistics: Theory and Methods, 44(11):2398–2415, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maeng:2017:BFI

[MS17a] Hyeyoung Maeng and Dong Wan Shin. Bootstrap forecast intervals for asymmetric volatilities via EGARCH model. Communications in Statistics: Theory and Methods, 46(3):1144–1157, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1014105.

Mathur:2017:NTT

[MS17b] S. K. Mathur and D. M. Sakate. A new test for two-sample location problem based on empirical distribution function. Communications in Statistics: Theory and Methods, 46(24):12345– 12355, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1295158.

Maji:2019:EPM

[MSB19a] Reba Maji, G. N. Singh, and Arnab Bandyopadhyay. Estimation of population mean in presence of random non response in two-stage cluster sampling. Communications in Statistics: Theory and Methods, 48(14):3586–3608, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1478101.

Misuraca:2019:BID

[MSB19b] Michelangelo Misuraca, Maria Spano, and Simona Balbi. BMS: An improved Dunn index for document clustering validation. Communications in Statistics: Theory and Methods, 48(20): 5036–5049, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1504968.

McClean:2019:SIC

[MSD19] Sally McClean, Christos H. Skiadas, and Guglielmo D’Amico. Special issue — communications in statistics — theory and methods 4th stochastic modeling techniques and data analysis international conference. Communications in Statistics: Theory and Methods, 48(1):1–2, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2019.
1587831.

Moslemi:2018:RPP

[MSEN18] Amir Moslemi, Mirmehdi Seyyed-Esfahani, and Seyed Taghi Akhavan Niaki. A robust posterior preference multi-response optimization approach in multistage processes. Communications in Statistics: Theory and Methods, 47(15):3547–3570, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1359301.

Muneer:2017:EFP

[MSK17] Siraj Muneer, Javid Shabbir, and Alamgir Khalil. Estimation of finite population mean in simple random sampling and stratified random sampling using two auxiliary variables. Communications in Statistics: Theory and Methods, 46(5):2181–2192, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1035394.

Maansson:2018:PSR

[MSK18] Kristofer Månsson, Ghazi Shukur, and B. M. Golam Kibria. Performance of some ridge regression estimators for the multinomial logit model. Communications in Statistics: Theory and Methods, 47(12):2795–2804, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.784996.

Mei:2017:EIP

[MSL17] Jingning Mei, Q. Shao, and R. Liu. Efficient inference for parameters of unobservable periodic autoregressive time series. Communications in Statistics: Theory and Methods, 46(15): 7390–7408, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1152484.

Midhu:2014:CDL

[MSN14] N. N. Midhu, P. G. Sankaran, and N. Unnikrishnan Nair. A class of distributions with linear hazard quantile function. Communications in Statistics: Theory and Methods, 43(17):3674– 3689, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monteiro:2012:IVS

[MSP12] Magda Monteiro, Manuel G. Scotto, and Isabel Pereira. Integer-valued self-exciting threshold autoregressive processes. Communications in Statistics: Theory and Methods, 41(15): 2717–2737, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mirjalili:2017:NMR

[MSR17] Ahmad Mirjalili, Mohammad Khanjari Sadegh, and Majid Rezaei. A note on the mean residual life of a coherent system with a cold standby component. Communications in Statistics: Theory and Methods, 46(20):10348–10358, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1235194.

Mishura:2011:P

[MSS11] Yuliya Mishura, Lyudmyla Sakhno, and Georgiy Shevchenko. Preface. Communications in Statistics: Theory and Methods, 40(19–20):3383–3384, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maansson:2014:NAI

[MSS14a] Kristofer Månsson, Ghazi Shukur, and Pär Sjölander. A new asymmetric interaction ridge (AIR) regression method. Communications in Statistics: Theory and Methods, 43(3):616–643, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maansson:2014:NRR

[MSS14b] Kristofer Månsson, Ghazi Shukur, and Pär Sjölander. A new ridge regression causality test in the presence of multicollinearity. Communications in Statistics: Theory and Methods, 43(2): 235–248, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Makkar:2014:BSA

[MSSU14] Puja Makkar, Puneet K. Srivastava, R. S. Singh, and S. K. Upadhyay. Bayesian survival analysis of head and neck cancer data using lognormal model. Communications in Statistics: Theory and Methods, 43(2):392–407, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehreyan:2019:NNM

[MST19] S. Zamani Mehreyan, A. Sayyareh, and D. Thomakos. Nonnested model selection based on the quantiles and it’s application in time series. Communications in Statistics: Theory and Methods, 48(2):332–353, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1410714.

Muller:2013:NSB

[MSW13] U. U. Müller, A. Schick, and W. Wefelmeyer. Non standard behavior of density estimators for functions of independent observations. Communications in Statistics: Theory and Methods, 42(16):2991–3000, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielniczuk:2012:SRA

[MT12] Jan Mielniczuk and Paweł Teisseyre. Selection of regression and autoregression models with initial ordering of variables.

Communications in Statistics: Theory and Methods, 41(24): 4484–4502, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madadi:2014:SIR

[MT14] M. Madadi and M. Tata. Shannon information in k-records. Communications in Statistics: Theory and Methods, 43(15): 3286–3301, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mohammadi:2016:EBE

[MT16] Zohreh Mohammadi and Mina Towhidi. Empirical Bayes estimation for the mean of the selected normal population when there are additional data. Communications in Statistics: Theory and Methods, 45(12):3675–3691, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.904359.

Moslehian:2019:NCS

[MT19] Mohammad Sal Moslehian and Ali Talebi. Non-commutative Stein inequality and its applications. Communications in Statistics: Theory and Methods, 48(7):1611–1620, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1435813.

Mansoor:2019:MOL

[MTC^{+}19] M. Mansoor, M. H. Tahir, Gauss M. Cordeiro, Serge B. Provost, and Ayman Alzaatreh. The Marshall-Olkin logisticexponential distribution. Communications in Statistics: Theory and Methods, 48(2):220–234, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1414254.

Maaita:2013:HMS

[MTM13] J.-O. Maaita, G. Tsaklidis, and E. Meletlidou. The homogeneous Markov system (HMS) as an elastic medium. The threedimensional case. Communications in Statistics: Theory and Methods, 42(16):2959–2970, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudakkar:2016:CPD

[Mud16] Syeda Rabab Mudakkar. Construction of peculiar diffusion process having Gaussian marginals. Communications in Statistics: Theory and Methods, 45(15):4509–4516, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.922984.

Mukherjee:2010:SSO

[Muk10] Amitava Mukherjee. Semi-sequential one-shot monitoring of small disorders with controlled Type I error rate. Communications in Statistics: Theory and Methods, 39(15):2829–2847, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:2011:SJI

[Muk11] Nitis Mukhopadhyay. On sharp Jensen’s inequality and some unusual applications. Communications in Statistics: Theory and Methods, 40(7):1283–1297, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muller:2012:RCS

[Mül12] Werner G. Müller. Remarks on Communications in Statistics — Theory and Methods, Vol. 39, No. 15 (2010), pp. 2818– 2828. Communications in Statistics: Theory and Methods, 41(6):1126–1127, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [AW10].

Murat:2013:BED

[Mur13] Małgorzata Murat. Bayesian estimation for deformed modified power series distribution. Communications in Statistics: Theory and Methods, 42(2):365–384, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murat:2016:BEN

[Mur16a] Małgorzata Murat. Bayesian estimation for non zero inflated modified power series distribution under linex and generalized entropy loss functions. Communications in Statistics: Theory and Methods, 45(13):3952–3969, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.912057.

Murteira:2016:GLT

[Mur16b] José M. R. Murteira. Goodness-of-link tests for multivariate regression models. Communications in Statistics: Theory and Methods, 45(24):7367–7375, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
980518.

Marella:2013:OOB

[MV13] Daniela Marella and Paola Vicard. Object-oriented Bayesian networks for modeling the respondent measurement error. Communications in Statistics: Theory and Methods, 42(19): 3463–3477, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mardia:2014:SFP

[MV14] Kanti V. Mardia and Jochen Voss. Some fundamental properties of a multivariate von Mises distribution. Communications in Statistics: Theory and Methods, 43(6):1132–1144, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maruotti:2016:TVC

[MV16a] Antonello Maruotti and Maurizio Vichi. Time-varying clustering of multivariate longitudinal observations. Communications in Statistics: Theory and Methods, 45(2):430–443, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meena:2016:RDM

[MV16b] K. S. Meena and T. Vasanthi. Reliability design for a MANET with cluster-head gateway routing protocol. Communications in Statistics: Theory and Methods, 45(13):3904–3918, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.911908.

Muller:2010:ENR

[MW10] Ursula U. Müller and Wolfgang Wefelmeyer. Estimation in nonparametric regression with non-regular errors. Communications in Statistics: Theory and Methods, 39(8–9):1619–1629, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miao:2013:SLR

[MWM13] Yu Miao, Yanling Wang, and Xinwen Ma. Some limit results for Pareto random variables. Communications in Statistics: Theory and Methods, 42(24):4384–4391, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meng:2014:MCC

[MWW14] Xianhua Meng, Jinglong Wang, and Xianyi Wu. Multiple comparisons controlling expected number of false discoveries. Communications in Statistics: Theory and Methods, 43(13):2830– 2843, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meng:2018:CCC

[MWW18] Bing Meng, Dingcheng Wang, and Qunying Wu. Complete convergence and complete moment convergence for arrays of rowwise negatively superadditive dependent random variables. Communications in Statistics: Theory and Methods, 47(16): 3910–3922, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1364391.

Ma:2018:ORI

[MWY18] Jianjing Ma, Guojing Wang, and George Xianzhi Yuan. Optimal reinsurance and investment problem in a defaultable market. Communications in Statistics: Theory and Methods, 47(7): 1597–1614, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1321772.

Mu:2015:RSR

[MX15] Weiyan Mu and Shifeng Xiong. Robust sparse regression with high-breakdown value. Communications in Statistics: Theory and Methods, 44(5):1033–1043, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miao:2019:NAS

[MX19] Yu Miao and Xiaoyan Xu. A note on the almost sure central limit theorem for the product of partial sums of m-dependent random variables. Communications in Statistics: Theory and Methods, 48(9):2102–2112, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1459707.

Miao:2013:DBE

[MXD13] Yu Miao, Tianyu Xue, and Tian Du. The discounted Berry– Esséen analogue for autoregressive processes. Communications in Statistics: Theory and Methods, 42(15):2684–2693, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miao:2011:MNE

[MYM11] Yu Miao, Guangyu Yang, and Jianyong Mu. Minimum L_{p} norm estimator for simple linear regressive model. Communications in Statistics: Theory and Methods, 40(4):571–580, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mahmoud:2010:MAE

[MZ10a] Mahmoud A. Mahmoud and Alyaa R. Zahran. A multivariate adaptive exponentially weighted moving average control chart. Communications in Statistics: Theory and Methods, 39(4):606– 625, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mahmoudi:2010:GPL

[MZ10b] E. Mahmoudi and H. Zakerzadeh. Generalized Poisson–Lindley distribution. Communications in Statistics: Theory and Methods, 39(10):1785–1798, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mao:2011:JMC

[MZ11] Jie Mao and Zhongyi Zhu. Joint mean-covariance models with applications to longitudinal data in partially linear model. Communications in Statistics: Theory and Methods, 40(17): 3119–3140, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mytalas:2014:CLT

[MZ14] George C. Mytalas and Michael A. Zazanis. Central limit theorem approximations for the number of runs in Markovdependent multi-type sequences. Communications in Statistics: Theory and Methods, 43(7):1340–1350, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mladenovic:2015:UPM

[MZ15] Pavle Mladenović and Lenka Zivadinović. Uniform AR(1) processes and maxima on partial samples. Communications in Statistics: Theory and Methods, 44(12):2546–2563, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ma:2017:PLC

[MZ17] Huijuan Ma and Yong Zhou. Pseudo-likelihood for casecohort studies under length-biased sampling. Communications in Statistics: Theory and Methods, 46(1):28–48, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.983613.

Mahdizadeh:2018:SPR

[MZ18] M. Mahdizadeh and Ehsan Zamanzade. Stratified pair ranked set sampling. Communications in Statistics: Theory and Methods, 47(24):5904–5915, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1402053.

Nahla:2012:PLL

[NA12] Ben Salah Nahla and Masmoudi Afif. Poisson limit laws for exponential dispersion models. Communications in Statistics: Theory and Methods, 41(3):393–402, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noughabi:2013:TNU

[NA13] Hadi Alizadeh Noughabi and Naser Reza Arghami. Testing normality using transformed data. Communications in Statistics: Theory and Methods, 42(17):3065–3075, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nayabuddin:2017:RRS

[NA17] Nayabuddin and Haseeb Athar. Recurrence relations for single and product moments of generalized order statistics from Marshall–Olkin extended Pareto distributions. Communications in Statistics: Theory and Methods, 46(16):7820–7826, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.960579.

Noue:2018:DDG

[NA18] Kiyoshi I. Noue and Sigeo Aki. On discrete distributions generated from drawing balls with bivariate labels. Communications in Statistics: Theory and Methods, 47(14):3532–3546, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1359300.

Nofal:2019:NEF

[NA19] Zohdy M. Nofal and M. Ahsanullah. A new extension of the fréchet distribution: Properties and its characterization. Communications in Statistics: Theory and Methods, 48(9):2267– 2285, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1465080.

Nekoukhou:2013:FSG

[NAA13] V. Nekoukhou, M. H. Alamatsaz, and A. H. Aghajani. A flexible skew-generalized normal distribution. Communications in Statistics: Theory and Methods, 42(13):2324–2334, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nascimento:2016:IOC

[NAAC16] A. D. C. Nascimento, G. J. A. Amaral, B. B. Achic, and J. T. M. Cruz. Influential observation in complex normal data for problems in allometry. Communications in Statistics: Theory and Methods, 45(9):2714–2729, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.889165.

Nekoukhou:2012:DAG

[NAB12] V. Nekoukhou, M. H. Alamatsaz, and H. Bidram. A discrete analog of the generalized exponential distribution. Communications in Statistics: Theory and Methods, 41(11):2000–2013, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nekoukhou:2016:NEG

[NAB16] V. Nekoukhou, M. H. Alamatsaz, and H. Bidram. A note on exponentiated geometric distribution: Another generalization of geometric distribution. Communications in Statistics: Theory and Methods, 45(5):1575, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.864769.

Nekoukhou:2015:DBE

[NABA15] V. Nekoukhou, M. H. Alamatsaz, H. Bidram, and A. H. Aghajani. Discrete beta-exponential distribution. Communications in Statistics: Theory and Methods, 44(10):2079–2091, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadarajah:2019:DMM

[NAC19] Saralees Nadarajah, Emmanuel Afuecheta, and Stephen Chan. On the distribution of maximum of multivariate normal random vectors. Communications in Statistics: Theory and Methods, 48(10):2425–2445, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1465088.

Nadar:2010:LCR

[Nad10] Mustafa Nadar. Local convergence rate of mean squared error in density estimation. Communications in Statistics: Theory and Methods, 40(1):176–185, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noorossana:2015:CVS

[NAD15] Rassoul Noorossana, Maryam Shekary A., and Ali Deheshvar. Combined variable sample size, sampling interval, and double sampling (CVSSIDS) adaptive control charts. Communications in Statistics: Theory and Methods, 44(6):1255–1269, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagaraja:2017:TMP

[Nag17] Chaitra H. Nagaraja. Theory and methods for partitioned Gini coefficients computed on post-stratified data. Communications in Statistics: Theory and Methods, 46(10):4809–4823, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1088031.

Nagakura:2018:RBM

[Nag18] Daisuke Nagakura. On the relationship between the matrix operators, vech and vecd. Communications in Statistics: Theory and Methods, 47(13):3252–3268, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1353623.

Nakashima:2017:MGL

[Nak17] Eiji Nakashima. Modification of GEE1 and linear mixed-effects models for heteroscedastic longitudinal Gaussian data. Communications in Statistics: Theory and Methods, 46(22):11110– 11122, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1260737.

Nakashima:2019:IAC

[Nak19] Eiji Nakashima. Impact of additive covariate error on linear model. Communications in Statistics: Theory and Methods, 48(22):5517–5529, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1515361.

Nassar:2017:APW

[NAMAK17] M. Nassar, A. Alzaatreh, M. Mead, and O. Abo-Kasem. Alpha power Weibull distribution: Properties and applications. Communications in Statistics: Theory and Methods, 46(20):10236– 10252, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1231816. See comment

[Jon18].

Nassiri:2014:SNI

[NAMD14] Vahid Nassiri, Mina Aminghafari, and Ali Mohammad-Djafari. Solving noisy ICA using multivariate wavelet denoising with an application to noisy latent variables regression. Communications in Statistics: Theory and Methods, 43(10–12):2297–2310, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nkurunziza:2016:SLS

[NAML16] Sévérien Nkurunziza, Marwan Al-Momani, and Eric Yu Yin Lin. Shrinkage and LASSO strategies in high-dimensional heteroscedastic models. Communications in Statistics: Theory and Methods, 45(15):4454–4470, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
921305.

Nagar:2017:BVG

[NAN17a] D. K. Nagar, M. Arashi, and S. Nadarajah. Bimatrix variate gamma-beta distributions. Communications in Statistics: Theory and Methods, 46(9):4464–4483, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1085562.

Najafabadi:2017:CPR

[NAN17b] Amir T. Payandeh Najafabadi, Fatemeh Atatalab, and Maryam Omidi Najafabadi. Credibility premium for ratemaking systems. Communications in Statistics: Theory and Methods, 46(1):415–426, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
995823.

Narukawa:2013:STF

[Nar13] Masaki Narukawa. On semiparametric testing of I(d) by FEXP models. Communications in Statistics: Theory and Methods, 42(9):1637–1653, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nastic:2012:SGM

[Nas12] Aleksandar S. Nastić. On shifted geometric INAR(1) models based on geometric counting series. Communications in Statistics: Theory and Methods, 41(23):4285–4301, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nassar:2018:CES

[Nas18] Hiba Nassar. A consistent estimator of the smoothing operator in the functional hodrick-prescott filter. Communications in Statistics: Theory and Methods, 47(12):3029–3042, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1346806.

Navarro:2016:VSP

[Nav16] Jorge Navarro. A very simple proof of the multivariate Chebyshev’s inequality. Communications in Statistics: Theory and Methods, 45(12):3458–3463, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
873135.

Nofal:2017:GTF

[NAYC17] Zohdy M. Nofal, Ahmed Z. Afify, Haitham M. Yousof, and Gauss M. Cordeiro. The generalized transmuted-G family of distributions. Communications in Statistics: Theory and Methods, 46(8):4119–4136, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1078478.

Nadarajah:2015:NFM

[NB15] S. Nadarajah and S. A. A. Bakar. New folded models for the log-transformed Norwegian fire claim data. Communications in Statistics: Theory and Methods, 44(20):4408–4440, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagatsuka:2016:CME

[NB16] Hideki Nagatsuka and N. Balakrishnan. A consistent method of estimation for the three-parameter lognormal distribution based on Type-II right censored data. Communications in Statistics: Theory and Methods, 45(19):5693–5708, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.948205.

Nekoukhou:2017:NGW

[NB17] Vahid Nekoukhou and Hamid Bidram. A new generalization of the Weibull-geometric distribution with bathtub failure rate. Communications in Statistics: Theory and Methods, 46(9): 4296–4310, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1081949.

Nagatsuka:2014:CME

[NBK14] Hideki Nagatsuka, N. Balakrishnan, and Toshinari Kamakura. A consistent method of estimation for the three-parameter gamma distribution. Communications in Statistics: Theory and Methods, 43(18):3905–3926, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadaraya:2013:ANE

[NBS13] E. Nadaraya, P. Babilua, and G. Sokhadze. About the nonparametric estimation of the Bernoulli regression. Communications in Statistics: Theory and Methods, 42(22):3989–4002, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nduka:2014:OCC

[NC14] U. C. Nduka and P. E. Chigbu. On the optimal choice of cube and star replications in restricted second-order designs. Communications in Statistics: Theory and Methods, 43(19):4195– 4214, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadarajah:2018:DDB

[NC18a] Saralees Nadarajah and Stephen Chan. Discrete distributions based on inter arrival times with application to football data. Communications in Statistics: Theory and Methods, 47(1):147– 165, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1300284.

Nadarajah:2018:EEM

[NC18b] Saralees Nadarajah and Stephen Chan. Elementary expressions for moments of truncated negative binomial random variables. Communications in Statistics: Theory and Methods, 47(15): 3734–3743, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1361991.

Noughabi:2018:TVE

[NC18c] H. Alizadeh Noughabi and M. Chahkandi. Testing the validity of the exponential model for hybrid Type-I censored data.

Communications in Statistics: Theory and Methods, 47(23): 5770–5778, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1402046.

Niu:2019:ERM

[NC19a] Xiaomeng Niu and Hyunkeun Ryan Cho. Efficient regression modeling for correlated and overdispersed count data. Communications in Statistics: Theory and Methods, 48(24):6005–6018, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1524486.

Niu:2019:SEI

[NC19b] Xiaomeng Niu and Hyunkeun Ryan Cho. Simultaneous estimation and inference for multiple response variables. Communications in Statistics: Theory and Methods, 48(11):2734–2747, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1472791.

Nadarajah:2016:DSO

[NCA16] Saralees Nadarajah, Stephen Chan, and Emmanuel Afuecheta. On the distribution of some ordered variables. Communications in Statistics: Theory and Methods, 45(7):2137–2142, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.875578.

Nadarajah:2017:TVR

[NCA17] Saralees Nadarajah, Stephen Chan, and Emmanuel Afuecheta. Tabulations for value at risk and expected shortfall. Communications in Statistics: Theory and Methods, 46(12):5956–5984, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1116572.

Nadarajah:2015:ZBF

[NCO15] Saralees Nadarajah, Gauss M. Cordeiro, and Edwin M. M. Ortega. The Zografos–Balakrishnan–G family of distributions: Mathematical properties and applications. Communications in Statistics: Theory and Methods, 44(1):186–215, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadar:2017:LCR

[NE17] Mustafa Nadar and Elif Erçelik. Local convergency rate of MSE in density estimation using the second-order modulus of smoothness. Communications in Statistics: Theory and Methods, 46(7):3164–3173, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1056367.

Negri:2009:ECU

[Neg09] Ilia Negri. Efficiency of a class of unbiased estimators for the invariant distribution function of a diffusion process. Communications in Statistics: Theory and Methods, 39(1):177–185, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nenes:2018:NMS

[Nen18] George Nenes. New measures of statistical performance of process control charts. Communications in Statistics: Theory and Methods, 47(18):4590–4608, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1377255.

Newcombe:2011:MLC

[New11a] Robert G. Newcombe. Measures of location for confidence intervals for proportions. Communications in Statistics: Theory and Methods, 40(10):1743–1767, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Newcombe:2011:PIC

[New11b] Robert G. Newcombe. Propagating imprecision: Combining confidence intervals from independent sources. Communications in Statistics: Theory and Methods, 40(17):3154–3180, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagode:2011:RAU

[NF11a] Marko Nagode and Matija Fajdiga. The REBMIX algorithm and the univariate finite mixture estimation. Communications in Statistics: Theory and Methods, 40(5):876–892, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagode:2011:RAM

[NF11b] Marko Nagode and Matija Fajdiga. The REBMIX algorithm for the multivariate finite mixture estimation. Communications in Statistics: Theory and Methods, 40(11):2022–2034, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ning:2011:UDE

[NFZ11] Jian-Hui Ning, Kai-Tai Fang, and Yong-Dao Zhou. Uniform design for experiments with mixtures. Communications in Statistics: Theory and Methods, 40(10):1734–1742, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:2012:BIP

[Ng12] Vee Ming Ng. Bayesian inference for the precision matrix for scale mixtures of normal distributions. Communications in Statistics: Theory and Methods, 41(24):4407–4412, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ngerng:2011:RNE

[Nge11] M. H. Ngerng. Recursive nonparametric estimation of local first derivative under dependence conditions. Communications in Statistics: Theory and Methods, 40(7):1159–1168, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Niaki:2013:EDV

[NGT13] Seyed Taghi Akhavan Niaki, Fazlollah Masoumi Gazaneh, and Moslem Toosheghanian. Economic design of variable sampling interval X-bar control charts for monitoring correlated non normal samples. Communications in Statistics: Theory and Methods, 42(18):3339–3358, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagaraja:2015:FIC

[NH15a] H. N. Nagaraja and Qinying He. Fisher information in censored samples from the Marshall–Olkin bivariate exponential distribution. Communications in Statistics: Theory and Methods, 44(19):4172–4184, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nkurunziza:2015:SEM

[NH15b] Sévérien Nkurunziza and Abdulkadir Hussein. Shrinkage estimation of the memory parameter in stationary Gaussian processes. Communications in Statistics: Theory and Methods, 44(8):1580–1591, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanda:2017:SGA

[NHAMG17] Asok K. Nanda, Nil Kamal Hazra, D. K. Al-Mutairi, and M. E. Ghitany. On some generalized ageing orderings. Communications in Statistics: Theory and Methods, 46(11):5273–5291, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1100738.

Nakagawa:2016:RFJ

[NHN16] Shigekazu Nakagawa, Hiroki Hashiguchi, and Naoto Niki. Recurrence formula on the joint distribution of sample skewness and kurtosis under normality. Communications in Statistics: Theory and Methods, 45(4):1148–1155, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.859704.

Nam:2017:MIS

[NHV17] Trinh Hoai Nam, Tien-Chung Hu, and Andrei Volodin. Maximal inequalities and strong law of large numbers for sequences of m-asymptotically almost negatively associated random variables. Communications in Statistics: Theory and Methods, 46(6):2696–2707, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1048885.

Noughabi:2016:QPL

[NI16] M. Shafaei Noughabi and S. Izadkhah. On the quantile past lifetime of the components of the parallel systems. Communications in Statistics: Theory and Methods, 45(7):2130–2136, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.875573.

Nievergelt:2017:LSL

[Nie17] Yves Nievergelt. Least-squares logistic curves with initial conditions exist. Communications in Statistics: Theory and Methods, 46(8):4016–4030, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1076480.

Nurunnabi:2011:DMI

[NIN11] A. A. M. Nurunnabi, A. H. M. Rahmatullah Imon, and M. Nasser. A diagnostic measure for influential observations in linear regression. Communications in Statistics: Theory and Methods, 40(7):1169–1183, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nduka:2017:FPT

[NIN17] U. C. Nduka, S. I. Iwueze, and E. C. Nwogu. Fitting polynomial trend to time series by the method of Buys-Ballot estimators. Communications in Statistics: Theory and Methods, 46(9): 4520–4538, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1085569.

Nishida:2015:VSM

[NK15] Kiheiji Nishida and Yuichiro Kanazawa. On variance-stabilizing multivariate non parametric regression estimation. Communications in Statistics: Theory and Methods, 44(10):2151–2175, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanda:2019:MIT

[NK19] Phalguni Nanda and Suchandan Kayal. Mean inactivity time of lower record values. Communications in Statistics: Theory and Methods, 48(20):5145–5164, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1508714.

Najarzadeh:2016:PEL

[NKG16] D. Najarzadeh, M. Khazaei, and M. Ganjali. Performance evaluation of likelihood-ratio tests for assessing similarity of the covariance matrices of two multivariate normal populations.

Communications in Statistics: Theory and Methods, 45(5): 1435–1452, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2013.863933.

Niaki:2015:SMA

[NKS15] Seyed Taghi Akhavan Niaki, Majid Khedmati, and Mir Emad Soleymanian. Statistical monitoring of autocorrelated simple linear profiles based on principal components analysis. Communications in Statistics: Theory and Methods, 44(21):4454–4475, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadarajah:2017:EPM

[NLRD17] Saralees Nadarajah, Rui Li, Rasool Roozegar, and Ali Dolati. On estimation of population mean based on two measurements. Communications in Statistics: Theory and Methods, 46(21):10758–10767, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1242744.

Nguyen:2017:PCR

[NM17] Hien D. Nguyen and Geoffrey J. McLachlan. Progress on a conjecture regarding the triangular distribution. Communications in Statistics: Theory and Methods, 46(22):11261–11271, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1263742.

Nguyen:2018:STR

[NM18] Hien D. Nguyen and Geoffrey J. McLachlan. Some theoretical results regarding the polygonal distribution. Communications in Statistics: Theory and Methods, 47(20):5083–5095, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1386312.

Nguyen:2019:ACD

[NM19] Hien D. Nguyen and Geoffrey McLachlan. On approximations via convolution-defined mixture models. Communications in Statistics: Theory and Methods, 48(16):3945–3955, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1487069.

Naderi:2019:VKF

[NMAA19] H. Naderi, P. Matuła, M. Amini, and H. Ahmadzade. A version of the Kolmogrov–Feller weak law of large numbers for maximal weighted sums of random variables. Communications in Statistics: Theory and Methods, 48(21):5414–5418, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1513146.

Nourbakhsh:2015:SWD

[NMJY15] Mohammadreza Nourbakhsh, Yaser Mehrali, Ahad Jamalizadeh, and Gholamhoseein Yari. On a selection Weibull distribution. Communications in Statistics: Theory and Methods, 44(8):1640–1652, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Newcombe:2011:DSI

[NN11a] Robert G. Newcombe and Markku M. Nurminen. In defence of score intervals for proportions and their differences. Communications in Statistics: Theory and Methods, 40(7):1271–1282, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nezhad:2011:MST

[NN11b] Mohammad Saber Fallah Nezhad and Seyed Taghi Akhavan Niaki. A multi-stage two-machines replacement strategy using mixture models, Bayesian inference, and stochastic dynamic programming. Communications in Statistics: Theory and Methods, 40(4):702–725, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nezhad:2013:NAS

[NN13] Mohammad Saber Fallah Nezhad and Seyed Taghi Akhavan Niaki. A new acceptance sampling policy based on number of successive conforming items. Communications in Statistics: Theory and Methods, 42(8):1542–1552, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nenes:2015:NMR

[NN15] George Nenes and Yiannis Nikolaidis. A new model for the representation of the ISO 2859 standard. Communications in Statistics: Theory and Methods, 44(1):159–170, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelson:2016:SPG

[NN16a] David Nelson and Siamak Noorbaloochi. Splice plots for generalized linear models. Communications in Statistics: Theory and Methods, 45(12):3524–3540, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
895842.

Nooghabi:2016:EPD

[NN16b] M. Jabbari Nooghabi and E. Khaleghpanah Nooghabi. On entropy of a Pareto distribution in the presence of outliers. Communications in Statistics: Theory and Methods, 45(17): 5234–5250, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.941495.

Nezhad:2018:DER

[NN18] Mohammad Saber Fallah Nezhad and Tahereh Jafari Nodoushan. Designing an economic rectifying sampling plan in the presence of two markets. Communications in Statistics: Theory and Methods, 47(5):1256–1272, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1317807.

Niavarani:2012:TNM

[NNA12] Mohammad R. Niavarani, Rassoul Noorossana, and Babak Abbasi. Three new multivariate process capability indices. Communications in Statistics: Theory and Methods, 41(2):341–356, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noorossana:2015:SMN

[NNI15] R. Noorossana, S. T. A. Niaki, and H. Izadbakhsh. Statistical monitoring of nominal logistic profiles in phase II. Communications in Statistics: Theory and Methods, 44(13):2689–2704, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nooghabi:2010:DOG

[NNN10] M. Jabbari Nooghabi, H. Jabbari Nooghabi, and P. Nasiri. Detecting outliers in gamma distribution. Communications in Statistics: Theory and Methods, 39(4):698–706, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See comments [LF17].

Ngaruye:2017:SAE

[NNvRS17] Innocent Ngaruye, Joseph Nzabanita, Dietrich von Rosen, and Martin Singull. Small area estimation under a multivariate linear model for repeated measures data. Communications in Statistics: Theory and Methods, 46(21):10835–10850, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1248784.

Naidoo:2016:MSB

[NNZA16] L. R. Naidoo, D. North, T. Zewotir, and R. Arnab. Multiplestart balanced modified systematic sampling in the presence of linear trend. Communications in Statistics: Theory and Methods, 45(14):4307–4324, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1091080.

Naidoo:2018:RMS

[NNZA18] L. R. Naidoo, D. North, T. Zewotir, and R. Arnab. Remainder modified systematic sampling in the presence of linear trend. Communications in Statistics: Theory and Methods, 47(10): 2469–2481, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1295076.

Namba:2012:MPH

[NO12] Akio Namba and Kazuhiro Ohtani. MSE performance of a heterogeneous pre-test ridge regression estimator. Communications in Statistics: Theory and Methods, 41(9):1692–1700, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Najera:2017:FG

[NO17] Edilberto Nájera and Federico O’Reilly. On fiducial generators. Communications in Statistics: Theory and Methods, 46(5): 2232–2248, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1040505.

Namba:2018:MPW

[NO18] Akio Namba and Kazuhiro Ohtani. MSE performance of the weighted average estimators consisting of shrinkage estimators. Communications in Statistics: Theory and Methods, 47(5): 1204–1214, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1316860.

Nooghabi:2016:SEE

[Noo16] Mehdi Jabbari Nooghabi. Shrinkage estimation of P (Y < X) in the exponential distribution mixing with exponential distribution. Communications in Statistics: Theory and Methods, 45(5):1477–1486, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
864768.

Normark:2016:MDF

[Nor16] Sofia Normark. Minimax designs for 2^{k} factorial experiments for generalized linear models. Communications in Statistics: Theory and Methods, 45(16):4788–4797, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.927502.

Najera:2019:FDP

[NORV19] Edilberto Nájera, Federico O’Reilly, and Silvia Ruiz-Velasco. Fiducial distribution in a power series family. Communications in Statistics: Theory and Methods, 48(23):5796–5808, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1522347.

Nosek:2014:CPD

[Nos14] K. Nosek. Change-point detection in two-phase regression with inequality constraints on the regression parameters. Communications in Statistics: Theory and Methods, 43(5):932–946, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noughabi:2017:GIB

[Nou17] Hadi Alizadeh Noughabi. Gini index based goodness-of-fit test for the logistic distribution. Communications in Statistics: Theory and Methods, 46(14):7114–7124, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1144767.

Nogales:2015:UEG

[NP15] A. G. Nogales and P. Pérez. Unbiased estimation for the general half-normal distribution. Communications in Statistics: Theory and Methods, 44(17):3658–3667, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Navarro:2017:CBG

[NP17a] Jorge Navarro and Georgios Psarrakos. Characterizations based on generalized cumulative residual entropy functions. Communications in Statistics: Theory and Methods, 46(3):1247–1260, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1014111.

Nematollahi:2017:ELP

[NP17b] N. Nematollahi and A. Pagheh. Estimation of the location parameter and the average worth of the selected subset of two parameter exponential populations under LINEX loss function. Communications in Statistics: Theory and Methods, 46(8): 3901–3914, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1076472.

Nyman:2017:SGG

[NPC17] Henrik Nyman, Johan Pensar, and Jukka Corander. Stratified Gaussian graphical models. Communications in Statistics: Theory and Methods, 46(11):5556–5578, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1105979.

Neammanee:2012:IUB

[NR12] K. Neammanee and N. Rerkruthairat. An improvement of a uniform bound on a combinatorial central limit theorem. Communications in Statistics: Theory and Methods, 41(9):1590– 1602, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Noughabi:2014:PRL

[NRB14] M. Shafaei Noughabi, A. H. Rezaei Roknabadi, and G. R. Mohtashami Borzadaran. On the percentile residual lifetime of parallel systems. Communications in Statistics: Theory and Methods, 43(18):3836–3847, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagar:2016:MVM

[NRCG16] Daya K. Nagar, Alejandro Roldán-Correa, and Arjun K. Gupta. Matrix variate Macdonald distribution. Communications in Statistics: Theory and Methods, 45(5):1311–1328, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.861494.

Nastic:2017:GTS

[NRI17] Aleksandar S. Nastić, Miroslav M. Ristić, and Ana V. Miletić Ilić. A geometric time-series model with an alternative dependent Bernoulli counting series. Communications in Statistics: Theory and Methods, 46(2):770–785, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1005100.

Nelson:2017:ECM

[NRN^{+}17] S. L. Prince Nelson, V. Ramakrishnan, P. J. Nietert, D. L. Kamen, P. S. Ramos, and B. J. Wolf. An evaluation of common methods for dichotomization of continuous variables to discriminate disease status. Communications in Statistics: Theory and Methods, 46(21):10823–10834, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1248783.

Nastic:2016:EBI

[NRP16] Aleksandar S. Nastić, Miroslav M. Ristić, and Predrag M. Popović. Estimation in a bivariate integer-valued autoregressive process. Communications in Statistics: Theory and Methods, 45(19):5660–5678, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
948203.

Nielsen:2014:GBC

[NRS^{+}14] J. D. Nielsen, J. S. Rosenthal, Y. Sun, D. M. Day, I. Bevc, and T. Duchesne. Group-based criminal trajectory analysis using cross-validation criteria. Communications in Statistics: Theory and Methods, 43(20):4337–4356, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nguimkeu:2015:IES

[NRW15] Pierre Nguimkeu, Marie Rekkas, and Augustine Wong. Interval estimation of the stress-strength reliability with independent normal random variables. Communications in Statistics: Theory and Methods, 44(6):1210–1221, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nair:2015:OFO

[NS15a] N. Unnikrishnan Nair and P. G. Sankaran. Odds function and odds rate for discrete lifetime distributions. Communications in Statistics: Theory and Methods, 44(19):4185–4202, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nayak:2015:ABL

[NS15b] Tapan K. Nayak and Bimal Sinha. An appreciation of balanced loss functions via regret loss. Communications in Statistics: Theory and Methods, 44(3):607–616, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nadaraya:2016:IFD

[NS16a] Elizbar Nadaraya and Grigol Sokhadze. On integral functionals of a density. Communications in Statistics: Theory and Methods, 45(23):7086–7102, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
974825.

Nair:2016:NZC

[NS16b] K. R. Muraleedharan Nair and N. Sreelakshmi. The New Zenga curve in the context of reliability analysis. Communications in Statistics: Theory and Methods, 45(22):6540–6552, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.763510.

Nezhad:2017:DOD

[NS17] Mohammad Saber Fallah Nezhad and Sina Seifi. Designing optimal double-sampling plan based on process capability index. Communications in Statistics: Theory and Methods, 46(13): 6624–6634, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1132325.

Nezhad:2018:NPD

[NS18] Mohammad Saber Fallah Nezhad and Faeze Zahmatkesh Saredorahi. A new policy for designing acceptance sampling plan based on Bayesian inference in the presence of inspection errors. Communications in Statistics: Theory and Methods, 47(21): 5307–5318, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1390130.

Nguyen:2019:OMS

[NS19] H. Nguyen and Q. Shao. Oracle model selection for correlated data via residuals. Communications in Statistics: Theory and Methods, 48(16):4067–4081, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1485946.

Navarro:2014:CBM

[NSL14] J. Navarro, S. M. Sunoj, and M. N. Linu. Characterizations of bivariate models using some dynamic conditional information divergence measures. Communications in Statistics: Theory and Methods, 43(9):1939–1948, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nair:2012:RAD

[NSP12] N. Unnikrishnan Nair, P. G. Sankaran, and M. Preeth. Reliability aspects of discrete equilibrium distributions. Communications in Statistics: Theory and Methods, 41(3):500–515, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nair:2017:SPD

[NSR17] N. Unnikrishnan Nair, P. G. Sankaran, and Nidhi P. Ramesh. Some properties of discrete bathtub-shaped distributions. Communications in Statistics: Theory and Methods, 46(11):5384– 5393, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1100743.

Nawaz:2019:PGN

[NSR^{+}19] Faisal Nawaz, Faisal Shahzad, Ijaz Ur Rehman, Muhammad Shujahat, Shabir Hyder, and Sameer Al Barghouthi. The performance of Gaussian and non Gaussian dynamic models in assessing market risk: The implications for risk management. Communications in Statistics: Theory and Methods, 48(13): 3360–3376, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1476707.

Nair:2018:PHM

[NSS18] N. Unnikrishnan Nair, P. G. Sankaran, and S. M. Sunoj. Proportional hazards model with quantile functions. Communications in Statistics: Theory and Methods, 47(19):4710–4723, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1445858.

Nidhi:2017:CAE

[NSSS17] Nidhi, B. V. S. Sisodia, Subedar Singh, and Sanjay K. Singh. Calibration approach estimation of the mean in stratified sampling and stratified double sampling. Communications in Statistics: Theory and Methods, 46(10):4932–4942, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1091083.

Nair:2012:GDS

[NSV12] N. Unnikrishnan Nair, P. G. Sankaran, and B. Vineshkumar. The Govindarajulu distribution: Some properties and applications. Communications in Statistics: Theory and Methods, 41(24):4391–4406, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Niglio:2015:TVA

[NV15] Marcella Niglio and Cosimo Damiano Vitale. Threshold vector ARMA models. Communications in Statistics: Theory and Methods, 44(14):2911–2923, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ngaruye:2019:MSE

[NVS19] Innocent Ngaruye, Dietrich Von Rosen, and Martin Singull. Mean-squared errors of small area estimators under a multivariate linear model for repeated measures data. Communications in Statistics: Theory and Methods, 48(8):2060–2073, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1444178.

Nguyen:2011:AES

[NY11] Dung Tien Nguyen and G. Yin. Asymptotic expansions for solutions of systems of Kolmogorov backward equations of two-time-scale switching jump diffusions. Communications in Statistics: Theory and Methods, 40(19–20):3425–3439, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ning:2016:TDS

[NY16] Jiannan Ning and Wende Yi. Tail dependence for skew Laplace distribution and skew Cauchy distribution. Communications in Statistics: Theory and Methods, 45(17):5224–5233, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.941494.

Nourbakhsh:2017:WRE

[NY17] Mohammadreza Nourbakhsh and Gholamhoseein Yari. Weighted Renyi’s entropy for lifetime distributions. Communications in Statistics: Theory and Methods, 46(14):7085–7098, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1148729.

Ozkale:2015:LIO

[ÖA15] M. Revan Özkale and Tugba Söküt Açar. Leverages and influential observations in a regression model with autocorrelated errors. Communications in Statistics: Theory and Methods, 44(11):2267–2290, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ortega:2016:CMM

[OAOQ16] Eva-Maria Ortega, José Alonso, and Encarnacion Ortega-Quiles. Comparisons of multistate models with discrete-time pure-birth process for recurrent events and uncertain parameters. Communications in Statistics: Theory and Methods, 45(6):1733–1746, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
960581.

Ortega:2016:GJR

[OB16] Francisco J. Ortega and Jesús Basulto. A generalization of Jeffreys’ rule for non regular models. Communications in Statistics: Theory and Methods, 45(15):4433–4444, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.921303.

Ong:2015:CDG

[OBPL15] S. H. Ong, Atanu Biswas, S. Peiris, and Y. C. Low. Count distribution for generalized Weibull duration with applications. Communications in Statistics: Theory and Methods, 44(19): 4203–4216, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ong:2012:GNC

[OCIS12] S. H. Ong, S. Chakraborty, T. Imoto, and K. Shimizu. Generalizations of non central negative binomial, Charlier series distributions, and their extensions by Lagrange expansion. Communications in Statistics: Theory and Methods, 41(4):571–590, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozel:2014:SRE

[ÖÇO14] Gamze Özel, Hülya Çingi, and Merve Oguz. Separate ratio estimators for the population variance in stratified random sampling. Communications in Statistics: Theory and Methods, 43 (22):4766–4779, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozmen:2010:BAZ

[ÖD10] Ilknur Özmen and Haydar Demirhan. A Bayesian approach for zero-inflated count regression models by using the reversible jump Markov chain Monte Carlo method and an application. Communications in Statistics: Theory and Methods, 39(12): 2109–2127, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oral:2019:MBR

[OD19] Evrim Oral and Denise M. Danos. Modeling binary responses with stochastic covariates. Communications in Statistics: Theory and Methods, 48(11):2619–2640, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472782.

Ormoz:2016:VSF

[OE16] Ehsan Ormoz and Farzad Eskandari. Variable selection in finite mixture of semi-parametric regression models. Communications in Statistics: Theory and Methods, 45(3):695–711, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ogura:2013:VSC

[OFS13] Toru Ogura, Yasunori Fujikoshi, and Takakazu Sugiyama. A variable selection criterion for two sets of principal component scores in principal canonical correlation analysis. Communications in Statistics: Theory and Methods, 42(12):2118–2135, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ogasawara:2011:AEM

[Oga11] Haruhiko Ogasawara. Asymptotic expansions in multi-group analysis of moment structures with an application to linearized estimators. Communications in Statistics: Theory and Methods, 40(9):1701–1716, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ogasawara:2015:BAM

[Oga15] Haruhiko Ogasawara. Bias adjustment minimizing the asymptotic mean square error. Communications in Statistics: Theory and Methods, 44(16):3503–3522, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ogasawara:2016:MPI

[Oga16] H. Ogasawara. Matching pseudocounts for interval estimation of binomial and Poisson parameters. Communications in Statistics: Theory and Methods, 45(17):5166–5178, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.941492.

Ogasawara:2019:ABI

[Oga19] Haruhiko Ogasawara. Asymptotic biases of information and cross-validation criteria under canonical parametrization. Communications in Statistics: Theory and Methods, 48(4):964–985, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1422759.

Otiniano:2017:MEV

[OGD17] C. E. G. Otiniano, C. R. Gonçalves, and C. C. Y. Dorea. Mixture of extreme-value distributions: identifiability and estimation. Communications in Statistics: Theory and Methods, 46(13):6528–6542, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1129423.

Ozonur:2016:GFT

[ÖGGB16] Deniz Özonur, Fikri Gökpinar, Esra Gökpinar, and Hülya Bayrak. Goodness of fit tests for Nakagami distribution based on smooth tests. Communications in Statistics: Theory and Methods, 45(7):1876–1886, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
867998.

Ogundimu:2016:UAM

[OH16] Emmanuel O. Ogundimu and Jane L. Hutton. A unified approach to multilevel sample selection models. Communications in Statistics: Theory and Methods, 45(9):2592–2611, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.887108.

Orihara:2019:DRE

[OH19] Shunichiro Orihara and Etsuo Hamada. Double robust estimator in general treatment regimes based on covariate-balancing. Communications in Statistics: Theory and Methods, 48(3):462– 478, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1414259.

Oskouei:2017:BSG

[OHKK17] Leili Hassani Oskouei, Parisa Hasanalipour, Hossein Jabbari Khamnei, and Kavoos Khorshidian. The beta skewgeneralized normal distribution. Communications in Statistics: Theory and Methods, 46(5):2270–2276, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1041980.

Omar:2015:CVR

[OJR15] M. Hafidz Omar, Anwar H. Joarder, and Muhammad Riaz. On a correlated variance ratio distribution and its industrial application. Communications in Statistics: Theory and Methods, 44 (2):261–274, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozbey:2015:EPP

[ÖK15] Fela Özbey and Selahattin Kaçiranlar. Evaluation of the predictive performance of the Liu estimator. Communications in Statistics: Theory and Methods, 44(10):1981–1993, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozbay:2017:PAO

[ÖK17] Nimet Özbay and Selahattin Kaçiranlar. The performance of the adaptive optimal estimator under the extended balanced loss function. Communications in Statistics: Theory and Methods, 46(22):11315–11326, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1267760.

Odaki:2019:JCR

[OL19a] Mitsuhiro Odaki and Min Li. Johansen cointegration rank tests under local alternative hypotheses when related drift or linear trend is not dependent upon sample size in VARs. Communications in Statistics: Theory and Methods, 48(23):5839–5849, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1522350.

Ortega:2019:ARI

[OL19b] Eva María Ortega and Xiaohu Li. Analytical results for IGFR and DRPFR distributions, with actuarial applications. Communications in Statistics: Theory and Methods, 48(2):178–202, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1082591.

Olive:2013:PGA

[Oli13] David J. Olive. Plots for generalized additive models. Communications in Statistics: Theory and Methods, 42(18):3310–3328, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ollis:2017:SCM

[Oll17] M. A. Ollis. Simple constructions for minimal neighbor and GN_{2} designs in linear blocks. Communications in Statistics: Theory and Methods, 46(4):1547–1550, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1021014.

Ozturk:2013:IBG

[OM13] Omer Ozturk and Steven N. MacEachern. Inference based on general linear models for order restricted randomization. Communications in Statistics: Theory and Methods, 42(18):3243– 3266, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Olivier:2017:RES

[OMB17] Jake Olivier, Warren L. May, and Melanie L. Bell. Relative effect sizes for measures of risk. Communications in Statistics: Theory and Methods, 46(14):6774–6781, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1134575.

Olmos:2017:BBS

[OMFB17] Neveka M. Olmos, Guillermo Martínez-Flórez, and Heleno Bolfarine. Bimodal Birnbaum–Saunders distribution with applications to non negative measurements. Communications in Statistics: Theory and Methods, 46(13):6240–6257, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1133824.

Ouadhi:2019:MCS

[OOT19] Kahina Ouadhi and Megdouda Ourbih-Tari. Monte Carlo simulation of ordinary least squares estimator through linear regression adaptive refined descriptive sampling algorithm. Communications in Statistics: Theory and Methods, 48(4):865–875, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1419265.

Olenko:2011:ASR

[OP11] Andriy Olenko and Tibor Pogány. Average sampling restoration of harmonizable processes. Communications in Statistics: Theory and Methods, 40(19–20):3587–3598, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oh:2017:ESP

[OPK17] Jin H. Oh, Sung H. Park, and Soon S. Kwon. Extended scaled prediction variance optimality for modified central composite design. Communications in Statistics: Theory and Methods, 46(19):9614–9624, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1213292.

Oh:2018:GER

[OPK18] Jin H. Oh, Sung H. Park, and Soon S. Kwon. Graphical evaluation of robust parameter designs based on extended scaled prediction variance and extended spherical average prediction variance. Communications in Statistics: Theory and Methods, 47(14):3523–3531, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1359299.

Olivares-Pacheco:2012:AEF

[OPSGB12] Juan F. Olivares-Pacheco, Estrella Salas, Héctor W. Gómez, and Heleno Bolfarine. An asymmetric extension for the family of elliptical slash distributions. Communications in Statistics: Theory and Methods, 41(6):1000–1012, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ou:2017:ACB

[OQ17] Zujun Ou and Hong Qin. Analytic connections between a double design and its original design in terms of different optimality criteria. Communications in Statistics: Theory and Methods, 46(15):7630–7641, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1158836.

Ou:2014:LBW

[OQC14] Zujun Ou, Hong Qin, and Xu Cai. A lower bound for the wraparound L_{2}-discrepancy on combined designs of mixed two- and three-level factorials. Communications in Statistics: Theory and Methods, 43(10–12):2274–2285, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ou:2017:SNL

[OQC17] Zujun Ou, Hong Qin, and Kashinath Chatterjee. Some new lower bounds to various discrepancies on combined designs. Communications in Statistics: Theory and Methods, 46(7): 3244–3254, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1060339.

Ou:2012:CAD

[OQL12] Zujun Ou, Hong Qin, and Hongyi Li. Connections among different criteria for optimal factor assignments. Communications in Statistics: Theory and Methods, 41(2):241–250, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Doric:2011:NGC

[– Dor11] Dragan – Dorić. New generalizations of Cauchy distribution. Communications in Statistics: Theory and Methods, 40(21): 3764–3776, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oliveira:2016:MHL

[OSX^{+}16] José Oliveira, Jéssica Santos, Cleber Xavier, Daniele Trindade, and Gauss M. Cordeiro. The McDonald half-logistic distribution: Theory and practice. Communications in Statistics: Theory and Methods, 45(7):2005–2022, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.873131. See correction [AE23c].

Ourbih-Tari:2017:SFE

[OTA17] Megdouda Ourbih-Tari and Mahdia Azzal. Survival function estimation with non parametric adaptive refined descriptive sampling algorithm: a case study. Communications in Statistics: Theory and Methods, 46(12):5840–5850, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1065328.

Otranto:2015:CSE

[Otr15] Edoardo Otranto. Capturing the spillover effect with multiplicative error models. Communications in Statistics: Theory and Methods, 44(15):3173–3191, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ogbonnaya:2016:EMD

[OU16] Chibueze E. Ogbonnaya and Emeka C. Uzochukwu. Estimation of missing data in analysis of covariance: A leastsquares approach. Communications in Statistics: Theory and Methods, 45(7):1902–1909, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
868000.

Omar:2019:NRM

[OW19] Kurdistan M. Taher Omar and Bo Wang. Nonparametric regression method with functional covariates and multivariate response. Communications in Statistics: Theory and Methods, 48 (2):368–380, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1410716.

Ozgul:2019:LE

[Özg19a] Nilgün Özgül. Letter to the Editor. Communications in Statistics: Theory and Methods, 48(2):177, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1410718.

Ozgul:2019:NCE

[Ozg19b] Nilgun Ozgul. New calibration estimator based on two auxiliary variables in stratified sampling. Communications in Statistics: Theory and Methods, 48(6):1481–1492, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1433852.

Paige:2018:SSC

[PA18] Robert L. Paige and Emad Abdurasul. Small sample confidence intervals for survival functions under the proportional hazards model. Communications in Statistics: Theory and Methods, 47(24):6108–6124, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1406514.

Pak:2014:MDP

[Pak14] Ro Jin Pak. The minimum density power divergence estimation for the lognormal density. Communications in Statistics: Theory and Methods, 43(21):4582–4588, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pak:2017:IFO

[Pak17] Ro Jin Pak. The influence function of the optimal bandwidth for kernel density estimation. Communications in Statistics: Theory and Methods, 46(2):602–608, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.1000501.

Pan:2019:EPA

[Pan19] Baoguo Pan. Estimation for periodic ARMA models with unspecified noises. Communications in Statistics: Theory and Methods, 48(12):2948–2961, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473597.

Paolino:2010:EIQ

[Pao10] D. S. Paolino. Exact inference for the p-th-quantile and the reliability of the two-parameter exponential distribution with singly Type II censoring: a standard approach. Communications in Statistics: Theory and Methods, 39(14):2561–2572, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papadopoulou:2013:SRM

[Pap13] Aleka A. Papadopoulou. Some results on modeling biological sequences and Web navigation with a semi Markov chain. Communications in Statistics: Theory and Methods, 42(16): 2853–2871, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2012:ABK

[Par12] Hyun Suk Park. Asymptotic behavior of the kernel density estimator from a geometric viewpoint. Communications in Statistics: Theory and Methods, 41(19):3479–3496, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2013:EDB

[Par13a] Changsoon Park. An economic design of the bounded reset chart using the integral controller for batch-oriented processes. Communications in Statistics: Theory and Methods, 42(18): 3296–3309, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2013:ACI

[Par13b] Dong Joon Park. Alternative confidence intervals on variance components in a simple regression model with a balanced twofold nested error structure. Communications in Statistics: Theory and Methods, 42(13):2281–2291, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2015:QRL

[Par15a] Jinho Park. Quantile regression with left-truncated and rightcensored data in a reproducing kernel Hilbert space. Communications in Statistics: Theory and Methods, 44(7):1523–1536, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2015:PCP

[Par15b] Mingue Park. Partially calibrated poststratified weights for handling unit non response. Communications in Statistics: Theory and Methods, 44(14):2984–3000, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2016:KLI

[Par16] Sangun Park. On the Kullback–Leibler information of hybrid censored data. Communications in Statistics: Theory and Methods, 45(15):4486–4493, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
921307.

Park:2017:TDD

[Par17a] Hyo-Il Park. On tests detecting difference in means and variances simultaneously under normality. Communications in Statistics: Theory and Methods, 46(20):10025–10035, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1228964.

Parker:2017:FSD

[Par17b] Thomas Parker. Finite-sample distributions of the Wald, likelihood ratio, and Lagrange multiplier test statistics in the classical linear model. Communications in Statistics: Theory and Methods, 46(11):5195–5202, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1096397.

Patterson:2016:TPC

[Pat16] David Patterson. A three-population constrained discrimination procedure. Communications in Statistics: Theory and Methods, 45(16):4771–4787, 2016. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
927501.

Peng:2019:CTM

[PAvS^{+}19] Ho-Lan Peng, Andrew Aschenbrenner, Kirk von Sternberg, Patricia D. Mullen, and Wenyaw Chan. A continuoustime Markov chain approach with the analytic likelihood in studies of behavioral changes. Communications in Statistics: Theory and Methods, 48(23):5756–5765, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1520886.

Panahbehagh:2017:GBI

[PB17] Bardia Panahbehagh and Jennifer Brown. Gap-based inverse sampling. Communications in Statistics: Theory and Methods, 46(19):9651–9661, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1217022.

Puza:2018:STU

[PB18] Borek Puza and Andre Bonfrer. A series of two-urn biased sampling problems. Communications in Statistics: Theory and Methods, 47(1):80–91, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1300282.

Parvardeh:2017:CRL

[PBA17] A. Parvardeh, N. Balakrishnan, and Azam Arshadipour. Conditional residual lifetimes of coherent systems under double monitoring. Communications in Statistics: Theory and Methods, 46(7):3401–3410, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1060348.

Parvardeh:2018:NCR

[PBA18] A. Parvardeh, N. Balakrishnan, and Azam Arshadipour. A note on the conditional residual lifetime of a coherent system under double monitoring. Communications in Statistics: Theory and Methods, 47(10):2373–2378, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.988267.

Panda:2017:SDR

[PBC17] Gopinath Panda, A. D. Banik, and M. L. Chaudhry. Stationary distributions of the R^{[X]} /R/1 cross-correlated queue. Communications in Statistics: Theory and Methods, 46(17):8666–8689, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1186192.

Pandey:2019:SFM

[PBL19] Arvind Pandey, Shashi Bhushan, and Ralte Lalpawimawha. Shared frailty models with baseline generalized Pareto distribution. Communications in Statistics: Theory and Methods, 48(17):4425–4447, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1500597.

Pereira:2012:TIB

[PBS12] Gustavo H. A. Pereira, Denise A. Botter, and Mônica C. Sandoval. The truncated inflated beta distribution. Communications in Statistics: Theory and Methods, 41(5):907–919, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peng:2014:SAP

[PBY^{+}14] Huiming Peng, Jiawen Bian, Diwei Yang, Zhihui Liu, and Hongwei Li. Statistical analysis of parameter estimation for 2-D harmonics in multiplicative and additive noise. Communications in Statistics: Theory and Methods, 43(22):4829–4844, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paula:2010:LIU

[PC10a] Gilberto A. Paula and Francisco José A. Cysneiros. Local influence under parameter constraints. Communications in Statistics: Theory and Methods, 39(7):1212–1228, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prudente:2010:GWL

[PC10b] Andrea A. Prudente and Gauss M. Cordeiro. Generalized Weibull linear models. Communications in Statistics: Theory and Methods, 39(20):3739–3755, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pereira:2012:SAP

[PC12] Luis N. Pereira and Pedro S. Coelho. A small area predictor under area-level linear mixed models with restrictions. Communications in Statistics: Theory and Methods, 41(13–14):2524– 2544, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pan:2016:QME

[PC16a] Baoguo Pan and Min Chen. Quasi-maximum exponential likelihood estimation for a non stationary GARCH(1, 1) model. Communications in Statistics: Theory and Methods, 45(4): 1000–1013, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2013.851225.

Pereira:2016:MNH

[PC16b] André G. C. Pereira and Viviane S. M. Campos. Multistage non homogeneous Markov chain modeling of the non homogeneous genetic algorithm and convergence results. Communications in Statistics: Theory and Methods, 45(6):1794–1804, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.997358.

Parrish:2016:YWE

[PCL16] Joan Parrish, Steven M. Crunk, and Bee Leng Lee. The Yule–Walker equations as a weighted least-squares problem and the association with tapering. Communications in Statistics: Theory and Methods, 45(17):5112–5122, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.936941.

Pinho:2015:HEE

[PCN15] Luis Gustavo Bastos Pinho, Gauss Moutinho Cordeiro, and Juvêncio Santos Nobre. The Harris extended exponential distribution. Communications in Statistics: Theory and Methods, 44(16):3486–3502, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pires:2018:SWR

[PCS18] Magda C. Pires, Enrico A. Colosimo, and Arlaine A. Silva. Survival Weibull regression model for mismeasured outcomes. Communications in Statistics: Theory and Methods, 47(3):601– 614, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1309434.

Petersen:2010:MLE

[PD10] Martin R. Petersen and James A. Deddens. Maximum likelihood estimation of the log-binomial model. Communications in Statistics: Theory and Methods, 39(5):874–883, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paraiba:2017:TLS

[PDdHNMR17] Carolina Costa Mota Paraíba, Carlos Alberto Ribeiro Diniz, Aline de Holanda Nunes Maia, and Lineu Neiva Rodrigues. Truncated location-scale non linear regression models. Communications in Statistics: Theory and Methods, 46(15):7355–7374, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1152485.

Pradhan:2009:PEQ

[PDS09] Biswabrata Pradhan, Anup Dewanji, and Debasis Sengupta. Parametric estimation of quality adjusted lifetime (QAL) distribution in simple illness-death model. Communications in Statistics: Theory and Methods, 39(1):77–93, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Psarrakos:2017:GCR

[PE17] Georgios Psarrakos and Polychronis Economou. On the generalized cumulative residual entropy weighted distributions. Communications in Statistics: Theory and Methods, 46(22):10914–10925, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1252402.

Pegoraro:2012:TCN

[Peg12] Italo Pegoraro. A transformation characterizing the normal distribution. Communications in Statistics: Theory and Methods, 41(16–17):3060–3067, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2012.
693238.

Peltier:2017:WIA

[Pel17] Corey Peltier. “What if” analysis: Benefits of utilizing a “What If” analysis in Excel. Communications in Statistics: Theory and Methods, 46(12):6119–6129, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1118511.

Peng:2017:CEE

[Pen17] Weicai Peng. Conditional entropy, entropy density, and strong law of large numbers for generalized controlled tree-indexed Markov chains. Communications in Statistics: Theory and Methods, 46(23):11880–11891, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1285935.

Pesarin:2015:SET

[Pes15] Fortunato Pesarin. Some elementary theory of permutation tests. Communications in Statistics: Theory and Methods, 44 (22):4880–4892, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pesta:2017:BBD

[Pes17] Michal Pesta. Block bootstrap for dependent errors-in-variables. Communications in Statistics: Theory and Methods, 46(4):1871–1897, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1030423.

Petropoulos:2010:CIE

[Pet10] Constantinos Petropoulos. A class of improved estimators for the scale parameter of a mixture model of exponential distribution with unknown location. Communications in Statistics:

Theory and Methods, 39(17):3153–3162, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pereira:2019:SMA

[PF19] Luísa Pereira and Cecília Fonseca. Statistical methods for assessing the contagion of spatial extreme events among regions. Communications in Statistics: Theory and Methods, 48(13): 3208–3218, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1473612.

Perez-Fernandez:2019:SMA

[PFD19] Raúl Pérez-Fernández and Bernard De Baets. The sample monomode and an associated test for discrete monomodality. Communications in Statistics: Theory and Methods, 48(21): 5419–5426, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1513148.

Polansky:2016:UOC

[PG16] Alan M. Polansky and Santu Ghosh. Using observed confidence levels to perform principal component analyses. Communications in Statistics: Theory and Methods, 45(12):3596–3611, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.904356.

Pak:2018:BIB

[PG18] Abbas Pak and Arjun Kumar Gupta. Bayesian inference on P (X > Y) in bivariate Rayleigh model. Communications in Statistics: Theory and Methods, 47(17):4095–4105, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1367814.

Paredes-Garcia:2019:ODE

[PGCT19] Wilfrido J. Paredes-García and Eduardo Castaño-Tostado. Optimal designs for estimating a choice hierarchy by a general nested multinomial logit model. Communications in Statistics: Theory and Methods, 48(23):5877–5888, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1523429.

Parsi:2009:SCI

[PGF09] S. Parsi, M. Ganjali, and N. Sanjari Farsipour. Simultaneous confidence intervals for the parameters of Pareto distribution under progressive censoring. Communications in Statistics: Theory and Methods, 39(1):94–106, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parsi:2011:CML

[PGF11] S. Parsi, M. Ganjali, and N. Sanjari Farsipour. Conditional maximum likelihood and interval estimation for two Weibull populations under joint Type-II progressive censoring. Communications in Statistics: Theory and Methods, 40(12):2117–2135, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pervez:2018:UNI

[PHH18] Zahid Pervez, Zawar Hussain, and Maryam Hafeez. On using non identical and independent Bernoulli trials in human surveys. Communications in Statistics: Theory and Methods, 47(8):1856–1867, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1328511.

Phillips:2018:QML

[Phi18] Robert F. Phillips. Quasi maximum likelihood estimation of dynamic panel data models. Communications in Statistics: Theory and Methods, 47(16):3970–3986, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1366521.

Peng:2011:RPD

[PHW11] Jiangyan Peng, Jin Huang, and Dingcheng Wang. The ruin probability of a discrete-time risk model with a one-sided linear claim process. Communications in Statistics: Theory and Methods, 40(24):4387–4399, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pakgohar:2019:LWD

[PHY19] A. Pakgohar, A. Habibirad, and F. Yousefzadeh. Lin–Wong divergence and relations on type I censored data. Communications in Statistics: Theory and Methods, 48(19):4804–4819, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1494839.

Piccolo:2015:IIC

[Pic15] Domenico Piccolo. Inferential issues on CUBE models with covariates. Communications in Statistics: Theory and Methods, 44(23):5023–5036, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pieczynski:2011:ESH

[Pie11] Wojciech Pieczynski. Exact smoothing in hidden conditionally Markov switching linear models. Communications in Statistics: Theory and Methods, 40(16):2823–2829, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pizzinga:2013:DRK

[Piz13] Adrian Pizzinga. Diffuse restricted Kalman filtering. Communications in Statistics: Theory and Methods, 42(17):3215–3233, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandya:2010:BEC

[PJ10] Mayuri Pandya and Prabha V. Jadav. Bayes estimation of change point in non standard mixture of left-truncated exponential with unknown proportions. Communications in Statistics: Theory and Methods, 39(15):2725–2742, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pourmousa:2014:BOS

[PJ14] Reza Pourmousa and Ahad Jamalizadeh. On bivariate order statistics from elliptical distributions. Communications in Statistics: Theory and Methods, 43(10–12):2183–2198, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pettyjohn:2013:ECI

[PJL13] Jeff Pettyjohn, Daniel R. Jeske, and Jun Li. Estimation and confidence intervals for clock offset in networks with bivariate exponential delays. Communications in Statistics: Theory and Methods, 42(6):1024–1041, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pettere:2011:RMF

[PK11] Gaida Pettere and Tõnu Kollo. Risk modeling for future cash flow using skew t-copula. Communications in Statistics: Theory and Methods, 40(16):2919–2925, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pericleous:2012:CME

[PK12] Katerina Pericleous and Stratis Kounias. The concept of majorization in experimental design. Communications in Statistics: Theory and Methods, 41(16–17):2836–2850, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2011.607533.

Pratikno:2017:TET

[PK17] Budi Pratikno and Shahjahan Khan. Testing equality of two intercepts for the parallel regression model with non sample prior information. Communications in Statistics: Theory and Methods, 46(6):3099–3110, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1054941.

Park:2018:ODU

[PK18] Chun Gun Park and Inyoung Kim. Outlier detection using difference-based variance estimators in multiple regression. Communications in Statistics: Theory and Methods, 47(24): 5986–6001, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1404101.

Patra:2019:ECH

[PK19] Lakshmi Kanta Patra and Somesh Kumar. Estimating the common hazard rate of several exponential distributions. Communications in Statistics: Theory and Methods, 48(19):4861–4873, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1500599.

Pak:2014:IBR

[PKJ14] A. Pak, N. B. Khoolenjani, and A. A. Jafari. Inference on P (Y < X) in bivariate Rayleigh distribution. Communications in Statistics: Theory and Methods, 43(22):4881–4892, 2014.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pasha:2017:ADI

[PKM17] M. A. Pasha, Y. Khadem, and M. B. Moghadam. Application of decreasing integrated hazard in Rahim and Banerjee model on economic design of X̄-control charts for systems with increasing failure rate and early replacement. Communications in Statistics: Theory and Methods, 46(21):10481–10494, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1236962.

Park:2012:EIR

[PL12] Sangun Park and Johan Lim. On the effect of imperfect ranking on the amount of Fisher information in ranked set samples. Communications in Statistics: Theory and Methods, 41(19): 3608–3620, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pumi:2017:CRP

[PL17] Guilherme Pumi and Sílvia R. C. Lopes. Copulas related to piecewise monotone functions of the interval and associated processes. Communications in Statistics: Theory and Methods, 46(2):828–860, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1006781.

Peng:2017:RPP

[PLC17] Hongfeng Peng, Zhijie Liu, and Tsangyao Chang. Revisiting purchasing power parity in BRICS countries using more powerful quantile unit-root tests with stationary covariates. Communications in Statistics: Theory and Methods, 46(20):10051– 10057, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1228968.

Pradhan:2017:DAM

[PLDS17] Upendra Kumar Pradhan, Krishan Lal, Sukanta Dash, and K. N. Singh. Design and analysis of mixture experiments with process variable. Communications in Statistics: Theory and Methods, 46(1):259–270, 2017. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
990104.

Pal:2012:ODE

[PM12] Manisha Pal and Nripes Kumar Mandal. Optimum designs for estimation of parameters in a quadratic mixture-amount model. Communications in Statistics: Theory and Methods, 41 (4):665–673, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parkash:2016:CME

[PM16a] Om Parkash and Mukesh. Contribution of maximum entropy principle in the field of queueing theory. Communications in Statistics: Theory and Methods, 45(12):3464–3472, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.875574.

Priyanka:2016:SER

[PM16b] Kumari Priyanka and Richa Mittal. Searching effective rotation patterns for population median using exponential type estimators in two-occasion rotation sampling. Communications in Statistics: Theory and Methods, 45(18):5443–5460, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.944661.

Priyanka:2017:FAI

[PM17] Kumari Priyanka and Richa Mittal. A fresh approach for intercession of nonresponse in multivariate longitudinal designs. Communications in Statistics: Theory and Methods, 46(18): 9303–9323, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1208240.

Pasha:2018:EBD

[PM18] M. A. Pasha and M. Bameni Moghadam. An extension of Ben–Daya–Rahim (2000) and Rahim–Banerjee (1993) cost models for economic design of T^{2}-control charts for multivariate deteriorating processes with preventive maintenance and early replacement. Communications in Statistics: Theory and Methods, 47(9):2098–2109, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1295159.

Pal:2011:ODP

[PMA11] Manisha Pal, Nripes K. Mandal, and M. L. Aggarwal. Optimum designs for parameter estimation in linear mixture models with synergistic effects. Communications in Statistics: Theory and Methods, 40(11):2059–2073, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pal:2014:TEA

[PMA14] Aditi Pal, Murari Mitra, and M. Z. Anis. On testing exponentiality against NBAFR alternatives. Communications in Statistics: Theory and Methods, 43(8):1834–1844, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Porcu:2010:NCS

[PMC10] E. Porcu, J. Mateu, and C. Comas. A note on continuous spatial-temporal dynamics of stochastic processes. Communications in Statistics: Theory and Methods, 39(19):3472–3484, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pereira:2013:MBE

[PMC13] Luis N. Pereira, Jorge M. Mendes, and Pedro S. Coelho. Modelbased estimation of unemployment rates in small areas of Portugal. Communications in Statistics: Theory and Methods, 42(7):1325–1342, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pasha:2017:GAE

[PMF17] M. A. Pasha, M. B. Moghadam, and Sh. Fani. A generalized approach to economic design of control charts: a discussion on implementing. Communications in Statistics: Theory and Methods, 46(21):10495–10506, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1236963.

Pasha:2018:EQC

[PMFK18] M. A. Pasha, M. B. Moghadam, S. Fani, and Y. Khadem. Effects of quality characteristic distributions on the integrated model of Taguchi’s loss function and economic statistical design of X̄-control charts by modifying the Banerjee and Rahim economic model. Communications in Statistics: Theory and Methods, 47(8):1842–1855, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1328512.

Pasha:2017:ITL

[PMKF17] M. A. Pasha, M. Bameni Moghadam, Y. Khadem, and S. Fani. An integration of Taguchi’s loss function in Banerjee–Rahim model for the economic and economic statistical design of X̄- control charts under multiple assignable causes and Weibull shock model. Communications in Statistics: Theory and Methods, 46(24):12113–12129, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1291975.

Pan:2017:PEC

[PMW17] Bai Jing Pan, Ma Ming, and Yang Ya Wen. Parameter estimation of the censored δ-shock model on uniform interval. Communications in Statistics: Theory and Methods, 46(14): 6936–6946, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1140780.

Pescim:2016:EGD

[PN16] Rodrigo R. Pescim and Saralees Nadarajah. An extension of the gamma distribution. Communications in Statistics: Theory and Methods, 45(18):5495–5516, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.944663.

Park:2018:GCK

[PNK18] Sangun Park, Hadi Alizadeh Noughabi, and Ilmun Kim. General cumulative Kullback–Leibler information. Communications in Statistics: Theory and Methods, 47(7):1551–1560, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321767.

Pham:2019:BAP

[PNPB19] Hoa Pham, Darfiana Nur, Huong T. T. Pham, and Alan Branford. A Bayesian approach for parameter estimation in multi-stage models. Communications in Statistics: Theory and Methods, 48(10):2459–2482, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1465090.

Popovic:2012:TSM

[Pop12] Bozidar V. Popović. On an Ar(1) time series model with marginal two parameter Wright inverse-gamma distribution. Communications in Statistics: Theory and Methods, 41(1):166– 177, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Popinski:2016:TRE

[Pop16] Waldemar Popiński. Trigonometric regression estimation for observations with additive and multiplicative errors. Communications in Statistics: Theory and Methods, 45(3):804–812, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pordzik:2011:OLS

[Por11] Paweł R. Pordzik. Orthogonality and linear sufficiency in partitioned and reduced linear models. Communications in Statistics: Theory and Methods, 40(6):1124–1130, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Porro:2015:ZDI

[Por15] Francesco Porro. Zenga distribution and inequality ordering. Communications in Statistics: Theory and Methods, 44(18): 3967–3977, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pescim:2019:NDP

[POS^{+}19] Rodrigo R. Pescim, Edwin M. M. Ortega, Adriano K. Suzuki, Vicente G. Cancho, and Gauss M. Cordeiro. A new destructive Poisson odd log-logistic generalized half-normal cure rate model. Communications in Statistics: Theory and Methods, 48(9):2113–2128, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1459709.

Pourtaheri:2017:TLC

[Pou17] Reza Pourtaheri. Three-level control charts with variable sample size, sampling interval, and control limits. Communications in Statistics: Theory and Methods, 46(4):1927–1940, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030426.

Park:2010:ECC

[PP10] Hee J. Park and Sung H. Park. Extension of central composite design for second-order response surface model building. Communications in Statistics: Theory and Methods, 39(7): 1202–1211, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Palmitesta:2012:GDM

[PP12] Paola Palmitesta and Corrado Provasi. GSH dependence modeling with an application to risk management. Communications in Statistics: Theory and Methods, 41(16–17):3030–3042, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/03610926.2012.686646.

Plunkett:2017:TST

[PP17] Amanda Plunkett and Junyong Park. Two-sample tests for sparse high-dimensional binary data. Communications in Statistics: Theory and Methods, 46(22):11181–11193, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1260743.

Park:2018:GCF

[PP18] Sangun Park and Jiwhan Park. A general class of flexible Weibull distributions. Communications in Statistics: Theory and Methods, 47(4):767–778, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1118509.

Parui:2019:BDI

[PPMD19] Shyamsundar Parui, Rajender Parsad, B. N. Mandal, and Sukanta Dash. Block designs for incomplete factorial treatment structures with two factors. Communications in Statistics: Theory and Methods, 48(24):6038–6053, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1524911.

Patil:2011:FWC

[PR11] H. S. Patil and R. N. Rattihalli. Fixed width confidence interval for the parameter of U(θ, mθ) distribution. Communications in Statistics: Theory and Methods, 40(5):907–915, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pavlenko:2017:SCH

[PR17] Tatjana Pavlenko and Anuradha Roy. Supervised classifiers for high-dimensional higher-order data with locally doubly exchangeable covariance structure. Communications in Statistics: Theory and Methods, 46(23):11612–11634, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1275695.

Prescott:2011:OBM

[Pre11] Philip Prescott. Orthogonally blocked mixture experiments in ellipsoidal restricted regions. Communications in Statistics: Theory and Methods, 40(5):763–784, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Princy:2015:MMK

[Pri15] T. Princy. Mixture models and the krätzel integral transform. Communications in Statistics: Theory and Methods, 44(2):390– 405, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Proia:2018:TRC

[Pro18] Frédéric Proïa. Testing for residual correlation of any order in the autoregressive process. Communications in Statistics: Theory and Methods, 47(3):628–654, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1310240.

Patil:2011:TEI

[PS11] M. K. Patil and D. T. Shirke. Tests for equality of inflation parameters of two zero-inflated power series distributions. Communications in Statistics: Theory and Methods, 40(14):2539– 2553, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2012:CNE

[PS12] Sangun Park and Dong Wan Shin. On the choice of nonparametric entropy estimator in entropy-based goodness-of-fit test statistics. Communications in Statistics: Theory and Methods, 41(5):809–819, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pap:2013:CDP

[PS13] Gyula Pap and Tamás T. Szabó. Change detection in INAR(p) processes against various alternative hypotheses. Communications in Statistics: Theory and Methods, 42(7):1386–1405, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paroissin:2014:FTN

[PS14] Christian Paroissin and Ali Salami. Failure time of non homogeneous gamma process. Communications in Statistics: Theory and Methods, 43(15):3148–3161, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parmet:2017:SDV

[PS17a] Y. Parmet and E. Schechtman. Simultaneous decomposition of the variance by sources and subgroups — new insights. Communications in Statistics: Theory and Methods, 46(15):7479–7494, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1152489.

Patil:2017:EDN

[PS17b] S. H. Patil and D. T. Shirke. Economic design of non parametric sign control chart. Communications in Statistics: Theory and Methods, 46(18):8987–8998, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1200095.

Pan:2018:ALL

[PS18] Juming Pan and Junfeng Shang. Adaptive LASSO for linear mixed model selection via profile log-likelihood. Communications in Statistics: Theory and Methods, 47(8):1882–1900, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1332219.

Pan:2016:RMS

[PSF16] Zhengqiang Pan, Quan Sun, and Jing Feng. Reliability modeling of systems with two dependent degrading components based on gamma processes. Communications in Statistics: Theory and Methods, 45(7):1923–1938, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.870201.

Pillai:2012:SPG

[PSP12] Thulasyammal R. Pillai, Mahendran Shitan, and Shelton Peiris. Some properties of the generalized autoregressive moving average (GARMA(1, 1; δ_{1}, δ_{2})) model. Communications in Statistics: Theory and Methods, 41(4):699–716, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:2015:LRT

[PSS^{+}15] Junyong Park, Bimal Sinha, Arvind Shah, Dihua Xu, and Jianxin Lin. Likelihood ratio tests for interval hypotheses with applications. Communications in Statistics: Theory and Methods, 44(11):2351–2370, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papadakis:2010:MPD

[PT10] Emmanuel N. Papadakis and Efthymios G. Tsionas. Multivariate Pareto distributions: Inference and financial applications. Communications in Statistics: Theory and Methods, 39(6):1013–1025, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Priya:2015:NUS

[PT15] R. S. Priya and P. Yageen Thomas. A note on the use of some functions of spacings in the estimation of common scale parameter of several symmetric distributions. Communications in Statistics: Theory and Methods, 44(14):2924–2932, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Priya:2016:ARS

[PT16] R. S. Priya and P. Yageen Thomas. An application of ranked set sampling when observations from several distributions are to be included in the sample. Communications in Statistics: Theory and Methods, 45(23):7040–7052, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.974819.

Paul:2017:CRR

[PT17] Jerin Paul and P. Yageen Thomas. Concomitant record ranked set sampling. Communications in Statistics: Theory and Methods, 46(19):9518–9540, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1213286.

Priyanka:2019:CCE

[PT19] Kumari Priyanka and Pidugu Trisandhya. A composite class of estimators using scrambled response mechanism for sensitive population mean in successive sampling. Communications in Statistics: Theory and Methods, 48(4):1009–1032, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1422762.

Phuong:2018:MEH

[PTV18] Cao Xuan Phuong, Dang Duc Trong, and Tran Quoc Viet. On the mean L^{1}-error in the heteroscedastic deconvolution problem with compactly supported noises. Communications in Statistics: Theory and Methods, 47(16):3871–3892, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1364389.

Pulcini:2013:MMA

[Pul13] Gianpaolo Pulcini. Modeling the mileage accumulation process with random effects. Communications in Statistics: Theory and Methods, 42(15):2661–2683, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Punathumparambath:2013:NFS

[Pun13] Bindu Punathumparambath. A new family of skewed slash distributions generated by the Cauchy kernel. Communications in Statistics: Theory and Methods, 42(13):2351–2361, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papadopoulou:2014:VCD

[PV14] Aleka Papadopoulou and P.-C. G. Vassiliou. On the variances and convariances of the duration state sizes of semi-Markov systems. Communications in Statistics: Theory and Methods, 43(7):1470–1483, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pathak:2016:VMD

[PV16] A. K. Pathak and P. Vellaisamy. Various measures of dependence of a new asymmetric generalized Farlie–Gumbel– Morgenstern copulas. Communications in Statistics: Theory and Methods, 45(18):5299–5317, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
942428.

Priebe:2013:OQQ

[PVB13] Carey E. Priebe, Joshua Vogelstein, and Davi Bock. Optimizing the quantity/quality trade-off in connectome inference. Communications in Statistics: Theory and Methods, 42(19): 3455–3462, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pielaszkiewicz:2017:W

[PVS17] Jolanta Pielaszkiewicz, Dietrich Von Rosen, and Martin Singull. On E[∏_{i=0}^{k} Tr{W^{m_{i}}}], where ∼ W_{√}(I, \). Communi-

cations in Statistics: Theory and Methods, 46(6):2990–3005, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1053942.

Podgorski:2011:ESM

[PW11] Krzysztof Podgórski and Jörg Wegener. Estimation for stochastic models driven by Laplace motion. Communications in Statistics: Theory and Methods, 40(18):3281–3302, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Podgorski:2016:CIS

[PW16] Krzysztof Podgórski and Jonas Wallin. Convolution-invariant subclasses of generalized hyperbolic distributions. Communications in Statistics: Theory and Methods, 45(1):98–103, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pararai:2017:EPL

[PWLO17] Mavis Pararai, Gayan Warahena-Liyanage, and Broderick O. Oluyede. Exponentiated power Lindley–Poisson distribution: Properties and applications. Communications in Statistics: Theory and Methods, 46(10):4726–4755, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1076473. See correction [AE23d].

Park:2016:CIR

[PY16a] Dong Joon Park and Min Yoon. Confidence intervals for the regression coefficient in a simple regression model with a balanced two-fold nested error structure. Communications in Statistics: Theory and Methods, 45(17):5053–5065, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.935435.

Puza:2016:ICI

[PY16b] Borek Puza and Mo Yang. Improved confidence intervals for the exponential mean via tail functions. Communications in Statistics: Theory and Methods, 45(2):529–539, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peng:2015:SAN

[PYB^{+}15] Huiming Peng, Shaoquan Yu, Jiawen Bian, Yujie Zhang, and Hongwei Li. Statistical analysis of non linear least squares estimation for harmonic signals in multiplicative and additive noise. Communications in Statistics: Theory and Methods, 44 (2):217–240, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pang:2017:SSO

[PYL17] Shanqi Pang, Rong Yan, and Sen Li. Schematic saturated orthogonal arrays obtained by using the contractive replacement method. Communications in Statistics: Theory and Methods, 46(18):8913–8924, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1197253.

Zang:2013:KER

[pZ13] Qing pei Zang. A kind of exact rates in complete moment convergence for moving-average processes. Communications in Statistics: Theory and Methods, 42(13):2464–2471, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pang:2015:AIM

[PZ15] Tianxiao Pang and Danna Zhang. Asymptotic inferences for an AR(1) model with a change point and possibly infinite variance. Communications in Statistics: Theory and Methods, 44(22): 4848–4865, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qiu:2016:CWS

[QC16] DeHua Qiu and PingYan Chen. Convergence for weighted sums of widely orthant dependent random variables. Communications in Statistics: Theory and Methods, 45(15):4445–4453, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.921304.

Qiu:2019:AAG

[QC19] Qingan Qiu and Lirong Cui. Availability analysis for general repairable systems with repair time threshold. Communications in Statistics: Theory and Methods, 48(3):628–647, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1417430.

Qiu:2019:AAO

[QCK19] Qingan Qiu, Lirong Cui, and Dejing Kong. Availability analysis and optimal inspection policy for systems with neglected down time. Communications in Statistics: Theory and Methods, 48(11):2787–2809, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473425.

Qingguo:2017:MEF

[Qin17] Tang Qingguo. M-estimation for functional linear regression. Communications in Statistics: Theory and Methods, 46(8): 3782–3800, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1073312.

Qiu:2017:SIU

[Qiu17] Zhiping Qiu. Statistical inference under imputation for proportional hazard model with missing covariates. Communications in Statistics: Theory and Methods, 46(23):11575–11590, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1275696.

Qiu:2018:FRR

[Qiu18] Guoxin Qiu. Further results on the residual quantile entropy. Communications in Statistics: Theory and Methods, 47(13): 3092–3103, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1348519.

Qin:2017:SSM

[QJKC17] Hong Qin, Nabakumar Jana, Somesh Kumar, and Kashinath Chatterjee. Stress-strength models with more than two states under exponential distribution. Communications in Statistics: Theory and Methods, 46(1):120–132, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.988257.

Qian:2018:PDF

[QJWC18] Linyi Qian, Zhuo Jin, Wei Wang, and Lyu Chen. Pricing dynamic fund protections for a hyperexponential jump diffusion process. Communications in Statistics: Theory and Methods, 47(1):210–221, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1301475.

Qomi:2017:STI

[QK17] M. Naghizadeh Qomi and A. Kiapour. Shortest tolerance intervals controlling both tails of the exponential distribution based on record values. Communications in Statistics: Theory and Methods, 46(1):271–279, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
990106.

Qureshi:2019:ERC

[QKCH19] Muhammad Nouman Qureshi, Sadia Khalil, Chang-Tai Chao, and Muhammad Hanif. Estimation of rare and clustered population variance in adaptive cluster sampling. Communications in Statistics: Theory and Methods, 48(21):5387–5400, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1513144.

Qeadan:2012:JDS

[QKP12] Fares Qeadan, Tomasz J. Kozubowski, and Anna K. Panorska. The joint distribution of the sum and the maximum of IID exponential random variables. Communications in Statistics: Theory and Methods, 41(3):544–569, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2011:QEP

[QL11a] Yongsong Qin and Qingzhu Lei. Quantile estimation in the presence of auxiliary information under negatively associated samples. Communications in Statistics: Theory and Methods, 40(23):4289–4307, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2011:ELP

[QL11b] Yongsong Qin and Yinghua Li. Empirical likelihood for partially linear models with missing responses: The fixed design case. Communications in Statistics: Theory and Methods, 40 (10):1849–1865, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2015:ELL

[QL15] Yongsong Qin and Qingzhu Lei. Empirical likelihood for linear models under linear process errors. Communications in Statistics: Theory and Methods, 44(15):3218–3233, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2018:BBT

[QL18] Ruibing Qin and Yang Liu. Block bootstrap testing for changes in persistence with heavy-tailed innovations. Communications in Statistics: Theory and Methods, 47(5):1104–1116, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1316398.

Qin:2019:RDM

[QL19] Ruibing Qin and Weiqi Liu. Ratio detection for mean change in α mixing observations. Communications in Statistics: Theory and Methods, 48(7):1693–1708, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1438623.

Qiu:2014:BLU

[QLZ14] Hongbing Qiu, Ji Luo, and Shurong Zheng. On Bayes linear unbiased estimator under the balanced loss function. Communications in Statistics: Theory and Methods, 43(24):5175–5194, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qomi:2012:EAS

[QNP12] M. Naghizadeh Qomi, N. Nematollahi, and A. Parsian. Estimation after selection under reflected normal loss function. Communications in Statistics: Theory and Methods, 41(6):1040– 1051, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qiu:2014:WEL

[QNZ14] Zhiping Qiu, Makhija Neeta, and Yong Zhou. Weighted estimator for the linear transformation models with multivariate failure time data. Communications in Statistics: Theory and Methods, 43(16):3516–3535, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qomi:2017:BSE

[Qom17] M. Naghizadeh Qomi. Bayesian shrinkage estimation based on Rayleigh type-II censored data. Communications in Statistics: Theory and Methods, 46(19):9859–9868, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1222440.

Qin:2014:CIN

[QQL14] Yongsong Qin, Tao Qiu, and Qingzhu Lei. Confidence intervals for nonparametric regression functions with missing data. Communications in Statistics: Theory and Methods, 43(19): 4123–4142, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qi:2010:DCP

[QTD10] Peiyan Qi, Zheng Tian, and Xifa Duan. Detection of change point in nonparametric function with unit-root noise by wavelet. Communications in Statistics: Theory and Methods, 40(2):187–198, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qi:2015:SMV

[QTD15] Peiyan Qi, Zheng Tian, and Xifa Duan. Sequential monitoring variance changes in location model. Communications in Statistics: Theory and Methods, 44(24):5267–5284, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qumsiyeh:2012:CBB

[Qum12] Maher B. Qumsiyeh. An L^{2} comparison between the bootstrap and the empirical Edgeworth expansion. Communications in Statistics: Theory and Methods, 41(2):251–261, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quoreshi:2017:BIV

[Quo17] A. M. M. Shahiduzzaman Quoreshi. A bivariate integer-valued long-memory model for high-frequency financial count data. Communications in Statistics: Theory and Methods, 46(3): 1080–1089, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.997361.

Qian:2019:PHE

[QWWW19] Linyi Qian, Wei Wang, Ning Wang, and Shuai Wang. Pricing and hedging equity-indexed annuities via local riskminimization. Communications in Statistics: Theory and Methods, 48(6):1417–1434, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1433848.

Qian:2014:VEI

[QWZ14] Linyi Qian, Rongming Wang, and Qian Zhao. Valuation of equity-indexed annuities with stochastic interest rate and jump diffusion. Communications in Statistics: Theory and Methods, 43(14):2870–2885, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qi:2017:CNU

[QYLL17] Zong-Feng Qi, Jian-Feng Yang, Yan Liu, and Min-Qian Liu. Construction of nearly uniform designs on irregular regions. Communications in Statistics: Theory and Methods, 46(17): 8318–8327, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1179755.

Qin:2011:LLL

[QZY11] Xu Qin, Jiang-She Zhang, and Xiao-Dong Yan. Local linear least squares kernel regression for linear and circular predictors. Communications in Statistics: Theory and Methods, 40(21): 3812–3823, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:2012:FMT

[QZY12] Xu Qin, Jiang-She Zhang, and Xiao-Dong Yan. A finite mixture three-parameter Weibull model for the analysis of wind speed data. Communications in Statistics: Theory and Methods, 41 (12):2160–2171, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qian:2010:MPP

[QZZN10] Cunhua Qian, Dingzhou Zhang, Xufeng Zhao, and Toshio Nakagawa. A maintenance policy for a parallel system with cascading failure. Communications in Statistics: Theory and Methods, 39(13):2394–2403, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Razmkhah:2011:NPI

[RA11] M. Razmkhah and Jafar Ahmadi. Nonparametric prediction intervals for future order statistics in a proportional hazard model. Communications in Statistics: Theory and Methods, 40(10):1807–1820, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roozbeh:2016:SRR

[RA16] Mahdi Roozbeh and Mohammad Arashi. Shrinkage ridge regression in partial linear models. Communications in Statistics: Theory and Methods, 45(20):6022–6044, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.955115.

Rao:2017:ERM

[RAA17] G. Srinivasa Rao, Muhammad Aslam, and Osama H. Arif. Estimation of reliability in multicomponent stress-strength based on two parameter exponentiated Weibull Distribution. Communications in Statistics: Theory and Methods, 46(15):7495–7502, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1154155.

Rashid:2017:SNA

[RAAJ17] Fareeha Rashid, Muhammad Akram, Atif Akbar, and Anum Javed. Some new augmented fractional Box–Behnken designs. Communications in Statistics: Theory and Methods, 46(4): 2007–2012, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1032423.

Rabhi:2013:TCH

[Rab13] Anissa Rabhi. On threshold choice in hypothesis testing for dynamical systems with small noise. Communications in Statistics: Theory and Methods, 42(5):800–811, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rezapour:2015:DNO

[RAB15] M. Rezapour, M. H. Alamatsaz, and N. Balakrishnan. Distribution of the number of observations greater than the i-th dependent progressively Type-II censored order statistic and its use in goodness-of-fit testing. Communications in Statistics: Theory and Methods, 44(12):2517–2529, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roozbeh:2012:SUR

[RAG12] M. Roozbeh, M. Arashi, and M. Gasparini. Seemingly unrelated ridge regression in semiparametric models. Communications in Statistics: Theory and Methods, 41(8):1364–1386, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rahimov:2012:CLS

[Rah12] I. Rahimov. Conditional least squares estimators for the offspring mean in a subcritical branching process with immigration. Communications in Statistics: Theory and Methods, 41 (12):2096–2110, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rahimov:2016:EOM

[Rah16] I. Rahimov. Estimation of the offspring mean in a branching process with non stationary immigration. Communications in Statistics: Theory and Methods, 45(19):5619–5627, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.948200.

Rahman:2018:AEV

[Rah18] M. Rahman. Asymptotic estimate of variance with applications to stochastic differential equations arises in mathematical neuroscience. Communications in Statistics: Theory and Methods, 47(2):289–306, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1303729.

Raissi:2015:AOI

[Raï15] Hamdi Raïssi. Autoregressive order identification for VAR models with non constant variance. Communications in Statistics: Theory and Methods, 44(10):2059–2078, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rajarshi:2018:CPC

[Raj18] M. B. Rajarshi. Comparison of pivotals for confidence bounds and intervals for the mean of a stationary time series. Communications in Statistics: Theory and Methods, 47(1):18–27, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1300276.

Rao:2015:BXD

[RAK15] G. Srinivasa Rao, Muhammad Aslam, and Debasis Kundu. Burr–XII distribution parametric estimation and estimation of reliability of multicomponent stress-strength. Communications in Statistics: Theory and Methods, 44(23):4953–4961, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:2016:IAR

[RAM16] Mohammad Z. Raqab and D. K. Al-Mutairi. Inferential analysis for the reliability parameter based on the threeparameter Lindley distribution. Communications in Statistics: Theory and Methods, 45(16):4923–4936, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.932810.

Ramprasath:2017:RSF

[Ram17] L. Ramprasath. Role of stylized features in constructing better estimators. Communications in Statistics: Theory and Methods, 46(15):7612–7620, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1157191.

Rao:2010:NDE

[Rao10] B. L. S. Prakasa Rao. Nonparametric density estimation for functional data via wavelets. Communications in Statistics: Theory and Methods, 39(8–9):1608–1618, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2017:CIH

[Rao17a] B. L. S. Prakasa Rao. Chebyshev’s inequality for Hilbert-space valued random elements with estimated mean and covariance. Communications in Statistics: Theory and Methods, 46(19): 9407–9414, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1208243.

Rao:2017:WED

[Rao17b] B. L. S. Prakasa Rao. Wavelet estimation for derivative of a density in a GARCH-type model. Communications in Statistics: Theory and Methods, 46(5):2396–2410, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1044671.

Rao:2018:SAL

[Rao18] B. L. S. Prakasa Rao. On some analogues of lack of memory properties for the Gompertz distribution. Communications in Statistics: Theory and Methods, 47(18):4415–4421, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1376084.

Rezapour:2013:MAA

[RAP13] Mohsen Rezapour, Mohammad Hossein Alamatsaz, and Franco Pellerey. Multivariate aging with Archimedean dependence structures in high dimensions. Communications in Statistics: Theory and Methods, 42(11):2056–2070, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:2010:AUB

[Raq10] Mohammad Z. Raqab. Attainability of the upper bounds for the mean past lifetime of parallel system components. Communications in Statistics: Theory and Methods, 39(17):3072–3080, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raschke:2017:CPD

[Ras17] Mathias Raschke. Criticism of the predictive distribution. Communications in Statistics: Theory and Methods, 46(15):7621– 7629, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1157884.

Rajesh:2014:NEG

[RASMN14] G. Rajesh, E. I. Abdul-Sathar, R. Maya, and K. R. Muraleedharan Nair. Nonparametric estimation of the geometric vitality function. Communications in Statistics: Theory and Methods, 43(1):115–130, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rajesh:2017:DWS

[RASR17] G. Rajesh, E. I. Abdul-Sathar, and S. Nair Rohini. On dynamic weighted survival entropy of order α. Communications in Statistics: Theory and Methods, 46(5):2139–2150, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1033552.

Rajab:2018:SNC

[RAST18] Muhammad Rajab, Rashid Ahmed, Farrukh Shehzad, and M. H. Tahir. Some new constructions of circular balanced repeated measurements designs. Communications in Statistics: Theory and Methods, 47(17):4142–4151, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1367817.

Rattanawong:2013:NUC

[Rat13] Petcharat Rattanawong. A non-uniform concentration inequality for a random permutation sum. Communications in Statistics: Theory and Methods, 42(10):1879–1888, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ravichandran:2019:TPA

[Rav19] J. Ravichandran. Transition probabilities and ARL performance of Six Sigma zone control charts. Communications in Statistics: Theory and Methods, 48(16):3976–3991, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1483510.

Razzaghi:2014:HMS

[Raz14] Mehdi Razzaghi. A hierarchical model for the skew-normal distribution with application in developmental neurotoxicology. Communications in Statistics: Theory and Methods, 43(8): 1859–1872, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rashid:2019:SNT

[RAZ19] Fareeha Rashid, Atif Akbar, and Zahra Zafar. Some new third order designs robust to one missing observation. Communications in Statistics: Theory and Methods, 48(24):6054–6062, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1528362.

Rasouli:2010:ELI

[RB10] Abbas Rasouli and N. Balakrishnan. Exact likelihood inference for two exponential populations under joint progressive Type-II censoring. Communications in Statistics: Theory and Methods, 39(12):2172–2191, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2011:RCN

[RB11] Sugata Sen Roy and Sankha Bhattacharya. Rate of convergence to normality of estimators in a random coefficient ARMA(p, q) model. Communications in Statistics: Theory and Methods, 40(6):1081–1092, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2013:MCP

[RBB13] Josemar Rodrigues, N. Balakrishnan, and Patrick Borges. Markov-correlated Poisson processes. Communications in Statistics: Theory and Methods, 42(20):3696–3703, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramesh:2019:BPN

[RBK^{+}19] V. P. Ramesh, Priyanga Baskaran, Aarthika Krishnamoorthy, Divya Damodaran, and Preethi Sadasivam. Back propagation neural network based big data analytics for a stock market challenge. Communications in Statistics: Theory and Methods, 48(14):3622–3642, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1478103.

Ristic:2019:ZIN

[RBN19] Miroslav M. Ristić, Marcelo Bourguignon, and Aleksandar S. Nastić. Zero-inflated NGINAR(1) process. Communications in Statistics: Theory and Methods, 48(3):726–741, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1435808.

Rakitzis:2016:EPE

[RC16] Athanasios C. Rakitzis and Philippe Castagliola. The effect of parameter estimation on the performance of one-sided Shewhart control charts for zero-inflated processes. Communications in Statistics: Theory and Methods, 45(14):4194–4214, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.917188.

Rego:2012:SPB

[RCC12] Leandro C. Rêgo, Renato J. Cintra, and Gauss M. Cordeiro. On some properties of the beta normal distribution. Communications in Statistics: Theory and Methods, 41(20):3722–3738, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2014:MLS

[RCdC14] Josemar Rodrigues, Gauss M. Cordeiro, and Mário de Castro. Modeling lifetimes by a stochastic process hitting a critical point. Communications in Statistics: Theory and Methods, 43 (10–12):2570–2580, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2016:UCC

[RCdCN16] Josemar Rodrigues, Gauss M. Cordeiro, Mário de Castro, and Saralees Nadarajah. A unified class of compound lifetime distributions. Communications in Statistics: Theory and Methods, 45(8):2323–2331, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.881491.

Ramprasath:2010:LMT

[RD10] L. Ramprasath and T. M. Durairajan. Locally maximin test for the mixture proportion with Type-I censoring. Communications in Statistics: Theory and Methods, 39(12):2264–2270, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rahmani:2016:EIR

[RD16] Samir Rahmani and Abdelnasser Dahmani. Exponential inequalities for the Robbins–Monro’s algorithm under mixing condition. Communications in Statistics: Theory and Methods, 45(2):520–528, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:2013:P

[RdC13] Paulo C. Rodrigues and Miguel de Carvalho. Preface. Communications in Statistics: Theory and Methods, 42(7):1143–1144, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rykov:2010:DMR

[RE10] Vladimir Rykov and Dmitry Efrosinin. Degradation models with random life resources. Communications in Statistics: Theory and Methods, 39(3):398–407, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reed:2012:PLA

[Ree12] William J. Reed. Power-law adjusted survival models. Communications in Statistics: Theory and Methods, 41(20):3692–3703, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rejchel:2016:LCL

[Rej16] Wojciech Rejchel. Lasso with convex loss: Model selection consistency and estimation. Communications in Statistics: Theory and Methods, 45(7):1989–2004, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.870799.

Ren:2016:EEU

[Ren16] Xingwei Ren. The equalities of estimations under a general partitioned linear model and its stochastically restricted model. Communications in Statistics: Theory and Methods, 45(22): 6495–6509, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.960587.

Rezaei:2016:PMP

[Rez16] Majid Rezaei. On proportional mean past lifetimes model. Communications in Statistics: Theory and Methods, 45(14): 4035–4047, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.915039.

Rezapour:2017:PTI

[Rez17] Mohsen Rezapour. Progressively Type-II censored conditionally N-ordered statistics from a unified elliptically contoured copula. Communications in Statistics: Theory and Methods, 46(11):5595–5611, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1107584.

Roy:2012:LMD

[RF12] Anuradha Roy and Miguel Fonseca. Linear models with doubly exchangeable distributed errors. Communications in Statistics: Theory and Methods, 41(13–14):2545–2569, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rasay:2018:IME

[RFM18] Hasan Rasay, Mohammad Saber Fallahnezhad, and Yahia Zare Mehrjerdi. An integrated model for economic design of chisquare control chart and maintenance planning. Communications in Statistics: Theory and Methods, 47(12):2892–2907, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1343848.

Rezaei:2015:MPM

[RG15] Majid Rezaei and Behzad Gholizadeh. On the mixture proportional mean residual life model. Communications in Statistics: Theory and Methods, 44(20):4263–4277, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ram:2016:NSS

[RG16] Mangey Ram and Arpan Gupta. Numerical solution to stochastic partial differential equation with variable repair rate. Communications in Statistics: Theory and Methods, 45(21):6245– 6252, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.957863.

Robertson:2019:CUE

[RG19] David S. Robertson and Ekkehard Glimm. Conditionally unbiased estimation in the normal setting with unknown variances. Communications in Statistics: Theory and Methods, 48 (3):616–627, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1417429.

Reyes:2019:NCS

[RGBG19] Jimmy Reyes, Diego I. Gallardo, Heleno Bolfarine, and Héctor W. Gómez. A new class of slash-elliptical distributions. Communications in Statistics: Theory and Methods, 48(12): 3105–3121, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1473607.

Rezaei:2015:RRH

[RGI15] Majid Rezaei, Behzad Gholizadeh, and Salman Izadkhah. On relative reversed hazard rate order. Communications in Statistics: Theory and Methods, 44(2):300–308, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rifo:2012:FBA

[RGL12] Laura L. R. Rifo and Verónica González-López. Full Bayesian analysis for a model of tail dependence. Communications in Statistics: Theory and Methods, 41(22):4107–4123, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reyes:2015:RKF

[RGM15] Adriana Reyes, Ramón Giraldo, and Jorge Mateu. Residual kriging for functional spatial prediction of salinity curves. Communications in Statistics: Theory and Methods, 44(4):798–809, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reyes:2016:MSS

[RGV16] Jimmy Reyes, Héctor W. Gómez, and Ignacio Vidal. Modified skew-slash distribution. Communications in Statistics: Theory and Methods, 45(4):1070–1080, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.854913.

Richter:2013:GSR

[Ric13] Wolf-Dieter Richter. Geometric and stochastic representations for elliptically contoured distributions. Communications in Statistics: Theory and Methods, 42(4):579–602, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Richard:2017:HRT

[Ric17] Patrick Richard. Heteroskedasticity-robust tests with minimum size distortion. Communications in Statistics: Theory and Methods, 46(13):6463–6477, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1125497.

Rolls:2015:ITC

[RJ15] David A. Rolls and Owen D. Jones. An improved test for continuous local martingales. Communications in Statistics: Theory and Methods, 44(13):2674–2688, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roozegar:2016:CMO

[RJN16] Roohollah Roozegar, Ahad Jamalizadeh, and Alireza Nematollahi. Concomitants of multivariate order statistics from multivariate elliptical distributions. Communications in Statistics: Theory and Methods, 45(3):722–738, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rogerson:2016:CSU

[RK16] Peter Rogerson and Peter Kedron. A changepoint statistic with uniform Type I error probabilities. Communications in Statistics: Theory and Methods, 45(16):4663–4672, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.927495.

Rudiuk:2018:EDF

[RK18] Edmund Rudiuk and Aleksander Kowalski. The exact distribution function of the ratio of two dependent quadratic forms. Communications in Statistics: Theory and Methods, 47(21): 5227–5240, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1388400.

Reddy:2019:OSS

[RK19] Karuna G. Reddy and M. G. M. Khan. Optimal stratification in stratified designs using Weibull-distributed auxiliary information. Communications in Statistics: Theory and Methods, 48(12):3136–3152, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473609.

Ringham:2019:DSS

[RKMG19] B. M. Ringham, S. M. Kreidler, K. E. Muller, and D. H. Glueck. On the distribution of summary statistics for missing data. Communications in Statistics: Theory and Methods, 48(5): 1149–1165, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1425447.

Rong:2009:SER

[RL09] Jian-Ying Rong and Xu-Qing Liu. On superiority of estimation regarding Pitman closeness criterion in general mixed linear models. Communications in Statistics: Theory and Methods, 39 (2):256–265, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramos:2011:APP

[RL11a] Alexandra Ramos and Anthony Ledford. An alternative point process framework for modeling multivariate extreme values. Communications in Statistics: Theory and Methods, 40(12): 2205–2224, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2011:ETS

[RL11b] Anuradha Roy and Ricardo Leiva. Estimating and testing a structured covariance matrix for three-level multivariate data. Communications in Statistics: Theory and Methods, 40(11): 1945–1963, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2012:CRM

[RL12] Anuradha Roy and Ricardo Leiva. Classification rules for multivariate repeated measures data with equicorrelated correlation structure on both time and spatial repeated measurements. Communications in Statistics: Theory and Methods, 41(8):1411–1420, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rong:2012:QFM

[RLL12] Jian-Ying Rong, Zhi-Feng Lu, and Xu-Qing Liu. On quadratic forms of multivariate t distribution with applications. Communications in Statistics: Theory and Methods, 41(2):300–308, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramos:2019:IWL

[RLSL19] Pedro L. Ramos, Francisco Louzada, Taciana K. O. Shimizu, and Aline O. Luiz. The inverse weighted Lindley distribution: Properties, estimation and an application on a failure time data. Communications in Statistics: Theory and Methods, 48(10):2372–2389, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1465084.

Ren:2010:NNR

[RM10] Qunshu Ren and Majid Mojirsheibani. A note on nonparametric regression with β-mixing sequences. Communications in Statistics: Theory and Methods, 39(12):2280–2287, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2014:BAL

[RM14] Soumya Roy and Chiranjit Mukhopadhyay. Bayesian accelerated life testing under competing Weibull causes of failure. Communications in Statistics: Theory and Methods, 43(10–12): 2429–2451, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roli:2015:HBM

[RM15] Giulia Roli and Paola Monari. Hierarchical Bayesian models for the estimation of correlated effects in multilevel data: a simulation study to assess model performance. Communications in Statistics: Theory and Methods, 44(12):2644–2653, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehta:2016:MRP

[RM16] Nitu Mehta (Ranka) and V. L. Mandowara. A modified ratiocum-product estimator of finite population mean using ranked set sampling. Communications in Statistics: Theory and Methods, 45(2):267–276, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruiz-Medina:2011:MCP

[RMC11] María Dolores Ruiz-Medina and Rosa María Crujeiras. Minimum contrast parameter estimation for fractal random fields based on the wavelet periodogram. Communications in Statistics: Theory and Methods, 40(19–20):3599–3613, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rafie:2017:GVB

[RMM17] Saeedreza Rafie, Masoud Ghaderi Moghadam, and M. B. Moghadam. A generalized version of Banerjee and Rahim model in economic and economic statistical designs of multivariate control charts under generalized exponential shock model. Communications in Statistics: Theory and Methods, 46(16):8016–8026, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1171354.

Ribereau:2016:SGE

[RMN16] Pierre Ribereau, Esterina Masiello, and Philippe Naveau. Skew generalized extreme value distribution: Probability-weighted moments estimation and application to block maxima procedure. Communications in Statistics: Theory and Methods, 45(17):5037–5052, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
935434.

Rezaei:2017:NEC

[RMNA17] S. Rezaei, A. K. Marvasty, S. Nadarajah, and M. Alizadeh. A new exponentiated class of distributions: Properties and applications. Communications in Statistics: Theory and Methods, 46(12):6054–6073, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1116579.

Roli:2012:ECE

[RMP12] Giulia Roli, Fabrizia Mealli, and Barbara Pacini. Estimation of causal effects in latent strata with an encouragement for response. Communications in Statistics: Theory and Methods, 41(16–17):3150–3161, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
587628.

Rerkruthairat:2013:INU

[RN13] N. Rerkruthairat and K. Neammanee. An improvement of a non uniform bound on normal approximation of randomized orthogonal array sampling designs. Communications in Statistics: Theory and Methods, 42(1):118–134, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roozegar:2017:PSS

[RN17] Rasool Roozegar and Saralees Nadarajah. The power series skew normal class of distributions. Communications in Statistics: Theory and Methods, 46(22):11404–11423, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1267758.

Ristic:2012:EIV

[RNB12] Miroslav M. Ristić, Aleksandar S. Nastić, and Hassan S. Bakouch. Estimation in an integer-valued autoregressive process with negative binomial marginals (NBINAR(1)). Communications in Statistics: Theory and Methods, 41(4):606–618, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez:2016:ODH

[ROM16] Carmelo Rodríguez, Isabel Ortiz, and Ignacio Martínez. Aoptimal designs for heteroscedastic multifactor regression models. Communications in Statistics: Theory and Methods, 45(3): 757–771, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rebelo:2012:NTC

[RON12] Efigénio Rebelo, Patrícia Oom Do Valle, and Rui Nunes. Nonnested testing for competing autoregressive dynamic models estimated by instrumental variables. Communications in Statistics: Theory and Methods, 41(20):3799–3812, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roozegar:2017:LBR

[Roo17] Rasool Roozegar. Limiting behavior of randomly weighted averages of symmetric heavy-tailed random variables. Communications in Statistics: Theory and Methods, 46(9):4539–4544, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1085571.

Roodman:2018:BSC

[Roo18] David Roodman. Bias and size corrections in extreme value modeling. Communications in Statistics: Theory and Methods, 47(14):3377–3391, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1353630.

Rosa:2019:OBD

[Ros19] Samuel Rosa. E- and R-optimality of block designs for treatment-control comparisons. Communications in Statistics: Theory and Methods, 48(21):5233–5244, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1508720.

Roy:2016:AOP

[Roy16] Surupa Roy. Analysis of ordered probit model with surrogate response data and measurement error in covariates. Communications in Statistics: Theory and Methods, 45(9):2665–2678, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.887115.

Rozora:2018:SHT

[Roz18] Iryna V. Rozora. Statistical hypothesis testing for the shape of impulse response function. Communications in Statistics: Theory and Methods, 47(6):1459–1474, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321125.

Rattihalli:2010:GSD

[RP10] R. N. Rattihalli and P. Y. Patil. Generalized υ-spherical densities. Communications in Statistics: Theory and Methods, 39 (19):3568–3583, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rukhin:2012:LDG

[RP12] Andrey Rukhin and Carey E. Priebe. On the limiting distribution of a graph scan statistic. Communications in Statistics: Theory and Methods, 41(7):1151–1170, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ravi:2015:IMD

[RP15] S. Ravi and A. S. Praveena. On the intersection of max domains of attraction of p-max stable laws and the class of subexponential distributions. Communications in Statistics: Theory and Methods, 44(1):115–124, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Romero-Padilla:2017:EVB

[RPH17] Juan Manuel Romero-Padilla and Chein-Pai Han. Estimation of variance in bivariate normal distribution after the preliminary test of homogeneity. Communications in Statistics: Theory and Methods, 46(3):1290–1305, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1014114.

Rosco:2015:BMK

[RPJ15] J. F. Rosco, Arthur Pewsey, and M. C. Jones. On Blest’s measure of kurtosis adjusted for skewness. Communications in Statistics: Theory and Methods, 44(17):3628–3638, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rezakhah:2018:IMM

[RPM18] S. Rezakhah, A. Philippe, and N. Modarresi. Innovative methods for modeling of scale invariant processes. Communications in Statistics: Theory and Methods, 47(13):3178–3191, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1350273.

Rai:2013:EIR

[RPY^{+}13] Shesh N. Rai, Jianmin Pan, Xiaobin Yuan, Jianguo Sun, Melissa M. Hudson, and Deo K. Srivastava. Estimating incidence rate on current status data with application to a phase IV cancer trial. Communications in Statistics: Theory and Methods, 42(17):3117–3133, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2009:CBM

[RR09] Dilip Roy and Rishideep Roy. Characterizations of bivariate and multivariate life distributions based on reciprocal subtangent. Communications in Statistics: Theory and Methods, 39 (1):158–169, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2013:SCR

[RR13] Dilip Roy and Rishideep Roy. Stability of the characterization results in terms of hazard rate and mean residual life for the univariate and bivariate setups. Communications in Statistics: Theory and Methods, 42(9):1583–1598, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roy:2014:LMP

[RR14] Rishideep Roy and Dilip Roy. The lack of memory property in the density form for the bivariate setup. Communications in Statistics: Theory and Methods, 43(14):2859–2869, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rasheed:2018:MCS

[RRRA18] Uzma Rasheed, H. M. Kashif Rasheed, Muhammad Rasheed, and Rashid Ahmed. Minimal circular strongly balanced repeated measurements designs in periods of three different sizes. Communications in Statistics: Theory and Methods, 47(16): 4088–4094, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1361995.

Reid:2010:ASP

[RS10] N. Reid and Y. Sun. Assessing sensitivity to priors using higher order approximations. Communications in Statistics: Theory and Methods, 39(8–9):1373–1386, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rajarajan:2013:EDQ

[RS13] Jyoti Rajarajan and Chhaya Sonar. G-efficient designs for quadratic mixture models. Communications in Statistics: Theory and Methods, 42(24):4417–4430, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2014:NHR

[RS14] T. J. Rao and A. K. P. C. Swain. A note on the Hartley–Ross unbiased ratio estimator. Communications in Statistics: Theory and Methods, 43(15):3162–3169, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rackauskas:2016:CDS

[RS16] Alfredas Rackauskas and Charles Suquet. Computing the distribution of sequential Hölder norms of the Brownian motion. Communications in Statistics: Theory and Methods, 45(15): 4378–4391, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.911903.

Ryu:2017:IIV

[RS17] Hang Keun Ryu and Daniel J. Slottje. Income inequality versus utility inequality. Communications in Statistics: Theory and Methods, 46(8):3631–3640, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1069348.

Razie:2019:JHR

[RS19] F. Razie and E. Bahrami Samani. Joint (k, l)-hurdle random effects models for mixed longitudinal power series and normal outcomes. Communications in Statistics: Theory and Methods, 48(12):3025–3043, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473601.

Rezapour:2013:SCR

[RSA13] M. Rezapour, E. T. Salehi, and M. H. Alamatsaz. Stochastic comparison of residual and past lifetimes of (n − k + 1)-out-of-n systems with dependent components. Communications in Statistics: Theory and Methods, 42(12):2185–2199, 2013.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rabia:2019:RSC

[RSA19] Mahwish Rabia, Nadia Saeed, and Muhammad Aslam. Reliability and sensitivity comparisons and average run lengths of CUSUM scale chart. Communications in Statistics: Theory and Methods, 48(9):2147–2162, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1459712.

Rezaei:2017:TLG

[RSAN17] Sadegh Rezaei, Behnam Bahrami Sadr, Morad Alizadeh, and Saralees Nadarajah. Topp–Leone generated family of distributions: Properties and applications. Communications in Statistics: Theory and Methods, 46(6):2893–2909, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1053935.

Raman:2016:CAE

[RSC16] Rohan Kumar Raman, U. C. Sud, and Hukum Chandra. Calibration approach for estimating population total with subsampling of non respondents under single- and two-phase sampling. Communications in Statistics: Theory and Methods, 45(10): 2842–2856, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.889164.

Razie:2017:LVM

[RSG17] F. Razie, E. Bahrami Samani, and M. Ganjali. Latent variable model for mixed correlated power series and ordinal longitudinal responses with non ignorable missing values. Communications in Statistics: Theory and Methods, 46(12):5738–5753, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1105980.

Racugno:2010:BAR

[RSV10] Walter Racugno, Alessandra Salvan, and Laura Ventura. Bayesian analysis in regression models using pseudo-likelihoods. Communications in Statistics: Theory and Methods, 39(19):

3444–3455, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rios:2011:SEA

[RSV11] Armando J. Rios, James R. Simpson, and José A. Vázquez. Sequential experimentation approach for augmenting of resolution III fractions. Communications in Statistics: Theory and Methods, 40(13):2337–2357, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rajesh:2017:EIM

[RSV17] G. Rajesh, E. I. Abdul Sathar, and K. V. Viswakala. Estimation of inaccuracy measure for censored dependent data. Communications in Statistics: Theory and Methods, 46(20):10058–10070, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1228969.

Rong:2017:LDA

[RTT17] Yaohua Rong, Manlai Tang, and Maozai Tian. Longitudinal data analysis based on generalized linear partially varyingcoefficient models. Communications in Statistics: Theory and Methods, 46(4):1983–2001, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1032421.

Rublik:2010:NSC

[Rub10] Frantisek Rublík. A note on simultaneous confidence intervals for ratio of variances. Communications in Statistics: Theory and Methods, 39(6):1038–1045, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ranganai:2014:MCH

[RVD14] Edmore Ranganai, Johan O. Van Vuuren, and Tertius De Wet. Multiple case high leverage diagnosis in regression quantiles. Communications in Statistics: Theory and Methods, 43(16): 3343–3370, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raubenheimer:2014:BEL

[RvdM14] Lizanne Raubenheimer and Abrie van der Merwe. Bayesian estimation for linear functions of Poisson rates. Communications in Statistics: Theory and Methods, 43(19):4025–4045, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruffa:2015:ACS

[RVP15] Suela Ruffa, Grazia Vicario, and Giovanni Pistone. Analysis of the covariance structure in manufactured parts. Communications in Statistics: Theory and Methods, 44(21):4540–4551, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rakhmawati:2016:PEP

[RWY16] Dwi Yuli Rakhmawati, Chien-Wei Wu, and Chao-Lung Yang. Performance evaluation of processes with asymmetric tolerances in the presence of gauge measurement errors. Communications in Statistics: Theory and Methods, 45(10):3011–3026, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.894068.

Rychlik:2011:SNB

[Ryc11] Tomasz Rychlik. Sharp negative bounds on differences of small quantiles and means of distributions from nonparametric families. Communications in Statistics: Theory and Methods, 40 (12):2155–2171, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rychlik:2014:MEE

[Ryc14a] Tomasz Rychlik. Maximal expectations of extreme order statistics from increasing density and failure rate populations. Communications in Statistics: Theory and Methods, 43(10–12): 2199–2213, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rychlik:2014:QIE

[Ryc14b] Tomasz Rychlik. Quantile inequalities for the expectations of generalized L-statistics. Communications in Statistics: Theory and Methods, 43(16):3443–3463, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ren:2013:MSV

[RZ13] Yunwen Ren and Xinsheng Zhang. Model selection for vector autoregressive processes via adaptive lasso. Communications in Statistics: Theory and Methods, 42(13):2423–2436, 2013.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rempel:2014:EOD

[RZ14] Michael F. Rempel and Julie Zhou. On exact K-optimal designs minimizing the condition number. Communications in Statistics: Theory and Methods, 43(6):1114–1131, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:2016:DPP

[RZHC16] M. Bhaskara Rao, Haimeng Zhang, Chunfeng Huang, and Fu-Chih Cheng. A discrete probability problem in card shuffling. Communications in Statistics: Theory and Methods, 45(3):612– 620, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shore:2010:SCG

[SA10] Haim Shore and Fatina A’wad. Statistical comparison of the goodness of fit delivered by five families of distributions used in distribution fitting. Communications in Statistics: Theory and Methods, 39(10):1707–1728, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:2012:BMO

[SA12] Samir M. Shaarawy and Sherif S. Ali. Bayesian model order selection of vector moving average processes. Communications in Statistics: Theory and Methods, 41(4):684–698, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stracqualursi:2014:RCT

[SA14] L. Stracqualursi and P. Agati. The role of classification trees and expert knowledge in building Bayesian networks: a case study in medicine. Communications in Statistics: Theory and Methods, 43(4):839–850, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaarawy:2015:BIS

[SA15a] Samir M. Shaarawy and Sherif S. Ali. Bayesian identification of seasonal multivariate autoregressive processes. Communications in Statistics: Theory and Methods, 44(4):823–836, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soltani:2015:ACD

[SA15b] A. R. Soltani and S. M. Aboukhamseen. An alternative cluster detection test in spatial scan statistics. Communications in Statistics: Theory and Methods, 44(8):1592–1601, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shahriari:2016:RCC

[SA16] Hamid Shahriari and Orod Ahmadi. A robust R control chart based on a two-step estimator of the process dispersion. Communications in Statistics: Theory and Methods, 45(9):2504– 2523, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.882952.

Shahriari:2017:REC

[SA17a] Hamid Shahriari and Orod Ahmadi. Robust estimation of complicated profiles using wavelets. Communications in Statistics: Theory and Methods, 46(4):1573–1593, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1024859.

Smallman:2017:SIS

[SA17b] Luke Smallman and Andreas Artemiou. A study on imbalance support vector machine algorithms for sufficient dimension reduction. Communications in Statistics: Theory and Methods, 46(6):2751–2763, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1048889.

Soltani:2017:LBR

[SA17c] Ahmad Reza Soltani and Kamel Abdollahnezhad. Limiting behavior of random Stieltjes partial sum: adjusted method of moments estimators. Communications in Statistics: Theory and Methods, 46(18):9053–9062, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1202283.

Shishebor:2019:ABP

[SA19a] Z. Shishebor and B. Amiri Aghbilagh. On the asymptotic behavior of the periodograms of periodically correlated spatial processes: Periodicity detection. Communications in Statistics: Theory and Methods, 48(8):1854–1870, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1440309.

Singh:2019:EBB

[SA19b] Saroja Kumar Singh and Sarat Kumar Acharya. Equivalence between Bayes and the maximum likelihood estimator in M/M/1 queue. Communications in Statistics: Theory and Methods, 48(19):4780–4793, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481971.

Sen:2015:GCT

[SAB15] Kanwar Sen, Manju L. Agarwal, and Sonali Bhattacharya. Geiger counter-type Pólya–Eggenberger distributions. Communications in Statistics: Theory and Methods, 44(23):4912–4926, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sani:2018:CCW

[SAB18] H. R. Nili Sani, M. Amini, and A. Bozorgnia. Complete convergence for weighted sums of arrays of APND random variables. Communications in Statistics: Theory and Methods, 47(10): 2425–2431, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1205623.

Saber:2019:PEB

[Sab19] Mohammad Mehdi Saber. Performance of extrapolation based on Pitman’s measure of closeness in spatial regression models with extended skew t innovations. Communications in Statistics: Theory and Methods, 48(2):282–299, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1408830.

Stavropoulos:2013:LRT

[SAC13] A. Stavropoulos, M. G. Akritas, and C. Caroni. Likelihood ratio type statistics for repeated measures designs with heterogeneous covariance matrices. Communications in Statistics: Theory and Methods, 42(6):1070–1086, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salehi:2016:CTS

[SAD16] Mahdi Salehi, Jafar Ahmadi, and Sanku Dey. Comparison of two sampling schemes for generating record-breaking data from the proportional hazard rate models. Communications in Statistics: Theory and Methods, 45(13):3721–3733, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.909938.

Saadi:2019:EQD

[SAD19] Nora Saadi, Smail Adjabi, and Lamia Djerroud. On the estimation of the quantile density function by orthogonal series. Communications in Statistics: Theory and Methods, 48(21): 5265–5289, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1510003.

Saghir:2015:PDS

[Sag15] Aamir Saghir. Phase-I design scheme for X̄-chart based on posterior distribution. Communications in Statistics: Theory and Methods, 44(3):644–655, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szczepanska-Alvarez:2017:EEM

[SÁHLvR17] Anna Szczepańska-Álvarez, Chengcheng Hao, Yuli Liang, and Dietrich von Rosen. Estimation equations for multivariate linear models with Kronecker structured covariance matrices. Communications in Statistics: Theory and Methods, 46(16): 7902–7915, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1165852.

Sajeevkumar:2011:UQR

[Saj11] N. K. Sajeevkumar. Use of quasi-ranges in the estimation of scale parameter of a symmetric distribution. Communications in Statistics: Theory and Methods, 40(10):1780–1786, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadooghi-Alvandi:2013:PBA

[SAJ13] S. M. Sadooghi-Alvandi and A. A. Jafari. A parametric bootstrap approach for one-way ANCOVA with unequal variances. Communications in Statistics: Theory and Methods, 42(14): 2473–2498, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadooghi-Alvandi:2012:OWA

[SAJMF12] S. M. Sadooghi-Alvandi, A. A. Jafari, and H. A. Mardani-Fard. One-way ANOVA with unequal variances. Communications in Statistics: Theory and Methods, 41(22):4200–4221, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadooghi-Alvandi:2015:TTF

[SAKK15] S. M. Sadooghi-Alvandi and M. Kharrati-Kopaei. Testing three-factor interaction in unreplicated three-way layouts. Communications in Statistics: Theory and Methods, 44(3):587– 606, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salmaso:2015:CBP

[Sal15] Luigi Salmaso. Combination-based permutation tests: Equipower property and power behavior in presence of correlation. Communications in Statistics: Theory and Methods, 44(24):5225– 5239, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samuh:2011:EMR

[SAS11] Monjed Hisham Samuh and Mohammad Fraiwan Al-Saleh. The effectiveness of multistage ranked set sampling in stratifying the population. Communications in Statistics: Theory and Methods, 40(6):1063–1080, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadooghi-Alvandi:2013:SBC

[SASMF13] S. M. Sadooghi-Alvandi, Z. Shishebor, and H. A. Mardani-Fard. Sharp bounds on a class of copulas with known values at several points. Communications in Statistics: Theory and Methods, 42 (12):2215–2228, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singull:2012:MKS

[SAvR12] Martin Singull, M. Rauf Ahmad, and Dietrich von Rosen. More on the Kronecker structured covariance matrix. Communications in Statistics: Theory and Methods, 41(13–14):2512–2523, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sayyareh:2014:LSR

[Say14] Abdolreza Sayyareh. Linear signed rank test for model selection. Communications in Statistics: Theory and Methods, 43 (21):4492–4502, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sayyareh:2017:NPM

[Say17] Abdolreza Sayyareh. Non parametric multiple comparisons of non nested rival models. Communications in Statistics: Theory and Methods, 46(17):8369–8386, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1179759.

Stepien-Baran:2011:SEC

[SB11] Agnieszka Stepień-Baran. Sequential estimation of a continuous distribution function from delayed observations. Communications in Statistics: Theory and Methods, 40(10):1717–1733, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scutari:2012:BNS

[SB12] Marco Scutari and Adriana Brogini. Bayesian network structure learning with permutation tests. Communications in Statistics: Theory and Methods, 41(16–17):3233–3243, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2011.593284.

Sharma:2014:VUB

[SB14] R. Sharma and R. Bhandari. On variance upper bounds for unimodal distributions. Communications in Statistics: Theory and Methods, 43(21):4503–4513, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salah:2016:SCR

[SB16a] Khardani Salah and Thiam Baba. Strong consistency result of a non parametric conditional mode estimator under random censorship for functional regressors. Communications in Statistics: Theory and Methods, 45(7):1863–1875, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.867997.

Shekhar:2016:IBD

[SB16b] Shashi Shekhar and Lalmohan Bhar. Incomplete block designs for asymmetric parallel line assays. Communications in Statistics: Theory and Methods, 45(8):2413–2425, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.881493.

Sabo:2017:OLA

[SB17] Roy T. Sabo and Ghalib Bello. Optimal and lead-in adaptive allocation for binary outcomes: a comparison of Bayesian methods. Communications in Statistics: Theory and Methods, 46(6): 2823–2836, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1053929.

Shafay:2017:BPO

[SBA17] A. R. Shafay, N. Balakrishnan, and Jafar Ahmadi. Bayesian prediction of order statistics with fixed and random sample sizes based on k-record values from Pareto distribution. Communications in Statistics: Theory and Methods, 46(2):721–735, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1004093.

Sharafi:2013:DFC

[SBK13] Maryam Sharafi, N. Balakrishnan, and Baha-Eldin Khaledi. Distribution-free comparison of hazard rates of two distributions under Type-II censoring. Communications in Statistics: Theory and Methods, 42(10):1889–1898, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saidane:2013:GLF

[SBL13] Mohamed Saidane, Xavier Bry, and Christian Lavergne. Generalized linear factor models: a new local EM estimation algorithm. Communications in Statistics: Theory and Methods, 42(16):2944–2958, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Suzuki:2017:GLT

[SBLC17] Adriano K. Suzuki, Gladys D. C. Barriga, Francisco Louzada, and Vicente G. Cancho. A general long-term aging model with different underlying activation mechanisms: Modeling, Bayesian estimation, and case influence diagnostics. Communications in Statistics: Theory and Methods, 46(6):3080–3098, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1053945.

Sekeh:2014:SRB

[SBR14] S. Yasaei Sekeh, G. R. Mohtashami Borzadaran, and A. H. Rezaei Roknabadi. Some results based on a version of the generalized dynamic entropies. Communications in Statistics: Theory and Methods, 43(14):2989–3006, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sastry:2016:ZDG

[SBRGD16] D. V. S. Sastry, Deepesh Bhati, R. N. Rattihalli, and E. Gómez-Déniz. On zero-distorted generalized geometric distribution. Communications in Statistics: Theory and Methods, 45(18): 5427–5442, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.942437.

Sabo:2010:EMA

[SC10] Roy T. Sabo and N. Rao Chaganty. Estimation methods for an autoregressive familial correlation structure. Communications in Statistics: Theory and Methods, 39(6):973–991, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Strydom:2018:MNE

[SC18] Nina Strydom and Nico Crowther. Multivariate normal estimation: the case (n > p). Communications in Statistics: Theory and Methods, 47(5):1071–1090, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1316405.

Sharma:2019:QBF

[SC19] Dreamlee Sharma and Tapan Kumar Chakrabarty. The quantile-based flattened logistic distribution: Some properties and applications. Communications in Statistics: Theory and Methods, 48(14):3643–3662, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481966.

Sheu:2010:GPP

[SCC10] Shey-Huei Sheu, Chin-Chih Chang, and Yen-Luan Chen. A generalized periodic preventive maintenance model with virtual age for a system subjected to shocks. Communications in Statistics: Theory and Methods, 39(13):2379–2393, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sheu:2011:ARP

[SCC11] Shey-Huei Sheu, Chin-Chih Chang, and Chaun-Hung Chiu. Age replacement policy with a safety constraint via the Bayesian method. Communications in Statistics: Theory and Methods, 40(23):4151–4164, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sheu:2012:ESI

[SCC12] Shey-Huei Sheu, Chin-Chih Chang, and Yen-Luan Chen. An extended sequential imperfect preventive maintenance model with improvement factors. Communications in Statistics: Theory and Methods, 41(7):1269–1283, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silveira:2010:GLN

[SCdSP10] Liciana V. A. Silveira, Enrico A. Colosimo, and José Raimundo de S. Passos. A generalized log-normal model for grouped survival data. Communications in Statistics: Theory and Methods, 39(15):2659–2666, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schott:2012:ATE

[Sch12] James R. Schott. An approximation for the test of the equality of the smallest eigenvalues of a covariance matrix. Communications in Statistics: Theory and Methods, 41(24):4439–4443, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schmiedt:2016:DAA

[Sch16] Anja B. Schmiedt. Domains of attraction of asymptotic distributions of extreme generalized order statistics. Communications in Statistics: Theory and Methods, 45(7):2089–2104, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.870206.

Schott:2017:STA

[Sch17] James R. Schott. Some tests for the allometric extension model in regression. Communications in Statistics: Theory and Methods, 46(12):6112–6118, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1118510.

Saez-Castillo:2015:TCS

[SCMR15] A. J. Sáez-Castillo and A. M. Martínez-Rodríguez. Two conditionally specified mixed binomial distributions. Communications in Statistics: Theory and Methods, 44(14):3056–3072, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:2009:ULD

[Sco09] W. F. Scott. The uniform-Laplace distribution and the Dtest. Communications in Statistics: Theory and Methods, 39 (2):249–255, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scollnik:2011:BIC

[Sco11] David P. M. Scollnik. Bayesian inference for a class of start-up demonstration tests with rejection of units upon the observation of d failures. Communications in Statistics: Theory and Methods, 40(14):2528–2538, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shafiee:2011:TDW

[SCSMN11] Mahmood Shafiee, Stefanka Chukova, Mohammad Saidi-Mehrabad, and Seyed Taghi Akhavan Niaki. Two-dimensional warranty cost analysis for second-hand products. Communications in Statistics: Theory and Methods, 40(4):684–701, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samawi:2017:KDE

[SCYR17] Hani Samawi, Arpita Chatterjee, JingJing Yin, and Haresh Rochani. On kernel density estimation based on different stratified sampling with optimal allocation. Communications in Statistics: Theory and Methods, 46(22):10973–10990, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1257714.

Samawi:2019:QEB

[SCYR19] Hani Samawi, Arpita Chatterjee, Jingjing Yin, and Haresh Rochani. On quantiles estimation based on different stratified sampling with optimal allocation. Communications in Statistics: Theory and Methods, 48(6):1529–1544, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1433856.

Sudhagoni:2012:MCP

[SD12] Ramu G. Sudhagoni and Gemechis D. Djira. Multiple comparisons of parametric models in longitudinal studies. Communications in Statistics: Theory and Methods, 41(4):729–740, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shafiei:2014:BSD

[SD14] S. Shafiei and M. Doostparast. Balakrishnan skew-t distribution and associated statistical characteristics. Communications in Statistics: Theory and Methods, 43(19):4109–4122, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Si:2012:JII

[SDC^{+}12] Shubin Si, Hongyan Dui, Zhiqiang Cai, Shudong Sun, and Yingfeng Zhang. Joint integrated importance measure for multi-state transition systems. Communications in Statistics: Theory and Methods, 41(21):3846–3862, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shafiei:2015:PVT

[SDJ15] Sobhan Shafiei, Mahdi Doostparast, and Ahad Jamalizadeh. Prediction of variables via their order statistics in bivariate elliptical distributions with application in the financial markets. Communications in Statistics: Theory and Methods, 44 (3):627–643, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2017:MBT

[SDS17] P. G. Sankaran, Isha Dewan, and E. P. Sreedevi. A martingalebased test for independence of time to failure and cause of failure for competing risks models. Communications in Statistics: Theory and Methods, 46(16):8178–8186, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1175631.

Deng:2014:NFP

[sDsWcCjK14] Wen shuenn Deng, Jyh shyang Wu, Li ching Chen, and Shun jie Ke. A note on the frequency polygon based on the weighted sums of binned data. Communications in Statistics: Theory and Methods, 43(8):1666–1685, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sosa-Escudero:2013:TPE

[SE13] Walter Sosa-Escudero. Testing for persistence in the error component model: A one-sided approach. Communications in Statistics: Theory and Methods, 42(14):2601–2616, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Santos:2015:IMQ

[SE15] Bruno R. Santos and Silvia N. Elian. Influence measures in quantile regression models. Communications in Statistics: Theory and Methods, 44(9):1842–1853, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:2017:TSH

[SE17] Ché L. Smith and Lloyd J. Edwards. A test of separate hypotheses for comparing linear mixed models with non nested fixed effects. Communications in Statistics: Theory and Methods, 46(11):5487–5500, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1104352.

Shams:2015:MIW

[SEA15] M. Shams, M. Emadi, and N. R. Arghami. Maximal invariant and weakly equivariant estimators. Communications in Statistics: Theory and Methods, 44(6):1293–1317, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sei:2013:JIG

[Sei13] Tomonari Sei. A Jacobian inequality for gradient maps on the sphere and its application to directional statistics. Communications in Statistics: Theory and Methods, 42(14):2525–2542, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Selvitella:2017:MAE

[Sel17a] Alessandro Selvitella. The Monge–Ampère equation in transformation theory and an application to 1/α-probabilities. Communications in Statistics: Theory and Methods, 46(4):2037–2054, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1040509.

Selvitella:2017:CFA

[Sel17b] Alessandro Selvitella. On 1/α-characteristic functions and applications to asymptotic statistical inference. Communications in Statistics: Theory and Methods, 46(4):1941–1958, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030427.

Seneta:2014:IMC

[Sen14] Eugene Seneta. Inhomogeneous Markov chains and ergodicity coefficients: John Hajnal (1924–2008). Communications in Statistics: Theory and Methods, 43(7):1296–1308, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sengupta:2016:AVE

[Sen16a] S. Sengupta. Admissibility of a variance estimator in finite population sampling under Warner’s randomized response plan with multiple responses. Communications in Statistics: Theory and Methods, 45(19):5745–5750, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.948208.

Sengupta:2016:AOS

[Sen16b] S. Sengupta. Admissible and optimal sampling strategy for estimating finite population mean in randomized response surveys with multiple responses. Communications in Statistics: Theory and Methods, 45(21):6223–6228, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.957859.

Sengupta:2017:CHT

[Sen17a] S. Sengupta. On the comparison of Horvitz–Thompson and Murthy’s sampling strategies for estimating finite population proportions in direct and randomized response surveys. Communications in Statistics: Theory and Methods, 46(3):1456– 1461, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1019149.

Sengupta:2017:CWU

[Sen17b] S. Sengupta. On the comparison of Warner’s and unrelated question randomized response plans for estimating sensitive finite population proportions. Communications in Statistics: Theory and Methods, 46(5):2370–2374, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1044669.

Sen:2018:LMU

[Sen18a] Amit Sen. Lagrange multiplier unit root test in the presence of a break in the innovation variance. Communications in Statistics: Theory and Methods, 47(7):1580–1596, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321771.

Sengupta:2018:AUE

[Sen18b] S. Sengupta. Admissible unbiased estimation of finite population variance under a randomized response model. Communications in Statistics: Theory and Methods, 47(20):5077–5082, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1386311.

Schaalje:2012:OSN

[SF12a] G. Bruce Schaalje and Paul J. Fields. Open-set nearest shrunken centroid classification. Communications in Statistics: Theory and Methods, 41(4):638–652, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schluter:2012:TQA

[SF12b] Stephan Schlüter and Matthias Fischer. A tail quantile approximation for the Student t distribution. Communications in Statistics: Theory and Methods, 41(15):2617–2625, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shafipour:2016:SAS

[SF16] Gholamreza Shafipour and Abdolvahhab Fetanat. Survival analysis in supply chains using statistical flowgraph models: Predicting time to supply chain disruption. Communications in Statistics: Theory and Methods, 45(21):6183–6208, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.957856.

Seif:2016:SAS

[SFSH16] Asghar Seif, Alireza Faraz, Erwin Saniga, and Cédric Heuchenne. A statistically adaptive sampling policy to the Hotelling’s T^{2} control chart: Markov chain approach. Communications in Statistics: Theory and Methods, 45(13):3919–3929, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.911910.

Shabbir:2010:EFP

[SG10a] Javid Shabbir and Sat Gupta. On estimating finite population mean in simple and stratified random sampling. Communications in Statistics: Theory and Methods, 40(2):199–212, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shabbir:2010:SEF

[SG10b] Javid Shabbir and Sat Gupta. Some estimators of finite population variance of stratified sample mean. Communications in Statistics: Theory and Methods, 39(16):3001–3008, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shakeri:2010:ESF

[SG10c] Nezhat Shakeri and Gurprit Grover. Estimation of survivorship function based on doubly censored data with application to HIV/AIDS. Communications in Statistics: Theory and Methods, 39(5):777–790, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szeto:2011:MPG

[SG11] K. Y. Szeto and Yunbo Guo. Multi-population genetic algorithm for locating multi-optima in noisy complex landscape. Communications in Statistics: Theory and Methods, 40(16): 3029–3048, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samani:2014:MCB

[SG14] E. Bahrami Samani and M. Ganjali. Mixed correlated bivariate ordinal and negative binomial longitudinal responses with nonignorable missing values. Communications in Statistics: Theory and Methods, 43(13):2659–2673, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:2015:RAE

[SG15] A. K. Md. Ehsanes Saleh and Amal F. Ghania. Ridge autoregression estimation: LS method. Communications in Statistics: Theory and Methods, 44(15):3303–3320, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:2016:NED

[SG16a] A. K. Md. Ehsanes Saleh and Amal Ghania. New estimators of distribution functions. Communications in Statistics: Theory and Methods, 45(11):3145–3157, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.897138.

Samani:2016:BRE

[SG16b] E. Bahrami Samani and M. Ganjali. A Bayesian random effects model for analyzing mixed negative binomial and continuous longitudinal responses. Communications in Statistics: Theory and Methods, 45(8):2354–2367, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.857865.

Schwartz:2016:BRM

[SG16c] Jacob Schwartz and David E. Giles. Bias-reduced maximum likelihood estimation of the zero-inflated Poisson distribution. Communications in Statistics: Theory and Methods, 45(2):465– 478, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sebastian:2016:TSM

[SG16d] Nicy Sebastian and Rudolf Gorenflo. Time series models associated with Mittag-Leffler type distributions and its properties. Communications in Statistics: Theory and Methods, 45(24): 7210–7225, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.978946.

Sun:2016:STH

[SG16e] Luna Sun and Alix I. Gitelman. Simultaneous tests for homogeneity of two zero-inflated (beta) populations. Communications in Statistics: Theory and Methods, 45(17):5137–5153, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.938828.

Shabbir:2017:EFP

[SG17a] Javid Shabbir and Sat Gupta. Estimation of finite population mean in simple and stratified random sampling using two auxiliary variables. Communications in Statistics: Theory and Methods, 46(20):10135–10148, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1231822.

Shabbir:2017:EPC

[SG17b] Javid Shabbir and Sat Gupta. Estimation of population coefficient of variation in simple and stratified random sampling under two-phase sampling scheme when using two auxiliary variables. Communications in Statistics: Theory and Methods, 46(16):8113–8133, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1175627.

Shabbir:2017:GEC

[SG17c] Javid Shabbir and Sat Gupta. On generalized exponential chain ratio estimators under stratified two-phase random sampling. Communications in Statistics: Theory and Methods, 46(6): 2910–2920, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1053936.

Shongwe:2017:LEA

[SG17d] S. C. Shongwe and M. A. Graham. A letter to the editor about “Machado, M. A. G. and Costa A. F. B. (2014), Some comments regarding the synthetic chart. Communications in Statistics — Theory and Methods, 43(14), 2897–2906”. Communications in Statistics: Theory and Methods, 46(21):10476–10480, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1236958. See [MC14].

Singh:2017:EEP

[SG17e] Housila P. Singh and Swarangi M. Gorey. Efficient estimation of population mean of sensitive variable in presence of scrambled responses. Communications in Statistics: Theory and Methods, 46(19):9557–9565, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1213288.

Singh:2017:ENR

[SG17f] Housila P. Singh and Swarangi M. Gorey. An efficient new randomized response model. Communications in Statistics: Theory and Methods, 46(19):9629–9635, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1213291.

Singh:2017:ERR

[SG17g] Housila P. Singh and Swarangi M. Gorey. An efficient randomized response model utilizing higher order moments ratios of scrambling variables. Communications in Statistics: Theory and Methods, 46(23):11712–11720, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1277755.

Singh:2017:NEU

[SG17h] Housila P. Singh and Swarangi M. Gorey. A new efficient unrelated randomized response model. Communications in Statistics: Theory and Methods, 46(24):12059–12074, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1291972.

Sadeghkhani:2018:NGB

[SG18a] Abdolnasser Sadeghkhani and Indranil Ghosh. A new generalized Balakrishnan type skewed-normal distribution: properties and associated inference. Communications in Statistics: Theory and Methods, 47(18):4483–4492, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1376090.

Sharma:2018:MAP

[SG18b] Kush Sharma and Davinder Kumar Garg. m-associate PBIB designs using Youden-m squares. Communications in Statistics: Theory and Methods, 47(7):1653–1662, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1324990.

Singh:2019:EPP

[SG19] Housila P. Singh and Swarangi M. Gorey. Estimation of population proportion of a qualitative character using randomized response technique in stratified random sampling. Communications in Statistics: Theory and Methods, 48(4):794–809, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1417436.

Shabbir:2019:GCE

[SGA19] Javid Shabbir, Sat Gupta, and Shakeel Ahmed. A generalized class of estimators under two-phase stratified sampling for non response. Communications in Statistics: Theory and Methods, 48(15):3761–3777, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481969.

Sarabia:2011:MPB

[SGD11] José María Sarabia and Emilio Gómez-Déniz. Multivariate Poisson–beta distributions with applications. Communications in Statistics: Theory and Methods, 40(6):1093–1108, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salmeron:2017:REE

[SGGdML17] Roman Salmeron, Catalina Garcia, Jose Garcia, and Maria del Mar Lopez. The raise estimator estimation, inference, and properties. Communications in Statistics: Theory and Methods, 46(13):6446–6462, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1125496.

Saba:2017:MCM

[SGKB17] R. Aliakbari Saba, Z. Rezaei Ghahroodi, K. Kiani, and D. M. Berridge. Mean and cell-mean imputation resulting in the use of a semicontinuous distribution and a mixture of semicontinuous distributions. Communications in Statistics: Theory and Methods, 46(23):11450–11465, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1267762.

Shalabh:2011:ERC

[SGM11] Shalabh, Gaurav Garg, and Neeraj Misra. Estimation of regression coefficients in a restricted measurement error model using instrumental variables. Communications in Statistics: Theory and Methods, 40(19–20):3614–3629, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Satpati:2012:CGC

[SGPA12] Subrata Kumar Satpati, V. K. Gupta, Rajender Parsad, and M. L. Aggarwal. Computer-generated change-over designs for correlated observations. Communications in Statistics: Theory and Methods, 41(20):3786–3798, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samuh:2012:UPT

[SGR^{+}12] Monjed H. Samuh, Leonardo Grilli, Carla Rampichini, Luigi Salmaso, and Nicola Lunardon. The use of permutation tests for variance components in linear mixed models. Communications in Statistics: Theory and Methods, 41(16–17):3020–3029, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/03610926.2011.587933.

Samawi:2011:DFR

[SH11] Hani M. Samawi and Amal Helu. Distribution-free runs test for conditional symmetry. Communications in Statistics: Theory and Methods, 40(15):2709–2718, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schilling:2017:LMP

[SH17a] Mark F. Schilling and Bret Holladay. Length minimization for Poisson confidence procedures. Communications in Statistics: Theory and Methods, 46(2):861–873, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1006782.

Shad:2017:TBC

[SH17b] Muhamamd Yousaf Shad and Ijaz Husain. Trade-off between cost and variance for a multi-objective compromise allocation in stratified random sampling. Communications in Statistics: Theory and Methods, 46(6):2655–2666, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1040507.

Shams:2010:CIP

[Sha10a] Sedigheh Shams. A convergent iterative procedure for constructing bivariate distributions. Communications in Statistics: Theory and Methods, 39(6):1026–1037, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shang:2010:MCP

[Sha10b] Junfeng Shang. A multiple comparison procedure based on a variant of the Schwarz information criterion in a mixed model. Communications in Statistics: Theory and Methods, 39(6): 1095–1109, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shang:2016:DIC

[Sha16] Junfeng Shang. A diagnostic of influential cases based on the information complexity criteria in generalized linear mixed models. Communications in Statistics: Theory and Methods, 45(13):3751–3760, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
911902.

Shafay:2017:BEP

[Sha17] A. R. Shafay. Bayesian estimation and prediction based on generalized Type-I hybrid censored sample. Communications in Statistics: Theory and Methods, 46(10):4870–4887, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1089292.

Shabbir:2018:EUT

[Sha18a] Javid Shabbir. Efficient utilization of two auxiliary variables in stratified double sampling. Communications in Statistics: Theory and Methods, 47(1):92–101, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1300280.

Shakhatreh:2018:NTP

[Sha18b] Mohammed K. Shakhatreh. A new three-parameter extension of the log-logistic distribution with applications to survival data. Communications in Statistics: Theory and Methods, 47(21):5205–5226, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1388399.

Sharma:2018:BAH

[Sha18c] Vikas Kumar Sharma. Bayesian analysis of head and neck cancer data using generalized inverse Lindley stress–strength reliability model. Communications in Statistics: Theory and Methods, 47(5):1155–1180, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1316858.

Shahrastani:2019:EBH

[Sha19] Shahram Yaghoobzadeh Shahrastani. Estimating E-Bayesian and hierarchical Bayesian of scalar parameter of Gompertz distribution under type II censoring schemes based on fuzzy data. Communications in Statistics: Theory and Methods, 48(4):831– 840, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1417438.

Shafqat:2018:ACC

[SHANA18] Ambreen Shafqat, Jaffer Hussain, Amjad D. Al-Nasser, and Muhammad Aslam. Attribute control chart for some popular distributions. Communications in Statistics: Theory and Methods, 47(8):1978–1988, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1335414.

Salehi:2017:MGR

[SHB17] E. Salehi and S. S. Hashemi-Bosra. The mean general residual lifetime of (n − k + 1)-out-of-n systems with exchangeable components. Communications in Statistics: Theory and Methods, 46(13):6382–6400, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1129413.

Shen:2010:NEB

[She10a] Pao-Sheng Shen. Nonparametric estimators of the bivariate survival function under left truncation and right censoring. Communications in Statistics: Theory and Methods, 39(16): 2877–2889, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2010:SAD

[She10b] Pao-Sheng Shen. Semiparametric analysis of doubly truncated data. Communications in Statistics: Theory and Methods, 39 (17):3178–3190, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2012:GTD

[She12a] Pao-Sheng Shen. A generalized time-dependent conditional linear model with left-truncated and right-censored data. Communications in Statistics: Theory and Methods, 41(1):128–137, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2012:NED

[She12b] Pao-Sheng Shen. Nonparametric estimators of the distribution function for one modified model of current status data. Communications in Statistics: Theory and Methods, 41(22):4096–4106, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2013:NEB

[She13] Pao-Sheng Shen. Nonparametric estimation of the bivariate distribution function with doubly truncated data. Communications in Statistics: Theory and Methods, 42(21):3805–3818, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shemehsavar:2014:BGM

[She14a] Soudabeh Shemehsavar. A bivariate gamma model for a latent degradation process. Communications in Statistics: Theory and Methods, 43(9):1924–1938, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2014:GTE

[She14b] Pao-Sheng Shen. A generalization of Turnbull’s estimator for interval-censored and doubly truncated data. Communications in Statistics: Theory and Methods, 43(14):2958–2972, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2015:ATM

[She15a] Pao-Sheng Shen. Additive transformation models for multivariate interval-censored data. Communications in Statistics: Theory and Methods, 44(5):1065–1079, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2015:EML

[She15b] Pao-Sheng Shen. Estimation of multiple linear regression model with twice-censored data. Communications in Statistics: Theory and Methods, 44(21):4631–4640, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2015:FTM

[She15c] Pao-Sheng Shen. On fitting transformation model to survey data. Communications in Statistics: Theory and Methods, 44 (18):3796–3811, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2016:GFT

[She16] Pao-Sheng Shen. On goodness-of-fit tests for Aalen’s additive model based on left-truncated right-censored or doubly censored data. Communications in Statistics: Theory and Methods, 45(16):4812–4823, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
930908.

Shen:2017:LTM

[She17] Pao-Sheng Shen. Linear transformation models for survival analysis with tumor growth information in cancer screening study. Communications in Statistics: Theory and Methods, 46(4):1916–1926, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1030425.

Sheena:2018:AER

[She18a] Yo Sheena. Asymptotic expansion of the risk of maximum likelihood estimator with respect to α-divergence. Communications in Statistics: Theory and Methods, 47(16):4059–4087, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1380828.

Shen:2018:NPA

[She18b] Pao-Sheng Shen. Non parametric analysis of gap times for bivariate recurrent event data with cure fraction. Communications in Statistics: Theory and Methods, 47(1):181–195, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1301473.

Sheu:2013:MCS

[SHH13] Shey-Huei Sheu, Chi-Jui Huang, and Tsung-Shin Hsu. Maximum chi-square generally weighted moving average control chart for monitoring process mean and variability. Communications in Statistics: Theory and Methods, 42(23):4323–4341, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2009:CBI

[SHHH09] Ruiguang Song, Kathleen McDavid Harrison, Debra L. Hanson, and H. Irene Hall. Correction of bias in imputing missing values of categorical variables. Communications in Statistics: Theory and Methods, 39(2):350–362, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shiraishi:2010:MCB

[Shi10] Taka-Aki Shiraishi. Multiple comparisons based on Studentized M-statistics in a randomized block design. Communications in Statistics: Theory and Methods, 39(8–9):1563–1573, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Son:2010:CTP

[SHLK10] Chang-Kyoon Son, Ki-Hak Hong, Gi-Sung Lee, and Jong-Min Kim. The calibration for two-phase randomized response estimator. Communications in Statistics: Theory and Methods, 39(6):992–1012, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2017:PLD

[SHLQ17] Lixin Song, Zhiqiang Hua, Dawei Lu, and Xiaomeng Qi. Precise large deviations for the difference of two sums of WUOD and non identically distributed random variables with dominatedly varying tails. Communications in Statistics: Theory and Methods, 46(4):2013–2028, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1032424.

Singh:2016:USAa

[SHM16] G. N. Singh, F. Homa, and S. Maurya. On the use of several auxiliary variables in estimation of current population mean in successive sampling. Communications in Statistics: Theory and Methods, 45(11):3114–3125, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.895838.

Shore:2015:GMR

[Sho15] Haim Shore. A general model of random variation. Communications in Statistics: Theory and Methods, 44(9):1819–1841, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shushi:2018:GEF

[Shu18] Tomer Shushi. The generalized exponential family of distributions and its characteristics. Communications in Statistics: Theory and Methods, 47(10):2520–2526, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1342833.

Shushi:2019:EFC

[Shu19] Tomer Shushi. Explicit formulas for the cumulants and the vector-valued odd moments of the multivariate linearly skewed elliptical distributions. Communications in Statistics: Theory and Methods, 48(12):3085–3091, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1473605.

Swishchuk:2013:NDP

[SI13] Anatoliy Swishchuk and Md Shafiqul Islam. Normal deviation and Poisson approximation of a security market by the geometric Markov renewal processes. Communications in Statistics: Theory and Methods, 42(8):1488–1501, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shinozaki:2017:VSM

[SI17] N. Shinozaki and T. Iida. A variable selection method for detecting abnormality based on the T^{2} test. Communications in Statistics: Theory and Methods, 46(17):8603–8617, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1185120.

Szymkowiak:2019:SRA

[SI19] Magdalena Szymkowiak and Maria Iwińska. Some results about bivariate discrete distributions through the vector of aging intensities. Communications in Statistics: Theory and Methods, 48(9):2175–2184, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1459714.

Silva:2018:CBF

[Sil18] Ivair R. Silva. On the correspondence between frequentist and Bayesian tests. Communications in Statistics: Theory and Methods, 47(14):3477–3487, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1359296.

Siman:2012:KA

[Sim12] Miroslav Siman. On Kendall’s autocorrelations. Communications in Statistics: Theory and Methods, 41(10):1733–1738, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Siman:2014:MPC

[Sim14a] Miroslav Siman. Multivariate process capability indices: a directional approach. Communications in Statistics: Theory and Methods, 43(9):1949–1955, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Siman:2014:PIM

[Sim14b] Miroslav Siman. Precision index in the multivariate context. Communications in Statistics: Theory and Methods, 43(2):377– 387, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2016:NSC

[Sin16] Parminder Singh. A note on successive comparisons between several ordered variances. Communications in Statistics: Theory and Methods, 45(11):3264–3273, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901371.

Siray:2014:MRC

[Šir14] Gülesen Üstündağ Širay. Modified and restricted r-k class estimators. Communications in Statistics: Theory and Methods, 43(24):5130–5155, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Siray:2015:CEU

[Šir15] Gülesen Üstündağ Širay. r-d class estimator under misspecification. Communications in Statistics: Theory and Methods, 44(22):4742–4756, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sivaganesan:2010:OBT

[SJ10] Siva Sivaganesan and Dongming Jiang. Objective Bayesian testing of a Poisson mean. Communications in Statistics: Theory and Methods, 39(11):1887–1897, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2019:BDT

[SJ19] Xiaohui Shen and Long Jiang. BSDEs driven by time-changed lévy noises with non-Lipschitz generators. Communications in Statistics: Theory and Methods, 48(3):523–536, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1414263.

Sun:2019:CPE

[SJJ19] Yuqin Sun, Bo Jiang, and Hong Jiang. Computations of predictors/estimators under a linear random-effects model with parameter restrictions. Communications in Statistics: Theory and Methods, 48(14):3482–3497, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1476714.

Sharma:2015:SPS

[SJS15] Suresh Kumar Sharma, Kanchan Jain, and Neetu Singla. A selection procedure for selecting the least increasing failure rate average distribution. Communications in Statistics: Theory and Methods, 44(9):1786–1796, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarabia:2017:LCL

[SJT17] José María Sarabia, Vanesa Jordá, and Carmen Trueba. The Lamé class of Lorenz curves. Communications in Statistics: Theory and Methods, 46(11):5311–5326, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.775306.

Shi:2017:CSD

[SJY17] Zhiyan Shi, Jinli Ji, and Weiguo Yang. A class of small deviation theorem for the sequences of countable state random variables with respect to homogeneous Markov chains. Communications in Statistics: Theory and Methods, 46(14):6823–6830, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1137594.

Singh:2010:SIM

[SK10a] G. N. Singh and Jaishree Prabha Karna. Some imputation methods to minimize the effect of non response in two-occasion rotation patterns. Communications in Statistics: Theory and Methods, 39(18):3264–3281, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Suzuki:2010:PSC

[SK10b] Taiji Suzuki and Fumiyasu Komaki. On prior selection and covariate shift of β-Bayesian prediction under α-divergence risk.

Communications in Statistics: Theory and Methods, 39(8–9): 1655–1673, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2011:UST

[sK11] Feng shou Ko. Using the score test to identify the longitudinal biomarker considering accelerate failure time model with the frailty in survival analysis. Communications in Statistics: Theory and Methods, 40(21):3824–3838, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2012:AAS

[sK12a] Feng shou Ko. Approaches to allocate sample size rationally into individual regions for a multi-regional trial under heterogeneous effect size. Communications in Statistics: Theory and Methods, 41(20):3648–3665, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2012:USM

[SK12b] G. N. Singh and Jaishree Prabha Karna. On the use of superpopulation model in estimation of population mean in twooccasion rotation patterns. Communications in Statistics: Theory and Methods, 41(4):619–637, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singull:2012:DMQ

[SK12c] Martin Singull and Timo Koski. On the distribution of matrix quadratic forms. Communications in Statistics: Theory and Methods, 41(18):3403–3415, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2012:BAR

[SK12d] Joon Jin Song and Jong-Min Kim. Bayesian analysis of randomized response sum score variables. Communications in Statistics: Theory and Methods, 41(10):1875–1884, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shabbir:2013:EFP

[SK13a] Javid Shabbir and Nasir Saeed Khan. On estimating the finite population mean using two auxiliary variables in two phase sampling in the presence of non response. Communications in Statistics: Theory and Methods, 42(22):4127–4145, 2013.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:2013:DSS

[sK13b] Feng shou Ko. A design for survival studies: a multi-regional trial with unrecognized heterogeneity. Communications in Statistics: Theory and Methods, 42(13):2301–2323, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2014:EPM

[SK14] G. N. Singh and J. P. Karna. Estimation of population mean under non response in two-occasion rotation patterns. Communications in Statistics: Theory and Methods, 43(16):3429–3442, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shoaee:2015:SSR

[SK15a] Shirin Shoaee and Esmaile Khorram. Stress-strength reliability of a two-parameter bathtub-shaped lifetime distribution based on progressively censored samples. Communications in Statistics: Theory and Methods, 44(24):5306–5328, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Son:2015:CER

[SK15b] Chang-Kyoon Son and Jong-Min Kim. Calibration estimator of a rare sensitive attribute with Poisson distribution. Communications in Statistics: Theory and Methods, 44(4):855–871, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saidi:2016:CCD

[SK16a] Ahmed Ait Saidi and Mecheri Kheira. The conditional cumulative distribution function in single functional index model. Communications in Statistics: Theory and Methods, 45(16): 4896–4911, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.932808.

Seneta:2016:UDL

[SK16b] Eugene Seneta and Simon Ku. Unique decomposition of low-order time series. Communications in Statistics: Theory and Methods, 45(11):3357–3366, 2016. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
904349.

Ko:2016:CAS

[sK16c] Feng shou Ko. Comparisons of allocating sample size rationally into individual regions under heterogeneous effect size in a multiregional trial by a fixed effect model and a random effect model. Communications in Statistics: Theory and Methods, 45(23):7060–7074, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.974823.

Shukla:2016:IMC

[SK16d] Diwakar Shukla and Uttam Kumar Khedlekar. Inventory model for convertible item with deterioration. Communications in Statistics: Theory and Methods, 45(4):1137–1147, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.859703.

Srinivas:2016:MUE

[SK16e] V. Srinivas and B. K. Kale. ML and UMVU estimation in the M/D/1 queuing system. Communications in Statistics: Theory and Methods, 45(19):5826–5834, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.950750.

Seo:2017:BIE

[SK17] Jung In Seo and Yongku Kim. Bayesian inference on extreme value distribution using upper record values. Communications in Statistics: Theory and Methods, 46(15):7751–7768, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1165848.

Seo:2018:RBA

[SK18a] Jung In Seo and Yongku Kim. Robust Bayesian analysis for exponential parameters under generalized type-II progressive hybrid censoring. Communications in Statistics: Theory and Methods, 47(9):2259–2277, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1339086.

Singh:2018:EES

[SK18b] G. N. Singh and M. Khalid. An effective estimation strategy for population mean under random non-response situations in two-phase successive sampling. Communications in Statistics: Theory and Methods, 47(19):4641–4660, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1444179.

Sankaran:2019:CDQ

[SK19a] P. G. Sankaran and M. Dileep Kumar. A class of distributions with the quadratic mean residual quantile function. Communications in Statistics: Theory and Methods, 48(19):4936–4957, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1504967.

Singh:2019:WRE

[SK19b] Shivangi Singh and Chanchal Kundu. On weighted Renyi’s entropy for double-truncated distribution. Communications in Statistics: Theory and Methods, 48(10):2562–2579, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1472779.

Singh:2012:NTH

[SKG12] Parminder Singh, Narendra Kumar, and Amar Nath Gill. A note on testing homogeneity of several exponential location parameters. Communications in Statistics: Theory and Methods, 41(3):474–484, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:2018:SES

[SKG18] A. K. Md. Ehsanes Saleh, B. M. Golam Kibria, and Florence George. Simultaneous estimation of several CDF’s: homogeneity constraint. Communications in Statistics: Theory and Methods, 47(12):2813–2826, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1129419.

Son:2010:CRR

[SKHL10] Chang-Kyoon Son, Jong-Min Kim, Ki-Hak Hong, and Gi-Sung Lee. Calibration for randomized response estimators. Communications in Statistics: Theory and Methods, 39(17):3163–3177, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2010:IEF

[SKK10] Housila P. Singh, Sunil Kumar, and Marcin Kozak. Improved estimation of finite-population mean using sub-sampling to deal with non response in two-phase sampling scheme. Communications in Statistics: Theory and Methods, 39(5):791–802, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smaili:2016:FPH

[SKK16] Khaled Smaili, Therrar Kadri, and Seifedine Kadry. Finding the PDF of the hypoexponential random variable using the Kad matrix similar to the general Vandermonde matrix. Communications in Statistics: Theory and Methods, 45(5):1542–1549, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.930910.

Safe:2018:MCA

[SKK18] H. Safe, R. B. Kazemzadeh, and Y. Gholipour Kanani. A Markov chain approach for double-objective economic statistical design of the variable sampling interval X̄ control chart. Communications in Statistics: Theory and Methods, 47(2):277– 288, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1235197.

Singh:2016:EUN

[SKM16a] G. N. Singh, M. Khetan, and S. Maurya. Estimation under non response when it occurs on both occasions in twooccasion successive sampling. Communications in Statistics: Theory and Methods, 45(13):3939–3951, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.911913.

Singh:2016:UMV

[SKM16b] G. N. Singh, M. Khetan, and S. Maurya. Use of multiauxiliary variables to cope with non response in estimation of current population mean in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 45(16):4840–4857, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.930914.

Singh:2017:UMA

[SKM17] G. N. Singh, M. Khetan, and S. Maurya. On the use of multi-auxiliary variables to neutralize the effect of non response in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 46(7):3501–3519, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1065331.

Singh:2017:SECb

[SKS17] G. N. Singh, M. Khalid, and A. K. Sharma. Some efficient classes of estimators of population mean in two-phase successive sampling under random non response. Communications in Statistics: Theory and Methods, 46(24):12194–12209, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1291973.

Singh:2016:FEP

[SKT16] Housila P. Singh, Jong M. Kim, and Tanveer A. Tarray. A family of estimators of population variance in two-occasion rotation patterns. Communications in Statistics: Theory and Methods, 45(14):4106–4116, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.915047.

Sheu:2013:GWM

[SL13] Shey-Huei Sheu and ShinLi Lu. The generally weighted moving average variance chart. Communications in Statistics: Theory and Methods, 42(17):3204–3214, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2014:TMS

[SL14] Yanfeng Shen and Zhengyan Lin. Tests for a multiple-sample problem in high dimensions. Communications in Statistics: Theory and Methods, 43(2):291–305, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2015:DVE

[SL15] Huihui Sun and Jinguan Lin. Diagnostics of variance of the error in mixed effects linear models based on M-estimation. Communications in Statistics: Theory and Methods, 44(9):1779– 1785, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvestrov:2016:SAM

[SL16] Dmitrii Silvestrov and Yanxiong Li. Stochastic approximation methods for American type options. Communications in Statistics: Theory and Methods, 45(6):1607–1631, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.915046.

Singh:2018:NSO

[SL18] Alok Kumar Singh and S. Lalitha. A novel spatial outlier detection technique. Communications in Statistics: Theory and Methods, 47(1):247–257, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1301477.

Slaoui:2017:RKD

[Sla17] Yousri Slaoui. Recursive kernel density estimators under missing data. Communications in Statistics: Theory and Methods, 46(18):9101–9125, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1205618.

Shen:2018:PEL

[SLF18] Yu Shen, Han-Ying Liang, and Guo-Liang Fan. Penalized empirical likelihood for quantile regression with missing covariates and auxiliary information. Communications in Statistics: Theory and Methods, 47(8):2001–2021, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1335413.

Sljivic:2017:CEM

[Slj17] Nusa Mikuljan Sljivić. Cross-entropy method for estimation of posterior expectation in Bayesian VAR models. Communications in Statistics: Theory and Methods, 46(23):11933–11947, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1288252.

Song:2014:UCC

[SLL14] Lixin Song, Dawei Lu, and Fang Liu. Using convex combination of copulas and EM algorithm for credibility. Communications in Statistics: Theory and Methods, 43(6):1222–1233, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:2017:OPR

[SLL17] William F. Scott, Sheng Lu, and Cheuk Ngai Lo. Orthogonal polynomials, repeated measures, and SPSS. Communications in Statistics: Theory and Methods, 46(7):3342–3364, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1060345.

Saggou:2017:PMM

[SLOT17] Hafida Saggou, Tassadit Lachemot, and Megdouda Ourbih-Tari. Performance measures of M/G/1 retrial queues with recurrent customers, breakdowns, and general delays. Communications in Statistics: Theory and Methods, 46(16):7998–8015, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1171352.

Song:2018:VRS

[SLSC18] Fengli Song, Peng Lai, Baohua Shen, and Guosheng Cheng. Variance ratio screening for ultrahigh dimensional discriminant analysis. Communications in Statistics: Theory and Methods, 47(24):6034–6051, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1406113.

Shih:2019:FCR

[SLSE19] Jia-Han Shih, Wei Lee, Li-Hsien Sun, and Takeshi Emura. Fitting competing risks data to bivariate Pareto models. Communications in Statistics: Theory and Methods, 48(5):1193–1220, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1425450.

Song:2014:SCB

[SLSY14] Q. Song, R. Liu, Q. Shao, and L. Yang. A simultaneous confidence band for dense longitudinal regression. Communications in Statistics: Theory and Methods, 43(24):5195–5210, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2011:PIR

[SLZ11] Yanfeng Shen, Zhengyan Lin, and Jun Zhu. Penalized independence rule for testing high-dimensional hypotheses. Communications in Statistics: Theory and Methods, 40(13):2424–2435, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2018:RPC

[SLZ18] Jingyun Sun, Yongjun Li, and Ling Zhang. Robust portfolio choice for a defined contribution pension plan with stochastic income and interest rate. Communications in Statistics: Theory and Methods, 47(17):4106–4130, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1367815.

Schwarz:2009:LWP

[SM09] Wolf Schwarz and Jeff O. Miller. Locking the Wiener process to its level-crossing time. Communications in Statistics: Theory and Methods, 39(2):372–381, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Skiadas:2013:P

[SM13] Christos H. Skiadas and Sally I. McClean. Preface. Communications in Statistics: Theory and Methods, 42(16):2851–2852, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sultan:2014:MOS

[SM14] Khalaf S. Sultan and Mohamed E. Moshref. Moments of order statistics from Weibull distribution in the presence of multiple outliers. Communications in Statistics: Theory and Methods, 43(10–12):2214–2226, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2016:IES

[SM16] Housila P. Singh and Vishal Mehta. Improved estimation of scale parameters of Morgenstern type bivariate uniform distribution using ranked set sampling. Communications in Statistics: Theory and Methods, 45(5):1466–1476, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.864767.

Sellers:2017:UMM

[SM17] Kimberly F. Sellers and Darcy S. Morris. Underdispersion models: Models that are “under the radar”. Communications in Statistics: Theory and Methods, 46(24):12075–12086, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1291976.

Shiraishi:2018:CTP

[SM18] Taka-Aki Shiraishi and Shin-Ichi Matsuda. Closed testing procedures for all pairwise comparisons in a randomized block design. Communications in Statistics: Theory and Methods, 47(15):3571–3587, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1359302.

Smith:2013:QPP

[Smi13] Andrew D. A. C. Smith. Quadratic programming and penalized regression. Communications in Statistics: Theory and Methods, 42(7):1363–1372, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:2019:HPM

[Smi19] Scott E. Smith. The hazard-product method of generalization: Application to the Gompertz and exponentiated halfnormal distributions. Communications in Statistics: Theory and Methods, 48(5):1177–1192, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1425449.

Singh:2017:SECa

[SMK17] G. N. Singh, S. Maurya, and M. Khetan. Some effective combinations of available information under non response in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 46(1):280–295, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.990105.

Shu:2010:IRF

[SMM10] Hong Shu, Xinsheng Ma, and Chunsheng Ma. Isotropic random fields with the von Kármán–Whittle type correlation structure. Communications in Statistics: Theory and Methods, 39(19): 3504–3512, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2015:SER

[SMMS15] G. N. Singh, D. Majhi, S. Maurya, and A. K. Sharma. Some effective rotation patterns in estimation of population mean in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 44(12):2571–2585, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarvate:2018:GDD

[SMN18] Dinesh Sarvate, Nutan Mishra, and Kasifa Namyalo. Group divisible designs with block size five from Clatworthy’s table. Communications in Statistics: Theory and Methods, 47(9): 2085–2097, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1179763.

Singh:2018:IEP

[SMP18] Housila P. Singh, Abhishek C. Mishra, and Surya K. Pal. Improved estimator of population total in PPS sampling. Communications in Statistics: Theory and Methods, 47(4):912–934, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1313987.

Singh:2015:ERP

[SMPH15] G. N. Singh, D. Majhi, S. Prasad, and F. Homa. Effective rotation patterns under non response in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 44(6):1182–1195, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvestrov:2014:CAM

[SMS14] Dmitrii Silvestrov, Raimondo Manca, and Evelina Silvestrova. Computational algorithms for moments of accumulated Markov and semi-Markov rewards. Communications in Statistics: Theory and Methods, 43(7):1453–1469, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saadatmelli:2019:EDC

[SMSF19] Aitin Saadatmelli, M. Bameni Moghadam, Asghar Seif, and Alireza Faraz. Economic design of X̄ control charts with multiple assignable causes under Burr XII shock model. Communications in Statistics: Theory and Methods, 48(3):500–522, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1414262.

Schaffrin:2012:RLE

[SN12] Burkhard Schaffrin and Gerhard Navratil. On reproducing linear estimators within the Gauss–Markov model with stochastic constraints. Communications in Statistics: Theory and Methods, 41(13–14):2570–2587, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soleimani:2014:MMS

[SN14] Paria Soleimani and R. Noorossana. Monitoring multivariate simple linear profiles in the presence of between profile autocorrelation. Communications in Statistics: Theory and Methods, 43(3):530–546, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Simcharoen:2016:NUB

[SN16] W. Simcharoen and K. Neammanee. Non uniform bound on a combinatorial central limit theorem. Communications in Statistics: Theory and Methods, 45(18):5517–5532, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.944828.

Saber:2017:CSI

[SN17] M. M. Saber and A. R. Nematollahi. Comparison of spatial interpolation methods in the first order stationary multiplicative spatial autoregressive models. Communications in Statistics: Theory and Methods, 46(18):9230–9246, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1205619.

Siripraparat:2018:NUB

[SN18] Tatpon Siripraparat and Kritsana Neammanee. A non uniform bound for half-normal approximation of the number of returns to the origin of symmetric simple random walk. Communications in Statistics: Theory and Methods, 47(1):42–54, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1300286.

Sayyad:2017:NPA

[SNAN17] Ali Sayyad, Seyed Taghi Akhavan Niaki, and Behrouz Afshar-Najafi. A non parametric approach to monitor simple linear profiles in phases I and II. Communications in Statistics: Theory and Methods, 46(11):5203–5222, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1099668.

Sohi:2012:NSI

[SNN12] Zahra Asadollahi Sohi, Rassoul Noorossana, and Seyed Taghi Akhavan Niaki. New statistic to increase correctness in simulation factor screening using frequency domain method. Communications in Statistics: Theory and Methods, 41(12): 2242–2255, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutar:2016:MLS

[SNN16] Santosh S. Sutar and U. V. Naik-Nimbalkar. A model for k-out-of-m load-sharing systems. Communications in Statistics: Theory and Methods, 45(20):5946–5960, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.953691.

Sumritnorrapong:2019:INU

[SNS19] Patcharee Sumritnorrapong, Kritsana Neammanee, and Jiraphan Suntornchost. An improvement of a non-uniform bound for combinatorial central limit theorem. Communications in Statistics: Theory and Methods, 48(9):2129–2146, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1459711.

Saadatmand:2017:EMV

[SNSA17] A. Saadatmand, A. R. Nematollahi, and S. M. Sadooghi-Alvandi. On the estimation of missing values in AR(1) model with exponential innovations. Communications in Statistics: Theory and Methods, 46(7):3393–3400, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1060347.

Shen:2019:ANL

[SNW19] Aiting Shen, Mingming Ning, and Caoqing Wu. The asymptotic normality of the linear weighted estimator in nonparametric regression models. Communications in Statistics: Theory and Methods, 48(6):1367–1376, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1429633.

Seck:2017:AIP

[SNWL17] Cheikh Tidiane Seck, Joseph Ngatchou-Wandji, and Gane Samb Lô. Asymptotic inference in poverty indices: an empirical processes approach. Communications in Statistics: Theory and Methods, 46(12):6192–6212, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1122060.

Stockhammar:2012:SHR

[SÖ12] Pär Stockhammar and Lars-Erik Öller. A simple heteroscedasticity removing filter. Communications in Statistics: Theory and Methods, 41(2):281–299, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sengupta:2014:UAC

[SO14a] Ashis Sengupta and S. H. Ong. A unified approach for construction of probability models for bivariate linear and directional data. Communications in Statistics: Theory and Methods, 43 (10–12):2563–2569, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:2014:UDD

[SO14b] Engin A. Sungur and Jessica M. Orth. Understanding directional dependence through angular correlation. Communications in Statistics: Theory and Methods, 43(19):4143–4155, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:2017:NDM

[SO17] Engin A. Sungur and Jessica M. Orth. A note on dependence modeling for Bernoulli variables. Communications in Statistics: Theory and Methods, 46(16):8217–8229, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1177080.

Song:2016:NPD

[Son16] Qiongxia Song. Non parametric derivative estimation with confidence bands. Communications in Statistics: Theory and Methods, 45(2):277–290, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soulakova:2017:GCI

[Sou17] Julia N. Soulakova. Generalized confidence intervals compatible with the Min test for simultaneous comparisons of one subpopulation to several other subpopulations. Communications in Statistics: Theory and Methods, 46(19):9441–9449, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1212072.

Shitan:2011:TSP

[SP11a] Mahendran Shitan and Shelton Peiris. Time series properties of the class of generalized first-order autoregressive processes with moving average errors. Communications in Statistics: Theory and Methods, 40(13):2259–2275, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shuran:2011:WBC

[SP11b] Zhao Shuran and Ren Peimin. Weighted bootstrap confidence intervals for generalized Behrens–Fisher problems. Communications in Statistics: Theory and Methods, 40(21):3780–3788, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shitan:2013:AAV

[SP13] Mahendran Shitan and Shelton Peiris. Approximate asymptotic variance-covariance matrix for the Whittle estimators of GAR(1) parameters. Communications in Statistics: Theory and Methods, 42(5):756–770, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Speevak:2015:NRC

[SP15] Ted Speevak and Takis Papaioannou. A note on rank correlation inequalities with missing data. Communications in Statistics: Theory and Methods, 44(7):1381–1386, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saboor:2016:MOG

[SP16a] Abdus Saboor and Tibor K. Pogány. Marshall–Olkin gamma-Weibull distribution with applications. Communications in Statistics: Theory and Methods, 45(5):1550–1563, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.953694.

Singh:2016:SEP

[SP16b] G. N. Singh and S. Prasad. Some estimation procedures using a linear model in successive sampling. Communications in Statistics: Theory and Methods, 45(9):2679–2698, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.887116.

Singh:2016:EER

[SP16c] Housila P. Singh and Surya K. Pal. An efficient effective rotation pattern in successive sampling over two occasions. Communications in Statistics: Theory and Methods, 45(17):5017–5027, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.935431.

Singh:2016:USAb

[SP16d] Housila P. Singh and Surya K. Pal. Use of several auxiliary variables in estimating the population mean in a two-occasion rotation pattern. Communications in Statistics: Theory and Methods, 45(23):6928–6942, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.972570.

Singh:2017:CET

[SP17a] Housila P. Singh and Surya K. Pal. A class of exponential-type ratio estimators of a general parameter. Communications in Statistics: Theory and Methods, 46(8):3957–3984, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1076477.

Singh:2017:IVE

[SP17b] Housila P. Singh and Surya K. Pal. Improvement over variance estimation using auxiliary information in sample surveys. Communications in Statistics: Theory and Methods, 46(15): 7732–7750, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1161799.

Singh:2017:SGR

[SP17c] Housila P. Singh and Surya K. Pal. Search of good rotation patterns using exponential method of estimation in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 46(11):5466–5486, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1104351.

Spanos:2019:NCL

[Spa19] Aris Spanos. Near-collinearity in linear regression revisited: The numerical vs. the statistical perspective. Communications in Statistics: Theory and Methods, 48(22):5492–5516, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1513147.

Speevak:2017:ISS

[Spe17] Ted Speevak. Inequalities for sums of squares of reranked differences involving ties and missing data. Communications in Statistics: Theory and Methods, 46(17):8419–8429, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1179762.

Singh:2017:NCE

[SPS17] Housila P. Singh, Surya Kant Pal, and Ramkrishna S. Solanki. A new class of estimators of finite population mean in sample surveys. Communications in Statistics: Theory and Methods, 46(6):2630–2637, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1030429.

Singh:2018:SCR

[SPY18] Housila P. Singh, Surya K. Pal, and Anita Yadav. A study on the chain ratio-ratio-type exponential estimator for finite population variance. Communications in Statistics: Theory and Methods, 47(6):1442–1458, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321124.

Shi:2019:ELE

[SQCL19] Jianhua Shi, Zhiping Qiu, Xiaoping Chen, and Haiqing Lin. The empirical likelihood estimation of the quantile function under the multiplicative intercept risk model. Communications in Statistics: Theory and Methods, 48(13):3377–3387, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1476709.

Soulakova:2011:AAA

[SR11] Julia N. Soulakova and Ananya Roy. Application of Anbar’s approach to hypothesis testing to detect the difference between two proportions. Communications in Statistics: Theory and Methods, 40(10):1866–1878, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shang:2017:NPB

[SR17] Kan Shang and Cavan Reilly. Non parametric Bayesian analysis of the two-sample problem with censoring. Communications in Statistics: Theory and Methods, 46(24):12008–12022, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1288249.

Sunoj:2018:QBS

[SR18] S. M. Sunoj and R. S. Rasin. A quantile-based study on ageing intensity function. Communications in Statistics: Theory and Methods, 47(22):5474–5484, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1395049.

Sreelakshmi:2018:ICB

[Sre18] N. Sreelakshmi. An introduction to copula-based bivariate reliability concepts. Communications in Statistics: Theory and Methods, 47(4):996–1012, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1316396.

Srinivas:2011:EMM

[SRK11] V. Srinivas, S. Subba Rao, and B. K. Kale. Estimation of measures in M/M/1 queue. Communications in Statistics: Theory and Methods, 40(18):3327–3336, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2017:MDR

[SRN17] P. G. Sankaran, Nidhi P. Ramesh, and N. Unnikrishnan Nair. Multivariate discrete reversed hazard rates. Communications in Statistics: Theory and Methods, 46(19):9817–9833, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1222430.

Shukla:2014:EEP

[SRT14] D. Shukla, Yashwant Singh Rajput, and Narendra Singh Thakur. Edge estimation in the population of a binary tree using node-sampling. Communications in Statistics: Theory and Methods, 43(13):2815–2829, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2010:CNC

[SS10a] Sanzhi Shi and Lixin Song. Comparison of nonparametric components in two partially linear models. Communications in Statistics: Theory and Methods, 39(10):1799–1816, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Skiadas:2010:DSA

[SS10b] Charilaos Skiadas and Christos H. Skiadas. Development, simulation, and application of first-exit-time densities to life table data. Communications in Statistics: Theory and Methods, 39 (3):444–451, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saporta:2011:P

[SS11a] Gilbert Saporta and Christos H. Skiadas. Preface. Communications in Statistics: Theory and Methods, 40(16):2821–2822, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:2011:TLM

[SS11b] W. F. Scott and B. Stewart. Tables for the Lilliefors and modified Cramér–von Mises tests of normality. Communications in Statistics: Theory and Methods, 40(4):726–730, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2012:PHM

[SS12a] P. G. Sankaran and V. N. Sreeja. A proportional hazards model for successive duration times under informative censoring. Communications in Statistics: Theory and Methods, 41(2): 262–280, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2012:NEB

[sS12b] Pao sheng Shen. Nonparametric estimation of the bivariate distribution function with doubly censored data. Communications in Statistics: Theory and Methods, 41(2):195–211, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Siray:2012:SCE

[ŠS12c] Gülesen Üstündağ Širay and Sadullah Sakallioğlu. Superiority of the r–k class estimator over some estimators in a linear model. Communications in Statistics: Theory and Methods, 41(15):2819–2832, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sreedevi:2012:SBA

[SS12d] E. P. Sreedevi and P. G. Sankaran. A semiparametric Bayesian approach for the analysis of competing risks data. Communications in Statistics: Theory and Methods, 41(15):2803–2818, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sunoj:2012:SRR

[SS12e] S. M. Sunoj and T. B. Sreejith. Some results on reciprocal subtangent in the context of weighted models. Communications in Statistics: Theory and Methods, 41(8):1397–1410, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sellers:2013:DDN

[SS13a] Kimberly F. Sellers and Galit Shmueli. Data dispersion: Now you see it … now you don’t. Communications in Statistics: Theory and Methods, 42(17):3134–3147, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2013:ECE

[SS13b] Housila P. Singh and Ramkrishna S. Solanki. An efficient class of estimators for the population mean using auxiliary information. Communications in Statistics: Theory and Methods, 42 (1):145–163, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2013:ERP

[SS13c] Housila P. Singh and Ramkrishna S. Solanki. Efficient ratio and product estimators in stratified random sampling. Communications in Statistics: Theory and Methods, 42(6):1008–1023, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2013:IEF

[SS13d] Housila P. Singh and Ramkrishna S. Solanki. Improved estimation of finite population variance using auxiliary information. Communications in Statistics: Theory and Methods, 42(15): 2718–2730, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2013:SCE

[SS13e] Housila P. Singh and Ramkrishna S. Solanki. Some classes of estimators for the population median using auxiliary information. Communications in Statistics: Theory and Methods, 42 (23):4222–4238, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Skiadas:2014:FET

[SS14a] Christos H. Skiadas and Charilaos Skiadas. The first exit time theory applied to life table data: The health state function of a population and other characteristics. Communications in Statistics: Theory and Methods, 43(7):1585–1600, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Solanki:2014:ECE

[SS14b] Ramkrishna S. Solanki and Housila P. Singh. An efficient class of estimators for the population mean using auxiliary information in stratified random sampling. Communications in Statistics: Theory and Methods, 43(16):3380–3401, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Subramani:2014:EPM

[SS14c] J. Subramani and Sarjinder Singh. Estimation of population mean in the presence of linear trend. Communications in Statistics: Theory and Methods, 43(15):3095–3116, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shukla:2015:PTR

[SS15a] D. Shukla and B. V. S. Sisodia. Preliminary test regression estimator when two prior values of population mean (μ_{x}) of an auxiliary variable (x) are available. Communications in Statistics: Theory and Methods, 44(17):3541–3548, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Solanki:2015:ICE

[SS15b] Ramkrishna S. Solanki and Housila P. Singh. An improved class of estimators for the general population parameters using auxiliary information. Communications in Statistics: Theory and Methods, 44(20):4241–4262, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Solanki:2015:SCE

[SS15c] Ramkrishna S. Solanki and Housila P. Singh. Some classes of estimators for median estimation in survey sampling. Communications in Statistics: Theory and Methods, 44(7):1450–1465, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankaran:2016:PHM

[SS16a] P. G. Sankaran and E. P. Sreedevi. A proportional hazards model for the analysis of doubly censored competing risks data. Communications in Statistics: Theory and Methods, 45(10): 2975–2987, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.894064.

Shi:2016:ARG

[SS16b] Jianhong Shi and Weixing Song. Asymptotic results in gamma kernel regression. Communications in Statistics: Theory and Methods, 45(12):3489–3509, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
890225.

Singh:2016:IEP

[SS16c] G. N. Singh and A. K. Singh. An improved estimation procedure of population mean in two-occasion successive sampling. Communications in Statistics: Theory and Methods, 45(13): 3930–3938, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.911911.

Singh:2016:TSC

[SS16d] Sarjinder Singh and Stephen Andrew Sedory. Two-step calibration of design weights in survey sampling. Communications in Statistics: Theory and Methods, 45(12):3510–3523, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.892137.

Siripanthana:2016:SRM

[SS16e] S. Siripanthana and E. C. Stillman. Sufficient reduction methods for multivariate time-dependent health surveillance data. Communications in Statistics: Theory and Methods, 45(4):950– 966, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2013.853793.

Skiadas:2016:HSC

[SS16f] Christos H. Skiadas and Charilaos Skiadas. The health state curve and the health state life table: Life expectancy and healthy life expectancy estimates. Communications in Statistics: Theory and Methods, 45(6):1682–1692, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.944662.

Solanki:2016:GFE

[SS16g] Ramkrishna S. Solanki and Housila P. Singh. The generalized family of estimators of population mean using auxiliary information in double sampling. Communications in Statistics: Theory and Methods, 45(14):4086–4105, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.882951.

Solanki:2016:IES

[SS16h] Ramkrishna S. Solanki and Housila P. Singh. An improved estimation in stratified random sampling. Communications in Statistics: Theory and Methods, 45(7):2056–2070, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.826367.

Srivastava:2016:CTR

[SS16i] Manoj Kumar Srivastava and Namita Srivastava. Chain-type ratio estimator in successive sampling using multi-auxiliary information. Communications in Statistics: Theory and Methods, 45(16):4798–4811, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
930907.

Sankaran:2017:QBC

[SS17a] P. G. Sankaran and S. M. Sunoj. Quantile-based cumulative entropies. Communications in Statistics: Theory and Methods, 46 (2):805–814, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1006779.

Shen:2017:CAR

[SS17b] Aiting Shen and Deping Shi. L_{r} convergence for arrays of rowwise asymptotically almost negatively associated random variables. Communications in Statistics: Theory and Methods, 46(23):11801–11812, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1285932.

Srivastava:2017:TMM

[SS17c] Muni S. Srivastava and Martin Singull. Test for the mean matrix in a Growth Curve model for high dimensions. Communications in Statistics: Theory and Methods, 46(13):6668–6683, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1132328.

Singh:2018:CTP

[SS18] Parminder Singh and Navdeep Singh. A closed testing procedure for comparing successive exponential populations with respect to location parameter. Communications in Statistics: Theory and Methods, 47(3):548–561, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1307407.

Samanthi:2019:BEB

[SS19] Ranadeera G. M. Samanthi and Jungsywan Sepanski. A bivariate extension of the beta generated distribution derived from copulas. Communications in Statistics: Theory and Methods, 48(5):1043–1059, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1429626.

Satheesh:2010:LDR

[SSA10] S. Satheesh, E. Sandhya, and T. Lovely Abraham. Limit distributions of random sums of Z_{+}-valued random variables. Communications in Statistics: Theory and Methods, 39(11):1979– 1984, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sharafi:2017:NGA

[SSB17a] Maryam Sharafi, Zahra Sajjadnia, and Javad Behboodian. A new generalization of alpha-skew-normal distribution. Communications in Statistics: Theory and Methods, 46(12):6098–6111, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1117639.

Shirke:2017:DGD

[SSB17b] D. T. Shirke, S. R. Supanekar, and Deepesh Bhati. On kdistorted generalized discrete family of distributions. Communications in Statistics: Theory and Methods, 46(23):11591– 11603, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1275697.

Singh:2017:EVE

[SSB17c] G. N. Singh, A. K. Sharma, and A. Bandyopadhyay. Effectual variance estimation strategy in two-occasion successive sampling in presence of random non response. Communications in Statistics: Theory and Methods, 46(14):7201–7224, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1146769.

Sang:2018:AFB

[SSC18] Liheng Sang, Guangjun Shen, and Qiangqiang Chang. Approximation of fractional Brownian sheet by Wiener integral. Communications in Statistics: Theory and Methods, 47(6):1423– 1441, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1321123.

Sen:2018:DIP

[SSD18] Sumen Sen, Pooja Sengupta, and Norou Diawara. Doubly inflated Poisson model using Gaussian copula. Communications in Statistics: Theory and Methods, 47(12):2848–2858, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1342831.

Sreedevi:2019:CCI

[SSD19] Engakkattu Purushothaman Sreedevi, Paduthol Godan Sankaran, and Isha Dewan. Comparison of cumulative incidence functions of current status competing risks data with discrete observation times. Communications in Statistics: Theory and Methods, 48(23):5766–5776, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1520887.

Sun:2010:BIM

[SSH10] Xiaoqian Sun, Dongchu Sun, and Zhuoqiong He. Bayesian inference on multivariate normal covariance and precision matrices in a star-shaped model with missing data. Communications in Statistics: Theory and Methods, 39(4):642–666, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shim:2017:MSV

[SSH17] Jooyong Shim, Kyungha Seok, and Changha Hwang. Monotone support vector quantile regression. Communications in Statistics: Theory and Methods, 46(10):5180–5193, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1096395.

Singh:2018:SEP

[SSKP18] G. N. Singh, S. Suman, M. Khetan, and C. Paul. Some estimation procedures of sensitive character using scrambled response techniques in successive sampling. Communications in Statistics: Theory and Methods, 47(8):1830–1841, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1327073.

Sun:2012:RAA

[SSM12] Liuquan Sun, Xinyuan Song, and Xiaoyun Mu. Regression analysis for the additive hazards model with covariate errors. Communications in Statistics: Theory and Methods, 41(11): 1911–1932, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Symeonaki:2016:IEM

[SSM16] M. Symeonaki, G. Stamatopoulou, and C. Michalopoulou. Intergenerational educational mobility in Greece: Transitions and social distances. Communications in Statistics: Theory and Methods, 45(6):1710–1722, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
957857.

Sugita:2011:FWI

[SSON11] Tomonori Sugita, Kunio Shimizu, S. H. Ong, and C. M. Ng. A family which integrates the generalized Charlier series and extended noncentral negative binomial distributions. Communications in Statistics: Theory and Methods, 40(11):1977–1998, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2012:RPU

[SSP^{+}12] G. N. Singh, V. K. Singh, Priyanka, Shakti Prasad, and Jaishree Prabha Karna. Rotation patterns under imputation of missing data over two-occasion. Communications in Statistics: Theory and Methods, 41(10):1857–1874, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sreejith:2019:NPE

[SSR19] T. B. Sreejith, S. M. Sunoj, and G. Rajesh. Non-parametric estimation of reciprocal coordinate subtangent for right censored dependent scheme. Communications in Statistics: Theory and Methods, 48(13):3177–3190, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473610.

Shariati:2014:PEA

[SSS14a] Nima Shariati, Hamid Shahriari, and Rasoul Shafaei. Parameter estimation of autoregressive models using the iteratively robust filtered fast-τ method. Communications in Statistics: Theory and Methods, 43(21):4445–4470, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2014:EBC

[SSS14b] Housila P. Singh, Ramkrishna S. Solanki, and Sarjinder Singh. Estimation of Bowley’s coefficient of skewness in the presence of auxiliary information. Communications in Statistics: Theory and Methods, 43(22):4867–4880, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2016:GRP

[SSS16] Housila P. Singh, Ramkrishna S. Solanki, and Alok K. Singh. A generalized ratio-cum-product estimator for estimating the finite population mean in survey sampling. Communications in Statistics: Theory and Methods, 45(1):158–172, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:2017:MBS

[SSS17a] Ajeet Kumar Singh, Priyanka Singh, and V. K. Singh. Model-based study of families of exponential-type estimators in presence of non response. Communications in Statistics: Theory and Methods, 46(13):6478–6490, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1129417.

Singh:2017:EUMb

[SSS17b] Housila P. Singh, Manoj K. Srivastava, and Namita Srivastava. Efficient use of multi-auxiliary information in search of good rotation patterns in successive sampling. Communications in Statistics: Theory and Methods, 46(5):2249–2269, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1041979.

Su:2017:AKM

[SSS17c] Shu-Ching Su, Stephan A. Sedory, and Sarjinder Singh. Adjusted Kuk’s model using two non sensitive characteristics unrelated to the sensitive characteristic. Communications in Statistics: Theory and Methods, 46(4):2055–2075, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1040503.

Singh:2018:EPM

[SSS18] G. N. Singh, A. K. Sharma, and A. K. Singh. Estimation of population mean on the most recent occasion under non response in h-occasion successive sampling using several auxiliary variables. Communications in Statistics: Theory and Methods, 47 (4):839–861, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1236957.

Salinas:2019:CET

[SSS19a] Veronica I. Salinas, Stephen A. Sedory, and Sarjinder Singh. Calibrated estimators in two-stage sampling. Communications in Statistics: Theory and Methods, 48(6):1449–1469, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1433850.

Singh:2019:RRR

[SSS19b] G. N. Singh, C. Singh, and S. Suman. Revisit of a randomized response model for estimating a rare sensitive attribute under probability proportional to size sampling using Poisson probability distribution. Communications in Statistics: Theory and Methods, 48(11):2766–2786, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1508718.

Singh:2018:TSU

[SSSK18] G. N. Singh, C. Singh, S. Suman, and A. Kumar. A twostage unrelated randomized response model for estimating a rare sensitive attribute in probability proportional to size sampling using Poisson distribution. Communications in Statistics: Theory and Methods, 47(15):3744–3766, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1361992.

Sharma:2016:GIL

[SSSM16] Vikas Kumar Sharma, Sanjay Kumar Singh, Umesh Singh, and Faton Merovci. The generalized inverse Lindley distribution: a new inverse statistical model for the study of upside-down bathtub data. Communications in Statistics: Theory and Methods, 45(19):5709–5729, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.948206.

Singh:2014:HRT

[SST14] Housila P. Singh, Balkishan Sharma, and Rajesh Tailor. Hartley–Ross type estimators for population mean using known parameters of auxiliary variate. Communications in Statistics: Theory and Methods, 43(3):547–565, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2018:CAM

[sSW18] Pao sheng Shen and Li Ning Weng. The Cox-aalen model for left-truncated and right-censored data. Communications in Statistics: Theory and Methods, 47(21):5357–5368, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1390135.

Statti:2018:HBP

[SSY18] Florencia Statti, Mariela Sued, and Victor J. Yohai. High breakdown point robust estimators with missing data. Communications in Statistics: Theory and Methods, 47(21):5145–5162, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1388396.

Sajeevkumar:2010:USE

[ST10] N. K. Sajeevkumar and P. Yageen Thomas. Use of spacings in the estimation of common scale parameter of several distributions. Communications in Statistics: Theory and Methods, 39(11):1951–1959, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2013:CBM

[ST13] Xiangjin Shen and Hiroki Tsurumi. Comparison of Bayesian model selection criteria and conditional Kolmogorov test as applied to spot asset pricing models. Communications in Statistics: Theory and Methods, 42(9):1599–1617, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanchez:2016:SMC

[ST16a] Juan Fernández Sánchez and Wolfgang Trutschnig. Some members of the class of (quasi-)copulas with given diagonal from the Markov kernel perspective. Communications in Statistics: Theory and Methods, 45(5):1508–1526, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.864856.

Sharma:2016:OBD

[ST16b] M. K. Sharma and Mekonnen Tadesse. Optimal block designs for double cross experiments. Communications in Statistics: Theory and Methods, 45(15):4392–4396, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.911909.

Shen:2016:PLD

[ST16c] Xinmei Shen and Hailan Tian. Precise large deviations for sums of two-dimensional random vectors with dependent components of heavy tails. Communications in Statistics: Theory and Methods, 45(21):6357–6368, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.839794.

Singh:2016:ENP

[ST16d] Housila P. Singh and Tanveer A. Tarray. An efficient new “Partial” randomized response model. Communications in Statistics: Theory and Methods, 45(22):6611–6624, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.963618.

Singh:2016:MOO

[ST16e] Housila P. Singh and Tanveer A. Tarray. A modified optimal orthogonal new additive model. Communications in Statistics: Theory and Methods, 45(22):6663–6669, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.963622.

Singh:2016:RWI

[ST16f] Housila P. Singh and Tanveer A. Tarray. Role of weights in improving the efficiency of Kim and Warde’s mixed randomized response model. Communications in Statistics: Theory and Methods, 45(4):1014–1030, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.853795.

Singh:2016:STO

[ST16g] Housila P. Singh and Tanveer A. Tarray. A stratified Tracy and Osahan’s two-stage randomized response model. Communications in Statistics: Theory and Methods, 45(11):3126–3137, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.895839.

Singh:2017:EUMa

[ST17a] Housila P. Singh and Tanveer A. Tarray. An efficient use of moment’s ratios of scrambling variables in a randomized response technique. Communications in Statistics: Theory and Methods, 46(2):521–531, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
997363.

Singh:2017:TSL

[ST17b] Housila P. Singh and Tanveer A. Tarray. A two-stage Land et al.’s randomized response model for estimating a rare sensitive attribute using Poisson distribution. Communications in Statistics: Theory and Methods, 46(1):389–405, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.995821.

Sobel:2018:BBS

[ST18] Marc Sobel and Ibrahim Turkoz. Bayesian blinded sample size re-estimation. Communications in Statistics: Theory and Methods, 47(24):5916–5933, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1404097.

Staudte:2017:IQM

[Sta17] Robert G. Staudte. Inference for quantile measures of kurtosis, peakedness, and tail weight. Communications in Statistics: Theory and Methods, 46(7):3148–3163, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1056366.

Son:2018:CMA

[STD18] Ta Cong Son, Dang Hung Thang, and Le Van Dung. On convergence of moving average series of martingale differences fields taking values in Banach spaces. Communications in Statistics: Theory and Methods, 47(22):5590–5603, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1397172.

Stepanov:2011:RBR

[Ste11] A. Stepanov. Runs based on the ratios of consecutive order statistics. Communications in Statistics: Theory and Methods, 40(18):3252–3268, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stepanov:2015:SC

[Ste15] Alexei Stepanov. On strong convergence. Communications in Statistics: Theory and Methods, 44(8):1615–1620, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stiburek:2017:SID

[Sti17] David Stiburek. Statistical inference on the drift parameter in fractional Brownian motion with a deterministic drift. Communications in Statistics: Theory and Methods, 46(2):892–905, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1006784.

Stoimenova:2011:PET

[Sto11] Eugenia Stoimenova. The power of exceedance-type tests under Lehmann alternatives. Communications in Statistics: Theory and Methods, 40(4):731–744, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stoyanov:2012:IPI

[Sto12] Jordan Stoyanov. Inference problems involving moment determinacy of distributions. Communications in Statistics: Theory and Methods, 41(16–17):2864–2878, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/03610926.2011.605237.

Singh:2017:EPB

[STRD17] Devendra Pratap Singh, Yogesh Mani Tripathi, Manoj Kumar Rastogi, and Nikhil Dabral. Estimation and prediction for a Burr type-III distribution with progressive censoring. Communications in Statistics: Theory and Methods, 46(19):9591–9613, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1213290.

Stylianou:2010:FCD

[Sty10] S. Stylianou. Foldover conference designs for screening experiments. Communications in Statistics: Theory and Methods, 39(10):1776–1784, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sengupta:2011:CMD

[SU11] Ashis Sengupta and Fidelis I. Ugwuowo. A classification method for directional data with application to the human skull. Communications in Statistics: Theory and Methods, 40 (3):457–466, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spoljaric:2013:SPP

[SU13] Drago Spoljarić and Ivo Ugrina. On statistical properties of palindromes in DNA. Communications in Statistics: Theory and Methods, 42(7):1373–1385, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srinivas:2014:BUE

[SU14] V. Srinivas and H. Jayakrishna Udupa. Best unbiased estimation and CAN property in the stable M/M/1 queue. Communications in Statistics: Theory and Methods, 43(2):321–327, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spoljaric:2017:LDN

[SU17a] Drago Spoljarić and Ivo Ugrina. Limiting distribution of the number of clumps of palindromes in DNA. Communications in Statistics: Theory and Methods, 46(17):8748–8759, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1189573.

Su:2017:APD

[Su17b] Chun-Lung Su. Asymptotic posterior distributions for balanced nested multi-way models with a large number of main effect levels. Communications in Statistics: Theory and Methods, 46(19):9425–9440, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1212076.

Subramani:2014:STS

[Sub14] J. Subramani. Star-type systematic sampling schemes for finite populations. Communications in Statistics: Theory and Methods, 43(1):175–190, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugimoto:2013:CPL

[Sug13] Tomoyuki Sugimoto. On consistencies of the profile likelihood estimators and their derivatives of the distribution function in doubly censored data. Communications in Statistics: Theory and Methods, 42(15):2741–2757, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimamura:2019:BGF

[SUKK19] Kaito Shimamura, Masao Ueki, Shuichi Kawano, and Sadanori Konishi. Bayesian generalized fused lasso modeling via NEG distribution. Communications in Statistics: Theory and Methods, 48(16):4132–4153, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1489056.

Sung:2011:CMA

[Sun11] Soo Hak Sung. Convergence of moving average processes for dependent random variables. Communications in Statistics: Theory and Methods, 40(13):2366–2376, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sung:2012:CCA

[Sun12] Soo Hak Sung. On complete convergence for arrays of dependent random variables. Communications in Statistics: Theory and Methods, 41(9):1663–1674, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sung:2013:SLT

[Sun13] Soo Hak Sung. Strong limit theorems for pairwise NQD random variables. Communications in Statistics: Theory and Methods, 42(21):3965–3973, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sung:2015:SLT

[Sun15] Soo Hak Sung. A strong limit theorem for weighted sums of negatively dependent random variables. Communications in Statistics: Theory and Methods, 44(2):428–439, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2017:THE

[Sun17a] Hui-Hui Sun. Testing for heteroscedasticity of exponential correlation mixed-effects linear models based on M-estimation. Communications in Statistics: Theory and Methods, 46(9): 4620–4630, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.955116.

Sun:2017:PDA

[Sun17b] Huihui Sun. Perturbation diagnostics of autocorrelation coefficients in non linear mixed-effects models with AR(1) errors based on M-estimation. Communications in Statistics: Theory and Methods, 46(16):8269–8277, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1177084.

Sun:2017:WQA

[Sun17c] Jing Sun. Weighted quantile average estimation for general linear models with missing covariates. Communications in Statistics: Theory and Methods, 46(17):8706–8722, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1189570.

Sun:2019:UBR

[Sun19] Zhongyang Sun. Upper bounds for ruin probabilities under model uncertainty. Communications in Statistics: Theory and Methods, 48(18):4511–4527, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1491991.

Suresh:2012:TCH

[Sur12] R. P. Suresh. A test for constant hazard against a changepoint alternative. Communications in Statistics: Theory and Methods, 41(9):1583–1589, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:2019:SEH

[SUS19] Richa Srivastava, S. K. Upadhyay, and V. K. Shukla. Subjective elicitation of hyperparameters of a conjugate Dirichlet prior and the corresponding Bayes analysis. Communications in Statistics: Theory and Methods, 48(19):4874–4887, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1500600.

Singh:2010:EMU

[SV10] Housila P. Singh and Gajendra K. Vishwakarma. Estimation of mean using auxiliary information and post-stratification. Communications in Statistics: Theory and Methods, 39(19):3456– 3471, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szatmari-Voicu:2011:RES

[SV11] D. Szatmari-Voicu. Robust M- and L-estimators of scale parameter. Communications in Statistics: Theory and Methods, 40(22):4065–4077, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shafiq:2017:EPS

[SV17] Muhammad Shafiq and Reinhard Viertl. On the estimation of parameters, survival functions, and hazard rates based on fuzzy life time data. Communications in Statistics: Theory and Methods, 46(10):5035–5055, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1099667.

Seaman:2014:IPW

[SW14a] Shaun Seaman and Ian White. Inverse probability weighting with missing predictors of treatment assignment or missingness. Communications in Statistics: Theory and Methods, 43(16): 3499–3515, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2014:AMC

[SW14b] Haifang Shi and Dehui Wang. An approximation model of the collective risk model with INAR(1) claim process. Communications in Statistics: Theory and Methods, 43(24):5305–5317, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shu:2014:SES

[SW14c] Ming-Hung Shu and Hsien-Chung Wu. Supplier evaluation and selection based on stochastic dominance: a quality-based approach. Communications in Statistics: Theory and Methods, 43(14):2907–2922, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:2019:DAF

[SW19a] Jiao Song and Jiang-Lun Wu. A detection algorithm for the first jump time in sample trajectories of jump-diffusions driven by α-stable white noise. Communications in Statistics: Theory and Methods, 48(19):4888–4902, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1500602.

Song:2019:GLL

[SW19b] Pingfan Song and Shaochen Wang. Generalized log-Lindley distribution and its applications in stochastic comparison. Communications in Statistics: Theory and Methods, 48(16):4008– 4018, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1483513.

Song:2019:MGF

[SW19c] Pingfan Song and Shaochen Wang. On the moment generating functions for integer valued random variables. Communications in Statistics: Theory and Methods, 48(20):5169–5174, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1508716.

Shen:2014:CCN

[SWL14] Aiting Shen, Xinghui Wang, and Jimin Ling. On complete convergence for nonstationary φ-mixing random variables. Communications in Statistics: Theory and Methods, 43(22):4856– 4866, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2014:RCC

[SWLW14] Aiting Shen, Xinghui Wang, Xiaoqin Li, and Xuejun Wang. On the rate of complete convergence for weighted sums of arrays of rowwise φ-mixing random variables. Communications in Statistics: Theory and Methods, 43(13):2714–2725, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2012:PRR

[SWW^{+}12] Jian Shi, Shaoping Wang, Kang Wang, John P. Sullivan, and Chin-Yin Chen. Performance-related reliability analysis methodology of computer networks. Communications in Statistics: Theory and Methods, 41(21):4000–4022, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2016:CPW

[SWZ16] Aiting Shen, Xinghui Wang, and Huayan Zhu. Convergence properties for weighted sums of NSD random variables. Communications in Statistics: Theory and Methods, 45(8):2402– 2412, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.881492.

Shen:2018:DCR

[SX18] Yan Shen and Ancha Xu. On the dependent competing risks using Marshall-Olkin bivariate Weibull model: Parameter estimation with different methods. Communications in Statistics: Theory and Methods, 47(22):5558–5572, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1397170.

Shen:2017:CCW

[SXW17] Aiting Shen, Mingxiang Xue, and Wenjuan Wang. Complete convergence for weighted sums of extended negatively dependent random variables. Communications in Statistics: Theory and Methods, 46(3):1433–1444, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019147.

Shen:2017:SCT

[SXZ17] Guangjun Shen, Liangwen Xia, and Dongjin Zhu. A strong convergence to the tempered fractional Brownian motion. Communications in Statistics: Theory and Methods, 46(8):4103–4118, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1078477.

Sakamoto:2010:AEF

[SY10] Yuji Sakamoto and Nakahiro Yoshida. Asymptotic expansion for functionals of a marked point process. Communications in Statistics: Theory and Methods, 39(8–9):1449–1465, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Said:2011:ANK

[SY11] Elias Ould Saïd and Djabrane Yahia. Asymptotic normality of a kernel conditional quantile estimator under strong mixing hypothesis and left-truncation. Communications in Statistics: Theory and Methods, 40(14):2605–2627, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2014:ASB

[SY14a] Guangjun Shen and Litan Yan. An approximation of subfractional Brownian motion. Communications in Statistics: Theory and Methods, 43(9):1873–1886, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2014:EDS

[SY14b] Guangjun Shen and Litan Yan. Estimators for the drift of subfractional Brownian motion. Communications in Statistics: Theory and Methods, 43(8):1601–1612, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2016:SLL

[SY16] Zhiyan Shi and Weiguo Yang. The strong law of large numbers and the Shannon–McMillan theorem for the m th-order nonhomogeneous Markov chains indexed by an m rooted Cayley tree. Communications in Statistics: Theory and Methods, 45(7): 2045–2055, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2013.791371.

Sepehrifar:2017:TME

[SY17a] Mohammad Sepehrifar and Shantia Yarahmadian. Testing monotonic equilibrium residual entropy of N-State Random Evolution. Communications in Statistics: Theory and Methods, 46(20):10088–10096, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1228967.

Shi:2017:DTI

[SY17b] Zhiyan Shi and Weiguo Yang. The definition of tree-indexed Markov chains in random environment and their existence. Communications in Statistics: Theory and Methods, 46(16): 7934–7941, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1167907.

Shi:2017:SLL

[SY17c] Zhiyan Shi and Weiguo Yang. Strong laws of large numbers for the m th-order asymptotic odd-even Markov chains indexed by an m-rooted Cayley tree. Communications in Statistics: Theory and Methods, 46(4):1855–1870, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030422.

Sun:2017:WCC

[SY17d] Xichao Sun and Litan Yan. Weak convergence to a class of multiple stochastic integrals. Communications in Statistics: Theory and Methods, 46(17):8355–8368, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1179758.

Sadeghi:2019:SNS

[SYC19] F. Sadeghi, F. Yousefzadeh, and M. Chahkandi. Some new stochastic orders based on quantile function. Communications in Statistics: Theory and Methods, 48(4):942–953, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1422757.

Symeonaki:2019:RCF

[Sym19] Maria Symeonaki. Rate of convergence in fuzzy non homogeneous Markov systems. Communications in Statistics: Theory and Methods, 48(1):40–49, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1395044.

Su:2013:APN

[SYP13] Rung-Hung Su, Dong-Yuh Yang, and W. L. Pearn. Assessing profitability of a newsboy-type product with normally distributed demand based on multiple samples. Communications in Statistics: Theory and Methods, 42(18):3401–3419, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:2016:CSD

[SYW16] Zhiyan Shi, Weiguo Yang, and Bei Wang. A class of small deviation theorems for the random variables associated with m th-order asymptotic circular Markov chains. Communications in Statistics: Theory and Methods, 45(23):7027–7039, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.972746.

Sun:2013:ELB

[SZ13] Guohong Sun and Xuemin Zi. An empirical-likelihood-based multivariate EWMA control scheme. Communications in Statistics: Theory and Methods, 42(3):429–446, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:2012:SAI

[SZD12] Zhimeng Sun, Zhongzhan Zhang, and Jiang Du. Semiparametric analysis of isotonic errors-in-variables regression models with missing response. Communications in Statistics: Theory and Methods, 41(11):2034–2060, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szewczak:2017:BET

[Sze17] Zbigniew S. Szewczak. Berry–Esséen theorem for sample quantiles of asymptotically uncorrelated non reversible Markov chains. Communications in Statistics: Theory and Methods, 46(8):3985–4003, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1076478.

Sun:2014:LCT

[SZL14] Chen Sun, Ming Zhao, and Yangjing Long. Learning concepts and taxonomic relations by metric learning for regression. Communications in Statistics: Theory and Methods, 43(14):2938– 2950, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:2016:RCS

[SZV16] Aiting Shen, Ying Zhang, and Andrei Volodin. On the rate of convergence in the strong law of large numbers for negatively orthant-dependent random variables. Communications in Statistics: Theory and Methods, 45(21):6209–6222, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.957858.

Takada:2010:TSS

[Tak10a] Yoshikazu Takada. A three-stage selection procedure with an exact consistency. Communications in Statistics: Theory and Methods, 39(8–9):1574–1584, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takagi:2010:CSO

[Tak10b] Yoshiji Takagi. On characterization of second-order asymptotic optimality under asymmetric loss functions: Loss coefficient vector and admissibility. Communications in Statistics: Theory and Methods, 39(8–9):1359–1372, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takagi:2014:ALR

[Tak14] Yoshiji Takagi. Asymptotics of likelihood ratio tests for general one-sided hypotheses in the two-sample normal model. Communications in Statistics: Theory and Methods, 43(17):3614–3628, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takai:2018:USM

[Tak18] Keiji Takai. On the use of the selection matrix in the maximum likelihood estimation of normal distribution models with missing data. Communications in Statistics: Theory and Methods, 47(14):3392–3407, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1353631.

Tan:2010:CMM

[Tan10a] Frans E. S. Tan. Conditions for D_{A}-maximin marginal designs for generalized linear mixed models to be uniform. Communications in Statistics: Theory and Methods, 40(2):255–266, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:2010:SCA

[Tan10b] Hidekazu Tanaka. Sufficient conditions for the admissibility under LINEX loss function in regular case. Communications in Statistics: Theory and Methods, 39(8–9):1477–1489, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taneja:2013:GSD

[Tan13] Inder Jeet Taneja. Generalized symmetric divergence measures and the probability of error. Communications in Statistics: Theory and Methods, 42(9):1654–1672, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:2018:AMS

[Tan18] Zhongquan Tan. Asymptotics of maxima and sums for a type of strongly dependent isotropic Gaussian random fields. Communications in Statistics: Theory and Methods, 47(20):5013–5028, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1383430.

Tarasinska:2017:SSN

[Tar17] Joanna Tarasińska. Sample size needed to get given ratio of endpoints for confidence interval of standard deviation in a normal distribution. Communications in Statistics: Theory and Methods, 46(23):11480–11484, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1271428.

Tattar:2016:BTH

[Tat16] Prabhanjan N. Tattar. Bootstrap test of hypothesis for the multi-state models in survival analysis. Communications in Statistics: Theory and Methods, 45(5):1270–1277, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.861490.

Tavangar:2016:SRT

[Tav16] Mahdi Tavangar. Some representation theorems in terms of mean residual lifetime and mean inactivity time of coherent systems. Communications in Statistics: Theory and Methods, 45 (2):226–235, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tuncer:2011:FCD

[TAY11] Yalcin Tuncer, Fahrettin Arslan, and Mehmet Yilmaz. Functions of concordance and dependence with related measures. Communications in Statistics: Theory and Methods, 40(18): 3349–3365, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tahmasebi:2012:IPC

[TB12] S. Tahmasebi and J. Behboodian. Information properties for concomitants of order statistics in Farlie-Gumbel-Morgenstern (FGM) family. Communications in Statistics: Theory and Methods, 41(11):1954–1968, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Turkmen:2013:IFA

[TB13] Asuman S. Turkmen and Nedret Billor. Influence function analysis for the robust partial least squares (RoPLS) estimator. Communications in Statistics: Theory and Methods, 42(15): 2818–2836, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Teles:2015:MIT

[TB15] Paulo Teles and Paula Brito. Modeling interval time series with space-time processes. Communications in Statistics: Theory and Methods, 44(17):3599–3627, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tailor:2014:RPT

[TC14] Rajesh Tailor and Sunil Chouhan. Ratio-cum-product type exponential estimator of finite population mean in stratified random sampling. Communications in Statistics: Theory and Methods, 43(2):343–354, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tablada:2017:MFD

[TC17] Claudio J. Tablada and Gauss M. Cordeiro. The modified Fréchet distribution and its properties. Communications in Statistics: Theory and Methods, 46(21):10617–10639, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1239115.

Tahir:2016:LFD

[TCA^{+}16] M. H. Tahir, Gauss M. Cordeiro, Ayman Alzaatreh, M. Mansoor, and M. Zubair. The logistic-X family of distributions and its applications. Communications in Statistics: Theory and Methods, 45(24):7326–7349, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.980516.

Tchatoka:2019:NEW

[Tch19] Firmin Doko Tchatoka. Near exogeneity, weak identification and specification testing: Some asymptotic results. Communications in Statistics: Theory and Methods, 48(13):3191–3207, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1473611.

Tahir:2016:WDD

[TCM^{+}16] M. H. Tahir, Gauss M. Cordeiro, M. Mansoor, M. Zubair, and Morad Alizadeh. The Weibull–Dagum distribution: Properties and applications. Communications in Statistics: Theory and Methods, 45(24):7376–7398, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.983610.

Tuntapthai:2015:BEB

[TCN15] Mongkhon Tuntapthai, Nattakarn Chaidee, and Kritsana Neammanee. Berry–Esséen bounds for random index non linear statistics via Stein’s method. Communications in Statistics: Theory and Methods, 44(16):3464–3485, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Theobald:2014:GTP

[TD14] C. M. Theobald and A. M. Davie. Group testing, the pooled hypergeometric distribution, and estimating the number of defectives in small populations. Communications in Statistics: Theory and Methods, 43(14):3019–3026, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tahmasebi:2018:MIR

[TD18] S. Tahmasebi and S. Daneshi. Measures of inaccuracy in record values. Communications in Statistics: Theory and Methods, 47(24):6002–6018, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1404102.

Talebi:2011:WSP

[TE11] Hooshang Talebi and Nabaz Esmailzadeh. Weighted searching probability for classes of equivalent search designs comparison.

Communications in Statistics: Theory and Methods, 40(4):635– 647, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tuncel:2018:SRU

[TE18a] Altan Tuncel and Serkan Eryilmaz. System reliability under δ-shock model. Communications in Statistics: Theory and Methods, 47(19):4872–4880, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1459708.

Tuzen:2018:CCD

[TE18b] Muhammed Fatih Tüzen and Semra Erbas. A comparison of count data models with an application to daily cigarette consumption of young persons. Communications in Statistics: Theory and Methods, 47(23):5825–5844, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1402050.

Teerapabolarn:2010:NUB

[Tee10] K. Teerapabolarn. Non uniform bounds on geometric approximation via Stein’s method and w-functions. Communications in Statistics: Theory and Methods, 40(1):145–158, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Teerapabolarn:2012:LUB

[Tee12] K. Teerapabolarn. The least upper bound on the Poissonnegative binomial relative error. Communications in Statistics: Theory and Methods, 41(10):1833–1838, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Teerapabolarn:2014:IPA

[Tee14] K. Teerapabolarn. An improvement of Poisson approximation for sums of dependent Bernoulli random variables. Communications in Statistics: Theory and Methods, 43(8):1758–1777, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsagris:2012:APC

[TEF12] Michael Tsagris, Ioannis Elmatzoglou, and Christos C. Frangos. The assessment of performance of correlation estimates in discrete bivariate distributions using bootstrap methodology. Communications in Statistics: Theory and Methods, 41 (1):138–152, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tenreiro:2017:WLS

[Ten17] C. Tenreiro. A weighted least-squares cross-validation bandwidth selector for kernel density estimation. Communications in Statistics: Theory and Methods, 46(7):3438–3458, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1062108.

Tenreiro:2018:NCB

[Ten18] Carlos Tenreiro. A new class of boundary kernels for distribution function estimation. Communications in Statistics: Theory and Methods, 47(21):5319–5332, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1390131.

Tang:2012:ODT

[TGXX12] Yincai Tang, Qiang Guan, Peirong Xu, and Haiyan Xu. Optimum design for Type-I step-stress accelerated life tests of twoparameter Weibull distributions. Communications in Statistics: Theory and Methods, 41(21):3863–3877, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:2015:FDP

[TH15] Rodrigo Tsai and Luiz K. Hotta. Fitting distributions with the polyhazard model with dependence. Communications in Statistics: Theory and Methods, 44(9):1886–1895, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsou:2017:LIC

[TH17] Tsung-Shan Tsou and Wei-Cheng Hsiao. Likelihood inference for correlated binary data without any information about the joint distributions. Communications in Statistics: Theory and Methods, 46(5):2151–2160, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1033553.

Theodossiou:2019:STI

[The19] Panayiotis Theodossiou. Skewed type III generalized logistic distribution. Communications in Statistics: Theory and Methods, 48(23):5809–5819, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1522348.

Tang:2015:GPM

[THHL15] Xingyu Tang, Zhaoping Hong, Yuao Hu, and Heng Lian. Gaussian process models for non parametric functional regression with functional responses. Communications in Statistics: Theory and Methods, 44(16):3428–3445, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thomas:2014:RPR

[Tho14] Naiju M. Thomas. On the ratios of pathway random variables. Communications in Statistics: Theory and Methods, 43(23): 4972–4987, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taheri:2016:CTF

[THV16] S. M. Taheri, G. Hesamian, and R. Viertl. Contingency tables with fuzzy information. Communications in Statistics: Theory and Methods, 45(20):5906–5917, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.953688.

Tahir:2011:CPD

[TIA11] M. H. Tahir, Ijaz Iqbal, and M. L. Aggarwal. Cyclic polygonal designs with block size 3 and joint distance α = 2. Communications in Statistics: Theory and Methods, 40(14):2583–2590, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2012:CRB

[Tia12] Yongge Tian. Characterizing relationships between estimations under a general linear model with explicit and implicit restrictions by rank of matrix. Communications in Statistics: Theory and Methods, 41(13–14):2588–2601, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ting:2010:OTL

[Tin10] Chao-Ping Ting. D-optimal two-level fractional factorial designs of resolution III with two dispersion factors. Communications in Statistics: Theory and Methods, 39(14):2515–2532, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tung:2013:GMD

[TJ13] David D. Tung and S. Rao Jammalamadaka. On the Gini mean difference test for circular data. Communications in Statistics: Theory and Methods, 42(11):1998–2008, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takai:2013:AII

[TK13] Keiji Takai and Yutaka Kano. Asymptotic inference with incomplete data. Communications in Statistics: Theory and Methods, 42(17):3174–3190, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Triantafyllou:2014:FRA

[TK14] Ioannis S. Triantafyllou and Markos V. Koutras. Failure rate and aging properties of generalized beta- and gamma-generated distributions. Communications in Statistics: Theory and Methods, 43(19):4046–4061, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Toker:2017:REC

[TK17a] Selma Toker and Selahattin Kaçiranlar. Ridge estimator with correlated errors and two-stage ridge estimator under inequality restrictions. Communications in Statistics: Theory and Methods, 46(3):1407–1421, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019145.

Tschumitschew:2017:EDN

[TK17b] Katharina Tschumitschew and Frank Klawonn. Effects of drift and noise on the optimal sliding window size for data stream regression models. Communications in Statistics: Theory and Methods, 46(10):5109–5132, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1096388.

Turner:2019:IUD

[TK19] Ronna C. Turner and Elizabeth A. Keiffer. Impact of unbalanced DIF item proportions on group-specific DIF identification. Communications in Statistics: Theory and Methods, 48(15):3746–3760, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481968.

Tripathy:2017:EQS

[TKJ17] Manas Ranjan Tripathy, Somesh Kumar, and Adarsha Kumar Jena. Estimating quantiles of several normal populations with a common mean. Communications in Statistics: Theory and Methods, 46(11):5656–5671, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1109663.

Tripathi:2017:MEL

[TKP17] Yogesh Mani Tripathi, Somesh Kumar, and C. Petropoulos. Minimax estimators for the lower-bounded scale parameter of a location-scale family of distributions. Communications in Statistics: Theory and Methods, 46(18):9185–9193, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1205611.

Teh:2012:MPM

[TKW12] S. Y. Teh, Michael B. C. Khoo, and Zhang Wu. Monitoring process mean and variance with a single generally weighted moving average chart. Communications in Statistics: Theory and Methods, 41(12):2221–2241, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taraldsen:2015:FPS

[TL15a] Gunnar Taraldsen and Bo H. Lindqvist. Fiducial and posterior sampling. Communications in Statistics: Theory and Methods, 44(17):3754–3767, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:2015:FCE

[TL15b] Yun-Cheng Tsai and Yuh-Dauh Lyuu. Faster convergence to the estimation of quadratic variation with microstructure noise.

Communications in Statistics: Theory and Methods, 44(13): 2827–2841, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taraldsen:2016:CPI

[TL16] Gunnar Taraldsen and Bo Henry Lindqvist. Conditional probability and improper priors. Communications in Statistics: Theory and Methods, 45(17):5007–5016, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.935430.

Tsai:2017:NRK

[TL17] Yun-Cheng Tsai and Yuh-Dauh Lyuu. A new robust Kalman filter for filtering the microstructure noise. Communications in Statistics: Theory and Methods, 46(10):4961–4976, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1096390.

Tian:2017:BCQ

[TLT17] Yuzhu Tian, Heng Lian, and Maozai Tian. Bayesian composite quantile regression for linear mixed-effects models. Communications in Statistics: Theory and Methods, 46(15):7717–7731, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1161798.

Turcicova:2019:MML

[TME19] Marie Turcicová, Jan Mandel, and Krystof Eben. Multilevel maximum likelihood estimation with application to covariance matrices. Communications in Statistics: Theory and Methods, 48(4):909–925, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1422755.

Tan:2017:OCS

[TMZ^{+}17] Jiyang Tan, Yuhui Ma, Hanjun Zhang, Ziqiang Li, and Xiangqun Yang. Optimal control strategies for dividend payments and capital injections in compound Markov binomial risk model with penalties for deficits. Communications in Statistics: Theory and Methods, 46(10):5072–5092, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1096385.

Tong:2019:EPD

[TMZZ19] Jinying Tong, Xuan Ma, Zhenzhong Zhang, and Enwen Zhu. Ergodicity for population dynamics driven by stable processes with Markovian switching. Communications in Statistics: Theory and Methods, 48(10):2446–2458, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1465089.

Tabrizi:2010:MEB

[TN10] N. Jafari Tabrizi and N. Nematollahi. Minimax estimation of the bounded parameter of some discrete distributions under LINEX loss function. Communications in Statistics: Theory and Methods, 39(15):2701–2710, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thongtha:2013:CUB

[TN13] D. Thongtha and K. Neammanee. Constants on a uniform Berry–Esséen bound on some Borel sets in r^{k} via Stein’s method. Communications in Statistics: Theory and Methods, 42(23):4247–4258, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsukuda:2015:ATT

[TN15] Koji Tsukuda and Yasushi Nagata. Asymptotic theory of Taguchi’s natural estimators of the signal to noise ratio for dynamic robust parameter design. Communications in Statistics: Theory and Methods, 44(22):4734–4741, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Teimouri:2016:SBS

[TN16] Mahdi Teimouri and Saralees Nadarajah. Simulation of Balakrishnan skew-normal order statistics. Communications in Statistics: Theory and Methods, 45(11):3310–3322, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901381.

Taghipour:2017:DTS

[TN17] M. Taghipour and A. R. Nematollahi. On discrete-time stable multiple Markov processes. Communications in Statistics: Theory and Methods, 46(4):1694–1708, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1026993.

Tamatani:2019:HDA

[TN19] Mitsuru Tamatani and Kanta Naito. High dimensional asymptotics for the naive Hotelling T^{2} statistic in pattern recognition. Communications in Statistics: Theory and Methods, 48(22): 5637–5656, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1517217.

Toomaj:2017:REP

[Too17] A. Toomaj. Renyi entropy properties of mixed systems. Communications in Statistics: Theory and Methods, 46(2):906–916, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1006785.

Torabi:2012:GVN

[Tor12] Hamzeh Torabi. A generalized version of Neyman–Pearson lemma for testing fuzzy hypotheses based on r-level sets. Communications in Statistics: Theory and Methods, 41(24):4379– 4390, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:2015:SOA

[TOT15] Hidekazu Tanaka, Chie Obayashi, and Yoshiji Takagi. On second order admissibilities in two-parameter logistic regression model. Communications in Statistics: Theory and Methods, 44(9):1958–1966, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:2014:TPR

[TP14] Tiejun Tong and Heng Peng. Tail probability ratios of normal and Student’s t distributions. Communications in Statistics: Theory and Methods, 43(18):3797–3811, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thomas:2019:DDP

[TP19] P. Yageen Thomas and Jerin Paul. On diagnostic devices for proposing half-logistic and inverse half-logistic models using generalized (k) record values. Communications in Statistics: Theory and Methods, 48(5):1073–1091, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1423697.

Tripathi:2018:ESP

[TPJ18] Yogesh Mani Tripathi, Constantinos Petropoulos, and Mayank Jha. Estimation of the shape parameter of a Pareto distribution. Communications in Statistics: Theory and Methods, 47(18):4459–4468, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1376088.

Tafakori:2018:NLM

[TPN18] Laleh Tafakori, Armin Pourkhanali, and Saralees Nadarajah. A new lifetime model with different types of failure rate. Communications in Statistics: Theory and Methods, 47(16):4006–4020, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1367811.

Trong:2014:TRD

[TPTT14] Dang Duc Trong, Cao Xuan Phuong, Truong Trung Tuyen, and Dinh Ngoc Thanh. Tikhonov’s regularization to the deconvolution problem. Communications in Statistics: Theory and Methods, 43(20):4384–4400, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tennekoon:2016:SMB

[TR16a] Vidhura Tennekoon and Robert Rosenman. Systematically misclassified binary dependent variables. Communications in Statistics: Theory and Methods, 45(9):2538–2555, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.887105.

Tripathi:2016:EUH

[TR16b] Yogesh Mani Tripathi and Manoj Kumar Rastogi. Estimation using hybrid censored data from a generalized inverted exponential distribution. Communications in Statistics: Theory and Methods, 45(16):4858–4873, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.932805.

Tran:2011:COP

[Tra11] Minh-Ngoc Tran. A criterion for optimal predictive model selection. Communications in Statistics: Theory and Methods, 40 (5):893–906, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tran:2017:LPE

[Tra17] Ngoc Khue Tran. LAN property for an ergodic Ornstein– Uhlenbeck process with Poisson jumps. Communications in Statistics: Theory and Methods, 46(16):7942–7968, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1167908.

Triacca:2017:PID

[Tri17] Umberto Triacca. Predictive information and distance between past and future of a time series. Communications in Statistics: Theory and Methods, 46(16):8230–8235, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1177081.

Tiwari:2012:OPS

[TS12] Neeraj Tiwari and U. C. Sud. An optimal procedure for sample coordination using multiple objective functions and nearest proportional to IPPS size sampling designs. Communications in Statistics: Theory and Methods, 41(11):2014–2033, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tarray:2016:SSU

[TS16] Tanveer A. Tarray and Housila P. Singh. A sinuous stratified unrelated question randomized response model. Communications in Statistics: Theory and Methods, 45(22):6510–6520, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.960588.

Tarray:2017:ASK

[TS17a] Tanveer A. Tarray and Housila P. Singh. An adept-stratified Kuk’s randomized response model using Neyman allocation. Communications in Statistics: Theory and Methods, 46(6): 2870–2881, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1053933.

Tarray:2017:ORR

[TS17b] Tanveer A. Tarray and Housila P. Singh. An optional randomized response model for estimating a rare sensitive attribute using Poisson distribution. Communications in Statistics: Theory and Methods, 46(6):2638–2654, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1040506.

Tarray:2017:SRR

[TS17c] Tanveer A. Tarray and Housila P. Singh. A stratified randomized response model for sensitive characteristics using the negative hypergeometric distribution. Communications in Statistics: Theory and Methods, 46(6):2607–2629, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1010007.

Tarray:2017:SUQ

[TS17d] Tanveer A. Tarray and Housila P. Singh. A stratified unrelated question randomized response model using Neyman allocation. Communications in Statistics: Theory and Methods, 46(1):17– 27, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.983612.

Tiwari:2017:PCI

[TS17e] Vidhika Tiwari and N. K. Singh. Process capability index for bivariate exponentially distributed quality characteristics and its sampling properties. Communications in Statistics: Theory and Methods, 46(22):11099–11109, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1257717.

Towhidi:2017:TFC

[TS17f] M. Towhidi and M. Salmanpour. A test of fit for a continuous distribution based on the empirical convex conditional mean function. Communications in Statistics: Theory and Methods, 46(3):1490–1505, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1019151.

Teerapabolarn:2019:NUB

[TS19] K. Teerapabolarn and C. Soponpimol. A non uniform bound on geometric approximation with w-functions. Communications in Statistics: Theory and Methods, 48(16):4119–4131, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1487983.

Tsai:2010:HAB

[Tsa10] Miao-Yu Tsai. A hybrid approximation Bayesian test of variance components for longitudinal data. Communications in Statistics: Theory and Methods, 39(16):2849–2864, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tabrizi:2019:AMC

[TSG19] E. Tabrizi, E. Bahrami Samani, and M. Ganjali. Analysis of mixed correlated bivariate zero-inflated count and (k, l)-inflated beta responses with application to social network datasets. Communications in Statistics: Theory and Methods, 48(7): 1651–1681, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1435815.

Tsionas:2010:BAP

[Tsi10] Efthymios G. Tsionas. Bayesian analysis of Poisson regression with lognormal unobserved heterogeneity: With an application to the patent-R&D relationship. Communications in Statistics: Theory and Methods, 39(10):1689–1706, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2014:BAL

[Tsi14] Efthymios G. Tsionas. Bayesian analysis of least absolute relative error regression. Communications in Statistics: Theory and Methods, 43(23):4988–4997, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2015:LBI

[Tsi15] Efthymios G. Tsionas. Likelihood-based inference in Sdistributions. Communications in Statistics: Theory and Methods, 44(1):153–158, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:2018:NPI

[Tsi18] Mike G. Tsionas. Note on posterior inference for the Bingham distribution. Communications in Statistics: Theory and Methods, 47(12):3022–3028, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1346805.

Terfas:2019:TSQ

[TSOT19] Imane Terfas, Hafida Saggou, and Megdouda Ourbih-Tari. Transient study of a queueing system with one unreliable server, batch arrivals, two types of verification, loss and vacation. Communications in Statistics: Theory and Methods, 48(10):2580– 2601, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1472780.

Tailor:2015:IPE

[TSS15] Rajesh Tailor, Balkishan Sharma, and Housila P. Singh. An improved procedure of estimating the finite population mean using auxiliary information in the presence of random non response. Communications in Statistics: Theory and Methods, 44(6):1196–1209, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsukada:2014:UEC

[Tsu14] Shin-Ichi Tsukada. Unbiased estimator for a covariance matrix under two-step monotone incomplete sample. Communications in Statistics: Theory and Methods, 43(8):1613–1629, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:2010:ALE

[TT10] Hidekazu Tanaka and Masashi Tatsukawa. On the admissibility of linear estimators in a multivariate normal distribution under LINEX loss function. Communications in Statistics: Theory and Methods, 39(17):3011–3020, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Toque:2011:TSF

[TT11] Carole Toque and Virginie Terraza. Time series factorial models with uncertainty measures: Applications to ARMA processes and financial data. Communications in Statistics: Theory and Methods, 40(9):1533–1544, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tran:2017:SEF

[TT17] Kien C. Tran and Mike G. Tsionas. Series estimation of functional-coefficient partially linear regression model. Communications in Statistics: Theory and Methods, 46(15):7593–7602, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1157189.

Tailor:2017:IRP

[TTC17] Rajesh Tailor, Ritesh Tailor, and Sunil Chouhan. Improved ratio- and product-type exponential estimators for population mean in case of post-stratification. Communications in Statistics: Theory and Methods, 46(21):10387–10393, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2012.732180.

Tian:2014:EFM

[TTZ14a] Yuzhu Tian, Maozai Tian, and Qianqian Zhu. Estimating a finite mixed exponential distribution under progressively Type-II censored data. Communications in Statistics: Theory and Methods, 43(17):3762–3776, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2014:LQR

[TTZ14b] Yuzhu Tian, Maozai Tian, and Qianqian Zhu. Linear quantile regression based on EM algorithm. Communications in Statistics: Theory and Methods, 43(16):3464–3484, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tu:2017:EEN

[Tu17] Y. Tu. Efficient estimation of non parametric simultaneous equations models. Communications in Statistics: Theory and Methods, 46(7):3411–3416, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1062104.

Tunno:2015:BAT

[Tun15] Ferebee Tunno. Bounded area tests for comparing the dynamics between ARMA processes. Communications in Statistics: Theory and Methods, 44(18):3921–3941, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thomas:2011:ACO

[TV11] P. Yageen Thomas and T. G. Veena. On an application of concomitants of order statistics in characterizing a family of bivariate distributions. Communications in Statistics: Theory and Methods, 40(8):1445–1452, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:2013:EVD

[TW13] Zhongquan Tan and Yuebao Wang. Extremes values of discrete and continuous time strongly dependent Gaussian processes. Communications in Statistics: Theory and Methods, 42(13): 2451–2463, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:2014:ABM

[TW14] Mingtian Tang and Yunyan Wang. The asymptotic behavior of INAR(p) models. Communications in Statistics: Theory and Methods, 43(14):3047–3056, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tian:2018:NFM

[TWG18] Weizhong Tian, Tonghui Wang, and Arjun K. Gupta. A new family of multivariate skew slash distribution. Communications in Statistics: Theory and Methods, 47(23):5812–5824, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1402049.

Taylor:2010:POB

[TWH10] Douglas J. Taylor, Mark A. Weaver, and Natalie Cheung Hall. A proportional odds beta-binomial model for evaluating the effect of treatment in cross-over studies with baseline covariates: an application to condom failure data. Communications in Statistics: Theory and Methods, 39(16):2991–3000, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:2014:TSN

[TWTG14] Man-Lai Tang, Qin Wu, Guo-Liang Tian, and Jian-Hua Guo. Two-sample non randomized response techniques for sensitive questions. Communications in Statistics: Theory and Methods, 43(2):408–425, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tan:2016:CCN

[TWZ16] Xili Tan, Hang Wang, and Yong Zhang. Complete convergence of the non-identically distributed pairwise NQD random sequences. Communications in Statistics: Theory and Methods, 45(9):2626–2637, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
804569.

Tian:2014:GEL

[TX14] Ruiqin Tian and Liugen Xue. Generalized empirical likelihood inference in generalized linear models for longitudinal data. Communications in Statistics: Theory and Methods, 43(18): 3893–3904, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:2017:GSL

[TY17] Ying Tang and Weiguo Yang. The generalizations of strong law of large numbers for asymptotic even-odd Markov chains indexed by a homogeneous tree. Communications in Statistics: Theory and Methods, 46(19):9415–9424, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1212073.

Tounsi:2017:WFM

[TZ17] Mariem Tounsi and Raoudha Zine. Wilks’ factorization of the matrix-variate Dirichlet–Riesz distributions. Communications in Statistics: Theory and Methods, 46(9):4494–4509, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1085565.

Tang:2016:LEC

[TZW16] Linjun Tang, Zhangong Zhou, and Changchun Wu. The LAD estimation of the change-point linear model with randomly censored data. Communications in Statistics: Theory and Methods, 45(2):479–491, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Usha:2017:DMC

[UB17] M. Usha and S. Balamurali. Designing of a mixed-chain sampling plan based on the process capability index with chain sampling as the attributes plan. Communications in Statistics: Theory and Methods, 46(21):10456–10475, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1236961.

Ugbe:2014:NOS

[UC14] T. A. Ugbe and P. E. Chigbu. On non overlapping segmentation of the response surfaces for solving constrained programming problems through super convergent line series. Communications in Statistics: Theory and Methods, 43(2):306–320, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ukaegbu:2015:CPC

[UC15] Eugene C. Ukaegbu and Polycarp E. Chigbu. Comparison of prediction capabilities of partially replicated central composite designs in cuboidal region. Communications in Statistics: Theory and Methods, 44(2):406–427, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Urbano:2012:WRG

[UDC12] Mariana Ragassi Urbano, Clarice Garcia Borges Demétrio, and Gauss Moutinho Cordeiro. On Wald residuals in generalized linear models. Communications in Statistics: Theory and Methods, 41(4):741–758, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Udom:2018:SAR

[Udo18] Akaninyene Udo Udom. Stochastic approximation results for variational inequality problem using random-type iterative schemes. Communications in Statistics: Theory and Methods, 47(22):5435–5444, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1395046.

Um:2016:SCS

[UK16] Sang Joon Um and Jong Gurl Kim. A simultaneous control scheme for mean and fraction non conforming in small process variation. Communications in Statistics: Theory and Methods, 45(16):4745–4759, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
784989.

Upadhyay:2009:BAC

[UM09] S. K. Upadhyay and Bhaswati Mukherjee. Bayes analysis and comparison of accelerated Weibull and accelerated Birnbaum– Saunders models. Communications in Statistics: Theory and Methods, 39(2):195–213, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ullah:2013:AIL

[UPA13a] M. A. Ullah, G. R. Pasha, and M. Aslam. Assessing influence on the Liu estimates in linear regression models. Communications in Statistics: Theory and Methods, 42(17):3100–3116, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ullah:2013:LID

[UPA13b] M. A. Ullah, G. R. Pasha, and M. Aslam. Local influence diagnostics in the modified ridge regression. Communications in Statistics: Theory and Methods, 42(10):1851–1869, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ugwuowo:2011:MSS

[UU11] F. I. Ugwuowo and E. F. Udoumoh. Multi-stage semi-Markov models for recovery process. Communications in Statistics: Theory and Methods, 40(16):3011–3019, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ushakov:2017:PTH

[UU17] N. G. Ushakov and V. G. Ushakov. Permutation tests for homogeneity based on some characterizations. Communications in Statistics: Theory and Methods, 46(15):7692–7702, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1158838.

Unal:2013:CIS

[ÜY13] Deniz Ünal and Güzin Yüksel. Comparison of the iterative Stein-rule and the usual estimators of the disturbance variance under the Pitman nearness criterion in a linear regression model with proxy variables. Communications in Statistics: Theory and Methods, 42(8):1531–1541, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Uyanto:2017:CRT

[Uya17] Stanislaus S. Uyanto. Coefficient of relationship for two symmetric alpha-stable variables when alpha is in the interval (1, 2]. Communications in Statistics: Theory and Methods, 46(14): 6874–6881, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1137599.

Udom:2019:GSS

[UZ19] Akaninyene Udo Udom and Quanxin Zhu. Global stability of stochastic systems with Poisson distributed random time-delay. Communications in Statistics: Theory and Methods, 48(6): 1305–1315, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1429629.

VanZyl:2016:RIN

[Van16] J. Martin Van Zyl. Regression with an infinite number of observations applied to estimating the parameters of the stable distribution using the empirical characteristic function. Communications in Statistics: Theory and Methods, 45(11):3323–3331, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.901382.

VanDerwerken:2017:EGS

[Van17] Douglas VanDerwerken. Not every Gibbs sampler is a special case of the Metropolis–Hastings algorithm. Communications in Statistics: Theory and Methods, 46(20):10005–10009, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1228961.

Vassiliou:2013:EPN

[Vas13] P.-C. G. Vassiliou. Exotic properties of non homogeneous Markov and semi-Markov systems. Communications in Statistics: Theory and Methods, 42(16):2971–2990, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vasiliadis:2014:TAM

[Vas14a] G. Vasiliadis. Transient analysis of the M/M/k/N/N queue using a continuous time homogeneous Markov system with finite state size capacity. Communications in Statistics: Theory and Methods, 43(7):1548–1562, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vassiliou:2014:MSG

[Vas14b] P. C. G. Vassiliou. Markov systems in a general state space. Communications in Statistics: Theory and Methods, 43(7): 1322–1339, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vassiliou:2014:P

[Vas14c] P.-C. G. Vassiliou. Preface. Communications in Statistics: Theory and Methods, 43(7):1293–1295, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vasiliadis:2016:TAF

[Vas16] George Vasiliadis. Transient analysis of a finite source discretetime queueing system using homogeneous Markov system with state size capacities (HMS/c). Communications in Statistics: Theory and Methods, 45(5):1403–1423, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.863931.

Vock:2012:NPI

[VB12] Michael Vock and N. Balakrishnan. Nonparametric prediction intervals based on ranked set samples. Communications in Statistics: Theory and Methods, 41(12):2256–2268, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vock:2013:CBT

[VB13] Michael Vock and N. Balakrishnan. A connection between two nonparametric tests for perfect ranking in balanced ranked set sampling. Communications in Statistics: Theory and Methods, 42(2):191–193, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VanNiekerk:2015:SBA

[VBAR15] J. Van Niekerk, A. Bekker, M. Arashi, and J. J. J. Roux. Subjective Bayesian analysis of the elliptical model. Communications in Statistics: Theory and Methods, 44(17):3738–3753, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Volkovich:2011:APN

[VBATK11] Zeev Volkovich, Zeev Barzily, Renata Avros, and Dvora Toledano-Kitai. On application of a probabilistic K-nearest neighbors model for cluster validation problem. Communications in Statistics: Theory and Methods, 40(16):2997–3010, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Verrill:2015:AEE

[VEKH15] Steve P. Verrill, James W. Evans, David E. Kretschmann, and Cherilyn A. Hatfield. Asymptotically efficient estimation of a bivariate Gaussian–Weibull distribution and an introduction to the associated pseudo-truncated Weibull. Communications in Statistics: Theory and Methods, 44(14):2957–2975, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Velilla:2014:PSD

[Vel14] Santiago Velilla. On the properties of the SAVE directions. Communications in Statistics: Theory and Methods, 43(21): 4612–4627, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Veres-Ferrer:2017:EFD

[VFP17a] Ernesto J. Veres-Ferrer and Jose M. Pavía. Elasticity function of a discrete random variable and its properties. Communications in Statistics: Theory and Methods, 46(17):8631–8646, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1186190.

Veres-Ferrer:2017:PEC

[VFP17b] Ernesto J. Veres-Ferrer and Jose M. Pavía. Properties of the elasticity of a continuous random variable. A special look at its behavior and speed of change. Communications in Statistics: Theory and Methods, 46(6):3054–3069, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1053943.

Vijay:2011:RBF

[Vij11] Vivek Vijay. Relationships between full and layer models with applications to level merging. Communications in Statistics: Theory and Methods, 40(4):745–761, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vinogradov:2013:SUL

[Vin13] Vladimir Vinogradov. Some utilizations of Lambert W function in distribution theory. Communications in Statistics: Theory and Methods, 42(11):2025–2043, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vishwakarma:2018:ARE

[Vis18] Gajendra K. Vishwakarma. An alternative to ratio estimator in post-stratification. Communications in Statistics: Theory and Methods, 47(8):1989–2000, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1335419.

Varghese:2014:NBR

[VJV14] Eldho Varghese, Seema Jaggi, and Cini Varghese. Neighborbalanced row-column designs. Communications in Statistics: Theory and Methods, 43(6):1261–1276, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vieira:2015:NMB

[VLD15] Caroline C. Vieira, Rosangela H. Loschi, and Denise Duarte. Nonparametric mixtures based on skew-normal distributions: an application to density estimation. Communications in Statistics: Theory and Methods, 44(8):1552–1570, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Votsi:2014:HSM

[VLTP14] I. Votsi, N. Limnios, G. Tsaklidis, and E. Papadimitriou. Hidden semi-Markov modeling for the estimation of earthquake occurrence rates. Communications in Statistics: Theory and Methods, 43(7):1484–1502, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Volaufova:2014:P

[VM14] Julia Volaufova and Augustyn Markiewicz. Preface. Communications in Statistics: Theory and Methods, 43(5):919–920, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vanderperre:2017:STR

[VM17] E. J. Vanderperre and S. S. Makhanov. On the survival time of a repairable duplex system sustained by a cold standby unit subjected to a priority rule. Communications in Statistics: Theory and Methods, 46(16):7872–7886, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1165847.

Vangeneugden:2010:MCL

[VML^{+}10] Tony Vangeneugden, Geert Molenberghs, Annouschka Laenen, Helena Geys, Caroline Beunckens, and Cristina Sotto. Marginal correlation in longitudinal binary data based on generalized linear mixed models. Communications in Statistics: Theory and Methods, 39(19):3540–3557, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vangeneugden:2014:MCL

[VMVD14] Tony Vangeneugden, Geert Molenberghs, Geert Verbeke, and Clarice G. B. Demétrio. Marginal correlation from logit- and probit-beta-normal models for hierarchical binary data. Communications in Statistics: Theory and Methods, 43(19):4164– 4178, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

V:2017:BAM

[VN17] Sergio A. Calderón V. and Fabio H. Nieto. Bayesian analysis of multivariate threshold autoregressive models with missing data. Communications in Statistics: Theory and Methods, 46 (1):296–318, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.990758.

VoVan:2018:SCC

[VN18] Tai VoVan and Thao NguyenTrang. Similar coefficient for cluster of probability density functions. Communications in Statistics: Theory and Methods, 47(8):1792–1811, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1327075.

Voinov:2010:DPF

[Voi10] Vassilly Voinov. A decomposition of Pearson–Fisher and Dzhaparidze–Nikulin statistics and some ideas for a more powerful test construction. Communications in Statistics: Theory and Methods, 39(4):667–677, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Voinov:2015:NWT

[Voi15] Vassilly Voinov. A note on Wald’s type chi-squared tests of fit. Communications in Statistics: Theory and Methods, 44(21): 4622–4630, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Voinov:2010:NVV

[VP10] Vassilly Voinov and Natalie Pya. A note on vector-valued goodness-of-fit tests. Communications in Statistics: Theory and Methods, 39(3):452–459, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Valero:2016:LTM

[VPCG16] Jordi Valero, Marta Pérez-Casany, and Josep Ginebra. On left-truncating and mixing Poisson distributions. Communications in Statistics: Theory and Methods, 45(5):1424–1434, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.863932.

Viseu:2014:MUI

[VPMF14] C. Viseu, L. Pereira, A. P. Martins, and H. Ferreira. On the multivariate upcrossings index. Communications in Statistics: Theory and Methods, 43(6):1277–1292, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Voinov:2016:NIC

[VPMV16] Vassilly Voinov, Natalie Pya, Rashid Makarov, and Yevgeniy Voinov. New invariant and consistent chi-squared type goodness-of-fit tests for multivariate normality and a related comparative simulation study. Communications in Statistics: Theory and Methods, 45(11):3249–3263, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901370.

VandenBerg:2013:BGG

[VRB13] J. Van den Berg, J. J. J. Roux, and A. Bekker. A bivariate generalization of gamma distribution. Communications in Statistics: Theory and Methods, 42(19):3514–3527, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Villandre:2012:SCT

[VRC12] Luc Villandré, Benjamin Rich, and Antonio Ciampi. Soft classification trees. Communications in Statistics: Theory and Methods, 41(16–17):3244–3258, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.1080/03610926.2011.
632103.

Vanegas:2013:ARI

[VRC13] Luis Hernando Vanegas, Luz Marina Rondon, and Francisco José A. Cysneiros. Assessing robustness of inference in symmetrical nonlinear regression models. Communications in Statistics: Theory and Methods, 42(9):1692–1711, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ventosa-Santaularia:2010:SIV

[VS10] Daniel Ventosa-Santaulària. Spurious instrumental variables. Communications in Statistics: Theory and Methods, 39(11): 1997–2007, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Viswakala:2019:CEH

[VS19] K. V. Viswakala and E. I. Abdul Sathar. Classical estimation of hazard rate and mean residual life functions of Pareto distribution. Communications in Statistics: Theory and Methods, 48(17):4367–4379, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1494289.

vanStaden:2016:KLE

[vSK16] Paul J. van Staden and Robert A. R. King. Kurtosis of the logistic-exponential survival distribution. Communications in Statistics: Theory and Methods, 45(23):6891–6899, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.972566.

Ventosa-Santaularia:2016:SSS

[VSN16] Daniel Ventosa-Santaulària and Antonio E. Noriega. A simple solution for spurious regressions. Communications in Statistics: Theory and Methods, 45(19):5561–5583, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.948196.

Vazler:2012:WMD

[VSS12] Ivan Vazler, Kristian Sabo, and Rudolf Scitovski. Weighted median of the data in solving least absolute deviations problems. Communications in Statistics: Theory and Methods, 41(8):1455–1465, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Veena:2015:ACO

[VT15] T. G. Veena and P. Yageen Thomas. Application of concomitants of order statistics of independent non identically distributed random variables in estimation. Communications in Statistics: Theory and Methods, 44(12):2530–2545, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Veena:2017:RCO

[VT17] T. G. Veena and P. Yageen Thomas. Role of concomitants of order statistics in determining parent bivariate distributions. Communications in Statistics: Theory and Methods, 46(16): 7976–7997, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1171351.

Vilar:2012:BTE

[VV12] Juan M. Vilar and José A. Vilar. A bootstrap test for the equality of nonparametric regression curves under dependence. Communications in Statistics: Theory and Methods, 41(6):1069– 1088, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VanderWielen:2015:GSD

[VV15] Michael Vander Wielen and Ryan Vander Wielen. The general segmented distribution. Communications in Statistics: Theory and Methods, 44(10):1994–2009, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Varghese:2017:VBR

[VV17] Eldho Varghese and Cini Varghese. Variance-balanced row– column designs involving diallel crosses incorporating specific combining abilities for comparing test lines with a control line. Communications in Statistics: Theory and Methods, 46(5): 2123–2131, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1032428.

Villa:2017:MJL

[VW17] Cristiano Villa and Stephen Walker. On the mathematics of the Jeffreys–Lindley paradox. Communications in Statistics: Theory and Methods, 46(24):12290–12298, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1295073.

Varathan:2018:OGL

[VW18] Nagarajah Varathan and Pushpakanthie Wijekoon. Optimal generalized logistic estimator. Communications in Statistics: Theory and Methods, 47(2):463–474, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1307406.

ValdEs:2014:BET

[VYZ14] JosÉ E. ValdÉs, Yoel G. Yera, and Leonel Zuaznabar. Bounds for the expected time to extinction and the probability of extinction in the Galton–Watson process. Communications in Statistics: Theory and Methods, 43(8):1698–1707, 2014. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

vanZyl:2017:EEW

[vZ17] J. Martin van Zyl. Exact expressions for the weights used in least-squares regression estimation for the log-logistic and Weibull distribution. Communications in Statistics: Theory and Methods, 46(4):1720–1730, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1026995.

Vellaisamy:2019:EOS

[VZ19] Palaniappan Vellaisamy and Aklilu Zeleke. Exponential order statistics and some combinatorial identities. Communications in Statistics: Theory and Methods, 48(20):5099–5105, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1508710.

vanZyl:2017:BCC

[vZvdM17] R. van Zyl and A. J. van der Merwe. A Bayesian control chart for a common coefficient of variation. Communications in Statistics: Theory and Methods, 46(12):5795–5811, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1112914.

vanZyl:2019:BPM

[vZvdM19] R. van Zyl and A. J. van der Merwe. Bayesian process monitoring schemes for the two-parameter exponential distribution. Communications in Statistics: Theory and Methods, 48(7): 1766–1797, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1440307.

Wu:2017:WSR

[WA17] Jibo Wu and Yasin Asar. A weighted stochastic restricted ridge estimator in partially linear model. Communications in Statistics: Theory and Methods, 46(18):9274–9283, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1206936.

Wu:2018:RAU

[WAA18] Jibo Wu, Yasin Asar, and Mohammad Arashi. On the restricted almost unbiased Liu estimator in the logistic regression model. Communications in Statistics: Theory and Methods, 47(18): 4389–4401, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1376082.

Wu:2017:DVT

[WAJ17] Chien-Wei Wu, Muhammad Aslam, and Chi-Hyuck Jun. Developing a variables two-plan sampling system for product acceptance determination. Communications in Statistics: Theory and Methods, 46(2):706–720, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1004092.

Wang:2010:SMO

[Wan10a] P. C. Wang. A simple method for obtaining minimum aberration designs. Communications in Statistics: Theory and Methods, 39(18):3363–3370, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2010:TPL

[Wan10b] Zhu Wang. Tests for paired lifetime data with frailty models. Communications in Statistics: Theory and Methods, 39(17):

3122–3139, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2012:EIE

[Wan12] Bing Xing Wang. Exact interval estimation for the scale family under general progressive Type-II censoring. Communications in Statistics: Theory and Methods, 41(24):4444–4452, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:UAO

[Wan13a] Kaiyong Wang. The uniform asymptotics of the overshoot of a random walk with light-tailed increments. Communications in Statistics: Theory and Methods, 42(5):830–837, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:CPD

[Wan13b] Lichun Wang. Consistency of posterior distributions for heteroscedastic nonparametric regression models. Communications in Statistics: Theory and Methods, 42(15):2731–2740, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:QME

[Wan13c] Litong Wang. Quasi-minimax estimation in the linear model with measurement errors. Communications in Statistics: Theory and Methods, 42(8):1563–1571, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:SKI

[Wan13d] Litong Wang. Several Kantorovich inequalities involving Tracy–Singh product. Communications in Statistics: Theory and Methods, 42(2):194–200, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:CCE

[Wan15] Hongxia Wang. Conditional Choquet expectation. Communications in Statistics: Theory and Methods, 44(18):3782–3795, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2016:IVG

[Wan16a] Jiantian Wang. Improved versions of Greenwood estimators under Koziol–Green model. Communications in Statistics: Theory and Methods, 45(1):142–157, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2016:EED

[Wan16b] Liang Wang. Estimation for exponential distribution based on competing risk middle censored data. Communications in Statistics: Theory and Methods, 45(8):2378–2391, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.879893.

Wang:2016:IEL

[Wan16c] Liang Wang. Interval estimation for a lower-truncated distribution based on the double Type-II censored sample. Communications in Statistics: Theory and Methods, 45(19):5679–5692, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.948204.

Wang:2016:IPL

[Wan16d] Ran Wang. Inferences on partial linear models with right censored data by empirical likelihood. Communications in Statistics: Theory and Methods, 45(21):6342–6356, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.839042.

Wang:2016:CSC

[Wan16e] Yashi Wang. Conditional k-out-of-n systems with a cold standby component. Communications in Statistics: Theory and Methods, 45(21):6253–6262, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
958354.

Wang:2017:BCT

[Wan17a] Lei Wang. Bartlett-corrected two-sample adjusted empirical likelihood via resampling. Communications in Statistics: Theory and Methods, 46(22):10941–10952, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1257720.

Wang:2017:RIG

[Wan17b] Liang Wang. Reliability inference for a general lower-truncated family of distributions under records. Communications in Statistics: Theory and Methods, 46(12):6151–6173, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1122057.

Wang:2017:DTL

[Wan17c] P. C. Wang. Design two-level factorial experiments in blocks of size four. Communications in Statistics: Theory and Methods, 46(1):9–16, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.983611.

Wang:2017:OME

[Wan17d] P. C. Wang. Orthogonal main-effect plans in row-column designs for two-level factorial experiments. Communications in Statistics: Theory and Methods, 46(21):10685–10691, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1242743.

Wang:2017:ETS

[Wan17e] Qing Wang. Extrapolation techniques in U-statistic variance estimation. Communications in Statistics: Theory and Methods, 46(17):8387–8400, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1179760.

Wang:2017:SLT

[Wan17f] Z.-Z. Wang. Some limit theorems of delayed sums for rowwise conditionally independent stochastic arrays. Communications in Statistics: Theory and Methods, 46(11):5265–5272, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1099674.

Wang:2017:SHE

[Wan17g] Zhi Wang. Stochastic heat equation and martingale differences. Communications in Statistics: Theory and Methods, 46(4): 1650–1660, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1024866.

Wang:2018:LRO

[Wan18] Jiantian Wang. Likelihood ratio order of parallel systems under multiple-outlier models. Communications in Statistics: Theory and Methods, 47(1):55–63, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1300278.

Wang:2019:MRE

[Wan19a] Lei Wang. Multiple robustness estimation in causal inference. Communications in Statistics: Theory and Methods, 48(23): 5701–5718, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1520881.

Wang:2019:TVL

[Wan19b] Lihong Wang. Time varying long memory parameter estimation for locally stationary long memory processes. Communications in Statistics: Theory and Methods, 48(10):2529–2547, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1472777.

Watkins:2011:IRB

[Wat11] Alan J. Watkins. On an integral related to the Burr XII distribution. Communications in Statistics: Theory and Methods, 40(21):3777–3779, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Worz:2018:NML

[WB18] Sascha Wörz and Heinz Bernhardt. A note on multivariate linear regression. Communications in Statistics: Theory and Methods, 47(19):4785–4790, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1445863.

Worz:2019:MHT

[WBGS19] Sascha Wörz, Heinz Bernhardt, Anja Gräff, and Huber Stefan. Multiple hypothesis tests based on conditional differences in means. Communications in Statistics: Theory and Methods, 48(4):1033–1041, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1425452.

Wallenstein:2010:CRP

[WBH10] Sylvan Wallenstein, Carol Bodian, and Robin Herbert. A capture-recapture problem when information is obtained from two qualitatively different sources. Communications in Statistics: Theory and Methods, 39(15):2688–2700, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2016:NHE

[WC16] Chunlin Wang and Gemai Chen. A new hybrid estimation method for the generalized Pareto distribution. Communications in Statistics: Theory and Methods, 45(14):4285–4294, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.919399.

Wu:2018:UWT

[WCF18] Yuh-Jenn Wu, Li-Hsueh Cheng, and Wei-Quan Fang. Using Wald-type estimator to combat outliers and berkson-type uncertainties with mixture distributions in linear regression models. Communications in Statistics: Theory and Methods, 47(14):3324–3337, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1353627.

Wang:2012:UAD

[WCGZ12] Xueli Wang, Hua Chen, Zhi Geng, and Xiaohua Zhou. Using auxiliary data for binomial parameter estimation with nonignorable nonresponse. Communications in Statistics: Theory and Methods, 41(19):3468–3478, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2016:RRB

[WCK16] Shu Wu, Philippe Castagliola, and Michael B. C. Khoo. Run rules based phase II c and np charts when process parameters are unknown. Communications in Statistics: Theory and Methods, 45(4):1182–1197, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
861493.

Wang:2010:ELI

[WCL10] Xiuli Wang, Fang Chen, and Lu Lin. Empirical likelihood inference for the parameter in additive partially linear EV models. Communications in Statistics: Theory and Methods, 39(19): 3513–3524, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2011:ACI

[WCL11] Dja-Shin Wang, Chao-Yu Chou, and Yu-Chang Lin. On the asymptotic confidence intervals of multiple-stream yield index. Communications in Statistics: Theory and Methods, 40(22): 3948–3958, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2017:SNR

[WCW17] Shijie Wang, Cen Chen, and Xuejun Wang. Some novel results on pairwise quasi-asymptotical independence with applications to risk theory. Communications in Statistics: Theory and Methods, 46(18):9075–9085, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1202287.

Wang:2016:RES

[WCZP16] Liying Wang, Lirong Cui, Jie Zhang, and Jing Peng. Reliability evaluation of a semi-Markov repairable system under alternative environments. Communications in Statistics: Theory and Methods, 45(10):2938–2957, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.894062.

Wang:2016:VCK

[WDLZ16] Jinliang Wang, Changshou Deng, Jiangfeng Li, and Mingfa Zhou. On variances and covariances of a kind of extreme order statistics. Communications in Statistics: Theory and Methods, 45(11):3274–3282, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
901373.

Wang:2014:CMF

[WDM14] Renxiang Wang, Juan Du, and Chunsheng Ma. Covariance matrix functions of isotropic vector random fields. Communications in Statistics: Theory and Methods, 43(10–12):2081–2093, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wen:2017:UAN

[WdX17] Yuzhuo Wen and Guo dong Xing. On the uniformly asymptotic normality of frequency polygons for ψ-mixing samples. Communications in Statistics: Theory and Methods, 46(21):10416– 10425, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1205617.

Wan:2011:GFT

[WDZ11] Shuwen Wan, Xin Deng, and Biao Zhang. A goodness-of-fit test for logistic regression models in stratified case-control studies via empirical likelihood. Communications in Statistics: Theory and Methods, 40(22):3988–4000, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Webb:2011:SEM

[Web11] Steve Webb. Screening experiments with many factors. Communications in Statistics: Theory and Methods, 40(10):1879– 1892, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Webel:2012:SND

[Web12] Karsten Webel. A short note on the dependence structure of a stochastic process generated by the asymmetric cusp map. Communications in Statistics: Theory and Methods, 41(2):357– 362, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weerahandi:2012:GPE

[Wee12] Samaradasa Weerahandi. Generalized point estimation with application to small response estimation. Communications in Statistics: Theory and Methods, 41(22):4069–4095, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:2012:SIP

[Wei12] Chuanhua Wei. Statistical inference in partially linear varyingcoefficient models with missing responses at random. Communications in Statistics: Theory and Methods, 41(7):1284–1298, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:2014:SSR

[Wei14] Laisheng Wei. A summary of some research on PC and Bayesian PC criterion in China. Communications in Statistics: Theory and Methods, 43(18):3866–3892, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ward:2019:ICS

[WER^{+}19] Ralph C. Ward, Leonard Egede, Viswanathan Ramakrishnan, Lewis Frey, Robert Neal Axon, Clara Libby E. Dismuke, Kelly J. Hunt, and Mulugeta Gebregziabher. An improved comorbidity summary score for measuring disease burden and predicting mortality with applications to two national cohorts. Communications in Statistics: Theory and Methods, 48(18): 4642–4655, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1498896.

Wu:2017:CEA

[WF17] Yuh-Jenn Wu and Wei-Quan Fang. Consistent estimation approach to tackling collinearity and berkson-type measurement error in linear regression using adjusted Wald-type estimator. Communications in Statistics: Theory and Methods, 46(11): 5501–5516, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1104353.

Wang:2014:VWS

[WG14] Xiao Wang and Joseph Glaz. Variable window scan statistics for normal data. Communications in Statistics: Theory and Methods, 43(10–12):2489–2504, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:DEJ

[WG15] Dan Wang and Pengjiang Guo. Detection and estimation of jump points in non parametric regression function with AR(1) noise. Communications in Statistics: Theory and Methods, 44(6):1097–1110, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weerahandi:2016:GMI

[WG16a] Sam Weerahandi and Jinadasa Gamage. A general method of inference for two-parameter continuous distributions. Communications in Statistics: Theory and Methods, 45(9):2612–2625, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.887109.

Wu:2016:CWS

[WG16b] Yongfeng Wu and Mingle Guo. Convergence of weighted sums for sequences of pairwise NQD random variables. Communications in Statistics: Theory and Methods, 45(20):5977–5989, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.953695.

Wei:2019:RWB

[WGG^{+}19] Wenhui Wei, Shesheng Gao, Bingbing Gao, Yongmin Zhong, Chengfan Gu, and Zhaohui Gao. Random weighting-based quantile estimation via importance resampling. Communications in Statistics: Theory and Methods, 48(19):4820–4833, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1496256.

Wang:2019:CPC

[WGW19] Shijie Wang, Duo Guo, and Wensheng Wang. Closure property of consistently varying random variables based on precise large deviation principles. Communications in Statistics: Theory and Methods, 48(9):2218–2228, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1459717.

Wu:2010:GCI

[WH10] Chien-Wei Wu and P. H. Huang. Generalized confidence intervals for comparing the capability of two processes. Communications in Statistics: Theory and Methods, 39(13):2351–2364, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2011:FIL

[WH11] Yanhua Wang and Shuyuan He. Fisher information in lefttruncated data. Communications in Statistics: Theory and Methods, 40(9):1525–1532, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weng:2013:SEP

[WH13] Ruby C. Weng and Cheng-Hung Hsu. A study of expansions of posterior distributions. Communications in Statistics: Theory and Methods, 42(2):346–364, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2014:SBP

[WH14] Dongliang Wang and Alan D. Hutson. A smooth bootstrap procedure towards deriving confidence intervals for the relative risk. Communications in Statistics: Theory and Methods, 43(9):1979–1990, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:ITB

[WH15a] Dongliang Wang and Alan D. Hutson. Inversion theorem based kernel density estimation for the ordinary least squares estimator of a regression coefficient. Communications in Statistics: Theory and Methods, 44(8):1571–1579, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:SRTa

[WH15b] Gaowen Wang and Nan-Wei Han. Spurious regressions in time series with long memory. Communications in Statistics: Theory and Methods, 44(4):837–854, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2016:LEQ

[WH16a] Hongjun Wang and Paul S. Horn. A low-end quantile estimator from a right-skewed distribution. Communications in Statistics: Theory and Methods, 45(10):2810–2833, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.889162.

Wang:2016:AMC

[WH16b] Tianhua Wang and Shuguang Huang. An adaptive multivariate CUSUM control chart for signaling a range of location shifts. Communications in Statistics: Theory and Methods, 45(16): 4673–4691, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.927494.

Wiedermann:2016:APP

[WH16c] Wolfgang Wiedermann and Michael Hagmann. Asymmetric properties of the Pearson correlation coefficient: Correlation as the negative association between linear regression residuals. Communications in Statistics: Theory and Methods, 45(21): 6263–6283, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.960582.

Wang:2017:WHD

[WH17a] Dongliang Wang and Alan D. Hutson. A weighted Harrell– Davis distance test with applications to censored data. Communications in Statistics: Theory and Methods, 46(10):5022–5034, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1096396.

Wang:2017:SCE

[WH17b] Xinghui Wang and Shuhe Hu. The strong consistency of M-estimates in linear models with extended negatively dependent errors. Communications in Statistics: Theory and Methods, 46(10):5093–5108, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1096386.

Watts:2018:MTP

[WH18] David D. Watts and Joshua D. Habiger. A multiple testing protocol for exploratory data analysis and the local misclassification rate. Communications in Statistics: Theory and Methods, 47(15):3588–3604, 2018. CODEN CSTMDC.

ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1361982.

Wang:2016:AEL

[WHC16] Shanshan Wang, Tao Hu, and Hengjian Cui. Adjusted empirical likelihood inference for additive hazards regression. Communications in Statistics: Theory and Methods, 45(24):7294–7305, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.978948.

Wang:2017:RWS

[WHHW17] Shijie Wang, Yiyu Hu, Jijiao He, and Xuejun Wang. Randomly weighted sums and their maxima with heavy-tailed increments and dependence structure. Communications in Statistics: Theory and Methods, 46(21):10851–10863, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1248785.

Wang:2016:GRC

[WHV16] Xuejun Wang, Shuhe Hu, and Andrei Volodin. General results of complete convergence and complete moment convergence for weighted sums of some class of random variables. Communications in Statistics: Theory and Methods, 45(15):4494–4508, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.921308.

Wang:2013:CCW

[WHVH13] Xuejun Wang, Tien-Chung Hu, Andrei Volodin, and Shuhe Hu. Complete convergence for weighted sums and arrays of rowwise extended negatively dependent random variables. Communications in Statistics: Theory and Methods, 42(13):2391–2401, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2018:DER

[WHY18] Kaiyuan Wu, Wei Hou, and Hongbo Yang. Density estimation via the random forest method. Communications in Statistics: Theory and Methods, 47(4):877–889, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1285929.

Wang:2018:RWS

[WHYW18] Shijie Wang, Yiyu Hu, LianQiang Yang, and Wensheng Wang. Randomly weighted sums under a wide type of dependence structure with application to conditional tail expectation. Communications in Statistics: Theory and Methods, 47(20):5054– 5063, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1386309.

Wen:2017:SSS

[WHZ17] Miin-Jye Wen, Li-Ching Huang, and Junjiang Zhong. Singlestage sampling procedure of the t best populations under heteroscedasticity. Communications in Statistics: Theory and Methods, 46(18):9265–9273, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1206935.

Wiedermann:2018:NFM

[Wie18] Wolfgang Wiedermann. A note on fourth moment-based direction dependence measures when regression errors are non normal. Communications in Statistics: Theory and Methods, 47(21):5255–5264, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1388403.

Willink:2009:SFD

[Wil09] R. Willink. A single form for t-distributions and symmetric beta distributions. Communications in Statistics: Theory and Methods, 39(1):170–176, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:2013:BTS

[WILL13] Heung Wong, Wai Cheung Ip, Jin Shan Liu, and Jian Yan Long. Bayesian time series analysis of structural changes in level and trend. Communications in Statistics: Theory and Methods, 42 (21):3949–3964, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Witt:2014:SDS

[Wit14] Gary Witt. A simple distribution for the sum of correlated, exchangeable binary data. Communications in Statistics: Theory and Methods, 43(20):4265–4280, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2011:SCE

[WJ11] Qunying Wu and Yuanying Jiang. The strong consistency of M estimator in a linear model for negatively dependent random samples. Communications in Statistics: Theory and Methods, 40(3):467–491, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:2015:SIP

[WJH15] Chuan-Hua Wei, Xu-Jie Jia, and Hong-Sheng Hu. Statistical inference on partially linear additive models with missing response variables and error-prone covariates. Communications in Statistics: Theory and Methods, 44(4):872–883, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:PIE

[WJY15] Xiaoqin Wang, Yin Jin, and Li Yin. Point and interval estimations of marginal risk difference by logistic model. Communications in Statistics: Theory and Methods, 44(17):3703–3722, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2012:SEA

[WK12] Hong Wang and John Klein. Semiparametric estimation for the additive inverse Gaussian frailty model. Communications in Statistics: Theory and Methods, 41(12):2269–2278, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williams:2013:TAC

[WK13] Matthew R. Williams and Dong-Yun Kim. A test for an abrupt change in Weibull hazard functions with staggered entry and Type I censoring. Communications in Statistics: Theory and Methods, 42(11):1922–1933, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weerahandi:2019:NRA

[WK19] S. Weerahandi and K. Krishnamoorthy. A note reconciling ANOVA tests under unequal error variances. Communications in Statistics: Theory and Methods, 48(3):689–693, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1419264.

Wang:2010:EEE

[WL10] Fu-Kwun Wang and Chih-Wen Lee. An M-estimator for estimating the extended Burr Type III parameters with outliers. Communications in Statistics: Theory and Methods, 40(2):304– 322, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2011:LLE

[WL11] Hanchao Wang and Zhengyan Lin. Local linear estimation of second-order diffusion models. Communications in Statistics: Theory and Methods, 40(3):394–407, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:EBL

[WL13] Qing-Wen Wang and Xin Liu. The equalities of BLUPs for linear combinations under two general linear mixed models. Communications in Statistics: Theory and Methods, 42(19): 3528–3543, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2014:DFA

[WL14] Nan Wang and Wei Liu. A distribution-free approach in statistical modelling with repeated measurements and missing values. Communications in Statistics: Theory and Methods, 43(8): 1686–1697, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:VSS

[WL15] Kangning Wang and Lu Lin. Variable selection in semiparametric quantile modeling for longitudinal data. Communications in Statistics: Theory and Methods, 44(11):2243–2266, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2017:ESH

[WL17a] Chao Wang and Hong Liu. Estimation for the scaled halflogistic distribution under Type-I progressively hybrid censoring scheme. Communications in Statistics: Theory and Methods, 46(24):12045–12058, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1291968.

Wang:2017:GSF

[WL17b] Houchun Wang and Nengxiang Ling. On the Gerber–Shiu function with random discount rate. Communications in Statistics: Theory and Methods, 46(1):210–220, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.988265.

Wang:2017:ODH

[WL17c] Tao Wang and Zhonghua Li. Outlier detection in highdimensional regression model. Communications in Statistics: Theory and Methods, 46(14):6947–6958, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1140783.

Wu:2017:ETS

[WL17d] Xing Wu and Tian Liu. Estimation and testing for semiparametric mixtures of partially linear models. Communications in Statistics: Theory and Methods, 46(17):8690–8705, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1189569.

Wang:2019:WEH

[WL19] Jinru Wang and Wei Liu. Wavelet estimations for heteroscedastic super smooth errors. Communications in Statistics: Theory and Methods, 48(10):2356–2371, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1465086.

Wang:2011:MSC

[WLC11] Shuyun Wang, Yihui Luan, and Qin Chang. On model selection consistency of Bayesian method for normal linear models. Communications in Statistics: Theory and Methods, 40(22): 4021–4040, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2017:SIT

[WLC17] Lu Wang, Yong Lin, and John T. Chen. Simultaneous inference for treatment regimes. Communications in Statistics: Theory and Methods, 46(19):9679–9690, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1217017.

Wang:2011:APC

[WLF11] Jiang-Feng Wang, Han-Ying Liang, and Guo-Liang Fan. Asymptotic properties of conditional quantile estimator under left-truncated and α-mixing conditions. Communications in Statistics: Theory and Methods, 40(14):2462–2486, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2014:CCE

[WLHW14] Xuejun Wang, Xiaoqin Li, Shuhe Hu, and Xinghui Wang. On complete convergence for an extended negatively dependent sequence. Communications in Statistics: Theory and Methods, 43(14):2923–2937, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:TLO

[WLL13] Yue-Wen Wang, Li-Yu Liu, and Chen-Tuo Liao. Two-level orthogonal and saturated designs for estimating main effects and certain important interactions. Communications in Statistics: Theory and Methods, 42(24):4411–4416, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2018:RIS

[WLS18] Xiaoguang Wang, Dawei Lu, and Lixin Song. Robust image segmentation via Bayesian type criterion. Communications in Statistics: Theory and Methods, 47(17):4215–4228, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1371756.

Wang:2015:LLE

[WLW15] Hongxia Wang, Jinguan Lin, and Jinde Wang. Local linear estimation for spatiotemporal models based on least absolute deviation. Communications in Statistics: Theory and Methods, 44(7):1508–1522, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2014:CFC

[WLY14a] Dabuxilatu Wang, Pinghui Li, and Masami Yasuda. Construction of fuzzy control charts based on weighted possibilistic mean. Communications in Statistics: Theory and Methods, 43 (15):3186–3207, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2014:ATP

[WLY14b] Kaiyong Wang, Jinguan Lin, and Yang Yang. Asymptotics for tail probability of random sums with a heavy-tailed number and dependent increments. Communications in Statistics: Theory and Methods, 43(10–12):2595–2604, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:2017:PAW

[WLZ17] Zhengyuan Wei, Juan Li, and Xiaoyang Zheng. A probabilistic approach to Wallis’ formula. Communications in Statistics: Theory and Methods, 46(13):6491–6496, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1129418.

Withers:2016:CRE

[WN16a] Christopher S. Withers and Saralees Nadarajah. Conditioning out rare events for exponential families has negligible effect on inference. Communications in Statistics: Theory and Methods, 45(16):4886–4895, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
932807.

Withers:2016:DMF

[WN16b] Christopher S. Withers and Saralees Nadarajah. The distribution of the maximum of a first-order moving average:

The discrete casex. Communications in Statistics: Theory and Methods, 45(16):4729–4744, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
927499.

Withers:2017:DRV

[WN17a] Christopher S. Withers and Saralees Nadarajah. Discrete random vectors generated by Taylor expansions. Communications in Statistics: Theory and Methods, 46(18):9137–9149, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1205622.

Withers:2017:OPG

[WN17b] Christopher S. Withers and Saralees Nadarajah. Orthogonal polynomials generated by random vectors. Communications in Statistics: Theory and Methods, 46(12):6130–6136, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1118512.

Wang:2018:BRC

[WNL18] Xiaoguang Wang, Yi Niu, and Xiaofang Li. Bayesian ROC curve estimation under binormality using an ordinal category likelihood. Communications in Statistics: Theory and Methods, 47(18):4628–4640, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1380830.

Wojtys:2011:PMS

[Woj11] Małgorzata Wojtyś. Post-model-selection method for density estimation. Communications in Statistics: Theory and Methods, 40(17):3082–3098, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2011:WAG

[WP11] Cuiling Wang and Myunghee Cho Paik. A weighting approach for GEE analysis with missing data. Communications in Statistics: Theory and Methods, 40(13):2397–2411, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2016:NPR

[WP16] Litong Wang and Guobing Pan. The non parametric regression estimate with dependent measurement errors. Communications in Statistics: Theory and Methods, 45(10):2998–3010, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.894067.

Wiangnak:2018:GLA

[WP18] Piyachart Wiangnak and Suvra Pal. Gamma lifetimes and associated inference for interval-censored cure rate model with COM-Poisson competing cause. Communications in Statistics: Theory and Methods, 47(6):1491–1509, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321769.

Wu:2019:CCS

[WP19] Yongfeng Wu and JiangYan Peng. Complete convergence and the strong laws of large numbers for pairwise NQD random variables. Communications in Statistics: Theory and Methods, 48(15):3864–3875, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481979.

Wang:2014:ELB

[WQ14] Binhuan Wang and Gengsheng Qin. Empirical likelihood-based confidence intervals for the sensitivity of a continuous-scale diagnostic test with missing data. Communications in Statistics: Theory and Methods, 43(15):3248–3268, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2018:UPR

[WQ18] Zhenghong Wang and Hong Qin. Uniformity pattern and related criteria for mixed-level designs. Communications in Statistics: Theory and Methods, 47(13):3192–3203, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1350274.

Wang:2016:HCC

[WQW16] Wei Wang, Linyi Qian, and Wensheng Wang. Hedging of contingent claims written on non traded assets under Markovmodulated models. Communications in Statistics: Theory and Methods, 45(12):3577–3595, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
904355.

Wang:2010:ENE

[WR10] Yuanjia Wang and Daniel Rabinowitz. Efficient nonparametric estimation from kin-cohort data. Communications in Statistics: Theory and Methods, 39(20):3622–3634, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2019:MVP

[WRZ19] Suxin Wang, Ximin Rong, and Hui Zhao. Mean-variance problem for an insurer with default risk under a jump-diffusion risk model. Communications in Statistics: Theory and Methods, 48(17):4221–4249, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1490432.

Wang:2011:ALV

[WS11] Xiaoguang Wang and Lixin Song. Adaptive lasso variable selection for the accelerated failure models. Communications in Statistics: Theory and Methods, 40(24):4372–4386, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2012:AMP

[WS12a] Lichun Wang and Radhey S. Singh. An application of matrix power series to linear models. Communications in Statistics: Theory and Methods, 41(8):1315–1333, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2012:SMP

[WS12b] Qiang Wu and Allan R. Sampson. Structured modeling for post-mortem brain tissue data. Communications in Statistics: Theory and Methods, 41(7):1194–1213, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:OPP

[WS13] Guojun Wang and Siva Sivaganesan. Objective priors for parameters in a normal linear regression with measurement error. Communications in Statistics: Theory and Methods, 42(15): 2694–2713, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2014:BFC

[WS14] Min Wang and Xiaoqian Sun. Bayes factor consistency for oneway random effects model. Communications in Statistics: Theory and Methods, 43(23):5072–5090, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:TPS

[WSC13] Xiaoguang Wang, Lixin Song, and Jie Cui. Tuning parameter selector for the penalized likelihood method in multivariate generalized linear models. Communications in Statistics: Theory and Methods, 42(21):3873–3888, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2014:PLI

[WSK14] Xiaoguang Wang, Lixin Song, and Xiaoning Kang. Profile likelihood inferences on the partially linear model with a diverging number of parameters. Communications in Statistics: Theory and Methods, 43(1):13–27, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:SPL

[WSlT15] Mingqiu Wang, Lixin Song, and Guo liang Tian. SCADpenalized least absolute deviation regression in high-dimensional models. Communications in Statistics: Theory and Methods, 44(12):2452–2472, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2017:SAC

[WSS17] Min Wu, Yimin Shi, and Yudong Sun. Statistical analysis for competing risks model from a Weibull distribution under progressively hybrid censoring. Communications in Statistics: Theory and Methods, 46(1):75–86, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.985838.

Watanabe:2010:CAS

[WST10] Tsutomu Watanabe, Hiroshi Shiraishi, and Masanobu Taniguchi. Cluster analysis for stable processes. Communications in Statistics: Theory and Methods, 39(8–9):1630–1642, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2011:BEP

[WSW11] Mingqiu Wang, Lixin Song, and Xiaoguang Wang. Bridge estimators in the partially linear model with high dimensionality. Communications in Statistics: Theory and Methods, 40(24): 4325–4346, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2016:CFL

[WSW16] Suxin Wang, Shiyu Song, and Yongjin Wang. Closed-form likelihood estimation for one type of affine point processes. Communications in Statistics: Theory and Methods, 45(19):5818– 5825, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.950749.

Wang:2015:LEC

[WT15] Yunyan Wang and Mingtian Tang. Local M-estimation for conditional variance in heteroscedastic regression models. Communications in Statistics: Theory and Methods, 44(1):48–62, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2018:ASP

[WT18] Fu-Kwun Wang and Yeneneh Tamirat. Acceptance sampling plan based on an exponentially weighted moving average statistic with the yield index for autocorrelation between polynomial profiles. Communications in Statistics: Theory and Methods, 47(19):4859–4871, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1454960.

Wang:2011:CTE

[WTB11] Xikui Wang, Qier Tan, and Mikelis G. Bickis. Clinical trials with exponential survival times. Communications in Statistics: Theory and Methods, 40(8):1510–1524, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:EPO

[WTZ15] Yinfeng Wang, Yanlin Tang, and Xinsheng Zhang. Estimation of parameters of the Ornstein–Uhlenbeck type processes with continuum of moment conditions. Communications in Statistics: Theory and Methods, 44(24):5189–5203, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:RAL

[WTZS15] Peijie Wang, Xingwei Tong, Shishun Zhao, and Jianguo Sun. Regression analysis of left-truncated and Case I intervalcensored data with the additive hazards model. Communications in Statistics: Theory and Methods, 44(8):1537–1551, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2011:JTC

[Wu11] Jianhong Wu. A joint test for conditional heteroscedasticity in dynamic panel data models. Communications in Statistics: Theory and Methods, 40(8):1434–1444, 2011. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2013:REM

[Wu13] Jianhong Wu. Robust estimation of moments in dynamic panel models with potential intercorrelation. Communications in Statistics: Theory and Methods, 42(23):4199–4209, 2013. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2016:RRM

[Wu16a] Chien-Hua Wu. Randomized response model versus mixture model in the crossover design for clinical trials. Communications in Statistics: Theory and Methods, 45(15):4611–4627, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.927490.

Wu:2016:PPA

[Wu16b] Jibo Wu. On the predictive performance of the almost unbiased Liu estimator. Communications in Statistics: Theory and Methods, 45(17):5193–5203, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
941498.

Wu:2016:SCE

[Wu16c] Jibo Wu. Superiority of the r–k class estimator over some estimators in a misspecified linear model. Communications in Statistics: Theory and Methods, 45(5):1453–1458, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.863934.

Wu:2016:IAT

[Wu16d] Yanhong Wu. Inference after truncated one-sided sequential test. Communications in Statistics: Theory and Methods, 45(10):3076–3094, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
894768.

Wu:2017:QME

[Wu17a] Jibo Wu. Quasi-minimax estimation in the partial linear model. Communications in Statistics: Theory and Methods, 46(6): 2982–2989, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1053941.

Wu:2017:SSP

[Wu17b] Jibo Wu. The small sample properties of the restricted principal component regression estimator in linear regression model. Communications in Statistics: Theory and Methods, 46(4): 1661–1667, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1024867.

Wu:2017:ASC

[Wu17c] Qunying Wu. Almost sure central limit theorem for selfnormalized products of partial sums of negatively associated sequences. Communications in Statistics: Theory and Methods, 46(6):2593–2606, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1006786.

Wu:2017:EAS

[Wu17d] Qunying Wu. An extension of almost sure central limit theorem for the maximum of stationary Gaussian random fields. Communications in Statistics: Theory and Methods, 46(8):3667– 3675, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1069353.

Wu:2018:IGD

[Wu18a] Jibo Wu. Improvement of generalized difference-based mixed Liu estimator in partially linear model. Communications in Statistics: Theory and Methods, 47(18):4433–4442, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1376086. See comments [Li19].

Wu:2018:PDB

[Wu18b] Jibo Wu. Performance of the difference-based Liu-type estimator in partially linear model. Communications in Statistics: Theory and Methods, 47(12):2979–2987, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1346804.

Wu:2019:CSC

[Wu19] Yanhong Wu. A combined SR-CUSUM procedure for detecting common changes in panel data. Communications in Statistics: Theory and Methods, 48(17):4302–4319, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1494285.

Wu:2016:CMC

[WV16] Yongfeng Wu and Andrei Volodin. Complete moment convergence of weighted sums for arrays of negatively dependent random variables and its applications. Communications in Statistics: Theory and Methods, 45(11):3185–3195, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901365.

Wang:2010:PVO

[WW10] Wei Wang and Wensheng Wang. Pricing vulnerable options under a Markov-modulated regime switching model. Communications in Statistics: Theory and Methods, 39(19):3421–3433, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wen:2011:CEG

[WW11] Limin Wen and Xianyi Wu. The credibility estimator with general dependence structure over risks. Communications in Statistics: Theory and Methods, 40(10):1893–1910, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:LLR

[WW13a] Hongxia Wang and Jinde Wang. Local linear regression for non grid spatiotemporal models with autoregressive errors. Communications in Statistics: Theory and Methods, 42(23):4259–4275, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:TFD

[WW13b] P. C. Wang and Soushan Wu. Trend-free designs in blocked factorial experiments. Communications in Statistics: Theory and Methods, 42(10):1870–1878, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:PLD

[WW13c] Shijie Wang and Wensheng Wang. Precise large deviations for sums of random variables with consistent variation in dependent multi-risk models. Communications in Statistics: Theory and Methods, 42(24):4444–4459, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:BIA

[WW15a] Hai-Bin Wang and Ping Wu. Bayesian inference of autoregressive and functional-coefficient moving average models. Communications in Statistics: Theory and Methods, 44(3):453–467, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:SRTb

[WW15b] Qin Wang and Rongning Wu. Semiparametric regression for time series of counts. Communications in Statistics: Theory and Methods, 44(5):983–995, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Worms:2015:TCT

[WW15c] Julien Worms and Rym Worms. A test for comparing tail indices for heavy-tailed distributions via empirical likelihood. Communications in Statistics: Theory and Methods, 44(15): 3289–3302, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2017:EUD

[WW17a] Lichun Wang and Haocheng Wang. On estimating uniform distribution via linear Bayes method. Communications in Statistics: Theory and Methods, 46(19):9450–9458, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1212075.

Wesolowski:2017:EAO

[WW17b] Jacek Wesolowski and Robert Wieczorkowski. An eigenproblem approach to optimal equal-precision sample allocation in subpopulations. Communications in Statistics: Theory and Methods, 46(5):2212–2231, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1040501.

Wang:2019:CCA

[WW19] Wenjuan Wang and Qunying Wu. Complete convergence for arrays of row-wise ND random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 48(13):3165–3176, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1476711.

Wang:2015:BEB

[WWH15a] Wensheng Wang, Ka-Ho Wu, and Kyo-Shin Hwang. The Berry–Esséen bounds for sample rescaled poly-variograms. Communications in Statistics: Theory and Methods, 44(17): 3639–3652, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:2015:IQF

[WWH15b] Lai Wei, Dongliang Wang, and Alan D. Hutson. An investigation of quantile function estimators relative to quantile confidence interval coverage. Communications in Statistics: Theory and Methods, 44(10):2107–2135, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:ELC

[WWL13] Guannan Wang, Zhizhong Wang, and Xiaohui Liu. Empirical likelihood for censored partial linear model based on imputed value. Communications in Statistics: Theory and Methods, 42 (4):644–659, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:ELI

[WWT15] Guannan Wang, Zhizhong Wang, and Ye Tian. The empirical likelihood inference of a regression parameter in censored partial linear models based on a piecewise polynomial. Communications in Statistics: Theory and Methods, 44(12):2431–2451, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wen:2013:NCC

[WWW13] Limin Wen, Jiangfeng Wang, and Xianyi Wu. A new class of credibility estimators under the generalized weighted premium principle. Communications in Statistics: Theory and Methods, 42(3):447–465, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2017:EIS

[WWW17] Xuejun Wang, Shijie Wang, and Rui Wang. Exponential inequalities for sums of unbounded φ-mixing sequence and their applications. Communications in Statistics: Theory and Methods, 46(1):457–464, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
995827.

Wang:2019:OIC

[WWWX19] Hao Wang, Rongming Wang, Jiaqin Wei, and Shaosheng Xu. Optimal investment-consumption-insurance strategy in a continuous-time self-exciting threshold model. Communications in Statistics: Theory and Methods, 48(14):3530–3548, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1477161.

Wang:2017:IKD

[WWY17] Bing Xing Wang, Xiu Kun Wang, and Keming Yu. Inference on the Kumaraswamy distribution. Communications in Statistics: Theory and Methods, 46(5):2079–2090, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1032425.

Wan:2018:EDI

[WWZC18] Qiang Wan, Yongzhong Wu, Wenhui Zhou, and Xiaohong Chen. Economic design of an integrated adaptive synthetic X̄ chart and maintenance management system. Communications in Statistics: Theory and Methods, 47(11):2625–2642, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1271425.

Wang:2015:BTP

[WX15] Li Wang and Xingzhong Xu. Bonferroni-type plug-in procedure controlling generalized familywise error rate. Communications in Statistics: Theory and Methods, 44(14):3042–3055, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2017:CVS

[WX17] Li Wang and Xingzhong Xu. Consistent variable selection via the optimal discovery procedure in multiple testing. Communications in Statistics: Theory and Methods, 46(13):6303–6322, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1069351.

Wang:2019:VES

[WX19] Zhaoliang Wang and Liugen Xue. Variance estimation for sparse ultra-high dimensional varying coefficient models. Communications in Statistics: Theory and Methods, 48(5):1270– 1283, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1429627.

Wang:2013:DSW

[WXPL13] Lei Wang, Dongdong Xiang, Xiaolong Pu, and Yan Li. A double sequential weighted probability ratio test for one-sided composite hypotheses. Communications in Statistics: Theory and Methods, 42(20):3678–3695, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2012:PIS

[WXPS12] Ronghua Wang, Xiaoling Xu, Rong Pan, and Naijun Sha. On parameter inference for step-stress accelerated life test with geometric distribution. Communications in Statistics: Theory and Methods, 41(10):1796–1812, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2018:RAR

[WXS18] Xin Wang, Xiaoming Xue, and Liuquan Sun. Regression analysis of restricted mean survival time based on pseudoobservations for competing risks data. Communications in Statistics: Theory and Methods, 47(22):5614–5625, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1397174.

Wang:2013:GCI

[WXZ13] Bing Xing Wang, Min Xie, and Jun Xing Zhou. Generalized confidence interval for the scale parameter of the power-law process. Communications in Statistics: Theory and Methods, 42 (5):898–906, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wan:2016:ULR

[WXZ16] Shuwen Wan, Binrong Xu, and Biao Zhang. Using logistic regression for semiparametric comparison of population means and variances. Communications in Statistics: Theory and Methods, 45(9):2485–2503, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.882169.

Wu:2018:QVS

[WXZ18] Jian Wu, Liugen Xue, and Peixin Zhao. Quickly variable selection for varying coefficient models with missing response at random. Communications in Statistics: Theory and Methods, 47(10):2327–2336, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
922989.

Wang:2011:MDR

[WY11] Yinfeng Wang and Chuancun Yin. Minimum of dependent random variables with convolution-equivalent distributions. Communications in Statistics: Theory and Methods, 40(18):3245– 3251, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2012:CGM

[WY12] Jiaping Wang and Qiqing Yu. Consistency of the generalized MLE with interval-censored and masked competing risks data. Communications in Statistics: Theory and Methods, 41(23): 4360–4377, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2013:TSR

[WY13] Jibo Wu and Hu Yang. Two stochastic restricted principal components regression estimator in linear regression. Communications in Statistics: Theory and Methods, 42(20):3793–3804, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2016:EMG

[WY16] Xing Wu and Conglian Yu. Estimation of the mixtures of GLMs with covariate-dependent mixing proportions. Communications in Statistics: Theory and Methods, 45(24):7242–7257, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.975822.

Wu:2017:DMD

[WYG17] Jianzhang Wu, Xiao Yu, and Wei Gao. Disequilibrium multidividing ontology learning algorithm. Communications in Statistics: Theory and Methods, 46(18):8925–8942, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1197254.

Wu:2012:NMR

[WYL12] Xiaoyue Wu, Hua Yan, and Lirong Li. Numerical method for reliability analysis of Phased-mission system using Markov chains. Communications in Statistics: Theory and Methods, 41(21):3960–3973, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wylupek:2013:DDT

[Wyl13] Grzegorz Wylupek. Data-driven tests for trend. Communications in Statistics: Theory and Methods, 42(7):1406–1427, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wen:2017:CPB

[WYMZ17] Limin Wen, Jun Yu, Guoping Mei, and Yi Zhang. The credibility premiums based on estimated moment-generating function. Communications in Statistics: Theory and Methods, 46(3): 1090–1106, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1010010.

Wysocki:2019:SAC

[Wys19] Włodzimierz Wysocki. Sampling from Archimedean n-copulas. Communications in Statistics: Theory and Methods, 48(23): 5682–5700, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1520880.

Wei:2016:LSD

[WYY16] Meng Wei, Guangyu Yang, and Lingling Yang. The limiting spectral distribution for large sample covariance matrices with unbounded m-dependent entries. Communications in Statistics: Theory and Methods, 45(22):6651–6662, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.963621.

Wichitchan:2019:SRS

[WYY19] Supawadee Wichitchan, Weixin Yao, and Guangren Yang. A simple root selection method for univariate finite normal mixture models. Communications in Statistics: Theory and Methods, 48(15):3778–3794, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1481972.

Wang:2015:FFS

[WYZ15] Jianxin Wang, Ye Yuan, and Shengli Zhao. Fractional factorial split-plot designs with two- and four-level factors containing clear effects. Communications in Statistics: Theory and Methods, 44(4):671–682, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2010:NDP

[WZ10] Yashi Wang and Peng Zhao. A note on DRHR preservation property of generalized order statistics. Communications in Statistics: Theory and Methods, 39(5):815–822, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2013:IEF

[WZ13] Haiying Wang and Sherry Z. F. Zhou. Interval estimation by frequentist model averaging. Communications in Statistics: Theory and Methods, 42(23):4342–4356, 2013. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2015:SRA

[WZ15] Xueli Wang and Xiao-Hua Zhou. Some results about standardization for a non confounder in estimators of (log) relative risk. Communications in Statistics: Theory and Methods, 44(7):1497–1507, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2016:GOD

[WZ16] Xiaodi Wang and Yingshan Zhang. General orthogonal designs for parameter estimation of ANOVA models under weighted sum-to-zero constraints. Communications in Statistics: Theory and Methods, 45(21):6229–6244, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.957860.

Wu:2019:CDM

[WZ19] Sheng Wu and Yi Zhang. A class of distortion measures generated from expectile and its estimation. Communications in Statistics: Theory and Methods, 48(10):2390–2408, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1465085.

Wang:2019:PLL

[WZJH19] Yige Wang, Nan Zhang, Zhuo Jin, and Tin Long Ho. Pricing longevity-linked derivatives using a stochastic mortality model. Communications in Statistics: Theory and Methods, 48(24): 5923–5942, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1563171.

Wang:2017:WED

[WZK17] Jinru Wang, Qingqing Zhang, and Junke Kou. Wavelet estimators for the derivatives of the density function from data contaminated with heteroscedastic measurement errors. Communications in Statistics: Theory and Methods, 46(15):7337–7354, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1148735.

Wei:2010:SIC

[WZL10] Zhengyuan Wei, Xinsheng Zhang, and Taifu Li. On Stein identity, Chernoff inequality, and orthogonal polynomials. Communications in Statistics: Theory and Methods, 39(14):2573–2593, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2015:REC

[WZT15] Qingtai Wu, Jin Zhang, and Jiashan Tang. Reliability evaluation for a class of multi-unit cold standby systems under Poisson shocks. Communications in Statistics: Theory and Methods, 44(15):3278–3288, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:2017:NPC

[WZX17] Dabuxilatu Wang, Liang Zhang, and Qiang Xiong. A non parametric CUSUM control chart based on the Mann–Whitney statistic. Communications in Statistics: Theory and Methods, 46(10):4713–4725, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1073314.

Wang:2014:WAC

[WZY14] Jingle Wang, Ming Zheng, and Wen Yu. Wavelet analysis of change points in nonparametric hazard rate models under random censorship. Communications in Statistics: Theory and Methods, 43(9):1956–1978, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:2017:NTS

[WZZ17] Xinqi Wu, Sanguo Zhang, and Qingzhao Zhang. A note on the two-sample mean problem based on jackknife empirical likelihood. Communications in Statistics: Theory and Methods, 46(16):7827–7836, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1024864.

Xu:2019:PDB

[XDW19] Chao Xu, Yinghui Dong, and Guojing Wang. The pricing of defaultable bonds under a regime-switching jump-diffusion model with stochastic default barrier. Communications in Statistics: Theory and Methods, 48(9):2185–2205, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1459715.

Xi:2018:CCC

[XDWC18] Mengmei Xi, Xin Deng, Xuejun Wang, and Zhaoyang Cheng. L^{p} convergence and complete convergence for weighted sums of AANA random variables. Communications in Statistics: Theory and Methods, 47(22):5604–5613, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1397173.

Xu:2012:MCS

[XF12] Hai-Yan Xu and Heliang Fei. Models comparison for step-stress accelerated life testing. Communications in Statistics: Theory and Methods, 41(21):3878–3887, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2011:DBA

[XFC11] Haiyan Xu, Heliang Fei, and Zhenmin Chen. Design of bivariate accelerated life tests with one main effect and one interaction effect for log-location-scale distributions. Communications in Statistics: Theory and Methods, 40(15):2781–2798, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2019:SIV

[XFWC19] Hong-Xia Xu, Guo-Liang Fan, Cheng-Xin Wu, and Zhen-Long Chen. Statistical inference for varying-coefficient partially linear errors-in-variables models with missing data. Communications in Statistics: Theory and Methods, 48(22):5621–5636, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1517216.

Xia:2011:CMM

[XGZ11] Zhiming Xia, Pengjiang Guo, and Wenzhi Zhao. CUSUM methods for monitoring structural changes in structural equations. Communications in Statistics: Theory and Methods, 40(6):1109–1123, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xie:2015:RMD

[XGZ15] Jie-Hua Xie, Jian-Wei Gao, and Wei Zou. On a risk model with delayed claims under stochastic interest rates. Communications in Statistics: Theory and Methods, 44(14):3022–3041, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2017:MGF

[XHH17] Nuo Xu, Xuan Huang, and Samuel Huang. A measure of general functional dependence between two continuous variables. Communications in Statistics: Theory and Methods, 46(9): 4327–4352, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1081951.

Xiaozhi:2018:APE

[XHL18a] Peng Xiaozhi, Wu Hecheng, and Ma Ling. Asymptotic properties of the estimators of the semi-parametric spatial regression model. Communications in Statistics: Theory and Methods, 47(7):1663–1678, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1324983.

Xiong:2018:SIS

[XHL18b] Qiang Xiong, Zhiyong Hu, and Yuan Li. Statistic inference for a single-index ARCH-m model. Communications in Statistics: Theory and Methods, 47(1):102–117, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1300283.

Xian:2016:OCD

[XHY16] Jin-Guo Xian, Dong Han, and Jian-Qi Yu. Online change detection of Markov chains with unknown post-change transition probabilities. Communications in Statistics: Theory and Methods, 45(3):597–611, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xia:2017:VSL

[Xia17a] Dong Xia. Variable selection of linear programming discriminant estimator. Communications in Statistics: Theory and Methods, 46(7):3321–3341, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1060344.

Xiao:2017:GCR

[Xia17b] Qing Xiao. Generating correlated random vector involving discrete variables. Communications in Statistics: Theory and Methods, 46(4):1594–1605, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1024860.

Xie:2016:LTM

[Xie16] Junshan Xie. A limit theorem on moment convergence of centered spectral statistics of random matrices. Communications in Statistics: Theory and Methods, 45(24):7119–7129, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.978022.

Xing:2010:RPC

[Xin10] Yang Xing. Rates of posterior convergence for iid observations. Communications in Statistics: Theory and Methods, 39(19): 3389–3398, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xing:2015:UPD

[Xin15] Yang Xing. A universal prior distribution for Bayesian consistency of non parametric procedures. Communications in Statistics: Theory and Methods, 44(5):972–982, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiong:2016:OGS

[Xio16] Wenjun Xiong. The optimal group size using inverse binomial group testing considering misclassification. Communications in Statistics: Theory and Methods, 45(15):4600–4610, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.923461.

Xia:2017:DQL

[XJW17] Tian Xia, Xuejun Jiang, and Xueren Wang. Diagnostics for quasi-likelihood nonlinear models. Communications in Statistics: Theory and Methods, 46(18):8836–8851, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1193201.

Xia:2019:APA

[XJW19] Tian Xia, Xuejun Jiang, and Xueren Wang. Asymptotic properties of approximate maximum quasi-likelihood estimator in quasi-likelihood nonlinear models with random effects. Communications in Statistics: Theory and Methods, 48(8):1890–1901, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1440311.

Xu:2018:ODA

[XK18] Xiaojian Xu and Mark Krzeminski. Optimal designs for accelerated life tests with multiple stress factors and heteroscedasticity. Communications in Statistics: Theory and Methods, 47(6): 1273–1306, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1317809.

Xue:2019:PES

[XKW19] Hongqi Xue, Arun Kumar, and Hulin Wu. Parameter estimation for semiparametric ordinary differential equation models.

Communications in Statistics: Theory and Methods, 48(24): 5985–6004, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1523433.

Xiao:2013:FTR

[XLL13] Hong Min Xiao, Ze Hui Li, and Wei Wei Liu. The finite time ruin probability of a new risk model based on entrance process. Communications in Statistics: Theory and Methods, 42(2):336– 345, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xia:2015:BAA

[XLL15a] Qiang Xia, Rubing Liang, and Jinshan Liu. A Bayesian analysis of autoregressive models with exogenous variables and powertransformed and threshold GARCH errors. Communications in Statistics: Theory and Methods, 44(9):1967–1980, 2015. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2015:PRM

[XLL15b] Zhibin Xu, Luqin Liu, and Yanyan Liu. Proportional rate model with incomplete covariate and complete auxiliary information. Communications in Statistics: Theory and Methods, 44(24): 5285–5305, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xia:2012:BAT

[XLPL12] Qiang Xia, Jinshan Liu, Jiazhu Pan, and Rubing Liang. Bayesian analysis of two-regime threshold autoregressive moving average model with exogenous inputs. Communications in Statistics: Theory and Methods, 41(6):1089–1104, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiao:2019:ESP

[XLS19] Juxia Xiao, Xu Li, and Jianhong Shi. Estimation in a semiparametric partially linear errors-in-variables model with inverse Gaussian kernel. Communications in Statistics: Theory and Methods, 48(17):4394–4424, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1496255.

Xu:2012:SSR

[XLZ^{+}12] Ying Xu, K. F. Lam, Feifei Zhou, Paul S. F. Yip, and Ray Watson. A study of suicide risk using a Cox cure model via a retrospective sampling and multiple imputation. Communications in Statistics: Theory and Methods, 41(18):3389–3402, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2014:ASC

[XM14] Shoufang Xu and Yu Miao. Almost sure convergence of weighted sums for negatively associated random variables. Communications in Statistics: Theory and Methods, 43(10–12): 2581–2594, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2017:FFM

[XM17a] Jing Xu and Jun Ma. Fitting finite mixture models using iterative Monte Carlo classification. Communications in Statistics: Theory and Methods, 46(13):6684–6693, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1132329.

Xudong:2017:CIS

[XM17b] Huang Xudong and Li Mengmeng. Confidence intervals for sparse precision matrix estimation via Lasso penalized Dtrace loss. Communications in Statistics: Theory and Methods, 46(24):12299–12316, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1295074.

Xu:2016:QSF

[XML16] Jie Xu, Yu Miao, and Jicheng Liu. Quasi-sure functional limit theorem for increments of a fractional Brownian sheet in Hölder norm. Communications in Statistics: Theory and Methods, 45(5):1564–1574, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1109662.

Xu:2017:NPH

[XMP17] Jing Xu, Jun Ma, and Tania Prvan. Non parametric hazard estimation with dependent censoring using penalized likelihood and an assumed copula. Communications in Statistics: Theory and Methods, 46(22):11383–11403, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1267757.

Xu:2019:MPW

[XN19] Haifeng Xu and Akio Namba. MSE performance of the weighted average estimators consisting of shrinkage estimators when each individual regression coefficient is estimated. Communications in Statistics: Theory and Methods, 48(13):3280– 3290, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1475569.

Xu:2019:PPT

[XO19] Haifeng Xu and Kazuhiro Ohtani. PMSE performance of two different types of preliminary test estimators under a multivariate t error term. Communications in Statistics: Theory and Methods, 48(17):4320–4338, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1494286.

Xu:2018:LTM

[XP18] Cheng Xu and Tianxiao Pang. Limit theory for moderate deviations from a unit root with a break in variance. Communications in Statistics: Theory and Methods, 47(24):6125–6143, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1406515.

Xin:2017:MRR

[XQ17a] Xin Xin and Feng Qiu. Multiple-response repeated measurement or multivariate growth curve model with distributionfree errors. Communications in Statistics: Theory and Methods, 46(24):12370–12386, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1300273.

Xing:2017:BER

[XQ17b] Ji-Ji Xing and Xi-Yuan Qian. Bayesian expectile regression with asymmetric normal distribution. Communications in Statistics: Theory and Methods, 46(9):4545–4555, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1088030.

Xia:2019:VSS

[XQS19] Yafeng Xia, Yarong Qu, and Nailing Sun. Variable selection for semiparametric varying coefficient partially linear model based on modal regression with missing data. Communications in Statistics: Theory and Methods, 48(20):5121–5137, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1508712.

Xu:2011:NBA

[XT11] Ancha Xu and Yincai Tang. Nonparametric Bayesian analysis of competing risks problem with masked data. Communications in Statistics: Theory and Methods, 40(13):2326–2336, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2012:OBA

[XT12] Ancha Xu and Yincai Tang. Objective Bayesian analysis for linear degradation models. Communications in Statistics: Theory and Methods, 41(21):4034–4046, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2017:PBIb

[XT17] Liwen Xu and Maozai Tian. Parametric bootstrap inferences for panel data models. Communications in Statistics: Theory and Methods, 46(11):5579–5594, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1105981.

Xiao:2015:ELP

[XTG15] Yanting Xiao, Zheng Tian, and Wenyan Guo. Empirical likelihood for partially non linear models with missing response variables at random. Communications in Statistics: Theory and Methods, 44(16):3523–3540, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xi:2017:CMC

[XTL^{+}17] Mengmei Xi, Xufei Tang, Junjie Lin, Ming Wang, and Xuejun Wang. Complete moment convergence and L_{r} convergence for asymptotically almost negatively associated random variables. Communications in Statistics: Theory and Methods, 46(20):10373–10386, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1235198.

Xiao:2016:ELD

[XTS16] Yanting Xiao, Zheng Tian, and Jin Sun. Empirical likelihood dimension reduction inference for partially non-linear error-in-responses models with validation data. Communications in Statistics: Theory and Methods, 45(24):7103–7118, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.978021.

Xiong:2017:RQR

[XTT17] Wei Xiong, Maozai Tian, and Man-Lai Tang. Randomized quantile regression estimation for heteroskedastic non parametric model. Communications in Statistics: Theory and Methods, 46(10):5147–5179, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1096393.

Xu:2010:AEN

[Xu10] Jin Xu. Asymptotic expansion of the non null distribution of the two-sample t-statistic under non normality with application in power comparison. Communications in Statistics: Theory and Methods, 39(11):1915–1921, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2012:EGF

[Xu12] Haifeng Xu. The exact general formulas for the moments of a ridge regression estimator when the regression error terms follow a multivariate t distribution. Communications in Statistics: Theory and Methods, 41(15):2788–2802, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2013:MPM

[Xu13] Haifeng Xu. MSE performance and minimax regret significance points for a HPT estimator when each individual regression coefficient is estimated. Communications in Statistics: Theory and Methods, 42(12):2152–2164, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2016:FSP

[Xu16] Haifeng Xu. Finite sample properties of an HPT estimator when each individual regression coefficient is estimated in a misspecified linear regression model. Communications in Statistics: Theory and Methods, 45(2):506–519, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2017:MPM

[Xu17a] Haifeng Xu. MSE performance and minimax regret significance points for a HPT estimator under a multivariate t error distribution. Communications in Statistics: Theory and Methods, 46(7):3123–3134, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1056364.

Xu:2017:PBIa

[Xu17b] Liwen Xu. Parametric bootstrap inferences for the growth curve models with intraclass correlation structure. Communications in Statistics: Theory and Methods, 46(7):3308–3320, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1060343.

Xuemei:2017:IMS

[Xue17] Hu Xuemei. The indirect method for stochastic logistic growth models. Communications in Statistics: Theory and Methods, 46(3):1506–1518, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1019152.

Xiang:2016:SHC

[XW16] Fei Xiang and Stephen G. Walker. Sup-Hellinger consistency for local density regression. Communications in Statistics: Theory and Methods, 45(5):1329–1344, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.863927.

Xia:2010:APM

[XWJ10] Tian Xia, Xue-Ren Wang, and Xue-Jun Jiang. Asymptotic properties of the MLE in nonlinear reproductive dispersion models with stochastic regressors. Communications in Statistics: Theory and Methods, 39(15):2800–2810, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xie:2017:LRT

[XX17] Junshan Xie and Nannan Xiao. The likelihood ratio test for high-dimensional linear regression model. Communications in Statistics: Theory and Methods, 46(17):8479–8492, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1183785.

Xiao:2018:ARP

[XX18] Hongmin Xiao and Lin Xie. Asymptotic ruin probabilities of a dependent renewal risk model based on entrance processes with constant interest rate. Communications in Statistics: Theory and Methods, 47(22):5396–5417, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1390137.

Xing:2018:CMG

[XXMZ18] Yanchun Xing, Lili Xu, Wenqing Ma, and Zhichuan Zhu. Conditional mix-GEE models for longitudinal data with unspecified random-effects distributions. Communications in Statistics: Theory and Methods, 47(4):862–876, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1267763.

Xing:2009:EIS

[XY09] Guodong Xing and Shanchao Yang. An exponential inequality for strictly stationary and negatively associated random variables. Communications in Statistics: Theory and Methods, 39 (2):340–349, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xing:2010:PRL

[XY10] Guo-Dong Xing and Shan-Chao Yang. Precise rates in the law of the iterated logarithm for associated random variables. Communications in Statistics: Theory and Methods, 40(1):130– 144, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2011:MBV

[XY11a] Jianwen Xu and Hu Yang. More on the bias and variance comparisons of the restricted almost unbiased estimators. Communications in Statistics: Theory and Methods, 40(22):4053–4064, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2011:AIC

[XY11b] Xiaojian Xu and Wai Kong Yuen. Applications and implementations of continuous robust designs. Communications in Statistics: Theory and Methods, 40(6):969–988, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2012:STL

[XY12] Jianwen Xu and Hu Yang. On the Stein-type Liu estimator and positive-rule Stein-type Liu estimator in multiple linear regression models. Communications in Statistics: Theory and Methods, 41(5):791–808, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:2013:PPR

[XY13] Jianwen Xu and Hu Yang. Performances of the positive-rule Stein-type ridge estimator in a linear regression model with spherically symmetric error distributions. Communications in Statistics: Theory and Methods, 42(3):543–560, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xing:2014:RCS

[XY14] Guo-Dong Xing and Shan-Chao Yang. Rate of convergence of strong law of large numbers for positively associated sequences. Communications in Statistics: Theory and Methods, 43(13): 2752–2765, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xing:2019:BRS

[XY19] Guodong Xing and Shanchao Yang. On the Bahadur representation of sample quantiles for ψ-mixing sequences and its application. Communications in Statistics: Theory and Methods, 48(5):1060–1072, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1423696.

Xu:2012:GCB

[XYH12] Ying Xu, Paul S. F. Yip, and Richard M. Huggins. A generalized chain binomial model with aggregated data. Communications in Statistics: Theory and Methods, 41(18):3325–3338, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xing:2014:SCC

[XYL14] Guo-Dong Xing, Shan-Chao Yang, and Yong-Ming Li. Strong consistency of conditional value-at-risk estimate for φ-mixing samples. Communications in Statistics: Theory and Methods, 43(23):5105–5113, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiuyun:2016:BEP

[XZ16] Peng Xiuyun and Yan Zaizai. Bayesian estimation and prediction for the inverse Weibull distribution under general progressive censoring. Communications in Statistics: Theory and Methods, 45(3):621–635, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xie:2017:EDP

[XZ17] Jie-Hua Xie and Wei Zou. On the expected discounted penalty function for a risk model with dependence under a multi-layer dividend strategy. Communications in Statistics: Theory and Methods, 46(4):1898–1915, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030424.

Xiao:2019:MCC

[XZ19] Qing Xiao and Shaowu Zhou. Matching a correlation coefficient by a Gaussian copula. Communications in Statistics: Theory and Methods, 48(7):1728–1747, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1439962.

Xiong:2018:ANL

[XZA18] Xianzhu Xiong, Peiqin Zhou, and Chen Ailian. Asymptotic normality of the local linear estimation of the conditional density for functional time-series data. Communications in Statistics: Theory and Methods, 47(14):3418–3440, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1359292.

Xu:2010:HTC

[XZD10] Xingzhong Xu, Qian Zhang, and Xiaobo Ding. Hypothesis testing and confidence regions for the mean sojourn time of an M/M/1 queueing system. Communications in Statistics: Theory and Methods, 40(1):28–39, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xia:2019:MGC

[XZD19] Rui Xia, Qiuyue Zhang, and Xiaoyan Deng. Multiscale Gaussian convolution algorithm for estimate of Gaussian mixture model. Communications in Statistics: Theory and Methods, 48(23):5889–5910, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1523431.

Xie:2015:MEW

[XZNL15] Hongmei Xie, Weiwei Zhuang, Keshe Ni, and Wenyu Liu. Multivariate excess wealth ordering of generalized order statistics. Communications in Statistics: Theory and Methods, 44(24): 5146–5160, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yata:2010:IDE

[YA10] Kazuyoshi Yata and Makoto Aoshima. Intrinsic dimensionality estimation of high-dimension, low sample size data with Dasymptotics. Communications in Statistics: Theory and Methods, 39(8–9):1511–1521, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yadav:2013:ESE

[YA13] Manoj K. Yadav and Girdhar G. Agarwal. On estimation of standard error of intra-class correlation coefficient in unbalanced nested designs. Communications in Statistics: Theory and Methods, 42(1):88–97, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yee:2017:SBL

[YA17] Chia Chye Yee and Yves F. Atchadé. On the sparse Bayesian learning of linear models. Communications in Statistics: Theory and Methods, 46(15):7672–7691, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1158837.

Yanez:2017:AIM

[YAA^{+}17] N. David Yanez, Ibrahim Aljasser, Mose Andre, Chengcheng Hu, Michal Juraska, and Thomas Lumley. Assessing the impact of measurement error in modeling change in the absence of auxiliary data. Communications in Statistics: Theory and Methods, 46(6):2667–2680, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1040508.

Yadav:2016:ESE

[Yad16] Manoj K. Yadav. Estimating standard error of intra-class correlation coefficients up to three level unbalanced nested clinical trials. Communications in Statistics: Theory and Methods, 45(22):6688–6699, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
966837.

Yamada:2012:HDE

[Yam12] Takayuki Yamada. High-dimensional Edgeworth expansion of LR statistic for testing circular symmetric covariance structure and its error bound. Communications in Statistics: Theory and Methods, 41(11):1887–1910, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamada:2017:FWL

[Yam17a] Hiroshi Yamada. The Frisch–Waugh–Lovell theorem for the lasso and the ridge regression. Communications in Statistics: Theory and Methods, 46(21):10897–10902, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1252403.

Yamada:2017:SPU

[Yam17b] Hiroshi Yamada. A small but practically useful modification to the Hodrick–Prescott filtering: a note. Communications in Statistics: Theory and Methods, 46(17):8430–8434, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1179764.

Yamada:2018:SLS

[Yam18a] Hiroshi Yamada. Several least-squares problems related to the Hodrick–Prescott filtering. Communications in Statistics: Theory and Methods, 47(5):1022–1027, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1285934.

Yamada:2018:IET

[Yam18b] Takayuki Yamada. Interval estimation in two-group discriminant analysis under heteroscedasticity for large dimension. Communications in Statistics: Theory and Methods, 47(23): 5717–5728, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1400060.

Yang:2012:ANL

[Yan12] Qinglong Yang. Asymptotic normality of LS estimators in the simple linear EV regression model with PA errors. Communications in Statistics: Theory and Methods, 41(23):4276–4284, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2014:EFT

[Yan14] Yang Yang. Estimate for the finite-time ruin probability in the discrete-time risk model with insurance and financial risks.

Communications in Statistics: Theory and Methods, 43(10–12): 2094–2104, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2015:FSO

[Yan15] Fan Yang. First- and second-order asymptotics for the tail distortion risk measure of extreme risks. Communications in Statistics: Theory and Methods, 44(3):520–532, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yan:2016:RGB

[Yan16] Ting Yan. Ranking in the generalized Bradley–Terry models when the strong connection condition fails. Communications in Statistics: Theory and Methods, 45(2):340–353, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yanagisawa:2017:TCP

[Yan17a] Yukio Yanagisawa. Tests for compound periodicities and estimating a non linear function. Communications in Statistics: Theory and Methods, 46(8):3676–3689, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1069354.

Yang:2017:USC

[Yan17b] Xin Yang. Uniform strong consistency of histogram density estimation for dependent process. Communications in Statistics: Theory and Methods, 46(4):1709–1719, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1026994.

Yan:2018:ASC

[Yan18] Jigao Yan. Almost sure convergence for weighted sums of WNOD random variables and its applications to non parametric regression models. Communications in Statistics: Theory and Methods, 47(16):3893–3909, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1364390.

Yan:2019:CCM

[Yan19] Jigao Yan. On complete convergence in Marcinkiewicz– Zygmund Type SLLN for END random variables and its applications. Communications in Statistics: Theory and Methods, 48(20):5074–5098, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1508709.

Yao:2012:BML

[Yao12] Weixin Yao. Bayesian mixture labeling and clustering. Communications in Statistics: Theory and Methods, 41(3):403–421, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yarovaya:2011:SBR

[Yar11] E. B. Yarovaya. Supercritical branching random walks with a single source. Communications in Statistics: Theory and Methods, 40(16):2926–2945, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yarovaya:2013:BRW

[Yar13] Elena Yarovaya. Branching random walks with heavy tails. Communications in Statistics: Theory and Methods, 42(16): 3001–3010, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yarovaya:2014:OSS

[Yar14] E. B. Yarovaya. Operators satisfying the Schur condition and their applications to the branching random walks. Communications in Statistics: Theory and Methods, 43(7):1523–1532, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2010:NTP

[YC10] Hu Yang and Xinfeng Chang. A new two-parameter estimator in linear regression. Communications in Statistics: Theory and Methods, 39(6):923–934, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2011:SRT

[YC11a] Hu Yang and Juan Cui. A stochastic restricted two-parameter estimator in linear regression model. Communications in Statistics: Theory and Methods, 40(13):2318–2325, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yi:2011:DAU

[YC11b] Seongbaek Yi and Taeryon Choi. A direct approach to understanding posterior consistency of Bayesian regression problems. Communications in Statistics: Theory and Methods, 40(18): 3315–3326, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yan:2012:CNE

[YC12] Li Yan and Xia Chen. Consistency and normality of Mestimators in partly linear models with stochastic adapted errors. Communications in Statistics: Theory and Methods, 41 (9):1557–1568, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2014:CSL

[YC14] Weiming Yang and N. Rao Chaganty. A contrasting study of likelihood methods for the analysis of longitudinal binary data. Communications in Statistics: Theory and Methods, 43(14): 3027–3046, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2017:RWS

[YC17] Changjun Yu and Dongya Cheng. Randomly weighted sums of linearly wide quadrant-dependent random variables with heavy tails. Communications in Statistics: Theory and Methods, 46 (2):591–601, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.1000500.

Yan:2018:PAP

[YC18] Bo Yan and Zhuo Chen. A prediction approach for precise marketing based on ARIMA-ARCH model: a case of China mobile. Communications in Statistics: Theory and Methods, 47(16): 4042–4058, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1380827.

Yen:2018:MDS

[YCAJ18] Ching-Ho Yen, Chia-Hao Chang, Muhammad Aslam, and Chi-Hyuck Jun. Multiple dependent state repetitive sampling plans based on one-sided process capability indices. Communications in Statistics: Theory and Methods, 47(6):1403–1412, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1321121.

Yang:2017:ELC

[YCH17] Guangren Yang, Xia Cui, and Sumin Hou. Empirical likelihood confidence regions in the single-index model with growing dimensions. Communications in Statistics: Theory and Methods, 46(15):7562–7579, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1157190.

Yang:2014:BAS

[YCL14] Ming-Chung Yang, Lee-Shen Chen, and Tachen Liang. A Bayesian approach for selecting the best exponential population with interval censored samples. Communications in Statistics: Theory and Methods, 43(10–12):2357–2369, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yiqi:2017:BEI

[YCL17] Bao Yiqi, Vicente G. Cancho, and Francisco Louzada. On the Bayesian estimation and influence diagnostics for the Weibull-negative-binomial regression model with cure rate under latent failure causes. Communications in Statistics: Theory and Methods, 46(3):1462–1489, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019150.

Yang:2017:ELI

[YCZ17] Yiping Yang, Lifang Chen, and Peixin Zhao. Empirical likelihood inference in partially linear single-index models with endogenous covariates. Communications in Statistics: Theory and Methods, 46(7):3297–3307, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1060341.

Yuan:2014:AIP

[YD14] Ao Yuan and Jan G. De Gooijer. Asymptotically informative prior for Bayesian analysis. Communications in Statistics: Theory and Methods, 43(14):3080–3094, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamada:2019:MWH

[YD19a] Hiroshi Yamada and Ruixue Du. A modification of the Whittaker–Henderson method of graduation. Communications in Statistics: Theory and Methods, 48(15):3795–3800, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1481974.

Yuan:2019:WEL

[YD19b] Xiaohui Yuan and Xiaogang Dong. Weighted empirical likelihood for quantile regression with non ignorable missing covariates. Communications in Statistics: Theory and Methods, 48(12):3068–3084, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473604.

Yu:2015:RAE

[YDC15] Kai Yu, Xin Dang, and Yixin Chen. Robustness of the affine equivariant scatter estimator based on the spatial rank covariance matrix. Communications in Statistics: Theory and Methods, 44(5):914–932, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yi:2014:CAE

[YDL14] Min Yi, Guohua Deng, and Hua Liang. Cramér asymptotic efficiency in a semiparametric model. Communications in Statistics: Theory and Methods, 43(13):2618–2628, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ye:2019:AMM

[Ye19] Qingqing Ye. The analysis of M^{X} /M/1 queue with twostage vacations policy. Communications in Statistics: Theory and Methods, 48(18):4492–4510, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1483514.

Yeh:2012:CGM

[Yeh12] Hsiaw-Chan Yeh. Characterizations of the general multivariate Weibull distributions. Communications in Statistics: Theory and Methods, 41(1):76–87, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yeh:2013:GMW

[Yeh13] Hsiaw-Chan Yeh. General multivariate Weibull processes. Communications in Statistics: Theory and Methods, 42(4):692– 715, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yeh:2015:MWP

[Yeh15] Hsiaw-Chan Yeh. Multivariate Weibull and Pareto distributions in Hilbert space. Communications in Statistics: Theory and Methods, 44(14):3073–3086, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yeh:2017:MSL

[Yeh17] Hsiaw-Chan Yeh. Multivariate semi-α-Laplace distributions. Communications in Statistics: Theory and Methods, 46(15): 7661–7671, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1158835.

Yen:2012:FTO

[Yen12] Ching-Ho Yen. Fuzzy testing for one-sided process capability indices. Communications in Statistics: Theory and Methods, 41(9):1603–1616, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yazdi:2017:CBC

[YF17] Ahmad Ahmadi Yazdi and Mohammad Saber Fallahnezhad. Comparison between count of cumulative conforming sampling plans and Dodge–Romig single sampling plan. Communications in Statistics: Theory and Methods, 46(1):189–199, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.988263.

Yera:2019:VGG

[YFV19] Yoel G. Yera, Carlos A. Fernández, and José E. Valdés. A variant of the Geo/G/1 queues with disasters and general repair times. Communications in Statistics: Theory and Methods, 48(24):6119–6133, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1528368.

Yang:2016:PIP

[YGL16] Hu Yang, Chaohui Guo, and Jing Lv. Penalized inverse probability weighted estimators for weighted rank regression with missing covariates. Communications in Statistics: Theory and Methods, 45(5):1388–1402, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.863930.

Yang:2019:RRM

[YGM19] Tao Yang, Colin M. Gallagher, and Christopher S. McMahan. A robust regression methodology via m-estimation. Communications in Statistics: Theory and Methods, 48(5):1092–1107, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1423698.

Yan:2017:RCM

[YGWQ17] Hua Yan, Li Gao, Kui Wang, and Lei Qi. Reliability computing method combined with compressed storage and Krylov subspace for phased-mission system. Communications in Statistics: Theory and Methods, 46(20):9972–9984, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1228966.

Yumiba:2011:CBF

[YH11] Hiromu Yumiba and Yoshifumi Hyodo. Characterization of balanced fractional 3^{m} factorial designs of resolution III. Communications in Statistics: Theory and Methods, 40(11):2074–2080, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2016:FRP

[YH16] Hu Yang and Jiewu Huang. Further research on the principal component two-parameter estimator in linear model. Communications in Statistics: Theory and Methods, 45(3):566–576, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2019:BLR

[YH19] Youjiao Yu and Jane L. Harvill. Bootstrap likelihood ratio test for Weibull mixture models fitted to grouped data. Communications in Statistics: Theory and Methods, 48(18):4550–4568, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1494838.

Yang:2010:STZ

[YHA10a] Zhao Yang, James W. Hardin, and Cheryl L. Addy. Score tests for zero-inflation in overdispersed count data. Communications in Statistics: Theory and Methods, 39(11):2008–2030, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2010:SRT

[YHA10b] Zhao Yang, James W. Hardin, and Cheryl L. Addy. Some remarks on testing overdispersion in zero-inflated Poisson and binomial regression models. Communications in Statistics: Theory and Methods, 39(15):2743–2752, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2014:ELG

[YHC14] Zhuoxi Yu, Bing He, and Min Chen. Empirical likelihood for generalized partially linear single-index models. Communications in Statistics: Theory and Methods, 43(19):4156–4163, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamada:2013:AAE

[YHS13] Takayuki Yamada, Masashi Hyodo, and Takashi Seo. The asymptotic approximation of EPMC for linear discriminant rules using a Moore–Penrose inverse matrix in high dimension. Communications in Statistics: Theory and Methods, 42(18): 3329–3338, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamada:2017:IED

[YHS17] Takayuki Yamada, Tetsuto Himeno, and Tetsuro Sakurai. Interval estimation in discriminant analysis for large dimension. Communications in Statistics: Theory and Methods, 46(18): 9042–9052, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1202282.

Yang:2017:AAI

[YHW17] Wenzhi Yang, Shuhe Hu, and Xuejun Wang. On the asymptotic approximation of inverse moment for non negative random variables. Communications in Statistics: Theory and Methods, 46(16):7787–7797, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2013.
781648.

Yilmaz:2012:HRP

[Yil12] Mehmet Yilmaz. Hazard rate properties of systems with two components. Communications in Statistics: Theory and Methods, 41(12):2111–2132, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yilmaz:2016:FSL

[Yil16] Abdullah Yilmaz. The flexible skew Laplace distribution. Communications in Statistics: Theory and Methods, 45(23):7053– 7059, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.974821.

Yildiz:2018:PJL

[Yil18] Nilgün Yildiz. On the performance of the jackknifed Liutype estimator in linear regression model. Communications in Statistics: Theory and Methods, 47(9):2278–2290, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1339087.

Yemao:2017:NFD

[YJ17] Xia Yemao and Gou Jianwei. A note on the finite-dimensional Dirichlet prior. Communications in Statistics: Theory and Methods, 46(19):9388–9396, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1208239.

Yamada:2019:EFS

[YJ19] Hiroshi Yamada and Fatima Tuj Jahra. Explicit formulas for the smoother weights of the Whittaker-Henderson graduation of order 1. Communications in Statistics: Theory and Methods, 48(12):3153–3161, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1476713.

Yang:2019:CRW

[YJS19] Rui Yang, Zhaojing Jing, and Yan Shen. On convergence rate for weighted sums of arrays of rowwise ANA random variables. Communications in Statistics: Theory and Methods, 48(4):986– 994, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1422760.

Yang:2019:SBK

[YJSL19] Aijun Yang, Xuejun Jiang, Lianjie Shu, and Pengfei Liu. Sparse Bayesian kernel multinomial probit regression model for high-dimensional data classification. Communications in Statistics: Theory and Methods, 48(1):165–176, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1463385.

Yunus:2011:IPT

[YK11] Rossita M. Yunus and Shahjahan Khan. Increasing the power of the test through pre-test — a robust method. Communications in Statistics: Theory and Methods, 40(4):581–597, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

You:2017:ODE

[YKL^{+}17] H. W. You, Michael B. C. Khoo, M. H. Lee, P. Castagliola, and S. Saha. Optimal design of exponentially weighted moving average-X̄ chart for the mean with estimated process parameters. Communications in Statistics: Theory and Methods, 46(22):11077–11090, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1257716.

Yin:2010:NIL

[YL10] Xiangrong Yin and Bing Li. A note on the invariance law for surrogate dimension reduction. Communications in Statistics:

Theory and Methods, 39(15):2721–2724, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:2015:CIC

[YL15] Demei Yuan and Shunjing Li. From conditional independence to conditionally negative association: Some preliminary results. Communications in Statistics: Theory and Methods, 44(18): 3942–3966, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:2016:SRF

[YL16] Demei Yuan and Lan Lei. Some results following from conditional characteristic functions. Communications in Statistics: Theory and Methods, 45(12):3706–3720, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.906614.

Ye:2017:PAQ

[YL17] Qingqing Ye and Liwei Liu. Performance analysis of queue with two-stage vacation policy. Communications in Statistics: Theory and Methods, 46(17):8647–8665, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1186191.

Yan:2018:OAA

[YL18a] Rongfang Yan and Tianqi Luo. On the optimal allocation of active redundancies in series system. Communications in Statistics: Theory and Methods, 47(10):2379–2388, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1054942.

Yang:2018:WLM

[YL18b] Hu Yang and Ning Li. WLAD-LASSO method for robust estimation and variable selection in partially linear models. Communications in Statistics: Theory and Methods, 47(20):4958– 4976, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1383427.

Ye:2018:AMM

[YL18c] Qingqing Ye and Liwei Liu. Analysis of MAP/M/1 queue with working breakdowns. Communications in Statistics: Theory and Methods, 47(13):3073–3084, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1346808.

You:2018:OSI

[YL18d] Yinping You and Xiaohu Li. Ordering k-out-of-n systems with interdependent components and one active redundancy. Communications in Statistics: Theory and Methods, 47(19):4772– 4784, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2018.1445862.

Yang:2016:REV

[YLG16a] Hu Yang, Jing Lv, and Chaohui Guo. Robust estimation and variable selection for varying-coefficient single-index models based on modal regression. Communications in Statistics: Theory and Methods, 45(14):4048–4067, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.915043.

Yang:2016:RVS

[YLG16b] Hu Yang, Jing Lv, and Chaohui Guo. Robust variable selection and parametric component identification in varying coefficient models. Communications in Statistics: Theory and Methods, 45(18):5533–5549, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
948199.

Yan:2019:RAS

[YLL19] Rongfang Yan, Bin Lu, and Xiaohu Li. On redundancy allocation to series and parallel systems of two components. Communications in Statistics: Theory and Methods, 48(18):4690–4701, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1500603.

Yang:2010:SET

[YLS10] Jun-Qiang Yang, Xuewen Lu, and Radhey S. Singh. Semiparametric estimation for two-sample location-scale models under Type I censorship. Communications in Statistics: Theory and Methods, 39(11):1932–1950, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

You:2017:OAC

[YLS17] Yinping You, Xiaohu Li, and Jairo Santanilla. Optimal allocations of coverage limits for two independent random losses of insurance policy. Communications in Statistics: Theory and Methods, 46(1):497–509, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
997360.

Yan:2017:ARP

[YLW17] Litan Yan, Yumiao Li, and Di Wu. Approximation of the Rosenblatt process by semimartingales. Communications in Statistics: Theory and Methods, 46(9):4556–4578, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1088032.

Yang:2010:CTE

[YLX10] Hu Yang, Wenxue Li, and Jianwen Xu. Comparison of two estimators of parameters under Pitman nearness criterion. Communications in Statistics: Theory and Methods, 39(17):3081– 3094, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yemao:2017:BAC

[YM17] Xia Yemao and Pan Maolin. Bayesian analysis for confirmatory factor model with finite-dimensional Dirichlet prior mixing. Communications in Statistics: Theory and Methods, 46(9): 4599–4619, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1083110.

Yang:2018:SDF

[YM18] Mengta Yang and Reza Modarres. β-skeleton depth functions and medians. Communications in Statistics: Theory and Methods, 47(20):5127–5143, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1386320.

Yao:2015:ACF

[YMF15] Feng Yao and Carlos Martins-Filho. An asymptotic characterization of finite degree U-statistics with sample size-dependent kernels: Applications to nonparametric estimators and test statistics. Communications in Statistics: Theory and Methods, 44(15):3251–3265, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamaguchi:2010:VBA

[YOD10] Y. Yamaguchi, H. Okamura, and T. Dohi. A variational Bayesian approach for estimating parameters of a mixture of Erlang distribution. Communications in Statistics: Theory and Methods, 39(13):2333–2350, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yousefzadeh:2017:BHB

[You17] F. Yousefzadeh. E-Bayesian and hierarchical Bayesian estimations for the system reliability parameter based on asymmetric loss function. Communications in Statistics: Theory and Methods, 46(1):1–8, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
968736.

Yi:2017:CMC

[YQC17] Yanchun Yi, Dehua Qiu, and Pingyan Chen. Complete moment convergence for weighted sums of extended negatively dependent random variables. Communications in Statistics: Theory and Methods, 46(20):10189–10202, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1231814.

Yamada:2012:TMA

[YS12] Takayuki Yamada and Muni S. Srivastava. A test for multivariate analysis of variance in high dimension. Communications in Statistics: Theory and Methods, 41(13–14):2602–2615, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamashita:2015:MAP

[YS15a] Haruka Yamashita and Hideo Suzuki. The methods for approximation of principal points for binary distributions on the basis of submodularity. Communications in Statistics: Theory and Methods, 44(11):2291–2309, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yao:2015:MLR

[YS15b] Weixin Yao and Weixing Song. Mixtures of linear regression with measurement errors. Communications in Statistics: Theory and Methods, 44(8):1602–1614, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2016:PLD

[YS16] Yang Yang and Liwei Sha. Precise large deviations for aggregate claims. Communications in Statistics: Theory and Methods, 45(10):2801–2809, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
887111.

Yan:2019:AEN

[YS19] Ailing Yan and Fengli Song. Adaptive elastic net-penalized quantile regression for variable selection. Communications in Statistics: Theory and Methods, 48(20):5106–5120, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1508711.

Yang:2017:ASD

[YSC17a] Guangren Yang, Yanqing Sun, and Xia Cui. Automatic structure discovery for varying-coefficient partially linear models. Communications in Statistics: Theory and Methods, 46(15): 7703–7716, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1161796.

Yu:2017:PEO

[YSC17b] Qian Yu, Guangjun Shen, and Mingxiang Cao. Parameter estimation for Ornstein–Uhlenbeck processes of the second kind driven by α-stable Lévy motions. Communications in Statistics: Theory and Methods, 46(21):10864–10878, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1248786.

Yang:2016:NSL

[YSD^{+}16] Zheng-Ling Yang, Yan-Wen Song, Zhi-Feng Duan, Teng Wang, and Jun Zhang. New sigmoid-like function better than Fisher z transformation. Communications in Statistics: Theory and Methods, 45(8):2332–2341, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.771750.

Yaqub:2017:EPM

[YSG17] Mazhar Yaqub, Javid Shabbir, and Sat Narian Gupta. Estimation of population mean based on dual use of auxiliary information in non response. Communications in Statistics: Theory and Methods, 46(24):12130–12151, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1291969.

Yu:2011:ENC

[YSHL11] Qiqing Yu, Chingfu Sen, Jinlong Huang, and Chinsan Lee. Is each NPMLE of a continuous bivariate distribution function with singly right-censored data really inconsistent? Communications in Statistics: Theory and Methods, 40(5):844–862, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yan:2015:PEI

[YSK15] Huahui Yan, Huisheng Shu, and Xiu Kan. Pricing equityindexed annuities when discrete dividends follow a Markovmodulated jump diffusion model. Communications in Statistics: Theory and Methods, 44(11):2207–2221, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yenigun:2011:TIT

[YSR11] C. D. Yenigün, G. J. Székely, and M. L. Rizzo. A test of independence in two-way contingency tables based on maximal correlation. Communications in Statistics: Theory and Methods, 40(12):2225–2242, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yagi:2018:TEM

[YSS18] Ayaka Yagi, Takashi Seo, and Muni S. Srivastava. Testing equality of mean vectors in a one-way MANOVA with monotone missing data. Communications in Statistics: Theory and Methods, 47(22):5534–5546, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1397168.

Yennum:2019:ISC

[YSS19] Niharika Yennum, Stephen A. Sedory, and Sarjinder Singh. Improved strategy to collect sensitive data by using geometric distribution as a randomization device. Communications in Statistics: Theory and Methods, 48(23):5777–5795, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1522346.

Yin:2014:SLE

[YTX14] G. Yin, Yousef Talafha, and Fubao Xi. Stochastic liénard equations with random switching and two-time scales. Communications in Statistics: Theory and Methods, 43(7):1533–1547, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2010:ELM

[Yu10] Wen Yu. Empirical likelihood method for general additivemultiplicative hazard models. Communications in Statistics: Theory and Methods, 39(16):2977–2990, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2012:EMS

[Yu12] Hong-Fwu Yu. The effect of mis-specification between the lognormal and Weibull distributions on the interval estimation of a quantile for complete data. Communications in Statistics: Theory and Methods, 41(9):1617–1635, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2019:SSI

[Yu19] Xianye Yu. Smoothness of self-intersection local time of multidimensional fractional Brownian motion. Communications in Statistics: Theory and Methods, 48(17):4278–4293, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1493508.

Yadav:2013:GFT

[YUSC13] Rohini Yadav, Lakshmi N. Upadhyaya, Housila P. Singh, and S. Chatterjee. A generalized family of transformed ratioproduct estimators for variance in sample surveys. Communications in Statistics: Theory and Methods, 42(10):1839–1850, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yadav:2014:CEP

[YUSC14a] Rohini Yadav, Lakshmi N. Upadhyaya, Housila P. Singh, and S. Chatterjee. A class of estimators of population variance using auxiliary information in sample surveys. Communications in Statistics: Theory and Methods, 43(6):1248–1260, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yadav:2014:IRP

[YUSC14b] Rohini Yadav, Lakshmi N. Upadhyaya, Housila P. Singh, and S. Chatterjee. Improved ratio and product exponential type estimators for finite population mean in stratified random sampling. Communications in Statistics: Theory and Methods, 43 (15):3269–3285, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yadavalli:2017:AQE

[YVCA17] V. S. S. Yadavalli, V. S. Vaidyanathan, P. Chandrasekhar, and S. Abbas. Applications of quadrivariate exponential distribution to a three-unit warm standby system with dependent structure. Communications in Statistics: Theory and Methods, 46(14):6782–6790, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1134576.

Yang:2011:ESL

[YW11a] Hu Yang and Jibo Wu. Estimation in singular linear models with stochastic linear restrictions and linear equality restrictions. Communications in Statistics: Theory and Methods, 40 (24):4364–4371, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2011:SMN

[YW11b] Hu Yang and Jibo Wu. Some matrix norm Kantorovich inequalities and their applications. Communications in Statistics: Theory and Methods, 40(22):4078–4085, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:2013:NMB

[YW13] Gonglin Yuan and Zengxin Wei. Non monotone backtracking inexact BFGS method for regression analysis. Communications in Statistics: Theory and Methods, 42(2):214–238, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yin:2015:MET

[YW15] Li Yin and Xiaoqin Wang. Measuring and estimating treatment effect on count outcome in randomized trial and observational studies. Communications in Statistics: Theory and Methods, 44(5):1080–1095, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2017:BEF

[YW17a] Kai Yang and Dehui Wang. Bayesian estimation for first-order autoregressive model with explanatory variables. Communications in Statistics: Theory and Methods, 46(22):11214–11227, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1260736.

Yuan:2017:TSE

[YW17b] Min Yuan and Laisheng Wei. Two-sided empirical Bayes test for location parameter in the gamma distribution. Communications in Statistics: Theory and Methods, 46(9):4215–4225, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1080842.

Yu:2018:CCW

[YW18] Donglin Yu and Qunying Wu. Complete convergence for weighted sums of extended negatively dependent random variables under sub-linear expectations. Communications in Statistics: Theory and Methods, 47(19):4741–4750, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1445860.

Yang:2014:NBE

[YWH14] Wenzhi Yang, Xuejun Wang, and Shuhe Hu. A note on the Berry–Esséen bound of sample quantiles for φ-mixing sequence. Communications in Statistics: Theory and Methods, 43(19): 4187–4194, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2017:SPI

[YWH17] Wenzhi Yang, Yiwei Wang, and Shuhe Hu. Some probability inequalities of least-squares estimator in non linear regression model with strong mixing errors. Communications in Statistics: Theory and Methods, 46(1):165–175, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.988261.

Yao:2017:ANN

[YWL17] Mei Yao, Jiang-Feng Wang, and Lu Lin. Asymptotic normality for a non parametric estimator of conditional quantile with left-truncated data. Communications in Statistics: Theory and Methods, 46(13):6280–6292, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1124120.

Yao:2018:VSW

[YWLW18] Mei Yao, Jiang-Feng Wang, Lu Lin, and Yu-Xin Wang. Variable selection and weighted composite quantile estimation of regression parameters with left-truncated data. Communications in Statistics: Theory and Methods, 47(18):4469–4482, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1376089.

Yeh:2011:EED

[YWLY11] Li Lun Yeh, Peng Kai Wang, Feng Chia Li, and Yu Mei Yeh. An extension of economic design of x-bar control charts for non normally distributed data under Weibull shock models. Communications in Statistics: Theory and Methods, 40(21):3879–3902, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2017:EAM

[YWLZ17] Shi-Hang Yu, De-Hui Wang, Kun Li, and Zhi-Wen Zhao. Estimation in autoregressive models with surrogate data and validation data. Communications in Statistics: Theory and Methods, 46(3):1532–1545, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1019154.

Yu:2015:LMT

[YWO^{+}15] Qiqing Yu, George Y. C. Wong, Michael P. Osborne, Yuting Hsu, and Xiaosong Ai. The Lehmann model with timedependent covariates. Communications in Statistics: Theory and Methods, 44(20):4380–4395, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2014:SLL

[YWS14] Weiguo Yang, Bei Wang, and Zhiyan Shi. Strong law of large numbers for countable asymptotic circular Markov chains. Communications in Statistics: Theory and Methods, 43(18): 3943–3954, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yao:2017:OIC

[YWX17] Dingjun Yao, Rongming Wang, and Lin Xu. Optimal impulse control for dividend and capital injection with proportional reinsurance and exponential premium principle. Communications in Statistics: Theory and Methods, 46(5):2519–2541, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.901374.

Ye:2018:CPS

[YWZ18] Shili Ye, Dongying Wang, and Runchu Zhang. 2^{m−p} compromise plans with strongly clear specified main effects and two-factor interactions. Communications in Statistics: Theory and Methods, 47(12):2827–2834, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1303739.

Yang:2015:VSS

[YX15] Hu Yang and Xiaochao Xia. Variable selection for semiparametric varying coefficient partially linear errors-in-variables (EV) model with missing response. Communications in Statistics: Theory and Methods, 44(21):4521–4539, 2015. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yan:2017:NCI

[YX17] Liang Yan and Xingzhong Xu. A new confidence interval in measurement error model with the reliability ratio known. Communications in Statistics: Theory and Methods, 46(19): 9636–9650, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1217018.

Yi:2018:HDL

[YX18] Linqi Yi and Junshan Xie. A high-dimensional likelihood ratio test for circular symmetric covariance structure. Communications in Statistics: Theory and Methods, 47(6):1392–1402, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1319484.

Yang:2011:ELG

[YXC11] Yiping Yang, Liugen Xue, and Weihu Cheng. The empirical likelihood goodness-of-fit test for a regression model with randomly censored data. Communications in Statistics: Theory and Methods, 40(3):424–435, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2011:CLB

[YXG11] Jun Yang, Min Xie, and Thong Ngee Goh. Control limits based on the narrowest confidence interval. Communications in Statistics: Theory and Methods, 40(12):2172–2181, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:2017:MHL

[YXZ17] Wei Yu, Wangli Xu, and Lixing Zhu. A modified Hosmer– Lemeshow test for large data sets. Communications in Statistics: Theory and Methods, 46(23):11813–11825, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1285922.

Yang:2014:LIA

[YY14] Lian Yang and Hu Yang. Local influence analysis for the ridge regression under stochastic linear restrictions. Communications in Statistics: Theory and Methods, 43(17):3580–3595, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2016:ALI

[YY16] Hu Yang and Lian Yang. Assessing local influence for elliptical linear models under equality constraints. Communications in Statistics: Theory and Methods, 45(15):4517–4527, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2013.773351.

Yan:2017:BAP

[YYL17] Rongfang Yan, Yinping You, and Xiaohu Li. On bivariate ageing properties of exchangeable Farlie–Gumbel–Morgenstern distributions. Communications in Statistics: Theory and Methods, 46(23):11843–11853, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1285927.

Yongjin:2018:CBR

[YYLM18] Zhang Yongjin, Sun Youchao, Li Longbiao, and Zhao Ming. Copula-based reliability analysis for a parallel system with a cold standby. Communications in Statistics: Theory and Methods, 47(3):562–582, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1309432.

Yang:2018:SLL

[YYS^{+}18] Jie Yang, Weiguo Yang, Zhiyan Shi, Yiqing Li, Bei Wang, and Yue Zhang. Strong law of large numbers for generalized sample relative entropy of non homogeneous Markov chains. Communications in Statistics: Theory and Methods, 47(7):1571–1579, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1321770.

Yang:2019:LBA

[YYS19] Xue Yang, Gui-Jun Yang, and Ya-Juan Su. Lower bound of average centered L_{2}-discrepancy for U-type designs. Communications in Statistics: Theory and Methods, 48(4):995–1008, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1422761.

Yang:2017:SLL

[YYW17] Guoqing Yang, Weiguo Yang, and Xiaotai Wu. The strong laws of large numbers for countable non homogeneous hidden Markov models. Communications in Statistics: Theory and Methods, 46(17):8808–8819, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1193203.

Yang:2019:LLN

[YYWH19] Wenzhi Yang, Lei Yang, Da Wei, and Shuhe Hu. The laws of large numbers for Pareto-type random variables with infinite means. Communications in Statistics: Theory and Methods, 48(12):3044–3054, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1473602.

Yang:2014:FSP

[YYX14] Hu Yang, Huiliang Ye, and Kai Xue. A further study of predictions in linear mixed models. Communications in Statistics: Theory and Methods, 43(20):4241–4252, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:2017:SCN

[YYY17] Xiutao Yang, Shanchao Yang, and Xin Yang. Strong consistency of non parametric kernel regression estimator for strong mixing samples. Communications in Statistics: Theory and Methods, 46(21):10537–10548, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1239110.

Yan:2017:AEL

[YZ17a] Zhen Yan and Junjian Zhang. Adjusted empirical likelihood for value at risk and expected shortfall. Communications in Statistics: Theory and Methods, 46(5):2580–2591, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.1002933.

Yu:2017:GAQ

[YZ17b] Chang Yu and Daniel Zelterman. A general approximation to quantiles. Communications in Statistics: Theory and Methods, 46(19):9834–9841, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1222433.

Yue:2018:RIV

[YZ18] Rong-Xian Yue and Xiao-Dong Zhou. Robust integer-valued designs for linear random intercept models. Communications in Statistics: Theory and Methods, 47(17):4338–4354, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1373820.

Yu:2017:URM

[YZFM17] Chang Yu, Sanguo Zhang, Christine Friedenreich, and Charles E. Matthews. Using repeated measures to correct correlated measurement errors through orthogonal decomposition. Communications in Statistics: Theory and Methods, 46(23):11604–11611, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1275693.

Yang:2017:AEC

[YZG17] Weiguo Yang, Daying Zhu, and Rong Gao. Almost everywhere convergence for sequences of pairwise NQD random variables. Communications in Statistics: Theory and Methods, 46(5): 2494–2505, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1048883.

Yan:2017:ODA

[YZP17] Fei Yan, Chongqi Zhang, and Heng Peng. Optimal designs for additive mixture model with heteroscedastic errors. Communications in Statistics: Theory and Methods, 46(13):6401–6411, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1129414.

Zaier:2014:TDE

[ZA14] L. Hedhili Zaier and M. Abed. Temporal disaggregation of economic time series using artificial neural networks. Communications in Statistics: Theory and Methods, 43(8):1824–1833, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zadlo:2015:PCD

[Zad15] Tomasz Zadlo. On prediction for correlated domains in longitudinal surveys. Communications in Statistics: Theory and Methods, 44(4):683–697, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zang:2017:ASL

[Zan17] Qing-Pei Zang. Almost sure local central limit theorem for sample range. Communications in Statistics: Theory and Methods, 46(3):1050–1055, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2012.
763092.

Zare:2019:MCD

[Zar19] Karim Zare. Multiple cases deletion measures in linear measurement error models. Communications in Statistics: Theory and Methods, 48(4):954–963, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1422758.

Zafar:2014:PVC

[ZARH14] Raja Fawad Zafar, Nasir Abbas, Muhammad Riaz, and Zawar Hussain. Progressive variance control charts for monitoring process dispersion. Communications in Statistics: Theory and Methods, 43(23):4893–4907, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zamanzade:2014:PTP

[ZAV14] Ehsan Zamanzade, Nasser Reza Arghami, and Michael Vock. A parametric test of perfect ranking in balanced ranked set sampling. Communications in Statistics: Theory and Methods, 43(21):4589–4611, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2012:BRC

[ZB12a] Yudong Zhao and Bruce M. Brown. A boundary-respecting confidence interval method for a general effect size measure in paired experiments. Communications in Statistics: Theory and Methods, 41(2):363–379, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zoia:2012:SCD

[ZB12b] Maria Grazia Zoia and Laura Barbieri. A short-cut derivation for the solution of autoregressive models from sharp algebraic arguments. Communications in Statistics: Theory and Methods, 41(20):3704–3721, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:SPP

[ZB17] Zhengcheng Zhang and N. Balakrishnan. Stochastic properties and parameter estimation for a general load-sharing system. Communications in Statistics: Theory and Methods, 46(2):747– 760, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2015.1004095.

Zaman:2019:MRE

[ZB19] Tolga Zaman and Hasan Bulut. Modified ratio estimators using robust regression methods. Communications in Statistics: Theory and Methods, 48(8):2039–2048, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1441419.

Zamani:2017:NPD

[ZBA17] Z. Zamani, G. R. Mohtashami Borzadaran, and M. Amini. On a new positive dependence concept based on the conditional mean inactivity time order. Communications in Statistics: Theory and Methods, 46(4):1779–1787, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1030417.

Zhang:2018:MST

[ZBHW18] Chao Zhang, Zhidong Bai, Jiang Hu, and Chen Wang. Multisample test for high-dimensional covariance matrices. Communications in Statistics: Theory and Methods, 47(13):3161–3177, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1350272.

Zhang:2017:STM

[ZBTL17] Shuxia Zhang, Gejun Bao, Boping Tian, and Yijun Li. Semiparametric test for multiple change-points based on empirical likelihood. Communications in Statistics: Theory and Methods, 46(7):3574–3585, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1066815.

Zheng:2011:NRB

[ZC11] Yaohui Zheng and Nanlian Cai. A note on the relations between some concepts of positive dependence. Communications in Statistics: Theory and Methods, 40(8):1335–1341, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2014:WCL

[ZC14] Defei Zhang and Zengjing Chen. A weighted central limit theorem under sublinear expectations. Communications in Statistics: Theory and Methods, 43(3):566–577, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:ATB

[ZC17a] Yanfang Zhang and Fengyang Cheng. Asymptotic tail behavior of a random sum with conditionally dependent subexponential summands. Communications in Statistics: Theory and Methods, 46(12):5888–5895, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1112916.

Zhao:2017:LDE

[ZC17b] Shoujiang Zhao and Ting Chen. Large deviation expansion for maximum-likelihood estimator of α Brownian bridge. Communications in Statistics: Theory and Methods, 46(15):7313–7326, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1148734.

Ziram:2016:SAN

[ZCC16] Ferhat Ziram, Dominique Cellier, and François Charlot. Statistical analysis of the number of self-overlapping leftmost repeats in an homogeneous stationary Markov chain on finite states. Communications in Statistics: Theory and Methods, 45(20): 6087–6101, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.957851.

Zhao:2018:CSP

[ZCCW18] Jing Zhao, Weihu Cheng, Haiqing Chen, and Mixia Wu. Comparisons of several Pareto distributions based on record values. Communications in Statistics: Theory and Methods, 47(10): 2456–2468, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1256410.

Zeng:2019:MSM

[ZCHR19] Jie Zeng, Weihu Cheng, Guozhi Hu, and Yaohua Rong. Model selection and model averaging for semiparametric partially linear models with missing data. Communications in Statistics: Theory and Methods, 48(2):381–395, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1410717.

Zhu:2017:OFD

[ZDDH17] Huiming Zhu, Chao Deng, Yingchun Deng, and Ya Huang. Optimal financing and dividend policy with Markovian switching regimes. Communications in Statistics: Theory and Methods, 46(5):2161–2180, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1035393.

Zhou:2016:EPF

[ZDS16] Jianjun Zhou, Jiang Du, and Zhimeng Sun. M-estimation for partially functional linear regression model based on splines. Communications in Statistics: Theory and Methods, 45(21): 6436–6446, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.921309.

Zheng:2017:ANE

[ZDW17] Lulu Zheng, Yang Ding, and Xuejun Wang. Asymptotic normality for the estimator of non parametric regression model under φ-mixing errors. Communications in Statistics: Theory and Methods, 46(14):6764–6773, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1134574.

Zhang:2017:JDF

[ZDZ17] Tao Zhang, Pengjie Dai, and Qingzhao Zhang. Joint detection for functional polynomial regression with autoregressive errors. Communications in Statistics: Theory and Methods, 46(16): 7837–7854, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1096384.

Zeng:2017:IPL

[Zen17a] Guoping Zeng. Invariant properties of logistic regression model in credit scoring under monotonic transformations. Communications in Statistics: Theory and Methods, 46(17):8791–8807, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1193200.

Zeng:2017:EML

[Zen17b] Guoping Zeng. On the existence of maximum likelihood estimates for weighted logistic regression. Communications in Statistics: Theory and Methods, 46(22):11194–11203, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1260742.

Fu:2016:SEA

[zF16] Ying zi Fu. Stochastic EM algorithm of a finite mixture model from hurdle Poisson distribution with missing responses. Communications in Statistics: Theory and Methods, 45(20):5918– 5932, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.953689.

Zhang:2017:PLS

[ZF17] Jun Zhang and Zhenghui Feng. Partial linear single-index models with additive distortion measurement errors. Communications in Statistics: Theory and Methods, 46(24):12165–12193, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1291971.

Zhai:2018:LOD

[ZF18] Yi Zhai and Zhide Fang. Locally optimal designs for some dose-response models with continuous endpoints. Communications in Statistics: Theory and Methods, 47(16):3803–3819, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1361996.

Zhou:2017:QRI

[ZFD17] Xiuqing Zhou, Yanqin Feng, and Xiuli Du. Quantile regression for interval censored data. Communications in Statistics: Theory and Methods, 46(8):3848–3863, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1073317.

Zhang:2013:CAL

[ZFL13] Shangli Zhang, Zhide Fang, and Gang Liu. Characterization of admissible linear estimators in multivariate linear model with respect to inequality constraints under matrix loss function. Communications in Statistics: Theory and Methods, 42(15): 2837–2850, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2017:TNP

[ZFZ17] Yan Zhou, Liya Fu, and Baoxue Zhang. Two non parametric methods for change-point detection in distribution. Communications in Statistics: Theory and Methods, 46(6):2801–2815, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1048891.

Zhang:2017:FCM

[ZG17a] Sumei Zhang and Junhao Geng. Fourier-cosine method for pricing forward starting options with stochastic volatility and jumps. Communications in Statistics: Theory and Methods, 46(20):9995–10004, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1228960.

Zhao:2017:SSD

[ZG17b] Bo Zhao and Joseph Glaz. Scan statistics for detecting a local change in variance for two-dimensional normal data. Communications in Statistics: Theory and Methods, 46(11):5517–5530, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1104354.

Zadeh:2019:ATI

[ZG19a] A. Badamchi Zadeh and M. H. Ghahramani. Analysis of a two incongruent arrivals and services M/G/1 queue with random feedback. Communications in Statistics: Theory and Methods, 48(10):2509–2520, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1470644.

Zombade:2019:STN

[ZG19b] D. M. Zombade and V. B. Ghute. Shewhart-type nonparametric control chart for process location. Communications in Statistics: Theory and Methods, 48(7):1621–1634, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1435811.

Zheng:2017:RAL

[ZGK17] Qi Zheng, Colin Gallagher, and K. B. Kulasekera. Robust adaptive lasso for variable selection. Communications in Statistics: Theory and Methods, 46(9):4642–4659, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1019138.

Zhibin:2016:MFP

[ZH16] Liu Zhibin and Liu Haohan. Multi-factor prediction of water flooding efficiency based on a time-varying system. Communications in Statistics: Theory and Methods, 45(20):5873–5883, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.950754.

Zhang:2011:ONC

[Zha11] Zhengcheng Zhang. Ordering new conditional residual lifetimes of k-out-of-n systems. Communications in Statistics: Theory and Methods, 40(9):1591–1600, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2013:AGS

[Zha13] Qiang Zhang. Asymmetric group sequential designs under fractional Brownian motion. Communications in Statistics: Theory and Methods, 42(12):2172–2184, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2014:TSC

[Zha14] Haibing Zhao. Two-sided confidence bands with the inferential sensitivity of one-sided bands. Communications in Statistics: Theory and Methods, 43(6):1179–1191, 2014. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2015:GCR

[Zha15a] Bojuan Barbara Zhao. Generalized concept of relative risk and wider applications of the proportional hazards model and the Kaplan–Meier estimator. Communications in Statistics: Theory and Methods, 44(24):5254–5266, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2015:TNN

[Zha15b] Guolong Zhao. A test of non null hypothesis for linear trends in proportions. Communications in Statistics: Theory and Methods, 44(8):1621–1639, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2016:PUR

[Zha16a] Lingxiang Zhang. Performance of unit-root tests for non linear unit-root and partial unit-root processes. Communications in Statistics: Theory and Methods, 45(15):4528–4536, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.922985.

Zhang:2016:EAS

[Zha16b] Yong Zhang. An extension of almost sure central limit theorem for self-normalized products of sums for mixing sequences. Communications in Statistics: Theory and Methods, 45(22): 6625–6640, 2016. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2014.963619.

Zhang:2017:SLL

[Zha17a] Guohui Zhang. The strong laws of large numbers for weighted sums of extended negatively dependent random variables. Communications in Statistics: Theory and Methods, 46(20):9881– 9891, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1222435.

Zhang:2017:ILQ

[Zha17b] Jin Zhang. On the independence of linear and quadratic forms in normal variates. Communications in Statistics: Theory and Methods, 46(17):8493–8496, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1183786.

Zhang:2017:PUR

[Zha17c] Lingxiang Zhang. Partial unit root and surplus-lag Granger causality testing: a Monte Carlo simulation study. Communications in Statistics: Theory and Methods, 46(24):12317–12323, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1295077.

Zhang:2017:BRV

[Zha17d] Ying-Ying Zhang. The Bayes rule of the variance parameter of the hierarchical normal and inverse gamma model under Stein’s loss. Communications in Statistics: Theory and Methods, 46(14):7125–7133, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1148733.

Zhang:2017:EPS

[Zha17e] Yuanqing Zhang. Estimation of partially specified spatial panel data models with random-effects and spatially correlated error components. Communications in Statistics: Theory and Methods, 46(3):1056–1079, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.972567.

Zheng:2018:IBI

[Zhe18] Songfeng Zheng. An improved Bennett’s inequality. Communications in Statistics: Theory and Methods, 47(17):4152–4159, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2017.1367818.

Zhou:2015:KBP

[ZHL15] Jie Zhou, Lu Han, and Sanyang Liu. Kernel-based profile estimation for ordinary differential equations with partially measured state variables. Communications in Statistics: Theory and Methods, 44(16):3446–3463, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2014:DDA

[ZHLZ14] Yanli Zhang, Lei Huo, Lu Lin, and Yunhui Zeng. The Dantzig discriminant analysis with high dimensional data. Communications in Statistics: Theory and Methods, 43(23):5012–5025, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2009:WML

[Zho09] Tianyue Zhou. Weighting method for a linear mixed model. Communications in Statistics: Theory and Methods, 39(2):214– 227, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2010:MIM

[Zho10] Ying Zhou. Multiple interval mapping for quantitative trait loci via EM algorithm in the presence of crossover interference. Communications in Statistics: Theory and Methods, 39(17): 3041–3057, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2017:DSD

[Zho17] Likai Zhou. Double-smoothed drift estimation of jumpdiffusion model. Communications in Statistics: Theory and Methods, 46(8):4137–4149, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1078479.

Zhu:2010:ESI

[Zhu10] Li-Ping Zhu. Extending the scope of inverse regression methods in sufficient dimension reduction. Communications in Statistics: Theory and Methods, 40(1):84–95, 2010. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:BRU

[ZHW17] J. Zhang, J. Huang, and X. Wu. Bayesian ratemaking under Dirichlet process mixtures. Communications in Statistics: Theory and Methods, 46(22):11327–11340, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1267761.

Zougab:2018:MGB

[ZHZA18] N. Zougab, L. Harfouche, Y. Ziane, and S. Adjabi. Multivariate generalized Birnbaum-Saunders kernel density estimators. Communications in Statistics: Theory and Methods, 47(18): 4534–4555, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1377252.

Zamani:2012:FFG

[ZI12] Hossein Zamani and Noriszura Ismail. Functional form for the generalized Poisson regression model. Communications in Statistics: Theory and Methods, 41(20):3666–3675, 2012. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zamani:2014:FFZ

[ZI14] Hossein Zamani and Noriszura Ismail. Functional form for the zero-inflated generalized Poisson regression model. Communications in Statistics: Theory and Methods, 43(3):515–529, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ziegel:2013:SMR

[Zie13] Johanna Ziegel. Stereological modelling of random particles. Communications in Statistics: Theory and Methods, 42(7): 1428–1442, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zilinskas:2011:SSE

[Zil11] Antanas Zilinskas. Small sample estimation of parameters for Wiener process with noise. Communications in Statistics: Theory and Methods, 40(16):3020–3028, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zine:2012:MVB

[Zin12] Raoudha Zine. On the matrix-variate beta distribution. Communications in Statistics: Theory and Methods, 41(9):1569– 1582, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zinodiny:2018:BMR

[Zin18] S. Zinodiny. Bayes minimax ridge regression estimators. Communications in Statistics: Theory and Methods, 47(22):5519– 5533, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2017.1397167.

Zhang:2010:NBH

[ZJC10] Lingyun Zhang, Wenjiang Jiang, and Gemai Chen. A note on Bartlett’s MTest for homogeneity of variances. Communications in Statistics: Theory and Methods, 39(16):2890–2900, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:BMA

[ZJLO17] Zhiyong Zhang, Kaifeng Jiang, Haiyan Liu, and In-Sue Oh. Bayesian meta-analysis of correlation coefficients through power prior. Communications in Statistics: Theory and Methods, 46(24):11988–12007, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1288251.

Zhao:2017:EPR

[ZJZ17] Mu Zhao, Hongmei Jiang, and Yong Zhou. Estimation of percentile residual life function with left-truncated and rightcensored data. Communications in Statistics: Theory and Methods, 46(2):995–1006, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1010006.

Zareifard:2013:EBE

[ZK13a] Hamid Zareifard and Majid Jafari Khaledi. Empirical Bayes estimation in regression models with generalized skew-slash errors. Communications in Statistics: Theory and Methods, 42(6):1105–1123, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2013:PPF

[ZK13b] Juan Zhang and John E. Kolassa. A practical procedure to find matching priors for frequentist inference. Communications in Statistics: Theory and Methods, 42(15):2758–2767, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2015:AOD

[ZK15a] Yongli Zhang and Sergio Koreisha. Adaptive order determination for constructing time series forecasting models. Communications in Statistics: Theory and Methods, 44(22):4826–4847, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zocchi:2015:GCT

[ZK15b] Silvio S. Zocchi and Célestin C. Kokonendji. On general continuous triangular and two-sided power distributions. Communications in Statistics: Theory and Methods, 44(14):3099–3106, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2016:UAT

[ZK16] Chenhua Zhang and Shaobai Kan. Uniform approximation of the tail probability of weighted sums of subexponential random variables. Communications in Statistics: Theory and Methods, 45(21):6290–6298, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.960585.

Zafeiris:2019:RMT

[ZK19] Konstantinos N. Zafeiris and Anastasia Kostaki. Recent mortality trends in Greece. Communications in Statistics: Theory and Methods, 48(1):112–126, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1353625.

Zeebari:2018:SUR

[ZKS18] Zangin Zeebari, B. M. Golam Kibria, and Ghazi Shukur. Seemingly unrelated regressions with covariance matrix of crossequation ridge regression residuals. Communications in Statistics: Theory and Methods, 47(20):5029–5053, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1383431.

Zheng:2016:MQF

[ZKW16] Shimin Zheng, Jeff Knisley, and Kesheng Wang. Moments and quadratic forms of matrix variate skew normal distributions. Communications in Statistics: Theory and Methods, 45(3):794– 803, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2012:ANE

[ZL12] Jing-Jing Zhang and Han-Ying Liang. Asymptotic normality of estimators in heteroscedastic semi-parametric model with strong mixing errors. Communications in Statistics: Theory and Methods, 41(12):2172–2201, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2014:OPC

[ZL14a] Peng Zhao and Xiaohu Li. Ordering properties of convolutions from heterogeneous populations: a review on some recent developments. Communications in Statistics: Theory and Methods, 43(10–12):2260–2273, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2014:WEN

[ZL14b] Xing-Cai Zhou and Jin-Guan Lin. Wavelet estimator in nonparametric regression model with dependent error’s structure. Communications in Statistics: Theory and Methods, 43(22): 4707–4722, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhuang:2015:ARS

[ZL15] Jinsen Zhuang and Xiaohu Li. Allocating redundancies to kout-of-n systems with independent and heterogeneous components. Communications in Statistics: Theory and Methods, 44 (24):5109–5119, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2016:CDC

[ZL16a] Huiming Zhang and Bo Li. Characterizations of discrete compound Poisson distributions. Communications in Statistics: Theory and Methods, 45(22):6789–6802, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.901375.

Zhang:2016:BET

[ZL16b] Jing-Jing Zhang and Han-Ying Liang. Berry–Esséen type bounds in heteroscedastic errors-in-variables model. Communications in Statistics: Theory and Methods, 45(13):3993–4017, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.915042.

Zhao:2016:DBE

[ZL16c] Haibing Zhao and Rui Li. Difference-based estimation and model identification for panel data semiparametric models with cross-section dependence. Communications in Statistics: Theory and Methods, 45(4):1099–1117, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.857417.

Zhao:2016:SCE

[ZL16d] Kaifeng Zhao and Heng Lian. Sparsistent and constansistent estimation of the varying-coefficient model with a diverging number of predictors. Communications in Statistics: Theory and Methods, 45(21):6385–6399, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.890224.

Zhao:2016:CMA

[ZL16e] Sheng-Li Zhao and Peng-Fei Li. Construction of minimum aberration blocked two-level regular factorial designs. Communications in Statistics: Theory and Methods, 45(17):5028–5036, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.935433.

Zhou:2016:LLV

[ZL16f] Xiuqing Zhou and Guoxiang Liu. LAD-lasso variable selection for doubly censored median regression models. Communications in Statistics: Theory and Methods, 45(12):3658–3667, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2014.904357.

Zhou:2016:LTR

[ZL16g] Zhiyong Zhou and Zhengyan Lin. Limit theory for random coefficient autoregressive process under possibly infinite variance error sequence. Communications in Statistics: Theory and Methods, 45(12):3562–3576, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.904354.

Zou:2017:CRW

[ZL17] Yu-Ye Zou and Han-Ying Liang. Convergence rate of wavelet density estimator with data missing randomly when covariables are present. Communications in Statistics: Theory and Methods, 46(2):1007–1023, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1010008.

Zhou:2018:ELI

[ZL18] Hui Zhou and Jie Li. Empirical likelihood inference for error density estimators in first-order autoregression models. Communications in Statistics: Theory and Methods, 47(10):2351– 2361, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.974820.

Zhao:2011:RSO

[ZLD11] Peng Zhao, Xiaohu Li, and Gaofeng Da. Right spread order of the second-order statistic from heterogeneous exponential random variables. Communications in Statistics: Theory and Methods, 40(17):3070–3081, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2015:SCT

[ZLD15] Qiang Zhang, Dejian Lai, and Barry R. Davis. Stochastically curtailed tests under fractional Brownian motion. Communications in Statistics: Theory and Methods, 44(5):1053–1064, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2016:RCI

[ZLD16] Qiang Zhang, Dejian Lai, and Barry R. Davis. Repeated confidence intervals and prediction intervals using stochastic curtailment under fractional Brownian motion. Communications in Statistics: Theory and Methods, 45(14):4295–4306, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.919400.

Zhang:2014:NFU

[ZLDT14] Zhong-Yuan Zhang, Tao Li, Chris Ding, and Jie Tang. An NMF-framework for unifying posterior probabilistic clustering and probabilistic latent semantic indexing. Communications in Statistics: Theory and Methods, 43(19):4011–4024, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2014:LAR

[ZLF14] Hui Zhou, Jie Li, and Ke-Ang Fu. LIL for the adjusted range of partial sums in AR(1) models with possibly infinite variance.

Communications in Statistics: Theory and Methods, 43(17): 3690–3697, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2011:TTB

[ZLL11] Caiya Zhang, Yan Luo, and Zhengyan Lin. On two types of breakdown points of weighted L^{2}-median. Communications in Statistics: Theory and Methods, 40(7):1131–1141, 2011. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2011:APQ

[ZLN11] Sanguo Zhang, Yuan Liao, and Wei Ning. Asymptotic properties of quasi-maximum likelihood estimates in generalized linear models. Communications in Statistics: Theory and Methods, 40(24):4417–4430, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:EDB

[ZLT17] Xing Zhang, Jian Liu, and Chao Tan. Economic design based on a multivariate Bayesian VSI chart with a dual control limit. Communications in Statistics: Theory and Methods, 46(14): 6808–6822, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1136418.

Zou:2010:ANC

[ZLWZ10] Changliang Zou, Yukun Liu, Zhaojun Wang, and Runchu Zhang. Adaptive nonparametric comparison of regression curves. Communications in Statistics: Theory and Methods, 39(7):1299–1320, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:CSM

[ZLYZ17] Tian-Fang Zhang, Zhi-Ming Li, Jian-Feng Yang, and Run-Chu Zhang. Construction of some mixed two- and fourlevel regular designs with GMC criterion. Communications in Statistics: Theory and Methods, 46(17):8497–8509, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1183787.

Zhang:2015:LSE

[ZLZ15] Shibin Zhang, Zhengyan Lin, and Xinsheng Zhang. A least squares estimator for lévy-driven moving averages based on discrete time observations. Communications in Statistics: Theory and Methods, 44(6):1111–1129, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2019:ODM

[ZLZ19] Zhibin Zhu, Guanghui Li, and Chongqi Zhang. A-optimal designs for mixture central polynomial model with qualitative factors. Communications in Statistics: Theory and Methods, 48(10):2345–2355, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1472783.

Zhang:2013:MRR

[ZM13] Zhengcheng Zhang and William Q. Meeker. Mixture representations of reliability in coherent systems and preservation results under double monitoring. Communications in Statistics: Theory and Methods, 42(3):385–397, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zenga:2016:CPT

[ZMCM16] Mariangela Zenga, Adele H. Marshall, Franca Crippa, and Hannah Mitchell. The Coxian phase-type distribution as a contribution to the multilevel model of in-hospital mortality. Communications in Statistics: Theory and Methods, 45(6):1819–1830, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1038394.

Zhang:2017:CCW

[ZMMX17] Lin Zhang, Yu Miao, Jianyong Mu, and Jie Xu. Complete convergence for weighted sums of mixingale sequences and statistical applications. Communications in Statistics: Theory and Methods, 46(21):10692–10701, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1242738.

Zhao:2017:RPA

[ZMN17] Xufeng Zhao, Satoshi Mizutani, and Toshio Nakagawa. Replacement policies for age and working numbers with random life cycle. Communications in Statistics: Theory and Methods, 46(14):6791–6802, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1136416.

Zheng:2017:RBM

[ZMV^{+}17] Shimin Zheng, Eunice Mogusu, Sreenivas P. Veeranki, Megan Quinn, and Yan Cao. The relationship between the mean, median, and mode with grouped data. Communications in Statistics: Theory and Methods, 46(9):4285–4295, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1081948.

Zhang:2018:RBV

[ZN18] Y. Zhang and S. Nadarajah. A review of backtesting for value at risk. Communications in Statistics: Theory and Methods, 47(15):3616–3639, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1361984.

Zoia:2009:TGL

[Zoi09] Maria Grazia Zoia. Tailoring the Gaussian law for excess kurtosis and skewness by Hermite polynomials. Communications in Statistics: Theory and Methods, 39(1):52–64, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zornig:2014:GBN

[Zör14] Peter Zörnig. On generalized binomial and negative binomial distributions for dependent Bernoulli variables. Communications in Statistics: Theory and Methods, 43(9):1887–1906, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:2017:CMI

[ZP17] Qi Zheng and Limin Peng. Consistent model identification of varying coefficient quantile regression with BIC tuning parameter selection. Communications in Statistics: Theory and Methods, 46(3):1031–1049, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1010009.

Zhang:2010:TST

[ZPZ10] Chongqi Zhang, Heng Peng, and Jin-Ting Zhang. Two samples tests for functional data. Communications in Statistics: Theory and Methods, 39(4):559–578, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:ANI

[ZQQY17] Yong Zhang, Xiaodi Qian, Hong Qin, and Ting Yan. Affiliation networks with an increasing degree sequence. Communications in Statistics: Theory and Methods, 46(22):11163–11180, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2016.1260741.

Zhang:2018:OAC

[ZQZY18] Xue-Ru Zhang, Zong-Feng Qi, Yong-Dao Zhou, and Feng Yang. Orthogonal-array composite design for the third-order models. Communications in Statistics: Theory and Methods, 47(14): 3488–3507, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1359297.

Zhang:2019:EBE

[ZRL19a] Ying-Ying Zhang, Teng-Zhong Rong, and Man-Man Li. The empirical Bayes estimators of the mean and variance parameters of the normal distribution with a conjugate normal-inverse-gamma prior by the moment method and the MLE method. Communications in Statistics: Theory and Methods, 48(9): 2286–2304, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1465081.

Zhang:2019:EIB

[ZRL19b] Ying-Ying Zhang, Teng-Zhong Rong, and Man-Man Li. Expectation identity for the binomial distribution and its application in the calculations of high-order binomial moments. Communications in Statistics: Theory and Methods, 48(22):5467–5476, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1435818.

Zinodiny:2017:MEM

[ZRN17] S. Zinodiny, S. Rezaei, and S. Nadarajah. Minimax estimation of the mean of the multivariate normal distribution. Communications in Statistics: Theory and Methods, 46(3):1422–1432, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1019146.

Zhou:2010:MSI

[ZS10] Xiuqing Zhou and Ningzhong Shi. Methods of statistical inference for median regression models with doubly censored data. Communications in Statistics: Theory and Methods, 39(17): 3140–3152, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2012:ISM

[ZS12] Hong Zhang and Ping Sa. An improved solution for multiple comparisons with a control in response surface methodology with two predictor variables. Communications in Statistics: Theory and Methods, 41(7):1228–1241, 2012. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:2015:SAC

[ZS15] Guangyu Zheng and Yimin Shi. Statistical analysis in constantstress accelerated life tests for generalized exponential distribution with progressive Type-I hybrid censoring. Communications in Statistics: Theory and Methods, 44(23):4962–4982, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2017:CGM

[ZS17a] Sheng-Li Zhao and Qing Sun. On constructing general minimum lower order confounding two-level block designs. Communications in Statistics: Theory and Methods, 46(3):1261–1274, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1014112.

Zhou:2017:LTT

[ZS17b] Xingwu Zhou and Martin Solberger. An LM-type test for idiosyncratic unit roots in the exact factor model with integrated factors. Communications in Statistics: Theory and Methods, 46(10):4888–4914, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1091079.

Zhao:2015:CPE

[ZSD^{+}15] Xibin Zhao, Shubin Si, Hongyan Dui, Zhiqiang Cai, Junbo Wang, and Xiaoyu Song. Compositional performance evaluation with importance measures. Communications in Statistics: Theory and Methods, 44(24):5240–5253, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2016:OSE

[ZSJ16] Qin Zhou, Lianjie Shu, and Wei Jiang. One-sided EWMA control charts for monitoring Poisson processes with varying sample sizes. Communications in Statistics: Theory and Methods, 45(20):6112–6132, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
957853.

Zeebari:2012:MRP

[ZSK12] Z. Zeebari, G. Shukur, and B. M. G. Kibria. Modified ridge parameters for seemingly unrelated regression model. Communications in Statistics: Theory and Methods, 41(9):1675–1691, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:AHR

[ZSL17] Shuo Zhang, Zhan-Jie Song, and Ying Li. Approximation of homogeneous random field from local averages. Communications in Statistics: Theory and Methods, 46(8):3864–3877, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1073318.

Zhang:2015:TSM

[ZSS15] Ying Zhang, Jagbir Singh, and Ramalingam Shanmugam. A tutorial of survival modeling to capture covariate effect. Communications in Statistics: Theory and Methods, 44(13):2788–2797, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2018:SIH

[ZSS18] Fanrong Zhao, Weixing Song, and Jianhong Shi. Statistical inference for heteroscedastic semi-varying coefficient EV models. Communications in Statistics: Theory and Methods, 47(10): 2432–2455, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1242735.

Zhang:2019:EMF

[ZSWA19] Zhongwen Zhang, Lixin Song, Xiaoguang Wang, and Muhammad Amin. Estimation of multivariate frailty models with varying coefficients. Communications in Statistics: Theory and Methods, 48(3):450–461, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1414257.

Zheng:2013:ELB

[ZSY13] Ming Zheng, Yunting Sun, and Wen Yu. Empirical likelihood based synthetic data method for censored regression analysis. Communications in Statistics: Theory and Methods, 42(24): 4365–4383, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2013:ILS

[ZT13] Fang Zhu and Gong Tang. An index of local sensitivity to nonignorability for a pseudolikelihood method. Communications in Statistics: Theory and Methods, 42(6):954–973, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:IRC

[ZT17] Shuxia Zhang and Boping Tian. Inference for random coefficient INAR(k) with the occasional level shift random noise based on dual empirical likelihood. Communications in Statistics: Theory and Methods, 46(14):6994–7006, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1139134.

Zhang:2018:SMD

[ZT18] Shuxia Zhang and Boping Tian. Semiparametric method for detecting multiple change points model in financial time series.

Communications in Statistics: Theory and Methods, 47(11): 2664–2683, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1316401.

Zhao:2015:NCS

[ZTSL15] Shi-Shun Zhao, Jian Tao, Ning-Zhong Shi, and Nan Lin. A note on the comparison of the Stein estimator and the James–Stein estimator. Communications in Statistics: Theory and Methods, 44(16):3363–3374, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2013:EEL

[ZTZ^{+}13] Bin Zhang, Xingwei Tong, Jing Zhang, Chunjie Wang, and Jianguo Sun. Efficient estimation for linear transformation models with current status data. Communications in Statistics: Theory and Methods, 42(17):3191–3203, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zeller:2019:IDS

[ZVG19] Camila Borelli Zeller, Filidor Vilca, and Manuel Galea. Influence diagnostics for the structural Sharpe model under normal/independent distributions. Communications in Statistics: Theory and Methods, 48(8):1815–1835, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1444176.

Zhao:2014:NCM

[ZVS14] Hui Zhao, Kate Virkler, and Jianguo Sun. Nonparametric comparison for multivariate panel count data. Communications in Statistics: Theory and Methods, 43(3):644–655, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2009:LEM

[ZW09] Xiuqing Zhou and Jinde Wang. LAD estimation for multiple regression models with doubly censored data. Communications in Statistics: Theory and Methods, 39(1):117–129, 2009. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2011:ELA

[ZW11] Zhi-Wen Zhao and De-Hui Wang. Empirical likelihood for an autoregressive model with explanatory variables. Communications in Statistics: Theory and Methods, 40(3):559–570, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2013:RNS

[ZW13] Chao Zhou and Mingzhe Wang. Research on a new system with neglected or delayed failure impact. Communications in Statistics: Theory and Methods, 42(1):1–10, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:EGP

[ZW17a] Yuan-Lin Zhang and Guan-Jun Wang. An extended geometric process repair model with delayed repair and slight failure type. Communications in Statistics: Theory and Methods, 46(1):427– 437, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.995824.

Zhang:2017:GPR

[ZW17b] Yuan Lin Zhang and Guan Jun Wang. A geometric process repair model for a cold standby repairable system with imperfect delay repair and priority in use. Communications in Statistics: Theory and Methods, 46(16):8046–8058, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1171356.

Zhang:2017:OAR

[ZW17c] Yuan Lin Zhang and Guan Jun Wang. An optimal agereplacement policy for a simple repairable system with delayed repair. Communications in Statistics: Theory and Methods, 46(6):2837–2850, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1053930.

Zhang:2018:EGP

[ZW18] Y. L. Zhang and G. J. Wang. An extended geometric process repair model with imperfect delayed repair under different objective functions. Communications in Statistics: Theory and Methods, 47(13):3204–3219, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1353620.

Zang:2019:SEN

[ZW19a] Qingpei Zang and Tao Wang. Sequential estimation for nonhomogeneous Ornstein–Uhlenbeck processes. Communications in Statistics: Theory and Methods, 48(8):1955–1962, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1440596.

Zhang:2019:GPW

[ZW19b] Yuan Lin Zhang and Guan Jun Wang. A geometric process warranty model using a combination policy. Communications in Statistics: Theory and Methods, 48(6):1493–1505, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1433853.

Wang:2019:ASC

[zWDD19] Zhong zhi Wang, Yun Dong, and Fangqing Ding. On almost sure convergence for sums of stochastic sequence. Communications in Statistics: Theory and Methods, 48(14):3609–3621, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1478102.

Zhang:2016:FCG

[ZWL16] Xingfa Zhang, Heung Wong, and Yuan Li. A functional coefficient GARCH-M model. Communications in Statistics: Theory and Methods, 45(13):3807–3821, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2014.906615.

Zhang:2013:AGM

[ZWLI13] Xingfa Zhang, Heung Wong, Yuan Li, and Wai-Cheung Ip. An alternative GARCH-in-mean model: Structure and estimation. Communications in Statistics: Theory and Methods, 42(10): 1821–1838, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2018:QRE

[ZWNS18] Jun Zhang, Jiefei Wang, Cuizhen Niu, and Ming Sun. Quantile regression estimation for distortion measurement error data.

Communications in Statistics: Theory and Methods, 47(20): 5107–5126, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1386319.

Zhao:2017:CHT

[ZWP17] Zhiwen Zhao, Dehui Wang, and Cuixin Peng. Conditional heteroscedasticity test for Poisson autoregressive model. Communications in Statistics: Theory and Methods, 46(9):4437–4448, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1085560.

Zhao:2017:NLP

[ZWPB17] Zhi-Wen Zhao, De-Hui Wang, Cui-Xin Peng, and Li Bi. A note on the limiting properties of the least squares estimation for the random coefficient autoregressive model. Communications in Statistics: Theory and Methods, 46(4):2002–2006, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1032422.

Zhao:2015:ELB

[ZWPZ15] Zhi-Wen Zhao, De-Hui Wang, Cui-Xin Peng, and Mei-Li Zhang. Empirical likelihood-based inference for stationaryergodicity of the generalized random coefficient autoregressive model. Communications in Statistics: Theory and Methods, 44(12):2586–2599, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2016:ODE

[ZWY16] Yongxia Zhao, Rongming Wang, and Dingjun Yao. Optimal dividend and equity issuance in the perturbed dual model under a penalty for ruin. Communications in Statistics: Theory and Methods, 45(2):365–384, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2011:ELF

[ZWZ11] Haixiang Zhang, Dehui Wang, and Fukang Zhu. The empirical likelihood for first-order random coefficient integer-valued autoregressive processes. Communications in Statistics: Theory and Methods, 40(3):492–509, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2012:GPS

[ZWZ12] Haixiang Zhang, Dehui Wang, and Fukang Zhu. Generalized RCINAR(1) process with signed thinning operator. Communications in Statistics: Theory and Methods, 41(10):1750–1770, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2017:EOS

[ZWZ17] Peng Zhao, Lei Wang, and Yiying Zhang. On extreme order statistics from heterogeneous beta distributions with applications. Communications in Statistics: Theory and Methods, 46(14):7020–7038, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1143007.

Zhou:2017:JAN

[ZWZwZ17] Wenhui Zhou, Qiang Wan, Yanfang Zheng, and Yong wu Zhou. A joint-adaptive np control chart with multiple dependent state sampling scheme. Communications in Statistics: Theory and Methods, 46(14):6967–6979, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2015.
1132323.

Zhao:2015:REA

[ZWZZ15] Xiaobing Zhao, Jinglong Wang, Xian Zhou, and Zhongyi Zhu. Recurrent events analysis in the presence of terminal event and zero-recurrence subjects. Communications in Statistics: Theory and Methods, 44(4):710–725, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2010:ELI

[ZX10] Peixin Zhao and Liugen Xue. Empirical likelihood inferences for semiparametric varying-coefficient partially linear models with longitudinal data. Communications in Statistics: Theory and Methods, 39(11):1898–1914, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2016:AQM

[ZX16] Xiaorui Zhu and Li Xie. Adaptive quasi-maximum likelihood estimation of GARCH models with Student’s t likelihood. Communications in Statistics: Theory and Methods, 45(20):6102– 6111, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2014.957852.

Zhang:2018:MSM

[ZX18] Jiamao Zhang and Jianwen Xu. A model selection method based on the adaptive LASSO-penalized GEE and weighted Gaussian pseudo-likelihood BIC in longitudinal robust analysis. Communications in Statistics: Theory and Methods, 47(23): 5779–5794, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1402047.

Zhao:2010:NGF

[ZXD10] Jianxin Zhao, Xingzhong Xu, and Xiaobo Ding. New goodness of fit tests based on stochastic EDF. Communications in Statistics: Theory and Methods, 39(6):1075–1094, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2017:SES

[ZXF17] Yang Zhao, Liugen Xue, and Sanying Feng. Semiparametric estimation of the single-index varying-coefficient model. Communications in Statistics: Theory and Methods, 46(9):4311–4326, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2015.1081950.

Zhang:2016:GVA

[ZXM16] Qian Zhang, Xingzhong Xu, and Shiyun Mi. A generalized p-value approach to inference on the performance measures of an M/E_{k} /1 queueing system. Communications in Statistics: Theory and Methods, 45(8):2256–2267, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.879181.

Zhang:2018:TSI

[ZXSZ18] Ying-Ying Zhang, Yu-Han Xie, Wen-He Song, and Ming-Qin Zhou. Three strings of inequalities among six Bayes estimators. Communications in Statistics: Theory and Methods, 47(8): 1953–1961, 2018. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2017.1335411.

Zhao:2016:ACT

[ZXW16] Guimei Zhao, Xingzhong Xu, and Li Wang. The adaptive calibration of testing p-values. Communications in Statistics: Theory and Methods, 45(4):922–932, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.853791.

Xiong:2013:ELI

[zXyL13] Xian zhu Xiong and Zheng yan Lin. Empirical likelihood inference for nonparametric regression functions with functional stationary ergodic data. Communications in Statistics: Theory and Methods, 42(19):3421–3431, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2010:LTR

[ZY10] Yong Zhang and XiaoYun Yang. Limit theory for random coefficient first-order autoregressive process. Communications in Statistics: Theory and Methods, 39(11):1922–1931, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2013:ADP

[ZY13] Yong Zhang and Xiao-Yun Yang. Asymptotic distribution for products of sums of linear processes under dependence. Communications in Statistics: Theory and Methods, 42(13):2292– 2300, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2014:BOD

[ZY14] Ying Zhang and Keying Ye. Bayesian D-optimal designs for Poisson regression models. Communications in Statistics: Theory and Methods, 43(6):1234–1247, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2016:RWA

[ZY16] Zai-Ying Zhou and Ying Yang. A random weighting approach for posterior distributions. Communications in Statistics: Theory and Methods, 45(12):3441–3457, 2016. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2013.835412.

Zhang:2017:OIC

[ZY17a] Weijia Zhang and Po Yang. Optimal investment and consumption with unknown parameters. Communications in Statistics: Theory and Methods, 46(12):5985–5993, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1116573.

Zhou:2017:NGC

[ZY17b] Ming Zhou and Zhao Yang. A note on generating correlated matched-pair binary data through conditional linear family. Communications in Statistics: Theory and Methods, 46(16): 8059–8068, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2016.1171357.

Zhou:2019:AEV

[ZY19] Zhiyong Zhou and Jun Yu. Adaptive estimation for varying coefficient models with nonstationary covariates. Communications in Statistics: Theory and Methods, 48(16):4034–4050, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1484483.

Zhou:2017:ROI

[ZYH17] Jieming Zhou, Xiangqun Yang, and Ya Huang. Robust optimal investment and proportional reinsurance toward joint interests of the insurer and the reinsurer. Communications in Statistics: Theory and Methods, 46(21):10733–10757, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1242734.

Zheng:2018:SMR

[ZYH18] Yanting Zheng, Jingping Yang, and Jianhua Z. Huang. Shuffle of min’s random variable approximations of bivariate copulas’ realization. Communications in Statistics: Theory and Methods, 47(10):2337–2350, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
964808.

Zhang:2017:CRD

[ZYLZ17] Tian-Fang Zhang, Jian-Feng Yang, Zhi-Ming Li, and Run-Chu Zhang. Construction of regular 2^{n}4^{1} designs with general minimum lower-order confounding. Communications in Statistics: Theory and Methods, 46(6):2724–2735, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1048887.

Zhu:2011:RPP

[ZYN11] Jinxia Zhu, Hailiang Yang, and Kai Wang Ng. Ruin probabilities for the perturbed compound Poisson risk process with investment. Communications in Statistics: Theory and Methods, 40(21):3917–3934, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2012:PWV

[ZYX12] Junlong Zhao, Yuliang Yin, and Xingzhong Xu. Penalized weighted variance estimate for dimension reduction. Communications in Statistics: Theory and Methods, 41(3):453–473, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2011:DRR

[ZYZ11] Li-Ping Zhu, Ya-Ni Yang, and Li-Xing Zhu. Dimension reduction in regressions through weighted variance estimation. Communications in Statistics: Theory and Methods, 40(11): 1929–1944, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2010:SET

[ZZ10] Xiaobing Zhao and Xian Zhou. Semiparametric estimation in transformation models with cure fraction. Communications in Statistics: Theory and Methods, 39(18):3371–3388, 2010. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2011:ELI

[ZZ11] Tao Zhang and Zhongyi Zhu. Empirical likelihood inference for longitudinal data with missing response variables and errorprone covariates. Communications in Statistics: Theory and Methods, 40(18):3230–3244, 2011. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2015:APS

[ZZ15] Miao Zhang and Gaofeng Zong. Almost periodic solutions for stochastic differential equations driven by G-Brownian motion. Communications in Statistics: Theory and Methods, 44(11): 2371–2384, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2017:LTD

[ZZ17a] Yang Zhang and Lixin Zhang. Limit theorems for dependent Bernoulli variables with statistical inference. Communications in Statistics: Theory and Methods, 46(4):1551–1559, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1021015.

Zhang:2017:MHG

[ZZ17b] Yiying Zhang and Peng Zhao. On the maxima of heterogeneous gamma variables. Communications in Statistics: Theory and Methods, 46(10):5056–5071, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2015.1091084.

Zhao:2017:DRB

[ZZ17c] Junlong Zhao and Xiuli Zhao. Dimension reduction boosting. Communications in Statistics: Theory and Methods, 46(9): 4151–4162, 2017. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2015.1079637.

Zhong-Zhi:2017:KAP

[ZZ17d] Wang Zhong-Zhi. A kind of asymptotic properties of moving averages for Markov chains in Markovian environments. Communications in Statistics: Theory and Methods, 46(22):10926– 10940, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-

415X (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/03610926.2016.1252404.

Zhang:2018:NPQ

[ZZ18a] Feipeng Zhang and Yong Zhou. Non-parametric quantile estimate for length-biased and right-censored data with competing risks. Communications in Statistics: Theory and Methods, 47(10):2407–2424, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2016.
1200092.

Zhang:2018:NPM

[ZZ18b] Yi Zhang and Qingle Zheng. Non parametric mixture of strictly monotone regression models. Communications in Statistics: Theory and Methods, 47(2):415–426, 2018. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2017.1303730.

Zhang:2018:SMA

[ZZ18c] Yi Zhang and Qingle Zheng. Semiparametric mixture of additive regression models. Communications in Statistics: Theory and Methods, 47(3):681–697, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1310243.

Zhao:2018:VSE

[ZZ18d] Jun Zhao and Yi Zhang. Variable selection in expectile regression. Communications in Statistics: Theory and Methods, 47(7):1731–1746, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1324989.

Zhao:2018:SRC

[ZZ18e] Sheng-Li Zhao and Qian-Qian Zhao. Some results on constructing two-level block designs with general minimum lower order confounding. Communications in Statistics: Theory and Methods, 47(9):2227–2237, 2018. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2017.
1337148.

Zeng:2019:TWE

[ZZ19a] Guoping Zeng and Edward Zeng. On the three-way equivalence of AUC in credit scoring with tied scores. Communications in Statistics: Theory and Methods, 48(7):1635–1650, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1435814.

Zeng:2019:NRM

[ZZ19b] Peng Zeng and Yu Zhu. A novel regularization method for estimation and variable selection in multi-index models. Communications in Statistics: Theory and Methods, 48(12):3055–3067, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1473603.

Zhang:2019:AAM

[ZZ19c] Sumei Zhang and Guangdong Zhang. An analytical approximation method for pricing barrier options under the double Heston model. Communications in Statistics: Theory and Methods, 48(22):5657–5671, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1549257.

Zhao:2019:MLE

[ZZ19d] Huiyan Zhao and Chongqi Zhang. Maximum likelihood estimation for reflected Ornstein–Uhlenbeck processes with jumps. Communications in Statistics: Theory and Methods, 48(5): 1221–1233, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1425451.

Zhao:2019:MLD

[ZZ19e] Qian-Qian Zhao and Sheng-Li Zhao. Mixed-level designs with multi block variables containing clear two-factor interaction components. Communications in Statistics: Theory and Methods, 48(13):3402–3412, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2018.
1476715.

Zardasht:2012:NER

[ZZA12] V. Zardasht, P. Zeephongsekul, and M. Asadi. On nonparametric estimation of a reliability function. Communications in Statistics: Theory and Methods, 41(6):983–999, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2013:DLM

[ZZC13] Xiao-Xin Zhu, Bin Zhu, and Chun-Zheng Cao. Diagnostics for a linear model with first-order autoregressive symmetrical errors. Communications in Statistics: Theory and Methods, 42(13):2335–2350, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2019:RCI

[ZZJ19] Jin Zhang, Junmei Zhou, and Fen Jiang. A reformulation of the criteria for the independence of quadratic functions in normal variables. Communications in Statistics: Theory and Methods, 48(5):1234–1238, 2019. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2018.1425453.

Zhu:2012:SSS

[ZZL12] Yanyang Zhu, Aofu Zhang, and Baoliang Liu. Sequential series systems and their risk assessments. Communications in Statistics: Theory and Methods, 41(21):3903–3914, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:2016:LAR

[ZZL16] Tao Zhang, Qingzhao Zhang, and Naixiong Li. Least absolute relative error estimation for functional quadratic multiplicative model. Communications in Statistics: Theory and Methods, 45(19):5802–5817, 2016. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). URL http://

www.tandfonline.com/doi/full/10.1080/03610926.2014.
950748.

Zhu:2019:ESF

[ZZL19] Hanbing Zhu, Riquan Zhang, and Huiying Li. Estimation on semi-functional linear errors-in-variables models. Communications in Statistics: Theory and Methods, 48(17):4380–4393, 2019. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/03610926.2018.1494836.

Zhao:2015:SVC

[ZZLL15] Weihua Zhao, Riquan Zhang, Jicai Liu, and Yazhao Lv. Semi varying coefficient zero-inflated generalized Poisson regression model. Communications in Statistics: Theory and Methods, 44 (1):171–185, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zeng:2010:BMN

[ZZP10] Peng Zeng, Yu Zhu, and Kevin T. Phelps. Bounds on the maximum number of factors in designs with two distinct groups of factors. Communications in Statistics: Theory and Methods, 39(12):2302–2310, 2010. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:2013:MAR

[ZZS13] Hui Zhao, Jie Zhou, and Liuquan Sun. A marginal additive rates model for recurrent event data with a terminal event. Communications in Statistics: Theory and Methods, 42(14): 2567–2583, 2013. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:2014:CAR

[ZZSP14] Liang Zhu, Hui Zhao, Jianguo Sun, and Stanley Pounds. A conditional approach for regression analysis of longitudinal data with informative observation time and non-negligible observation duration. Communications in Statistics: Theory and Methods, 43(23):4998–5011, 2014. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:2019:OGL

[ZZZ19] Shengbin Zhou, Jingke Zhou, and Bo Zhang. Overlapping group lasso for high-dimensional generalized linear models. Communications in Statistics: Theory and Methods, 48(19): 4903–4917, 2019. CODEN CSTMDC. ISSN 0361-0926 (print),

1532-415X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/03610926.2018.1500604.